A Bibliography on Data Compression

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

and

David Salomon Professor Emeritus

Department of Computer Science California State University Northridge

Northridge, CA, USA

31 August 2024 Version 1.212

Title word cross-reference

#40 [SD06].

(m) [Haf95]. + [Wir76]. 1 [CD94, LLM89, Sal13]. 11 [LLM89]. 15 × 15 [RC09]. $160.00 [Ano98]. 2 [ASG99, AAR11, GB98, HESF11, KC91, LK91, LA96, MSR04, NC97, SCV11, SSM02, TRS03, YL05]. 2^{k} [Sun16]. 2: 1 [HOR05, IF07]. 3 [Adj06, AH02, BIP01, BS95, BSS09, BF07, BPT06, BA04, CPD^{+}15, FOhC09, FO10, GVSSBRAL08, GSW02, GB06, Hal11, JTC^{+}12, KL97, LKRR91, LAP07, LLS^{+}21, LMH^{+}09, MLDH15, MZP06, MSR04, McC03, MOK00, PST97, PK10, PA07, TCP03, TFV05, TMM96, TCCT02, TPM20, VDPL00, WJDZ10, YKLK99, ZWZ^{+}12, ZGB04, Z^{+}13]. 4 [LP07]. $40 [Ma96]. $49.50 [Hoc95]. 4 × 4 [SMS08]. 5 [Cm10, XXZ11]. $50 [Ma96]. $64.95 [Hoc96]. 9/7 [DZZ08]. = [Wir76]. ^{∗} [DH97, DH98]. ^{2} [Mor12]. E_{8} [GCBV95]. f [LG78]. ∞ [ETT15]. k [MRS05, RS00]. L [AMS^{+}03]. l_{1} [DØJJ09]. L_{∞} [WCB97]. l_{p} [CR06]. m [JMnRS19, KM16]. n [ILRS03, Ken61, ØJH05, PPR02, SDJ04b]. O(1) [FHS08]. O(ND) [Mye86]. ω [LM91]. P [AF88a, DG17, ETT15]. Q [PMLA88, PM88, KS96a, KS98]. QR [BTLK18]. R [CWW20]. R^{N} [BLO01]. × [Lio91]. Z [BHB98, GCBV95, LKA99].

* [Ant97].

-160 [SW12a]. -ary [KM16]. -bit [SDJ04b]. -Bonacci [JMnRS19]. -channel [ØJH05, PPR02]. -coder [BHB98, PMLA88, PM88]. -coding [LKA99]. -Compressed [AF88a]. -constrained [WCB97]. -D [BA04, HESF11, LLM89, MSR04, AH02, GB98, KC91, LKRR91, LK91, LA96, LP07, MOK00, Sal13, SCV11, SSM02, TMM96]. -D/ [MSR04]. -dimensional [BS95, ILRS03]. -Dimensions [Ken61]. -Encoded [DG17]. -error [RS00]. -Grams [KS96a, KS98]. -infinite-constrained [AMS^{+}03]. -Laplacian [ETT15]. -orbit [LM91]. -Range [FHS08]. -type [GLL04]. -WFA [Haf95].

/Data [Hil75]. /Intra [BZL^{+}12]. /MPI [BKK20]. /post [DLT05].

0-387-94211-4 [Hoc95]. 0-89186-6876-4 [Ma96]. ’04 [IEE04].

1 [BY90, Ano95, Ano08i, BS94, Cra98c, GVSSBRAL08, ISO00, ITU94, LK08, LW08, Lum13, SD02, SD06, Shl94]. 1-55860-346-8 [Hoc96]. 1’-Ended [BY90]. 1.0 [BLFF96]. 1.1 [FGM^{+}99]. 100 [Ben96]. 108.80 [Ano98]. 10918-2 [ISO95]. 10918-3 [ISO97, ISO99a]. 10918-4 [ISO99b]. 11172 [ISO93]. 11172-3 [ISO93]. 1134 [Per89]. 11496-10 [ITU02]. 1171 [Per90]. 1172 [PH90]. 12/90 [ITU90]. 128/ [SW12a]. ’13 [ACM13]. 1331 [Sim92]. 13818-7 [ISO03b]. 151 [ECM91a]. 15444 [ISO00]. 15444-1 [ISO00]. 1548 [Sim93a]. 1549 [Sim93b]. 1570 [Sim94a]. 159 [ECM91b]. 16 [Ano06w, PS93]. 16-bit [Ano06w]. 160 [SW12a]. 1661 [Sim94b]. 1662 [Sim94c]. 1742 [Gol06]. 18th [WS12]. 1945 [BLFF96]. 1950 [DG96]. 1951 [Deu96a]. 1952 [Deu96b]. 1962 [Ran96]. 1979 [Woo96]. 1985 [IEE85]. 1986 [Ano88a, Ano88b]. 1987 [Ano88c]. 1988 [WGM88]. 1991 [IEE91]. 1993 [KRW^{+}93]. 1997 [ACM97]. 1997/Amd [ISO99a]. 19th [vL94].

2 [Kro00, MRG99, NDN13a, Nga95]. 2-Channel [SGLT98]. 2-dimensional [GL03a]. 2000 [BSG10, Bar11, ISO00, JPE00, MGBB00, MF98c, MFD^{+}09, NU05, RK09a, Ric12, SEC01, TY04, TM02, Yu04]. 2001 [SM01a]. 2003 [ACM03, FLA^{+}03]. 2004 [IEE04, SC04]. 2005 [IEE05, SC05]. 2006 [Ano06c]. 2008 [LL08, Neu10]. 2009 [IEE09, ISO09]. 2010 [IEE10]. 2011 [IEE11a, IEE11b, SM09, SM11b]. 2012 [CLA12]. 2013 [Bus09]. 2017 [ISO17]. 20th [BJZ94, WGM88]. 2153 [Sim97]. 222 [ECM95]. 25th [IEE84]. 29 [RHC87]. 29th [FLA^{+}03]. 2K [HSP^{+}13].

3 [ISO93, NDN13b]. 300 [Ano95]. 300-to-1 [Ano95]. 32 [Ano06i]. 32-bit [Ano06i]. 321 [ECM01]. 3284 [KMMV02]. 32nd [IEE91]. 3309 [ISO84]. 37th [IEE96]. 3D [SFT03]. 3rd [Weg92].

4 [KL03, LR04, L^{+}05, NU05, PE02, Rez04, Ric03, Sch04, Sym03, TY04, VRHL07, VRL09]. 4.als [Ano06s]. 421M [Ano08i]. 4K [HSP^{+}13].

5 [Kro00, Sch03]. 5/3 [Ara13]. 51 [Sim94d].

64 [LK91]. 64-bit [HD88, Pat94]. 673 [Vit89].

7 [NHHS02]. 749 [SM95a]. 75 [RBR10]. 754 [FOP12, IEE85]. 754-1985 [IEE85]. 78 [AF18]. 781 [BS98].

87e [RHC87]. ’88 [IEE88a, IEE88b, Jur88].

’91 [SR91b]. ’92 [SC92]. ’93 [SC93, vL94]. ’94 [SC94]. ’95 [IEE95b, SC95]. ’96 [BDLS96, SC96]. ’97 [SC97, Car97b]. ’98 [SC98]. ’99 [SC99].

= [GLM10, Sta09].

A-Posteriori [BF95b, XTH09]. A-TDB [CHW04]. A/D [CD00, CDL07, DDGV06]. AAC [ISO03b, Bra99]. Abbildung [Hil91]. ABI [MC13]. Abis [Xu06]. above [EANG04]. absence [Hat95]. absolutely [Tit02]. Abstract [Dun80, KW01, SS78, Sto83]. Acad [Sie90]. Academic [Ano98]. ACB [Lam99]. Accelerated [WKC94, HSP^{+}13]. Accelerating [HLL^{+}20, Sau95, ZZL^{+}15, CC98]. Acceleration [ZRZ^{+}19, KMA^{+}20]. Accelerator [ZZJB13]. Accelerators [AW15]. Access [FOhC09, FN09, Gag97, GE07, IFK12, KN10, RSWW01, DLT16, Jac92, LW99, RSWW02, SYP04, YK20, ZE01]. Accessible [CH09, KCL06]. Accessing [Mul87, MM00]. Accounting [BSSU18]. Accuracy [Vuc05, Cve99, CD00, CDL07, Lin97]. Accuracy-optimized [Vuc05]. Accurate [ALMSS09, Cve98, FG07, LZWL11a]. ACELP [PFAK10]. achieve [MW06]. Achieves [NR08]. ACITY [NDN13a, NDN13b]. acknowledgments [CG04]. ACM [ACM97, ACM03, ACM13, KRW^{+}93, LL08, RHC87, SM01a, ACM78, YR87]. ACM-SIAM [ACM97]. Acoustic [BT93]. Acoustics [IEE95b, IEE04]. Acquisition [SF11, YX12, EW96]. Acrobat [Ado06]. acronyms [YBW00]. across [BFMX03, SW17, Tun04, WLL^{+}20]. Activations [ZRZ^{+}19]. Active [CGC11, NcC02]. activity [SV00]. Acts [HL17]. Acyclic [EEH17]. adaptable [KK01, LGM21]. Adaptation [CC06a, HL17, GZV97, Lan91, PRM97, SSM02, VFE00]. Adapted [WFG^{+}94, CAC97]. Adapting [CFG12, DS04]. Adaption [Zan92]. Adaptive [Abr89, AW04, And12, Ano06e, ALM11, Bar81, BS17, BSTW86, Bör18, Ces91, CW97, CSCW00, CHB86, CDSB98, CW84, CS94a, CS94b, DLT05, Dav95b, DGG10, DY12, ECM91a, ECM95, EGLS21, Fal73, FLS20, Fel74, FK95, FSC^{+}11, FSL^{+}12, FM12, FA97, Fow97a, Fow97b, Fra97, FNI17, GKN09, Ged14, GA08, GMC^{+}06, HB09, HF97, HF99, Hou79, HKKW13, HSL^{+}20, HM93, Jak85, Kar12, KA98, Kno98, KB99, KZ02, KSY95, KS06b, KH04, Lan83b, LS95a, LHK03, Lin95a, LLP96, LC91, LDZ^{+}20, MWCG07, Mal06, MWC01, MAS00, MSC99, MSC00b, Nek07, Nus76, OHSS12, PMLA88, PB98b, PCD17, PW86, Prz98, Ram87, Rez07b, RD82, RA04a, SYZO15, SW12b, SK09, SST^{+}12, Stu94a, Tad98, Tan93, TCKM10, UHB09, WLL06, WZMG07, Wil91a, WB91, WB82, WZW09, Xio11, XW10, YHM20]. Adaptive [YXPM11, YLZ11, YMK10, Yok96, YKO98, ZW07, ZJW^{+}12, AF99, ABH08, ATL^{+}88, Att95, ARH03, BZ97, BW98, BH93b, BHB98, BAGCGC^{+}06, CPD^{+}15, CC93, CZS92, CCMW03, CZZR19, CHW04, CS93, CS95, CW03, DP76, DGG09, DDL98, For96, Fow98, HKA06, HGH05, Huy94, JOC97, JU10, JV04, KLJ12, Lee99, LMS94, Mäk89, Mal01, MNBC18, MM97, MSR04, Mod03, MM05, MMA00, NPW97a, OG04, PX12, PB98a, Ram91, RV94, RA03, RK93b, RSV^{+}00, RSC99, RS01, RHC87, Sai05b, ST97, SW97, SH94b, SM11a, SCC12, Sta07, SRG^{+}21, TT94, THG03, TC06b, TM97, WHH^{+}99, WS00, WSS94, WC88, WY91, WCB97, Wu99a, Yok93, FSC^{+}11, ITU90]. adaptive-boosting [OG04]. Adaptive-CoMPI [FSC^{+}11]. Adaptive-Length [FLS20]. Adaptive-rate [KA98]. Added [MN08a]. Adding [YL99, Eve98]. Additions [Cap89]. Additive [Fen02, HS07, BAL02]. Address [Bec93, MMB07, FG97, SJ94]. addressable [Bet94]. Addressing [MVJ17, CKW93]. adds [Ude93]. Adelson [Sie90]. Adjeroh [Neu10]. Adjustment [SjWL08]. Adjustments [Lak98]. Admissible [LZWL11b]. ADPCM [ITU90]. Advanced [AH14, FIP01, ISO03b, ITU02, NDPL00, Ric06, Wak07, B^{+}13]. Advances [KM11, NDN13a, NDN13b, FGC94, KT13, P^{+}13, Sch05, WS12]. advantage [Abe99, MLS03]. Advantages [GF10, RY90]. aerial [LK96b]. AES [FIP01, HIMM20]. Affine [SG09, MZ95, RK93a]. afforded [WC88]. AFIS [Zir07]. Africa [Kes91]. After [SM97]. against [MR00, SDFH00]. aggregate [FFFM13]. Aggregation [BSSU18, ABA14, ASK12, HBSASA19, IMB^{+}13, KLJ12]. aggressive [Wak05]. Aho [TM05b]. AI [GLM^{+}21]. aid [DSSR95b]. Air [BRNMG^{+}12, CPT^{+}20]. Aire [Pea90, Sie12, Sie90]. AKA [Whi06]. Algebra [EBH^{+}17, Dav12, EBH^{+}16, EBH^{+}18, EBH^{+}19]. Algorithm [AMPF09, ABR10, A^{+}87, BH07, Bar00a, BŁ01, BGPW96b, BY76, Bi08, Bla09, Blo96b, BF07, BJ95, BA04, BBBP10, Bre87, BW94b, BR10, BF95b, CDAM07, Car97a, CBK96, lCcYxFpW09, CSF77, Che05, CYL^{+}08, CRN08, CWY^{+}10, CHMW10, CZZZ11, CZT^{+}11, CC06b, CK13, CIS08, CK94b, CK94a, DZZ08, DD98, ECM91a, ECM91b, ECM95, ECM01, Fel03, Gag94, GKSB17, GWSR12, GK88a, GK88b, HB09, HC06, HZKL11, HNC^{+}16, HM76, Ign98, Jen99, JLJ08, Joh89, KKS00, Kie99, KRK09, KKJ11, Kno98, KU95, Kur83, LZWL11b, LV03, Lan83b, LA12, LV95, LBG80, LST99, LZCF10, Mai95, MTA^{+}20, MYS20, MEIS09, MSW10, McC76, MSV98, Miu07, Mye86, NZ07, NK16, PM91, PB98b, Pig99, PSM01, PS06, Pyl90, Rez11a, RPE81, RD82, Sad96b, SMS08, Sch72, SSP06, Sco02, SW12b, SAS10, SK09, SPS98]. Algorithm [Tan93, TW92, Tra87, VRHL07, Vol97, WHZ12, WSS96, WSS00, Wil91b, WY10, YS10, YYZ12, YZLH16, YK10, ZCW^{+}11, ZL77, Abe05, Abe07a, Abe10, AM95, AAJ04, AGJA06, ABH08, BKS99, BY00, BL97, BPMA02, BH92, BH93b, BDCC97, BM03, CS92, Cha87, CWL99, Che15, CDC96, CHW04, CLD02, CL95, Cra98a, Cre96, DS04, DB02, Deo00, Deo02, DY91, Dum06, FWG98, FW93, Ghi04, Ghi05, GRG05, GLL04, Hor91b, Hos02, JT96, JS99a, Jia95, Jon91, KY99, KM95, Lak00b, LLCC95, LW21, LAPL07, LPR97, LHY12, LA96, LNhCW16, MB00b, MLP99a, MLP99b, Mis91, NLW96, NT91, dlCNLA^{+}95, OT03, OPS91, OG16, PBC05, PGS10, PMK91, PB98a, PM95, RL95, Sad98a, Sad96a, DP91, Sch97, SH94b, SMRR04, SPS99, SKCB95a, SKCB95b, SK99, Sta07, Sub99, TFRH98, TM05b, TN95]. algorithm [TO04, TLLB06, TVS^{+}03, TR88, VO91, VFE00, VRR^{+}16, Vuc06, WZL^{+}17, Woo94, WF03, WZ91b, YL99, YG03, ZDY97, ZTSM05, ZLN06, ZC91, BS98, SM95a, Vit89]. Algorithmic [AA71, CD09, Cha77, EEH17, PL90, Wu96, Zur89, Cra98a]. Algorithms [ACM97, Ach94, AKSV15, AM96, AM98, Ano06e, AM93, BM90, BGPW97, Bet94, Bil20, BYR06, Bud76, CFG12, CKC09, CMP94, CDL09, CLR01, DUH^{+}19, DLOM08, FM95b, FL90, FVW10, GKPR96, GS85, GGG^{+}09, HC92, KKP16a, KKP16b, KN05, Kno98, Knuta, KM99a, LMJ^{+}21, LLM89, LE01, LS94, LXG^{+}18, Mar93, Mat96, MMB07, MT02, NLW99, Pas76, PC08, RY90, Rez08, SB97a, SM14, SR93, Uhl96, VRL09, Wil93, Wir76, Zal88, Zha90, ZZL^{+}15, ZBMM97, DS92, AHCI^{+}12, Api91, AB97, BK74, BTLK18, Bun97a, CGSS15, CM91, Cha00, CDDM05, CB97, CN96, CL96, Cus90, Deo02, DQ97, DWB04, DW05, EV14, EGS91, Eve98, FFW98, Fow97b, FU98, HS94, INI^{+}16, JKL13, KJP99, KWLG92, KH00, KG95, KB74, KM97a, KMS98, Lan91, Leh02, LBH91, LZ05, LM08]. algorithms [Mac12, MV20, MR93a, MPL99, Mod03, NLW95, PO00, Per21, PST05, Ran88, RMB91, RL09, Sai05a, Sew00, Sha94, SHD02, Tad98, VW98, Vol02, Wak05, Wak06, WF95, Yok98, You98b, Zir07]. Aliasing [Rob91b]. Aligned [AM05, Ant95b, CK06]. Alignment [CK13, Wol99a, Wol99c, NKP^{+}16]. Alignment-Based [CK13]. Alignments [HDCH09, MPASM^{+}09]. All-Match [DB08]. Alleviating [ZLT^{+}18]. Alley [Kie99, Kno98, PB98b, Pig99, Sco02]. Allocated [ZG02b]. Allocation [BH07, EJ07, HB09, LZX^{+}09, LMH^{+}09, LCG11, MFT07, QB09, Ric11, WB82, FZ05, Hoa99, HN02, HAH05, KUCV04, MS02, Moh02, Ort96, PBC05, Raj04, RLR04, SH04, SCHB05, SLR05, SY04, Use96, Wu93, WBM03, ZAA00]. allow [ISO99a]. Allows [Man97]. Almost [BY76, De 03, GR99, NZC09, OS90, Wan88, YK10, KN05]. Almost-optimal [GR99, KN05]. Alpha [SWGZ09, WPA08, WS09]. Alpha-Plane [SWGZ09]. Alphabet [BN14, Dun80, SLZ08, AdlF06, CCMW03, Gu05, LF95, LNV97, RA04c, Sai05a]. Alphabet-Independent [BN14]. alphabetical [She92]. Alphabets [GvV75, MSWB93, OS03, RA04a, XG97b, ZIL93]. Also [lGAR03a, Ano06k]. Alternative [Del95, CC00]. Alternatives [FK85b, KBN08, FTL03]. Alto [ACM13, BBKF95]. Amar [Neu10]. Amd [ISO99a]. America [Ano09c]. American [Ano09a, Ano06c, dABdSFNA08, Fou85]. Amiga [Ano09e]. Amplitude [Kra49, PFAK10, Smi79]. Analog [ARR10, ECBL12, HGFL09, ME10, PR01, TFA93, WAL05]. Analyses [AM96, AM98, FLZW10, GPM02, Say92]. Analysis [BGPW96a, Bel11, BT93, BRD10, CCM96a, CCM96b, Cre94, DFH^{+}19, DH18, DSV00b, DF08, DSG12, Fis95b, Fow06, Fow08, Fow09, FDH^{+}20, HW98, HTS^{+}18, HV91a, HV94a, KHCS08, KD18, Lew95, LHD^{+}20, LJZ^{+}08, LMH^{+}09, Liu10, LS94, MSV98, MR93b, MFT07, MP06, PFAK10, Por73, Raj20, RRMG07, RH07, SQ16, Sam90b, SB85, SS05a, SZ13, SPS98, SMB98, Tei98, TWC^{+}09, Vit87, WJDZ10, WZY^{+}11, Wil70, WFLZ09, You98a, ZKK09, ZDR06, AF04, BDV00, B^{+}13, BG96, CME06, Cre02, DSSR95b, Den01, DY91, DVDD98, How93, HV93a, Ish89, JT96, KLV07, KAB11, KL05, KK08, KGR06, LCLX04, LS95b, MV20, Man01, Mar02, M^{+}09, MTTH13, MR06, RG02, Sai04, SO94, SGA^{+}16, SPS99, SY04, TBKM05, TH03, WV99a, Zhu02]. Analytical [LZWL11a, CS13, GFV97]. Analytics [JAL^{+}12, PZZ^{+}22, XCK18, ZZD21]. Analytics-Driven [JAL^{+}12]. analyze [QLQ11]. anatomical [ZGB04]. angiogram [GSW02]. animated [PA07, ZO04]. Animating [Kom96]. Animation [GYM93, Hal11, HKMS08, Kom96, YGM97, VB14]. animations [Cha96, COMF99]. annealing [BWS95]. Annotated [KT07b]. Announcing [FIP01]. Annual [ACM97, ACM13, IEE96, KRW^{+}93, YR87, ACM78, ACM79, ACM95, IEE84, IEE91]. ANSI [Ano06c, IEE85]. ANSI/IEEE [IEE85]. antenna [ZDR06]. anti [QZ14]. anti-forensics [QZ14]. Antidictionaries [CMRS00, DI04, DI05]. Antidictionary [CN02, OM07]. Antisequential [BB05]. antisymmetric [YSZ11]. Apache [Pie03, WZL^{+}17]. Apparatus [BT93, ELZC84, Wel84a]. Appearance [RK09b, WDR11]. Apple [Ano92c]. Applet [Bar00a]. Applicability [Str99c]. applicable [JdVL03, She90]. Application [Bak07, BKRSR09, BFCP09, CS95, DXS08, FSC^{+}11, Fis95a, Fry85, Goo03, HL05, Hou79, JAC19, Jon88, KA95, KJ10, LLS^{+}21, LB81, MMZ12, Mun91, PR98, Ryt02, Ryt03, SPS98, Suz11, TR93, TPM20, WWY^{+}07, WYM10, WS12, WMNP12, Yok93, YMYS99, ZBMM97, ZZPB09, ALRT16, BT17, Ber99, Bör80, CD95, ECM98, EEKB15, Gin95, Goo91, Kes91, KUCV04, KNTG08, Lar96, MS97, NAIP03, Nat94, NC97, NM91b, PM98, PMH95, RK15, RA04c, Sad99, Sal91, SPS99, Tro96, TMM96]. Application-Driven [WYM10, WMNP12]. Applications [AM96, AM98, ACD02, ABM10, BIP01, BY00, BCH02, Bil20, BBHB94, CNPC98, CNCC98, CMP94, ETT15, Eln84, JSS15, Jog73, KW03, LSH12, LV08, L^{+}05, LE01, LLYH14, MOT^{+}03, McC92, MR93b, MN08c, MN09, MJZMA08, NK20, Per06, PR11, RY90, RB98, RC09, RKL16, RV09, Sam90a, SSP06, Sho08, Spi93, SDV11, Tar79, Tsu72, Wil96, XSW07, ARZG03, ABF^{+}91, Ano88a, Ano88b, Ano88c, BTLK18, Cap85, CM97, CLA12, Dav12, FLMM09, GLI16, HM87, HM91, JC95, Laf00, Le 91, LJ05, LG78, Lu92, MTTH13, NT05b, PHM^{+}14, P^{+}13, PB99, PR00, R^{+}06, R^{+}07, Ruf95, Sch03, Sch04, SS04, SM11a, Sou95, Wak05, XWC^{+}17, YSZ11, You98b]. Applied [BG96, KU95, Lan10, Li86, Wan73, BKV05, BHT19, BH93a, Mar02, Nad83, Prz00, Reb12, Wan06]. Applying [HN07, LGLK05, YHM20, MLS03]. APPn [ISO99b]. Approach [Ach76, BW10, BK90, DP01, Fon81, FM12, GGM12, HDK^{+}12, IK00, JBG17, JGM^{+}20, KTSA99, KK07, Kra91, LF03, Liu10, MW10, MŠH12, MKM^{+}08, NR99, NYJ08, NYJ09, PB11, SA93, San89, Sha10, SMB98, TY08, TCN^{+}17, Teu01, WZY^{+}11, WGZW14, Wit08, WK93, XMRF^{+}13, YSM08, ZZL^{+}15, ZYT^{+}15, AMBC19, BZ97, BBG10, BK92, CPT^{+}20, CKS95, DLT16, DHW^{+}17, GB94, JWK98, Kes91, LS04, Li02a, MV20, MZP06, MZ91, MB02, Nov98, OY06, OWS98, OHRS21, PH91, PD07, QT03, SW06, SS00, Str99b, TPY95, TCOR05, TH01, TW00, Tse05, Von04, XWC^{+}17, YZZ^{+}14, ZIIK05, ZLN06]. Approaches [Bel87, MV16, NK20, AN10, CLME04, MSE17, Pat94, Tad98, Whi06]. Appropriately [ABS^{+}16]. Approximant [DV01]. approximants [Gul04]. Approximate [AKL18, BKS^{+}18, FLS20, HWY10, JSCM17, RNOM09, Sad96b, ZMAB03, BFG09, NKT^{+}01, Sad93, TM97, YGSR01, ZA05]. Approximated [LWZI12]. Approximating [DV01, dRV12]. Approximation [DGG10, FK09, KB93, Leh02, PM05, Roe77, Ryt02, AAJ04, AGJA06, Bar04, CKW91, CDGO02, CS13, DPS93, DGG09, FGC93, Ind91, LW14, Ryt03, Sam85]. Approximations [BYR06, LW21, PGS10, WV00]. April [ACM79, Ano03, Auc98, SR91b, SC93, SC96, SC98, SC02, SM12, WGM88]. Arbitrarily [TOL11, FM95a, KA99, SRP04]. Arbitrary [BRALSSMP08, BK96, Fre61, LZR^{+}10, LZM^{+}10a, Mor76, BPZ99, CVDL16, GD02, GM99]. Arc [KG03, MZT12, Nel87]. Arc-length [KG03]. Archetype [BJ95]. architected [KMA^{+}20]. Architecting [KHHK21]. Architectural [FGC94, MV16]. Architecture [Bar11, BRV08, KGG^{+}14, LZX^{+}08, NMN99, Nga95, PB11, Ano06i, BDS95, BDCC97, DB03, GAS16, GCK^{+}96, MV20, MS05, RH91, SKCB95a, SKCB95b, SR91a, WL99]. Architectures [Ach94, HAM17, KB92, MRH11, SFT03, KSCE16, Mar93, MBBA91, SHC^{+}16, SJPB17]. Archival [BLC^{+}16, Pec82]. Archive [Del93, Gil08].

Archiver [Mah08a, Ano09e]. archivers [Ano08e, Ros06]. archives [NT91]. Archiving [Nel87, Ano09f, Dia16, HWS06, HHG07, HHW05]. Area [MCM^{+}17, Pea90, Sie12, JB95, JKV15]. Arithmetic [AFL96, Car97a, DOK88, ECM91b, HV92c, HV92d, HV94a, IEE85, JCF93, Kar12, Kle89, LaJR81, LHW00, MWC01, MŠH12, MSFZ08, MCRKC09, MNW95, MNW98, PMLA88, Ris76, Rub79a, SG92, Sch98, Sco02, SMFZ07, WNC87, XG97a, ZZPB09, BM98, BH92, BH93b, BHS06, Blo96a, CKW91, CIRM95, DHS01a, DTZZ12, DHS02, FGC93, FGC94, FN93, HS98, HL05, HV91a, HV91b, HV92a, HV93a, HV96, Huy94, JYHC03, KB06, KCR98, LW99, LW04, MW99, NYC05, PMK91, PSH00, PS93, Sai04, SM98, TT94, TR91, Tom06, Ton93, XWL03, XHS05, XG97b, van99]. Arithmetic-Type [MSFZ08]. ARQ [CMWM00]. Array [GPM02, MS12, NZC09, SLZ08, ZN08, ZRZ^{+}19, BBQ^{+}95, KC91, Lee91, LCL03, RK15, Sad99, SH92]. Arrays [ABM08, FOF09, FOF11, FH08, GMV17, GNF15, LMSE18, Lin14, MM93, Neu10, RL99, Mul87, Sad98a]. Art [Knuta, PF^{+}88, SGLT98]. artefacts [RKSM03]. articles [Hop09]. Artifact [MGD08]. Artifacts [QLH12, Cre02, vB97]. Artificial [Abb93]. Arts [IMA06]. ary [KM16, LLN^{+}04]. ASCII [Ano09a]. ASI [Fis95b]. ASIACRYPT [WS12]. ASIC [BL98, MCM^{+}17]. ASIC-Based [MCM^{+}17]. Asilomar [M^{+}98]. Aspects [BGK95, Boy91, Rub79b, BGK91]. Assembly [Kut02]. assess [BHT19]. Assessing [LLW^{+}14, YTY12, TDG^{+}19]. Assessment [GMSK09, IRB97, MC13, Mor12, RL08, WGZ^{+}18, MSE17]. Assignment [WMWW10, BP00, CCV04, HW06, LBS98, RF05]. assignments [MZ95]. Assisted [BF12, JR12, Lev95, Sha11, SWGZ09]. Associated [ISO93, ISO03b]. Associative [Buy94, Lea78, RKSM03, Lam99]. Assorted [Per06]. Asteroid [CHB86]. Astronomical [PWS10]. Astronomy [KKT10]. Asymmetric [AJN08, Cen09, Haf95, MS95a, MTD08, YSXZ04, DSV00a, Gag06]. Asymmetrical [LLW11]. Asymptotic [Fre91, HO19, JS95, KS00b, Rez05, DTZZ12, IF06, WZL91]. Asymptotically [ARR01, GL09, MGL11, Wan88, YO91, Ghi03, Sub99]. Asymptotics [Ger91, ZKD05]. Asynchronous [ITU94, PN91]. Asynchronous-to-Synchronous [ITU94]. Atlanta [ACM79]. ATM [Pap99, SMS91, YH00]. atmospheric [KNTG08]. attack [BH93b]. attacks [LRH^{+}21]. Attention [ZCG98]. attraction [BW99]. Attribute [DKB^{+}16, JLJ^{+}20]. attributes [Kes91]. Audio [BG03b, BS94, dABdSFNA08, Coa06a, HS01, Haw90a, Haw90b, ISO93, IRB97, ITU91, JC95, Jay97, KY08, Kno98, KMS96, L^{+}05, LLYH14, Mal07, NR14, Pan93, Pan95, Poh85, RH96, Rat92, Rei94, Rez04, S^{+}09, Shl94, SGLT98, Wav06, YM08, dMGdC10, Ano06a, Ano06b, Ano06m, CMWM00, Coa06b, Cre02, DYW^{+}96, DHS^{+}01b, Ghi03, Ghi04, Ghi06, HS98, ISO03b, KC06, KL05, LR04, Mer91, RK95, SH94b, TYD^{+}20, hyd06, vB98, Ano06m]. Audio-Visual [dABdSFNA08]. AudioPaK [HS98]. Audiovisual [ITU06a]. Audit [LYC16]. Auditing [TCN^{+}17]. August [HWS06, HHG07, HHSS08, HHW05, HPV09, H^{+}10, HPT11, IEE95a, R^{+}06, R^{+}07, Sch03, Sch04]. Authentication [WHZ12, TG17]. authenticity [Kir04, Kir05]. Authorities [ISO99b]. Auto [LLS^{+}21, ZXM^{+}10]. Auto-encoder [LLS^{+}21]. Autobiography [Gri73]. autocorrelation [CA02]. Automata [CCRCK09, CK93a, CK94b, CV96, CV97, Dac01, JP07, MCRKC09, AFHU97, BBC95, BBZ14, CK93b, Haf96, Hau12a, HB99, JdVL03, KF96, Laf00, LM91, OY06, Zip91, Zip92]. Automated [Per03a, Ano92c, Kes91]. Automatic [Cha89, Fra67, HZ95, JEEL16, KD18, MdM67, Rai87, TDL^{+}19, Zhe03, MP95]. Automatically [Ano06q]. Automaton [LdV95]. Autoregressive [GH08]. availability [Bli91]. Available [Ano06p, Ano11b, Ano09e]. AVC [MHKK03, MHKK06, MFT07, SMS08, SMW07, SS07, TY04, Ten04, VRHL07, VRL09, WXC^{+}10]. Average [AM96, AM98, AL10, JST01, LS96, LST99, MSV98, BSS05, DHS02, JLS04, LS95b, RS00, RA03]. Average-Based [AL10]. Average-Case [MSV98]. averages [BS04, WBS05]. Averaging [MD95]. AVHRR [THM91]. AVIRIS [RC96]. AVL [Yok93]. Aware [AS14a, ABS^{+}16, BS17, BLSM19, CRN08, DH18, GS12, PBO^{+}15, ASK12, BA06, BS06, GHSV06, IMB^{+}13, JCS^{+}08, SRG^{+}21]. awareness [Li10]. AXECHOP [LDM05].

B [GVSSBRAL08, DH97, DH98, BL97, BDHJ04, FV16, KB06, LZWL11a, LZWL11b, Sew01]. B-coder [KB06]. B-Spline [BDHJ04, BL97]. B-Spline-Based [LZWL11a, LZWL11b]. B-Tree [FV16]. B-W [Sew01]. Back [Ano18, BR05]. back-to-front [BR05]. background [DSSR95a, SMRR04]. Backup [ZYF^{+}20, Mia90, SHWH12]. backups [Bre04, SS04]. Backward [MEMS06, RAE16, YGNM07, GZV97, GMNK06, NO95]. Backward-Compatible [RAE16]. Balanced [GF10, Tar79, FH02, RF05]. baleine [Sta09]. ball [CR06]. Band [EY81, MP06, Tan93, Tat94, Wic90, PHG86, SZ05, TCOR05, ZF92]. band-limited [ZF92]. Band-Limiting [EY81]. Bandelet [PM05]. bandlimited [CDL02]. Bands [OVCS^{+}09]. Bandwidth [Bec93, Fou85, HGH05, IF07, KR10, SKY10, SF09, USY17, YNQ17, Cam79, CT98b, KB99, Luc96, MB91, THG03]. bang [VN90a]. Bangalore [PBPT95]. Banks [SN96b, WL08, DSV00b, DKG01, HMS92, PRM97]. Bar [Maa94, NC97]. Barbara [SM01a]. BARF [Mah08a]. barotropic [Wil02]. Base [BSSU18, GAS16, LB81, LAPL07, Sno76, WZ91b]. Base-victim [GAS16]. Baseband [Tra87]. Based [AKS10, AR08, AK08, AAL^{+}21, ABR10, AH14, ASLB20, ANS12, ALMSS09, AL10, BZZ08, Bar12, BRNMG^{+}12, BS17, BZ08, BLC^{+}16, BH12, BES07, BC10, Car97a, CD08, CW91, CFY^{+}10, CCKH21, CXZD12, CXYQ12, CHMW10, CLS19, CK13, CN02, CV96, CVK97, CV97, DS08, Dav95b, DGG10, DDT^{+}12, DZ08, FWS^{+}08, FJ07, FNP08, Fen02, FSC^{+}11, FM12, Fu03, FLZW10, FGM21, GWT10, GWSR12, GJ20, GK88a, GK88b, HU99a, Hal11, HY10, HU99b, HV08, HTS^{+}18, HDK^{+}12, HC92, HV94a, HHLW20, HSL^{+}20, HH08, JA07, JAC19, JFC12, JEK98, JR12, Kid09, KKMS99, KT09, KU95, LZWL11a, LZWL11b, LMBN08, LZX^{+}08, LZX^{+}09, LZM^{+}10a, LM00, LE07, LLA08, LYC16, LLT^{+}16, LSC10, LHS05, LSW08, LWZI12, LG78, LT09, LS94, LDZ^{+}20, MGD08, MYS20, MEIS09, MWC01, MMZ12, MRS01, MW97]. Based [MCM^{+}17, MCRKC09, MM03, MKM^{+}08, MCL12, Mof91, MFT07, MO11, MF11, MJZMA08, MPGMD19, MP06, NM91a, PGW^{+}17, PB11, PP18, Pop11, PS06, Pro86, QWH12, RS16, Raj20, Ram85, RJL11, Ric11, Ryt02, SI99, Sad96b, SP96, SQ16, SKY10, SB85, SP11, SSP21, SWW12, SHW^{+}01, SK09, SSRM09, SR16, Suz12, TV09, TSS99, Tao82, TW92, Tod89, TW01a, TPM20, USY17, VST11, VS10, VN08, WZMG07, WJDZ10, WHW10, WW12, WHB16, WZY^{+}11, WSS96, WN10, XG10, XXZ11, YHM20, YZLH16, Yok96, Yok97, ZSP^{+}19, ZD95a, ZD95b, ZXM^{+}10, ZB12, ZRZ^{+}19, ZLC^{+}15, ZZL^{+}15, ZYT^{+}15, ZHX^{+}19, ZX11, ABA14, AM95, AF99, AF04, ANT95a, AB08, AHCI^{+}12, AMS^{+}03, ADCU08, Ant97, Ara13, Arn97, Arn99, AR06, AN10, BGR01a, BGR01b, BDV00, BK00, BL97, BW98, BJCO03, BB05, BF95a, BP00]. based [Ber99, BBC95, BHLY20, BH10, BKR92, BM96, Car00, CPD^{+}15, Cha96, Cha87, CSCW00, CZAR20, CB97, Chi99, CPR03, CO97, CO98, CAC96, CAC97, CDGO02, CLM14, CKV97, Cus90, DLT05, DUH^{+}19, DS06, DB02, DB03, Dav95a, DGG09, DSM07, DYW^{+}96, DKS21, DPS99, EMS03, EGNA06, EW96, FC98, FSC91, FFW98, FW95, FWG98, FAEY95, FDZ00, FJ96, GK99, GKPS05, GZV97, GYLC09, GD07, GO96, Gul04, GMNK06, HGS97, HSX02, HKMS08, Hoa99, HM07, HV93a, IW94a, JJM18, JG05, JSC20, JV04, KLV07, KG95, KL03, Kir05, KK08, KCR98, KPY96, KM95, LS04, LGM21, LK96b, LM99, LS95a, Leh02, LDM05, LMV03, LO99, LSD97, LN02, LAP07, LMO97, LDK99, Lin95a, LCL03, LW11a, LHY12, Lin98, LW04, Liu07, LPZJ15, LRO97, LV14, MB00a, MPL02, MRS99, MSR04]. based [MS05, MAC^{+}04, Mis91, MP95, MM05, MSYY98, MBBA91, Nam91, NC97, Niu02, NO95, NYC05, OS97, OR94, OTW^{+}99, OG06, PBC05, PX12, PGS10, PRM97, PBEA95, PA07, PH91, PD16, PHG86, PR01, Prz98, QT03, RK93a, Raj04, RH97a, RH91, RAFH92, RB01, Rez05, RS04, Ryt03, SAR00, SMC99, Sch05, Sch06, Sha06, SLR05, SCC12, SGA^{+}16, SHK99, SGD05, Ski05, SN94, SZ05, SHC^{+}16, SKCB95a, SKCB95b, Sub99, SM95b, SZM03, Sun16, TKR00, TPAK96, TCOR05, TM04, TC96, TC98, TW96a, TWMG93, TYD^{+}20, VL97, VL98, VN90a, WHH^{+}99, WCH03, WLZW04, WS05, WAL05, WBS05, WSK^{+}11, WC02a, WV00, WS09, Wil00, Wil02, Woo00, WF93, XWL03, XSW07, Xue99b, Xue99a, YP03, YR96, YS07, YZZ^{+}14, YGNM07, YMYS99, YM97, YML97, ZO04, ZCG98, Zha99, Z^{+}13, ZZJB13]. based/segment [CAC96]. basée [LG78]. baseline [CL91, Lak00b]. Bases [BJZ94, Dau88, FLA^{+}03, HGR11, WFG^{+}94, BK96, Eff97, ZD81]. Basic [Fra97, ITU91, PMLA88, WST95]. Basis [AAR11, LE02, Wic90, Ber71, CWYW91, Che15, LW14, SWA94, WSA94]. Batch [TH10]. Batch-Pipelining [TH10]. Batched [BTLK18, GMC^{+}06]. batching [SRG^{+}21]. Battery [KL07]. Battery-powered [KL07]. Batting [Bar00a]. Bauded [Tra87]. Bayes [PGOO94, WG94]. Bayesian [Bun98, LZZ^{+}12, Suz11, Suz12, TTW07, VL98, WB17]. Bayesian/MDL [Suz12]. Bazaar [Ano00]. BC [LL08]. BCH [FSL^{+}12, KK07]. BCJR [Nov00]. BDAM [GMC^{+}06]. BDDs [PV06]. Be [SW17, WPXL09, ZW08, BKR^{+}93, FS03, Kro00]. beaked [Sta09]. beamlets [HCD04]. Beauty [PR86, PL90]. bec [Sta09]. been [Per21]. Behavior [HL17, KS00b, JS95, LRH^{+}21, Szp91, Ven17, WV00]. behaviour [RS02]. Beijing [WS12, Yan10]. being [FS03]. Belief [Xio11]. Bell [Neu10]. below [EANG04]. Belur [Ma96]. BEM [HO19, XTH09]. Benchmark [LN17]. Benchmarking [Swa08]. benchmarks [Ano08d]. benefits [BSF16]. Benford [AOT13, QZH10]. Bentley [RHC87]. BeOS [Kro00]. Berlin [FLA^{+}03]. Bernoulli [GL09]. Besov [DGG09, DGG10, KP03]. Besov-Type [DGG10, DGG09]. Best [AA71, MD95, Wic90]. Best-Basis [Wic90]. Better [AOAAH17, BS88a, Bel86, WPXL09, ZW08, KSCE16, LKG12, LL96, WZ91a, Mah08a]. Between [Col86, ER78, Abe99, BW94a, Cha00, KFSM04, TDL^{+}19, VW98, WZY^{+}11]. Beyond [Dun81, Joh89, Mil81, Pet81, KL16, RTK^{+}16, SW12a, Tom06, LKG12]. Bézier [ML98]. bezstratnej [Swa04b]. BFS [AD09]. Bi [CC01, CV96, CV97, FVW10, IW96, JAC19, MF96, NPW97a, RMB07, Huy94, Mal01, MF97b, WW92]. Bi-dimensional [JAC19]. Bi-directional [FVW10]. Bi-Level [CV96, CV97, RMB07, CC01, IW96, MF96, NPW97a, Huy94, Mal01, MF97b, WW92]. biased [FA93]. biasing [BJ18]. Bibliographic [LB81]. bibliography [Sab94]. Bicriteria [FFFV19]. Bidirectional [FK90, MJZMA08, SK10b, KSY95]. Bidirectionally [Gir99]. Big [DH18, VPS17, Win06, ZLC^{+}15, ARA91, XWC^{+}17, YZZ^{+}14]. Big-Memory [ZLC^{+}15]. BIIF [Car97b]. bilevel [ATL^{+}88, CS95, HS97, PB99]. Bill [Bar00b]. Binary [AJ95, AJ96b, A^{+}87, AM05, BY90, BdBN12, CD94, Cha66, CY92, Che77, CJDL09, CDL13, DZ03, DRR10, ECM91b, FLZW10, GM59, Hea72, IEE85, JEK98, KTMS10, Kno80, LK96a, Law77, LLW11, LKR12, MCRKC09, Mof91, MS00, MA05, MA06, OM07, PMLA88, Per03a, Per05, PB99, Rez07b, Rob97, Roe77, Sal00b, SYO94, SWGZ09, Vas07, WCY^{+}10, WFLZ09, ZCW^{+}11, AM95, AJ96a, AF99, AF04, BBF02, BAL02, BPMA02, BHB98, CM04, EA95, FGC94, FN93, HL02, JWK98, KB06, KK01, KGR06, Kri02, Lan91, LW99, LXG03, Maß15, MBBA91, NL03, NHHS02, OY06, RF00, She92, SV98, WF95, Zan92, ZGF02, vB98]. Binary-Image-Manipulation [A^{+}87]. binning [MGM13]. bintree [WF94]. bintree-structured [WF94]. Biography [Mah90]. Biological [CDAM07, YDDB15, AL00]. biologically [LSD97]. biomedical [NACM08, PWM18]. biorthogonal [BCP20, Use96]. bisection [KYNC96]. BISK [BF07]. Bit [ABR10, BLS11, CDW11, DR83, DB07, DB08, FK85a, HB09, ITU91, KSCE16, LSC10, LCG11, MFT07, Moh02, MMA99, OBGB11, QB09, TV09, Teu78, WB82, ZWZ11, ZX11, lGPL^{+}00, Ano06i, Ano06w, BK94, CL06, CAC96, CDMW02, CF97, CD00, CDL07, DB03, FC98, FZ05, Fen93, FA93, GC16, HLV94, HLV96, Hoa99, HAH05, HD88, JOC97, Jak78, KUCV04, KCS93, Li02c, MGG05, MB00a, MM97, MF98b, MRG99, NZ95, Ort96, Pat94, Raj04, RLR04, SC01a, SDJ04b, SH04, SCHB05, SLR05, SY04, TKYS16, TFV05, TEGM03, Use96, Use03, Wu93, Wu99b, YA13, YJE04]. Bit-Level [ABR10]. Bit-plane [KSCE16, DB03]. Bit-Rate [BLS11, ITU91, ZWZ11, ZX11, BK94, CD00, FC98, FA93, MF98b, TEGM03, YA13]. Bit-Strings [FK85a]. bit-vector [Jak78]. Bit-Vectors [Teu78, GC16]. Bitext [MPASM^{+}09]. Bitgroup [VT95]. bitmap [Ant95b, BK91, KL16, NCB15, WOS06]. bitmaps [LSYKK16, PW03b]. Bitplane [AL10, ALM11]. bitrate [CAC97]. bitrates [KC06]. Bits [GL09, S^{+}88, Win06]. Bitstream [Kut02, QZ12]. Bitstreams [BH07, XD11, WX00]. Bitwise [DX11]. Black [LaJR81, Mor76]. Black-White [Mor76]. Blainville [Sta09]. Blass [Bar00b]. BLAST [CW07]. blending [Bun97b, May99]. blendshape [SILN11]. Block [Ach76, AM97, ALA98, BZL^{+}12, BFT11, BW94b, lCcYxFpW09, CAL03, CD07, CPL09, Das95, DR83, Dun80, EC16, HLC07, JWK98, KHJ^{+}09, KMM^{+}12, Kru92, LP75, Ma96, MK03, MYS20, MOT^{+}03, MF10, MF11, PH11, Rea72, Rez07b, RD82, SM14, SH97, TSS99, Bod95, BB95, CS92, CVC95, CHW04, CAC96, CLD02, CW03, DLT05, DB02, De 06, EHAN05, ENA05, FWG98, FNK94, KW95, KPY96, Lak97, Lar98, LPR97, LMS94, LPZJ15, MB00b, NL03, dlCNLA^{+}95, PELA02, RA03, RA04c, Sad98b, Sch97, SO94, SCC12, SZ05, Ten04, TIY97, TEGM03, WC02a, YM97]. Block-Based [MF11, TSS99, CAC96, DLT05, FWG98, SCC12, SZ05, YM97]. block-based/segment-based [CAC96]. block-class-specific [SO94]. Block-Level [KMM^{+}12]. block-matching [CS92, CHW04, MB00b]. Block-noise-free [MOT^{+}03]. Block-Oriented [LP75, PH11]. Block-Sorting [BW94b, Sch97]. blockcipher [CMMS17]. blocking [MZ91, VBK89]. blocks [Lak02]. blocksorting [BM03]. Bloom [FB16, Mul87, Sal18]. BMP [Cha95a, Cha95b, Lus94, Mia99]. board [TLLB06, lGPL^{+}00]. BOCU [SD02, SD06]. BOCU-1 [SD02, SD06]. Body [Eva79]. BOF [GHD^{+}14]. Bonacci [JMnRS19]. Book [Ano98, Ano00, Neu10, PE02, Nel92, NG96a]. Boolean [MI91]. Boost [Dub11]. Boosters [FNP08]. Boosting [AMPHTSM10, FGMS05, OG04, YL05]. Boston [Ano98]. Both [Suz12]. Bottlenecks [BKS^{+}18]. Bound [MSFZ08, DC03, MW06, MF94]. boundary [AF04, BL97, LA91, Sch06, vB97]. Bounded [Ban09, De 05, DC18, FHS08, LMJ^{+}21, TTW07, ZLX^{+}20, CR95a, KGR06, LJPE21, NMW98]. bounded-error [KGR06]. Bounding [MLP99a, Sea90, UUiN12, ZKØ10, ZK11, KY99, YG02]. Bounds [CBHC08, DRR10, Sad98c, SMFZ07, SC11, TW01b, Vas07, WLL^{+}20, WMWW10, AL81, CH06, FV20, GL03a, KM97a, Nat94, RTK^{+}16, RBR09, RBR10]. Box [Sea90]. Boyer [BPMA02, CFG12, NT05a]. BP [XXZ11]. Braille [Ame06]. Branch [CCM96b, FFW98, CCM96a]. Brandenburg [Ano06j]. Breaking [RSMA19, Sau94]. Breakline [Cha89]. Breakpoint [TBMM98]. Breakpoints [MZT12]. bred [Whe97]. Brief [Lyo09]. Broad [Hoc96]. Broadband [LK08, SMS91]. Broadcast [FWZ^{+}12, ITU91, KJ10, LJ07, CL06, HKA06, Tun04, ZE00]. Broadcasting [DVB06, FWI99, HM07]. Broadside [Pom16]. Brotli [AKSV15, AFF^{+}19]. BTC [Das95]. BTF [GMSK09, KCL18, RK09b, WDR11]. BTF-CIELab [GMSK09]. BTFs [GMSK09]. BTTC [PSM01]. buck [VN90a]. Bucket [CC19]. Buffer [AMAM13, Dos72, LC03, SY04]. buffer-constrained [SY04]. Building [GBK16, SS96]. built [RSK88]. built-in [RSK88]. bulk [PM98]. Bull [Sie90]. Bunka [Ano03]. Buri [ZLC^{+}15]. Burlington [IEE96]. Burrows [Neu10, Abe05, Abe07a, Abe10, ABM08, BKS99, BK00, CT98a, Cha00, CJDL09, CHSS08, Deo00, Deo02, EV14, Eff99, FTL03, Ghi05, KLV07, KM99b, Kül12, MRS05, Man01, NSA06, NT20, NZ06, Sad98a, SI99, Sad99, WM01, ZMAB03]. Burst [EKT12, PS05]. bursty [BAL02, RG02]. Buyanovsky [Lam99]. BWT [Abe05, Abe07a, Abe10, AMB^{+}02, Arn02, BB05, BPMA02, Eff00a, IM01, KKP12, LCV07, PP18, Sew00, WS05, ZIIK05]. BWT-based [BB05]. BWT-stage [Abe05, Abe07a]. BWT-transformed [AMB^{+}02]. Byte [Ant95b, TDS^{+}21, VN90a, LKR17]. Byte-aligned [Ant95b]. byte-oriented [LKR17]. Byte-Select [TDS^{+}21]. bytecode [LF05]. Bzip2 [AKSV15].

C [BKK20, GMC^{+}06, Nel92, PF^{+}88]. C-BDAM [GMC^{+}06]. CA [BBKF95, Sal13]. Cache [AW04, ALYK12, AS14a, CYL^{+}08, Dou93, FV16, KHHK21, MV16, Ric12, RS21, SSW16, VMP^{+}16, AS14b, AR18, FGG^{+}08, FG97, GAS16, TK07]. Cache-Oblivious [FV16]. cache-obliviously [FGG^{+}08]. Cacheline [PBL^{+}17]. Caches [LHK03, TSFS21, YNQ17]. Caching [JSC20, CG91a, CG91b]. Calculating [GHV05, MPL99]. Calculation [Koh72, LM02]. calculus [LPS02]. Calgary [Bro08]. calibrated [TFC99, Vov06]. California [ACM78, ACM03, ACM13, HWS06, HHG07, HHSS08, HHW05, HPV09, H^{+}10, HPT11, IEE95a, R^{+}06, R^{+}07, Sch03, SM01a, USE93, M^{+}98]. Calligraphic [BL97]. Calligraphy [CC01]. CAM [LW00]. Camera [MOT^{+}03, GD06, Sch13, vB97]. Cameras [WZW11]. Campbell [THG03]. Can [Ano92a, BKR^{+}93, FS03, WM01, WPXL09, ZW08, Hal94, VN90b, Zir07]. Canada [IEE04, LL08]. candidate [Sal91]. Cannes [ZD81]. canonical [Nek00]. Canterbury [BPHA06]. capabilities [MR91]. Capable [MOT^{+}03]. Capacity [DRT81, SMLS15, YNQ17, GL03a, VBK89, YG02]. Capocelli [Ano04a, Ano07a, Ano08a, Ano09b, Ano10, Ano11a, Ano12a]. Capture [Hal11, VB14]. captures [Sch13]. Capturing [KR97, Tsa91]. caractères [Har04]. cardiac [ASM96]. cardinality [CCH94]. cardiology [Ng10]. Cards [Bar00b, DQ97]. carry [XG97a]. carry-over [XG97a]. cartoon [WRCB02]. Cascaded [NR14]. Cascading [PM91, Sal18, CZ99, PMK91]. Case [AM96, AM98, AS14a, BGPW96a, EKT12, MSV98, PV06, RK15, ZBMM97, FV20, GN01, RMG05, Sad99, VKG01, WLCZ15]. case-insensitive [Sad99]. cases [CDL^{+}19]. Cassini [WWA^{+}95]. casting [KA95]. cat [Per21]. categorization [FCW00]. Cathedral [Ano00]. Causal [HKW10, SSP06]. CBC [HIMM20]. CCD [DSSR95a]. CCIGS [CY92]. CCITT [ITU89]. CCP [Ran96]. CCSDS [GVSSBRAL08]. CCSDS-122-B-1 [GVSSBRAL08]. CD [KBD89]. CD-ROM [KBD89]. CDF [Ara13]. Cell [JB74, LE94, JCS^{+}08, VG95]. Cellular [CCRCK09, Laf00, MCRKC09, BBC95, Hau12a]. Center [Ano03, G^{+}75]. Centimeter [PHG86]. Centimeter-band [PHG86]. Central [ZKØ10]. centroid [Bon95]. CEO [BS06, Dic92, ZK11]. Cepstral [BLS11]. Certain [Nyq28]. certificates [Kir04, Kir05]. certified [Che15]. Chain [LWZI12, Mor76, WK92, EA95, TPAK96, CL16]. chain-code [TPAK96]. Chain-Link [Mor76]. Challenge [Ano95, Bro08]. Challenges [MF98c, Jur91]. chance [Man77]. changed [Mac12]. Changes [PC08]. Channel [ARR10, BFCP09, CT09, ECBL12, GE07, HS07, Hoo78, HGFL09, HKW10, Ing08, Kar61, SL08, SL10, SF09, SGLT98, VSG00, ZKK09, ZZPB09, AJRK98, BHS06, BF99c, CF98, CF99, CRR98, CS05, CT98b, CLME04, DS98, DHS01a, FZ02, FR06, GFV97, GF01, GL03a, GKV99, GG04, GCS00, HR06a, HSX02, HOR05, HM93, LA05, LLZ96, LS03, LNA99, MZ95, ØJH05, PM99, PSH00, PPR02, QJRA99, Ruf95, SCM05, SLXG04, Tun04, WS05, WAL05, WBM03, WJBM05, WWW06, XSX05, XHS05, YH00, YG02]. Channel-Decoding [BH01, BH00]. channel-induced [GFV97]. Channels [BH07, BM11a, Cat84, EEJ07, ECBL12, EKT12, HS07, HB09, KY08, KJ10, QB09, Sch70, SC11, BBF02, BAL02, CL16, EGNA06, GF01, GKV99, HR06a, HKA06, KH04, LNA99, MZ99, MRL99, MR06, PBC05, PAB03, PCRLSB00, RRG97, SG92, SZ97, SV98, Tun04, Vos06, YP03, YJE04, ZAA00]. Character [Jak85, ABH08, BL97, Hos02, Sab94]. Characteristics [Li86, Lle87, SHK99]. Characterization [LT08a, Sha11, FC98, JCS^{+}08]. Characters [CW97, Har04]. Chart [Bus09, JFC12]. Chart-Based [JFC12]. Charts [AB10]. Check [WY10, WS05]. checker [TDG^{+}19]. Checking [KKJ11, LZWL11b, PSR04, BCR98]. checkpointing [IMB^{+}13]. chemical [JU10]. Chengde [CLA12]. Chennai [NDN13a, NDN13b]. Chervonenkis [War97]. chest [Che91]. Chile [BJZ94]. China [CLA12, WS12, Yan10]. Chinese [CY92, CW97, Gu05, WT05, YMYS99]. Chip [Bec93, LHK03, ABF^{+}91, ATL^{+}88, CC03, CDC96, CR07, Gov94, KLAH99, Sch13, Tsa91, VL91, ZRZ^{+}19]. Chips [Ano03, IEE95a]. choice [Moh02]. Choices [LD10, Bry95]. choosable [Maß15]. chosen [BH93b]. Chroma [Poy08]. chromatic [JB99]. chrominance [Won91]. Chunking [CNPC98, CNCC98, BBG10]. Chunkless [BRV08]. Churn [BSSU18]. CID [KHHK21]. CIELab [GMSK09]. Ciphers [CCF^{+}18, KHJ^{+}09]. Ciphertext [CCF^{+}18]. Circuit [Chu92, YS10, PGS10, Zhu02]. cities [XWC^{+}17]. claim [Ano95]. clarity [Den01]. Class [Kun82, SK10a, BBQ^{+}95, RS98, SO94, WO00]. Classes [MY16, Pig02, FM95a, Rub88, Sav03]. classical [HJW02]. Classification [BF99a, BD18, BJ95, CY92, Cha04, CC19, DSM07, DMS08, DP01, Eva79, G^{+}75, HU99a, HU99b, HOG04, Kra91, LJ07, MSC00b, BE96, BYJ20, CPG96, DC03, Goo03, GPO98, JM03, Kes91, LGO99, Li02a, OG93, OG04, OG05, OG06, SO02, TYD^{+}20, WB00, YWPG03, Fow97b]. classification/reconstruction [JM03]. Classified [BH93a, DK95]. classifier [Lin95a]. classifying [NYOY94, RKSM03]. classroom [Hop09]. cleaning [DKP03]. Clef [ACP04]. Client [YNXC07, CSM00, CSM01, CS02, FSY01]. client/server [FSY01]. clients [CCA02]. Cliff [SR91b, SC92]. Climate [HNC^{+}16, KKL16, BHT19]. clinical [Ng10]. Clique [FM95b]. Clock [EK16]. Closure [Jag90, MSW04]. Cloud [CPD^{+}15, LLT^{+}16, PBGA19, PB11, YK20, YZZ^{+}14]. Cloud-based [CPD^{+}15]. Clouds [HKKW13]. Cluster [BS17, JKV15, MPGMD19, ADCU08, HM07, UH96]. Cluster-Based [BS17, ADCU08, HM07]. cluster-skipping [ADCU08]. Cluster-to-cluster [JKV15]. Clustered [BSS09, SSP06, Teu78, JG05, VG95]. clustered-dot [VG95]. Clustering [CCKH21, CLS19, Hil75, XJ14, ABA14, Ham95b, HOG04, MSE17, MNBC18, MS96, NT20, OHRS21, Sno76, Sou95, TPY95]. Clustering-Based [CLS19]. clusters [BKR94, KNTG08]. CMOS [TFA93]. CMYK [NDPL00]. CNN [ZRZ^{+}19]. Co [KHHK21, GLM^{+}21]. Co-Architecting [KHHK21]. co-design [GLM^{+}21]. Coarse [LMSE18]. Coarse-Grain [LMSE18]. Cobalt [Yua02]. Cobalt-60 [Yua02]. Cochlear [BT93]. CoCoPIE [GLM^{+}21]. codages [Har04]. Code [Ach76, Ano09a, BMM99, BA83, DV01, Fra97, Gir99, GRG08, Gra53, Hea72, ITU89, ITU90, KM11, Kro00, MGC07, NM91a, Per03b, RS04, TW92, TW01a, TW01b, WY10, XWL03, YO91, Yam00, YK10, Ano06v, BP00, BS95, BR04, BH10, Bry95, CC00, CGS93, CX04, CMP99, CLR03, CLR05, DWW02, DW03, Lak02, LDK99, LY91, LCLX04, LS96, MP85, MS05, MSOY96, NC97, PSSS03, RS00, Ric86, Sav97, SRP04, SLXG04, TPAK96, Tja05, VM98, YSXZ04, YSXZ05, ZE01, ZE03]. code-decode [MP85]. Code-Words [Gir99]. Codebook [CKC09, CS94b, Ham95b, HC06, CGS93, CS93, Del95, DE92, LL96, RRG95a, SS92, Wak05, ZDY97]. Codebooks [BŁ01, BF95a, FA93, GK95, LG99, WSS94]. Codec [Coa06a, FWS^{+}08, Haf01, SP96, YNXC07, YGNM07, Ano08i, CSM00, Coa06b, GMNK06, HS98, KLAH99, LK91, TW96a, WHH^{+}99]. Codecell [EM02]. Codecs [Ano08j, RL08, SGLT98, WZMG07, BCP04]. Coded [Abo06, IFK12, JCF93, SR09, Abo95, BV96, BF98, DHS01a, FW91, GFV97, Hal04, Jac92, JT98, LA05, LMS94, LCLX04, LBS98, PBC05, SV98, TM91, TVS^{+}03]. Codelet [Mod03]. COder [ZCG98, AL10, Lam99, Lan91, MCRKC09, MO11, RLG09, Rez07b, WZMG07, ALSSMPBR06, BF99b, BHB98, Cam06, Chi99, KB06, KK01, KP97, Lin98, NPW97b, PMLA88, PM88, Prz00, RK95, Sha93b, SM01b, SN94, WG00, WLZ92, Wit97, WF94]. coders [CF99, DvOL97, JV04, sKJ01, RV94, SLR05, TEGM03, Yao03]. Codes [AKS10, Abr93, AN08, AM05, BH01, BF11, BY90, BP85, BRD10, Big08, Bis08, BP09, BKYK94, BV04b, BV05, BC10, BF95b, But95, CD92, CD94, Cat84, Cen09, Cm10, CKP85, Col85, DRR10, DB07, Fel74, Fen96b, Fen02, FK85b, FK96, Fre91, Fre93, FVP08, GvV75, GPV11, GL09, Gün96, HY10, HDK^{+}12, HL90, How98, Huf52, KTMS10, KN99, KK07, KMSW18, KBN08, LV02, LV03, LK96a, LTO11, LKR12, Maa94, MGL11, MA06, NM10, OG12, PSR04, PS05, Pig01b, Pig02, RJL11, Rez07a, Rub79b, Rud71, Sal07b, Sar10, Sav09, SK10a, Sie88, TV09, TY07, TWM95, TW01a, Tun67, Vit87, WV98, Wil70, WK92, YK11, ZZPB09, AG02, Abr94, AMM00, AJ96a, Adj06, AB08, BCSV06, BH00, BHS06, BF99c, CC03, CH06, De 94b, DS98, DHS01a, Dum06]. codes [FR84, GN01, GKV99, GG04, GCS00, Hag07, HCL99, HL02, KN05, KB03, LLN^{+}04, LS95a, LS91, LKH^{+}00, LTB04, LM02, LM03, LW04, LXG03, MW06, MZ99, MF98b, MPL99, MLP99b, MLMD03, MTK95, MT96, MA05, MBBA91, Nas13, NSA06, Nek00, Nov00, OS01, OS03, OG16, PSH00, RG02, RA03, RA04c, SAR00, Sai05b, DP91, SF05, Sav98, Sav99, SDJ04a, SW97, She92, SLXG04, Stu94a, Stu94b, TKR00, TGFZ03, TW00, WO01, WS05, WAL05, WV99b, WS06, Whi06, XS05, XSX05, XHS05, YKLK99, YQ05, YR02, ZE00, ZGF02, ZE03]. codestream [TBKM05]. codestreams [BWM03, WBM03]. Codeword [AMPF09, Eli75, LBS98, Abr93, AMM00]. Codewords [LK96a]. Coding [Abo06, Abr63, AJ95, ASH87, AJN08, AHK97, AR09b, ARR10, AR12b, Ams86, AsM^{+}10, ABS^{+}16, ALM11, ALMJR^{+}12, BZZ08, BZL^{+}12, BFT11, BF99a, BGBV09, BRALSSMP08, BRNMG^{+}12, BF12, BT93, BSS09, BF07, BK93c, BK93d, BG03b, BLS11, BS94, BFCP09, BC10, Buy94, Cap59, CCRCK09, Car97a, CT09, Cen09, CCKH21, Che99, CTD07, CJDL09, CTC^{+}09, CDW11, CSP05, Cie98, CW84, CA93, CA97, DX11, Das95, Dav66, D^{+}92, DMC^{+}09, DOK88, DR83, Dun80, DG17, ECM91a, ECM91b, Eng93, ECBL12, EC09, EJ07, EKT12, FWS^{+}08, FWS11, FLS20, Fen96b, FR95, FSL^{+}12, FK90, FDU80, FVP08, GKN09, GMR95, GVSSBRAL08, Gha87, GK88a, GK88b, Haf95, HDCH09, HCMGB^{+}12, HN07, Hou97, HV92c, HV92d, HV94a, HGFL09, HR80, ISO93, ISO00, Ing08, IK11, II02, Jay97]. Coding [JAC19, JFC12, JLL^{+}10, JCF93, KMH10, Kar61, Kar12, KTMS10, KKN07, Kid09, Kie04, KH08, KT09, KP95b, Kle89, KW01, KS08, KdF12, Knu85, KØZ08, Kra49, KT81, KU95, LZPB11, LLZ09, LZR^{+}10, LZWL11a, LZWL11b, LZX^{+}08, LZX^{+}09, LZM^{+}10a, LaJR81, Lan83b, Law77, Lee81, LLA08, L^{+}05, Lio91, Liu97, LT09, Lyn66b, Lyn66a, Ma96, MTA^{+}20, MWC01, MEO11, MŠH12, MZT12, MSFZ08, MCRKC09, MN09, Moa86, MZ92, MNW98, MS00, MT02, Mof19, MTD08, MKH10, MSC00b, MGBRMSS11, Nek07, NL80, Nov00, NK16, Nus76, ØZ07, PM17, Pas76, PFAK10, PM91, Phi92a, PB98b, Poe83, Ram85, RR08, RH96, Rea72, RP88, RN10, Rez04, Rez08, RC09, Rez11b, Ric79, Ric91, Ric03, Ris76, Rob97, RD82, Rub79a, SYP04, SQ16, Sal05, SL08, SEC01, Sar10, SR08, Sch72]. Coding [Sch98, SMW07, SS07, SCV11, Sha93a, SSP21, SMFZ07, SF09, Sha10, Sha11, SW12b, SCW11, SWW12, Sho08, SSRM09, Smi79, Sof05, SR05, SjWL08, TY08, Tan93, TWW^{+}10, TW96b, Uhl95, VMFG07, VO08, VK95, Vit89, WWSY07, WWY^{+}07, WL08, WW10, WJDZ10, WHW10, WSDTO11, WV98, Whi06, Wic90, Wic94, Wil70, WNC87, Wu95, XWLZ08, XW10, YHM20, YXPM11, YLZ11, YGNM07, YMK10, Yok91, YM08, YDZL11, ZW07, Zha11, ZWZ11, ZJW^{+}12, ZWZ^{+}12, ZB12, ZLRZ09, ZX11, ZL78, ARZG03, Abe99, Adj06, ARR01, AF04, AMS^{+}03, AECG93, Ano06k, ABMD90, ABMD92, AL01, Arn99, Arn00, ARH03, BG06, BK74, BDV00, BL97, BH93a, BW95, Bar04, BS95, BM97, BS06, BM98, BR04, Ben96, BH92, BH93b, BR06, BK94, BHS06, BZM98, BK93b, BBQ^{+}95, CM97, CF98, CF99, CL91, Cha04]. coding [CCV05, CF96, CCMW03, CK95a, CKW91, CKCW95, CP03, CPR03, CO97, CAC97, CKS95, CT98b, CDSB98, CIRM95, CDGO02, CLME04, CPP95, Cre96, Cre02, CF97, Cve98, Cve99, DZ03, DLT05, DB02, DB03, DC97, DH00, DVS^{+}14, DS98, DHS01a, DDL98, DTZZ12, DKCS95, DHS^{+}01b, DSV00b, DHS02, DK95, ECG94, Eff97, Eff98, Eff99, EFZ03, ENA02, ENA03, EANG04, EGNA06, EGCK98, EA95, FC98, Faw96, FM95a, FZE03, FK95, FGC93, FGC94, FG01, FR06, For99, FJ00, FS02, FA93, Fow98, Fow00, FW02, FJ96, FN93, FNK94, FU98, FH96, Gag04, GK99, Gha84, Ghi06, GFH00, GZV97, GK98, Gra97, GE05, GE06, GMNK06, HR06a, HIMM20, HSX02, HW98, HL05, HZZY02, HOR05, HLV94, HV91a, HV91b, HV92a, HV93a, HV96, Hsi01, HCD04, Huy94, ISO95, ISO97, ISO99b, ISO03b]. coding [JR02, Jak78, JM03, JYHC03, JWK98, KC06, KLMXL05, KCB98, KA99, KP03, KJP99, KWLG92, KP95a, KL03, KW95, KA98, KB74, KCS93, KMS96, KCR98, Kri02, KM95, Lak00b, Lak02, LS04, LLN^{+}04, LCB03, Lee99, LW99, LHW00, Li02b, Li02c, LN02, Li02a, LTB04, LMO97, LM01, LR04, LAPL07, LSC91, LPR97, LG01b, LA05, LLZ93, LZZ98, LCLX04, LRO97, LS03, LG78, LNA99, LLY00, LKA99, Mal99, Mal00, MK03, MSW04, MWZ04, MF96, MF97b, MS03, MYRD04, Maß15, MM97, MZ95, MSR04, MF98b, MF04, MAC^{+}04, MM06, MZ91, MB02, MSWB93, MNW95, MP95, MW99, MSYY98, MSC99, Muk04, MOK00, MR06, NYOY94, NZ95, NM91b, NT20, Nov98, NYC05, OSMY95, OR94, OWS98, OTW^{+}99, PM98, PO00, PX12, PRM97, PAB03, PMK91, PSH00, PB98a, PR99, PR00]. coding [PR01, PS93, QLQ11, uR96, RF00, RH97a, RMZG03, RMG05, RAFH92, RRG97, RB01, RLR04, R^{+}07, RH97b, RA04a, RA04b, Sab94, SV00, Sai04, Sai05a, Sai05b, SW06, SF05, SG92, SNR06, Say92, Say98, SMC99, SC01a, Sch03, Sch04, SA03, SDJ04a, SLDJ05, SH03, SH04, SCHB05, SH94b, SCM05, SZ97, SHK99, SSM02, SH92, SM98, SM95b, Tai95, TT94, TY04, TCOR05, Tar95, Tei98, TN96, TR91, TW96a, TC06a, Tom06, TK99, Ton90, TM91, Ton93, TFV05, Tsa98, TKRR02, Tun04, Tur95, TM97, Use96, Use03, Use05, VG06, VL97, VKG01, VJB^{+}95, VW98, WS00, WW92, WXCM99, WO01, WC02a, WLZ92, WS06, WKM^{+}98, WY91, WF93, WWW97, Wu99a, WJBM05, XWL03, XSX05, XHS05, XG97a, XG97b, YKLK99, YR96, YSXZ04, YW04, YS91, YM97, Yu04, ZF92, ZRR00]. coding [ZE01, ZFE04, ZWS92, van99, ITU02, Ma96, Ric06]. coding-optimal [AL01]. Coefficient [JZL^{+}12, KMH10, TCKM10, CL91, Cre98, QMCX19].

Coefficients [BLS11, CBHC08, Sha93a, The89, CDR93, Gul04, Hsi01, KA99, Lak98, LRO97, MM97, MS95b, OWS98, Sha93b]. CoGI [XZG15]. Cognitive [YZLH16]. Cographs [UUiN12]. Cohen [Ara13]. Coherence [Liu08b]. Coherent [ALYK12, MM00]. COIN [SLZ^{+}19]. Collaboration [RVG11a]. Collaborative [HM07, KCCW17, ZZZ^{+}12]. collapsible [AC06]. Collected [Hae96]. collection [CZZR19]. Collections [MSW10, NP16, Wic03, CW02, FGG^{+}08, HPZ11, Mat15, WL15]. College [Bar00a]. Color [AB10, ASJ20, CYH94, CV97, DS08, FH98, GK88a, GK88b, HGR11, IK00, KA12, LE94, MS12, Moa98, NDPL00, TSS^{+}98, TEA08, Tsa98, TPM20, ANT95a, AKF06, Bar90a, BF99b, CWYW91, DH97, DH98, DKCS95, ER87, JG05, JB99, MM97, RAFH92, TM91, TCCT02, VPL98, YS91, van02b]. color-depth [TCCT02]. Color-Filter-Array [MS12]. Color-Mapped [KA12]. Colorado [Sch04]. Colored [HKW10, Ing08, KØZ08]. Coloring [DSMJ07]. colorization [BT17]. Colour [AC07, LRMPGR01, ISO99b]. column [VV04]. columnar [JLJ^{+}20]. columnist [Pou93]. Com [Lia84]. Com-pu-ter [Lia84]. Combination [Faw96, Zha98]. Combinative [DZZ08]. Combinatorial [SDV11, CDDM05, DZ03]. Combined [Ghi06, PP10, SA93]. Combining [BE96, CG91a, CG91b, CGR91, KT02, OG93, OSI02, TH01, ZYF^{+}20, B^{+}13, Bun98, CDGO02]. Comm [RHC87]. Commas [ER78]. Comment [TM98]. Comments [Cap89, Dav66, RHC87, Sha75]. Committee [Ano07d]. Common [IP06, MJ75, BM99]. Communicating [PHG86, SV09]. Communication [AJN08, BKS^{+}18, BLSM19, Bi08, Big08, Che99, CDL10, Fou85, Gra53, Gün96, HAM17, ISO84, LXG^{+}18, Moa86, Nus76, Sha48a, Sha48b, UPR99, Yok99b, BAL02, Fre98, Gag06, HPV09, KB99, MB91, Mer91, Mot07, OS01, QMCX19, Sha94, Vos06, WAL05]. Communication-Aware [BLSM19]. Communications [Ano03, IEE88a, IEE11b, AJRK98, Buy02, Cra96, HWS06, HHG07, HHSS08, HHW05, H^{+}10, HPT11, LKH^{+}00, MDPM00, Sal05]. Compact [BA83, HAR10, LB90, NPV15, PM18, TKYS16, Tou06, Uhl95, ZWZ^{+}12, CR97, ZRZ^{+}19]. compactage [LG78]. CompactDFA [BBHK14]. Compacting [BS88b, Tin77, Pro93]. Compaction [Col87, Sam89, YYO04]. Compactly [ABF^{+}02, Dau88]. companding [Sim98]. comparaison [LG78]. Comparative [BYJ20, GGG^{+}09, Sai04, RBD13]. compare [Bry95]. Comparing [CW07, Ven17, BFMX03]. Comparison [AKSV15, Ber99, CGSS15, EW96, Gil08, HM76, LCV07, LG78, Oh10, TCP03, Wic94, WFG^{+}94, SM97, AHCI^{+}12, Aro77, DvOL97, Fow97b, GCBV95, GLI16, KFN99, LZ05, Lyn69, SS92, THG03, You98a]. Comparisons [MLDH15, Ano08d, JM97]. COMPASSION [CNPC98, CNCC98]. Compatible [Cen09, KTMS10, RAE16, SD02, SD06, CZAR20, MEMS06]. Compensated [FG07, GKSB17, GYM93, LMBN08, Man98, MWC01, WL08, BK94, Cha96, FWG98, HLV94, LPR97, Won91]. Compensation [CA06, LJZ^{+}08, LZM10b, VST11, CCV04, Cre97, Fow06, HLV96, KG98, MB00a, MF98a, MM99, NO95, WXCM99, Xue99b, Xue99a]. competitive [Say98, Wan95]. CompEx [PM17]. CoMPI [FSC^{+}11]. compilation [Duv88, GLM^{+}21]. Complete [FK85b, FK96, NTF06, Raj20, Sal98, Sal00a, Sal04, Sal07a, De 94a, De 00b]. Completely [Pea90, Sie12]. Completion [CXYQ12, MW10]. Complex [LLJ09, MDH06, SILN11]. complexes [GD02]. Complexity [BGK91, BGK95, BK92, CD09, CDL10, CI07, DF08, EC09, KMH10, KYPY98, LZ76, LO99, LV08, MGD08, MTA^{+}20, MHKK03, MHKK06, Moa98, Mun91, OM07, PO00, Sar12, Sau94, SH94a, SG09, SG12, Sho08, VRHL07, VRL09, WSS96, Wei15, ZZPB09, Zur89, AR06, BLO01, CSP05, Cre97, DS06, DB02, DYW^{+}96, Eff97, Eff00a, EGCK98, FGC93, Ger91, GS96, GMNK06, JLS04, LN02, LSC91, LLWC08, MAC^{+}04, Mon93, ND04, OWS98, RGF06, RCMS04, RCH04, Rub88, Sai04, Sai05a, SO02, TW00, Wu99b, YA13, YWB01, ZR02, ZE03, vK00, van02b]. Complexity-distortion [LO99]. Complexity-scalable [PO00]. Compliance [ISO95]. compliant [Use05]. Component [AH14, DF08, Fow08, Fow09, SHW^{+}01, KF02]. Component-Based [AH14]. components [Won91]. Composite [Ram85, ZL19]. Composition [AC09, DGG10, Liu08a, DGG09, LCY07]. Comprehensive [DUH^{+}19, HSZ17]. Compress [ABP09, BS88a, DC18, GGM12, Ano06q, BST96, GC16, LGLK05, NN97, VV04, HSZ17]. compress/uncompress [Ano06q]. Compressed [AF88a, AF88b, ANS12, AF18, BCG^{+}14, BS20, BA16, BN14, BBH11, BBHS12, BdBN12, CW07, CCKH21, CXZD12, CHSV10, CWRL11, CEA18, Cot85, DCP17, Deu96a, DG96, DG17, EBH^{+}16, EBH^{+}17, EBH^{+}18, EBH^{+}19, EY81, FT98, FMMN07, FV16, FT04, FWZ08, FJD11, GMV17, GMC^{+}06, GNF15, GO15, GHSV06, HLS^{+}04, HSV^{+}08, HWY10, HKS^{+}11, HPST11, JSCM17, Jez15, JGL02, KL16, KTSA99, Kid09, KNB10, KE11, KS05, KS06a, KS07, KS11b, KS16, Kül11, LHK03, LKK^{+}17, LWW10, LLT^{+}16, Lin14, LKR12, Maa94, Man97, Mia99, MHM^{+}01, MF10, MF11, NCB15, NR99, Nav01, NM07, NP16, Oh10, PBL^{+}17, PGW^{+}17, RNOM09, SSW16, SF11, SLM^{+}17, Sen06, Sha09, TF11, TSFS21, VCW09, WB00, YK11, ZZZ^{+}12, ZL11, dMGdC10, ADCU08, AB92, ABF94, AN10, ACM^{+}15, BCD98, BPMA02, BR04, BD18, BFG09]. compressed [BEGV18, BR05, Bod95, BB95, CR97, CL06, DSSR95a, DSSR95b, De 03, FT95, FGG^{+}08, FV20, Fra06, GR99, GKPS05, GLM^{+}03, HKS^{+}13, INI^{+}16, KLR^{+}95, KIHO95, KTS^{+}98, KS00a, KS02, KSS06, LF05, LSYKK16, LZ05, LT03, MN08b, MI91, Mos98, Mul87, NKP^{+}16, NKT^{+}01, NT05a, Nos99, OW03a, OS97, OTY^{+}97, PW03b, RCMS04, SNZBY00, SR93, SLZ^{+}19, TM04, TM05b, TM05a, Tur96, VFK15, WF01, WC02b, ZG02a, ZTSM05]. compressed-image [OS97]. compressibility [AN06]. Compressing [AA71, BMM99, BMH08, Che01, FLMM09, FG97, Lan83a, LF04, MSW10, Mof00, NMWO96, Pap99, Pie03, PV06, RTK^{+}16, Ric12, SY01, Tho92, VK17, Whi87, WMB94, WMB99, WB03, XZG15, YNQ17, AM01, Ano09f, Ber99, DLT16, DK02, ELZC84, FN93, RL02, Sav03, WL15]. Compression [AG02, Abo95, Abr89, Ach94, Ach76, AMPHTSM10, ABM08, ABR10, AOAAH17, AB10, AKSV15, AW04, AM96, AM98, AKL18, AAR11, Amj08, Ams86, ALYK12, AH14, ASLB20, AMAM13, AC07, And12, AM05, AM10, Ano86a, Ano91, Ano92a, Ano92b, Ano92c, Ano94, Ano08b, Ano08d, Api91, AL00, AP03, ACK08, AD09, ACD02, ABM10, ASJ20, AVE18, BLNK14, BGU96, BM90, BIP01, BC08, BKK20, BRP19, BFT11, Ban09, BJCO03, Bar90a, Bar00a, Bar81, BW10, BŁ01, BS17, Bas85, BA08, Bec93, Bee79, Beh08, BGPW96b, BGPW96a, BGPW97, BD82, Bel86, Bel87, BWC89, BCW90, BK93a, BW94a, BM98, BZ08, BT93, BMM02, Ben97, BSTW86, BDHJ04, BLSM19, BK94, BRD10, Bi08, BCC^{+}18, Bil20, BBZ14, BMG18, Blo96b, BKYK94, BGK95, BV04a, BM11b, BK90]. Compression [BKR97a, BJ95, Bow70, Boy91, BBH93, Bre87, BBOH96, BES07, Bro08, Bru69, Bry95, Bud76, B^{+}92, BW94b, BR07, BR10, CA06, CCF^{+}18, CDAM07, Cap78, Car97a, CWS00, Car20, CNPC98, CNCC98, Cat84, CD08, Cha89, Cha91b, CW91, CY92, CBK96, CW97, CC01, lCcYxFpW09, CKC09, CC19, Che77, CYH94, CCM96b, Che99, Che05, CFC05, CC06a, CN07, CYL^{+}08, CRN08, CWY^{+}10, CTC^{+}10, CZZZ11, CZT^{+}11, CXF12, CHB86, CWLS15, CMK^{+}16, CLS19, CST11, CM02, CGC97, Chu92, CC06b, CAL03, CDLW05, CD07, CDL09, CDL10, CDL13, CW84, CMQW90, CMQW93, CMQW94, CK06, CK13, CA97, CGC11, CH87, CRA10, CH09, CH11, CFGL12, Cre94, CMRS00, CN02, CK93a, CK94b, CK94a, CK95b, CV96, CKV97, CVK97, CV97, CW03, Dac01, DKB^{+}16, DØJJ09, DXS08, DUH^{+}19, Das95, DI04]. Compression [DI05, DRT81, Dav95b, DLY^{+}98, Dav64, Dav66, De 00b, De 11, DV01, DD98, D^{+}92, DGG10, Deg94, Del93, DY12, DOK88, DC18, DDT^{+}12, DP01, DGGP05, DFH^{+}19, Dio93, Dis88, DH18, DLOM08, DSMJ07, Dos72, Dou93, DF08, DJCM09, DL01, DB10, Dub11, Dun80, DPdS08, ECM91a, ECM91b, ECM95, ECM01, EEH17, EGLS21, EP65, EP67, Eln84, EY81, EGF08, Eva79, Eva98, EK16, Fed93, FC94, Fal73, FNP08, FFFV19, FM95b, FL90, FPR95, FP20, Fel74, Fen95, Fen96a, FB98, FNV11, FOF09, FG89, FSC^{+}11, FCL13, Fis95a, FH98, Fon81, FL98, Fou85, FM12, FM89, Fow08, FDH^{+}20, FMP83, FK85a, FK96, FNK94, Fra97, Fra67, FN09, Fry85, Fu03, FB16, FGM21, Gad91, Gag94, Gag97, lGAR03a, Gai98, GO84, GHLN18, Gan94, GWT10, GPV11, Ged14, GS12, GWSR12]. Compression [GG92, GW73, GPM02, GBL^{+}98, GA08, Gil08, GB09, GJ20, GRG05, GMC^{+}06, GK88a, GK88b, Gol92, GS85, GRG08, G^{+}75, GLS99, GGG^{+}09, GT93, GM99, GHD^{+}14, GYM93, Gün96, GMSK09, HU99a, H^{+}98, Haf95, Hah74, HB96a, Hal11, Hal94, HY10, HU99b, HV08, HSZ17, HLL^{+}20, HH99, HS01, HKM10, Haw90a, Haw90b, HCBMSS12, HTS^{+}18, HI12, HKMS08, Hil75, HB08, HF97, HF99, Hoc96, HŘŠ11, Hoo78, HP92, HC92, HGR11, Hou79, HKKW13, HV92b, HV92c, HV93b, HV94a, HV94b, HHLW20, HC06, HLC07, HSL^{+}20, HG07, Hua01, Hua03, Hua11, HNC^{+}16, HO19, IEE09, IEE10, IEE11a, IEE11b, ILRS03, Ign98, IP06, IW94a, Ise01, IKM82, IK00, IW94b, Jag90, JJM18, JA07, JBG17, Jak85, JMW09, JP07, JSS15, Jay97, JAC19, Jen99, JAL^{+}12, JCS^{+}17, JSC20, JLL^{+}10]. Compression [JTC^{+}12, JS99b, Jog73, Joh89, Jon88, JM97, JEK98, JR12, JGM^{+}20, KKS00, KPM17, KXPZ12, Kar12, KB93, KL07, KKN07, Kel02, KL01, KCL06, KWC^{+}16, KN11, KKMS99, KT09, KJ95, KBD89, KBN08, KHJ^{+}09, KMM^{+}12, Kno98, KCBM21, KA12, KLUZ08, Koh72, KL97, KB92, KMMV02, KT07b, KM99a, Kra91, KN10, KS06b, KKL16, KGG^{+}14, Kül11, Kun82, Kur83, Kut02, Kyl90, LN17, LMBN08, LBMN08, Lam99, LE94, LaJR81, Lan84, LK96b, LZ06, LZ07, LM00, LLPP19, Lea78, LE07, Lee81, LK08, LSH12, LKK^{+}17, LMSE18, LH87, LA12, Lev95, Li86, LF03, LAP07, LYC16, LL16, LJF19, LLS^{+}21, LV95, LW11b, LHD^{+}20, LMJ^{+}21, Lin78, Lin95b, Lin95c, LI08, LP75, LGK97, Liu97, LHS05, LP07, Liu07, Liu08a, LMH^{+}09, LPG10, Lle87, LE01, LE02, Log04, Log06, LMM11, LRMPGR01, LG78]. Compression [LS94, LLJ09, LLYH14, LXG^{+}18, LDZ^{+}20, LF01, Lyn66b, Lyn66a, LB81, Ma78, Ma96, MWCG07, MWLW09, MLDH15, Mah08b, MD95, MN08a, Mal07, MS12, Man97, Man98, MYS20, MEIS09, MWC01, Mar80, MdM67, MPASM^{+}09, MPAFF10, MMZ12, Mat96, MRS01, MOT^{+}03, Mat12, MP12, MJ75, MC13, MFCM96, MW97, MCM^{+}17, MSV98, MAS00, MKW^{+}06, MM03, MMB07, MNM04, MKK12, MN08c, MN09, ME10, Mis11, MPFL97, MMK98, McC03, MKM^{+}08, MV16, MVJ17, Miu07, MCL12, Moa86, Moa98, Mof90, Mof91, Mof97, MS00, MT02, Moh04, MF98c, MY16, Mor76, MAL10, MRS03, MRS06, MGC07, MOW08, Mul97, MCHAM08, MPGMD19, MP06, NMN99, NLW99, NACM08, NM91a, NM96a, NBIT07, NIB^{+}09, NR14, Nel91, NcC02, Neu10, NDPL00, NMWM94, NMW97b, NK20, Nga95, NS01, NT02, Nus76, OINC10, OSI02]. Compression [Oku98, OS90, Pae91, PM17, Pan93, Pan95, PMZ13, PP10, PK10, Pas76, PBGA19, PBEA95, Pec82, PBO^{+}15, PWS10, PGW^{+}17, PM92, PM05, PB11, PSR04, PS05, PM91, Per21, Phi92a, Phi92b, PV21, Pig99, PCD17, Poe83, PELA02, Pop05, Pop11, Por73, PR71, PP03, PSM01, PS06, PW86, Pro86, Prz98, PA97, Pyl90, QLH12, RJ91, Rai87, RT87, RS16, Raj20, Ram87, Ram85, RJL11, Rea95, Ran96, RYe00, RY01, RBD13, Rea72, Reg81, Rei94, RM21, Ric03, RKL16, RL99, Rob91a, Rob94b, Rob97, RPE81, RS21, Roe03, RC96, RH07, RS99, Ros92, Ros94, Ros98, RV93, Roy87, Rub76, Rub79b, RK09b, Rya19, Ryt02, SI99, Sad96b, Sad98c, SW91, Sal18, Sal13, Sal98, Sal00a, Sal00b, Sal02a, Sal02b, Sal04, Sal07a, Sal07b, Sal08, SM10a]. Compression [SA93, San89, SW17, SM96, Sar12, SW12a, Sau94, SH94a, SA92, Say96, Say00, Say03, Say06, SBE16, SYZO15, SD02, SD06, SYO94, Sch86, SSP06, Sch70, SYDR07, SP11, Sco02, SB06, Sea90, SDFH00, SS05a, SM14, SILN11, SS05b, SSP21, SAS10, SKF^{+}00, SMO98, Sho08, SHW^{+}01, SK09, SST^{+}12, SPS98, Ska98, Ski08, Smi79, Sof05, SCH^{+}14, SWWW15, SR16, Spi93, SMS91, Sta67, Ste81, SS82, Sto88, SR91b, SC92, SC93, SC94, SC95, SC96, SC97, SR97b, SH97, SC98, SC99, SC00, SC01c, SC02, SC03, SC04, SC05, SC06, SC07, SM08, SM09, SM10b, SM11b, SM12, SCxx, S^{+}99, SY18, Swa08, Tad20, Tan93, TSS99, TCN^{+}17, TSS^{+}98, Tao82, TDL^{+}19, Tar79, TGHL98, Tau00, TM02, THHS93, Ten00, Teu78, TR93, Teu01, Tha94, TS17]. Compression [The89, TRS03, TGFZ03, THM91, Tis87, TWMG93, TLR85, Tod89, TDS^{+}21, TWC^{+}09, Tra87, TMS02, TCCT02, TPM20, Tsu72, Tun67, USY17, Uhl96, UHB09, UUiN12, VCP09, VPS17, VS10, VN08, VPGB10, Vo07, Wak07, Wal91a, WKC94, WL13, Wan73, Wan08, WYM10, WW12, WHZ12, WHB16, WB17, WGZ^{+}18, WLL^{+}20, Wat94, Wav06, Weg07, WZY^{+}11, WGZW14, WSS96, WSS00, Wei15, Wel84b, Wel72, Wel99, WMNP12, WS09, WPA08, Wic90, Wic92, Wic94, WFG^{+}94, Wil70, Wil89, Wil91a, Wil91b, Wil06, WWA^{+}95, WNC87, WB91, WBH^{+}92, Wit08, W^{+}00, Wol99a, Wol99c, Won85, WK93, Woo00, WB82, Wu96, WS08, WN10, WZW11, WDR11, XDZ11, Xu06, Yac98, Yal78, YHM20, Yan03, YGSR01, YSM08, YS10, YYZ12, Yaz81, Ye01, YZ07, Yok96, Yok97, YKO98, Yor95, YTY12, Yu96a, Yu96b, Yua02, ZSP^{+}19, Zha90]. Compression [ZD95a, ZD95b, Zha98, ZL00, ZG02b, ZWZ^{+}18, ZRZ^{+}19, ZYF^{+}20, ZZS^{+}22, ZZL^{+}15, Zhe03, ZLT^{+}18, ZBMM97, ZL77, ZL78, ZdMNBY00, ZMVR14, ZWZW12, ZLX^{+}20, dMGdC10, dRV12, SM97, Str99c, ARZG03, Abb93, ABA14, Abe04, Abe05, AT05, Abe07a, Abe10, AS97, AK99, ANdlF04, ANdlF07, AAJ04, AGJA06, AF99, DS92, AH02, Ait86, ABF^{+}02, AKF06, ABH08, AB08, ASK12, ASM96, AFHU97, AL81, APC92, ASP03, ASG99, ALRT16, AALR12, ABF^{+}91, ARA91, Ano82, Ano86b, Ano88a, Ano88b, Ano88c, Ano95, Ano06a, Ano06b, Ano06c, Ano08e, Ano18, Ant95b, Ant97, ACP05, AC06, Ara13, AS14b, Arn96, AM97, Arn97, AB97, Aro77, AR06, ATL^{+}88, AC98, AC14, AGS96, AR18, Att95, AMBC19, ABG^{+}94, BGM99, BS04, BGR01a, BGR01b, BKV05, BW99, BK74, BC05, BT17, BI98, BPZ99, BHT19, Bak07]. compression [BKS99, BK00, BE95a, BE95b, BW98, BB05, BA06, BP00, Bar90b, BRM88, BMT91, Bas91, BFMX03, BSF16, BDS95, BM89, BMNM^{+}93, BM99, BCR98, BH92, BH93b, Ber71, Bet94, BNK93, BBC95, BDCC97, BKV89, BMA04, BBG10, BB02, BGK91, BHLY20, Bon95, BH10, BK91, BKR92, BKR^{+}93, BKR94, BKR^{+}95, BKR97b, Bör80, Bör18, BP98, BTLK18, BPT06, BWS95, BB93, BAGCGC^{+}06, BJ18, Bri91, BFNP10, BM96, BCP20, BB92, Bun95, BL98, BM03, BYJ20, CM99, CVDL16, CCO^{+}19, Cam79, CGSS15, CWZ99, CW01, CW02, Cap85, CDL^{+}19, CCD^{+}19, CPD^{+}15, CG91a, CG91b, Ces91, CPG96, Cha96, Cha87, CZ98b, CC03, CL91, Cha91a, CM91, Cha91c, CC93, CKW93, Cha93, CSCW00, CT98a, Cha00, CWYW91, Che91, CR91, CZS92, CCM96a, CR96, CCMW05, Che15]. compression [CMMS17, CZZR19, CPT^{+}20, CZAR20, Che00, Che06, CDC96, CKCW95, Chi98a, CB97, CB96, CW15, Cho96, CSM01, CS02, CO97, CO98, CCH94, CAC96, CC98, Chu02, CT98b, CLD02, CLD04, COMF99, CLME04, CA02, CD95, Coo05, CR07, Cor85, Cor86, CBC00, Cov96, Cra98a, Cra98b, Cra98c, Cra96, Cre98, CL96, CLM14, CWW20, CD91, CDR93, CK93b, CKV93, Cus90, CS13, DF93, DPS93, DP76, DH97, DH98, Dav95a, Dav96, Dav93, DG76, De 94a, DS96, De 00a, De 05, De 06, DJL91, DGG09, DTZZ12, DHW^{+}17, Deo00, Deo02, DQ97, Dic92, DYW^{+}96, DC03, DVDD98, DKV07, Du00, DKS21, DJID20, DBTNT00, Duv88, DPS99, ECM98, EV14, Eff97, Eff00a, EEKB15, ER87, Eks96, ES98, ES00, EGS91, EFF00b, EHSC92, EN84, EW96, FSY01, FOSS17, FSC91, FS04, FFFM13]. compression [FS98, FFW98, FP19, FDKB11, FW95, Fen12, FK95, FGMS05, FGC94, FWA01, For96, FGGV04, FY95, FAEY95, Fow00, FRT96, FLMPR06, FM96, FCW00, FU98, FW93, FNI17, FWI99, GN79, GSS01, GD02, GAS16, GBK16, GD06, GCL06, GZ91, Ghi03, Ghi04, Ghi05, Ghi06, GSW02, GB98, GB06, Gil92, GK94, Gin95, GO95, GSPR19, GFH00, Goo03, Goo91, Gov94, Gra99, Gra91, Gu05, GLM^{+}21, GLI16, GP95, GO96, GS05, GB94, Haf96, Hag07, HM97, HIMM20, Ham94, Ham95b, HL97, HS06, HBSASA19, HAY06, Har96, HL05, Hat95, Hau12a, HJW02, HM87, HM91, HS97, Hes08, HWS06, HHG07, HHSS08, HS94, HV06a, HLV95, HM07, Hoc95, Hof97, HBL^{+}17, HSP^{+}13, HL00, HLR01, Hor91b, Hor95, Hos02, HV91a, HV91b, HV92a, How93, HV93a, How96, HV96]. compression [HZ95, Hua91, HAH05, HHW05, HPV09, H^{+}10, HPT11, HD88, Hus70, HSV04, HV06b, ISO95, ISO97, ISO99a, ISO99b, IW96, IYY^{+}11, IM01, Ish89, IMB^{+}13, IY00, IE91, JT96, JM96, JEEL16, JU10, Jak78, Jak88, JC95, Jia95, JB99, JdVL03, JLJ^{+}20, Jon91, JCS^{+}08, JB95, JKV15, KM16, VG95, Kam98, KLV07, KF97, KN90, KLJ12, KK03, KK04, KA95, KKFL14, KP03, Kes91, KR97, KAB11, KH00, KYNC96, KNY96, KYPY98, KP95a, KG95, KL03, KSCE16, KCL18, KV19, Kir04, KL05, Kir05, KG03, KF94, KF95, KF96, Kle89, KS00a, Kli00, KB99, KB74, KS91, KLAH99, KK08, Kop97, KF02, KF03, KS89, KRT05, KT07a, Kor67, KGR06, Kra10, Kru92, KM97b, KM98, KM99b, Kül12, KMP05, KNTG08, KPY96, Laf00, LTZ91, Lak00b, LLW^{+}14, LGS92, Lan98]. compression [Lar96, Lar98, LM99, Lar99, LOS20, Le 91, LW14, LKRR91, Lee91, Lee00, LJ05, LL02, Leh02, Lel91, LN91, LH91, LKG12, LBK16, LKR17, LMV03, LC03, LSD97, LGO99, LCT03, Li10, LW21, LCJP21, LBH91, LDK99, LV96, LW00, LCL03, LCY07, LW11a, Lit95, Liu91, LM91, LY91, LM08, LPZJ15, LJPE21, LRH^{+}21, LXG03, LV14, LLP96, LPS02, LNV97, Lop12, LOMSS05, Lu92, LLY00, LLWC08, Luc00, Luc96, Lyn69, MSE17, MV20, MLS03, MKNG06, Mai95, Mäk89, MZP06, MF97a, MEMS06, MNBC18, MPL02, MR91, Mar93, MR93a, Mar02, MF96, MF98a, Mas91, MRS99, Mat15, May99, McG93, MP85, MMS91, MS95a, MBB08, MS05, Mer91, MT98, MASC98, MLP99a, MM06, Mil95, MI91, MPC03, Mis91, MBWP03, MK06, MB02, Mod03, MZ92, MSZ94, MS96, MS95b, MRG99, MS97, MM05, MSYY98, MM99]. compression [MMA00, MM00, MSOY96, NP00, NAIP03, Nad83, NLW95, NL03, NYOY94, Nam91, NY98, NEGF96, Nat93, NSA06, Nel89, Nel92, NG96a, NMW97a, NMW98, NR95, Ng96b, NC97, NPW97a, NPW97b, NL06, NSFO^{+}99, dlCNLA^{+}95, NT05b, NYC05, OG93, OS96, OT03, Oka05, OM03, OY06, OW03b, OHRS21, OSC14, PHM^{+}14, PM98, PST97, PZ12, PGS10, Pat94, PW03a, PKG08, PWM18, PA07, PH91, Per88, PMK91, Pes07, PM18, Pig01a, PN91, PHG86, Pou93, Pow93, PSSS03, PMH95, PB91, Pu06, QZH10, QT03, QZ14, RL04, RK93a, Ram91, Ram92, Ran88, RMB91, RKB06, Rat92, RTT02, RKSM03, RSK88, Reg90, RY92, RS98, RGF06, RSMA19, Ric86, RK93b, RSV^{+}00, RK15, R^{+}06, R^{+}07, RDDD96, RCMS04, Rob91b, Rob94a, RC98, RC92, RS04, RL09, Rub88, RBR09]. compression [RBR10, Ruf94, RRG95b, RHC87, RA04b, Ryt03, Sab94, Sad98b, Sad99, SOI00, Sad96a, SST^{+}98, SS99, Sal91, Sam85, SW04b, SMOY97, Sau95, Sau96, Say92, Say05, Say12, SRKS95, SMC99, SJ94, SCKS97, Sch97, Sch03, Sch04, Sch05, Sch06, SV95, SK04, SO94, SPL20, Sha94, Sha91, Sha06, SC01b, Sha04, SH94b, SS03, SS04, SM11a, SRW^{+}04, SLZ91, SHD02, She90, SCC12, SHWH12, SGA^{+}16, SMRR04, SPS99, SGD05, Ski05, SN96a, Sno76, SLJ^{+}15, SSZZ14, Sou95, Spa00, SK99, SKLA12, Sta09, Sta07, Sta94, Ste91, SRG^{+}21, SS78, Sto83, SR91a, Sto96, SR97a, Sto98, Str99b, Sub99, SM95b, SZM03, Swa04a, Swa04b, SDV11, Szp91, TKR00, TKR03, Tad98, TFRH98, TVLS08, TD91, TT94, TBMM98, TPAK96, TK07, TCP03, TCOR05, TDG^{+}19, Tat94, TR98, Taw93]. compression [TN95, TO04, TWCA02, TLLB06, Tho91a, Thy10, TIY97, Tin89, TFC99, Ton93, Top96, Tro96, TMM96, Tsa91, Tsa98, Tse05, TC06b, TYD^{+}20, TR88, Tur75, Ude93, UH96, VT95, VPL98, VDPL00, VO91, VW94, VB14, VD18, VL97, VN90a, VN90b, Ven17, VL91, Vil93, VBL94, VED96, VK91, VK96, Vol02, Vol04, Vos06, Vov06, Vuc06, Wak05, Wak06, WRCB02, Wal91b, WSS94, Wan95, WCH03, WLZW04, WBS05, WLL06, Wan06, WSK^{+}11, WLH^{+}17, War97, WC02a, Wei90, Wel84a, WS91, Wil00, Wil02, Wil96, WM01, WC88, WBN91, WBM^{+}94, WBE^{+}94, WBMT99, Won91, Woo94, WM94, WF92, WWW97, WCB97, WCM98, WT05, WLS06, WOS06, WZ91b, XTH09, XWC^{+}17, XWL03, XCK18, Xue99b, Xue99a, YA13, YL05, YS07, YWMS08, YSZ11, YZZ^{+}14, YBW00, Yok93, Yok98, YMYS99, YL99, You97, You98b, YS92, Yu95]. compression [Yu96c, YK20, YGM97, Zan92, ZIL93, ZDY97, ZZ98, ZO04, ZA05, ZZD21, ZCG99, Zha99, ZGF02, ZGB04, ZLN06, Zhu02, Zip91, Zip92, Zir07, ZC91, ZGS02, lGPL^{+}00, vB98, vK00, van02b, vB97, Hoc96, NACM08, Tou06, Weg92, ZASB95, Hau12b, Ano98, Nas13]. Compression-Based [DDT^{+}12, IW94a]. compression-boosting [YL05]. compression-compilation [GLM^{+}21]. Compression-Decompression [SHW^{+}01]. Compression-domain [McC03]. Compression-Expansion [PM17]. Compression-Induced [KWC^{+}16, LMH^{+}09]. compression-oriented [CKW93]. Compression-Then-Encryption-Based [ASLB20]. Compression/Decompression [Ano86a, LMSE18, PP10, SM96, VL91]. Compressions [Kla98a, Kla98b]. Compressive [EC16, Fow08, Fow09, FLZW10, KCCW17, LSW09, LLW11, VFK15, WWS11, WCY^{+}10, WFLZ09, WZW09, XJ14, YX12, YZLH16, ZCW^{+}11, ZYT^{+}15, ZHX^{+}19, LGM21, LLWD14]. Compressive-Projection [Fow08, Fow09]. Compressor [AR08, AFF^{+}19, BFL^{+}10, BR09a, BR09b, LS99, PH11, SKY10, SW12b, Tit02, Ano06m, LDM05, MR04, SZ05, XMi03]. Compressors [FNP08, ALM^{+}18, CZ99, Fen97, MM95]. CompressPoints [CEA18]. Compror [LL02]. Comput [De 00b, RBR10]. Computation [Giv58, HAM17, HDK^{+}12, PP18, WMWW10, Zur89, Ben10, DS96, LNhCW16]. Computational [CSF77, Cho05, Luc96, Sab94, ZR02, CSP05, Sai04]. Computationally [SR05, VST11]. Computations [RG88]. Computed [MGBRMSS11]. Computer [AFL96, Aro77, Bar00b, GYM93, Hae96, IEE84, IEE91, IEE96, Knuta, Ma96, Pet80, PD07, Sal99, Sal06, Sam90a, SDS95, SDS96, Tan02, Ano09f, BBQ^{+}95, Dav12, KT13, Sal05, YGM97, vL94, Knu86d, Pin99]. Computers [Knu86e, M^{+}98, Mac12, Kro00]. Computing [ACM78, ACM79, ACM95, ACM13, AKL18, Cha66, CCA02, CG04, CIS08, PF^{+}88, SCW11, SCH^{+}14, USE90, USY17, Wol99a, YNXC07, ZLC^{+}15, AFL96, B^{+}13, CSM00, CLA12, LSD97, LN02, NDN13a, NDN13b, WGM88]. concatenated [LXG03, SAR00]. Concatenation [BCC^{+}18, NM91a]. Concatenations [NM96a]. Concealment [CXZD12, CHMW10, KØJ^{+}12, LZZ^{+}12, ZXM^{+}10, CCA99, CCV05, Niu02, YML97]. Concentric [NM91b]. Concentric-shell [NM91b]. Concept [Hil75]. concepts [CG91a, CG91b, Lop12, vL94]. Concise [Hau12b, Sal08]. concordance [BKR92, BKR94]. Concordances [BKR97a, BKR^{+}95]. Concurrent [DCP17, EGS91, SCW11, Sav03]. Conditional [Gha84, Gha87, Say98, WWW97, BS04, DK95, GDV03, Gra97, KKA02, SRKS95, WBS05]. Conditioned [CD08]. conditioning [LC91]. Cones [Ost35]. Conference [ACM78, ACM79, ACM03, BJZ94, BBKF95, CLA12, FLA^{+}03, IEE88a, IEE95b, IEE95c, IEE04, IEE05, IEE09, IEE10, IEE11a, IEE11b, KRW^{+}93, M^{+}98, NDN13a, NDN13b, SM01a, SR91b, SC92, SC93, SC94, SC95, SC96, SC97, SC98, SC99, SC00, SC01c, SC02, SC03, SC04, SC05, SC06, SC07, SM08, SM09, SM10b, SM11b, SM12, SCxx, USE93, WS12, Yan10, ZD81, BDLS96, M^{+}98, YR87]. configurable [NYC05, XMi03]. configuration [PH90]. Configurations [Fre61]. Conflict [RVG11b]. conjecture [Ano06l, Fen02]. Conjectures [Cal06a, GL04]. Conjurer [Per21]. Connected [Gad91, Kla98a, Kla98b]. connections [KC91]. Connectivity [Ros98]. conscious [HAY06, IA00]. Conservation [ACD02]. conserving [LGM21]. consideration [CZ99]. Considerations [Cra96, KBD89, HM87, HM91]. Consistency [LZX^{+}11, PSR04]. Consistently [LSYKK16]. Consortium [Uni07a]. Constancy [TEA08]. Constant [Ric10, Sch70, Aus99, MRG99]. constant- [MRG99]. Constrained [AR12a, CCG96, Cie98, HHLW20, KJP99, MSFZ08, RRG95a, SMFZ07, ZMVR14, Abo95, AMS^{+}03, CR95b, EM02, FZE03, FDKB11, FW91, GL03a, GFH99, GL01, GL03b, HN02, JBF97, OS01, OR94, RV94, RMG94, SRKS95, SY04, Wu92, WCB97, YM97, ZFE04]. Constrained-storage [RRG95a]. Constraining [LF95]. Constraint [SG12]. constraints [GG06, Mil95, Ort96, SH03, SO02]. Constructing [FVW10, HC92, LP95, TW01a, MLP99b, OG16]. Construction [BLO01, CD92, DHN85, EK16, Fre93, HY10, Huf52, LV03, McC76, NZC09, OG12, OM07, Rud71, Wan08, YSZ11, ZN08, ZL19, FRT96, HC98, KN05, MTK95, PR99, DP91]. constructions [CPR03, PPR02, SRP04]. Consumer [Jur92]. Consumption [CNW^{+}17, KL07, MW10]. contact [LS04]. contact-based [LS04]. containing [Ano09d]. Contaminated [TV09]. Content [BBG10, DH18, FJ00, JOC97, JEK98, OS97, PMZ13, WHZ12, WGZ^{+}18, WZY^{+}11, XSW07, ZJW^{+}12, BWB^{+}20, Bet94, CAC97, DS06, DHH04, KL03, NcC02, RKSM03, TBKM05, VL98]. Content-adaptive [JOC97]. Content-Aware [DH18]. Content-Based [JEK98, WZY^{+}11, OS97, XSW07, CAC97, DS06, KL03, VL98]. Content-dependent [BBG10]. contents [CHW04]. Contest [Ano91, Bar00a, Nel91]. Context [AF04, Aus98, BJCO03, Blo98, CD10, DXS08, FVW10, FGM21, GWSR12, KF02, Kül11, MWC01, MRH11, MW97, Mof91, OHSS12, Prz98, RT95, Tar95, TWMG93, TLR85, VNHB12, Vol97, WZMG07, WSS96, WST95, Wu95, Wu99a, Yok96, Yok97, Yok98, ZW12, AS97, ASS98, AF99, AKF06, BW98, Blo96a, Bun97c, Bun98, CO97, DHH04, DDL98, Eks96, ES00, FWA01, HS06, HC98, Hsi01, IY00, JV04, JWK98, Lar98, LH91, Lin98, SOI00, TVW97, Von04, WBS05, WCB97, WCM98, Wu99b, YWPG03]. Context-Adaptive [WZMG07]. Context-Based [MWC01, MW97, WSS96, BJCO03, Prz98, TWMG93, AF99, CO97, JV04, Lin98, WBS05]. Context-Free [CD10, GWSR12]. Context-Tree [Vol97, WST95, TVW97, Von04]. Contexts [CTW95, CT97, AdlPN04, SG04]. contextual [GK94]. contiguity [EM02]. continued [Sha48b]. Continuity [JTC^{+}12]. Continuous [BW10, CWS00, Lew95, Roe77, Suz12, YNXC07, Hil91, ISO95, ISO97, ISO99b, KCR98]. Continuous-Tone [CWS00, YNXC07, ISO95, ISO97, ISO99b]. Contour [TW96b, WF01, Yal78, Tur96]. Contours [Wil70]. Contraction [WW12]. contrast [SWA94]. Contribution [BSG10]. Contributions [Pig01a]. contributory [GYLC09]. Control [CDW11, HL17, ISO84, IFK12, Lyn66b, MWCG07, Ran96, WKM^{+}98, ALSSMPBR06, Cap85, CB04, CB97, DLT16, FE06, ITU06b, HLV97, Hoa99, KKA02, KH04, LZ05, Lut88a, Lut88b, VD18, Yu04, ZG02a, lGPL^{+}00]. Controlled [Dos72, MTA^{+}20, SC11, KG98, MR91, Ruf95, WS05]. conventional [LXG03, Lyn69]. Convergence [Fre91, Kom95, SPS98, LLZ93, SPS99, SP98]. convergent [She90]. Conversations [CWRL11]. Conversion [ITU94, CD00, CDL07, DDGV06, Sab94]. Converting [Kro00]. convex [CMW99]. Convolutional [HH08, KWLG92, LXG03, Nov00]. Convolutionally [YWC09]. Cool [Ano03]. Cooperative [SI99, FRT96, LS95a]. coordinate [LPR97]. coordinates [BG06]. Coptic [Uni07b]. Corasick [TM05b]. Core [DJCM09, Koh72, ILRS03, KSCE16, NYC05]. Cores [SHW^{+}01, BL98]. corpora [SZM03]. Corporation [Uni03]. corpus [AB97, Bro08, TFC99, BPHA06, HŘŠ11]. correct [MS06]. Correcting [BF95b, But95, LV02, Pet80, SK10a, TICH98, AB08, Hag07, ITU94, LS95a]. Correction [DSSR95a, Dun81, Mil81, Pet81, RS81, ZW12, BEQ98, MRL99, RKSM03, Sun16]. Correlated [EGF08, FSL^{+}12, GE07, HS07, LTO11, RR08, SR08, SSRM09, SV09, TOL11, WWSY07, CS05, GF01, KT02, SRP04, TGFZ03, YQ05, ZGF02]. Correlation [BZL^{+}12, DMC^{+}09, LZPB11, Sal02a, WPXL09, JB99, PKG08, VSG00]. Correspondence [GW73]. Corrigenda [EP67]. Corrigendum [RBR10]. corrupt [BWM03]. Cosine [ANR74, BYR06, CL97, CSF77, CY91, FL90, KB93, Liu97, RY90, SFT03, Sha75, SM95a, Str99a, Wat94, Wic94, WFG^{+}94, KC91, MW92]. cosmology [M^{+}09]. Cost [DUH^{+}19, FLS20, PC08, SSW16, WMNP12, Zur89, CPT^{+}20, HLV96, SG92, SR97a, Tja05]. Cost-based [DUH^{+}19]. costs [Dum06]. CoTe [Swa08]. Could [SW17, PWM18]. count [Abe07a]. counts [NdR04, OSMY95]. Courbe [Pea90, Sie12, Sie90]. Course [Ken61, LN91, Reb12]. Covariance [RKL16, YWPG03]. Cover [LMC05]. COVQ [PCRLSB00]. CPM [LA05]. crack [Bro08]. Cracovie [Sie90]. CRC [RG88, Wil93]. creating [KS89]. Creation [Hal11]. CREW [ZASB95]. Criteria [BKRSR09, Suz11, Suz12]. Criterion [Dun80, Lev47, ENA02, ENA03, TKRR02]. criterions [LC91]. Critical [HLL^{+}20]. Crochemore [JMnRS19]. Cropping [SDFH00, FDKB11]. Cross [CD09, Dio93, RMG94]. Cross-Complexity [CD09]. Cross-Platform [Dio93]. Crossing [Cha89]. Crowned [Bar00a]. Cryptanalysis [TG17]. Crypto [MM06]. Crypto-compression [MM06]. Cryptographic [Mun91, Rub79b]. cryptographically [SK04]. Cryptography [Big08, Ben10, Way09]. Cryptology [WS12]. Cryptosystems [ZSP^{+}19]. CS [BP00]. cSHB [NCB15]. CSI [PR11]. CSU [SV95]. CSV [YYO04]. CT [LKRR91]. CTW [OSI02, Suz96]. Cube [Kla98a, Kla98b]. Cube-Connected [Kla98a, Kla98b]. Cubic [PW86]. Cuckoo [PS12, BJ18]. Culture [Zei93, Zei94]. curiosities [Ano08b]. Curve [Pea90, Sie12, Yal78, BS98]. Curve-To-Curve [Yal78]. Curves [Col87, HGFL09, PLF91, Sag94, Sal06, CT98b, FOSS17, HCD04]. curvilinear [Tro96, TMM96]. Custom [GRG08]. Customized [VW94]. Customizing [ZBMM97]. cut [MSC00a]. Cutset [RN10]. Cutting [CC19, GM99]. Cycles [Kla98a, Kla98b]. cyclic [GPV11]. Czip [HNC^{+}16].

D [BIP01, BA04, HESF11, LLM89, MSR04, RHC87, Gol06, Adj06, AH02, ASG99, AAR11, BF07, BPT06, CPD^{+}15, CD00, CDL07, DDGV06, FOhC09, FO10, GVSSBRAL08, GSW02, GB98, GB06, Hal11, JTC^{+}12, KC91, KL97, LKRR91, LK91, LAP07, LLS^{+}21, LA96, LP07, LM08, LMH^{+}09, LCG11, MLDH15, MZP06, McC03, MOK00, NC97, PST97, PK10, PA07, Sal13, SCV11, SSM02, TCP03, TRS03, TFV05, TMM96, TCCT02, TPM20, VDPL00, WJDZ10, XD11, XXZ11, YKLK99, YL05, ZWZ^{+}12, ZGB04, Z^{+}13, BSS09]. D-image [GB06]. D-optimized [WHH^{+}99]. D-pattern [ASG99]. D-SPIHT [TCP03]. D-TV [Z^{+}13]. D/ [MSR04]. DACLIC [ZCG99]. DANN [RK93b]. danych [Swa04b]. Dasarathy [Ma96]. Data [ACM03, Abr89, Ach94, Ach76, ABM08, AFF^{+}19, AM96, AM98, AA71, Amj08, Ams86, ALYK12, Ano82, Ano88a, Ano88b, Ano88c, Ano91, Ano92b, Ano94, Ano06a, Ano06b, Ano06c, Api91, AR18, AVE18, ABG^{+}94, BM90, BC08, BKK20, Ban09, Bar81, Bar90b, BW11, BS17, Bas85, Bec93, Bee79, Beh08, BGPW96b, BGPW96a, BGPW97, BMNM^{+}93, BT93, BMM02, Ben97, BSTW86, BM99, BH92, Bi08, BKV89, Bil20, Blo96b, BJZ94, BGK95, Bör80, Bow70, Boy91, Bre87, Bru69, Bud76, BL98, B^{+}92, BW94b, BR07, BR09a, BR09b, Cap78, Cap85, Car20, Cat84, CD08, Cha89, Cha91b, CY92, CW97, CCKH21, Che77, CCM96b, Che99, Che05, CHB86, CMK^{+}16, CWRL11, CM02, Chu92, CDL09, CW84, CGC11, Cor85, Cor86, CH87, Cot85, Cre94, CMRS00, CK94a]. Data [Cus90, DXS08, Das95, DI04, DI05, DRT81, Dav64, Dav66, DG76, De 00b, De 11, DD98, DJL91, DGG10, DOK88, Deu96a, DG96, DC18, DP01, DFH^{+}19, DH18, DVDD98, DLOM08, Dos72, Doz91, Du00, DB10, ECM91a, ECM91b, ECM95, ECM01, ER87, EP65, EP67, ETT15, Eln84, Eva79, Fal73, FLS20, FFFV19, FL90, Fel74, FOF09, FG89, FCL13, Fon81, Fou85, FM12, FM89, FDH^{+}20, FLMPR06, FM96, Fra67, FK09, FLA^{+}03, Fry85, Fu03, Gad91, Gag94, Gag97, GO84, Gan94, GPV11, Ged14, Gil92, GK88a, GK88b, GF10, GS85, G^{+}75, GGG^{+}09, Gün96, Hag07, Hah74, Hal11, Hal94, HAM17, HY10, Hau12b, Haw90a, Haw90b, HL17, HM87, HM91, Hes08, Hil75, Hoc96, Hof97, HŘŠ11, HP95, HN07, Hoo78, Hou79, HHLW20, Hua01, Hua03, Hua11, HNC^{+}16, Hus70, IEE09]. Data [IEE10, IEE11a, IEE11b, ISO84, Ign98, IKM82, IW94b, Jen99, JAL^{+}12, JCS^{+}17, JSC20, JS99b, Jog73, Joh89, Jon88, KHCS08, KPM17, KKN07, KN11, KE11, KJ95, KHJ^{+}09, KS91, Koh72, Kop97, KMMV02, Kor67, Kra91, KM97b, KKL16, Kun82, Kur83, Lam99, Lan84, LW14, LE07, Lee81, LSH12, LKK^{+}17, LH87, Lel91, Lew95, Li86, LKH^{+}00, LAP07, LS99, Lin78, LY91, LHS05, Lle87, LE02, Log04, Log06, LG78, Lu92, LS94, LXG^{+}18, LF01, Lyn66b, Lyn66a, LB81, Ma78, Ma96, MWLW09, Mah08b, Mar93, Mar80, MdM67, M^{+}09, Mat96, MRS01, Mat12, MC13, McG93, MP85, MBB08, MKSS12, MMB07, MNM04, MKK12, MN09, MSM09, ME10, MV16, MVJ17, Miu07, MCL12, Moa86, Mof90, Moh04, MRS06, MP06, NP00, Nad83, NM91a, NM96a, Nel91, Neu10, NS01, NSFO^{+}99, Nus76]. Data [OINC10, OSI02, Oku98, OS90, PZZ^{+}22, Pas76, PBGA19, Pec82, PM92, Per88, PSR04, PS05, Phi92a, Phi92b, Pig01a, Poe83, Pom16, Pop05, Por73, PR71, PP03, Pow93, PW86, Pro86, RMB07, Ram87, Ram85, Ram92, RYe00, RY01, Rea72, Reg81, Ric86, RL99, RPE81, Ros92, Ros94, Roy87, Rub79b, Rya19, Sad96b, Sad98c, Sal18, Sal13, Sal98, Sal00a, Sal02a, Sal02b, Sal03, Sal04, Sal07a, Sal07b, Sal08, SM10a, Sam85, Sam90a, Sam90b, San89, SKY10, SF11, Say96, Say00, Say06, Sch86, SSP06, Sch70, Sch11, Sea90, SM01a, SAS10, Smi79, SCH^{+}14, Sta09, Sta67, SMB98, Ste81, SS82, Sto88, SR91b, SC93, SC94, SC95, SC96, SC97, SR97b, SC98, SC99, SC00, SC01c, SC02, SC03, SC06, SC07, SM08, SM09, SM10b, SM11b, SM12, SCxx]. Data [SY18, TV09, Tad20, TD91, Tan93, TCN^{+}17, TSS^{+}98, Tao82, The89, Tho92, Tis87, Tod89, Ton93, Tra87, Tsu72, USY17, UHB09, VMP^{+}16, VB14, VN90b, Vo07, Wan73, WLZW04, Wan08, WYM10, WW12, Weg92, Wel84b, Whi87, WS91, Wil70, Wil89, Wil91a, Wil91b, WWA^{+}95, Wir76, WNC87, XDZ11, Xu06, Yac98, Yal78, YHM20, Yan03, Yaz81, Ye01, Yok96, Yok97, Yok99a, Yok10, YS92, Yu96a, YDDB15, Yua02, Zha90, ZG02b, ZGF02, Zhe03, ZYT^{+}15, ZCJ^{+}20, ZLT^{+}18, Zip91, Zip92, ZL77, ZWZW12, ZLX^{+}20, dRV12, Abb93, Abe04, AT05, Ait86, ABH08, AB08, ABF^{+}91, Ano86b, Ano98, Ano08d, Ano08e, Ano18, Ara13, Aro77, ATL^{+}88, Att95, AMBC19, BGM99, BGR01a, BGR01b, BK74, BC05, BHT19, BKS99, BK00, BY00, Bar04, BB05, BA06, BRM88, BMT91, BM89, Ber71, Bet94, BBC95, Bli91, Bod95]. data [BKR^{+}93, B^{+}13, Bre04, BB95, BB92, CM99, CDL^{+}19, CCD^{+}19, CPD^{+}15, Ces91, Cha87, CC03, Cha91a, CM91, Cha91c, CC93, Cha93, Cha00, Che91, CZS92, CCM96a, Che15, CZZR19, CPT^{+}20, CDC96, CW15, Cho96, Chu02, Cov96, Cra98a, Cra96, CDR93, CW03, CKV93, CS13, DF93, DP76, De 94a, De 00a, DGG09, DTZZ12, DHW^{+}17, DY91, DJID20, DBTNT00, EL03, ELZC84, Eff00a, ES98, ES00, EFF00b, EHSC92, ENA03, FFFM13, FS98, FFW98, FWA01, FY95, FW02, FG97, FNI17, GFC^{+}09, Gin95, GRG05, Goo03, Goo91, GLI16, GHSV06, HIMM20, Ham94, HL05, Hat95, Hau12a, HJW02, HWS06, HHG07, HHSS08, Hor91b, Hos02, HV91a, How93, Hua91, HAH05, HHW05, HPV09, H^{+}10, HPT11, HM93, HD88, ISO98, Ind91, Ish89, IMB^{+}13, Jia95, Jon91, JCS^{+}08, JKV15, KM16, KLR^{+}95]. data [KN99, KLJ12, KA99, KIHO95, KA95, Kes91, KNY96, KYPY98, KP95a, KSCE16, KV19, Kle89, KB74, KRT05, KUCV04, KGR06, KPY96, LLW^{+}14, Lan98, Lar96, Lar99, Lee91, LAB^{+}14, LL02, Leh02, LN91, LH91, LSD97, LCT03, LW00, LCL03, LCY07, Lit95, Liu91, LM91, LRH^{+}21, Lop12, Lum13, MV20, Mai95, Mäk89, MT92, MNBC18, MRS99, MR00, MKW^{+}06, Mil95, MPC03, Mis91, Mis11, MBWP03, MB02, MM00, MOK00, NTF06, NAIP03, Nam91, NY98, Nas13, Nat93, Nel89, Nel92, NG96a, NR95, Ng96b, NT91, NHHS02, OTY^{+}97, OHRS21, PM99, Pat94, PWM18, PBEA95, Per21, PN91, PHG86, PB91, Pu06, QZ12, uR96, Ran88, RMB91, RKB06, RSK88, Reg90, R^{+}06, R^{+}07, Rob91b, Rub88, RRG95b, RHC87, Sad96a, SW91, Sal05, Sal91, Say92, Say05, Say12, Sch97, Sch03]. data [Sch04, SV95, SO94, Sha94, Sha91, She90, SGA^{+}16, Sno76, SLJ^{+}15, Sou95, Spa00, SKLA12, Sta94, Ste91, SRG^{+}21, SS78, Sto83, SR91a, SR97a, Str99b, Sub99, Swa04b, Szp91, Tad98, TFRH98, TK07, TDG^{+}19, TWCA02, Tha94, Tho91a, TIY97, THM91, Tro96, TMM96, TC06b, TYD^{+}20, Tur75, Use03, Use05, VT95, VN90a, Ven17, VED96, VK91, VK96, VFK15, Vol02, Vol04, Vos06, Vuc05, Vuc06, Wan06, Wei90, WS93, Wel84a, Wil00, Wil02, Wil96, Woo94, Woo00, WLS06, WZ91b, XWC^{+}17, YKLK99, YGSR01, YL05, YWMS08, YZZ^{+}14, Yok93, YYO04, YL99, Yu96c, Zan92, ZIL93, ZD81, ZC99, Zha99, Zhu02, Zir07, ZC91, Kro00, SC92, SC04, SC05]. data-aware [GHSV06, IMB^{+}13]. Data-Compressing [Tho92]. Data-Compression [Fel74, FM89, Mat96, Ros92, SSP06, Mäk89]. Data-Parallel [MKSS12]. data-processing [YYO04]. Data-type [AR18]. data/image [R^{+}06, R^{+}07, Sch03, Sch04]. Database [DDT^{+}12, LL08, RSWW01, RSWW02, RV93, SI99, Che06, Cor85, MB91, OW03a]. Databases [Bas85, CN07, IW94b, CWZ99, KF97, OS97]. Datacompression.info [Ano07b]. Datagrams [Sim92, Per89, Per90]. Dataset [LN17]. Datasets [FOhC09, VPS17, SHWH12, WB03]. Date [MF98c]. Daubechies [Ara13, Anoxx, AC14, KAB11, KKT10, RV09]. David [Hau12b, Nas13, Bar05, Pin99]. DC [IEE95c, QMCX19]. DCA [FH08, KCCW17, ZIIK05]. DCAST [FWZ^{+}12]. DCC [SC98, SC04, SC05, SR91b, SC93, SC94, SC95, SC96, SC97, SC99, SC00, SC01c, SC02, SC03, SC06, SC07, SM08, SM09, SM11b, SM12, SCxx]. DCDedupe [ZWZ^{+}18]. DCEs [ITU94]. DCLZ [ECM91a]. DCT [HESF11, Abo95, AG09, AKL18, BS95, Bel11, Bli93, BA04, CBHC08, CWY^{+}10, GK99, Lak02, LF03, Liu08a, LLM89, MS93, RL95, RB01, SFT03, SHK99, SWA94, The89, TB11, Wat93, WSA94, YXPM11]. DCT-based [GK99, RB01]. DCT-II [BA04]. DCTs [Kie99]. DDC [SC92, SM10b]. DDJ [Ano91, Nel91]. DDoS [LRH^{+}21]. De-Clustering [CCKH21]. De-duplication [BC08]. de-noising [SCC12]. de-quantization [Lak98]. Dead [BK93c, BK93d, DSG12, BK93b]. Dead-Zone [DSG12]. deadlines [FE06]. Deadzone [Ric11, AMS^{+}03]. Deaf [Fou85]. Deal [Bli93]. deblocking [DSSR95b, Gra03]. Deblurring [NK20, Xio11]. Debut [Bar00b]. decades [Per21]. Decaying [Muk04]. December [PBPT95, WS12, Yan10]. Decentralized [GV09]. Decipherability [McM56]. Decision [CCA99, Fon81, HH08, SMS08, XSW07, PCRLSB00, WO00, XHS05]. Decisions [HKKW13, PM99]. declarative [EBH^{+}19]. Decodable [Gir99, KPM16, Oka05]. decode [MP85]. Decoder [BP09, CD07, HS07, LZPB11, SL10, ALSSMPBR06, CMP99, ND04, WO01]. decoders [De 00a, De 03, FZ02, IC98, TM98]. Decoding [BH01, BT93, BF95b, DCP17, EC16, HMR10, HL90, HHLW20, HGFL09, HH08, KK07, KW03, Kle07, LSW09, LLW11, MTA^{+}20, Mar90, Nek00, Rez07a, Sie88, TH10, WY10, XS05, YWC09, ZZPB09, BH00, BK05, CC00, CCA99, CWL99, CS05, CMW99, Chu97, CME05, GFV98, GF01, GG04, GCS00, KK03, KW00, KT02, LA05, LHY12, LW04, MLMD03, MSC00a, MBBA91, Nov00, PM99, RG02, Ruf95, SJPB17, Sri99, SV98, TM98, VSG00, VRR^{+}16, WL99, WS05, WV99b, WWW06, YP03]. Decoding-Complexity [MTA^{+}20]. decoding/encoding [MBBA91]. Decomposed [LJPE21, JF96]. Decomposition [AH92, BKK20, Mal89, MP12, PS06, RK09b, The89, Che15, JLJ^{+}20, KF94, KGR06, Kri02, MPL02, SCC12]. Decompositions [BRP19, GO15, SM97, KF95]. decompressed [ML98]. Decompression [Ano86a, BWM03, BH12, CC06a, DJCM09, GMC^{+}06, JBG17, KCL06, KHHK21, LHK03, LMSE18, MKSS12, OBGB11, PZZ^{+}22, PP10, PBL^{+}17, SM96, SYO15, SHW^{+}01, Zhe03, CLR03, CLR05, EV14, FNI17, HB96b, ILRS03, JJM18, JKL13, LOS20, LW99, LPS02, OW03a, PD16, RH91, SHC^{+}16, SJPB17, VL91, WZL^{+}17, Wel84a]. Decompressor [SKY10]. Deconstructing [CNW^{+}17]. Deconvolution [LSW09]. decorrelation [VPL98]. decoupled [BJ18]. Decoupling [FGM21, HAM17]. Dedicated [OINC10, WWA^{+}95]. Deduplicating [SMLS15]. Deduplication [CPL09, CGC11, ZWZ^{+}18, ZYF^{+}20]. Deep [BLSM19, COMF99, HLL^{+}20, HBL^{+}17, MDPM00, PHG86]. Defect [MOW08]. deferred [Yao03]. defined [KL97, SYP04]. Definition [JLL^{+}10, Tur96]. Definitive [Ano00, Roe99]. Deflate [Woo96, Deu96a, KMA^{+}20, AKSV15, Fel03]. degradation [CDMW02, MRL99]. degraded [Tun04]. Degree [KM16]. deinterlacing [BM97, KSY95]. Delaunay [BR05, PELA02]. Delay [AsM^{+}10, DY12, MSFZ08, SMFZ07, CL06, DHS^{+}01b, GLL04, OS01, WLZ92]. delay-constrained [OS01]. Delayed [HMR10]. Deletion [BBHS12, De 05, Sto98]. Delicately [lGAR03a]. Delta [Bar81, KS07, Kno98, ØZ07, Sha09, SK10b, TMS02, ZWZ^{+}18, ZYF^{+}20, KSS06, OT03, SHWH12, TO04]. Delta-Sigma [ØZ07]. DeltaFS [ZCJ^{+}20]. Deltas [Yor95]. Demand [JRALMSS11, MMA99]. Demonstrating [SHW^{+}01]. demonstration [VJB^{+}95]. Demystified [Sym03]. Denoising [QWH12, SWW12, dRV12, CDDM05, Gra03, Gul02, LK96b, YSZ11]. Dense [HO19, KBN08, PH11, VO08, KLMXL05]. Density [BRD10, WY10, GPO98, MW06, RL04, WS05]. dental [Lyn69]. Denver [Sch04]. Dependence [Yal78, HSV04]. Dependencies [KNB10]. Dependency [Abr89, VV04]. Dependent [LCG11, ZWZ11, BKV05, BBG10, BCP20, LG01b, SH04, Zee98]. Deplump [BW11]. Depth [AMAM13, CCRCK09, LLLZ12, LW11b, LMH^{+}09, MZT12, PK10, VN08, DL01, KMS96, PHM^{+}14, TCCT02, Z^{+}13]. depth-image-based [Z^{+}13]. derailment [GK95]. derailment-free [GK95]. Derivation [IFK12]. descent [DW05]. describing [Yok98]. Description [AK08, BZZ08, BZL^{+}12, BC10, CTD07, Cho56, DZ08, Dum08, FWS^{+}08, FWS11, ØZ07, SL08, TMD08, WL08, YLZ11, ZKK09, ZB12, BDV00, Car01, CMW99, DSV00a, DSV00b, DWB04, DW05, FE99, GK98, HW06, JR02, JV04, KGK00, Li02a, LZZ98, MR00, ØJH05, PAB03, RF05, SAR00, SVS99, Sha04, TH03, TH04, TH05, VFE00, WG00, WS06, WWW06]. Descriptions [BZL^{+}12, BZ09, KØZ08, RR08, YWC09, ZKØ10, BS02, FDZ00, JA03, LMC05, PPR02, Sri99, SB97b]. Descriptors [SP11]. Design [Ano06e, ALMJR^{+}12, BGM99, Bar00a, CA06, Cap78, Car01, Cen09, CYL^{+}08, CZ99, DSV00a, Dum08, FBX91, Gin95, HV93a, HV94a, HHCH11, Hua01, KM11, KMSW18, KMA^{+}20, KJ10, Kro00, LV02, Lev47, LS93, Lin78, LBG80, MNM04, Miu07, Pro86, RMZG03, Rez08, Sam90b, SL10, SCW11, SLXG04, VMP^{+}16, Vit87, Vol04, WO01, WAL05, Zhe03, Att95, BDV00, BWB^{+}20, Bli91, BF99c, CGS93, CRR98, CZZR19, CX04, CN96, Del95, DWB04, DW05, FWG98, GFH99, GLM^{+}21, How93, IF06, KR01, KG95, KF95, LHW00, LW00, LL96, LCLX04, ME02, OS01, PR99, RMG94, SNR06, SHD02, TH03, TH04, Wu92, WF95, YSXZ04, YSXZ05, YSCK00, ZE00, ZE01, ZE03]. Design-for-testability [Vol04]. designed [Zee98]. Designing [BLNK14, HDK^{+}12, LLZ96, Dum06]. Designs [EK16, BF95a]. Detail [Prz98]. Details [lGAR03b]. Detecting [Pet80, WCH03]. Detection [Cha89, CRA10, MGD08, TTW07, Wil93, WC02b, Ait86, AOT13, CS02, HL97, KCR98, LRH^{+}21]. detectors [ABL99]. determination [AL81]. Determine [AKS10]. Deterministic [KN11, DTZZ12]. Detroit [IEE95b]. Developers [Ano95, Rod95]. developing [EGS91]. Development [BFT11, IBM88, JBI03b, KRW^{+}93, Pav06, VPGB10, YR87, Ber99, Ish89, SV95, Sle74]. Developments [MW97]. Device [GZS^{+}96, GHD^{+}14, Kra49]. Devices [HAR10, GLM^{+}21, JCS^{+}08, TG17]. DFT [MBWP03, RG02]. DHV [Hua01]. Diagnostically [GSW02, TWCA02]. DIBs [Knu98]. DICE [YNQ17]. Dictionaries [AF18, Cha91b, HZKL11, KS11b, DS92, Cha91a, CK96, GHSV06, KS02, LCT03, Mit01, PX12, Sav99, SR97a, YR02]. Dictionary [Ant97, BES07, CDL09, ECM91a, GRG08, HY10, HSV^{+}08, HKS^{+}11, HGR11, Jak85, KS11a, LM00, MRS01, NLW99, NK20, Nix81, PBGA19, Pop11, SHD02, WHB16, YS07, BDS95, BCR98, BB92, CLM14, De 05, FRT96, HS94, HLV95, HKS^{+}13, Jak88, LTZ91, LM99, LOS20, Luc00, Mas91, MRS99, NLW95, NC97, SGD05, SP98, Sto98, SZM03, Yao03, YSCK00, TFD06]. Dictionary-Based [BES07, LM00, MRS01, Pop11, Ant97, YS07, CLM14, LM99, MRS99, SGD05, SZM03]. Diego [ACM78, ACM03, HWS06, HHG07, HHSS08, HHW05, HPV09, H^{+}10, HPT11, R^{+}06, R^{+}07, Sch03, USE93]. Dies [Bar00b]. Diet [Ano92c, WS93]. diff [Per05]. diff/patch [Per05]. Difference [GMSK09, Mye86, SV09, DLT16]. Differences [BMM99, Per03b]. Differencing [KMMV02]. Different [SB85]. Differential [Bis94, Bri91, Fen95, FA94, Gai98, HM76, ITU90, MGC07, SA92, SS05b, AAJ04, AGJA06, ABF^{+}02, BBG10, SS03, SS04]. Differentially [CNS12]. Differentiation [QT03]. Differentiation-based [QT03]. diffused [HL97]. Digital [BGU96, Bar00a, BBH97, Boo99, BG03b, BS94, Bud76, Car20, CGC97, Deg94, Eln84, Gha87, GBL^{+}98, GW92, HS01, Hou79, HR80, IEE88a, ISO93, ISO95, ISO97, ISO99b, ITU91, KS00b, LST99, MOT^{+}03, McC92, Pan93, Poh85, QLH12, RJ91, RS78, Roy87, She95, TSS^{+}98, Yaz81, Yua02, Che91, CF96, ER87, FJ00, FS02, FNK94, Fre98, GSW02, Ham95a, Hof97, HCD04, ISO98, JS95, Jur91, KA98, Lut88a, Lut88b, Mil95, Nat94, PR01, Spa00, WAL05, van02a, vB97, DVB06]. digital-analog [WAL05]. Digitising [HR06b]. Digitized [BBOH96, Fon81, Bör80]. Dimension [HV06a, KH08, TWW^{+}10, DRG94, ECG94, KFSM04, Man77, War97]. Dimensional [BGK95, BES05, BES07, DZZ08, LF03, LP07, Mar80, AB92, ABF94, BCSV06, BS95, CL97, Cho96, DJID20, Esk95, GL03a, HV06b, ILRS03, Ish89, JAC19, KP97, LBMN08, LGO99, MBB08, PHG86, Sau95, Tro96, VED96, Vuc05, WXCM99, WFG^{+}94, WB03]. dimensionality [LW21]. Dimensions [Ken61]. Direct [Col86, BW95, BF95a, Fra06, SILN11]. directed [CC03, HS06, LF04]. Direction [MWLW09, DHW^{+}17]. direction-preserving [DHW^{+}17]. Directional [LZR^{+}10, MF10, XWLZ08, XW10, ZX11, FVW10, LM08]. Directly [Man97, ZZD21]. Directories [ZCJ^{+}20]. directory [Ski05]. Dirichlet [LPG10]. disappear [Per21]. Disappearing [Way09]. Discovery [SSP06]. Discrete [ACM97, BYR06, CSF77, CY91, DV01, FL90, FCL13, HHCH11, JMW09, Kar61, KB93, Liu97, OWW^{+}04, PW86, RY90, Rez11a, RV09, SFT03, Sha75, SM95a, Str99a, Suz12, Wat94, GB98, Hop09, JJM18, KC91, LM08, MW92, MM00, SKCB95a, SKCB95b, TKR03, Vil93, ZGS02, ANR74, CL97]. discrete-tone [GB98]. discriminant [OG04, Wu99a]. discrimination [WZL91]. Discriminative [Ged14]. DISCUS [PR99]. discusses [Dic92]. DISE [CLR03, CLR05]. Disk [Ano92c, Pou93, VN90a, VN90b]. Disks [KJ95]. Disordered [HTS^{+}18]. Disparity [VMFG07]. Dispersion [IK11, MNBC18]. displacement [CS92, LKRR91]. Display [Hal11, JB74, Sab94, WSA94]. DisplayCast [CBR14]. Distance [AMAM13, HDK^{+}12, Sal13, WN10, YTY12, KA12, LO99, ML98]. distances [AGMP13]. distinct [AALR12]. Distinguisher [CWW20]. Distinguishers [LJF19, SW12a]. Distortion [Ach76, ALMSS09, ALM11, BFT11, BH07, BRALSSMP08, Cap78, CWY^{+}10, CKS08, DRR10, DØG08, DSG12, ECBL12, FE06, GH08, HM96, JLJ08, KRK09, LZR^{+}10, LZWL11a, LZWL11b, LT08b, LMH^{+}09, LWZI12, MWCG07, NYJ08, NYJ09, SS06, SjWL08, TDL^{+}19, TMD08, VRHL07, VRL09, Vas07, Ven07, WWY^{+}07, WJDZ10, WXC^{+}10, WFLZ09, YJE04, YDZL11, dRV12, Ber71, Bod95, CCA02, CG03, CG04, DC03, GK99, GFV97, GO96, HSX02, HB96b, JG05, LO99, LBH91, LG01a, MR91, MWZ04, MW06, MYRD04, MOG05, Max60, MS02, RMG05, RCH04, TKRR02, VB14, WV99a, WV00, WM94, WZ94]. distortion-based [HSX02]. Distortion-Complexity [VRHL07]. Distortion-Free [Cap78]. Distortion-limited [HM96]. Distortion-Optimal [NYJ08, YJE04]. Distortions [DØG08]. Distributed [AJN08, ARR10, BR06, BLSM19, CS05, De 11, DMC^{+}09, DSMJ07, EGF08, FWZ^{+}12, FVP08, GvV75, GPV11, GD06, KLMXL05, KH08, KD18, KJ10, LZPB11, LJ07, LGK97, LXG03, MMA00, PR99, PR00, PR11, RVG11a, RVG11b, SI99, SF05, Sar10, Sar12, SR08, SRP04, Sha91, SCW11, VMFG07, WWY^{+}07, XSX05, YXPM11, Zha11, ZWZ^{+}18, ES98, GDV03, JT96, LCY07, Pow93, RMZG03, SNR06, TCOR05]. Distribution [ITU91, IK00, OINC10, Yu96a, ANT95a, CCH94, Kri02, SS04, Yu95]. Distributions [BM11a, Rez11a, Abr94, BCSV06, CCMW03]. Disturbance [MVJ17, Yan03]. dither [Kli00]. Dithered [AR09a, AR09b, ØZ07, FZE03, ZF92, ZFE04]. Dithering [KRK09]. Divergence [Fre93, LRH^{+}21, Rez05]. Diverses [LG78]. Diversity [MTD08, Ton90, CG03]. division [Huy94]. DjVu [H^{+}98]. DLT2000 [Cre94]. DMC [BM89, Bun95, YL99]. DNA [AOAAH17, BLC^{+}16, Car20, CWLS15, CLS19, FS04, GT93, GLI16, HCMGB^{+}12, HCBMSS12, KT07a, LY97, MYS20, MR04, QLQ11, RDDD96, Sal18, TKR03]. DNA-Based [BLC^{+}16]. do [WM01]. Document [BBKF95, H^{+}98, MPGMD19, NPV15, SO94, Ado01, ZD95a, ZD95b, BB02, IW96, LT03, TBMM98, ZDY97]. documentation [ISO09, ISO17]. Documents [Moa98, WGZW14, WMB94, WMB99, ANdlF07, BD18, BHLY20, DPS99, EA95, HS06, Mul87, Ten00, YYO04]. Does [Ude93, Den01]. Dolby [Dol06b]. Domain [AG09, BS20, CA06, CFC05, DMC^{+}09, DMS08, DG17, HKM10, KOO09, KNB10, LZPB11, LGK97, LCG11, MWCG07, PS06, SDFH00, SWGZ09, YXPM11, YKO98, ABMD90, CR97, Gra03, LPZJ15, MBB08, McC03, PN91, RCMS04, Sau96, VFK15, Wan06, WB00, Zir07]. Domains [SW17, ZBMM97]. dominated [YG03]. Donald [Neu10]. DonationCoder [Ano06d]. données [LG78, Pig01a]. Don’t [Ano86b]. Doppler [Sch86]. dot [VG95]. Double [Ano92b, BR07, BR09a, FOP12, QLH12, ES98, Hal94]. Double-Precision [BR07, BR09a]. Down [Hua01, SS05a, WSK^{+}11]. down-sampling [WSK^{+}11]. Downsampling [SK09, You98a]. DPCM [Abo95, Hal04, PG72, ST97, You98b]. DPCM-coded [Hal04]. DPCM/DCT [Abo95]. DRACO [SMLS15]. Draft [SD06, Ano06f]. DRAM [BLNK14, YNQ17]. DRAM/PCM [BLNK14]. Drawings [JB74]. Drift [GKSB17, LT08a, RB01]. Drift-Compensated [GKSB17]. Drive [Cre94, BKV89, Mac12]. Driven [JAL^{+}12, VS10, WYM10, WZY^{+}11, WMNP12, CMWM00]. Drives [Ano92b]. dropout [Mos98]. Dropping [LSC10]. DSP [Rei94, SHW^{+}01]. DSU [SV95]. DSU/CSU [SV95]. Dual [BLNK14, BSG10, BF07, CCV04, Fre91, LJZ^{+}08, LZM10b, MWLW09, TC06a, CCV05, LC03, WL99]. Dual-Direction [MWLW09]. dual-path [WL99]. Dual-Phase [BLNK14]. Dual-Tree [Fre91]. Dual-TreeWavelet [BF07]. d’une [LG78]. Duplicate [YYZ12]. Duplication [MVJ17, BC08]. during [HB96b]. Duxbury [Ame06]. DVB [DVB06]. DVI [Gol92, Tin89]. DWT [LHS05, Niu02, PZ12]. DWT-Based [LHS05, Niu02]. DXPC [Gai98]. DXT [HSP^{+}13]. Dyadic [BM11a]. Dynamic [BCC^{+}18, BFNP10, BdBN12, CFC05, CHMW10, CH87, DLY^{+}98, Fra97, Gag04, Gag06, GMC^{+}06, JSS15, Knu85, KU95, LLA08, LRMPGR01, LDZ^{+}20, MN08b, MRS01, NT02, NK16, PBL^{+}17, SHC^{+}16, VS10, Vit87, Vit89, WKC94, WB17, Yok91, Yok99a, Yu96b, YG03, ABA14, BS95, Bun95, CLR03, CLR05, Cor86, DS96, FV20, HLS^{+}04, LM01, MZP06, MEMS06, MRS99, MSZ94, OSMY95, PZ12, RS98, RK15, SR97a, RK93b]. Dynamical [Sha10]. Dynamically [NMN99, ZG02b, IYY^{+}11]. dynamics [GSPR19].

Early [HP95, PBL^{+}17]. Earth [KCS93]. East [Kes91]. Easy [Pae91, Ude93, SCHB05]. EBCOT [MRH11, Tau00]. ECG [AC14, BMA04, HIMM20, HHLW20, Tai95]. echo [ITU06b]. ECMA-151 [ECM91a]. ECMA-159 [ECM91b]. ECMA-222 [ECM95]. ECMA-321 [ECM01]. ECOMP [Kyl90]. Economical [ER78, McC76, MPL99]. economically [AL81]. ECVQ [Car00]. EDAs [Tou06]. Edge [CK94a, HL97, JA07, JR12, LZR^{+}10, Lee99, LSW08, MPAFF10, Roe76, RD82, TS17, Yaz81, AMBC19, FC98, HB99, LW11a, MKNG06, Maß15, PHG86, Sun16, WRCB02, Xue99b, Xue99a]. Edge-adaptive [Lee99]. Edge-Based [JA07, LSW08]. Edge-Guided [MPAFF10]. Edge-Optimizing [CK94a]. edge-preserved [HL97]. edge-preserving [HB99]. Edgebreaker [Ros98, Szy02]. edges [Gul04]. edit [AGMP13]. edited [Hoc95]. Editing [WDR11, McC03]. Editor [Kro00, Dun81, Mil81, Pet81]. Editorial [Bil20, Mof97]. edn [Weg92]. Edward [Sie90]. Eero [Sie90]. Effect [CL91, MRS01, SV00, Che91, Hor95, MRS99]. Effective [HB08, MS00, OTY^{+}97, Ric08, SSW16, WGZW14, ZWZ^{+}18, AL81, CS13, Lit95, NT20, SPL20, Swa04a]. Effects [ASJ20, CFGL12, EY81, LLPP19, MAL10, vB97, LLW^{+}14, MZ91, SWA94, WSA94]. Efficiency [DG17, LKK^{+}17, LMSE18, PB11, WFG^{+}94, ZLT^{+}18, CKW91, WKM^{+}98]. Efficient [AB92, ALYK12, ALSSMPBR06, BCG^{+}14, BZM98, CCF^{+}18, CB04, CR96, CDDM05, CHSV10, CKP85, CO97, Chu97, CM04, CDR93, DFA^{+}10, DP01, DG17, EFPS16, EA95, FO10, FVW10, FOF09, FIKS18, FT04, GKPR96, GHD^{+}14, HESF11, HKA06, HL90, HLV96, HV93b, IKM82, JSC20, JTC^{+}12, JV04, Kle97, Kno80, KD18, KMS98, LV03, LMBN08, LMV03, LS99, LCL03, MRH11, MN09, NMN99, Nek07, Nix81, NZ07, OBGB11, PST97, PB98b, Pyl90, Rez07a, Rob97, SP96, Sai05a, SF11, SBE16, SSP06, SH04, SCW11, SR16, SR05, TPAK96, TCOR05, TMS02, TW96b, TM97, TM98, VST11, WV98, WZW11, XMRF^{+}13, XJ14, YGNM07, YK10, YDZL11, ZL95, ZN08, ZX11, Abe05, AMBC19, BBC95, BJ18, Bry95, CW01, CWL99, CZZR19, CK95a, CS13, DH00, DHW^{+}17, DE92, DJID20, Ghi04, HBSASA19, HPZ11]. efficient [How93, IC98, KK03, KCL18, KV19, LE94, LCY07, LLWC08, MB00a, MB00b, Mas91, May99, MLP99b, MK06, MTK95, NYOY94, Nel92, OS96, OM03, OTW^{+}99, OHRS21, PB98a, RL95, SNR06, SM11a, TVS^{+}03, VL97, WRCB02, WBM03, WF03, WOS06, YA13, YZZ^{+}14]. efficiently [NCB15]. Eigen [SK99]. Eigenvector [Fow09]. Eighth [ACM97]. EKMR [LCL03]. elastic [van02a]. electrical [Jur88]. Electrocardiogram [ABP09]. electrocardiograms [Bör80]. Electrocardiographic [Eva79]. electroencephalograph [Nad83]. electron [FW93]. Electronic [Bar00b, BBKF95, Weg07]. electronics [Jur92]. Electrophotographic [TSS^{+}98]. element [Hil91]. Elements [CT06, For96, Mar02]. Elevation [Roy87]. eleventh [ACM79]. Elias [AN08, Fen96b, Whi06]. Elimination [JLJ08, KRK09, KKJ11, Hos02, MKNG06, MB00b, Yao03]. Elizabeth [IEE04]. elliptic [FOSS17]. ELS [Wit97]. ELS-coder [Wit97]. Embedded [ALMJR^{+}12, BFT11, BH07, BW95, BMM02, BF98, ECM91a, HB09, Hsi01, KHHK21, KCBM21, LLA08, Li02b, MO11, Nos99, OINC10, PZZ^{+}22, QB09, SM14, Sha93a, XD11, YKO98, ZASB95, ZX11, AMS^{+}03, BG03a, BH10, CF99, Cha04, CP03, Cre96, Cre98, DB02, DH00, FW02, HB96b, KP97, LHW00, LMO97, LDK99, MS05, NY98, OT03, OWS98, RLR04, Sha93b, TO04, Wu99a, Wu99b, WX00, Yu96c]. Embedding [AG09, Sar10, Yok10]. Emerging [MLDH15, LR04]. EmilCont [Des08]. Emission [CFC05, ITU91]. Empirical [CRR98, PT10, MSWB93, ZE00]. empirically [Zee98]. Employed [GPM02]. enabled [Bar04]. Enabling [ZZS^{+}22, GLM^{+}21]. Encoded [CFG12, CNS12, DG17, HMR10, KS07, Kle07, PP18, RRMG07, YWC09, BK05, DS04, JR02, JA03, JRALMSS11, KS01, KT02, Kom95, LAB^{+}14, LLWD14, PM99, VSG00]. Encoder [ABP09, AAL^{+}21, BKRSR09, Dum08, HHCH11, Kro00, Nga95, RK09a, VRHL07, VRL09, LLS^{+}21, LG01a, Mal01, MWZ04, RV94, TCJ93]. encoders [FZ02, ZR02]. Encoding [AJ95, AJ96b, AOAAH17, AMAM13, BK74, BZZ08, Bar00b, BES05, Bud76, CR06, DLY^{+}98, EC09, ER78, Fis95b, Fre61, GKPR96, GJ20, Gra97, HAR10, KS07, Kno80, KD18, LZR^{+}10, LLCC11, MYS20, MG10, RS16, Raj20, SS99, Sha09, Wan88, Wic03, Yok10, ZHX^{+}19, AM95, AJ96a, ABF^{+}02, BE96, Blo96a, CM04, Cre97, Du00, FWY^{+}98, Fen93, FBS04, GB92, GVYZ97, IA00, KSS06, LY91, Mal06, MM97, MBBA91, NT05b, OHRS21, RA04c, Sch06, SB97b, Sun16, Szy02, TVLS08, TKYS16, TPAK96, TN99, VRR^{+}16, WLCZ15, WBE^{+}94, YG03, ZC99]. Encodings [GW73, GM59, Gol66, Gol72, Har04, Har07, PR98, PT10, Sto80, Wel72]. Encrypted [JSCM17, KXPZ12, KHJ^{+}09, SYDR07, FM96]. Encryption [ASLB20, FIP01, Joh89, KBD89, PMZ13, TCN^{+}17, WPA08, CMMS17, Cra96, DTZZ12, HIMM20, KV19, Kop97, KS89, KM97b, Laf00, LJ05, LPZJ15, R^{+}06, R^{+}07, Sab94, Sch03, Sch04]. Encryption/Compression [WPA08]. Encyclopedia [Mv94, Slo06]. end [GN01, KB99, TD91, Tsa98]. end-to-end [KB99, TD91]. Ended [BY90]. Ending [CD94]. endings [YS07]. Endpoint [CC19]. Endpoint-Cutting [CC19]. Energy [BA06, CNW^{+}17, HSP42, LGM21, LKK^{+}17, LMSE18, MW10, PM17, SF11, TCN^{+}17, WMNP12, WZW11, ZLT^{+}18, ZMVR14, AMBC19, BL97, FH96, JCS^{+}08, LLWC08, SM11a, TK07, YA13]. Energy-aware [BA06, JCS^{+}08]. Energy-Efficient [SF11, WZW11, SM11a]. Enforcement [ABS^{+}16]. Engine [TCJ93]. Engineer [Hou97]. Engineering [ACF^{+}21, Jur88, KT13, Lum13, S^{+}09]. Engines [ALM11, BCC10]. English [Ano09d, CW97, Coo05, Cus90, Gu05, NHK06, Sha51, TC96, TC97, TICH98, TC98, YS07]. Enhance [BMM02, MPASM^{+}09, SKY10]. Enhanced [EL03, FJ96, JR12, MPALSS^{+}12, MKH10, NR14, WZMG07, YMK10, ABA14, ASK12, KCL18]. Enhancement [LMS94, Roe76, RRG95b, ML98, Nos99, TVS^{+}03, WKM^{+}98]. Enhancements [Lak97, Pow93]. Enhancing [AJ95, AJ96a, BE95a, Boy91, FSC^{+}11, FH98, sKJ01, PR01, VMP^{+}16]. enlist [Sno76]. ENO [MMZ12]. ENO-Type [MMZ12]. Ensemble [WY10]. ensembles [HBL^{+}17]. enthusiast [hyd06]. entire [SRP04]. entity [Pow93]. entity-level [Pow93]. Entries [CS94b, CS93, LTZ91, LCT03]. entrophy [KS96b]. entropic [KM97a]. entropies [BFMX03]. Entropy [Cie98, Dun80, FW91, FHS08, Fre91, How98, JBF97, KM99a, Lan98, MN09, MPGMD19, NR08, PR71, RMG94, Sha51, Sha11, Tho91a, WZMG07, YHM20, YDDB15, Abo95, CR95a, CO97, CME05, CR95b, DK95, EM02, Faw96, FZE03, FGGV04, FA93, GW05, GFH99, GL01, GL03b, GG06, HV06a, JSS04, JV04, KK01, LBS98, LY97, LC91, MN08b, OR94, RTK^{+}16, Rez05, Sai05a, SRKS95, SDJ04a, SJPB17, SV98, TC96, VM98, Wit97, WWW97, YM97, Yu04, ZFE04, ZKD05]. Entropy-Based [MPGMD19]. entropy-biased [FA93]. Entropy-Bounded [FHS08]. entropy-compressed [MN08b]. Entropy-Constrained [Cie98, FW91, JBF97, RMG94, Abo95, CR95b, EM02, GFH99, GL01, OR94, SRKS95, YM97]. Entropy-Scaling [YDDB15]. entropy/memory [GG06]. Enumeration [DB10, Dub11]. Environment [KH08, FSY01, KS89, LKA02]. Environmental [DLY^{+}98]. Environments [CT09, HKKW13, LGK97, VK17, CS13]. EOF [Sco02]. EOS [Doz91, Sal91]. episodes [CM05]. Epitomic [WHW10]. Epsilon [PR71]. equal [PELA02]. equalization [CGR91]. equalizing [Vos06]. Equally [Roy87]. Equations [Jez16, PR98, DPS93, HSV04, HV06b, JU10]. equipped [OPS91]. Era [SCH^{+}14, Har04]. Erasure [DRR10, DZ08, EKT12, HKW10, CF97, GKV99, MRL99, SX04]. erasures [BBF02, DKG01, PK05, RG02, YJE04]. ère [Har04]. EREW [De 03]. Ergodic [CTD07, BST96, Cho05]. Erratum [De 00b]. Error [BF95b, But95, CXZD12, CCV05, CHMW10, De 94b, DC18, DFH^{+}19, HKS^{+}11, Hou97, HV92b, HHCH11, ITU94, KE11, KØJ^{+}12, LV02, Lev47, LMJ^{+}21, LZM10b, LZZ^{+}12, Lyn66b, PG72, RRMG07, Sad98c, SK10a, Sha11, Spa00, SR97b, SG11, TC06b, VD18, WLL^{+}20, Wil70, Wil93, WLS06, XWDY06, You97, ZXM^{+}10, ZLX^{+}20, ARZG03, ARR01, AB08, BEQ98, BM03, Cap85, CCA99, Chi99, CP03, CME06, DJL91, DWW02, EA95, FGC93, Hag07, Hal04, HL97, HGH05, KF95, KB03, KGR06, KCR98, LS95a, MYRD04, MRL99, MR06, Niu02, Nov98, RS00, SR97a, Sto98, TKRR02, YML97]. error- [HL97]. Error-Bounded [DC18]. Error-Control [Lyn66b]. Error-Correcting [BF95b, But95, LV02, ITU94]. error-correction [BEQ98]. error-propagation [SR97a]. error-protected [DWW02]. Error-recovery [De 94b]. Error-Resilient [SR97b, WLS06, BM03]. Error-Tolerance [HHCH11]. Errors [LWW10, PS05, Pet80, San89, ZW12, Mos98, YJE04]. escape [ASS97]. escapes [Fen12]. estimate [FS03]. estimates [Gul04, LY97]. Estimating [BFMX03, EN84, QZH10, WB91]. Estimation [ALMSS09, CA06, GV09, HMR10, JLJ08, Jog73, KKJ11, LCB03, LT08b, LLCC11, LZCF10, MN08c, MJZMA08, PM88, PT10, RKL16, SWGZ09, Suz12, TSS99, WLL^{+}20, Bör80, CS92, Chi98b, FWY^{+}98, FG01, GPO98, IDKW99, KS96b, LKRR91, LRO97, MB00b, Nos97, PM99, PGOO94, Ram91, Won91, ZR02]. Estimation-quantization [LCB03, LRO97]. Estimation-Theoretic [HMR10, LT08b]. Estimator [KMH10, KB06, KKA02, LA96]. Estimators [ALMSS09]. ETAOIN [Fan66]. ETAONI [Fan66]. étude [LG78]. Eudaemonic [Bas92]. Euler [Gol06]. European [BDLS96, dABdSFNA08]. EuroPVM [BDLS96]. Evaluating [BHT19, BCC10, CFGL12, HS06, Mat96, RS16, SW17]. Evaluation [CEA18, LLPP19, LMJ^{+}21, SW12b, SGLT98, TY08, XMRF^{+}13, AK99, AB97, CLR05, Jon91, Mai95, MSWB93, RBD13, Sha94, VBL94]. Even [Bel11]. Event [SSRM09, YCMR18]. Event-Based [SSRM09]. Events [WB91, Sav03]. ever [Bry95]. Everywhere [Bar88, Moc04]. evidence [Xue99a]. Evolution [BRD10]. Evolutionary [MSW10, Mat15]. Evolvable [TSS^{+}98, SST^{+}98]. Exact [Cha91a, GPM02, HDK^{+}12, FL13, KFN99, Ste91, Tai95]. example [Zir07]. Examples [Ano06e, Rea95, LS91, Tha94]. Exascale [SCH^{+}14]. exchange [ISO98]. Excitation [PFAK10]. Exclusive [DCP17]. ExCom [HŘŠ11]. Executable [BMM99, DXS08, MGC07, Per03b, Bun97a]. Executables [UPX03]. executing [LJ05]. Execution [LKK^{+}17, MCM^{+}17]. Exemplar [TPM20]. Exemplar-Based [TPM20]. exhaustive [BAGCGC^{+}06]. Existence [Maa94, Cm10]. Existing [Bel87, HN07]. exit [GYLC09, JFC12]. Exogenous [BGBV09]. Expansion [DD98, KH08, KR10, PM17, SB85, SF09, ZLC^{+}15, KF94]. expansions [BLO01, CMW99, Cve99, GVT95, GKV99, LKA02]. Expected [FP20, CH06]. expensive [Pou93]. Experience [Nix81]. Experimental [DUH^{+}19, Lew95, LMJ^{+}21, NL03, NT02, DvOL97]. Experiments [AVE18, BMT91, CT95, Dac01, GB92, LTZ91, RSC99, Rub76, dRV12, Ben10, THM91, Xue99a]. EXPERT [CM05]. Explained [Bla87, Bra99]. Explanation [Fel03, NMW97b, Api91]. Explicit [FWZ^{+}12, HLV94, YLZ11]. exploitation [GFC^{+}09]. Exploited [WPXL09, BKR^{+}93]. Exploiting [CS13, GFE07, KNB10, KWC^{+}16, MS96, VDPL00, Yal78, CA02]. Exploring [CCO^{+}19, PZZ^{+}22]. Exponent [ECBL12]. exponential [CD00, CDL07, ES98]. Exponentiation [Yac98]. exponents [CME06]. Expression [Nav01, BFG09]. Extend [Dou93]. Extended [HL02, KW01, LK96a, Lar96, SS78, TS17, SJ94, SW97, TO04, ZIL93, LE94]. Extending [CKCW95, KB03, MT98, RL96, ZCG99]. Extensible [Roe03, ACP05]. Extension [RKL16, SMW07, SS07, JdVL03, LAPL07, MRS05]. Extensions [Fre93, Sim94a, Sim97, Cra98a, KGK00, Use03, ISO97]. extensive [Tai95]. external [YZ07]. Extra [SMLS15]. extract [NN97]. Extracting [JGL02]. Extraction [DSM07, DMS08, WF01]. Extractors [FL98]. Extrapolation [Li86]. Extremely [RLG09, RM21, Wil91b, KC06]. eye [KK08]. eye-gaze-position [KK08]. EZW [WB00].

face [WM05, ZCG98]. face-to-face [WM05]. FaceCerts [KJ03]. faces [MP95]. Facilitate [Soc08b]. Facility [A^{+}87, Sal91]. FACOLA [ZCG98]. Facsimile [Ano07c, HR80, McC92, TH97]. Facsimile-images [TH97]. Factor [PC08, Por73, GLM10, LL02, Von04, Zip91, Zip92]. factor-graph [Von04]. Factorising [MJ75]. Factorization [CPS08, CIS08, FIKS18, KKP16a, KKP16b, Köp21, Ryt02, AHCI^{+}12, HPZ11, INI^{+}16, RS04, Ryt03]. factorizations [JMnRS19]. Fading [EEJ07, ECBL12, HR06a]. failed [TO04]. failure [BBZ14]. Fairfax [WGM88]. Fairmont [IEE04]. fall [Ano04b]. False [FB16]. False-Positive [FB16]. Families [LT09, BCSV06, MZ99]. Family [Ado06, RAE16, Ano09f, BCP20, SW97]. Fan [FM89]. Fano [AN08, Tja05]. Fast [ABA14, AAL^{+}21, ACM^{+}15, BR04, BK05, BBHS12, BA04, BYR06, CC19, CSF77, CR95a, CR97, CZZZ11, CMK^{+}16, CS02, CC06b, CA97, CA02, DI04, DI05, DJSS14, DWB04, Eff97, FOhC09, FGGV04, FK09, GWSR12, GO15, HESF11, HAR10, HV93b, HV94a, HV94b, HAH05, HNC^{+}16, JF96, JLJ08, KB93, Kie99, KOO09, KRK09, KKJ11, Kle07, KN10, KPY96, LV95, LLM89, Log04, LOMSS05, LKA99, Mal99, Mal00, Mal01, Man97, MŠH12, MGM13, MASC98, MAS00, MN09, Moa86, MZ96, NPW97b, OM03, OG12, Pas76, PFP99, RF05, RK95, RSWW01, RKB06, RC09, RA04c, SP96, SMS08, Sch98, SB97a, SM95a, Sie88, SNZBY00, SP98, SWGZ09, WHZ12, WXC^{+}10, Wil91b, WF94, XD11, Yac98, YNXC07, ZN08, ZGB04, Abe05, Ano06m, AC98, AM93, Bet94, CL97, CWL99, Che15, CDC96]. fast [Cod92, Cra98a, DPS93, DDL98, GMNK06, HV93a, LOS20, LF05, LAPL07, LHY12, LJPE21, LRO97, MR04, MLMD03, MM05, Nel92, OG16, PM98, PBC05, PD16, PSSS03, RSWW02, Sad98a, SYP04, Sch97, SKCB95a, SKCB95b, Sta07, SLZ^{+}19, TR88, VFK15, Vuc06, YZ07, YG03, CY91]. Faster [CW07, Dum06, FM89, HKS^{+}13, INI^{+}16, Jez15, LOS20, LKR17, MFCM96, NKT^{+}01, NP16, ZSP^{+}19, Bon95, BJ18, LSYKK16, LNhCW16]. fastq [GGM12]. FAT [NTF06]. Fault [Log06, Ait86, BEQ98]. fault-tolerant [BEQ98]. fax [CB97, CB96, McC92]. fax-based [CB97]. FBI [BBH93, BBOH96]. FDCT [CY91]. FDR [CC03]. feasibility [CCO^{+}19]. Feature [CZ98b, DSM07, DMS08, Eva79, JS99b, LPG10, SC01a, SP11, BE96, KS16, SB06]. Feature-Preserving [JS99b, SC01a]. Features [DS08, JGL02, HOG04]. featuring [Nel92, TCP03]. Feauveau [Ara13]. FEC [CMWM00, KT02]. Feed [SK09, Sha06]. Feed-Forward [SK09, Sha06]. Feedback [EEJ07, KR10, PR11, QB09, SF09, Lev96, LNA99, Rob91b, SCM05, ZRR00, ZDR06]. Feedforward [Sha10]. FFT [CY91, SKCB95a, SKCB95b]. FGK [MLP99a]. FIASCO [Haf01]. Fiber [HTS^{+}18, Kir05]. fiber-based [Kir05]. Fibonacci [AF87, Cap89, Kle07, KBN08, OR99, Vor61, Vor83, Zec72]. Fidelity [Dun80, WZW11, SM97, CPT^{+}20, SH94b, WCB97]. Field [HSL^{+}20, MPL02, YR96]. Fields [RN10, TRS03, ZL19, GO95, ILRS03]. fifth [ACM13]. Fifty [Ver98]. File [CC88a, CC88b, Bar12, Bee79, Bor95, B^{+}92, Cha95a, Cha95b, Del93, GW73, HM76, KL95, Lus94, Man97, Mia99, Mv94, NN97, Pae91, Pec82, Phi91, PC08, Rai87, Rod95, Rok85, Rub76, SS05b, Spi93, Swa93, Wel72, Yok10, You85, BGM99, BP00, Deu96b, Dic92, ISO98, ISO09, ISO17, KIHO95, MS96, Mur96, NTF06, OTY^{+}97, SS03, SS04, SM11a]. File-Size [Yok10]. Files [Car20, DCP17, GGM12, KW03, KS07, Kro00, LB90, Lia95, MZ96, SK10b, AAJ04, ADCU08, Ano06q, Ano09f, EN84, FB98, GKPS05, HZ95, Jac92, KSS06, LZ06, Mia99, NN97, RL04, SS04, TM05b, TM05a, Wei90]. Filesystem [Zhe03, CG91a, CG91b]. Filling [Col87, HGFL09, PLF91, Sag94, BS98, CT98b, Mia90]. Fills [Pea90, Sie12]. Filter [Ano06e, CHB86, DZZ08, Hua03, Lev47, MS12, SN96b, VBL94, WL08, DSV00b, DKG01, DKS21, HMS92, PRM97, SSM02]. Filtering [BRNMG^{+}12, MGBRMSS11, Nat93, SDFH00, DLT05, GO96, OWW^{+}04]. Filters [Ano06e, BT93, FB16, Sal18, WK92, YMK10, BK97, BJ18, BCP20, Mul87, Nat94, WMH96]. financial [Ano06c]. Finder [Yok91]. finding [CM05, CA02, EL03, Mit01]. finds [Pou93]. Fine [Ged14, HZZY02, van02b]. Fine-Grained [Ged14]. fine-granularity [van02b]. Fingerprint [BBH93, BBOH96, Fed93, RM21, HP92]. fingerprints [AGMP13]. Finite [BF11, Cha66, CJDL09, CK93a, CK94b, CK95b, CV96, CV97, EEH17, FG89, HB99, LZ76, LdV95, TR93, ZL19, AFHU97, BBZ14, CK93b, GD06, GK95, GS96, Haf96, IF06, IC98, JdVL03, KF96, LM91, MF04, OY06, RF00, TM98, ZDR06]. Finite-Length [BF11]. finite-rate [ZDR06]. Finite-State [TR93, BBZ14, GK95, GS96, IC98, TM98]. First [Fow09, HH08, IEE11b, NBIT07, NIB^{+}09, PBPT95, Ude93, FW02, NLW96, Reb12, RK15, She90, PBPT95]. first-generation [FW02]. first-order [She90]. Fisher [Hoc95, Wu99a]. Fitting [CYH94, DGG10, Dur60, VK17, DGG09]. Fix [AKS10, LV03, Sav09]. Fix-Free [AKS10, LV03, Sav09]. Fixed [AsM^{+}10, BF96, CI07, Lin14, Moa98, NM10, Rub79a, Sch70, Sho08, TW92, WZ91b, BH92, BF99c, DSSR95a, DW05, EM02, Gra05, LCT03, LBS98, RC98, Ruf94, Sav98, Sav99, Stu94b, TKR00, TW00, XWL03, YM97]. Fixed-Lag [AsM^{+}10]. fixed-length [LCT03, LBS98, RC98]. fixed-model [BH92]. Fixed-Rate [Lin14, Moa98, BF96, BF99c, DW05, EM02, Gra05, YM97]. Fixed-to-Variable [NM10, Stu94b]. FLAC [Coa06a, fd06, Coa06b]. FLAC-Free [Coa06a]. Flächenstück [Hil91]. Flag [Wan88]. flags [Fen93]. Flash [JCS^{+}17, JSC20, KJ95, JCS^{+}08]. Flash-based [JSC20]. Flash-Memory [KJ95]. flat [HR06a]. Flexible [BK91, BC10, MRS01, ZLRZ09, MRS99, SNZBY00, WC02a, ZTSM05]. Flicker [BCP04]. float [Ghi04]. Floating [BR07, BR09a, BR09b, DFH^{+}19, FM12, FDH^{+}20, IEE85, Kra10, Lin14, PWS10, SKY10, USY17, CDL^{+}19, Pat94, RKB06, Tro96, TMM96, Use03, Use05]. Floating-Point [BR07, BR09a, BR09b, DFH^{+}19, FM12, FDH^{+}20, IEE85, Lin14, SKY10, USY17, PWS10, CDL^{+}19, RKB06]. floats [LJPE21]. Florida [IEE84, IEE88b]. Flow [WFG^{+}94, DJID20, GN79, LZ05, MM99, ZG02a]. Flows [WW10]. fluctuation [BKR^{+}93]. fluorescence [Ish89]. Fly [Bal97, Eng93, Fet96, KS06b, BGM99, Hal94, RL09, Ano92c]. FM [NKP^{+}16]. FM-index [NKP^{+}16]. focal [Taw93]. focal-plane [Taw93]. Fock [MBWP03]. folding [BR06]. follow [GLL04]. followed [Ano09d]. Following [RD82]. fonction [LG78]. Font [Ano86a, Rok85, Won85]. Fontes [Har04]. Fonts [CY92, Har07, Har04]. Fookes [Kro00]. Footprint [FCL13]. Force [Kro00]. Foreground [CRA10, SMRR04]. foreground-background [SMRR04]. foreign [KIHO95]. Forensics [QLH12, QZH10, QZ14]. Forest [TGHL98]. Form [Giv58, Man77]. Formal [Hou79, Smi84]. formalism [Att95]. Format [CC88a, CC88b, Ano97, Ano06w, Cha95a, Cha95b, Cro95, Deu96a, DG96, KMMV02, Lus94, MNG03, NN97, Rod95, Roe03, Rok85, You85, Deu96b, Dia16, HN07, ISO98, ISO09, ISO17, KIHO95, Mur96, NHHS02, Ste12, fd06, vB97, Phi91, Ado01]. Formation [Koh72]. Formats [Bor95, KL95, Mia99, Mv94, Swa93]. Formulating [LLW11]. Formulation [LT08b, Gra05]. fortune [AP03]. Forty [ACM13]. Forty-fifth [ACM13]. Forums [Bar00b]. Forward [AF99, SK09, MW92, MRL99, Sha06]. Forward-adaptive [AF99]. forward-mapping [MW92]. Forwarding [BSSU18, RTK^{+}16]. Foundations [IEE84, IEE91, IEE96]. fountain [NSA06]. Four [Eve98, LP07, TSS99]. Four-Dimensional [LP07]. Fourier [CY91, Has18, BDV00, MWCG07, MM05, SMO98, THG03, Vil93]. Fourier-based [MM05]. foveation [EMS03]. Fowler [GHV05]. FPC [BR09a]. FPGA [CC19, GRG08, MCM^{+}17, MAC^{+}04, SKY10, SW12b, SDM02, USY17, WLCZ15, ZZJB13]. FPGA- [MCM^{+}17]. FPGA-Based [ZZJB13, SKY10, USY17]. FPS [HD88]. Fractal [BZ08, CGC97, DH97, DH98, Dav95b, DHN85, FC94, Fis95a, Fis95b, GK88a, GK88b, HU99a, Haf01, HU99b, Har96, Hoc95, LV95, Man82, MFCM96, MKM^{+}08, NMN99, RK93a, Sau94, SH94a, SP11, Wel99, APC92, BE95a, BE95b, BE96, Ber99, CSCW00, CC98, CD91, CDR93, Dav95a, Dav96, EW96, FU98, Ham95b, Ind91, JT96, JM96, KW95, Kom95, KFSM04, KPY96, LV96, LW11a, LLP96, Mon93, Nov98, PH91, RH97b, SS99, Sau95, Sau96, Sch06, WCH03, WLL06, Woo94, WM94, YW04, ZLN06]. fractal-based [Ber99, EW96, PH91]. fractal-like [CDR93]. Fractals [Ano95, Bar88, Fed88, Man77, PR86, Smi84, Zor88, Wil96, WB03, Hoc95]. Fractional [BGPW96b, LZCF10, KLAH99]. fragile [WHZ12]. fragment [NLW96]. Fragmentation [KWC^{+}16]. Frame [CXYQ12, CHMW10, CDW11, CZT^{+}11, EC09, ISO84, LHS05, LJZ^{+}08, LZM10b, MJZMA08, Nga95, ZWZ11, BDV00, CCV04, CCV05, GMR95, GKV99, LC03, LKA02, Str99b, Ten04, TC06a, Vil93, WM05, ZL00]. Frame-Level [CDW11]. frame-rate [WM05]. Frames [DG17, GYM93, HKW10, MOT^{+}03, PK05, VSG00, YSZ11, YGM97]. frames/s [MOT^{+}03]. Framework [ACD02, DZ08, DG17, HSZ17, HKMS08, LZR^{+}10, Ven07, Wic03, ZZS^{+}22, Col93, CR95b, GCL06, HLV97, LRO97, MR00, MS97, OR94, RT95, SF05, SSZZ14, TDG^{+}19, TEGM03, VL98, BV04a, BV04b]. frameworks [LM01]. Framing [Sim93b, Sim94c, vB97]. France [ZD81]. Franklin [IEE88a]. Fraunhofer [Bel06]. Free [AKS10, Cap78, CD10, CCKH21, Coa06a, FH96, lGAR03a, GWSR12, HESF11, HDK^{+}12, Kro00, LV03, Sav09, TY08, CKW91, Coa06b, DKCS95, Dum06, EA95, FGC93, GK95, GN01, Huy94, MOT^{+}03, MZ95, WCH03, YDLC08, van99, TFD06]. FreeBSD [Per03a]. French [Har04, NACM08, Pea90, Pig01a, Sie12, Zec72]. Frequencies [ITU89, KS08]. Frequency [BLS11, DMS08, FGM21, HO19, Liu10, OVCS^{+}09, WB91, Abe07a, BK96, CC03, CT95, Cus90, MOK00, OSMY95, YJGS00]. frequency-directed [CC03]. Frequent [WW12, ZZJB13]. Frequentative [NM91a]. FRESCO [WL13]. fringe [DP91]. Front [Kle89, Arn99, Arn00, BR05, FTL03]. fronthaul [YK20]. frontier [WBMT99]. FSM [MSW04]. ftc [Kra10]. FTL [SMLS15]. Full [Mor12, Vil93, WK93, Bar90a, BMNM^{+}93, CM91, CWW20, FMMN07, GCK^{+}96, MN08b, MZ92, NM07, VN90b, WS91, WBN91]. full-color [Bar90a]. Full-frame [Vil93]. full-motion [Bar90a]. Full-Reference [Mor12]. full-round [CWW20]. full-search [GCK^{+}96]. full-text [BMNM^{+}93, FMMN07, MN08b, MZ92, NM07, WBN91]. Fully [BCG^{+}14, Hou79, Jez15, FV20, GR99]. Fully-Compressed [BCG^{+}14]. fully-dynamic [FV20]. Function [ABB10, CKS08, DHN85, DØG08, LJF19, LG78, MMK98, TMD08, WWY^{+}07, CMMS17, CWW20, ES00, Haa10, Haa12, Har96, Kam98, KF96, LW14, RMG05, RA04b, ZLN06]. Functional [DSMJ07, MGV08, Pom16]. Functionality [Mah08a]. Functionals [SG12]. Functions [FCDH91, GH08, SW12a, SWWW15, TPM20, BR06, CDL02, GBK16, KL97, Li10, PGS10, PMH95, SWA94, WSA94, WO00]. Fundamental [Knuta, Pu06]. Fundamentals [TM02]. Funktionensysteme [Haa10, Haa12]. Fusion [Hua03, RR08, Vuc05]. future [Mac12, TH97]. Fuzzy [LPS02, PSM01, Tse05]. FWT [WK93].

G [Weg92]. G.131 [ITU06b]. G.711 [Ano72, ITU89]. G.711.0 [HKM10, KMH10, MKH10]. G.726 [ITU90]. Gadsby [Wri39]. gain [GMR95, HGS97]. gains [ARA91]. Gallager [WY10]. Gamal [LMC05]. game [LS95a, LPZJ15]. Games [Gar72]. Gaming [LLT^{+}16]. Gamma [Whi06]. Gapped [AC11]. gaps [Apo05]. Gates [Bar00b]. Gathering [LXG^{+}18, ZYT^{+}15]. Gauss [Bis94, HOG04, JG05, OG06]. Gaussian [BS06, DØG08, DSG12, DR06, GK98, GPO98, GH08, Ing08, JM03, KR10, KØZ08, Mal06, PA97, RMG05, SV09, TY07, TMD08, WWY^{+}07, ZKØ10, ZK11, Zha11]. gaze [KK08]. Gazette [Ano97]. GBFOS [VRHL07]. GBP [Ano98]. General [AFF^{+}19, BYR06, CW01, Giv58, Hou79, HO19, NR99, NPV15, Rob97, SW17, Suz11, Teu01, Woo94, ZZL^{+}15, CWZ99, CW02, CW15, LJ05, PM18, QJRA99, WLZ92]. General-Purpose [AFF^{+}19, Rob97, CW01, CWZ99, CW02, PM18]. Generalised [ST97, PA97]. Generalization [Arn02, Bun97b, Pop11, RMG05]. Generalized [AMAM13, BDHJ04, DSG12, FE99, GZ91, Gün96, HS97, KKFL14, LS95b, LST99, Maß15, Mon93, Ris76, SM95b, Ven07, FA97, GN79, KL05, LW21, Mal06, MR00, Sto96]. Generalizing [GN05]. generated [YGM97]. Generating [Fet96, HJW05, Vol97, BS98, TVW97]. Generation [Car20, Gol92, GHD^{+}14, LF05, SCV11, TPM20, VPS17, ZWZ^{+}12, ZdMNBY00, Ano06i, FW02, Mia90, NSFO^{+}99, Ric03, Sof05, SS92, Wak05, ZDY97]. generative [RK93b, WBM^{+}94]. Generator [CFY^{+}10]. generators [YSZ11]. Generic [BS94, GB09, ISO03b, KMMV02, LLZ09]. genetic [BL97, CN96, ZLN06]. GenLOT [DBTNT00]. Genome [IP06, LN17]. Genomes [XZG15, MBB08]. genomics [PWM18]. Genova [IEE05]. Geodesics [TWC^{+}09]. Geometric [CHSS08, Fre61, Sal99, TR98, BCSV06, Kri02, LLY00]. geometrical [TW00]. Geometrically [GvV75]. Geometries [HO19]. Geometry [AH02, Bar00a, BDHJ04, CVDL16, CC06a, CH11, DKB^{+}16, DGGP05, FDU80, GMC^{+}06, GK88a, GK88b, Ken61, KCL06, LPG10, Man82, Tad20, VS10, BKV05, CCMW05, CD91, LCB03, ZO04]. Geometry-Driven [VS10]. Georgia [ACM79]. German [Haa10, Haa12, Hil91]. Germany [AFL96, BDLS96, FLA^{+}03]. Gesellschaft [Bel06]. Get [Ano92a]. Getting [VN90a]. gFPC [BR10]. GFtoPK [Rok90]. GIF [Bal97, Bla87, Kom96, Lia95, Mia99, Rod95]. GIFs [Fet96, Knu98, Per21]. Gigabytes [WMB94, WMB99]. Gilchrist [Gil08]. Gipfeli [LA12]. GIS [Sam90a]. Given [Sad98c, DP91]. gives [VN90a]. GLA [KFN99]. Global [HB96b, LZX^{+}11, RLR04, ABL99, DW05, FWY^{+}98, KR97, RL96, WXCM99, Wu92]. Globally [DWW02, Wu93, ME02]. Glyphes [Har04]. Glyphs [Har04]. GML [WGZW14]. GMM [YWPG03]. GNN [HSL^{+}20]. GNN-Based [HSL^{+}20]. GNU [BGM99]. GNU/Linux [BGM99]. goes [Ham95a]. Going [HLNS17]. Golay [TV09]. Goldbach [Ano06l, Fen02, Gol06, Mah90]. Golomb [HL02, Kie04, Mal06, Sai05b, SW97, TY07]. Golomb-type [SW97]. good [GL03a]. GOST [CWW20, LJF19]. GPGPUs [AR18, OSC14]. GPS [CXF12]. GPU [CW15, FNI17, HSP^{+}13, JBG17, KPM16, MŠH12, OBGB11, PZZ^{+}22, TB11]. GPU-accelerated [HSP^{+}13]. GPU-decodable [KPM16]. GPUs [BTLK18, LKK^{+}17, PBO^{+}15]. GRAB [Les88]. Graceful [MRL99]. gradient [DW05]. Grading [SSP21]. Grafica [Hae96]. Grain [LMSE18]. Grained [Ged14]. Grammar [Ryt02, AR06, GKPS05, Leh02, LDM05, Ryt03]. Grammar-Based [Ryt02, AR06, GKPS05, Leh02, LDM05, Ryt03]. Grammars [BK90, CD10, EEH17, GHLN18, NMW97b, Lee91, NMWM94]. grammatical [Lak00a]. Grams [KS96a, KS98]. grand [KF96]. granular [HZZY02]. granularity [van02b]. Graph [AD09, DRR10, DSMJ07, FM95b, GA08, HTS^{+}18, LTO11, RJL11, Sha11, WW12, XG10, Xio11, YSM08, BHLY20, LF04, MKNG06, Pes07, SPL20, SY01, Von04, vL94]. Graph-Based [RJL11]. Graph-Structured [WW12]. Graph-theoretic [vL94]. Graphic [ISO98, Cro95]. Graphical [Fre98, YS91]. Graphics [Ano97, Bar90a, Col87, Dis88, Hae96, KL95, LLT^{+}16, MNG03, Mv94, Rod95, Sal99, Sal06, Sam90a, SDS95, SDS96, Mia99, SRG^{+}21, Roe03]. Graphs [AB10, ETT15, HMWC94, MEO11, RSWW01, UUiN12, dRV12, AM01, CR96, RSWW02]. Gray [Col85, KU95, TLR85, Gha84, Ric86]. Gray-code [Ric86]. Gray-Level [KU95, Gha84]. Gray-Scale [TLR85]. Grayscale [BBH93, BES05, BES07, Fed93, HC06, HLC07, VMFG07, CR91, Eks96, OS96, VW94]. Greedy [BGPW96b, BGPW97, BW94a, FP19, AAJ04, AGJA06, AL98, AL00, CLM14, Hor95, JYHC03, SP98, Yao03]. Greek [Uni07b]. Green [Bar00b]. Grey [CDLW05, Kno80, HP92]. Grey-Scale [Kno80, HP92]. greyscale [MT98, ZCG99]. Grid [Cha89, PLF91, CS13, EEKB15, FW02]. grids [Tro96, TMM96]. Gross [LWW10]. ground [Tur75]. Group [HL00, HLR01, Mar90, SRW^{+}04, WSDTO11, YZ09]. Grouping [Kra49, TFRH98]. Grove [M^{+}98]. Growth [FDU80]. Growth-Geometry [FDU80]. GSM [Xu06]. GST [KPM16]. Guaranteed [Bis08, BP09, TRS03, RDDD96]. Guest [Bar00b]. Guide [Ano00, Hoc96, Roe99, Sal02b, Shl94, Wil93, Mia99, Sch10]. Guided [Liu07, MPAFF10, WZW09, Eva98, TN96]. GZIP [Deu96b].

H [Sie90]. H.263 [EGCK98, GLM^{+}03]. H.264 [ITU02, Ano06f, Ano06g, Ano06h, CZT^{+}11, LK08, LT08a, LAPL07, LZCF10, MHKK03, MHKK06, MFT07, Ric03, Ric06, SMS08, SMW07, SS07, TY04, Ten04, TH10, VRHL07, VRL09, WXC^{+}10, XWDY06, Sof05]. H.264/AVC [MHKK03, MHKK06, MFT07, SMS08, SMW07, SS07, TY04, Ten04, WXC^{+}10]. H.264/MPEG [VRHL07, VRL09]. H.264/MPEG-4 [VRHL07, VRL09]. H.265 [GKSB17, HESF11]. H.265/HEVC [GKSB17]. Haar [CN07, Mul97, SDJ04b]. Haar-like [SDJ04b]. Hacks [Hae96]. Hadamard [Ish89, JLJ08]. Hadoop [DH18]. Haiku [Ano09c]. Half [FG07, BKV89]. half-inch [BKV89]. Half-Pel [FG07]. Halftone [Roe76, DDL98]. halftones [AC98, VG95]. Hammel [Ano00]. Hamming [AB08, Sal13, SK10a]. Handbook [Bor95, Hou97, RYe00, RY01, SM10a, Say03]. Handling [Sco02, Gil92, ZL95]. Handsets [KL07]. Hard [DC18, PCRLSB00, XHS05, GC16, RH97b, VN90a, VN90b]. Hard-decision [PCRLSB00]. Hard-to-Compress [DC18, GC16]. hardness [Mit01]. Hardware [AW15, MMB07, Ran88, SKY10, SM96, TSS^{+}98, VN90a, YHM20, ZL19, ZLC^{+}15, Ano82, BBC95, BB92, Cra98a, HM87, HM91, KA95, dlCNLA^{+}95, PZ12, Pat94, SST^{+}98, SJ94, fd06]. Hardware-Based [YHM20, ZLC^{+}15, VN90a]. Harmonic [LPG10, DVDD98, RC92]. harmonic-wavelet [RC92]. HARQ [DFA^{+}10]. Harten [ACD02]. Hartree [MBWP03]. Hash [Ano02, Ano12b, FCDH91, LJF19, LG78]. hash-coding [LG78]. Hashes [BC08]. Hashing [BD82, CW91, GSS01, PS12, RT87, HMWC94, Rez98, Sab94]. hassles [Hal94]. HBO [Ano18]. HCDC [ABR10]. HCI [VMP^{+}16]. HD [HSP^{+}13, TCKM10]. HDLC [Sim93b, Sim94c]. HDLC-like [Sim94c]. HDR [MCHAM08]. HDTV [BCD^{+}95, BK94, Jur91, WL99]. head [SLZ91]. headaches [Hal94]. header [ASP03, ISO99a]. Healthcare [ASLB20, CPD^{+}15]. Height [KS00b, TTW07]. Held [Weg92, AFL96]. help [VN90b]. Helper [Zha11]. Heterogeneous [HAM17, Mat15, HR06b, HZ95, MV20, SHC^{+}16, SJPB17]. Heuristic [CDLW05, Coo05, GPM02, LW04, BCD98, CLD04, De 05, Kir05]. HEVC [DG17, GKSB17, Kar12, MTA^{+}20]. hexagonal [HMS92]. Hexahedral [LI08]. Hidden [FVP08, WN10, EA98, FR95, For99, GFV98, JSS04, LGO99, Liu91, YP03, ZKD05]. hidden-Markov [YP03]. Hide [CC06b]. Hiding [CCKH21, LAP07, LHS05, Ara13, Auc98, CCD^{+}19, P^{+}13, QZ12, Way09]. Hierarchical [BJCO03, CVC95, Cha91b, CH09, EGLS21, HB96a, HV92b, IFK12, LV14, MPAFF10, NMW97b, OVCS^{+}09, SP96, Teu01, BTLK18, CCG96, CPG96, Cha91a, CR91, Che01, KP97, MMA99, NCB15, NMWM94, NMWO96, Sha93b, TN95, WS00, YGSR01]. Hierarchies [CH09, SBE16, VCP09, FM95a]. hierarchy [NMW97a, NMW98]. Hierarchyless [DGGP05]. High [AMPF09, AW04, Ano03, BCR98, BM11b, BS94, BR07, BR09a, CBR14, CYL^{+}08, CKW91, CKP85, DR06, DG17, GL03b, H^{+}98, HO19, ISO84, JLL^{+}10, KCB98, KGG^{+}14, LA12, MD95, MPAFF10, MNM04, MGV08, Moh04, MSOY96, MR06, NR08, NS01, NU05, RL04, RRMG07, RM21, RSV^{+}00, RCMS04, Sch86, TWW^{+}10, Tau00, Top96, USY17, Wel84a, Wel84b, Wic92, WZW11, XG97b, ZZS^{+}22, ZLT^{+}18, ABF^{+}91, Bis94, Che91, CZS92, CPT^{+}20, DJID20, EFF00b, FGGV04, Gin95, GL04, JB95, KAB11, KL03, LW21, LZZ98, LA96, MEMS06, dlCNLA^{+}95, RK93b, Ruf94, SST^{+}98, SDJ04a, SH94b, SLJ^{+}15, Sta94, SR91a, VL91, VED96, Vuc05, WF94, WCB97, Wu99b, WB03, ZDR06, Zir07]. high-dimensional [DJID20, VED96, Vuc05]. High-Efficiency [DG17]. High-Fidelity [WZW11, WCB97]. High-Frequency [HO19]. High-Level [ISO84]. High-Order [MPAFF10, NR08, FGGV04, Gin95, Wu99b]. High-Performance [AW04, CYL^{+}08, USY17, Wel84b, ZZS^{+}22, SDJ04a, SLJ^{+}15]. High-Quality [BM11b, BS94, H^{+}98]. High-Rate [RRMG07, DR06, MR06]. High-Resolution [MGV08, Wic92, Che91]. High-Speed [Ano03, BR09a, MNM04, KCB98, MSOY96, CZS92, JB95, Sta94, SR91a]. High-Throughput [WZW11]. high-volume [EFF00b]. Higher [CT98a, Cha00]. Highly [ALMSS09, FSL^{+}12, IKM82, MZT12, NYOY94, WL13, ANdlF07, CDC96, Gov94, HV06b]. highly-integrated [Gov94]. Highway [ZBMM97]. Hilbert [BS98, Col86, MO11]. HINT [CR91]. Histogram [BRP19, CTC^{+}09, DSSR95b, CGR91, ME02]. histograms [YS92]. historical [LGLK05]. History [Ano97, Ano06j, Ano07c, Bel06, HP95, LLYH14, Lyo09, Oku98, Hop09]. Hole [Hua01]. Holo [YX12]. Holo-Kronecker [YX12]. holographic [Luc96]. Home [SS96, Ano06d, Ano08f, Bus09, DVB06, Dol06a, Dol06b, IMA06, Lau09, Mer03, NTF06, PKW03, Pav06, Ros06, Soc08a, Tan08, Uni07a, Uni03, Wav06, Win03, XML03]. Homepage [JBI03a]. Homogeneous [MA06, Tin77]. Homomorphic [AGMP13, CCF^{+}18]. Homomorphic-Ciphertext [CCF^{+}18]. honors [LN91]. hop [GE06]. HoPE [PBL^{+}17]. Horizontal [TWW^{+}10]. Hot [IEE95a, PBL^{+}17]. Hot-Cacheline [PBL^{+}17]. Hotel [IEE88a, IEE04]. HPC [DC18, KV19]. HSI [DKS21]. HTML [Ski08]. HTTP [BLFF96, FGM^{+}99]. HTTP/1.0 [BLFF96]. HTTP/1.1 [FGM^{+}99]. Huffman [NK16, TM98, Abr94, AMM00, AOAAH17, Ams86, Bar00b, BP00, Bas91, Bis08, BP09, Bla09, BH10, BK93c, BK93d, BK93b, CFG12, Cat84, CWL99, CKCW95, CKP85, Chu97, DS04, DD98, Eng93, FR84, FT00, FK85b, FK90, FT04, Gal78, GYLC09, GCS00, HIMM20, HJW05, HV92a, HV96, HH08, IC98, Jac92, Jak78, KKN07, KCB98, KW00, KS01, KW03, KS08, KRSS21, Knu85, Lak00b, Lak02, LP95, LF05, LHY12, LY91, MK03, Maß15, MP85, MLP99b, Mof19, Nek00, NT05b, OSMY95, PM91, Phi92a, PB98b, PB98a, Pin99, QLQ11, RSV^{+}00, SK10a, Sie88, Sun16, Tja05, TW01a, VRR^{+}16, Vit87, Vit89, Wic03, Yok91]. Huffman-based [BH10, GYLC09]. Huffman-Code [TW01a]. Huffman-coded [Jac92]. Huffman-type [Abr94]. Huge [RL99]. Human [BMH08, DSM07, Ost35, EW96]. Hutter [Ano08c]. HVS [Top96]. Hy [Lia84]. Hy-phen-a-tion [Lia84]. Hybrid [BLNK14, BW98, BLS11, CAC96, GC16, HG07, KY10, LZWL11b, LM03, LZZ^{+}12, The89, VPS17, WHW10, Wav06, XMRF^{+}13, YW04, Abo95, FU98, GB94, MM99, TYD^{+}20, WAL05, WLL06, ZWS92]. hybrids [RC92].

Hypercube [KKS00, SYO94]. Hypergraphs [GB09, GB06, HMWC94]. Hyperspectral [Abo06, AAL^{+}21, BGBV09, CZZZ11, DF08, DPdS08, Fow08, GVSSBRAL08, JA07, MRS06, RH07, Abo95, GRG05, LLWD14, MAC^{+}04, MRS03, RCMS04, SZ05, SK99, TCP03, TCOR05, WBS05]. Hypertext [FGM^{+}99, BLFF96]. Hypothesis [RVG11a, RVG11b, Ye01, BT17, FWG98, Sha91].

I-Frame [Nga95]. I/O [Bec93, CNPC98, CNCC98, CHSV10, CD95, DLY^{+}98, HD88, SKY10]. I/O-Efficient [CHSV10]. IA [Ano06i]. IA-32 [Ano06i]. IBM [BWB^{+}20, IBM88, SLJ^{+}15]. ICASSP [IEE04, IEE95b]. ICICA [CLA12]. ICIP [IEE05]. ICIP’95 [IEE95c]. iCompression [Kro00]. ID [BCD98]. idealized [DHS02]. ideas [CBC00, Mac12]. Identification [Eva79, HLL^{+}20, Liu91, RM21, Goo91, SK04]. identifications [NMWO96]. identify [YBW00]. Identifying [LLYH14, Maa94, QLH12]. IDL [Kro00]. IEC [ITU02, ISO95, ISO97, ISO99a, ISO99b]. IEEE [IEE95b, IEE04, Ma96, FOP12, Ghi04, IEE85, Pat94, Yan10]. if [Bry95]. IH’98 [Auc98]. II [Auc98, HWS06, RK93b, BK74, BV04b, BA04, CLA12, DPS93, DGGP05, Fow97b, FWI99, KKT10, PHM^{+}14, VB14, Wil70, Xue99b, Xue99a]. IID [CK95a, MG10, MGL11]. III [HHG07, Ric91]. IJulia [VRR^{+}16]. IMA [IMA06]. Image [AR08, ASH87, AAL^{+}21, CC88a, CC88b, AKL18, AAR11, A^{+}87, AF88a, AF88b, ABMD90, ABMD92, ACD02, ABM10, ALM11, ALMJR^{+}12, BK74, BIP01, BF99a, BRNMG^{+}12, BSS09, BM11b, BJ95, BBH93, BES07, BA83, Car97a, CKC09, CBHC08, CYH94, CTC^{+}09, CWY^{+}10, CTC^{+}10, CZZZ11, CGC97, CCH94, CC06b, Cie98, CAL03, CD07, CDL13, CA97, CK93a, CK93b, CK94b, CK94a, DØJJ09, DX11, DS08, Das95, Dav95b, DC97, D^{+}92, DOK88, DR83, DDT^{+}12, DF08, Duv88, EC16, EMS03, ETT15, ENA02, Fed93, FC94, FO10, FLS20, FL90, FPR95, Fis95a, Fis95b, FH98, For99, FDU80, FWZ08, GVSSBRAL08, GWSR12, GB09, Gol92, GFH00, GW92, Gra91, HU99a, H^{+}98, Haf95, Haf01, HR06a, HU99b, HV08, HH99, HCMGB^{+}12, HF97, HF99, Hoc95, HGR11, HV92b, HV92c, HV93b, HV94b]. Image [HC06, HLC07, IEE95c, IEE05, ISO00, IKM82, Jay97, JZL^{+}12, JR12, JCF93, KF97, KB93, KL01, KF96, KT09, KB74, KB92, KKT10, KCR98, Kyl90, LLA08, LF03, LV95, LW11b, Lin95b, Lin95c, Liu97, Liu07, Liu08a, LZX^{+}11, LRO97, LRMPGR01, LF01, Ma96, MWCG07, MN08a, MGD08, MR91, MMZ12, MOT^{+}03, MW97, MAS00, Mia99, MCRKC09, MN08c, MN09, MMK98, MKM^{+}08, MO11, Mor12, MAL10, MSC00b, Mul97, MPGMD19, MP06, NMN99, NK20, Oh10, OLHL08, OG04, Pae91, PK10, PM92, PM05, Phi91, Pig99, PCD17, PSM01, PS06, QLH12, QWH12, RJ91, RMB07, RS16, Raj20, RH96, RP88, RLG09, RC09, RL08, RKL16, Rob97, Roe03, RH07, Ros81, SP96, Sam90a, SA93, SEC01, Sau94, SH94a, SB85, SA92, SP11, SDFH00, SM14, Sha93a, SHW^{+}01, SK09, SPS98, Ska98]. Image [SMB98, SH97, TVLS08, Tau00, TM02, TS17, The89, TWMG93, TEA08, Tod89, TW96b, Tur96, Uhl96, VMFG07, WRCB02, Wal91a, WHZ12, Wat94, WSS96, WSS00, Wel99, WV98, Whi87, WBH^{+}92, Wit08, WK93, WB82, WY91, Wu95, Wu96, WS08, WZW11, XZG15, XG10, Xio11, XWLZ08, XW10, Xue99b, Xue99a, YS10, YGNM07, YKO98, YWPG03, Zha90, ZW07, ZWS92, ZZPB09, Str99c, Abb93, AK99, ANT95a, ASM96, AMS^{+}03, ASG99, ALRT16, AECG93, AGS96, ARH03, BS04, BT17, BI98, BZ97, BHT19, BE95a, BE95b, BW98, BF99b, BH93a, Bar04, BAL02, Ber99, BDCC97, BZM98, Bon95, BAGCGC^{+}06, Bri91, BCP20, CC00, CM97, Ces91, CPG96, CZ98b, CF98, CRR98, CL91, CSCW00, Cha04, CWYW91, Che91, CF96, CO97, CO98, CLD02, CDSB98, CPP95, Cre96, Cre98, CD91, DLT05, DS06, DSSR95b, Dav95a]. image [Dav96, De 06, DH00, DvOL97, DDL98, DY91, DL01, Eff97, ES98, EGNA06, Esk95, EW96, FC98, FSC91, FW95, FGC94, FWA01, For96, FU98, GCBV95, GCL06, GZ91, GB98, GB06, GK94, GS96, Gra03, GP95, Gul02, GMNK06, Haf96, HM97, Ham94, Ham95b, Har96, HW98, HS97, HCL99, HL00, HLR01, HV91b, HV92a, HV96, Hsi01, Hua91, ISO98, IW96, JT96, JM96, JG05, JB99, JdVL03, JWK98, KLR^{+}95, KA99, KH00, KF94, KF95, KA98, sKJ01, KPY96, LGS92, LK96b, LKRR91, Lee99, Lee00, LGO99, Li02c, LW21, Lin95a, LV96, LCY07, LW11a, Lin98, LL96, LPZJ15, LLP96, Lop12, LLY00, LLWC08, MNG03, MB91, Mal00, MT92, MR97, MR93a, MF96, MM97, MMS91, MS95a, MASC98, Mis91, MS95b, MS97, MSC99, Muk04, MMA00, Ng10, NPW97a, NM91b, NL06, Nov00]. image [OG93, OS97, OM03, OY06, OR94, OTW^{+}99, OHRS21, PZ12, PW03a, PRM97, PH91, PB99, PN91, PELA02, PR01, PMH95, QZH10, QT03, RK93a, RH97a, RKSM03, RRG97, RY92, RK93b, R^{+}06, R^{+}07, RC98, RC92, Ruf94, RRG95b, SAR00, SV00, SYP04, SW91, SS99, Sau95, Sau96, SC01a, Sch03, Sch04, Sch05, Sch06, SLDJ05, SO94, Sha93b, SZ97, SCC12, SHK99, SSM02, SMRR04, SPS99, SN94, SN96a, SZ05, Sou95, SK99, Sta07, Sto96, Str99b, Sun16, SDV11, TT94, TPAK96, TCP03, Taw93, TN96, Thy10, TW96a, Tin89, Top96, Tsa91, Tsa98, Tse05, UH96, VT95, VBL94, Wak05, Wan95, WLZW04, WAL05, WBS05, Wan06, WSK^{+}11, WG00, Wil96, WBE^{+}94, Won91, WM94, WF93, WF94, WWW97, WCB97, WCM98, Wu99a, YA13, YR96, YSZ11, You98a, You98b, ZC99, ZDY97, ZZ98]. image [ZA05, ZAA00, ZLN06, Z^{+}13, ZC91, ZGS02, lGPL^{+}00, vK00]. image-based [WLZW04]. Image-Data [CK94a]. Image-Domain [YKO98]. image/video [KF94]. Imager [GGG^{+}09]. Imagery [Abo06, BFT11, Che77, MPALSS^{+}12, Poe83, Yaz81, Abo95, Ara13, MRS03, RCMS04, RS99, TCOR05]. Images [Bal97, BSG10, Bar11, BS88a, BGBV09, BBOH96, Cap78, CWS00, CY92, CCKH21, CDLW05, CRA10, CK95b, CV96, CVK97, CV97, DPdS08, Fon81, GK88a, GK88b, HCBMSS12, HI12, IFK12, JA07, JGL02, JEK98, KXPZ12, KS06a, KA12, KØJ^{+}12, KU95, LMBN08, LaJR81, Lev95, LP07, Maa94, MN08a, MS12, Mar90, Moa98, Mof91, Mor76, MFD^{+}09, MF10, NDPL00, PWS10, PC08, Ram85, Rea95, RM21, Roe77, RC96, SYO94, SSP21, TLR85, Uhl95, WGZ^{+}18, WPA08, WMB94, WMB99, YNXC07, YKO98, ZD95b, SM97, AF99, AKF06, Arn97, ATL^{+}88, Aus98, Aus00, BE96, BNK93, BP98, BM96, CR91, CLD04, CS95, CKV97, DSSR95a, DSSR95b, DvOL97, DKCS95, Eks96, EHSC92, FS04, FS02, FJ96, FN93, FNK94, Fun07, GB98, GVYZ97, HMS92, HL97, HP92, How96, HOG04, Huy94, ISO95]. images [ISO97, ISO99b, KR97, KW95, KS05, Kom95, KF02, KF03, KCS93, LK96b, LKRR91, Li02b, LMS94, LT03, Mal01, MF97b, ML98, May99, MSR04, MT98, MAC^{+}04, NACM08, Nos99, OS96, OY06, PBC05, QMCX19, RAFH92, Rob91a, Ruf95, SST^{+}98, SMC99, SM01b, SR93, SB97b, TBMM98, Tat94, TLLB06, TM91, TVS^{+}03, TH97, VPL98, VDPL00, Vil93, WW92, WCH03, Wat93, WF01, WBE^{+}94, Woo94, WF92, WC02b, YWMS08, YJE04, YS91, ZCG99, Zha99, ZGB04, van02a, van02b, PHM^{+}14]. Imaging [Car97b, FJD11, Sal91, ABF^{+}91, LLWD14]. immunity [FR06]. IMP [RRG95b]. Impact [KNTG08, YTY12, BHT19, NEGF96, PKG08]. Impairments [IRB97]. imperfect [DDGV06]. Implementable [JMW09]. Implementation [BPT06, CA06, Cap78, CYL^{+}08, CGC97, CAL03, GRG08, MEIS09, MNM04, NMN99, Nek07, PSM01, PS06, Pro86, SDM02, Tja05, WWA^{+}95, YHM20, Zhe03, BGM99, BY00, Ben96, CLR03, CLR05, FGC94, GCK^{+}96, Isr91, IC98, KCB98, Lak97, LMV03, MJB92, Mas91, MAC^{+}04, MT96, NPW97b, PGS10, Ram92, Swa04a, TCJ93, TM98]. Implementations [CDL13, DFA^{+}10, HV92d, TSS99, JT96, dlCNLA^{+}95, RSMA19, vB97]. Implementing [BCC10, DHH04, Kie99, KV19, Lam99, Mof90, OPS91, SOI00, Spi93, Mäk89]. Implications [LJ05]. Implicit [Dav96]. Implied [McM56]. importance [CDGO02]. Important [Lop12]. Impressions [Ude93]. Improve [FNV11, Cra98b, Cra98c, KR97, KM98, SMC99, YYO04]. Improved [BFG09, CKC09, CDW11, CZT^{+}11, CP03, CK13, CS94a, CN02, DDL98, DZ08, HZZY02, HC06, HD88, Jen99, Kar12, Kla98a, Kla98b, LLLZ12, LLA08, LJF19, PM99, PS12, QZ14, RMB07, RS02, Sal18, Sch86, SP11, SM14, SK09, Sto80, van02a, BC05, CR91, JLJ^{+}20, KB06, MR93c, TCP03, TN95, Ten04, XG97a, ZLN06]. Improvement [CBK96, Yok91, Bun97b]. Improvements [Bel87, Deo00, PM17, BWB^{+}20]. Improving [CWRL11, Hor91a, Hor91b, HZKL11, Jia95, KM99b, LKK^{+}17, LMSE18, LS92, LSW09, LNhCW16, MSC00a, PB11, Ski08, ZYF^{+}20, ASS97, BMT91, HBSASA19]. In-Line [ZYF^{+}20]. In-memory [UHB09, ACM^{+}15]. In-place [SS03, SS04, MGM13]. In-situ [JAL^{+}12]. inadequacy [Sai05b]. inadequate [Dia16]. inch [BKV89]. Including [CNPC98, CNCC98, IRB97]. inclusion [RBR09, RBR10]. incompressible [NMW99]. increase [BWB^{+}20, Bli91]. Increasing [Bec93]. Incremental [Abe07a, ADCU08, HB96a, LM02, MEIS09, VCP09, NMW97a, OWW^{+}04]. Independent [BN14, XD11, Bli91]. Index [AM05, BB02, CBHC08, CTC^{+}10, DØJJ09, HSV^{+}08, KS96a, LLA08, MI91, MS00, PV21, WMWW10, ZL11, AHCI^{+}12, DJSS14, FV20, HLS^{+}04, HW06, IE91, KS98, MZ95, NKP^{+}16, PM18, RF05, Sub99, TW96a, FV20, NKP^{+}16]. index-based [AHCI^{+}12]. Indexed [ZCJ^{+}20, AAS95, You98b]. Indexes [CHSV10, HWY10, Les88, RNOM09, Teu01, ACM^{+}15, FMMN07, GFH00, KL16, MN08b, Mul87, NCB15, NM07, WWW97]. Indexing [ANS12, BN14, CHSS08, FOP12, JAL^{+}12, MN08a, MZ96, Nav02, WMB94, WMB99, BEQ98, BEGV18, FLMM09, Tom06, VL98]. India [NDN13a, NDN13b, PBPT95]. Indicator [CCKH21]. Indicator-Free [CCKH21]. indices [AN10, LCT03, WOS06]. indifferentiable [GBK16]. Individual [Lan83a, ZKØ10, ZL78, dRV12, GLL04, IF06, MF97a, MF04, Wat93, WZL91]. Induced [HG07, KWC^{+}16, LMH^{+}09, NZC09, GFV97]. Induced-Sorting [NZC09]. Inducing [LS91]. induction [NMWM94, Pes07]. industrial [CZZR19, Zha99]. Inequalities [McM56]. Inequality [Ris76, GN05]. Inexact [AKL18]. Inference [BLSM19, CK94b, CK94a, LdV95, MEIS09, HB96b, NMW97a, NMW98]. Inferring [NM96b, XXZ11]. infinite [Abr94, AMS^{+}03, JMnRS19, OS03, RA04a]. Influence [MGBRMSS11]. info [NTF06, bru07]. Information [Abr63, Ano09a, BRD10, BSSU18, Big08, BK90, BCC10, Cac98, Cha77, CC06b, CLA12, CT06, DFA^{+}10, Doz91, ECM91a, ECM91b, Fan49, Fon81, FCDH91, FWZ08, GE07, ISO84, ISO93, ISO95, ISO97, ISO99b, ISO00, ISO03b, ISO09, ISO17, JFC12, JS99b, Kel02, KRW^{+}93, LP75, MJ75, MAL10, NM10, PWS10, Ric12, S^{+}88, Sar10, SL10, Sha11, SWW12, WW10, WS12, Wei15, Wol99a, Wol99c, Yan10, YR87, ZWZW12, Zur89, dCDV07, AG02, AG86, BSS05, BS06, Ben10, CPR03, GK94, Gin95, Goo03, IW96, JV04, KT13, LGLK05, Lyn69, MWZ04, NDN13a, NDN13b, P^{+}13, PR01, Rez05, Ser00, Sle74, TBKM05, TD91, Tit00, WO01, Way09, WV99b, YWPG03, ZE03, Auc98, Ano98]. Information-Preserving [Fon81]. Information-Theoretic [Cac98]. Informed [HKKW13, SHWH12]. infrared [GRG05]. ingenious [Mac12]. Ingrid [Anoxx]. Inheritance [CC01]. initial [PH90]. Inline [FDH^{+}20]. INMOS [OPS91]. innovation [Aus00, GD06, Yao03]. Innovations [KT13]. Inpainting [LSW08, XXZ11]. Inpainting-Based [XXZ11]. Input [Pom16, KT02, LW04, NEGF96, Sab94, Tha94]. input/output [NEGF96]. inputs [ABF^{+}02]. insecurity [CIRM95]. insensitive [Sad99]. Insertion [BBHS12, PS12, YWB01]. inspection [SLZ^{+}19, TWCA02]. Inspired [HL17, CLME04, LSD97]. Instantaneous [YK10, LF95]. Institute [IMA06]. Instruction [KHHK21, MMB07, YTY12, Fra06, KMA^{+}20, MS05]. instrument [Sal91]. Instruments [Hua01]. Integer [BYR06, DUH^{+}19, Fen02, GvV75, HESF11, LLCC11, LT09, Mal07, Mof00, Slo06, YZ09, AK99, Ano06g, BZM98, FZ05, HS98, LKR17, MSR04, SDJ04b, SM98]. Integer-Pel [LLCC11]. Integer-Reversible [Mal07]. integer-to-integer [AK99]. Integers [AHK97, Col86, Eli75, Fen96b, HB08, Wan88, YO91, Yam00, LBK16]. Integral [BRP19, Bee79, DV01, HD88, LJF19, FW93, HSV04, HV06b, MBWP03]. Integrated [LZCF10, SLJ^{+}15, Yu04, Gov94]. Integrating [HKMS08].

Integration [Kro00]. integrity [Ano06q, Bli91]. Intel [Ano06i, KGG^{+}14, Nel02]. Intelligent [HL17, LHD^{+}20]. Intelligible [Fou85]. Intensity [KG98]. intensive [COMF99]. Inter [BZL^{+}12, CWLS15, BF99b, LC03, SZ05, TCOR05, VSG00]. Inter- [BZL^{+}12, VSG00]. Inter-/Intra-Descriptions [BZL^{+}12]. inter-band [SZ05, TCOR05]. inter-color [BF99b]. inter-frame [GMR95]. Inter-Sequence [CWLS15]. Interactive [AJN08, BIP01, CH09, GO84, KH00, Wei15, Lut88a, Lut88b, PW03a, Pow93]. Interchange [Ano09a, ECM91a, ECM91b, Mur96, Sav03]. interdisciplinary [Sch10]. Interest [HV08, SSP21, FDKB11, MR00, SMC99]. Interest-Based [SSP21]. Interface [WGM88, Xu06, CT99, Zhu02]. Interference [MWCG07]. interframe [RLR04, VDPL00, WF92, WCM98]. interlayer [HZZY02]. Interleaving [How98, ZJW^{+}12]. intermittent [CL16]. Internal [KWC^{+}16]. International [ACM03, AFL96, Ano03, Auc98, BJZ94, BBKF95, CLA12, FLA^{+}03, HR80, IEE88a, IEE95b, IEE95c, IEE04, IEE05, IEE11b, KRW^{+}93, NDN13a, NDN13b, PBPT95, SM01a, WS12, Yan10, YR87, ZD81, vL94, ISO93, ISO00, ISO03a, ISO03b]. Internet [BY00, BMG18, Del93, ECM98, FWI99, Lum13, MMA00, XSW07]. Internet-adaptive [MMA00]. Interoperability [Soc08b]. Interpolation [CZT^{+}11, CDL02, JR12, LZCF10, MJZMA08, XG10, YMK10, ZW07, CR91, DF93, KSY95, LN02]. Interpolative [MS00, KWLG92, NT20, ZC91]. Interpretation [LT08b, BHS06, Fra06]. interpreters [LF05]. intersection [LBK16]. Intersections [ACK08, KL16]. interval [DTZZ12]. intervals [Gra97]. Intra [CWLS15, KOO09, LT08a, LSW08, SMS08, TWW^{+}10, WPXL09, GMR95, LC03, Ten04, VSG00]. intra-correlation [VSG00]. Intra-Descriptions [BZL^{+}12]. intra-frame [Ten04]. Intra-Sequence [CWLS15]. intra/inter [GMR95, LC03]. intra/inter-frame [GMR95]. Intractability [Rub88]. Intraframe [CF97, BCP04]. intranet [CBR14]. Introduction [Ano06r, Big08, BG03b, CLR01, Hau12b, Hoc96, LV08, RB98, Sal08, Say96, Say00, Say05, Say06, Say12, Wol99c]. introductory [Ben10]. intrusive [QLH12]. Invariant [DMS08, Li86]. Inventor [Bar00b]. Inverse [LSW09, NZ07, MW92, Sew01]. Inversion [KKP12, Arn99, Arn00, Arn02]. inversions [Arn96]. Inverted [AM05, CTC^{+}10, Les88, MZ96, PV21, ADCU08, MS96, PM18, Sub99, WLH^{+}17, Wei90]. Invertible [YZ09]. Investigating [SLDJ05]. investigation [CD95, Lit95, OR94, Sno76]. Invisible [FH98]. Invoking [ABS^{+}16]. IoT [AMBC19, BLSM19, TG17]. IP [ASP03, CRN08, RTK^{+}16, SHW^{+}01]. IPzip [CRN08]. Iris [SK04]. Irregular [BT93, CNPC98, CNCC98, CDL02]. ISBN [Hoc95, Hoc96, Ma96]. Island [IEE84]. Isn’t [Fan66]. ISO [ITU02, BS94]. ISO-MPEG-1 [BS94]. ISO/IEC [ITU02, ISO95, ISO97, ISO99a, ISO99b]. Isogeny [ZSP^{+}19]. Isogeny-Based [ZSP^{+}19]. Isolated [Gad91, YG02]. Isoline [Che05]. isovalue [MSE17]. Issue [Bil20, CS13, MVJ17, Mof97, Ska98]. Issued [LG78]. issues [BY00, Dic92, LG78, Mon93, You97]. Italy [IEE05]. Itemset [ZZJB13]. Iterated [BJ95, DHN85, Gul02, Har96, KF96, ZLN06]. iteration [WCH03]. iteration-free [WCH03]. Iterative [BH00, BH01, DF93, FDH^{+}20, Fra97, MMK98, Sri99, YWC09, ZX11, MF97a, She90, XHS05]. iteratively [ZC91]. ITU [HKM10, ITU02, ITU94, MKH10]. ITU-T [ITU06b, HKM10, ITU02, MKH10]. IV [HHSS08]. IWPP [PBPT95]. IWPP-94 [PBPT95]. IX [R^{+}06].

J [Ano00, RBR10, RHC87]. Jacob [Per21]. JafSoft [Kro00]. Jamming [And07]. January [ACM97, IEE09, IEE10, USE93]. Japan [Ano03, Oku98]. Japanese [MSYY98, SMOY97, YMYS99]. Java [Bar00a, ECM98, NN97, Wic03]. JBEAM [HCD04]. JBIG [FAEY95]. JBIG2 [Ano06r, TK99, YSCK00]. JBIG2.com [Ano06r]. Jeff [Gil08]. Jerzy [Ano98]. JHelioviewer [MFD^{+}09]. Joho [Ano03]. join [GYLC09]. join-exit-tree [GYLC09]. Joins [LAB^{+}14, OW03a]. Joint [AJRK98, BF99c, BFCP09, CPG96, CGS93, CF99, CT98b, DS98, DHS01a, EC16, ECBL12, FW02, FJD11, GF01, GG04, GCS00, Hil75, HGFL09, HH08, Ing08, LZX^{+}09, LJ07, LGO99, LA05, LCG11, LS03, MM99, PMZ13, PSH00, SF09, TCN^{+}17, WJBM05, WWW06, ZZPB09, AH02, BHS06, CF98, CS05, DC03, GG06, HR06a, HSX02, HOR05, JM03, LPR97, LNA99, PM99, PRM97, QJRA99, SW06, Tei98, WAL05, WBM03, XSX05, XHS05]. Journal [IBM88]. JParEnt [SJPB17]. JPEG [ISO99b, AOT13, ASJ20, BSG10, Bar11, BKRSR09, Ber99, BH12, CL91, CD95, CR95b, DSSR95a, DSSR95b, EFPS16, EW96, GWSR12, HSP^{+}13, HHCH11, HLNS17, ISO00, IFK12, JGL02, KB93, KW03, KS05, KS06a, Lak97, Lak98, Lak00b, LGS92, Lev95, Lin95b, Lin95c, Liu97, MD95, MGBB00, Mia99, MAL10, MFD^{+}09, Nel02, NU05, Oh10, OPS91, PM92, PC08, QZ12, QZ14, QMCX19, QLH12, RV94, RL96, RK09a, Ric10, Ric11, Ric12, RAE16, SEC01, SHW^{+}01, SHC^{+}16, SJPB17, SDM02, TY04, TM02, TCKM10, TEA08, TVS^{+}03, VW94, Wal91a, Wal91b, WHZ12, WLCZ15, WSS00, WF01, WC02b, Yu04, Zha90]. JPEG-2000 [MGBB00, JGL02]. JPEG-coded [TVS^{+}03]. JPEG-Like [Lin95c, Lin95b]. JPEG-LS [GWSR12, WSS00]. JPEG/MPEG [RV94]. JPEG2000 [ALSSMPBR06, ALMSS09, AL10, BFT11, BRALSSMP08, BWM03, BMA04, CB04, CE03, GFC^{+}09, JRALMSS11, KAB11, KUCV04, LBMN08, MRH11, MŠH12, MOT^{+}03, MPALSS^{+}12, Mor12, MGBRMSS11, Oh10, OLHL08, QZH10, RMB07, Ric08, TBKM05, TVLS08, TEA08, Use03, Use05, WBM03, WJBM05]. JPIP [MPALSSBR10, MPALSS^{+}12, Ric12]. Juan [IEE91]. July [Fis95b, HHW05, KRW^{+}93, NDN13a, NDN13b]. Junction [Kro00]. June [ACM95, ACM03, ACM13, IEE88a, KRW^{+}93, LL08, vL94]. Juni [Gol06]. Just [Zor88, ZL00]. just-in-time [ZL00].

K-Channel [ZKK09]. Kalman [Hua03]. Karhunen [EFZ03, GDV03, SB85]. Karhunen-Loeve [SB85]. Karlheinz [Ano06j]. Kaufmann [Hoc96]. Kb [LK91]. Kb/s [LK91]. kbit [ITU90]. kbit/s [ITU90]. kbits [Lio91]. kbits/s [Lio91]. KDE [FK09]. KDS [Pop05]. KDS-transformation [Pop05]. Kernel [VCW09, HSV04, HV06b]. Key [BD82, EC09, IFK12, JSC20, LHD^{+}20, Sle74, ZSP^{+}19, ZdMNBY00, GYLC09, KV19, Tho91a]. Key-Point [LHD^{+}20]. Key-value [JSC20]. Keypoints [TPM20]. Keyword [LT03, Mas91]. kgbarchiver.net [Ano08f]. Khalid [Hoc96]. King [Bar00a]. Kit [Kro00, Pav06]. KLT [BSS09, EHSC92]. KLT-wavelet [EHSC92]. Kluwer [Ano98]. knapsack [Moh02]. Knowledge [BMH08, GFE07]. Known [BP09, Ano06k]. Knuth [DS04]. Kohonen [GPO98]. Kolmogorov [LV08]. kompresji [Swa04b]. Kraft [GN05, Ris76]. Kronecker [YX12]. KungFQ [GGM12].

L [RHC87, Bel86, DH97, DH98]. L1 [VMP^{+}16]. labeled [FLMM09]. Laboratories [Dol06a]. Lag [AsM^{+}10]. Lagrangian [Gra05, LZX^{+}09, LT08b, LWZI12]. lakes [Kes91]. lambda [TKYS16]. lambda-terms [TKYS16]. LANDSAT [Cap78, EHSC92]. Landscape [MP06]. Lane [KCBM21]. langage [LG78]. Language [dABdSFNA08, Cho56, CWRL11, FNP08, FL98, LG78, RL99, SS00, Fou85]. Languages [CKP85, EEH17, Smi84, BFMX03, GN05]. LANs [PBEA95]. Laplacian [BA83, ETT15, LV14, VD18]. Laplacian-based [LV14]. Lapped [MSW96, Mal07, XWLZ08]. Large [AVE18, BKK20, BKS^{+}18, BJZ94, CMK^{+}16, DJCM09, FLA^{+}03, G^{+}75, LB81, MSW10, MFD^{+}09, SLZ08, SY18, VK17, WYM10, AAJ04, AF99, BHLY20, Bör18, CW02, DPS99, EBH^{+}16, EBH^{+}18, EBH^{+}19, FWY^{+}98, Gu05, ILRS03, KR97, Mat15, OY06, PB98a, RA04a, RA04c, SPL20, Szy02, YCMR18, YJGS00, YYO04, ZL95, ZD81, ZGB04, vB97]. large-alphabet-oriented [Gu05]. large-format [vB97]. Large-Scale [AVE18, BKS^{+}18, WYM10, BKK20, EBH^{+}16, EBH^{+}18, EBH^{+}19]. large-scaled [SPL20]. laser [MGM13, PHG86]. laser-based [PHG86]. LASIS [MWLW09]. Latency [HAM17, PM17, Sho08, SRG^{+}21, HSP^{+}13]. Latency-aware [SRG^{+}21]. Latency-Tolerant [HAM17]. Lattice [BF95a, Cat84, Ser00, ZW07, DSV00a, FDZ00, HW06, KGK00, ØJH05, SVS99, ZF94]. Lattice-based [BF95a, FDZ00]. lattices [GCBV95, TH05]. Lau [Lau09]. launch [PHG86]. Law [Bar00b, Ano07i, Bak07, AOT13, QZH10]. Lawrence [TW92]. Laws [ACD02]. Layer [BFCP09, CCKH21, HSL^{+}20, LCG11, NYJ09, Ost35, FJ00, KF03, LAPL07, WG00]. Layered [CZZR19, EJ07, OLHL08, PK10, CX04, CMWM00, DL01]. Layout [YS10, IW96]. layouts [GFC^{+}09]. Lazy [KKP16a, KKP16b, ZHX^{+}19]. LBG [Wak05]. LC [MKW^{+}06]. LC/MS [MKW^{+}06]. LCP [Sim94a]. LCT [RRG95b]. LDPC [BRD10, Cen09, KTMS10, LLN^{+}04, LKR12, SF05, Sar10, TGFZ03]. leader [GLL04]. leading [Bry95]. Leakage [Kel02, TK07]. Leaking [TSFS21]. Lean [Sau96]. learnability [War97]. Learned [HGR11, FV20]. Learning [BST96, BLSM19, EBH^{+}17, GV09, HLL^{+}20, II02, KdF12, KCBM21, KCCW17, SCV11, VMFG07, AMBC19, BKV05, BYJ20, EBH^{+}16, EBH^{+}18, EBH^{+}19, Fre98, GB92, HBL^{+}17, KH00, Lin95a, SB06, Tha94, Wan95, WG94]. Least [CXZD12, Llo82, PW86, ZXM^{+}10, LMV03, Pou93]. Leech [GCBV95].

LeGall [Ara13]. Lemon [Bar00b]. Lempel [ACP04, AJ96a, ANdlF04, ANdlF07, AHCI^{+}12, Apo05, AN10, ANS12, ACF^{+}21, AF18, BK93a, BFG09, BEGV18, BES05, CPS08, CI07, CIS08, De 11, DJSS14, FT95, FT98, Fen93, Fen95, FOF09, FIKS18, GKPR96, GVYZ97, HPZ11, Hor95, II02, JS95, JST01, JMnRS19, Jen99, KS96a, KS98, KKP16a, KKP16b, KY99, KW01, KS00b, Köp21, KM97a, KM99a, KVP20, Lan83a, Log04, LS95b, LS96, LST99, MW85, Mun91, NR99, Nav01, Nav02, NT05a, PFP99, PR98, RS00, RSV^{+}00, RL99, Ryt02, Ryt03, Sav97, Tho91b, WO00, Wil91b, WZ91b]. Lempel-Welch [ACP04]. Length [AJ95, AJ96b, BH01, BF11, BF95b, But95, Cha66, Cm10, CTW95, CT97, DLY^{+}98, FLS20, Fen96b, Fen02, Fre93, Gha87, GW73, GM59, GJ20, GK88a, GK88b, Gol66, Gol72, HDK^{+}12, KMSW18, LV02, Lan83b, LM01, NM10, PSR04, PS05, PM91, PP18, RJL11, Rez07a, SK10a, SDJ04a, Sho08, TWM95, TW92, TW01a, TW01b, Wel72, WV98, Wil70, AJ96a, BH00, Cap59, CC03, CZAR20, CH06, CK95a, CMP99, CM04, DS98, EA98, EN84, GB92, GFV97, Gha84, HCL99, JKL13, KN05, KG03, KT02, LKH^{+}00, Li02a, LCT03, LM02, LBS98, OHRS21, PM99, RC98, Sal07b, Sav98, Sav99, Sha04, SG04, Stu94a, Stu94b, TBKM05, TW00, WW92, WV99b, YKLK99, YM97, Nas13]. Length- [Cm10]. Length-limited [SDJ04a]. Length-restricted [LM01, KN05, LM02]. length/Golomb [Mal06]. lengths [Abr93, AMM00, Maß15]. Leonardo [Gri73]. Less [CPS08, YR02, FNI17]. Letter [CW91, Dun81, Mil81, Pet81, Dum06, Wri39]. Letter-oriented [CW91]. Lettre [Gol06]. Level [ABR10, BZL^{+}12, BZ09, Che77, CDW11, CKP85, CV96, CV97, ISO84, KMM^{+}12, KU95, Mof91, RMB07, SjWL08, BSF16, BCP20, CG91a, CG91b, CC01, Gha84, Huy94, IW96, Mal01, MF96, MF97b, MTD08, NPW97a, NPW97b, Pow93, Ski05, WW92]. level-dependent [BCP20]. Levenshtein [Ano06k]. Levenstein [Ano06k]. Lexical [BW99]. lexicographic [HLV97, Hoa99, LKG12]. Lexicographically [Yok99a]. Lexicon [CZ98a]. lexicons [RL02]. liar [LS95a]. libraries [PM18]. Library [lGAR03a, HWY10, Nel02, VL97, BBQ^{+}95, Fow00, Ish89]. Library-based [VL97]. Lie [WSDTO11, YZ09]. life [MBB08]. Lifetime [PM17]. Lifting [DZZ08, LMBN08, MEO11, ZW07, CDSB98]. Lifting-Based [LMBN08]. Light [HSL^{+}20]. Lightweight [DUH^{+}19, JCS^{+}17, KCBM21, BFNP10, LNhCW16, SM11a, TG17]. Like [CFG12, KVP20, KN10, ALM^{+}18, CDR93, CKV97, CVK97, Lin95b, Rob91a, SDJ04b, Sim94c, Lin95c]. Likelihood [DFA^{+}10, MN08c, KT07a, TKR03]. Limited [DRT81, HAR10, VK17, CR06, HM96, Lel91, SDJ04a, ZF92]. Limiting [EY81]. Limits [Cha97, Cha06, Cha07, ZE00]. Line [AJ96b, BGPW96a, BGPW97, CO98, JB74, LM00, LMJ^{+}21, MM93, Moh04, NBIT07, RLG09, Slo06, YGNM07, ZYF^{+}20, AJ96a, AL98, AL00, ACP05, Bun97a, B^{+}92, Cha00, CS93, DP76, DS96, Dou93, FWY^{+}98, Hil91, LL02, NLW96, OTW^{+}99, Pap99, SST^{+}98, SSM02, Vov06, WO00]. Line-based [YGNM07, OTW^{+}99]. Linear [ABL99, Beh08, BKYK94, Dav12, EBH^{+}17, HKM10, HC98, JAC19, Kun82, MSW04, Moh04, MSC00b, NBIT07, NIB^{+}09, NMW97a, NZC09, OM07, Poe83, RPE81, SH03, WK92, ZN08, BNK93, BdB06, Bre87, EBH^{+}16, EBH^{+}18, EBH^{+}19, FS98, FGMS05, FY95, HF97, KF95, LV92, LLY00, Moh02, MSC99, PGS10, PST05, RJL11, SN96a, VPL98, YQ05]. Linear-Time [NBIT07, NIB^{+}09, HC98, NMW97a]. lines [HCD04]. lingual [Chi98a, CZ99]. Linie [Hil91]. Link [ISO84, Mor76, RSWW02, AJRK98, RSWW01]. Linked [JSS15]. Linking [CHSS08]. Links [BSG10, Sim92, MB91, Per89, Per90]. Linux [Ano97, BGM99, Ven17]. List [AM96, AM98, Ano09d, DLOM08, MOW08, Cha00, Coo05, ITU06a, IY00, Ma96, WLH^{+}17]. Listless [SM14]. lists [DLT16, Say98]. literate [Sab94]. Literature [Mot98, Lyn69]. Lithography [YS10]. Little [RLG09, Smi09]. Live [SQ16]. LiveRender [LLT^{+}16]. LLCs [PBL^{+}17]. Lloyd [JG05]. Lloyd-clustered [JG05]. LMS [Ano06e]. Load [CNW^{+}17, Ind91]. Local [AM10, AL10, LZX^{+}11, VS10, Wic94, WFG^{+}94, BK96, BBG10, CGSS15, HB96b, JB95, SWW12, lGPL^{+}00]. Locality [BSSU18, GFC^{+}09]. Localization [YZLH16, MM00]. Localized [WMNP12]. Locally [BSTW86, CW97, GNF15, MA06, CC93, FWG98, RHC87, ST97, TC06b]. Location [CC01, GHD^{+}14, Yaz81, ZCG98]. LOCO [WSS96, WSS00]. LOCO-I [WSS96, WSS00]. Lodge [SR91b, SC92]. Loeve [EFZ03, GDV03, SB85]. Log [B^{+}92, DFA^{+}10, LA96, RL04, XCK18]. Log-Likelihood [DFA^{+}10]. Log-Structured [B^{+}92]. LogCA [AW15]. Logging [Hua01]. Logic [ABS^{+}16, SHW^{+}01, CMP99]. logical [BJ18]. Long [MWC01, NR14, BM99, Dia16, IDKW99]. Long-Term [MWC01, Dia16, IDKW99]. Longest [BK93a, KS00b, NBIT07, NIB^{+}09, BBHK14, NLW96, RSV^{+}00, RK15]. Longest-First [NBIT07, NIB^{+}09]. Longest-match [BK93a]. Look [BRP19, Nek00]. Look-Up [BRP19, Nek00]. lookahead [TC06a]. lookup [CCG96, CPG96, MMA99]. lookups [CMP99]. looming [PWM18]. Loop [ZW12, Cre97]. Lorentz [DGG09, DGG10]. Lorentz-Type [DGG10, DGG09]. Loss [BZZ08, CT09, CWRL11, LSC10, LHS05, PWS10, FNI17, GLM^{+}03, MR00, MLP99a, Sam89, WO00]. loss-less [FNI17]. Lossless [AR08, AKF06, AB10, AC07, Ano08d, Api91, Arn96, BA08, Ben97, BSS09, BWS95, BW94b, CCRCK09, Car97a, CWS00, CC01, CST11, Chu92, CAL03, CDLW05, CD07, CDL10, Coa06a, CA97, CRA10, DX11, De 06, DQ97, DJCM09, DB10, ECM95, ECM01, Eks96, EFF00b, EKT12, FS04, FH98, FWA01, FM12, Fre93, GWT10, GWSR12, GB09, GGG^{+}09, GE07, HS01, HKM10, Hau12a, HF97, HF99, HŘŠ11, HV92b, HV92c, HV93b, HV94b, How96, HNC^{+}16, JA07, JAL^{+}12, JB99, JLJ08, KMH10, KYNC96, KKJ11, KT09, Kno80, KCBM21, KA12, KRT05, LL16, L^{+}05, LW11b, LHD^{+}20, LI08, LE01, LE02, Log04, MWLW09, Mal07, MS12, MYS20, MF98a, MW97, MNM04, Mof97, MKH10, MSC00b, Ng96b, QZ12, RN10, Rez04, Rob94b, Roe03, RC96, SA92, Say03, SRKS95, SAS10, SST^{+}12]. Lossless [Sto96, SH97, S^{+}99, TT94, TWMG93, TK99, Tur95, TW96b, VCP09, VPL98, VPGB10, WBS05, Wav06, WSS96, WSS00, WF92, Wu95, Wu96, WCM98, XDZ11, YHM20, YNXC07, YS10, YS91, YGM97, ZE00, vB98, Abe99, AB08, Ano06m, Ano08e, Arn97, AB97, BC05, BF99b, BA06, BZM98, CGSS15, CCD^{+}19, Cha93, CR96, CW15, Chu02, CLD02, CLD04, Coa06b, Cre98, DvOL97, DDL98, DHS^{+}01b, DKS21, Eff99, Eff00a, ES98, ES00, FY95, GD02, GCL06, Ghi03, Ghi04, Ghi06, GSW02, GB98, GB06, GK94, GFH00, HM97, HIMM20, HS98, HW98, Hos02, HV91b, HV92a, How93, HV96, HAH05, KNY96, Kli00, Kül12, LGS92, Lee99, LL02, LR04, LV96, Lin98, MF97b, MS03, MTTH13, MMS91, MS95a, MT98, MSC99, Muk04, NR95, NYC05, Pes07, QT03, RKB06, RS98]. lossless [Rob94a, SV00, Sch97, SC01b, Sha04, SH94b, SN94, SN96a, SZ05, Sta07, Swa04b, SDV11, Tad98, TFRH98, Tat94, TN95, TWCA02, Tro96, TMM96, Use05, VDPL00, VL91, Vuc06, WS05, WBE^{+}94, WBMT99, ZA05, ZCG99, ZE01, ZE03, ZGB04, van02b]. Lossy [ASM96, AsM^{+}10, ALMJR^{+}12, Bar11, BW10, BGBV09, BSS09, BP98, CCRCK09, DC18, GL09, IK11, JMW09, JLJ08, VG95, KXPZ12, LL16, LS94, MF97b, NDPL00, PCD17, RRMG07, Sar10, SSP06, Sha10, TDL^{+}19, TPM20, WLL^{+}20, ZD95a, ZD95b, ZMVR14, ZLX^{+}20, dMGdC10, dRV12, SM97, BHT19, CCO^{+}19, CDL^{+}19, CZ98b, CZS92, Che15, CO97, FS04, How96, LLZ93, MF97a, Mod03, NL06, RRG95b, SV00, SLDJ05, TT94, TDG^{+}19, TW96a, Wil00, Wil02, WBM^{+}94, WBE^{+}94, YWMS08, ZE00]. Lossy-to-Lossless [BSS09]. Lossy/lossless [MF97b, WBE^{+}94]. Louisiana [ACM97]. Low [Ano03, AsM^{+}10, ALMJR^{+}12, BLS11, CDL10, DR83, DHS^{+}01b, DF08, EC09, FLS20, GKN09, GS96, KMH10, KL07, KM99a, KCS93, Les88, Li02c, LN02, LRH^{+}21, LLWC08, MHKK03, MHKK06, MW06, MŠH12, MM97, MF98b, Moa98, PC08, RGF06, SFT03, SSW16, Sar12, Sho08, SR97a, TV09, WSS96, Wu99b, WY10, YA13, ZWZ11, ZRZ^{+}19, ZE03, ZX11, ZZPB09, vK00, van02b, BLO01, BK94, CPT^{+}20, CL06, CSP05, CAC96, CAC97, CF97, Cve98, DRG94, DB02, DYW^{+}96, Eff97, FC98, GCK^{+}96, GMNK06, HLV96, HSP^{+}13, JOC97, KC06, LHW00, LW21, LW00, MB00a, MB91, NZ95, OWS98, Sai05a, SC01a, TW00, TFV05, TEGM03, TFA93, WS05, WM05, Xue99b, Xue99a]. low-bit [MB00a]. Low-Complexity [CDL10, DF08, KMH10, MHKK03, MHKK06, Moa98, Sar12, ZZPB09, LN02, LLWC08, RGF06, vK00, van02b, OWS98, Sai05a, TW00]. Low-Cost [FLS20, SSW16, SR97a]. Low-Delay [AsM^{+}10]. Low-Density [WY10]. low-latency [HSP^{+}13]. Low-Memory [GKN09]. low-pass [Xue99b, Xue99a]. Low-Power [Ano03, SFT03, LW00, TFA93]. Low-rate [LRH^{+}21, DRG94]. Lower [MSFZ08, ZW08, CH06, GL03a, LY97, MF94, OM03]. Lpc [Mar80, MKH10]. LRU [De 05]. LS [GWSR12, WSS00]. LSP [DK95]. Lucas [Zec72]. LUISA [FGM21]. Lumigraph [ZL00]. luminance [Won91]. LUTs [TPM20]. LVQ [DPdS08, LGLK05]. LVQ-SPECK [DPdS08]. Lyndon [INI^{+}16]. LZ [Cra98b, Cra98c, De 03, DCP17, JSS15, KL05, LW00, RH91, Yok10].

LZ-Compressed [DCP17]. LZ-Compression [JSS15]. LZ1 [Sto96]. LZ1-type [Sto96]. LZ2 [De 00a]. LZ’77 [LS03, AHCI^{+}12, Chi98a, DB08, Fra06, KN10, LNhCW16, Mos98, PP18]. LZ77-compressed [Fra06]. LZ77-Like [KN10]. LZ78 [ALM^{+}18, Chi98a, INI^{+}16]. LZ78-like [ALM^{+}18]. LZAC [Chu02]. LZB [Ban09]. LZFG [Jia95]. LZgrep [NT05a]. LZHAM [AKSV15]. LZMA [AKSV15, Pav06]. LZP [Blo96b]. LZRW1 [Rez98]. LZSS [lCcYxFpW09, CFY^{+}10, KS07, LD10, OSC14, Pyl90]. LZW [AJ95, AM95, AC11, Bla87, GR99, Hor91a, Hor91b, KTS^{+}98, KTSA99, Nel89, PM91, PMK91, Phi92b, Reg90, TM04, TM05b, TM05a, Tis87, WLS06, ZTSM05]. LZW-compressed [GR99]. LZX [Ano09e].

M [FR06, LLN^{+}04, SL08]. M-ary [LLN^{+}04]. M-Channel [SL08]. MAC [SV09, TOL11]. Mache [Sam89]. Machine [EBH^{+}17, GHV05, HKKW13, KdF12, KCBM21, NBK16, SYO94, AMBC19, BHS06, BDLS96, BYJ20, EBH^{+}16, EBH^{+}18, EBH^{+}19, Fre98, GB92, IC98, MT92, SB06, Tha94, TM98]. machines [KRT05, Lel91, OPS91, RF00, SLZ91]. Macintosh [Ano94, Ano06q, Ano09f]. Macro [BZL^{+}12, SS78]. Macroblock [SjWL08, XXZ11]. Macroblock-Level [SjWL08]. Macropixel [MS12]. MacWrite [You85]. Made [Pae91]. Magia [dP58]. Magnetic [CRA10]. Mail [NY98]. Main [BLNK14, MV16]. makes [ARA91]. Making [HKKW13, MD95, MS06, BBH97, Per21, Sad98a, WO00]. Management [ACM03, HWY10, KJ95, Kro00, PBGA19, SI99, SM01a, Bli91, BB92, CPD^{+}15, GYLC09, OTY^{+}97, Ude93, ZR02]. Managing [RB01, WMB94, WMB99]. Manifolds [DGGP05]. manipulating [SR93]. Manipulation [A^{+}87, BBKF95, CK95b, SILN11]. many [CBR14, KSCE16]. many-core [KSCE16]. many-to-many [CBR14]. MAP [BF95b, AA71, CCRCK09, SBE16, AKF06, BGK91, CPP95, Hil91, KF02, KF03, MBB08, RKSM03, SHD02, ENA02, ENA03]. Mapped [HKM10, KA12]. mapper [EHSC92]. Mapping [AMPF09, MP12, ES00, HOR05, MW92, SM98, ZCW^{+}11]. Mappings [ARR10, IF07, FR06, HR06b]. Maps [AB10, Eva79, Kun82, LLLZ12, MZT12, FJ00, FS02, Ham95b, TCCT02]. March [IEE11a, SC92, SC93, SC94, SC95, SC96, SC97, SC98, SC99, SC00, SC01c, SC03, SC04, SC06, SC07, SM08, SM09, SM10b, SM11b]. marginal [SY04]. Markers [FJ07, ISO99b]. market [Ano82, ABG^{+}94]. MARKOV [CM05, Bis94, BKR94, Bru69, Bun95, CS05, Cor86, CH87, EA98, EKT12, FR95, For99, FVP08, GFV98, HS07, HL02, JSS04, JMW09, LGO99, Liu91, LWZI12, Lle87, NYC05, OS96, RN10, Sch70, Sta67, TKR00, Tai95, Vas07, WWW06, WN10, YP03, Yu96b, ZKD05]. Markovian [JST01, SAS10]. Mars [MD95]. MASC [Tan93]. masked [BP98, SM01b]. Masking [KN11, LLPP19, Ric08, CE03, SWA94]. Maskless [YS10]. masks [YSZ11]. Massive [PGW^{+}17, YDDB15, ZCJ^{+}20, BGR01a, BGR01b, LF04]. Massively [lGAR03a, MRH11, BDS95, MT92, PM95]. master [JU10].

Match [DB08, LD10, BK93a, BBHK14, Lak00a, PFP99]. matched [CRR98]. Matching [ABM08, BM11a, CTC^{+}09, CTC^{+}10, CAL03, CD07, CW84, FT98, FT04, Gad91, HAR10, HSV^{+}08, HKS^{+}11, Jez15, KTSA99, Kid09, KRK09, KS06a, KS11b, KS11a, KØJ^{+}12, LS94, NR99, NZ95, Per06, RPE81, RNOM09, Sad96b, SF09, SKF^{+}00, SH97, YK11, ZL11, AdlPN04, AMB^{+}02, ASG99, AB92, ABF94, AJRK98, AGS96, BCD98, BA16, BFG09, BBHK14, CS92, Cha93, CR95a, CDDM05, CHW04, CLD02, CL96, DS04, De 06, FT95, FL13, GR99, HLV95, HKS^{+}13, How96, IW94a, KTS^{+}98, KS00a, KS01, KS05, KS16, KS96b, KPY96, LS04, Lar99, LO99, LPR97, LZ05, MB00b, MHM^{+}01, NKT^{+}01, NT05a, OW03b, QJRA99, RTT02, Sad93, SS06, TM04, TM05b, TM05a, TFV05, ZG02a, ZMAB03, ZA05, Neu10]. material [Ber99]. Materialize [Jag90]. Mathematical [Ano06l, Gar72, LG78, Sha48a, Sha48b, Zei93, Zei94, Ber71]. Mathematics [AG86, Cha97, Sch03, Sch04, GLM10, Hop09, Reb12, R^{+}06, R^{+}07]. Mathematik [GLM10]. mathématique [LG78]. MATLAB [Thy10]. Matrices [HO19, LKR12, BM98, GL03a, KGR06, LW21, Wat93, XTH09]. Matrix [CBK96, CXYQ12, Giv58, LWW10, MW10, BTLK18, HSV04, HV06b, NEGF96, ZGS02, DPS93]. Maximal [AMM00, HC97, PK05]. Maximally [Liu08b]. Maximizing [JTC^{+}12, VBK89]. Maximum [Ano08d, DFA^{+}10, MN08c, BBG10, KT07a, LC91, TKR03, TKRR02, BF95b]. Maxshift [TVLS08]. May [ACM78, ACM79, ACM95, Bar00b, IEE95b, IEE04, SM01a, PHG86]. Mbit [Ben96]. Mbits [ISO93]. Mbits/s [ISO93]. MBytes [PS93]. MBytes/sec [PS93]. MBZip [KPM17]. MCLT [YM08]. McMaster [Bar00a]. MCMC [BW10]. McMillan [GN05]. MCMM [EK16]. McrEngine [IMB^{+}13]. MD [VKG01]. MDL [MEIS09, RT95, Suz11, Suz12]. MDL-Based [MEIS09]. MDL/Bayesian [Suz11]. Mean [BBBP10, PG72, ARR01, ALRT16, AECG93, HGS97, KKA02, WF03]. mean-removed [AECG93]. Mean-Shift [BBBP10]. meaning [Ste12]. Means [SWW12, OW03b, PELA02]. Measure [CD08, CD10, CKS08, DØJJ09, GMSK09, LLZ09, Suz11, Yok97, Esk95, Tit00, WZ94]. Measurement [LZWL11a, LWW10, RK09b, LRH^{+}21]. Measurements [WWS11, Weg07, CR06]. measures [Ait86, BHT19, GHSV06, HLV96]. MEBAS [KS89]. Mechanism [LAP07, LYC16, HIMM20]. mechanisms [NY98]. Mechanizing [Bla09]. Media [AK08, Ano03, Ano08j, Hoc95, Hoc96, ISO93, KY10, Ma96, Bli91, CCA02, CG03, CG04, RCH04]. media-independent [Bli91]. Medical [LP07, Oh10, SSP21, BZM98, DvOL97, KF97, PW03a, SMC99, VDPL00, WF92]. medicaux [NACM08]. meet [Mil95]. Meets [Bar00b, SK10a]. mega [BKR97b]. mega-state [BKR97b]. Mel [BLS11]. Mel-Frequency [BLS11]. Members [LT09, Ma96]. Memoir [Bar05]. Memoizer [GWT10]. memories [Bet94, TN99]. Memory [BLNK14, BF12, BMM02, BLSM19, CJDL09, CEA18, DFA^{+}10, Dou93, FOF09, GKN09, KJ95, MWC01, MV16, OTW^{+}99, PB98b, PS12, Rez07a, SKY10, SW17, YGNM07, ZLC^{+}15, ZLT^{+}18, ACM^{+}15, CWL99, CO98, GCK^{+}96, GG06, JCS^{+}08, Lel91, LCY07, MLS03, MBBA91, PB98a, Pou93, RF00, TK07, Ude93, UHB09, Wak06, WLH^{+}17, YG02]. Memory- [BLSM19]. Memory-Assisted [BF12]. Memory-Efficient [FOF09, PB98b, Rez07a, YGNM07, OTW^{+}99, CWL99, PB98a]. Memoryless [Law77, SG92, SV98, XS05, YQ05]. Mention [lGAR03a]. Menus [Ano92c]. Merge [PB11, CS92]. Merged [GRG08, SBE16]. Merging [AdlPN04, YL99]. Meridian [Mer03]. Mesh [CH09, CH11, KCL06, LDZ^{+}20, MLDH15, MKSS12, MJZMA08, VCP09, VS10, AH02, BPT06, De 06, MZP06, VD18, YKLK99]. Mesh-Based [MJZMA08]. Meshes [CC06a, DJCM09, GMC^{+}06, JBG17, LI08, Ros98, BPZ99, BR05, CVDL16, CCMW05, LCB03, LV14, Szy02]. Message [AG09, CMMS17, EA98, JEEL16, KNTG08]. messages [ABA14, ASK12, ECM98, HBSASA19, RGF06]. book [Knu86c]. meteorological [KUCV04]. Method [AA71, Ara13, BT93, BM11b, Cha89, CXYQ12, Chu92, HY10, Huf52, IKM82, KE11, KCBM21, KJ10, Liu07, LF01, MM93, NMN99, OSI02, QWH12, Roe76, SI99, SY18, Tad20, Teu78, UHB09, VST11, WST95, XJ14, Yok96, AF99, ANT95a, AC98, BC05, CR91, CW15, DHH04, DYW^{+}96, ELZC84, Eff00a, FN93, FNI17, IY00, KNY96, KCL18, Kle89, KS00a, KPY96, Lak97, LMV03, LDK99, LPZJ15, LRH^{+}21, MPL02, MLMD03, MI91, NPW97b, QLQ11, QJRA99, RL02, SOI00, She90, SLZ^{+}19, Swa04b, TVW97, Von04, WLL06, Wel84a, YMYS99, ZAA00]. Methodology [CEA18]. Methods [BCH02, CK95b, FDH^{+}20, G^{+}75, GLS99, HU99a, HU99b, HH99, HV92c, Hua03, IRB97, KCCW17, LHK03, LT09, NT02, Sal02b, SH94a, SB85, Spi93, Sto88, Wic94, AALR12, Aro77, Ber99, CA02, DPS93, Del95, DJID20, DPS99, EANG04, Hor95, HV91b, KNY96, KP95a, LCL03, MR91, MBWP03, MSWB93, Sal91, SO94, Sno76, Sto96, TCP03, YGSR01, You98a, van99]. metoda [Swa04b]. Metric [GD07, Zur89]. Metrics [RL08, LLW^{+}14, Rez05, Top96]. Metro [CFGL12]. Michael [Ano00]. Michigan [IEE95b]. micro [SRG^{+}21]. micro-batching [SRG^{+}21]. Microarray [HCMGB^{+}12, HCBMSS12, FS04]. microcode [BS88b]. Microcoded [GRG08]. Microcomputer [DRT81]. MicroCT [BRNMG^{+}12]. Microprocessor [CYL^{+}08, Pro86, Ano06i]. Microsoft [Ano09f, Ude93]. Migration [NBK16]. MILC [WLH^{+}17]. MILP [HDK^{+}12]. MIME [SD02, SD06]. MIME-Compatible [SD02, SD06]. MIMO [KJ10]. MIMOLA [BS88b]. Minimal [FT00, FCDH91, Liu08b, EN84, GB92]. minimal-length [GB92]. Minimax [SW04a, MF94]. minimization [BL97, HLV94, LDK99]. Minimize [LHK03]. Minimizing [FGC93, MS02, Tin77]. Minimum [EA98, FHS08, Fou85, GPV11, Huf52, JG05, Kar61, PG72, Rud71, Sav09, ARR01, CME05, Li02a, LM02, Max60, MPL99, MT96, RMG94]. minimum-entropy [CME05]. Minimum-Redundancy [Kar61, Rud71, Sav09, LM02, MT96]. Mining [ZZJB13, Bar04, TH01, VFK15, WBMT99, YCMR18]. MINMAX [RLR04, SLR05]. misalignments [AGMP13]. mismatch [GL03b]. Mismatches [Per06]. Missions [VPGB10, PHG86, TD91]. mistake [RBR09, RBR10]. mistakes [MS06]. Mitigate [VMP^{+}16]. Mixed [Lan84, RA03]. Mixed-Order [Lan84]. Mixing [CGC11, Ged14, Mat12]. Mixture [WDR11, CMP99, DR06, GPO98, LRO97, OG06]. mixtures [HOG04, JM03, JG05]. MJPEG [CFGL12]. MLC [PM17]. MLC/TLC [PM17]. MLP [S^{+}99]. MMSE [HGFL09, PM99]. Mobile [CNW^{+}17, CTC^{+}09, Fra97, GHD^{+}14, JCS^{+}17, GLM^{+}21, SM11a]. mobile-to-mobile [SM11a]. mod gzip [Pie03]. modal [Sch06]. Modalities [Oh10]. Mode [EK16, SMS08, XSW07, LC03, MI91, PN91, Ten04]. Model [Abr89, AW15, AL10, Bar00b, BLSM19, BKR92, BH12, Cac98, Cap59, CD08, Cha96, CDW11, CST11, DX11, FGM21, GD07, HLL^{+}20, JLJ08, LZWL11a, Lan83a, Lan84, LSC10, MEIS09, MN09, MPALSS^{+}12, Raj04, Ram85, Roy87, Tad20, TR93, Tod89, WZW09, WDR11, YDZL11, ZXM^{+}10, dCDV07, ABA14, AdlPN04, ANT95a, Aus99, Aus00, BGR01a, BGR01b, BH92, CY91, DKP03, FM95a, FGGV04, HBL^{+}17, KG95, KT07a, LGS92, LF04, LGO99, LLP96, May99, MP95, NPW97a, Nov98, OG06, PMH95, RA04b, SS78, TKR03, WRCB02, Woo00, Xue99b, Yoo06, Zha98]. Model-based [Cha96, BGR01a, BGR01b, MP95, OG06, Woo00]. Model-Guided [WZW09]. model-residual [Nov98]. Modeling [AM10, ABB10, Beh08, BWC89, Blo98, BKR^{+}95, BKR97a, DXS08, DMC^{+}09, DSG12, GS12, HV92b, KSS06, Kül11, LCG11, MRH11, Ros81, Sal99, SG09, Sha06, SZ13, WLL^{+}20, Wit08, ZW12, AKF06, BW98, Blo96a, Bun97a, Hsi01, JWK98, KNTG08, Laf00, LRO97, LC91, NYC05, OWS98, Sch10, SS06, Tad98, Ton93, VT95, WCB97, WCM98, Wu99b]. Modelling [AMPHTSM10, VO08, WT05, Cor86, EA98, FN93, CH87]. Models [Cho56, Dur60, Hou79, LCV07, LAP07, LWZI12, Ram87, TC97, WSDTO11, WBN91, WN10, Abb93, ASS98, AH02, Aus98, BKR94, BB93, Bun97c, Bun98, CM91, Cha91c, Che00, Che01, CT99, Dav96, FR95, For99, FCW00, Fre98, Gin95, KPY96, LH91, McC03, MM95, OS96, PA07, RT95, RS99, SILN11, SN96a, TC96, TICH98, TC98, TH01, Vov06, WBM^{+}94]. Modems [Tho92]. Moderate [Sal91]. Modern [Knu86d]. Modification [BBH11, FH98, Pyl90, Lak00b]. Modifications [BKS99, Prz00, BY00]. Modified [Bar81, BF07, HESF11, JS99a, KA12, SPS98, SPS99, Tis87, TW01b, ZC99, CT98a, LKA99, Sad99, WB00, ZTSM05]. Modulated [Kra49, Mal07]. Modulation [ITU89, ITU90, Kno98, LLJ09, DHS01a, KA98]. modulations [BK97]. Modulator [Bar81]. Module [Ric91]. modules [Hop09]. Modulo [LMSE18]. Moiré [Roe76]. Molecular [Sch10, GSPR19, SGA^{+}16]. Monitor [Bar00a, Del93]. Monitoring [SSRM09, ZBMM97, CPT^{+}20, HIMM20]. Monkey [Ano06m]. Monochromatic [CDL13]. monochrome [RAFH92, TM91]. Monolithic [KB92]. Monotone [HB08]. Monotonic [Rez07a, Abr93]. Montreal [IEE04]. Moore [BPMA02, CFG12, NT05a]. Moore-Like [CFG12]. Morgan [Hoc96]. Morphisms [CI07]. Morphological [YR96, SS00].

Morris [DS04]. Morse [Ano06n]. Morton [BJ18]. Moscau [Gol06]. most [Ano09e, FL13, MF94]. mother [Anoxx, AC14]. Motif [AAR11, HTS^{+}18]. Motif-Based [HTS^{+}18]. Motifs [ACP04, ACP05]. Motion [CA06, CAC97, CFGL12, FWZ^{+}12, FG07, GYM93, Hal11, IDKW99, JLJ08, KKJ11, LJZ^{+}08, LZM10b, LLCC11, LZCF10, Man98, MWC01, MFT07, MJZMA08, NLU93, SMO98, SWGZ09, TSS99, TWC^{+}09, VST11, VB14, Won91, Bar90a, BK94, Cha96, CB04, CCV04, Chi98b, CHW04, CS02, Cre97, EGS91, FWY^{+}98, FWG98, FG01, Fow06, GO95, HLV94, HLV96, KG98, LPR97, LA96, MB00a, MB00b, MF98a, MM99, NO95, Nos97, WXCM99, WS91, YR96, ZR02, TBKM05]. Motion-adapted [CAC97]. Motion-Compensated [FG07, MWC01, Won91, FWG98, HLV94, LPR97]. Motivated [DSM07, LLW^{+}14]. Move [Arn99, Arn00, Kle89, FTL03]. Move-To-Front [Kle89, Arn99, Arn00]. Movies [VN08, CL06]. Moving [ISO93, ISO03b, KF94]. MP3 [Ano06j, Bel06, Bra99, Ste12]. MPC [RSMA19]. MPEG [BS94, Sof05, Ano95, Cre02, HLV97, Hoa99, KL03, Kro00, Le 91, Lev95, LR04, L^{+}05, LHS05, MEMS06, MPFL97, MRG99, Nga95, NHHS02, NU05, Pan95, PE02, RV94, Rez04, Ric03, Shl94, Sym03, TY04, TCJ93, WKC94, ZBMM97, Ano06s]. MPEG-1 [Shl94]. MPEG-2 [Kro00, MRG99, Nga95]. MPEG-4 [Sof05, KL03, LR04, L^{+}05, NU05, PE02, Rez04, Ric03, Sym03, TY04, VRHL07, VRL09]. mpeg-4.als [Ano06s]. MPEG-7 [NHHS02]. MPEG/Audio [Pan95]. MPI [BKK20, FSC^{+}11, VPS17]. MPI-Based [FSC^{+}11]. MPI-OpenMP [VPS17]. MPJ [Joh89]. MPM [KYPY98]. MR [RHC87, MSR04]. MRA [Fu03, Uhl96]. MRCSI [WL15]. MRF [ZL00]. MRI [GSPR19]. MS [MKW^{+}06, Nel87, RK09a, Ude93]. MS-DOS [Nel87, Ude93]. MS-SSIM [RK09a]. MSE [KUCV04, Top96]. MSEC [Xue99b]. MST [NL06]. Mukherjee [Neu10]. Muller [RP88]. mult [Chi98a]. mult-lingual [Chi98a]. Multi [BZ09, Bel11, BKRSR09, BBBP10, Che77, FJD11, GGG^{+}09, LBMN08, LZZ^{+}12, MC13, MTD08, MP06, SR08, Sch06, Sim92, Tan93, VFE00, WWSY07, ZLRZ09, ATL^{+}88, CZ99, Cho96, Eff98, Esk95, Fen93, FWG98, GE06, KF02, MS03, MSWB93, OW03b, Per89, Per90, PFP99, QT03, Sau95, SZM03, WD02, Xue99b, Xue99a, You98b]. Multi-Band [MP06]. multi-bit [Fen93]. multi-component [KF02]. multi-corpora [SZM03]. Multi-Criteria [BKRSR09]. Multi-dimensional [LBMN08, Cho96, Esk95, Sau95]. multi-fractal [Sch06]. multi-hypothesis [FWG98]. multi-indexed [You98b]. Multi-Level [Che77, MTD08]. multi-lingual [CZ99]. multi-match [PFP99]. Multi-modal [Sch06]. Multi-protocol [Sim92, Per89, Per90]. Multi-Pulse [Tan93]. multi-purpose [ATL^{+}88]. Multi-resolution [Bel11, BBBP10, VFE00, Eff98, GE06, OW03b, QT03, WD02]. Multi-scale [LZZ^{+}12, Xue99b, Xue99a]. Multi-Sided [MC13]. Multi-spectral [GGG^{+}09]. Multi-Stage [WWSY07, SR08, MS03]. multi-symbol [MSWB93]. Multi-view [FJD11, ZLRZ09]. Multialphabet [PS93]. Multiband [TY04]. Multibit [MBBA91]. Multiblock [KPM17]. multicast [CMWM00]. multicasting [SX04]. Multichannel [IRB97]. Multicomputers [HU99a, HU99b, LCY07]. Multicore [TH10, SHC^{+}16, SJPB17]. multidepth [KMP05]. Multidimensional [HM94, MM05, RKL16, LCL03, Sim98]. Multigigabit [PP10]. multihop [GE05]. Multihypothesis [FG01, TF11, Fow06]. Multilayer [Hua91]. multilevel [JCS^{+}08]. multilevel-cell [JCS^{+}08]. Multilingual [CK13, CKCW95, CK06]. multilist [CC93]. Multimedia [Bar12, BRV08, Che99, Col93, Fet96, FWI99, GBL^{+}98, Kut02, LSH12, PMZ13, Ric03, Spi93, Ber99, ITU06a, Laf00, Le 91, LSD97, Sof05]. multimedia-based [Ber99]. Multimedia-the [Col93]. Multimodal [AAL^{+}21]. Multimode [RRG97, Lin98]. Multiphase [EC16]. Multiple [AK08, BZZ08, BZL^{+}12, BT17, BZ09, BC10, CTD07, CWLS15, CMW99, CGC97, DZ08, Dum08, FWS^{+}08, FWS11, FOSS17, GE07, HDCH09, HB09, HLV95, Hoo78, JR02, KGK00, KTS^{+}98, KØZ08, MNG03, ØZ07, PAB03, SL08, SVS99, SB97b, TM05b, TMD08, WL08, WS06, Wol99a, Wol99c, YLZ11, YWC09, ZKK09, ZKØ10, ZB12, BDV00, Bar94, BS02, Car01, DSV00a, DSV00b, DW05, FZ05, FE99, FDZ00, GK98, HW06, JA03, Lak02, LLN^{+}04, LZZ98, MR00, Mit01, Moh02, NC97, Ort96, ØJH05, PPR02, SAR00, SH03, Sri99, Str99b, TH03, TH04, TH05, VFE00, WL15, Whe97, WJBM05, WWW06, ZL00, ZE01, ZDR06]. multiple-choice [Moh02]. Multiple-Description [ØZ07, ØJH05, SAR00]. Multiple-Description-Based [AK08]. Multiple-dictionary [HLV95]. Multiple-image [MNG03]. Multiple-pattern [TM05b]. multiple-resolution [Str99b]. Multiplexed [Hua01, YK10, Che01]. Multiplexing [EY81, DW03]. Multiplication [DKCS95, HESF11, Huy94, CKW91, FGC93, van99]. Multiplication-Free [HESF11, DKCS95, van99]. Multiplications [LLM89]. Multiplier [LWZI12]. Multiprocessor [ALYK12, LLCC95, Wak06]. Multiresolution [AH92, ACD02, ABM10, BCH02, CA06, Cie98, DMS08, KCL06, Mal89, PD16, SBE16, Tod89, ZFE04, ALRT16, Bon95, GZ91, PGS10, SW04a, Tun04]. multiresolution-based [PGS10]. Multiresolutional [KF95, KF94]. Multiscale [BCH02, MMZ12, Mil95, RS16, SZ13, SMB98, FC98, LW14]. Multisensor [Hua03, Ces91]. Multisensory [WMNP12]. multiset [Arn96]. Multispectral [AR08, EHSC92, MWCG07, MR93a, Rea72, TS17, Arn97, CM99, MT92, Tat94, TLLB06]. multistage [KMS96, MR93c, TCCT02]. Multistream [KLUZ08]. Multiterminal [AR12b, YSXZ04, YSXZ05]. multiuser [MS02]. Multivariate [WWY^{+}07, PGS10]. Multiview [EC16, JM97]. Multiwavelet [KL01, RM21]. multiwavelets [FH02]. Munich [BDLS96]. Music [Moc04, FK95]. Musical [BI98, IMA06]. Mutual [DFA^{+}10, BSS05, TG17]. MVP [Gov94]. mysteries [Ano06l].

N [FR06, RF05]. N-description [RF05]. Naïve [Per03b]. Narrowband [Bis95]. NASA [TD91]. National [Ano06c]. NATO [Fis95b]. Natural [FWS11, FNP08, FCL13, LG78, RL99, Sto80, BFMX03, KZ02, LY97, Rob91a, SS00, YWMS08, Zec72]. Naturalis [dP58]. Nature [Man82]. naturel [LG78]. naturels [Zec72]. Navigation [DDT^{+}12, Jog73, MPALSS^{+}12]. NC [IA00]. Near [BDS95, GB06, LMC05, Mal07, MTTH13, Rob94b, ZZS^{+}22, HW98, Kli00, MT98, Rob94a, ZCG99, ZE01, ZE03, ZGB04, van02b]. Near-Lossless [Mal07, Rob94b, GB06, MTTH13, HW98, Kli00, Rob94a, ZCG99, ZE03, ZGB04, van02b]. nearest [Goo03, Sau95, Sou95]. Nearly [FL98, LV92, OG12, Ric10, OG16, VN90b]. Nearly-Optimal [OG12]. Needed [Bar00b]. needs [Sch13]. Neighbor [NM91a, Goo03, Sau95, Sou95]. neighborhood [LW11a]. Neighbourhood [VS10, YW04]. nested [LCLX04]. Net [Bar00a]. Netherlands [vL94]. Netlist [YSM08]. Nets [VO08]. Network [Ano97, ABS^{+}16, BZZ08, Bi08, Cro95, EC16, FZE03, FE01, GN79, JGM^{+}20, KR10, LZX^{+}09, LTO11, LSC10, LE02, Log04, LXG^{+}18, MNG03, Roe03, SQ16, SY18, Suz12, TOL11, Wan08, WB17, Xu06, YLZ11, BSF16, BK94, BK92, CS13, EM02, FSC91, Goo91, GE05, GE06, HSP^{+}13, IA00, ME02, Nam91, OG05, PN91, RK93b, SW06, Sha94, Sha06, XWDY06, YA13, Zhu02]. Network-Adaptive [YLZ11]. Network-Based [LSC10]. Network-Coding-Based [SQ16]. network-conscious [IA00]. Network-Source [LZX^{+}09, SW06]. network-wide [BSF16]. Networked [Won85, ZHX^{+}19]. Networking [ZMVR14]. Networks [AH14, AsM^{+}10, DY12, HSZ17, HTS^{+}18, HP95, Kla98a, Kla98b, LL16, LRMPGR01, MW10, MKK12, MCL12, Nus76, SF11, Sar12, SSRM09, SMS91, TCN^{+}17, Tan02, TTW07, WHB16, WB17, XJ14, YZLH16, ZYT^{+}15, Abb93, AL81, CZZR19, GD06, Ham93, HR06b, Hes08, HM07, HN02, JB95, JKV15, KLMXL05, KLJ12, Lev96, LCJP21, LGLK05, LLWC08, OS01, OG06, PKG08, PD07, PR00, RBD13, SF05, SSZZ14, SKLA12, YK20, ZRR00, ZE01, CGSS15]. Neural [Bar00a, Ham93, JGM^{+}20, Kra91, LE02, Log04, Log06, Nam91, SAS10, Abb93, BK94, BK92, FSC91, Goo91, Lev96, RK93b]. neuro [ZGB04]. neuro-anatomical [ZGB04]. Nevada [ACM95]. News [Bar00a, Bar00b, Ano92c]. Newspaper [IKM82]. Next [Car20, Ric03, SCV11, Sof05, VPS17, ZdMNBY00, NSFO^{+}99]. Next-Generation [Car20, SCV11, VPS17, ZdMNBY00, Ric03, Sof05, NSFO^{+}99]. NGSC [GJ20]. Nine [Mac12]. NLG [Ano98]. NLTL [Moh04]. NMR [GSPR19]. no [Sam89]. no-loss [Sam89]. Node [RLG09, HS97]. Noise [AsM^{+}10, BT93, DMC^{+}09, FR06, KØZ08, MGBRMSS11, BAL02, sKJ01, LLZ96, MOT^{+}03, MB02, Nat93, ZF94]. noise-robust [MB02]. Noise-Shaped [KØZ08]. Noiseless [Cat84, Kar61, Ric79, Ric91, Tun67, ABF^{+}91, ECG94]. Noiselets [CWY^{+}10]. noising [SCC12]. Noisy [BH07, GFE07, HB09, KH08, QB09, SL10, ASM96, CF96, EGNA06, GF01, MZ99, PBC05, PAB03, RMG05, RRG97, SZ97, ZAA00]. noisy-varying [PAB03]. Nombres [Zec72]. Non [ASS98, Cm10, DXS08, DGG10, HF97, IF06, KTMS10, KS08, Moh04, QLH12, RJL11, SWW12, Uhl96, AAS95, DGG09, Hor95, KC91, MA05, PELA02, SF05, Suz96, XS05, ZGF02]. Non-asymptotic [IF06]. Non-Binary [KTMS10, ZGF02]. non-equal [PELA02]. Non-existence [Cm10]. non-greedy [Hor95]. Non-intrusive [QLH12]. Non-Linear [Moh04, HF97, RJL11]. Non-local [SWW12]. non-planar [KC91]. Non-sequential [DXS08]. Non-sorted [KS08]. Non-Stationary [Uhl96, AAS95]. non-systematic [XS05]. Non-threshold [DGG10]. non-tree [Suz96]. non-treshold [DGG09]. Non-uniform [ASS98, MA05, SF05]. nonaligned [SS04]. Noncausal [MR97]. Nonlinear [BG06, BT93, Fu03, ALRT16, BM97, CZ98b, CDGO02, FR06, Gul04, SN94, WV00]. Nonnegative [NK20]. Nonorthogonal [KGR06]. Nonstandard [SM97]. Nonstationary [GH08, Zan92]. Nonuniform [AR09a, BRD10, BM11b, CN07, LTB04]. normal [LCB03]. normalization [JYHC03]. Normalized [KT07a, YTY12, TKR03]. Norway [Fis95b]. notation [Poy08]. Note [Col85, CLM14, Lan83a, BM89, HSV04]. Notes [Pin99, ES00]. Nouvelle [Sie90]. Novel [CHMW10, CST11, FWS^{+}08, FK85a, SMS08, VST11, WHZ12, WB91, YDZL11, AB08, CSM00, DB03, OHRS21, SSZZ14, Sun16, TG17, WL99]. November [IEE88b, M^{+}98]. NP [RH97b]. NP-hard [RH97b]. NQ [LCLX04]. NTFS.com [NTF06]. nuclear [MM05]. Nucleotide [CW07, KT07b]. Number [CZ98a, CFY^{+}10, UUiN12, Wei07, HSV04]. Numbers [FOP12, Sto80, Vor61, Vor83, Zec72]. Numeric [Log04]. Numerical [DV01, PF^{+}88, USY17, EFF00b, Har04, Yok93]. numerics [AFL96]. numérique [Har04]. NVEs [BPT06]. NVMs [PM17]. Nyquist [Wik03].

O [Bec93, CNPC98, CNCC98, CD95, DLY^{+}98, HD88, SKY10]. O-Efficient [CHSV10]. Obfuscation [ZL19]. Object [BZ08, FOhC09, Fra97, HV08, PK10, Wic03, AF04, KL03, Sch05, Sch06, XWC^{+}17]. Object-Based [BZ08, HV08, Sch05, Sch06]. Object-Oriented [Wic03, XWC^{+}17]. Objective [RL08, WGZ^{+}18]. Objects [DHN85, SWGZ09, KL97, TFV05, TCCT02]. Oblique [FOhC09]. Oblivious [FV16]. obliviously [FGG^{+}08]. Obscura [Hae96]. observers [EW96]. obtained [BK97]. Occam [Nat94]. Occlusion [SC11]. Occupancy [MŠH12]. Occurrences [BKR97a, BKR^{+}95]. ocean [BB93]. October [BDLS96, IEE84, IEE91, IEE95c, IEE96, Ano88a, Ano88b, Ano88c]. Octopus [AC07]. Octree [ZO04, Bet94]. Octree-based [ZO04]. Odd [Bel11]. Off [ACP05, KWC^{+}16, LM00, NBIT07, AL98, AL00, DS96, GLM^{+}21, Ruf95]. Off-Line [LM00, NBIT07, ACP05, AL98, AL00, DS96]. off-the-shelf [GLM^{+}21]. Offline [BMM02, CR07, LM99]. Offs [KN11, MCM^{+}17, BEGV18]. Ogg [Ano06o]. Oja [SPS98, SPS99]. Oja-RLS [SPS98, SPS99]. old [Bal97, CBC00]. Omnidirectionally [FH02]. On-board [TLLB06, lGPL^{+}00]. On-Chip [LHK03, ZRZ^{+}19, CR07]. On-Device [GHD^{+}14]. on-ground [Tur75]. On-Line [AJ96b, BGPW96a, BGPW97, MM93, Slo06, B^{+}92, CS93, DS96, Dou93, FWY^{+}98, LL02, NLW96, SST^{+}98, SSM02, WO00, AJ96a, Bun97a, Cha00, Pap99, Vov06]. On-The-Fly [Eng93, KS06b, BGM99, RL09, Ano92c]. One [CCD^{+}19, HKS^{+}11, Jak88, RBR09, RBR10, Zha11, CH06, EANG04, MGG05, Shk02]. one-bit [MGG05]. One-Helper [Zha11]. One-inclusion [RBR09, RBR10]. One-pass [CCD^{+}19, Jak88]. one-to-one [CH06]. Online [Bar00b, BSSU18, CS94b, DY12, HKKW13, Ign98, KMM^{+}12, NHK06, NdR04, SAS10, TDL^{+}19, WSS94, CS13, DHW^{+}17, FL13]. only [BGU96, Sch13]. onto [CMW99]. Open [Cal06a, Haf01, Roe03, CT99, Fow00]. Open-Source [Haf01, Fow00]. OpenMP [NU05, VPS17]. Opens [Kro00]. operand [RS04]. operand-factorization [RS04]. operating [Ude93]. Operational [LZR^{+}10, LZWL11b]. Operations [KLR^{+}95, LCL03, Tur96]. operator [Bon95, NO95]. operator-based [NO95]. Operators [DV01, FP20, MI91]. OPM [Bel86]. OPM/L [Bel86]. opportunistic [GAS16]. OPT [RL95, RL96]. Optical [RY92, MM99]. Optimal [AKS10, Adj06, ABF94, AC07, BFT11, BH07, BGBV09, BCSV06, BY76, BNK93, CCMW03, CAL03, Dos72, DW03, EJ07, FJ07, FVW10, FL98, FY95, GvV75, GO95, GN01, GK98, GL09, HS07, HW06, JM03, KN90, KKN07, KY08, KM11, Kle97, KRSS21, Kli00, KF03, LZR^{+}10, LZWL11b, MGG05, MGL11, MOG05, MF04, NYJ08, NYJ09, Ort96, OG12, PWS10, RS98, RK09a, Ric10, RH97b, Sad96a, SK10a, Sco02, SR97b, SV98, SY04, TEGM03, Use96, VK91, VK96, Wan88, WX00, WMWW10, WY10, XD11, YO91, YX12, ZE01, ARR01, DS92, AALR12, AL01, ALM^{+}18, AR06, Att95, BDS95, BR04, CH06, De 00a, De 03, DSV00b, DWW02, DWB04, Dum06, Eff97, EM02, FGMS05, FWG98, FWA01, GSS01, GR99, Ghi03, GSPR19, HSX02, HC98, HAH05, KN05, KT02]. optimal [KRT05, KUCV04, KMS98, Lak02, LSC91, LV92, LPR97, Mit01, MTK95, ME02, OG16, PD16, Raj04, RMZG03, RCH04, DP91, Sub99, Tun04, VFK15, Wat93, WZL91, Wu93, WF95, WD02, YJE04, ZE00, ZAA00]. optimal-joint-coordinate [LPR97]. Optimality [AR12b, AMS^{+}03, CLM14, DW05, DBTNT00, Sad98b]. Optimally [FNV11]. Optimization [BRALSSMP08, CPL09, DC18, Dum08, FO10, Fra97, FB16, Kut02, LJ07, LT08b, LWZI12, RH07, SHK99, SSRM09, SjWL08, VRHL07, YDZL11, YK20, ZLX^{+}20, CR95b, GC16, HM93, LG01a, LG01b, MZ91, Pou93, RV94, RL96, SH03, SS06, WS00, Wu92]. Optimizations [NBK16, Wak07]. Optimized [ARR10, ALM11, CT09, CCMW05, Hat95, Knu98, Szy02, WXC^{+}10, BK92, CCA02, CG03, CG04, DC97, FZ02, FNI17, GK99, GK95, LM08, MYRD04, RSV^{+}00, SHWH12, Top96, VB14, Vuc05, WHH^{+}99]. Optimizing [CK94a, OSC14, TDL^{+}19, WOS06, RL95]. Optimum [BY90, Bru69, ECBL12, Hoo78, LD10, MS93, Rud71, VPGB10, WMH96, Stu94b, TR88]. options [PH90, TVLS08, Ven17]. oracle [LL02, PP03]. orbit [LM91]. Orchestra [Haw90a, Haw90b]. Order [ALM11, Bru69, CCKH21, FCDH91, KMH10, Lan84, MPAFF10, NR08, PR11, Tad20, XXZ11, Ant97, Bis94, FGGV04, Gin95, KT07a, LKG12, LOMSS05, ML98, NYC05, She90, Wu99b, ZIL93]. order-1 [KT07a]. order-preserving [Ant97, LOMSS05]. Ordering [BJCO03, BŁ01, KF03, Sab94, Tat94, WZL91]. Oregon [Auc98]. Organization [ISO03a, BK94]. Organized [JB74, MSOY96, OSMY95]. organizing [Ano92c, Goo91, Ham95b, RKSM03]. orientation [CSP05]. Oriented [Fon81, LP75, PH11, Ros94, Wic03, CW91, CKW93, Cha04, Gu05, HGS97, LKR17, XWC^{+}17]. Original [DR83]. Origins [Hea72]. Orlando [IEE88b]. Orleans [ACM97]. ORT [GJ20]. ortho [LK96b]. Orthogonal [ASH87, FG07, TOL11, ZZS^{+}22, Haa10, Haa12, MSW96, MS03]. Orthogonalen [Haa10, Haa12]. Orthonormal [BK97, Dau88, HGR11]. oscillatory [HSV04, HV06b]. Other [Fen02, Tho91b, Whi06]. Oulipo [Mot98]. Our [JPE00, Pou93, Bry95]. Out-of-Core [DJCM09, ILRS03]. Out-of-loop [Cre97]. Outlier [TV09]. Output [Pie03, BBQ^{+}95, LW04, NEGF96, Sab94]. Overcomplete [YM08, BLO01, CMW99, GVT95]. Overflow [NM10]. Overhead [HN02, LHK03, Les88, ZLT^{+}18, BKR97b]. Overhead-constrained [HN02]. Overlap [RS01]. Overlapped [FOhC09, FO10]. overload [JEEL16]. overlook [Ano86b]. oversampled [CD00, CDL07, DKG01]. Overview [Doz91, Lio91, LW08, Man98, PMLA88, Reg81, SMW07, Wol99c, MGBB00]. Own [SS96, FS03].

P [De 00b, Sie90, Lio91, AF88b, De 94a, Mor12]. P-complete [De 00b, De 94a]. P-Compressed [AF88b]. P2P [SQ16]. PA [KRW^{+}93, IEE88a]. Pacific [M^{+}98]. Package [BKK20]. packages [KH00]. Packed [Rok85, ZZ98, OTY^{+}97]. Packed-TS [ZZ98]. Packer [UPX03]. Packet [ALYK12, BZZ08, BSF16, CT09, CC19, CXZD12, HKW10, LSC10, MAS00, SX04, Wic94, WFG^{+}94, ZXM^{+}10, AL81, BBF02, CG03, CF97, HLR01, HV06b, MASC98, MRL99, PBEA95, SLZ^{+}19, YJE04, YML97, ZRR00]. packet-based [PBEA95]. Packet-level [BSF16]. packetised [Woo00]. Packetization [SMS91, XD11, DWW02, MYRD04, RC98, WX00]. packetized [RCH04]. Packets [CMQW90, CMQW93, CMQW94, Uhl95, XD11, KJP99]. Packing [Whi06]. Padding [ABP09]. Page [CNW^{+}17, Ano06d, Ano08f, Bus09, DVB06, Dol06a, Dol06b, Lau09, Mer03, NTF06, PKW03, Pav06, Ros06, Soc08a, Tan08, Uni07a, Uni03, Wav06, Win03, XML03]. pages [Ano98, IMA06]. Painless [Wil93]. pair [NR08]. Pairs [HGR11, NM96a]. Palette [FH98, Aus98, FS02, Li02b, Ste06]. Paletted [Sch11]. Palindromic [JMnRS19]. Palo [ACM13, BBKF95]. panoramic [TCCT02]. Papers [CS13, ACM78, ACM79, Sle74]. PAQ [Ano08e]. PAQ8 [KdF12]. paradigm [PR01]. paradigms [CS13]. paralinguistic [BYJ20]. Parallel [AMPHTSM10, DS92, BKK20, BKS^{+}18, BR09b, CNPC98, CNCC98, CGC97, CAL03, CD07, CDL09, CDL10, DLY^{+}98, De 11, DCP17, DKS21, EV14, ECBL12, FRT96, GS85, GP95, HV92a, HV96, JM96, JBG17, KW00, KW01, KW03, Kut02, LP95, Lea78, MRH11, MKSS12, MNM04, MBWP03, NLW95, NLW99, NEGF96, PS06, SKY10, SW12b, Sta94, TSS99, TN99, TB11, Uhl95, Uhl96, UH96, UUiN12, Wak07, You98b, BDS95, CS92, CY91, CLD02, CD95, De 00a, GLI16, LLCC95, LXG03, MV20, MT92, MLP99b, PBPT95, PM95, Sha06, SR91a, Tin89, Wak05, WLCZ15, WZL^{+}17, vB97, BDLS96, CDL13, SJPB17]. Parallelization [PGW^{+}17, Wak05, JKL13]. Parallelized [SCW11]. Parameter [Kie04, KL05, TLR85, VRL09]. Parameterization [MP12]. parameterizations [YSZ11]. Parameterized [LPG10, OS96, BA16, SCC12]. Parameters [Mar80, Abe99, BdB06, DK95, GFV98, GSPR19]. Parametric [BF11, CDW11, MN09, Nos97, SLM^{+}17, VCP09, VN08, Wit08, WDR11, KZ02]. PARCOR [KMH10]. Parentheses [GF10]. Parity [WY10, WS05]. Parity-Check [WY10]. parsable [Sav99]. Parse [YK10, Tar95]. parses [Apo05]. Parsing [BW94a, CK96, HS94, IM01, KVP20, Lan84, MRS01, NLW99, AALR12, ALM^{+}18, CLM14, Eva98, Hor95, JS95, JST01, LS95b, MRS99, Mod03, NLW96, OWW^{+}04, RSV^{+}00, WO00, Yao03]. parsings [FLMPR06, HC97]. Part [BF11, Cha95a, Cha95b, CLA12, ISO93, ISO03b, Lin95b, Lin95c, MEIS09, Phi91, SDS95, KUCV04, Fow97a, Ric91, Wil70, Xue99b, Xue99a]. Partial [BCC^{+}18, CW84, DH18, JLJ08, KRK09, SMS08, AdlPN04, GDV03, HLV95, KK03, LS04, LO99, ZA05]. partial-matching [LS04]. Partially [Oka05, Pom16, BP98, FR95]. particle [YGSR01]. Particular [Tsu72]. partition [NM91b, PELA02, WLL06]. Partitioned [Kur83, MMK98, LAB^{+}14]. Partitioning [EGLS21, FNV11, OVCS^{+}09, SP96, SM14, CC03, CCMW03, KP97, LKH^{+}00, OY06, Sai05a, SHC^{+}16, SLXG04, TCP03]. partitioning-based [SHC^{+}16]. Partitions [FM95b]. Pass [CS94a, DJID20, CCD^{+}19, CS95, Jak88, MRG99, RSC99, Xue99b, Xue99a]. Pass-efficient [DJID20]. password [BCR98]. Patch [NK20, Per05]. Patent [Rod95]. Patentable [Zal88]. Patents [lGAR03a, Wil06]. Path [GO15, JP07, SS05a, Tar79, ZG02a, Bar94, CG03, GN79, LZ05, WL99]. PathFinder [MD95]. Pattern [AMB^{+}02, ABM08, AGS96, BCD98, BM96, CDL13, CL96, HAR10, Jez15, KTSA99, Kid09, KRK09, KS01, KS06a, MP06, NR99, Neu10, OW03b, SSP06, SKF^{+}00, TM05a, WZY^{+}11, WS08, YK11, ZD95a, ZD95b, ASG99, BA16, Bon95, BBHK14, Cha93, CR95a, DS04, DSSR95a, GR99, GW05, GN01, Goo91, How96, KTS^{+}98, KS05, MHM^{+}01, R^{+}06, R^{+}07, TM04, TM05b, ZMAB03]. Pattern-Based [MP06, ZD95a, ZD95b, BM96]. Pattern-Driven [WZY^{+}11]. Pattern-matching [CL96]. Patterns [Ano86a, BHLY20, Sha04, YCMR18]. PC [Del93, Yu95, Yu96a]. PCA [KCCW17, SGA^{+}16]. PCA-based [SGA^{+}16]. PCA/DCA [KCCW17]. PCM [BLNK14, ITU89, Llo82, PG72]. PDE [CCO^{+}19]. PDEs [CZ98b]. PDF [Ado06, ALMSS09, FS03]. PDF-Based [ALMSS09]. Peak [Smi79, CA02]. Peano [Col85, SA93]. Pearl [Bla09]. PeaZip [Tan08]. pebbling [FP19]. Peer [BRV08, LZX^{+}08, YTY12].

Peer-to-Peer [BRV08, LZX^{+}08]. Pel [FG07, LLCC11]. Pennsylvania [Bar00b]. Perception [HKMS08]. Perception-based [HKMS08]. Perceptrons [VCW09, Hua91]. Perceptual [CB97, GMSK09, LLZ09, Mar02, Mor12, PHM^{+}14, SjWL08, WGZ^{+}18, YDZL11, Chi98b, Chi99, Cre02, KK08, PMH95, RH97a]. perceptual-based [Chi99, RH97a]. Perceptually [HM97, KC06, LV96, RAFH92, CVC95, DHS^{+}01b]. Perceptually-weighted [KC06]. percolating [Bun97c]. Perfect [CW91, Cm10, FCDH91]. Performance [AK99, AMPF09, Ait86, AW04, AW15, CYL^{+}08, CWY^{+}10, DY91, FSC^{+}11, GLI16, JT96, Kam98, KT81, KGG^{+}14, LZ05, LJZ^{+}08, LE01, MV20, MC13, MZ99, OS03, Raj20, SQ16, SG12, TY08, Tau00, USY17, Wel84b, WFLZ09, WMWW10, YK11, ZYF^{+}20, ZZS^{+}22, ZLT^{+}18, ZMVR14, ZLX^{+}20, ABA14, ASK12, AC14, CBR14, Eff00a, EGCK98, FBX91, HBSASA19, JLJ^{+}20, KFSM04, LBH91, LS93, LCLX04, MLS03, MOG05, MZ95, dlCNLA^{+}95, NU05, PST05, RG02, RK93b, RSV^{+}00, RCMS04, Ruf94, SV00, SDJ04a, Sew00, SLJ^{+}15, Tit02, TK99, Top96, WZL91, WM94, XG97b, YQ05, YYO04, Zee98, ZE00]. performances [Lan98]. Performing [MSM09, OW03a]. Periodic [Mun91, Pom16, SMO98, AC98]. periodicity [CA02]. periodicity-exploiting [CA02]. Permutation [Fel74, Arn99, FBS04, WS06]. permutations [Arn96, Arn97]. permuter [Ste06]. persistent [KV19]. person [SK04]. Personal [Bar05, Haw90a, Haw90b, Cra96]. Perspective [KdF12, Ma96, SB06]. perturbed [GLL04]. PEs [GRG08]. Peter [Ano00]. pFPC [BR09b]. PGM [FV20]. PGM-index [FV20]. PGV [GBK16]. Phase [BLNK14, ZWZW12, YYZ12]. phen [Lia84]. Philadelphia [IEE88a]. Phoneme [Li86]. Photo [TCKM10]. Phrase [KS00b, NMW98, Yao03, NMWO96, SP98]. PHY [GJ20]. PHY-NGSC-Based [GJ20]. Physical [BWB^{+}20, Dou93, VK17]. physically [LLW^{+}14]. Physics [AVE18]. pick [Bry95]. Picking [Wei07]. Pictorial [Lee91]. Picture [NL80, Ram85, Wic90, Wic92, Wic94, Wil70, Zor88, Mur96, Wal91b]. Pictures [Cap59, Gha87, ISO93, Kno80, NLU93, Wan73, ER87, Gha84, ISO03b, KF94]. PIDS [JLJ^{+}20]. Pie [Bas92]. Piecewise [LWZI12, LLY00, PGS10, SN96a, SM98, TR91, Aus99, BNK93, KF95, SC01b, XS05]. piecewise-constant [Aus99]. piecewise-linear [BNK93, KF95]. PIFS [LLP96]. Pin [Bec93]. Pipeline [BKRSR09, Car97b, Hal11, TWCA02]. pipelined [BDCC97, JM96]. Pipelining [BM90, TH10]. Pisano [Gri73]. Pittsburgh [KRW^{+}93]. Pixel [BJCO03, KXPZ12, LZPB11, LZCF10, Lyo09, MD95, Smi09, VST11, SYP04]. Pixel-Value [KXPZ12]. PK [Rok85]. Place [SS05b, MGM13, SS03, SS04]. Placement [AK08]. Plaine [Sie90]. Plaintext [Kel02, BH93b]. Planar [Pea90, Sie12, KC91]. Plane [AMAM13, FOhC09, Giv58, JTC^{+}12, SWGZ09, BK96, DB03, KSCE16, Pea90, Sie12, Taw93, Use03]. planes [GM99]. planning [PW03a]. Plants [PL90, Smi84]. Platform [Dio93, SHW^{+}01, JCS^{+}08]. platforms [Ano09f]. play [LS95a]. Playback [BGU96, HKA06]. Plaza [IEE88a]. Pleno [EFPS16]. PLHaar [SLDJ05]. Plots [FM89]. Plotting [Mul96]. plurally [Sav99]. Plus [DSG12, BG96]. PNG [Ano97, Ano00, Car97b, Cro95, Mia99, Roe99, Roe03, Ste06]. PNN [Wak05]. POCLib [ZZS^{+}22]. POCS [YML97]. PODS’08 [LL08]. Point [ABM10, BR07, BR09a, BR09b, DFH^{+}19, FM12, FDH^{+}20, IEE85, Kir04, LHD^{+}20, Lin14, PBGA19, Per89, Per90, PH90, SKY10, Sim92, Sim93a, Sim94b, USY17, CDL^{+}19, FOSS17, Kir05, MGM13, Pat94, PWS10, RKB06, Tro96, TMM96, Use03, Use05, Sim94d]. point-set [Kir05]. Point-to-Point [Per89, Per90, PH90, Sim92, Sim93a, Sim94b, Sim94d]. Point-Value [ABM10]. Points [CI07, JTC^{+}12, KKJ11]. pointwise [KY99]. Poisson [CKS08]. polar [BG06]. Polarimetric [ZWZW12]. policies [CCA02, CG04]. Policy [ABS^{+}16, QB09]. Policy-Aware [ABS^{+}16]. Polish [Swa04b]. Poly [PZ12]. Poly-DWT [PZ12]. Polygon [Lev95]. Polygon-Assisted [Lev95]. Polygonal [WKC94]. Polygons [Col85, Col86]. Polyharmonic [KKT10]. Polymorphic [PZ12]. Polynomial [AVE18, DV01, San89, DF93]. polynomials [ML98]. Polyphonic [NR14]. pools [Sau96]. Population [CLS19]. Portable [Ano97, Roe03, Ado01, Cro95]. portfolio [Cov96]. Portland [Auc98]. Portuguese [dABdSFNA08]. Position [BP09, DS08, KK08]. Positioning [FJ07]. Positive [FB16, San89, YO91, Yam00]. Positron [CFC05]. possibilities [Ber99]. Possible [RC92]. Post [Abe10, BFT11, FC98, ML98, Abe05, Abe07a, CSCW00, NT20, Nos99, WMH96]. Post-Compression [BFT11]. post-filtering [DLT05]. post-filters [WMH96]. Post-processing [FC98, ML98, CSCW00, Nos99]. post-transformation [NT20]. poster [BP00, Vol97]. Posterior [SF09, PGOO94]. Posteriori [BF95b, XTH09, BB95]. postfiltering [JOC97, LMS94]. postprocessing [ZCG99]. Potential [Mot98]. Power [Ano03, Bi08, EJ07, IF07, KL07, KWC^{+}16, MCM^{+}17, SFT03, Yan03, ZRZ^{+}19, Ind91, LHW00, LW00, NT05b, TFA93]. Power-Off [KWC^{+}16]. Powered [FIKS18, KL07]. Powerful [GGM12, Jez16, Ano06m, Ano09e]. pp [Hoc95, Hoc96, Ma96]. PPM [ASS97, ASS98, AdlF06, AMPF09, BFMX03, Bun97b, Che01, CTW95, CT97, DK02, DKP03, Eff00a, Fen12, HZKL11, KT09, KE11, Mof90, Oka05, OSI02, Shk02, SG04, TC96, TICH98, TH01]. PPM* [IY00]. PPM*-style [IY00]. PPM-Based [KT09, TC96]. PPMexe [DK02, DKV07]. PPP [PH90, Ran96, Sim92, Sim93b, Sim93a, Sim94c, Sim94a, Sim94b, Sim94d, Sim97, Woo96]. Practical [AN10, ACF^{+}21, Boy91, BB92, CCF^{+}18, CL06, CDL09, CKV93, Eff98, FHS08, FMP83, HV92d, JKL13, Jon91, LLM89, MCHAM08, NR99, Ng10, Rez07b, Ric79, Ric91, SKLA12, Wan73, BDS95, CLME04, CME05, LM03, PH91, SW06]. practicality [Shk02]. Practice [CHSV10, Has18, TM02, AL98, BK00]. Prague [HŘŠ11]. PRAM [De 03, HS94]. Pratt [DS04]. Pre [BA08, JRALMSS11, OLHL08, ZCW^{+}11, Cra98b, Cra98c, DLT05, VPL98, WMH96]. pre- [DLT05, WMH96]. pre-/post-filtering [DLT05]. Pre-encoded [JRALMSS11]. Pre-mapping [ZCW^{+}11]. pre-press [VPL98]. Pre-processing [BA08]. pre-processors [Cra98b, Cra98c]. Pre-Specifiable [OLHL08]. PRECI [BD82]. PRECIS [AC98]. Precise [DHS02]. Precision [BR07, BR09a, FOP12, Rez08, Rub79a, DE92, Kra10, SST^{+}98]. Precision-Redundancy [Rez08]. precomputed [HAH05]. Preconditioning [KL01]. predetermined [MSOY96]. Predicate [VO08]. Predicate/Transition [VO08]. Predicate/Transition-Nets [VO08]. predictable [Sav98]. Predicting [Gul04, PC08]. Prediction [AC07, CCM96b, CST11, CH11, DX11, FPR95, FM12, FGM21, HKM10, HLC07, HG07, JA07, JAC19, KMH10, KOO09, KT09, KB74, Lak00a, LT08a, Lev47, LSW08, MWLW09, MWC01, Mar80, MMS91, MSC00b, NR14, PBL^{+}17, Rez04, SMS08, Sha51, TWW^{+}10, VST11, WL08, XXZ11, ZA05, ZW12, AdlPN04, AH02, ABG^{+}94, CCM96a, CCMW05, FS98, FFW98, FWG98, FY95, Ghi06, HZZY02, IDKW99, VG95, Lee99, MR97, MSC99, PELA02, RA04a, TT94, TCOR05, Ten04, ZL00]. Prediction-Based [FM12, FGM21]. Prediction-Compensated [WL08]. Predictions [ZLRZ09, Vov06]. Predictive [AR08, AsM^{+}10, BS04, BKYK94, FU98, KdF12, KØZ08, KK08, MW97, NOG96, Poe83, RT87, Rob97, SL08, SAS10, TY08, WSDTO11, BK74, BdB06, Ger91, JA03, JS99a, KR01, KB74, KM95, MAC^{+}04, SA03, SCHB05, SN94, Swa04a, Swa04b, WKM^{+}98, WF03]. predictive-substitutional [Swa04a, Swa04b]. Predictively [HMR10, RRMG07, JR02]. Predictor [HF97, HF99, Log04, SKY10, SST^{+}12, Ghi03]. Predictors [ILR07, Log06, VS10, IF06]. Predykcyjno [Swa04b]. Predykcyjno-substytucyjna [Swa04b]. Preface [Ska98]. Prefetching [MPALSSBR10, CKV93, VK91, VK96]. Prefix [BY90, CD94, DB07, GKN09, Gir99, HAR10, HL90, LK96a, Nek07, OG12, Sal00b, Sto80, DS92, BCSV06, BK05, BBHK14, CK96, Dum06, GN01, IY00, LM03, MPL99, MLMD03, MM06, MTK95, MT96, OG16, DP91]. prefix-free [Dum06, GN01]. Prefixes [FMP83, Yok99a]. Preimage [Li10]. Preliminary [FK85a, LS94, LS95b]. Prepress [ISO98, NDPL00]. Preprocessing [HI12, KM98, PT10, Abe04, AT05, KL03, Mar02]. preprocessor [Chi98b, KYPY98]. Presence [CWRL11, LWW10, LLZ96]. Present [Suz12]. presented [ACM78, ACM79]. Preservation [LPG10, TRS03]. Preserve [BM11b]. Preserved [JR12, TS17, HL97]. Preserving [Fon81, FCDH91, IF07, JS99b, Prz98, SY18, Yok10, Ano06q, Ant97, CZ98b, DHW^{+}17, HB99, LOMSS05, SC01a, ZIL93]. preset [Mit01]. President [Dic92]. Press [Ma96, VPL98, SSZZ14]. Pretty [Zor88]. prevention [SR97a, Sto98]. Preview [NL80]. Price [Zei93, Zei94]. Primary [ITU91, YZLH16]. Prime [Cal06a]. Primer [Ano06r, Mot98, SDS95]. primitives [Rob91a]. Princeton [TCJ93]. Principal [AH14, DF08, Fow08, Fow09, TWC^{+}09]. principle [Li02a, RMG94]. Principles [GBL^{+}98, LL08, PMLA88, Poh85, Rea95, WSS00]. printability [Rub88]. Printed [Sha51]. Printer [TSS^{+}98, SST^{+}98]. printing [VW94].

Prior [GFE07, RVG11a, RVG11b]. priori [Hat95]. Priority [BZZ08, HH08, XXZ11]. Priority-First [HH08]. Privacy [KCCW17, Sal03, SY18, Kül12, WC88]. Privacy-Preserving [SY18]. Private [Buy02, Mot07, UPR99, Yok99b]. Prize [Ano08c, BMH08, Ano04a, Ano07a, Ano08a, Ano09b, Ano10, Ano11a, Ano12a]. proactive [BCR98]. Probabilistic [Cap59, Lan91, MKM^{+}08, Wol99c, CT99, LO99]. Probabilities [AL10, RVG11a, RVG11b, WB91, ASS97, Hat95, Rob91b]. Probability [FB16, JLJ08, NM10, OG12, PM88, Rez11a, Sad98c, WN10, Dav12, GN01, KB06, Lan10, YG03]. probably [VN90b]. probes [GSPR19]. Probing [AC09]. Problem [IP06, KS11a, LdV95, NZ07, WB91, ZK11, Zha11, BS06, CX04, CT95, CME06, FL13, Gra03, Moh02, PWM18, SA03, XG97a, YWPG03]. Problems [Blo98, De 00b, De 94a, GB92, GN01]. Procedure [ISO84, KKMS99]. Procedure-Frame [ISO84]. Procedures [ITU94, PW86, She90]. Proceedings [ACM95, ACM13, Ano06p, Ano11b, BBKF95, FLA^{+}03, IEE88b, IEE95b, IEE95c, SC00, SC04, SC05, SC06, SC07, SM09, USE93, WGM88, YR87, AFL96, BJZ94, IEE04, PBPT95, SC93, SC94, SC99, SC01c, SC02, SC03, SM10b, SM12, ZD81, ACM97, ACM03, Auc98, CLA12, IEE91, KRW^{+}93, LL08, NDN13a, NDN13b, SM01a, SC98, WS12, Yan10, BDLS96, SM08, vL94]. Process [CKS08, HHLW20, Liu97, SWW12, Bis94, Bis95, Du00, JSS04, ZKD05]. Processes [GH08, GW05, Liu91, SW04b, WV99a]. Processing [CMQW90, CMQW93, CMQW94, ETT15, GW92, G^{+}75, IEE95b, IEE95c, IEE04, IEE05, IEE11b, ISO84, JBG17, KB92, KKT10, Par87, Phi91, RS78, Sam90a, Sch86, She95, SHW^{+}01, SMB98, TSS99, TB11, VB14, Yua02, ZZS^{+}22, AG86, BA08, CSCW00, CS13, FC98, HHSS08, HPV09, H^{+}10, HPT11, Lop12, ML98, NCB15, Ng10, Nos99, PB99, PBPT95, Sab94, SS00, SRG^{+}21, Str99b, Tsa91, YZZ^{+}14, YYO04]. Processor [FC94, Lea78, MOT^{+}03, Isr91, KC91, WWA^{+}95]. Processors [AW04, CGC97, KGG^{+}14, BH10, Cra98b, Cra98c, HD88, IYY^{+}11, LJ05, MS05, PSSS03, RS04]. produce [Zir07]. Product [CC00, CGS93]. product-code [CGS93]. Production [GLM10]. Products [Kro00, Ano82]. Produktionsfaktor [GLM10]. Professional [Won85]. Profile [LST99, BSS05, JST01, KM16, LS95b, TY04]. profiles [ISO99b]. Profiling [BMM02]. Program [Ano07d, KKMS99, Knu86a, Knu86b, RS81, Luc00, OT03, DKV07]. Programmable [JGM^{+}20, SHW^{+}01]. programmed [CMP99]. Programmers [Kro00]. Programming [Fra67, Knuta, LV92, Sab94, SH03]. Programs [Cha66, Mah08b, Pet80, Wir76, CL16, ECM98, Fra06, OS01]. Progress [MR93b, DQ97, Lum13]. Progressive [BZ97, BAL02, BSS09, CVDL16, CC06a, DJCM09, FS02, FDU80, GD02, GVYZ97, HV94b, JTC^{+}12, JEK98, Kno80, LNA99, MS12, MT92, SWW12, SZ97, TGHL98, TFV05, VCP09, WWS11, BKV05, BF99c, CF99, FJ00, JF96, Mal99, Mal00, MDPM00, Tro96, TMM96, WF93, YJE04, ZAA00]. Progressive-to-Lossless [MS12]. Progressively [DVS^{+}14, PBC05]. Project [DVB06, FWI99, Kro00]. Projection [Fow08, Fow09]. Projections [EGF08, GFE07, YX12, CMW99]. projective [MDH06]. projects [Hop09]. Proof [Amj08, Bla09, FP19]. proofs [Li10]. Propagation [CC01, Xio11, KF95, SR97a, Sto98, VD18]. Properties [BYR06, EY81, KMSW18, OSI02, WST95, BPZ99, MW99, Rez05]. Property [HPST11, LLW11, Rud71, ZL11, DS92, LW11a, MM00]. property-based [LW11a]. proposal [Per89]. Proposed [SD06, Lit95]. Prospicient [EKT12]. protected [DWW02, HGH05]. Protection [DZ08, LZM10b, MR00, RRMG07, GLM^{+}03, Hal04, JEEL16, SX04]. Protein [NMW99, AN06]. Proteins [AC09]. proteomic [MKW^{+}06]. proto [Lit95]. Protocol [FGM^{+}99, Gai98, Kro00, Per89, Per90, PH90, Ran96, Sim92, Sim93a, Sim94b, Pow93, RS00, TG17, Zhu02, BLFF96, Sim94d, Woo96]. Protocols [SMC99]. Prototype [SW12b, BK97]. provable [FV20]. Provenance [BHLY20, WHB16, WB17, XMRF^{+}13]. provide [Kül12]. Provisions [ISO99a]. Proxy [MPALSSBR10, MPALSS^{+}12, VN08]. Proxy-Based [VN08]. pruned [MJB92, MMA99]. pruning [BZ97, LSC91]. Pseudo [CFY^{+}10, KA12, CMWM00]. pseudo-ARQ [CMWM00]. Pseudo-distance [KA12]. Pseudo-Random [CFY^{+}10]. pseudodifferential [DPS93]. pseudowavelets [PB99]. PSI14.K [Ric91]. Psychophysical [LLPP19]. pu [Lia84]. Publishers [Ano98, Hoc96]. Publishing [BBKF95]. Puerto [IEE91]. Pulse [Gra53, ITU89, ITU90, Moh04, Tan93]. pulsed [TC06a]. Pulses [Kra49]. Punctured [Cen09, Fen96b, ZGF02]. Pure [NZC09]. Purpose [AFF^{+}19, Hou79, Rob97, ATL^{+}88, CWZ99, CW01, CW02, CW15, LJ05, PM18]. pursuit [NZ95, SS06, TFV05]. Putting [Bar90a, WS93]. PVM [BDLS96, Uhl95, UH96]. PVQ [LBH91]. pyPcazip [SGA^{+}16]. Pyramid [ASH87, BA83, LV95, TPY95]. Pyramids [And12, LKA99].

Q [Lan91, Ano09d]. Q-Coder [Lan91]. QBF [PT10]. QccPack [Fow00]. QIC [VN90a]. QLFC [Ghi05]. QM [NPW97b]. QM-coder [NPW97b]. Quadratic [DØG08, BS06, Mod03, RMG05]. quadratic-Gaussian [RMG05]. quadratic-time [Mod03]. quadrisection [KNY96]. Quadtree [BF99a, FAEY95, HGS97, TN96, BZ97, Bet94, JM96, MB00a, Sam85]. Quadtree-guided [TN96]. quadtree-pruning [BZ97]. Quadtrees [AF88a, AF88b, Whi87, MO11]. qualitative [Ish89]. Quality [BS17, BM11b, BS94, GMSK09, H^{+}98, HHLW20, ITU91, JR12, LZM10b, MSE17, Mor12, NS01, Oh10, OLHL08, PC08, RL08, THHS93, WGZ^{+}18, Yan03, BHT19, Bry95, CCV04, CCV05, Che91, CPT^{+}20, Esk95, HGH05, KAB11, LL96, MSC00a, RK95, SMC99, TC06a, WF94, lGPL^{+}00]. Quality-Aware [BS17]. Quality-Constrained [HHLW20]. Quanta [HSP42]. Quantisation [Cie98, Prz98, FW95, YW04]. quantitative [Lyn69].

Quantization [Abo06, AR09a, AR12a, ASJ20, ABB10, BdB06, Bod95, BLS11, BB07, BBBP10, BBH93, CZZZ11, CM02, CS94a, CS94b, DFA^{+}10, Dav95b, DMS08, DMS10, DSG12, DZ08, FZ05, GG92, GVT95, GPO98, GG06, GV09, HLC07, JCF93, LT08a, LJ07, LBG80, Llo82, MWLW09, MHKK03, MHKK06, MSM09, MGV08, Moa86, MS95b, MDH06, ME02, ØZ07, PG72, PFAK10, PCD17, PR11, Ram85, Rea72, Rez11a, Ric10, SL08, SG12, SR09, SG11, TV09, Tao82, The89, Tra87, WWS11, WXC^{+}10, Wu95, WS08, YXPM11, ZKK09, ZW12, Abo95, AAS95, AECG93, ABMD90, AM93, BH93a, BW95, BF95a, BV96, Bis94, Bis95, BB93, BF98, BK92, CVC95, CCG96, CPG96, CF96, CL95, CS93, CGR91, CPP95, CR95b, Cve98, DS06, DRG94, DC97, DJL91, Del95, DY91, EMS03, ECG94, ENA05]. quantization [FZE03, FW91, FE99, FE01, FWA01, FA94, FA97, Fow97a, Fow97b, Fow98, Fow00, GCBV95, GCK^{+}96, GK94, GS96, GFH99, GZS^{+}96, GL01, GL03b, GL04, Gra05, GP95, GLL04, HMS92, HM96, HGS97, HW98, HW06, HM94, Isr91, JBF97, JT98, KGK00, KF95, KZ02, KFSM04, Lak98, LCB03, LLCC95, Lev96, LGO99, Li02a, LS92, LV92, Lin95a, LG99, LA91, LL96, LCLX04, LRO97, MJB92, MGG05, MSW96, MT92, Mis91, MS93, MMA99, MM99, MOK00, NOG96, NR95, NM91b, OG93, OR94, ØJH05, OG04, OG05, PGOO94, PN91, PM95, Prz00, RF05, RRG95a, RV94, RL95, RSC99, RS01, SW04a, SRKS95, SVS99, Ser00, SLR05, SHK99, Sou95, SS92, SO02, TPY95, TH03, TH05, TM91, Tse05, Vuc05, WHH^{+}99, WS00, Wan06, Wat93, WMH96, WZ94, Wu99a]. quantization [WD02, YP03, Yoo06, YM97, YG03, ZF92, ZF94, ZFE04, ZDR06, ZC91]. quantization-based [SLR05]. Quantized [CBHC08, DKG01, EEJ07, GKV99, LKA02, RVG11a, RVG11b, Tom06, Yor95, Cve99, CDL02, Dav95a, HM93, Mal06, WWW06]. Quantizer [ALMJR^{+}12, CKC09, Dum08, KMH10, KHCS08, KJ10, LLA08, SL10, CM97, CRR98, CN96, DWB04, DW05, FSC91, GK95, KR01, MOG05, MR93c, ME02, RMG94, SNR06, TFA93, WZ91a, Wu92, WF95, YM97]. Quantizers [AR09b, FZ02, SG09, AMS^{+}03, Bar94, BLO01, BF96, Car01, CGS93, DDGV06, DSV00a, DR06, EM02, EHAN05, FDZ00, FBX91, KMS96, LF95, LS93, LLZ96, MZ99, PRM97, RMZG03, TH04, Zee98]. Quantizing [Hoo78, Kra49, Max60]. Quantum [Ben10, VO08, HJW02]. Quasi [AALR12, BGBV09, GPV11, HB08, HV94a, HV93a]. Quasi-Arithmetic [HV94a, HV93a]. Quasi-cyclic [GPV11]. Quasi-distinct [AALR12]. Quasi-Monotone [HB08]. quasiarithmetic [GG04]. Quassi [KKJ11]. Quaternary [Liu07, The89]. Quebec [IEE04]. Queen [IEE04]. queriable [MPC03]. Queries [FHS08, FFFM13, LJPE21, YJGS00]. Query [XCK18, GC16, JLJ^{+}20, NCB15, Sno76]. Querying [JAL^{+}12]. Quest [Wei15, PWM18]. Question [Pou93]. Questions [Cal06a, LS95a]. Queueing [CKS08]. queues [FE06]. Queuing [MM95]. Qui [Pea90, Sie12, Sie90]. Quincunx [MMZ12, ZW07].

R [RHC87, LJF19, LM08, LCG11, PP10, WHH^{+}99, XD11]. R-D [LCG11, XD11]. R/D [WHH^{+}99]. R/D-optimized [WHH^{+}99]. RADARSAT [CW03]. Radial [LE02, LW14]. Radio [YZLH16, YK20]. Radiographic [Yua02]. radiography [Che91]. Radiology [Eln84]. RAID [Bar11, Bar12]. Railway [CFGL12]. RAM [CR07, DB03, MVJ17, RLG09, YDLC08]. Random [CFY^{+}10, CH09, DSG12, EGF08, FOhC09, FN09, Gag97, GFE07, Jac92, KCL06, KN10, LW99, RN10, ZYT^{+}15, Ara13, BKR^{+}93, Nat93, YJE04]. Randomized [AR12a, LW21]. Randomness [AC09, DTZZ12]. Range [Cam06, CHSS08, FHS08, LSH12, Sch86, SW12b, MEMS06, MGM13, NCB15, SYP04]. range-scanned [MGM13]. Range/Doppler [Sch86]. rank [LW21]. Ranking [Fen96a, Fen97]. ranks [Arn02, WM01]. Rapid [JAL^{+}12, NS01, Wit97]. rapidly [SH94b]. Raptor [FVP08, XSX05]. RAR [Ros06, TFD06]. rare [ABL99]. Rarissimo [Ano06q]. RARLAB [Ros06]. Raster [Col87, JB74]. RATCHET [SJ94]. Rate [Ach76, ALM11, ALMJR^{+}12, BFT11, BH07, BRALSSMP08, Ber71, BLS11, Cen09, CG03, CBHC08, CWY^{+}10, CDW11, CKS08, DRR10, DØG08, DR83, DC03, DSG12, EJ07, FSL^{+}12, GK99, GH08, GO96, HSX02, HB09, ITU91, KTMS10, KKA02, LZR^{+}10, LZWL11b, LZX^{+}09, LT08b, LBH91, LG01a, Lin14, LMH^{+}09, LWZI12, MTA^{+}20, Mar80, MYRD04, MN08c, Moa98, Moh04, OBGB11, RRMG07, Ric11, SL08, Sch70, Sho08, SjWL08, TV09, TDL^{+}19, TMD08, VB14, Vas07, Ven07, WWY^{+}07, WJDZ10, WC02a, WV99a, WV00, WXC^{+}10, WM94, WFLZ09, ZKØ10, ZK11, Zha11, ZWZ11, ZX11, ZL78, dRV12, Adj06, ALSSMPBR06, BH93a, BK94, BF96, BF99c, CCA02, CG04, CB04, CL06, CB97, CAC96, CF97, CD00, CDL07, DRG94, DW03, DW05, DR06, EM02, ENA05, EGNA06, FC98, FA93, GD06, GW05, GL03b]. rate [GL04, Gra05, HGS97, HLV97, Hoa99, HN02, JOC97, KA98, KUCV04, KCS93, KH04, Li02c, LZZ98, LRH^{+}21, LS96, MB00a, MW06, MM97, MF98b, Mil95, MRG99, MMA99, MR06, Ort96, PBC05, RMG05, RS00, RA03, RCH04, Ruf94, SRP04, TFV05, TEGM03, WM05, WBM03, YA13, YH00, YM97, Yu04, YK20, ZAA00, ZDR06, ZKD05, lGPL^{+}00]. Rate-Adaptive [FSL^{+}12]. rate-allocation [HN02, PBC05]. Rate-based [HSX02]. Rate-Compatible [Cen09, KTMS10]. Rate-Controlled [MTA^{+}20]. Rate-Distortion [ALM11, BFT11, BRALSSMP08, CKS08, DØG08, DSG12, GH08, LZR^{+}10, LZWL11b, SjWL08, TDL^{+}19, TMD08, WWY^{+}07, WJDZ10, WFLZ09, dRV12, CG03, DC03, GK99, GO96, LG01a, MYRD04, VB14, WV99a, WM94, CCA02, CG04, MW06, RMG05, RCH04]. Rateless [BC10]. Rates [LSC10, dCDV07, EANG04, HV06a, HLV96, LLZ93, NZ95, PR00, SC01a]. Ratio [BM11b, Sad98c, WLL^{+}20]. Rational [RKL16, BL97]. Ratios [DFA^{+}10, MD95, KM98, KM99b, Lit95, Zir07]. Raw [ZWZW12, GLI16]. Ray [HO19, SLM^{+}17, KA95, LKRR91]. ray-casting [KA95]. Raymond [Lau09]. RCPC [Adj06]. RD [RL95, RL96, WS00]. RD-OPT [RL95, RL96]. RD-optimization [WS00]. RDF [BHLY20]. Re [NR08, JYHC03, LNV97]. re-normalization [JYHC03]. Re-pair [NR08]. re-representation [LNV97].

Reaching [Liu07]. Read [MVJ17, YG02, Mis11]. Read-Disturbance [MVJ17]. read/write [YG02]. Reading [DCP17, Lia95, Phi91]. readings [CS13]. Real [EKT12, GKPS05, GCK^{+}96, Hoa99, JB95, KL03, KMP05, LYC16, MJB92, MMB07, PK05, SYZO15, SHW^{+}01, Tao82, Taw93, Weg07, Wei07, Ano04b, Bal97, FSC91, FA94, GLM^{+}21, KK08, LCJP21, Mer91, Mil95, MSC00a, SJ94]. Real-Time [EKT12, LYC16, MMB07, SYZO15, SHW^{+}01, Tao82, GKPS05, GCK^{+}96, Hoa99, JB95, KL03, KMP05, Taw93, Weg07, FSC91, FA94, GLM^{+}21, KK08, LCJP21, Mer91, Mil95, MSC00a, SJ94]. Reality [EFPS16]. Realization [THG03, BSF16, MW92]. reappear [Per21]. Reason [Cha06, Cha07]. Reasoning [Wol99c]. Recasting [Gra03]. receiver [CMWM00]. receiver-driven [CMWM00]. Receivers [ZKØ10]. Recipes [PF^{+}88]. Reciprocal [SM96]. Recognition [dABdSFNA08, DMS10, Gad91, GHD^{+}14, Li86, SC11, Yan03, Bon95, DHH04, Hua91, R^{+}06, R^{+}07]. Recommendation [ITU06b, GVSSBRAL08, HKM10, ITU02, ITU89, ITU90, ITU94]. recomposition [JM96]. Recompression [Jez15, Jez16, Kle97]. Reconfigurable [EK16, LMSE18, NMN99, PP10, IYY^{+}11]. reconnaissance [vB97]. Reconstruction [BT93, Fu03, FJD11, HSL^{+}20, MF11, TPM20, WCY^{+}10, WK92, ZCW^{+}11, Cve99, JF96, JM03, MR97, PW03a, ZWS92]. Record [Abe04, ACK08, Ros94, ACM78, ACM79, IEE88a, IEE95a, M^{+}98]. Record-Oriented [Ros94]. recording [Ham95a]. Recovery [FWZ08, GFE07, KWC^{+}16, KE11, MG08, Mos98, WZW09, ZZZ^{+}12, De 94b, Gul02, NTF06, QMCX19, Ste91, TO04]. Rectangular [MD95]. Recurrence [GL03a, You98b]. recurrent [CC00, Har96]. recursion [BS98, LHY12]. Recursive [CC19, HY10, Jog73, Mah08a, TIY97, Yam00, CCG96, KSY95]. Recursively [AAS95]. Recycling [Bar00b, DB07, DB08]. Red [Bar00a]. Reduce [MW10, Pom16, AN10, Du00]. Reduced [Che15, DS06, KFN99, KW95, MAC^{+}04, MSM09, ND04, RN10, SO02, UUiN12, CO98, Cre97, HD88, KRSS21, RCMS04]. Reduced-search [KW95]. Reducing [BKS^{+}18, DLY^{+}98]. Reduction [FCL13, Fou85, FK09, ITU91, Ind91, IF07, KHCS08, KH08, LSC10, SH94a, Tad20, TLR85, WMNP12, ZWZ11, BCP04, BKR97b, Kor67, LW21, LC91, NT05b, SS99, TK07, Wei90]. redundancies [VDPL00]. Redundancy [ASK12, BF11, Huf52, Kar61, MSFZ08, Rez08, Rud71, Sav97, Sav09, SMFZ07, SR16, DHS02, KF97, KR97, KY99, Kor67, LM02, LS96, MZ95, MF94, MPL99, MT96, RS00, RA03, Stu94b]. Redundancy-aware [ASK12]. Redundant [Liu08b, ABL99, Cve99, Fow06, YR02]. redundant-wavelet [Fow06]. Reed [KK07, RP88]. reel [BKV89]. reel-to-reel [BKV89]. Refactoring [Wic03]. Reference [Kro00, Mor12, Sal98, Sal00a, Sal04, Sal07a, MPE00, ZL00]. References [Ban09, OR99, WL15]. Referential [WL13]. refillable [Apo05]. Refinable [DZ08, SR09, BF96, JBF97, JT98]. refine [FS03]. refined [DVS^{+}14]. Refinement [JFC12, MMZ12, RJL11, SWW12, VCP09, CKW93, CX04, Cho96, Cre98, VKG01]. Refining [Haf96]. Refresh [ZLT^{+}18]. REGAUT [ISO99b]. Regenerating [GPV11]. Region [LMO97, SSP21, WJDZ10, ZKØ10, ZK11, Zha11, ZHX^{+}19, CSCW00, FDKB11, LMC05, LW11a, LZZ98, SMC99, SRP04, SM95b, Tur95, VKG01, YR96]. Region-Based [ZHX^{+}19, LMO97, CSCW00, YR96]. region-of-interest [FDKB11, SMC99]. Regions [HV08, MR00, PST97, SYP04]. Registers [Rub79a]. Registration [ISO99b, Tad20, ISO99a]. Regression [LZX^{+}11, LZZ^{+}12, SWWW15, NOG96, TKR03]. Regressive [ZXM^{+}10]. Regular [Cha89, Nav01, BLO01, BFG09, Szy02]. Regularization [SYO15, CA93]. Regularized [CXZD12, HB99, PX12]. Rehabilitation [Fou85]. Reinforcement [II02]. related [KNY96, Rez05]. Relation [KFSM04, Rez08, LPS02]. Relations [BdBN12, GL03a, LPS02, You98b]. Relationship [BW94a]. relationships [NO95]. Relative [BCC^{+}18, CD09, HPZ11, SG11, ZLX^{+}20, DJSS14]. Relative-Error-Bounded [ZLX^{+}20]. relaxation [GSPR19]. relaxing [De 05]. relaxivity [GSPR19]. Relay [LTO11, Sha11, TOL11]. Relay-Assisted [Sha11]. Relevance [LLZ09]. Reliability [LLLZ12]. Reliable [BRV08, Sch70, CDC96, CL16, KT02]. relighting [WLZW04]. remapping [Arn97]. Remark [NK16]. Remote [LMBN08, CM99, CCD^{+}19, HIMM20, uR96, YSXZ04]. removed [AECG93]. removing [MZ91]. Remplit [Pea90, Sie12, Sie90]. Render2MPEG [HKMS08]. Rendered [LMBN08]. Rendering [BGU96, GMC^{+}06, HB96a, HKMS08, LW11b, LGK97, LMH^{+}09, RK09b, SBE16, WDR11, CR97, LCY07, Mil95, PD16, ZL00, Z^{+}13]. Renormalisation [Sch98]. Reordering [LLA08, LKG12, Sch11, BB02, GFC^{+}09, Vuc06]. Reorganization [ZCJ^{+}20, Wei90]. RePair [LMM11]. Repetition [Wil89, Yok91]. Repetitive [NP16, RDDD96]. Replay [Amj08]. replenishment [FA97]. Replicant [KP03]. replication [SHWH12]. Report [FK85a, DQ97]. Reports [FWI99]. repository [Ano06h]. Representation [Cap89, Cot85, EFPS16, Eva79, LB90, Mal89, MMZ12, MSM09, NDPL00, SLM^{+}17, Xio11, ZWZ^{+}12, ZX11, Zec72, DE92, GO95, LNV97, PX12, PW03b, TBMM98, VL97, WSK^{+}11, YR96, Zec72]. Representations [AF87, BCG^{+}14, BB07, Eli75, YM08, FMMN07, Tou06]. representative [YCMR18]. Representatives [MPGMD19]. Represented [PK10]. Reproduction [SM97]. Reprogrammable [FC94]. repulsion [FW93]. requantization [CDMW02]. requirement [AN10]. Requirements [ITU91, Du00]. Research [BMNM^{+}93, Bi08, Che05, Fu03, Hua01, IBM88, Jak88, KRW^{+}93, LY91, Miu07, Wan06, Wan08, Xu06, Yan03, Zha99, Kro00, YR87]. researchers [PWM18]. Reserved [ZWZW12]. Residual [HKM10, MKH10, MF11, Rez04, Tra87, BH93a, Bar94, CM97, EGNA06, FBX91, Ghi06, GFH99, LG01b, LMS94, Nov98]. residuals [FU98, Nov00]. Residue [CA97, TY08]. Residue-Free [TY08]. resilience [ARZG03, Chi99, KB03]. resiliency [You97]. Resilient [LZM10b, SR97b, BM03, CP03, MYRD04, Spa00, TC06b, WLS06, XWDY06]. Resistant [WHZ12]. Resizing [WHW10]. Resolution [CXYQ12, LZX^{+}11, MGV08, RMB07, Sal91, Wic92, ZW08, BPZ99, Bel11, BBBP10, Che91, Cve98, DLT16, Eff98, GE06, LZZ98, OW03b, PX12, QT03, SMC99, SCKS97, Str99b, VFE00, WSA94, WD02, ZDR06]. Resonance [CRA10]. resource [CC03, MS02, hyd06]. Resources [And07, HAR10, Hop09, HJW02]. respect [BDS95, Sno76]. Responses [KN11, KN90]. Restoration [FWZ08, JZL^{+}12, BB95, DSSR95a, WHZ12, WC02b]. Restore [ZYF^{+}20]. restoring [ELZC84, WCH03]. Restricted [LA91, FLMPR06, GN05, KN05, LM01, LM02]. result [Mar02]. Results [GL04, Nel91, FL13, Ram91]. resurrecting [CBC00]. Resynchronization [FJ07, MW99]. Resynchronizing [HCL99]. Retina [Ost35]. Retrieval [BCC10, DS08, FO10, FCDH91, JEK98, KBD89, KRW^{+}93, LSH12, MZ96, NPV15, RR08, YR87, ZdMNBY00, ADCU08, BMNM^{+}93, CW01, DS06, Goo03, HPZ11, JG05, MZ92, OS97, VL97, VL98, WBN91, ZTSM05]. revealed [GSPR19]. Reverse [Yok99a]. Reversed [Köp21]. Reversible [BSS09, CCKH21, LV02, Mal07, MCRKC09, SDJ04b, TWM95, TW01a, TW01b, WZMG07, WV98, YKLK99, Yok10, ZASB95, AK99, BG03a, BH00, BZM98, CR91, GZS^{+}96, LKH^{+}00, Ram91]. Reversible-KLT [BSS09]. Reversing [BEQ98]. Review [Ano98, Hau12b, Nas13, Neu10, Sof05, Weg92, YTY12, FL13, Ram92]. Reviews [Ano00, Hoc96, Ma96, Hoc95]. Revision [CC88a, CC88b, ITU94]. Revisited [MNW98, BG06, Eff98, MNW95, Reg90]. Revisiting [SGD05]. RFC [BLFF96, Deu96a, Deu96b, DG96, FGM^{+}99, KMMV02, Per89, Per90, PH90, Ran96, Sim92, Sim93b, Sim93a, Sim94c, Sim94a, Sim94b, Sim97, Woo96]. RFID [FFFM13, TG17]. RFID-data [FFFM13]. RGB [BIP01, CLD04]. Rice [Kie04, Mal06, Sai05b, Whi06]. rich [CG04]. Rico [IEE91]. ridge [NOG96]. rift [Kes91]. Right [NBK16, Bry95]. Rigid [EEH17, HTS^{+}18]. Rigidity [HTS^{+}18]. Rigorous [Zei93, Zei94]. RIPEMD [SW12a]. RIPEMD-128 [SW12a]. RIPEMD-128/-160 [SW12a]. risk [PGOO94, WG94]. RL [NT05b]. RL-Huffman [NT05b]. RLS [DKS21, SPS98, SPS99]. RMS [Lev47]. RNN [TYD^{+}20]. RNN-based [TYD^{+}20]. Road [HSZ17, SY18, LCJP21, SSZZ14]. Roaring [LSYKK16]. robotic [Bli91]. Robust [AF87, BBF02, CM99, Cap89, CF96, CRA10, Cre96, FA93, FK85b, FK96, GKSB17, HDK^{+}12, KR01, LHS05, Liu08b, PSR04, PS05, RH97a, RC98, SP11, SDFH00, SA03, SCV11, WZMG07, WV98, XSW07, YLZ11, ZRR00, CF98, CRR98, CF97, HCL99, KN90, LKH^{+}00, MB02, Nov98, PBC05, SNR06, WMH96, YKLK99]. robustness [PK05]. Rods [Ost35]. ROI [BRALSSMP08, LZM^{+}10a, TVLS08]. ROI-Based [LZM^{+}10a]. role [HS06]. ROM [KBD89]. root [Nat95]. Rooting [WPA08]. rotated [WS06]. Rotation [DMS08]. Rotation-Invariant [DMS08].

Rotations [Giv58, BK97, HM94]. round [CWW20]. Rounds [SW12a]. route [Pou93]. Routing [ZWZ^{+}18, CG03, PKG08]. rows [LKG12]. Royalties [Rod95]. RPWS [WWA^{+}95]. RSHDLC [Fan66]. RTF [Kro00]. rules [Att95, MSOY96]. Run [Cap59, DLY^{+}98, Gha87, GJ20, GK88a, GK88b, Gol66, Gol72, Lan83b, PP18, Wil70, ARH03, CC03, CM04, GFV97, Mal06, OHRS21, Tsa98, WW92, YM97]. Run-Length [Gha87, GK88a, GK88b, Gol66, Gol72, PP18, Wil70, Cap59, CC03, CM04, Mal06, WW92, YM97]. run-length/Golomb [Mal06]. Runlength [KU95]. running [LNhCW16]. Runtime [CNPC98, CNCC98, LE94, GFC^{+}09]. Runtime-efficient [LE94]. RVLC [YKLK99].

s [ISO93, ITU90, Lio91, MOT^{+}03, BG96, LCG11, LK91]. S-Domain [LCG11]. S-Plus [BG96]. SA4 [RM21]. Safe [Hal94]. Salomon [Hau12b, Nas13]. Salus [Ano00]. same [BWB^{+}20]. Sample [MY16, War97, FS03, Rez05, RBR09, RBR10]. sample-based [Rez05]. samples [CDL02]. Samplify [Weg07]. Sampling [BT93, Hoo78, ZB12, GD06, LOS20, UHB09, WSK^{+}11, Wik03, ZF92, ZWZ11]. San [ACM78, ACM03, HWS06, HHG07, HHSS08, HHW05, HPV09, H^{+}10, HPT11, IEE91, R^{+}06, R^{+}07, Sch03, USE93]. Sandia [Bar00a]. Santa [SM01a]. Santiago [BJZ94]. SAR [RS99, Sch86, ZC99, ZWZW12]. SAT [AKS10]. SAT-Based [AKS10]. Satellite [Che77, Hua11, NT91, Uhl95, TLLB06, Tur75, lGPL^{+}00, HWS06, HHG07, HHSS08, HHW05, HPV09, H^{+}10, HPT11]. save [PWM18]. Saving [S^{+}88, VN90b, XDZ11]. Sayood [Hoc96]. SBHP [LP07]. SC2 [AS14b]. Scalability [CDL10, FSC^{+}11, RMB07, SS07, CAC97, GK99, KNTG08, RH97a, SMC99]. Scalable [Bar12, BBHK14, CCRCK09, CTC^{+}10, KXPZ12, LZX^{+}08, LP07, MZT12, MCRKC09, SMW07, SS07, Tau00, WHW10, ARR01, BM97, CKS95, CD95, Cre02, DWW02, DW03, Hal04, HZZY02, HN02, HGH05, IMB^{+}13, LAPL07, OM03, OTW^{+}99, PO00, RB01, VM98, WC02a, vK00, van02b]. Scalar [BLS11, BBH93, DSG12, Dum08, SL10, SG11, AMS^{+}03, BF96, DWB04, DW05, EM02, GS96, ILRS03, JBF97, ME02, RF05, SW04a, TH03, TH04, WMH96, WD02, ZFE04]. scalar-quantizers [AMS^{+}03]. Scalar/Vector [BLS11]. Scale [AVE18, BKS^{+}18, CDLW05, Kno80, TLR85, WYM10, ZCJ^{+}20, BKK20, EBH^{+}16, EBH^{+}18, EBH^{+}19, HP92, Li02c, LZZ^{+}12, Xue99b, Xue99a]. scaled [SPL20]. scales [KC06]. Scaling [EBH^{+}17, PC08, SDFH00, SC11, Won85, YDDB15, ZLC^{+}15, BSF16]. Scan [Col87, Pom16, SA93, NT05b, AFL96]. SCAN-95 [AFL96]. Scan-In [Pom16]. scanned [MGM13]. Scanning [ALM11, MO11, TCKM10, Ara13, GZ91]. Scattered [LF01]. Scenario [EK16, JRALMSS11]. Scenarios [AJN08]. Scene [GS12, MSC00a]. Scene-Aware [GS12]. Scenes [DKB^{+}16, WKC94, ZB12, GD06]. Schedule [LMSE18]. Scheduling [BKRSR09, LXG^{+}18, CG03]. Scheme [AKS10, AAL^{+}21, ABR10, BSTW86, BC10, CW91, DR83, GJ20, HDCH09, HHCH11, HSL^{+}20, IFK12, JMW09, JSC20, KY10, KS00b, LE94, Law77, Man97, MYS20, Moa98, Mof90, RD82, SA92, SF09, The89, Tis87, W^{+}00, XSW07, ZD95a, ZD95b, AB08, AS14b, BW98, BRM88, BM89, BEQ98, BHLY20, CWZ99, CW02, CF98, CC93, CKW93, CZAR20, CK95a, DDL98, ECM98, Gu05, GYLC09, HB99, JS95, JST01, KLJ12, KK03, KH04, LGM21, LCL03, LCY07, LS95b, LLWC08, MPL02, MSR04, MS05, NL03, NC97, PM98, PSSS03, RDDD96, Ruf94, RHC87, SM11a, Suz96, Swa04a, WBM03, YA13]. Schemes [AH14, ECBL12, FHS08, Ing08, Kno80, LZR^{+}10, MY16, ZMVR14, ALRT16, BDV00, BKR97b, HSP^{+}13, KF95, Kle89, Mäk89, PELA02, Szp91, TKYS16, Tur95]. Scheming [Mul96]. Schindler [NZ06]. Sci [De 00b, RBR10, Sie90]. Science [IEE84, IEE91, IEE96, WGM88, Aro77, Dav12, KCS93, PWM18, TD91, vL94]. sciences [KT13]. Scientific [AFL96, Bas85, PF^{+}88, UHB09, ZLX^{+}20, CDL^{+}19, RKB06, TDG^{+}19]. Scorecard [Bry95]. scorecards [Bry95]. scrambling [Kül12]. Screen [WGZ^{+}18, Win06, CBR14]. SDD [ZGS02]. SDK [Pav06]. SE [Pop11]. SE-Compression [Pop11]. sealing [KS91]. Search [AHK97, BBH11, BCC10, CCKH21, CNS12, HC06, KS00b, LSH12, LST99, SI99, Wol99a, Wol99c, YDDB15, ACM^{+}15, Bar94, BPMA02, Car00, GCK^{+}96, JS95, KM16, KFN99, KW95, LAPL07, LS92, Lin97, LW11a, LA96, LW04, MJB92, MR93c, RCMS04, Sad99, Sau95, Tou06, YG03]. Search-Order [CCKH21]. Searchable [BFL^{+}10, PW03b]. Searches [CW07, MM93, BAGCGC^{+}06]. Searching [BK93a, BPMA02, BY76, CFG12, CKC09, CHSS08, KS02, Man97, Nav01, SB97a, BR04, FGG^{+}08, LT03, SNZBY00, WL15]. sec [PS93]. Second [Auc98, Bru69, Deo02, Gol92, M^{+}98, NDN13a, NDN13b, Ben96]. Secondary [Sch86]. Secrecy [Boy91]. Secret [Ano06j, CC06b, Ara13]. Secrets [TSFS21]. Secure [ASLB20, WHB16, CPD^{+}15, SK04, Ano02, Ano12b]. Securing [PMZ13]. Security [Per03a, Sal03, WS12, Yan10, BH92, Wan06, XWC^{+}17]. SEED [HN07]. Seeks [Rod95]. Segment [SKY10, Chi98b]. segment-based [CAC96]. Segment-Parallel [SKY10]. Segmentation [LDZ^{+}20, WF93, ANT95a, HBL^{+}17, KL03, ME02, TPAK96, WY91, YWMS08]. Segmentation-based [WF93, TPAK96]. Segmenting [GS05]. Seidler [Ano98]. Seismic [DBTNT00, VED96, Kes91, KP95a]. Seismographic [HN07]. Select [TDS^{+}21]. Selecting [Kie04, NBK16]. Selection [AA71, DUH^{+}19, FWS11, MPGMD19, TDL^{+}19, TLR85, VRL09, Wu95, ZLRZ09, Adj06, BS95, CWYW91, Cov96, DDL98, KZ02, VBL94, YH00, Yoo06]. Selective [DJCM09, KKN07, MVJ17, RR08, SCKS97, ZWZ^{+}18, GZS^{+}96, LCT03]. selector [Bun97c]. Self [BR10, Dav95b, Dav95a, FR84, HL17, KF94, MN08a, MZ96, OSMY95, WHZ12, BK94, CHW04, DJSS14, Goo91, Ham95b, MSOY96, RKSM03, RSK88, SW04b]. Self-Adaptation [HL17]. self-index [DJSS14]. Self-Indexing [MN08a, MZ96]. self-organization [BK94]. Self-organized [OSMY95, MSOY96]. self-organizing [Goo91, Ham95b, RKSM03]. Self-Quantization [Dav95b]. Self-quantized [Dav95a]. Self-restoration [WHZ12]. self-similar [SW04b]. Self-similarity [KF94]. Self-Synchronizing [FR84]. self-test [RSK88]. Self-Tuning [BR10]. Semantic [MDPM00, PB91, WBM^{+}94, BGR01a, BGR01b, HBL^{+}17, MTTH13, CMK^{+}16]. Semi [WHZ12, Zei93, Zei94, ZGS02, CCMW03, NMWO96, Tur96]. semi-adaptive [CCMW03]. semi-compressed [Tur96]. Semi-discrete [ZGS02]. Semi-fragile [WHZ12]. Semi-Rigorous [Zei93, Zei94]. semi-structured [NMWO96]. Semidefinite [OS01]. semidirect [MBWP03]. Semilossless [KK04]. sensed [uR96]. Sensing [CXZD12, EC16, FLZW10, KNB10, LWW10, LSW09, LLW11, LKR12, MF10, MF11, PGW^{+}17, SF11, TF11, WWS11, WCY^{+}10, WFLZ09, WZW09, XJ14, YX12, YZLH16, ZCW^{+}11, ZZZ^{+}12, ZYT^{+}15, dMGdC10, CM99, CCD^{+}19, LGM21]. Sensitive [DY12, XWC^{+}17]. Sensor [AH14, CGSS15, DY12, EC16, KR10, LL16, Log06, MW10, MKK12, MCL12, RLG09, SF11, Sar12, SSRM09, WHB16, WB17, XJ14, ZYT^{+}15, ZMVR14, CZZR19, CS13, GD06, Hes08, HM07, KLMXL05, KLJ12, LGLK05, LLWC08, OG06, PKG08, PR00, RBD13, SF05, SKLA12, Tsa91, YA13]. sensors [DVS^{+}14]. Separable [SCV11]. Separate [YH00]. Separation [AB10, EK16, SMRR04]. September [AFL96, Ano88a, Ano88b, Ano88c, BJZ94, BBKF95, CLA12, FLA^{+}03, IEE05, IEE11b, ZD81]. Sequence [AC07, AC11, BC08, BCG^{+}14, CDAM07, CWLS15, Dav66, GWT10, Haf01, HDCH09, LN17, LHD^{+}20, LT09, Lyn66a, Wei06, FWA01, JLS04, ND04, TKR03, YR96]. Sequences [AOAAH17, BZ08, CW07, Cha66, CWLS15, CLS19, DØJJ09, FOP12, GF10, HMR10, HB08, KT07b, Lan83a, LZ76, LDZ^{+}20, Mof00, MA06, Mun91, PCD17, Slo06, WL13, ZL78, AN06, ASM96, AL00, DJSS14, ENA03, FMMN07, GT93, GLL04, HJW05, IF06, JR02, KT07a, LY97, MN08b, MF97a, MF04, QLQ11, RF00, Ram91, RDDD96, WSS94, WZL91, WJBM05, WWW06, XS05]. Sequencing [BS02, Car20, Sal18, VPS17]. Sequential [CDL13, Jen99, KØJ^{+}12, NM96b, Sha04, TH04, WWSY07, ZL77, Abe99, DXS08, GLI16, MR93c, Mod03]. SEQUITUR [MHM^{+}01, EL03, LS04]. Serial [Gün96, LXG03]. Serie [Sie90]. Series [BMG18, Dur60, UUiN12, Ano08e, CM05, EEKB15]. Series-Parallel [UUiN12]. Server [AK08, MPALSSBR10, FSY01, LN02]. server-based [LN02]. servers [NY98, Nat95]. Service [THHS93, Cra96]. Services [Ano06c, LK08, ABA14, CPD^{+}15, HBSASA19]. session [BP00]. Set [HLL^{+}20, LZ07, SP96, Sch11, GCL06, KF97, KP97, Kir05, MS05, TCP03, VL91, OVCS^{+}09]. Sets [Col86, CGC11, Eli75, FVW10, Fry85, Mof00, MFD^{+}09, Rez11b, BK97, BK91, CDL^{+}19, CMW99, Ind91, MGM13, Mis11, NL06, PB98a, Rub88, VG06, YJGS00, ZL95]. seven [Bry95]. Seventh [LL08, ACM95, ZD81]. Several [LE01, Bas91, DvOL97]. SHA [Bro08].

SHA-1 [Bro08]. Shadow [SBE16]. Shannon [Tja05, AN08, Gag04, HOR05, HR06b, THG03, Ver98, Wik03]. Shape [DMS10, JGL02, JEK98, LLZ09, LZR^{+}10, LZWL11a, LZWL11b, LPG10, MSR04, BAGCGC^{+}06, HGS97, MOK00]. shape-adaptive [BAGCGC^{+}06]. shape-gain [HGS97]. Shaped [KØZ08, KA99]. shapes [AF04]. Shaping [AsM^{+}10, sKJ01]. Shared [GF10, RR08, Wak06]. shared-memory [Wak06]. sharing [CBR14, CME06, SM11a, ZL95]. Sharper [Nat94]. Shearlet [TS17]. shelf [GLM^{+}21]. Shell [Hor06, NM91b]. Shells [TV09]. Shift [BBBP10, KTSA99, AGMP13]. Shift-And [KTSA99]. Shifted [QLH12]. Shifting [RBR09, RBR10]. Shirts [Bar00a]. Short [ENA05, KTMS10, ENA03, RGF06, Mis11]. Short-block [ENA05]. SHORTEN [Rob94a]. Shortest [IP06]. SHRDLU [Fan66]. Shrinkage [DGG10, DGG09]. SHS [Ano12b]. SIAM [ACM97]. SICLIC [BF99b]. Side [BS06, BRD10, GE07, JFC12, NM10, SL10, Sha11, SWW12, AG02, CPR03, JV04, MWZ04, PR01, Ser00, WO01, ZE03]. side-information [JV04]. Sided [MC13, Kri02]. SIGACT [LL08]. SIGART [LL08]. SIGIR [KRW^{+}93, YR87]. Sigma [ØZ07]. SIGMOD [ACM03, LL08, SM01a]. SIGMOD-SIGACT-SIGART [LL08]. Sign [CWRL11, Fou85, DH00]. Signal [AH92, CGC97, CMQW90, CMQW93, CMQW94, GG92, IEE95b, IEE04, Jay97, JAC19, Mal89, SZ13, She95, Tod89, Tsa91, Bör80, BYJ20, ELZC84, GSPR19, Lop12, Nad83, Ng10, RA04b, TYD^{+}20, VT95]. Signals [BT93, CT09, EY81, EGF08, Gad91, GFE07, MG08, M^{+}98, NR14, RS78, Tra87, Tsu72, WCY^{+}10, ZCW^{+}11, BK94, BMA04, CZAR20, DHS^{+}01b, ELZC84, JF96, KT02, NACM08, Ng96b, vB98]. signature [BSS05, Rob91b]. Signatures [Fow08]. signaux [NACM08]. significance [CPP95, DC97]. Significant [PWS10]. Significantly [LY97, BWB^{+}20, Cra98b, Cra98c]. Signs [Bar00a]. Silhouette [CVK97, CKV97]. Silhouette-like [CVK97, CKV97]. SIMD [CD95, GCK^{+}96, LLCC95, LBK16, MT92, ZZL^{+}15]. SIMD-Based [ZZL^{+}15]. Similar [WL13, DJSS14, GN01, NKP^{+}16, SW04b]. Similarities [CWLS15, Abe99]. Similarity [AC11, CD08, CD10, DØJJ09, GD07, Yok97, KF94, SS04]. Simoncelli [Sie90]. Simple [BŁ01, CDAM07, CC06b, CDLW05, Cra98b, Cra98c, CIS08, CV97, Dis88, FN09, GPM02, GGM12, Ing08, Jez16, KR10, Kno80, MSV98, NBIT07, Rob94b, Rob94a, Ros92, Sta07, Yok91, YKO98, BF99b, CLD04, KY99, MR04, OG05, PSSS03, TFV05]. simpler [KLV07]. simplicial [GD02]. Simplification [DGGP05, LMJ^{+}21]. simply [TKYS16]. simply-typed [TKYS16]. simulated [BWS95]. Simulation [Bar00b, Bi08, MG10, VO08, BHT19, BBQ^{+}95, KS89, Pow93, Sch10, Sha94, SGA^{+}16, Wil02]. Simulations [DLY^{+}98, CCO^{+}19, EFF00b, FG97, GSPR19, LLW^{+}14, Pow93]. Simultaneous [WPA08]. Sine [BYR06]. Singer [IEE84]. Single [BPZ99, CXYQ12, CS94a, CD00, CDL07, MYS20, MRG99, NYJ09, CS95, MS03, RSC99, Tsa91]. single- [MS03]. Single-bit [CD00, CDL07]. Single-Block [MYS20]. single-chip [Tsa91]. Single-Pass [CS94a, MRG99, CS95, RSC99]. Single-resolution [BPZ99]. singular [SCC12, YWPG03]. Sinogram [CFC05]. Site [Kom96]. situ [JAL^{+}12]. sixteenth [KRW^{+}93]. Sixth [BBKF95]. Size [CS94b, CPL09, PC08, Wei90, Yok10, CS93, De 05, EHAN05, Ind91, LF95, LDK99, PELA02, RSV^{+}00, VN90b, YR02]. Skeletal [Hal11, TBMM98]. skeleton [KRSS21]. sketching [AGMP13]. skinning [MZP06]. skip [Aus00]. skip-innovation [Aus00]. skipping [ADCU08]. Slashing [KKP12]. SLDC [ECM01]. Sleator [RHC87]. Slepian [Cen09, CJDL09, Cm10, CLME04, CME06, FWS^{+}08, FSL^{+}12, KTMS10, LLN^{+}04, LTB04, LCLX04, SRP04, SLXG04, YQ05]. slicing [MYRD04]. Sliding [Dun80, FOF11, KKS00, Yu95, Ben96, Gra99]. Sliding-Block [Dun80]. Sliding-window [Yu95]. SlimCache [JSC20]. slip [Ano04b]. Slowly [EEJ07, SC01b]. SLP [INI^{+}16]. Small [IRB97, KVP20, LZ06, LF04, Mul87, dlCNLA^{+}95, XG97b]. smaller [LSYKK16]. Smallest [CD10]. Smart [ASLB20, Bar00a, MPALSSBR10, DQ97, EEKB15, KF94, XWC^{+}17]. SMCA [HBSASA19]. SMIL [Ten00]. Smooth [TPM20, CDR93, May99]. Smoothing [AsM^{+}10, HGH05, Lak97]. smoothness [DJL91]. SMPs [NU05]. SMPTE [Ano08i, Soc08a]. Snowbird [SR91b, SC92, SC93, SC94, SC95, SC96, SC97, SC98, SC99, SC00, SC01c, SC02, SC03, SC04, SC06, SC07, SM08, SM09, SM10b, SM11b, SM12]. SNR [Hal04, Mor12]. SNR-scalable [Hal04]. SNS [sKJ01]. SOAP [ABA14, ASK12, HBSASA19]. social [XWC^{+}17]. Society [Ano09c, Ma96, Pin99]. Soft [FWZ^{+}12, HH08, YP03, B^{+}13, GCS00, How96, KT02, LA05, LW04, PM99, WV99b]. Soft-Decision [HH08]. Soft-decoding [YP03]. soft-input [KT02, LW04]. soft-output [LW04]. Software [BGU96, BKK20, CMP99, FSY01, Kro00, LE94, NLU93, Pat94, Pav06, Swa08, Yu96a, Ano82, Ano06q, Ano08d, Ano09f, DK02, Fow00, HM87, HM91, KCB98, dlCNLA^{+}95, Yu95, Kro00]. Software-only [BGU96]. Solar [MFD^{+}09, PHG86]. SOLiD [MC13]. Solomon [KK07]. Solution [CCF^{+}18, FP20, PR98, DPS93, XG97a, YWPG03]. solutions [BBHK14]. solve [GB92, PWM18]. Solving [Blo98, JU10]. Some [AL98, ES00, KM97a, Ram91, Ric79, Ric91, BCSV06, CLME04, Cra98a, GL03a, Hal94, Mar02, Szp91]. Somme [Zec72]. Sort [Yok97, ZIL93]. Sort-Based [Yok97]. Sorted [Gra99, Yok99a, KS08, LBK16]. Sorting [AN08, BW94b, NZC09, OG12, SB97a, Yok96, AM97, ALA98, BB05, CT98a, IY00, Lar98, LOMSS05, NL03, PST05, Ric86, Sad98b, Sch97, Sew00, YZ07]. Sound [IRB97, Yor95, Den01]. sounder [GRG05, HAH05]. Source [Abr89, AJN08, ARR10, BH01, BF12, BFCP09, CT09, CJDL09, CDMW02, Cve99, DRR10, DØG08, ECBL12, Fre93, FVP08, GvV75, GL09, GE05, Haf01, HS07, HDCH09, HGFL09, HKW10, Ing08, IK11, KM11, KH08, KØZ08, Kro00, LZX^{+}09, Law77, Lee81, Lle87, MGL11, MWZ04, MN09, Moa86, Nus76, Pas76, Ram87, Ram85, RJL11, RRMG07, Rez08, Sar10, Sch72, SF09, Sha10, Sha11, SSRM09, TW92, Vas07, WFLZ09, XD11, Zha11, ZZPB09, Abr94, AdlF06, Ano06v, AJRK98, BDV00, Bar04, BH00, BHS06, BF99c, CF98, CF99, CS05, CK95a, CPR03, CT98b, CME06, DS98, DHS01a, Eff98, Eff99, ENA03, EANG04, FZE03, FR06, Fow00, GF01, Gin95, GKV99, GG04, GCS00, HR06a, HSX02, HOR05, JST01, KLMXL05, KT02, KB03, Lan91, Li02a, LSC91, LA05, LLZ93, LLZ96]. source [LZZ98, LS03, LC91, LNA99, MZ95, MF98b, MB02, MR06, NTF06, PM99, PSH00, PR99, PR00, QJRA99, RMZG03, RS02, Ruf95, SW06, SF05, SNR06, Say98, TCOR05, TW00, Ton93, TKRR02, Tun04, VG06, VSG00, VW98, WO01, WS05, WAL05, WBM03, WJBM05, WWW06, XSX05, XHS05, YH00, YSXZ04, YSXZ05, ZE00, ZE01, ZE03, ZFE04]. Source-Channel [ARR10, BFCP09, CT09, ECBL12, HS07, Ing08, SF09, BF99c, CF98, CF99, CS05, CT98b, FR06, GF01, GKV99, GG04, GCS00, HSX02, HOR05, LS03, LNA99, MZ95, PM99, QJRA99, Tun04, WAL05, WBM03, WWW06, XSX05]. Source-Coding [Zha11]. source-controlled [Ruf95]. source-splitting [CME06]. Source-Word [Fre93]. Source/Channel [BH01, ZZPB09, BH00, DS98, PSH00, WJBM05]. Source/Channel-Decoding [BH01, BH00]. Sources [BW10, BF11, BRD10, Bud76, CTD07, Cm10, EKT12, FSL^{+}12, FVP08, FLZW10, GE07, HS07, Ing08, JMW09, LTO11, MG10, MGL11, RR08, Rez07a, SR08, Sch70, SV09, Suz11, TOL11, Vol97, WWSY07, WWY^{+}07, AAS95, BST96, CCMW03, CS05, CM04, FZ05, FM95a, FBS04, GFV98, GF01, HL02, Kri02, LLN^{+}04, LTB04, LXG03, Mal06, MSW04, MF94, MA05, PA97, RMG05, RS98, RA03, RA04a, RA04c, Sav98, SRP04, SC01b, SV98, Suz96, TKR00, TGFZ03, TVW97, YQ05, ZF92, Zan92, Zee98, ZGF02]. Space [Bar00a, BD82, CPS08, Col87, EP65, EP67, FIKS18, HGFL09, JSS15, KVP20, Lyn66b, McC76, MTK95, NPV15, Nix81, PLF91, Sag94, SYO15, Sew01, VK17, VN90b, VPGB10, XDZ11, ZN08, ANT95a, ABF^{+}91, AN10, BEGV18, BS98, BJ18, CT98b, DH97, DH98, FP19, Hal94, Li02b, Li02c, Lu92, MDPM00, MPL99, MDH06, MOK00, NMW98, NSFO^{+}99, PHG86, SB06, TD91, Tin77, TR88, Woo94]. Space-Economical [McC76, MPL99]. Space-efficient [MTK95, BJ18]. Space-Filling [Col87, HGFL09, PLF91, Sag94, BS98, CT98b]. space-frequency [MOK00]. Space-Time [SYO15, Sew01]. Spaced [GMV17, Roy87]. spaces [ISO99b, KP03, RA04b]. spanning [EN84, IEE88a]. Spark [WZL^{+}17]. Sparse [BB07, CBK96, DRR10, FP20, FK85a, FLZW10, LTO11, LKR12, MG08, PX12, RK09b, Sal00b, WSK^{+}11, WCY^{+}10, WFLZ09, WDR11, ZCW^{+}11, ZX11, Aus00, BM98, CCMW03, LCL03, NEGF96, PW03b, Rub88, CW91, Has18]. Sparsity [HG07, ZZZ^{+}12, BJ18]. SPARTAN [BGR01a, BGR01b]. Spatial [CFC05, DMC^{+}09, FPR95, GFC^{+}09, HG07, LZPB11, LLWD14, LCG11, LZZ^{+}12, MS12, MNBC18, Ric10, Sam90a, Sam90b, SDFH00, SS07, TWW^{+}10, CSP05, GK99, sKJ01, NCB15, OM03, PKG08]. Spatial-Domain [SDFH00]. Spatial-spectral [LLWD14]. Spatial-Temporal [LCG11]. Spatially [CCRCK09, WHW10, MM00]. Spatio [EC09, LDZ^{+}20, YX12, PX12]. Spatio-Temporal [EC09, YX12, LDZ^{+}20, PX12]. spatiotemporal [CPT^{+}20, YZZ^{+}14]. Special [Bil20, CS13, Mof97, Ska98, Zhu02, BMT91, VKG01]. specialities [Jur88]. species [BSS05]. Specifiable [OLHL08]. Specific [HCBMSS12, ZBMM97, AR18, KA95, MS97, PM98, SO94, Zir07]. Specification [CC88a, CC88b, Deu96a, DG96, Fed93, Deu96b]. Specified [BM11b, GN01]. SPECK [DPdS08]. spectra [DRG94, LMS94, MM05]. Spectral [EY81, FWZ08, ILR07, MS12, RKL16, RH07, Smi79, GGG^{+}09, Ish89, LLWD14]. Spectral-Spatial [MS12]. Spectrometer [Sal91]. Spectrum [AG09, Ton90]. specular [KCL18]. specular-enhanced [KCL18]. speculative [JKL13, WZL^{+}17]. Speech [BKYK94, BLS11, dABdSFNA08, Deg94, Gad91, HL17, IEE95b, IEE04, Jay97, KMH10, KP95b, Li86, LLJ09, PG72, Par87, RS78, Tao82, Yu96c, CM97, DRG94, Hua91, KM95, LMV03, Mer91]. Speed [Ano03, BR09a, Dum08, HU99a, HU99b, LA12, MNM04, ABF^{+}91, BE95a, Bry95, CZS92, JB95, KCB98, LA96, MSOY96, Sta94, SR91a, VL91, Wel84a]. Speed-Up [Dum08, HU99a, HU99b]. Speeding [BE95b, FM95b, JYHC03, SKF^{+}00, Reg90]. Speeding-Up [FM95b]. Speedup [HBL^{+}17, VPS17]. Speedy [Bre04]. Spelling [Dun81, Mil81, Pet80, Pet81, RS81]. sphere [LW14]. spherical [WLZW04]. SPHFB [OVCS^{+}09]. SPIFF [ISO99a, ISO99b, Mur96]. Spiffy [lGAR03a]. SPIHT [IA00, KP97, LG01a, LG01b, TCP03, VFE00, ZC99]. SPIHT-NC [IA00]. spike [WV99a]. Splay [Jon88]. Spline [BDHJ04, LZWL11a, LZWL11b, LF01, PW86, BL97]. Split [Luc00, PB11, TGHL98, CS92]. split-merge [CS92]. Split-stream [Luc00]. splitting [CME06, HS97]. Spread [AG09, Ton90]. Spread-Spectrum [AG09]. Spring [USE90]. Springer [Hoc95, Neu10]. Springer-Verlag [Hoc95]. Sprintz [BMG18]. sprite [FWY^{+}98]. Square [PLF91, Smi09, ARR01]. Squared [PG72]. Squares [CXZD12, Llo82, PW86, ZXM^{+}10, LMV03]. Squeeze [Ano92a, Bar90a, Bus09]. Squint [Sch86]. squish [Ano06o]. SSD [KWC^{+}16]. SSIM [CBHC08, RK09a]. SSTV [Ano07c]. ST [NZ07]. Stability [FDH^{+}20]. Stable [ABM10, KF95]. Stack [ARH03, Tsa98]. Stack-run [ARH03]. stack-run-end [Tsa98]. Stage [HI12, Log04, WWSY07, Abe05, Abe07a, FN93, MS03, SR08, WBE^{+}94]. stages [Abe10]. Staggered [SL08, TH05]. stagnate [PWM18]. Staircase [ABB10]. Stamping [JS99b]. Standard [Ano06c, Ano09a, Ano12b, BBH93, BS94, BBOH96, FIP01, HESF11, HCBMSS12, IEE85, ISO93, ISO00, ISO03b, JPE00, Lio91, MPFL97, PM92, Rez04, SEC01, SD06, SMW07, Shl94, SHW^{+}01, SDM02, Tho92, Wal91a, W^{+}00, SM97, Ano06f, Ano08i, BCD^{+}95, EGCK98, Le 91, LR04, SV95, Wal91b, YSCK00, Ano02, L^{+}05]. Standardization [WSS00, ISO03a]. Standards [BG03b, Car97b, GBL^{+}98, HR80, LK08, RH96, RAE16, Soc08b, TM02, THHS93, Ano06h, ITU06a, Mai95]. Stanford [IEE95a]. stars [Ano18]. Start [Pig01b]. Start/Stop [Pig01b]. State [BHS06, CK95b, Ric12, SSP06, SGLT98, TR93, Bis94, BBZ14, BKR97b, Bun97c, Bun98, GK95, GS96, IF06, IC98, Lev96, MF04, TM98, YL99]. State-of-the-Art [SGLT98]. States [Pom16, SM01a, For99, Boo99, HPV09, H^{+}10, HPT11]. Static [MSZ94, NLW99, BDS95, FGGV04, LM01, MP85, NLW95, Oka05, RK15]. Station [Bar00a]. Stationary [CTD07, KS96b, MG10, MGL11, Uhl96, AAS95, SC01b, XS05, YQ05]. Statistical [Bas85, CDAM07, Che00, Eva79, FNP08, AS14b, Gin95, MSOY96, NYC05]. Statistically [CD92, LSD97]. Statistics [EP65, EP67, Lee81, PR11, Sal02a, Ven07, WGM88, B^{+}13, CM04, Mal06, Ram91]. STD [Sim94d]. stealing [CDMW02]. Steganography [AG09, Cac98, MAL10, AOT13, Sun16, Way09]. Step [Dum08, JLL^{+}10, Deo02, Shk02]. Stereo [VMFG07, ZJW^{+}12, BdB06, Ghi03, JM97, PHM^{+}14]. stereopairs [Per88]. Stereoscopic [ZWZ^{+}12, SW91]. stetige [Hil91]. Still [BAGCGC^{+}06, HLNS17, Mur96, PM92, RL08, Thy10, Wal91a, Wic92, BP98, ISO95, ISO97, ISO99b, KF94, RAFH92, Ruf94, Ruf95, Wal91b, van02a]. stimulus [MK06]. STOC [ACM13]. Stochastic [AMAM13, Lev96, SPS98, dCDV07, Ger91, SPS99]. stock [ABG^{+}94]. Stop [Pig01b]. stopper [RTT02]. Storage [Bar11, Bar12, BLC^{+}16, FLS20, GPV11, Hah74, ISO93, JSCM17, JCS^{+}17, KMM^{+}12, Les88, RR08, Spi93, SY18, Won85, XDZ11, XMRF^{+}13, ZWZ^{+}18, ZHX^{+}19, CDC96, DE92, Du00, Hal94, HPZ11, HD88, JLJ^{+}20, Mer91, PST97, PM18, RRG95a, Say92, Tin77, van02a]. Store [Nix81, KV19]. stored [MBWP03]. stored-integral [MBWP03]. Storing [AF88a, AF88b, JAL^{+}12, KBD89]. Storm [Dic92]. Story [Ano18, Wri39]. straightness [AF04]. strange.rar [Ano08g]. Strategies [Mat12, MPALSSBR10, CDGO02, IM01, Say92, SW97, THG03]. Strategy [CW97, DUH^{+}19, VPS17, BS06, LW11a, Tou06]. Stream [CCF^{+}18, CRN08, GKSB17, MCM^{+}17, MM03, Rub79a, YHM20, KMP05, Luc00, SHWH12, WB00, LKR17]. Stream-Aware [CRN08]. Stream-Based [MM03, YHM20]. stream-informed [SHWH12]. Streamable [Hal11]. streamed [SRG^{+}21]. Streaming [AK08, BW11, BS17, BRV08, CSM01, ECM01, LLT^{+}16, SQ16, ZCJ^{+}20, Aus99, BPT06, CCA02, CG03, CG04, COMF99, DHW^{+}17, HGH05, KH04, LCJP21, RCH04, WM05]. streamline [MPL02]. streamline-based [MPL02]. Streamlining [LH91]. Streams [Ged14, Gir99, HB09, QB09, SKY10, USY17, Ano06q, DWW02, DW03, LBS98, Pap99, Wu99b, YCMR18]. String [AF18, BK93a, Cha93, CW84, FT95, FT98, FV16, FT04, GHLN18, KS11a, LD10, LS94, MM93, OM07, RTT02, RPE81, RNOM09, Sad96b, YYZ12, Ant97, BFG09, CDDM05, FL13, FGG^{+}08, KS96b, Lar99, LOS20, LOMSS05, NKT^{+}01, NT05a, RSV^{+}00, Sad93, SS04, WL15]. String-to-Dictionary [KS11a]. Strings [AA71, AF87, Cap89, ER78, FT98, FK85a, Jak85, KM99a, LZ07, MJ75, Sal00b, SB97a, BM99, FT95, HC97, NKP^{+}16]. Strip [Ise01]. Stripification [DGGP05]. Strong [HLNS17, LOS20]. Stronger [ANS12, Dub11]. Structural [DØJJ09, Rub88, AdlPN04]. Structure [Bar12, Cot85, FLZW10, ISO84, LMM11, NM96b, Pes07, Suz12, Wan73, Yok99a, Bun95, CC93, EL03, Fow97a, KM16, SY01, VO91, Wei90, WKM^{+}98]. Structured [B^{+}92, DX11, FB16, GA08, HC06, KM11, WW12, ANdlF04, ANdlF07, Car01, CPP95, ECM98, KJP99, LMV03, LS92, LS93, Lin97, LM08, MZ99, MMA99, NMWO96, OG04, PGOO94, RMG94, SN96a, TIY97, WS00, WZ91a, WY91, WF94]. Structures [GF10, Lan84, Lar99, MMB07, Sam90a, Sam90b, SST^{+}12, Wir76, ZG02b, BW95, GHSV06, PRM97, Sam85, YJGS00]. STT [MVJ17]. Student [YTY12]. Studies [Wil70, BYJ20]. Studio [Kom96, TY04]. Study [Che99, Chi98a, GGG^{+}09, KL01, Li86, LG78, MMK98, NT02, PT10, SMOY97, Wan73, Wu96, Yua02, Cam79, Kle89, NL03, WLCZ15]. StuffIt [Ano09f]. style [IY00]. Sub [Bru69, Tan93, UHB09, Wic90, BR04, VG95, Moh02]. Sub-Band [Tan93, Wic90]. sub-cell [VG95]. sub-linear [Moh02]. sub-optimal [BR04]. Sub-Optimum [Bru69]. Sub-sampling [UHB09]. Subband [HMS92, JCF93, KP95a, Liu97, Rea95, Sie90, SA90, VK95, BH93a, BK94, BBQ^{+}95, Cve98, DKCS95, FK95, GMR95, GCBV95, Hsi01, JF96, RK95, SRKS95]. subband/wavelet [Hsi01]. Subbands [LLJ09, GFV97, GS96]. subdictionaries [NYOY94]. Subdivision [BDHJ04]. Subdivision-Surface [BDHJ04]. Subgraph [WW12]. Subject [PC08]. Subjective [IRB97, RL08, SGLT98, WF94, WZ94]. Sublinear [JSS15, MG08]. Suboptimal [LS94, Sim98]. Suboptimality [EFZ03, MS03]. Subsampling [XW10, ZJW^{+}12, Poy08]. Subsampling-Adaptive [XW10]. Subsequence [AC09]. Substitution [CDL13, NBIT07, NIB^{+}09, SS82, AL98, AL00, RK15]. substitutional [Swa04a, Swa04b]. Substring [BCC^{+}18, DB10, Dub11, KKFL14]. substrings [ABL99]. substytucyjna [Swa04b]. Subtensor [EGLS21]. subtractive [Kli00]. Subtrees [Dav95b, Dav95a]. subword [Cha87, Jak88]. subword-tree [Cha87]. Successive [Bar04, CX04, Cho96, Cre98, Liu97, RJL11, VKG01, CDMW02, MB00b]. Successively [DZ08, JT98, BF96, JBF97]. Suffix [ABM08, AN08, FOF09, FOF11, FH08, FIKS18, GMV17, GNF15, HPST11, Kid09, Köp21, MM93, McC76, NP16, Neu10, NZC09, RL99, SLZ08, ZN08, BB05, Bun97c, CK96, Lar96, NP00, NAIP03, NdR04, PST05, RK15, Sad98a, Sad99, Sen06]. suffix-tree [Bun97c]. Suffixes [FMP83]. suitable [OT03, PGS10]. sum [BW95, BF95a, Zec72]. Summarization [BS20, MTTH13, SPL20]. Summary [YJGS00]. summation [CGS93]. Sums [BCC^{+}18]. Super [CXYQ12, LZX^{+}11, PX12, VBK89]. super-blocking [VBK89]. Super-Resolution [CXYQ12, LZX^{+}11, PX12]. supercompressed [KPM16]. supercomputer [BB93, TCJ93]. supercomputers [BRM88]. Supercomputing [IEE88b]. Superposition [EJ07]. Superstring [IP06].

Supervised [AAL^{+}21]. Supply [HAM17]. Support [IW94b, Kro00, KA99, PZ12]. Supported [Dau88]. Supporting [FDKB11, FFFM13, KK03]. Suppression [BT93, Roe76]. Sure [OS90]. Surface [BDHJ04, CYH94, Cot85, Eva79, WDR11, Hil91, LLY00]. Surfaces [LPG10, Sal06, SLM^{+}17, KL97]. surgery [TR98]. surgical [PW03a]. Surveillance [CFGL12, KD18, Ano04b, SCKS97, XWC^{+}17]. Survey [BS20, DUH^{+}19, Fry85, LE01, MLDH15, MV16, SCH^{+}14, Ben10, FNK94, RBD13, SKLA12]. SVD [ABP09, BTLK18]. SVM [dABdSFNA08]. Swap [CBC00]. SWC [LCLX04]. SWC-NQ [LCLX04]. SWIFT [YCMR18]. Switching [Cha00, VNHB12, VW98, ZB12, AL81, LC03]. syllables [LZ06]. Symbol [Fen96a, Fen97, NM96a, YSCK00, MSWB93, PM98, WB00]. Symbolic [Sha10]. symbols [AdlF06, PB98a]. Symbolwise [BW94a]. Symmetric [BV96, TMD08, BBF02, Cra96, MR06, ØJH05, PR00, SV98, YSZ11]. symmetric/antisymmetric [YSZ11]. Symmetrical [TW01b]. symmetrization [RS02]. Symmetrized [Bel11]. Symposium [ACM78, ACM79, ACM95, ACM97, ACM13, AFL96, Ano03, IEE84, IEE91, IEE96, LL08, WGM88, IEE95a]. Synchronizable [CD92]. Synchronization [Bis08, BP09, Dub11, PSR04, Rud71, KN99]. Synchronizing [FR84, LK96a]. Synchronous [ITU94, Lit95, Mai95, MM95, SV95]. Syndrome [RG02, TR88]. syndromes [PR99]. Synergy [WZY^{+}11]. syntactically [ECM98]. Syntax [Fra97, RV94, HS06, HAY06]. Syntax-constrained [RV94]. Synthesis [CHMW10, KP95b, LLLZ12, PLF91, SMO98, Tun67, CMP99, HZ95, Sab94]. Synthetic [Lev95, Cha96, GB98, GVYZ97]. System [AG09, ASLB20, BZ08, Bi08, BLC^{+}16, B^{+}92, CNPC98, CNCC98, CY92, DSM07, Dos72, Doz91, EY81, Fal73, Hou79, HHLW20, Hua01, ISO00, KHHK21, LLT^{+}16, Liu08b, MNM04, Per03a, Pro86, Rai87, RBR10, Sha10, Tao82, Wan08, Yua02, ZYF^{+}20, APC92, BGM99, BGR01a, BGR01b, CC03, CPT^{+}20, Cor85, DP76, Ind91, IMB^{+}13, KYPY98, KA98, LHW00, MLS03, MP95, OW03a, PHG86, RK93b, Sav03, Sno76, SZM03, Tai95, Tei98, Ude93, VJB^{+}95, Wak06, ZTSM05, ZE00, ZLN06, Z^{+}13]. system-on-a-chip [CC03]. Systematic [EP65, EP67, KJ10, LWW10, RRMG07, XS05]. Systems [ALYK12, BLNK14, BMM02, BKS^{+}18, CEA18, CPL09, DRT81, De 11, DHN85, Fra97, HLL^{+}20, Hoo78, ISO84, IRB97, ITU91, Jog73, KL01, KHHK21, KBD89, KCBM21, KD18, LL08, LJ07, Lin78, M^{+}98, MMK98, MV16, OINC10, PG72, PZZ^{+}22, SQ16, TSS99, TH10, USE90, VPGB10, Wak07, WMNP12, XDZ11, YS10, ZdMNBY00, Ano98, BMNM^{+}93, BdB06, Bli91, CDC96, GMR95, GL03a, Haa10, Haa12, Har96, Hof97, How93, Hus70, ITU06a, KIHO95, KT13, KF96, LLCC95, LDK99, Lin95a, MZ92, ME02, NTF06, dlCNLA^{+}95, OTY^{+}97, OT03, PD07, PHG86, PR01, PMH95, QMCX19, TD91, TO04, Ton90, Tsa91, Vos06, WBN91, Yu96c, ZDR06, ZWS92, vB97, Kro00]. systems-software [dlCNLA^{+}95]. Systolic [ZRZ^{+}19, RH91, SH92]. Syzygy [KK07]. SZ [TDL^{+}19].

T [ITU06b, Gün96, HY10, HKM10, ITU02, ITU94, MKH10]. T-Codes [Gün96, HY10]. Table [ACK08, CW91, SM96, WS09, CCG96, CPG96, CMP99, GSS01, SHK99]. Table-based [WS09]. table-lookup [CCG96, CPG96]. tables [BGR01a, BGR01b, Lak02, MP85, MSOY96, Nek00, RL95, RTK^{+}16, VV04, Vuc06, Whe97, Yok98]. Taboo [Pig02]. TADOC [ZZD21]. Tag [CC88a, CC88b, ISO98, TC98, YWB01, SP98, Phi91]. tag-phrase [SP98]. tagged [BR04]. tags [ISO99b, TG17]. tailbiting [Nov00]. take [Nat95]. taking [Abe99]. Tale [Gai98]. Talker [ITU06b]. tampered [WCH03, WC02b]. Tandem [KY08, SCM05]. Tangent [JTC^{+}12]. Tangent-Plane [JTC^{+}12]. Tangible [SMLS15]. Tanner [XG10, Xio11]. Tape [Cre94, Ano06n, BKV89, Mia90, VBK89]. target [EW96]. Tarjan [RHC87]. tarp [SSM02]. Task [DP01]. tasks [CL16]. Taxation [Wan08]. taxonomy [Bun97a]. Taylor [CH11]. TCQ [BF99a, JS99a]. TDB [CHW04]. TDSC [YYZ12]. teaching [Hop09]. Team [Bar00a, Bar00a]. Technical [lGAR03b, LW08, SD06]. Technique [ASLB20, CYH94, Dis88, Fou85, Hah74, Jag90, Jez16, JR12, KA12, LYC16, Liu97, MKK12, Por73, PSM01, Ros92, SSP21, Wel84b, Arn97, Bar94, BR04, BKR97b, EEKB15, Faw96, GB98, HBSASA19, HBL^{+}17, Ish89, KY99, LL96, MB00a, MF97a, MK06, PN91, RSK88, SPL20, WC02a]. Techniques [Ano88a, Ano88b, Ano88c, BRALSSMP08, BS20, BMM02, BKS^{+}18, BV04a, Cap78, CY92, Col87, CS94a, Das95, Fel74, Fry85, Gad91, HCBMSS12, IE91, MW97, Pec82, PV21, Poe83, RJ91, RS16, Raj20, RH96, Reg81, Ric79, Ric91, Ric08, SM96, Wan73, Wel99, Ait86, ABG^{+}94, Blo96a, BH10, BPT06, Cap85, CB96, CD95, Cra96, EGS91, Goo03, HM87, HM91, HZ95, Huy94, LO99, LSD97, LGLK05, Lyn69, Mar02, Nad83, Nel92, NcC02, Niu02, RY92, Sha04, Sno76, SS92, SRG^{+}21, Tad98, Vol02, Woo94, YZ07, YG02, YS91]. Technologies [Fra97, Bar90a]. Technology [ISO93, ISO00, Jur88, Jur92, Kar12, L^{+}05, McC92, Moh04, OINC10, Wan08, Aro77, Ben10, BWB^{+}20, BL98, Col93, IEE88a, ISO95, ISO97, ISO98, ISO99b, ISO03b, JC95, Lum13, NDN13a, NDN13b, hyd06, Dic92]. Telecom [Kro00]. Telecommunications [Fou85, She95, Ano06c]. teleconference [WM05]. Telegraph [Nyq28]. Telegraphy [Ano07e, Tet07]. Telemetry [EP65, EP67, LE01, Log04, Log06, Lyn66b]. telephone [RK95]. Teleportation [VO08]. telescope [NSFO^{+}99]. Television [Cap78, Wan73, BBH97]. Template [Gad91, KØJ^{+}12, DDL98, IW94a]. Temporal [CFC05, CZT^{+}11, EC09, Goo91, HG07, LL16, LCG11, PA07, SR16, YX12, ZB12, CAC97, GO96, LDZ^{+}20, PX12]. Temporally [BM97]. Tensor [BKK20, BRP19, EGLS21, LLS^{+}21, NK20, RK09b, XXZ11]. Tenth [YR87, ACM78]. ter [Lia84]. Term [MWC01, NR14, Che00, Dia16, IDKW99]. Terminal [ZK11]. terms [TKYS16]. Ternary [GHV05]. Terrain [GMC^{+}06, Mil95]. Test [CC03, Fun07, Gil08, IYY^{+}11, KKN07, KN11, LCT03, Pom16, RT87, HL05, KN90, MK06, NT05b, RSK88, Rob91b, Vol04]. Test-Data [KKN07]. test.rar [Ano08h]. testability [Vol04]. Testbed [HŘŠ11]. Tested [Koh72]. Testing [RVG11a, RVG11b, Ye01, HL00, HLR01, Sha91, SRW^{+}04, TFC99, ISO95]. Tests [Pom16]. tetrahedra [BR05]. Tetris [GSS01]. Tetris-Hashing [GSS01]. TetStreamer [BR05]. TEXbook [Knu84]. Text [AS97, AJ96b, ABR10, ANS12, Bas91, BA08, BN14, Bel86, Bel87, BWC89, BCW90, BW94a, CHSS08, CHSV10, CK13, Dis88, Fen95, Fen96a, FNV11, FMP83, FCW00, GA08, HC92, HV94a, Jay97, KTSA99, KBD89, Kro00, Kül11, LB90, LCV07, Lea78, LNV97, Man97, MPAFF10, MZ96, NBIT07, NIB^{+}09, NR99, Nav01, Nav02, Rub76, SI99, SKF^{+}00, TKR00, TR93, Vol97, WB91, WBMT99, ZZD21, ZdMNBY00, AT05, AM95, ANdlF04, AMB^{+}02, AN10, Att95, BW99, Bak07, BCD98, BMNM^{+}93, BPMA02, BR04, BD18, BFNP10, BM96, CR95a, CKCW95, Chi98a, CZ99, CT99, CL96, CLM14, Cus90, Dav93, DS96, De 03, DPS99, Fen97, FMMN07, FM96, Gil92, Gu05, GB94, HLS^{+}04, HV93a, How96, INI^{+}16, Jak88, KK04, KTS^{+}98, KG95, KM97b, KM98, KM99b, LZ06, LDK99, MN08b, MZ92]. text [MSYY98, Mos98, NYOY94, NKT^{+}01, NT05a, NM07, NSA06, NMWO96, NC97, RK15, Sab94, SOI00, SMOY97, SNZBY00, SZM03, TC97, TICH98, TC98, TH01, TN95, Tin77, TVW97, WC88, WBN91, WBM^{+}94, WT05, YS07, YBW00, Yok98, YMYS99, ZTSM05]. text-compression [CL96]. text-compression-based [LDK99]. Textbook [Bla09]. textes [LG78]. Texts [AMPHTSM10, BBH11, BBHS12, CFG12, FT04, Kle07, LG78, BK05, BFG09, CK06, DS04, KS01, MSZ94, SS00, YMYS99]. Textstream [PGW^{+}17]. Textual [AC06, CA97, DH18, SS82, WBH^{+}92, WBE^{+}94, AL98, AL00, FGMS05, Mul87, NR95]. Texture [And12, CC06a, CHMW10, DSM07, DMS08, IK00, LLPP19, LMH^{+}09, MP12, MCHAM08, OBGB11, SWWW15, COMF99, Kra10, LGM21, RS99, WRCB02, WB00]. texture-based [LGM21]. Texture/Depth [LMH^{+}09]. textured [CVDL16]. Textures [OBGB11, WS09, KPM16]. Texturing [DKB^{+}16]. Their [AC09, Bil20, BBHB94, Fry85, WK92, Du00, Gin95, G^{+}75, NAIP03, Sha04]. thematic [EHSC92]. Theme [Gal78, MW85]. theorem [GL01, Wik03]. Theorems [Zei93, Zei94]. Theoret [De 00b]. Theoretic [Ach76, Cac98, HMR10, LT08b, MAL10, vL94]. Theoretical [Boy91, KHCS08, TW01b, ZE00]. Theoretically [WY10]. Theorie [Haa10, Haa12]. Theory [ACM78, ACM79, ACM95, ACM13, Abr63, AC07, BCH02, Bel87, BP85, BK90, Cat84, Cha77, CHSV10, CMP94, CT06, Dav95b, Dav64, Fis95a, Has18, Mal89, Moa86, Nyq28, RB98, S^{+}88, Sha48a, Sha48b, Sto88, Ven07, Vol02, WS12, Yan10, APC92, AL98, BK00, Ben10, Ber71, Cho05, CA93, Dav95a, Haa10, Haa12, Hoc95, Ind91, Laf00, LPZJ15, PPR02, Sle74, Sto83, VG06, Ver98, ZDR06]. theory-based [LPZJ15]. there [Lut88a, Lut88b]. Thermodynamic [Zur89]. Thin [YNXC07, CSM00, CSM01, CS02]. Things [BMG18]. third [BDLS96, CLA12]. Thirty [M^{+}98]. Thirty-Second [M^{+}98]. Thomas [Dol06b, GHV05]. thread [Rob91a]. thread-like [Rob91a]. Three [CKC09, Cm10, Cho56, DZZ08, LF03, PW86, SW12a, WXCM99, CL97, Ish89, KP97, PHG86, Tro96, WG00]. Three-Dimensional [LF03, WXCM99, Ish89, KP97, PHG86, Tro96]. three-layer [WG00]. Threshold [DSG12, DGG10, FA97, PMH95, WV00]. threshold-based [WV00]. Thresholding [DGG10, JZL^{+}12, ZX11, RV94, RL96]. Thresholding-Based [ZX11]. Throughput [BR07, MCM^{+}17, WZW11]. TIFF [ISO98, CC88a, CC88b]. TIFF/IT [ISO98]. tight [PK05, YSZ11]. tightness [LMC05]. tiles [Lee00]. Tiling [FOhC09, FO10]. tilings [BK96]. Time [BEGV18, BMG18, CPS08, CME06, Dav66, Dur60, EY81, EKT12, HB09, JSS15, KKP12, LYC16, Liu10, Lyn66a, MMB07, NBIT07, NIB^{+}09, PB11, PS12, Rya19, Sau94, SYZO15, SYO15, SHW^{+}01, Tao82, WYM10, BKV05, Bal97, BK96, CM05, EEKB15, FSC91, FGMS05, FA94, GKPS05, GCK^{+}96, GLM^{+}21, HC98, Hoa99, JB95, KL03, KK08, KSY95, KMP05, LCJP21, MJB92, MSW04, Mer91, Mil95, Mod03, Moh02, MSC00a, NMW97a, PST05, SS99, SJ94, Sew01, SCM05, Taw93, Weg07, ZL00]. Time-Compressed [EY81]. Time-Compression [EY81]. time-dependent [BKV05]. Time-Frequency [Liu10, BK96]. time-recursive [KSY95]. Time-Series [Dur60, EEKB15]. Time-sharing [CME06]. Time-space [BEGV18]. Time-Universal [Rya19]. Time-Varying [HB09, WYM10]. Timed [JP07]. Times [Wil89]. timing [KN99]. Timothy [Neu10]. tion [Lia84]. Titan [PP10]. Titan-R [PP10]. TLC [PM17]. TMW [MT98]. today [Mac12]. Toggle [PBO^{+}15]. Toggle-Aware [PBO^{+}15]. Tolerable [Sad98c]. Tolerance [HHCH11]. Tolerant [HAM17, Log06, BEQ98]. tomographic [Vil93]. Tomography [CFC05, FCL13, MGBRMSS11]. Tomorrow [Zei93, Zei94]. Tone [CWS00, Roe77, YNXC07, GB98, ISO95, ISO97, ISO99b]. Tonnage [JS99b]. Tool [Kro00, Swa08, TMS02, CE03, NT05a, Yok98]. Toolbox [Ano06e]. Toolkit [Del93, Kro00, SGA^{+}16]. Tools [DDT^{+}12, Kro00, MKH10]. Top [SS05a]. Top-Down [SS05a]. Topics [Nyq28]. Topography [Ost35]. topological [Li02b, TR98, YKLK99]. Topology [LRMPGR01, MKSS12, MCL12, TRS03]. Topology-Based [MCL12]. Total [BH12, LMV03]. Toute [Pea90, Sie12, Sie90]. Trace [MM03, FG97, Sam89, SJ94, LCJP21]. Traceback [ASP03]. Traces [MMB07, PV06, LZ05, SJ94, ZG02a]. Tracing [HO19, SLM^{+}17]. Tracking [CFGL12, Vos06]. Tracks [Bar00a]. Trade [KN11, MCM^{+}17, Ruf95, BEGV18]. Trade-off [Ruf95]. Trade-Offs [KN11, MCM^{+}17, BEGV18]. Tradeoff [VRL09, Vas07]. Tradeoffs [Che06, Cam79, LO99, Sew01]. Trading [HJW02]. traffic [PBEA95, SLZ^{+}19]. training [DR06, Ind91]. Trajectories [CXF12, LCJP21]. Trajectory [HSZ17, LMJ^{+}21, SY18, DHW^{+}17, SSZZ14]. transaction [YJGS00]. Transductive [LZX^{+}11]. Transfer [BLFF96, FGM^{+}99, GL03a, JKV15, PN91]. Transfom [ANR74]. Transform [ABM08, AR09b, AR12b, BF07, CSF77, CJDL09, CHSS08, CY91, DZZ08, D^{+}92, FL90, GVSSBRAL08, HH99, HHCH11, Ish89, JLJ08, KB93, KOO09, Liu97, Liu08a, LSW09, MHKK03, MHKK06, MCRKC09, Mul97, Neu10, RS16, RY90, RYe00, RY01, Rea72, RP88, RM21, RC09, SFT03, Sha75, SM95a, Str99a, TS17, TM91, Wat94, Wic94, WK93, XW10, Ano06g, ABMD90, ABMD92, AL01, BG06, BW98, BK94, Bod95, BB93, BB95, CT98a, CR97, Cod92, Eff97, Eff99, EFZ03, Ghi05, GZV97, GK98, Ham94, HB99, HV06b, JJM18, JM03, KC91, KM99b, Kül12, LK91, LMS94, LM08, MRS05, Man01, MS03, ML98, MW92, MSR04, MZ91, MM00, PO00, PRM97, PN91, SLDJ05, Sew01, SM95b, Vil93, WM01, Xue99b, Xue99a, YL05, ZZ98, ZMAB03, CL97, Has18]. transform-compressed [Bod95, BB95]. transform-vector [BB93]. Transformation [Ano06w, BJ95, HESF11, MS12, RV09, SI99, SR16, Vo07, WSDTO11, Arn02, BK00, NT20, Pop05, Sad98a, Sad99, Zha99]. Transformations [Col86, TEA08, ZG02b, ALA98, Pro93, RK93a]. transformed [AMB^{+}02, Cha00]. Transforming [Giv58, KSCE16]. Transforms [ASH87, BGBV09, Bel11, BYR06, DSM07, FG07, Mal07, MEO11, MF10, RB98, SCV11, Sie90, SA90, SR05, XWLZ08, YZ09, Str99c, AK99, BZM98, CVC95, CDSB98, FW02, FH02, GDV03, Laf00, MS03, Mon93, MM05, NZ06, SDJ04b, SKCB95a, SKCB95b, SDV11, VED96, WM94, ZGS02]. Transition-Nets [VO08]. transitions [BBZ14]. Transitive [Jag90, Sha09]. Translation [Fel74]. Translation-Permutation [Fel74]. Translator [Ame06]. Transmission [AF87, BZZ08, BH07, Bow70, Cap89, DR83, EEJ07, EJ07, Fan49, Fou85, FK96, FDU80, GE07, HB09, HKW10, JRALMSS11, KY08, KY10, Kno80, Moh04, MPALSS^{+}12, NYJ08, NYJ09, Nov98, Nyq28, OLHL08, QB09, Sim92, XJ14, ZZPB09, BZ97, BY00, BBF02, BK94, BR05, BdB06, CC00, CF98, CF99, CRR98, CF96, CPT^{+}20, HR06a, HCL99, Hes08, HN02, JF96, KA98, KT02, KMS98, KCR98, MDPM00, MYRD04, PBC05, Per89, Per90, PR01, QJRA99, Ruf94, SAR00, Say92, Tro96, TMM96, WS91, XWDY06, YJE04, ZAA00]. Transmission-Efficient [XJ14]. transmissions [HSP^{+}13, PHG86]. transmitted [FS03]. Transparent [KMM^{+}12]. transport [BG03a, HR06b, Woo00]. Transpose [SH92]. transputer [OPS91]. traversal [GKPS05]. treatment [GFV97]. Tree [Car00, CTC^{+}09, CPP95, EEH17, FV16, Fre91, FLZW10, FB16, GHLN18, HC06, Kid09, KS00b, LSH12, Lin78, LMM11, LST99, McC76, OHSS12, Pro93, Rob97, TTW07, VNHB12, Vol97, WST95, YK10, AS97, ASS98, AM95, AKF06, BE96, Bun97c, Car01, Cha87, CM91, Cha91c, CSP05, DC97, DHH04, Eks96, GYLC09, KJP99, KLAH99, KF02, KRT05, LSC91, LS92, LS93, Lin97, LHY12, LM08, LC91, MJB92, MSW04, MF96, Muk04, MMA99, NPW97a, OG04, PRM97, PGOO94, RMG94, SOI00, Sen06, SN96a, Suz96, TVW97, Tur96, VO91, Von04, WZ91a, WY91]. Tree-based [Car00]. Tree-Structured [FB16, HC06, CPP95, Car01, KJP99, LS92, LS93, Lin97, LM08, MMA99, OG04, PGOO94, RMG94, SN96a, WZ91a, WY91]. Trees [AHK97, CNS12, EK16, FIKS18, Fra97, HPST11, Jon88, Köp21, LP95, Lee00, MSW10, NP16, SP96, Tar79, UUiN12, YGNM07, Bas91, CCA99, CR96, De 06, EN84, FLMM09, FT00, GMNK06, HC98, JS95, KM16, KYNC96, KP97, KRSS21, Lar96, Lar98, LMO97, Maß15, Mat15, MMS91, NP00, NAIP03, NdR04, OM03, Tar95, VKG01, Yok93, ZL95]. TreeWavelet [BF07]. Treewidth [PT10]. TreeZip [MSW10]. Trellis [Abo06, BKYK94, HH08, JCF93, Lin78, Ste81, Abo95, BV96, BF98, CM97, CPR03, DHS01a, ENA03, EANG04, EHAN05, ENA05, FW91, HW98, JT98, LF95, LA05, MF98b, PBC05, TM91, WO01, EGNA06, MG10, SR09]. Trellis-Based [HH08, CPR03, PBC05]. Trellis-Coded [Abo06]. Trends [MLDH15, Mer91]. treshold [DGG09]. tresholding [DGG09]. Triangle [DJCM09, Ise01, JBG17, MKSS12, Ros98, CCMW05, LV14, Szy02]. Triangular [Giv58, BPZ99]. Triangulated [DGGP05]. triangulation [LLY00, PELA02]. Trie [Nav02, Mas91]. Tries [ACF^{+}21, GO15, JSS15, NT02, RL99, AC06, RS02]. Trinary [ZCW^{+}11]. tristate [CZAR20]. TRLE [LCY07]. Trondheim [Fis95b]. True [BES07]. TRUEW [BG03a]. Truncated [NAIP03, NP00]. Truncation [BFT11, Das95, Ma96, RD82, FNK94, Kru92]. TS [ZZ98]. TSC [BR04]. TSVQ [Eff98]. Tucker [BKK20]. TuckerMPI [BKK20]. Tuning [BR10]. Tunstall [HN07]. tuples [ZL95]. Turbo [CPR03, GFV98, MB02, ZZPB09, AG02, GCS00, LTB04, XS05, ZGF02]. turbo-codes [GCS00]. turbulence [Wil00, Wil02]. Turbulent [WFG^{+}94, DJID20]. Tutorial [Hor06, KNY96, Lew95, Pan95, RG88]. TV [Boo99, EY81, Zha99, Z^{+}13]. Twenty [LL08, ACM95, LS95a]. Twenty-Seventh [LL08, ACM95]. Two [BF11, BGK95, BES05, BES07, CCKH21, DGGP05, Fra97, HSL^{+}20, JLL^{+}10, LTO11, Liu97, Log04, MEIS09, Mar80, McM56, MPL99, Mof91, RLG09, SL08, Ski05, TOL11, WFG^{+}94, YYZ12, ZK11, AB92, ABF94, Api91, BCSV06, BAGCGC^{+}06, CG91a, CG91b, Cha00, DWB04, FN93, GE06, HV06b, Kle89, Kri02, LMC05, LGO99, Li10, LG99, Mäk89, MBB08, NPW97b, SLZ91, VW98, Wak05, WG00, WBE^{+}94, YSZ11]. two-description [DWB04]. Two-Dimensional [BGK95, BES05, BES07, Mar80, WFG^{+}94, AB92, ABF94, BCSV06, HV06b, MBB08]. two-hop [GE06]. Two-Layer [CCKH21, HSL^{+}20]. Two-Level [Mof91, Ski05, CG91a, CG91b, NPW97b]. Two-Line [RLG09]. Two-Manifolds [DGGP05]. Two-Part [BF11]. Two-phase [YYZ12]. Two-Rate [SL08]. two-sided [Kri02]. Two-Stage [Log04, FN93, WBE^{+}94]. Two-Step [JLL^{+}10]. Two-Terminal [ZK11]. two-valued [BAGCGC^{+}06]. Two-Way [LTO11, TOL11, SLZ91]. Type [DGG10, FWS11, LE07, MMZ12, MSFZ08, Abr94, AR18, DGG09, GLL04, KZ02, KKT10, SW97, Sto96]. Type-Based [LE07]. typed [TKYS16]. Typefaces [Knu86d]. types [BMT91, ISO99a, ISO99b]. Typesetting [Knu86e]. Typical [CZ98a, Szp91]. Typography [BBKF95].

U [Ano09d]. U.S [Ano82]. ubiquitous [QMCX19]. UEP [CRR98]. Ulam [Wei06]. Ultimate [UPX03]. ultra [TG17]. ultra-lightweight [TG17]. Ultracompression [Des08]. ultrasound [RS99]. Ultraspectral [HI12, HAH05]. Unbounded [AHK97, AF87, BY76, Cap89, CTW95, CT97]. uncompress [Ano06q]. uncompressed [BD18]. Uncorrelated [DØG08].

Understanding [RS21, ASS97]. Unencumbered [lGAR03a]. Unequal [GLM^{+}03, Hal04, Dum06, HGH05]. Uneven [DZ08, LZM10b, CCV04, CCV05, DWW02]. Unicode [Ano06w, FB98, SD02, SD06, Uni07a, W^{+}00]. Unidirectional [BSG10]. Unification [JdVL03, Wol99a, Wol99c, KF96]. Unified [BK90, GCL06]. Uniform [DSG12, ASS98, CY91, DH97, DH98, FZ02, FBS04, GN01, MA05, Prz00, SF05]. Unifying [Bel87, NZ06, SI99, Fow97a]. unions [KL16]. Unique [McM56]. Unisys [Rod95]. Unit [PP10]. Unitary [Giv58]. United [HPV09, H^{+}10, HPT11, SM01a, Boo99]. units [SRG^{+}21]. Universal [AT05, AHK97, BK00, BF12, CM91, Cha91c, Cov96, Eff99, Eli75, FWS11, FM95a, FS98, FK85b, FK96, FDZ00, GZV97, JMW09, KT81, Law77, OS90, PRM97, Pig02, RF00, Ric79, Ric91, RA04b, Rya19, Sad96b, SC01b, SR09, Suz11, TH03, TW92, Wil89, Yam00, ZF92, ZL77, AECG93, CDDM05, CME05, ECG94, Eff97, FFW98, IF06, KIHO95, Liu91, MF97a, MSW04, MF04, NYOY94, OWW^{+}04, OS03, RRG95a, Sha04, VW98, WS05, WZL91]. Universality [LLZ93]. universe [IEE88a]. University [IEE95a]. Unknown [Lee81, AT-96, ATS06, Abe07b, AA01, Ano06t, Ano06u, Ano07f, Ano07g, Ano07h, App06, BOC01, Cal06b, CCMW03, CM04, GFV98, LBE07, Lan09, Mal06, Ohi01, Rat08, Sof07, UCL08, UDA08, Wol99b, amu06, cci01, suz06, ukl07]. Unless [PWM18]. unmanned [PHG86]. Unobtrusive [lGAR03a, MMB07]. UNRAR [Ano06v]. Unreliable [EEJ07]. unreversible [BG03a]. unstructured [ABF^{+}02]. Unsupervised [BKV05, VMFG07, YWMS08]. Up-sampling [ZWZ11]. Update [AM96, AM98, DLOM08, FOF11, Per03a, Cha00, Sto98]. updating [OT03, TO04]. Upgrading [Whe97]. uplink [YK20]. Upon [Tao82, CW91, Hoa99]. upper [AL81, CH06, GL03a]. upsampling [You98a]. USA [Auc98, KRW^{+}93, SC04, SM09, SM11b, ACM13, BBKF95, HWS06, HHG07, HHSS08, HHW05, IEE95c, R^{+}06, R^{+}07, Sch03, Sch04, USE93]. USB [Kro00]. Use [Ano06e, CDL^{+}19, Dub11, FOF09, MPASM^{+}09, SW17, AdlF06, APC92, BSS05, Bli91, Ish89, KS89, LM03, Pow93]. Used [EP65, SB85, EP67, FS03]. Useful [PT10]. Usefulness [G^{+}75]. Usenet [Ano92c]. User [BM11b, MPALSS^{+}12, YZLH16, XMi03]. user-configurable [XMi03]. User-Navigation [MPALSS^{+}12]. User-Specified [BM11b]. Using [AJ95, AOAAH17, AKL18, AM05, BZL^{+}12, Bel11, BT93, BRD10, Bon95, dABdSFNA08, CA06, Cap89, CD10, Cha89, Cha91b, CY92, CC01, CKC09, CPS08, CJDL09, CWY^{+}10, CWLS15, CMK^{+}16, CGC97, CC06b, CW84, Col87, CH87, CMRS00, CD91, CK93a, CK94a, DØJJ09, DZZ08, DS08, DRR10, DOK88, DSM07, DMS08, DMS10, Dou93, EEJ07, ECBL12, EGF08, EK16, FPR95, Fel74, FOF11, FH08, FL98, FDU80, FVP08, GHLN18, GVSSBRAL08, GW73, GV09, HB96a, HY10, HOR05, Hoc95, HGR11, HV92c, HGFL09, HZKL11, HSL^{+}20, ITU94, JS99b, JLJ08, JR12, JCF93, KHCS08, KK07, KRK09, KR10, KKJ11, KBN08, LZPB11, LZWL11a, Lak02, LS04, LBMN08, LLA08, LF03, LSC10, Liu97, LP07, LWZI12, Log04, Mal07, MS12, MWC01, MSM09, Moh04, Mul97, MFD^{+}09, MF10, MJZMA08, Nav02, NN97, NMW97b, PMZ13, Pec82]. Using [Phi92a, Pie03, Poe83, PSM01, QLQ11, RS16, RJL11, RM21, RL99, RK15, RD82, Rub79a, SF11, Sha93a, SSP21, SHW^{+}01, TY08, Tan93, TS17, The89, Tra87, Uhl96, VO08, VV04, WB17, Wat94, Wel72, WCY^{+}10, WPA08, XDZ11, XJ14, XXZ11, Yac98, YGNM07, YBW00, Yok97, YM08, Yor95, YK10, YTY12, YDZL11, ZCW^{+}11, SM97, AG02, Abo95, AM95, AJ96a, Adj06, DS92, AH02, ASP03, ADCU08, ALRT16, AMAM13, AC07, ABMD90, ABMD92, Arn96, ACM^{+}15, BS04, BL97, Bas91, BS95, BH00, BM97, BFMX03, BR04, BM99, BK96, Bet94, BK94, BZM98, BBQ^{+}95, BBHK14, CPG96, CZ98b, CC03, Cha91a, CM91, Cha91c, CC93, CC19, CMP99, CB96, CN96, CMW99, CT98b, CDSB98, CLR05, Cor86, Cre96, CR95b, CK93b, DS06, DC97, DJL91, DHS01a, DDL98, DTZZ12, DBTNT00]. using [FC98, FZ05, FZE03, FSC^{+}11, For99, FWA01, FA93, FA97, Fow98, FW02, FM96, FCW00, GB92, Ghi05, GB06, GFH00, GP95, Gul04, HMS92, Hag07, HR06a, HIMM20, HKA06, Hau12a, HLV95, HBL^{+}17, HV91b, HV92a, HV96, Hsi01, HOG04, Ind91, IW96, IMB^{+}13, IDKW99, IY00, JM96, JBG17, JYHC03, JB99, VG95, KF97, KAB11, KP97, KKA02, KC91, KRT05, KMS96, KCR98, Kri02, KM97b, sKJ01, LLN^{+}04, LLW^{+}14, LZ06, LOS20, LCB03, LW14, LKRR91, Lee91, LW99, LF04, Lev96, LK91, LCT03, LTB04, LG01b, LA05, LMS94, LA96, LY91, LRH^{+}21, LXG03, LOMSS05, LC91, LLY00, MSE17, MT92, MF98a, Mas91, MP12, ML98, MP85, MZ99, MM05, MSYY98, MBBA91, MMA99, MM99, Mul87, MSOY96, MOK00, NMWO96, NPW97a, NPW97b, NU05, Nov98, OG93, OM03, OY06]. using [OR94, OG04, PO00, PM99, PM18, PSH00, PELA02, PR99, PR01, QZH10, RV94, RK93b, RSV^{+}00, Rob91a, RS99, RRG95b, SAR00, SF05, SRKS95, SLDJ05, SO94, Sha94, Sha93b, SHWH12, Sto96, SY04, TBKM05, TKR03, TT94, TY04, TM05b, TC96, TICH98, TH01, TG17, TGFZ03, TM91, TWC^{+}09, Tsa91, Tur96, Uhl95, UH96, VPL98, VRHL07, VSG00, VRR^{+}16, Vil93, VED96, WRCB02, WSS94, WBS05, Wan06, Wei90, WS06, Wic03, Won91, Woo94, Wri39, WF03, XWL03, XSX05, XHS05, YS07, YWPG03, YM97, ZMAB03, ZTSM05, ZGF02, ZFE04, ZIIK05, ZDR06, ZLN06, AF87]. Utah [SC93, SC04, SM08, SM09, SM11b, SR91b, SC92, SC94, SC95, SC96, SC97, SC98, SC99, SC00, SC01c, SC02, SC03, SC06, SC07, SM10b, SM12]. UTF [Ano06w]. UTF-16 [Ano06w]. Utilities [Del93, Nel87, Ano09f]. Utility [Kro00, Ano92c, BP00, Per05]. Utilization [PS12]. Utilizing [BLNK14, WV99b]. Utrecht [vL94].

V [HPV09, Ma96, RHC87, Rok90]. V.42 [ITU94]. V.42bis [Tho91b, Tho92]. Vague [Dav93]. validated [AFL96]. Value [ABM10, AS14a, KXPZ12, MN08a, Eve98, JSC20, KV19, SCC12, WF03]. value-adding [Eve98]. Value-Aware [AS14a]. valued [BAGCGC^{+}06]. Vancouver [LL08]. Vapnik [War97]. Variable [AJ95, AJ96b, BH01, BFL^{+}10, BF95b, But95, CMMS17, CS94b, DRG94, DE92, ECG94, Fen96b, Fen02, Fre93, Gha87, GW73, GM59, HDK^{+}12, KMSW18, LV02, NM10, OBGB11, PSR04, PS05, PM91, RJL11, Rez07a, Sal07b, Sav98, Sav99, SK10a, Sho08, SG04, TWM95, TW92, TW00, TW01a, TW01b, Wel72, WV98, ZL78, AJ96a, BH93a, BH00, CZAR20, CL06, CK95a, CMP99, CS93, DS98, DTZZ12, DW03, ENA05, EGNA06, FA93, Gha84, HGS97, HCL99, JKL13, KT02, KB99, KMS96, LKH^{+}00, ML98, MRG99, NYC05, PO00, PM99, SDJ04a, Stu94a, Stu94b, TKR00, WV99b, XWL03, YKLK99, YR02, Nas13]. variable-bandwidth [KB99]. variable-bit-rate [MRG99]. variable-depth [KMS96]. Variable-dimension [DRG94]. variable-interval [DTZZ12]. Variable-Length [AJ95, BH01, BF95b, But95, Fen96b, Fen02, Gha87, HDK^{+}12, RJL11, SK10a, TWM95, TW01b, Sal07b, SG04, CZAR20, CMP99, Gha84, HCL99, JKL13, KT02, PM99, Nas13]. Variable-Rate [ZL78, ENA05]. Variable-to-Fixed [TW92, Sav98, Sav99, TW00, TKR00, XWL03]. Variable-to-Variable [BFL^{+}10, KMSW18, SDJ04a, Stu94a]. Variable-to-Variable-Length [Fre93]. Variables [DSG12, Suz12]. Variant [PS12]. Variants [CKP85, Sad98b, Tho91b]. Variation [BH12, AECG93]. Variation-Based [BH12]. Variations [Gal78, MW85, Mye86, KGK00]. Various [LG78, You98a]. Varying [HB09, LSC10, SSP21, WYM10, CHW04, FM95a, PAB03, SC01b, SCM05]. VBR [Hoa99]. VByte [LKR17]. VC [Ano08i, LK08, LW08, MY16]. VC-1 [Ano08i, LK08, LW08]. VCDIFF [KMMV02]. Vector [BŁ01, BLS11, BBBP10, CKC09, CZZZ11, CN96, CM02, Cie98, CS94a, CS94b, DFA^{+}10, DMS08, DMS10, FW95, FBS04, GG92, GK94, GV09, HLC07, LLA08, LG99, LBG80, MWLW09, MSM09, Moa86, MFT07, NR95, OG05, PCD17, Ram85, SG09, SG12, SWWW15, Sou95, TV09, TPY95, TSS99, Tao82, The89, TRS03, Tra87, WW92, Wu92, WS08, Yoo06, ZKØ10, ZK11, Zha11, AAS95, AECG93, ABMD90, AM93, BH93a, Bar94, BF95a, BV96, BB93, BK92, Car01, CM97, CVC95, CCG96, CPG96, CGS93, CL95, CS93, CGR91, CPP95, DS06, DRG94, Del95, DY91, DSV00a, DR06, EMS03, ECG94, FSC91, FE99, FE01, FA94, FA97, Fow97a, Fow97b, Fow98, FH02, GCK^{+}96, GK95, GFH99, GP95, HMS92, HM96, HGS97, HW06, HM93, IDKW99, Isr91, Jak78, KGK00, KR01]. vector [KZ02, KFSM04, LLCC95, LO99, Lev96, LGO99, Li02a, LS92, LV92, LS93, LLZ96, LA91, LL96, MJB92, MSW96, MT92, MPL02, MR93c, Mis91, MMA99, MM99, NOG96, NM91b, OG93, ØJH05, OG04, PGOO94, PN91, PM95, RRG95a, RSC99, RS01, RMG94, SVS99, SHD02, SS92, Tse05, TFA93, WHH^{+}99, WS00, Wan06, WZ91a, WZ94, WF95, YP03, YW04, YG03, Zee98, ZC91]. vectorized [LOS20]. Vectors [Teu78, Bör18, GK98, GC16, KS16, SB06]. Vegas [ACM95]. vehicle [KM16]. Vendor [Sim97]. venerable [Ude93]. ver [Fed93]. verification [KMA^{+}20]. Verlag [Hoc95]. Vermont [IEE96]. versatile [BP00]. Version [MNG03, Ado01, Deu96a, Deu96b, DG96, WZ91b]. versions [ISO99a]. versus [DS96, HSX02]. Vertex [CM02]. Very [BJZ94, FLA^{+}03, MD95, PC08, WM05, ZD81, AAJ04, ABF^{+}91, CAC96, CAC97, JOC97, NZ95, PB98a, PW03b, SC01a, VL91, YA13, Zir07]. VF [Kid09, YK10, YK11]. VF-Coding [Kid09]. VI [H^{+}10, Ano03, Sch03]. Via [CCM96b, VCP09, AN08, AB10, BKK20, BR06, BJ18, CYH94, CCM96a, CDL13, CKV93, DB10, Duv88, Eks96, EBH^{+}17, Eva98, FCL13, FK09, GLM^{+}21, HLL^{+}20, HCD04, KLJ12, KYNC96, KS96b, LV92, LSW08, LZZ^{+}12, LNV97, MNBC18, MSW04, MS02, MN09, MSC00b, Nat93, RPE81, RK09b, SS82, VK91, VK96, WO00, WCB97, WCM98, Xio11, YWMS08, ZZZ^{+}12, ZLT^{+}18, ZL78, dMGdC10]. victim [GAS16]. Video [AFHU97, AC07, ARA91, BGU96, BZZ08, BZL^{+}12, Bar81, BS20, BS95, BZ08, CA06, Cam79, CT09, CXZD12, CDW11, CZT^{+}11, CB96, CST11, CWRL11, DVB06, DMC^{+}09, DG17, EC09, FWS^{+}08, FWZ^{+}12, FG07, Fow98, GKSB17, GS12, GJ20, Ham95a, HMR10, HKMS08, HG07, ITU02, Jay97, JFC12, JLL^{+}10, JRALMSS11, Kar12, KØJ^{+}12, KD18, LZPB11, LE94, LZX^{+}08, LZX^{+}09, LZM^{+}10a, LK08, LC03, LSD97, LT08b, LLS^{+}21, LG01b, LHD^{+}20, Lio91, LHS05, LMH^{+}09, LLCC11, MTA^{+}20, Man98, MWC01, MEO11, MN08c, MN09, MPFL97, MF11, Nat95, NYJ08, NYJ09, Nga95, PB11, RH96, RC09, Ric03, Ric06, SYZO15, SYO15, SYDR07, SMW07, SS07, SCV11, SCM05, SCW11, SWW12, Sof05, SWGZ09, SR16, SR05, SjWL08, TSS99, THHS93, TF11, Tsu72, WZMG07, WWY^{+}07, WW10, WJDZ10, WHW10, WSDTO11, WV98, WS09]. Video [YXPM11, YMK10, YX12, YDZL11, ZXM^{+}10, ZWZ11, ZJW^{+}12, ZWZ^{+}12, ZB12, ZLRZ09, ZX11, ZBMM97, dCDV07, Str99c, ARZG03, Adj06, APC92, ASG99, Ano04b, Ano08i, BBF02, Bar90a, BM97, CS92, CCA99, CCV05, CB97, Chi98b, Chi99, CP03, CMWM00, CAC96, CAC97, CKS95, Cre97, CF97, DVS^{+}14, EGS91, EGCK98, FG01, FA94, GK99, GSW02, GO95, GLM^{+}03, Gov94, GO96, Hal04, HKA06, HLV94, Hoa99, HSP^{+}13, HN02, HGH05, JOC97, KJP99, KP97, KL03, KF94, KLAH99, KMS98, KH04, Le 91, LK91, LKH^{+}00, LMO97, LAPL07, LG99, LBS98, Lut88a, Lut88b, MEMS06, Mar02, MF98a, MYRD04, MTTH13, MF98b, MRG99, MSC00a, Muk04, MM99, Nat94, NZ95, NcC02, PX12, QJRA99, RB01, RLR04, SCKS97, SH03, SH04, SCHB05, SRW^{+}04, SLR05, Spa00, SM95b, TY04, TN99, TCJ93, Tei98, Thy10]. video [TC06a, VL97, VJB^{+}95, WHH^{+}99, WS00, WSS94, WL99, WXCM99, WM05, WLL06, WC02a, WS91, Won91, Woo00, WKM^{+}98, WJBM05, XWDY06, XSX05, YH00, YW04, YML97, ZRR00, ZR02, ZWS92, ZGS02, FWI99]. Video-bandwidth [Cam79]. Video-Coding [MTA^{+}20]. Video-on-Demand [JRALMSS11]. Video/Images [KØJ^{+}12]. Videoconferencing [MJZMA08]. Videophone [XSW07]. Videos [JSCM17, SST^{+}12]. VideoSave [Ano04b]. View [A^{+}87, LLLZ12, LMBN08, FJD11, KS91, ZLRZ09]. Views [Bar00a, Bar00b]. VII [HPT11, IEE95a, Sch04]. Viral [IP06]. Virginia [WGM88]. Virtex [SDM02]. Virtual [And07, HKKW13, NBK16, VK17, BDLS96]. Visas [Bar00b]. Visibility [LSC10, SWA94, WSA94]. Vision [HSP42]. Visual [dABdSFNA08, DSM07, EC16, EFPS16, GHD^{+}14, Ric08, RL08, CE03, DVS^{+}14, Rob91a, SHK99, WS08]. Visualization [CH09, GLS99, LMBN08, NS01, PW03a, Sou95]. visualizations [MPL02]. visualize [QLQ11]. Visualizing [MFD^{+}09, RL09, WYM10]. Visually [Wat93]. Vital [Bar00a]. Viterbi [JS99a]. VLC [ND04, WZMG07]. VLDB [FLA^{+}03]. VLIW [PSSS03, RS04]. VLR [FJ07]. VLR-Based [FJ07]. VLSI [Ach94, ATL^{+}88, Bec93, Ben96, Ben97, FSC91, FGC93, FGC94, HL05, Isr91, JB95, Nga95, RMB91, SFT03]. vocalizations [Sta09]. vocoders [DRG94]. Voice [ITU89, Par87, Smi79, SMS91, Den01]. voiceband [Ng96b]. VoIP [KY10]. Volume [FM12, HB96a, Knuta, LMBN08, LGK97, NDN13a, NDN13b, NS01, Pom16, CR97, EFF00b, FY95, LCY07, MSE17, Lum13]. volumes [BZM98]. Volumetric [PW03a]. Volunteers [Bar00b]. Voting [XXZ11]. Voxel [DKB^{+}16]. VP [LSH12]. VP-Tree [LSH12]. VPVQ [WS08]. VQ [CCKH21, CS95, DS08, DSM07, DE92, GFH00, HOG04, HAH05, KG95, Lev96, LBH91, Lin97, SA93, SN94, Wak05, Wak06, Wak07, WSS94, WAL05, WG94, WWW97, WF03]. VQ-based [KG95, SN94, WAL05]. VQ-Compressed [CCKH21]. VR [VK17]. VRML [Car97b]. vs [CME06, DCP17, Lan84]. VSPC [Fry85]. VTC [NU05].

W [MMZ12, Sew01]. Walk [ZYT^{+}15]. WAN [SHWH12]. WAN-optimized [SHWH12]. WARC [ISO09, ISO17]. warping [Nos97]. Wars [Nel87]. Was [Ano04b]. Washington [IEE95c]. Wasserstein [MOG05]. Watermarking [ASLB20, DG17, GKSB17, SDFH00, WHZ12, Way09]. Wave [LLYH14, HSV04]. Waveform [Rob94b, Ste91, PCRLSB00, Rob94a]. Wavelet [AR08, BT93, BMA04, BBH93, BB93, BBQ^{+}95, CA06, CN07, CMQW90, CMQW93, CMQW94, DPS93, DZZ08, Dav95b, D^{+}92, DGG09, DGG10, Ham94, HHLW20, HHCH11, JR12, KAB11, KNB10, Kro00, LZM^{+}10a, LLA08, Lew95, LP07, Liu07, Liu08a, LSW09, LLJ09, MG08, Mal89, MWC01, MP12, MR93b, MAS00, Mul97, MOK00, Prz98, QWH12, uR96, Raj20, RB98, RLG09, SMC99, Sha93a, SWGZ09, SR16, Uhl95, Wel99, Wic94, WFG^{+}94, Wil00, Wil02, WK93, XW10, YGNM07, SM97, Str99c, AK99, AMS^{+}03, Anoxx, ABMD90, ABMD92, Ara13, ARH03, BW98, BW95, BZM98, BG96, BCP20, CSP05, CO97, CO98, CDSB98, Cod92, CDGO02, CPP95, Cre96, CDR93, DB02, Dav95a, DH00, DYW^{+}96, EHSC92, FC98, FW95, FK95, For96, FW02, FH02, Fow06, FJ96, GFC^{+}09, GSW02, GSPR19, Gul04, GMNK06, GS05, HB99]. wavelet [HLR01, Hsi01, HB96b, HSV04, HV06b, JU10, KA99, KJP99, KP97, KLAH99, Lee00, LK91, Lin98, LM08, LRO97, LKA99, Mal99, MPL02, MM97, MSR04, MASC98, MAC^{+}04, MS95b, MM00, OM03, OR94, OWS98, PZ12, PA07, PR01, Prz00, RC98, RC92, SW04b, Sha93b, SM01b, SKCB95a, SKCB95b, SK99, TCOR05, TN96, Top96, Use96, VBL94, VED96, WXCM99, WC02a, Wu99a, Wu99b, XTH09, Xue99b, Xue99a, YSZ11, You97, Zha99, vB97]. Wavelet-Based [AR08, Dav95b, HHLW20, LLA08, MWC01, Raj20, Liu07, Prz98, SMC99, Wil00, Wil02, AMS^{+}03, CDGO02, Dav95a, FC98, FJ96, MAC^{+}04, OR94, PA07, PR01]. wavelet-BEM [XTH09]. Wavelet-Domain [CA06, KNB10]. wavelet-packet [HV06b]. Wavelet/Scalar [BBH93].

Wavelets [Ano95, BDHJ04, BBHB94, CMP94, Dau88, JS99b, LF01, Mul96, Pig99, SDS95, SDS96, SN96b, SS96, VK95, AC14, BG03a, CW03, DJL91, GZS^{+}96, KKT10, Mal00, Tro96, TMM96, WLZW04, ZASB95]. WavPack [Wav06]. Way [BS88a, LTO11, Nix81, TOL11, SLZ91]. Weather [Bar00b]. web [HBSASA19, PD16, ABA14, AM01, BY00, BV04b, BV05, CMK^{+}16, HPZ11, Kom96, MKNG06, Mia99, Pie03, RSWW01, RSWW02, SY01]. web-based [PD16]. Web-data [BY00]. Weber [KHCS08]. WebGraph [AM10, BV04a, BV04b]. website [MPE00]. Wei [RHC87]. Weighted [CK93a, CK94b, KØJ^{+}12, LdV95, ZXM^{+}10, ARR01, AFHU97, AECG93, CVC95, CK93b, ECG94, Eff97, Haf96, JJM18, JdVL03, KC06, OY06, PGOO94, Wan95, WG94]. Weighting [ES98, OHSS12, Vol97, Vol02, WST95, AS97, DHH04, Eks96, SOI00, TVW97, Von04]. Weiner [Lev47]. Welch [ACP04, Apo05, Sav97]. Welcome [Ros06]. Well [Koh72, Vov06]. Well-calibrated [Vov06]. Well-Tested [Koh72]. WENO [ABM10]. WES [MC13]. WFA [CK94a, CKV97, CVK97, Haf95]. WG [vL94]. whale [Sta09]. Whatever [VN90b]. wheel [AP03]. Wheeler [Neu10, Abe05, Abe07a, Abe10, ABM08, BKS99, BK00, Bar05, CT98a, Cha00, CJDL09, CHSS08, Deo00, Deo02, EV14, Eff99, FTL03, Ghi05, KLV07, KM99b, Kül12, MRS05, Man01, NSA06, NT20, NZ06, Sad98a, SI99, Sad99, WM01, ZMAB03]. Wheeler-based [KLV07]. White [LaJR81, Mor76]. Whole [LN17]. wide [BSF16, JKV15, BV04b, BV05]. wide-area [JKV15]. Widely [LSC10]. Width [CC01]. WinAce [Win03]. Window [FOF11, KKS00, KU95, Ben96, Gra99, RSV^{+}00, TC06a, Yu95, YG03]. Window-Based [KU95]. windowed [BDV00]. Windows [Ano08j, Lee00, Ano09f, FG89, Hor06, Knu98, Swa93]. Winning [Wan95]. Winning-weighted [Wan95]. WinRAR [Ros06]. Wins [Bar00a]. Winter [USE93]. WinZips [Per21]. Wireless [AH14, DY12, KY08, KY10, KD18, LL16, MW10, MKK12, MCL12, NYJ08, NYJ09, RLG09, SAR00, SF11, Sar12, TCN^{+}17, VJB^{+}95, WHB16, XJ14, ZYT^{+}15, AJRK98, CC00, CCA02, CZZR19, HKA06, Hes08, HM07, JB95, KLJ12, KH04, LKA02, LLWC08, PKG08, QJRA99, RBD13, SF05, SKLA12, XWDY06, YA13, CGSS15]. Within [SS07, VK17]. Without [HHLW20, HO19, PZZ^{+}22, PWS10, Fen12, Hal94, LLCC11, Mal00, OW03a, OSMY95, OG12, Rez98, WM01, Wri39]. Wolf [Cen09, CJDL09, Cm10, CLME04, CME06, FWS^{+}08, FSL^{+}12, KTMS10, LLN^{+}04, LTB04, LCLX04, SRP04, SLXG04, YQ05]. Word [AMPF09, AM05, BKR97a, DPS99, FNP08, Fre93, Gad91, HC92, Jez16, Lia84, MPASM^{+}09, NC97, PR98, BKR^{+}95, Coo05, Cus90, LY91, MSYY98, SNZBY00, YS07, YMYS99]. Word-Aligned [AM05]. Word-Based [FNP08, HC92, DPS99, MSYY98, YMYS99]. Word-Codeword [AMPF09]. wordlength [ABH08]. Words [Gir99, Rez11b, AdlF06, Ano09d, JMnRS19, MRS05, Wri39]. Work [CAL03, De 00a, BBH97, De 03, Ham93, MLP99b]. work-efficient [MLP99b]. Work-Optimal [CAL03, De 00a, De 03]. Workflows [BKRSR09]. workload [XCK18]. workloads [NCB15]. Works [Dis88].

Workshop [Auc98, PBPT95, vL94]. workstation [UH96]. Workstations [Won85]. world [LF04, BV04b, BV05]. World-Wide [BV04b]. Worst [BGPW96a, FV20]. Worst-Case [BGPW96a, FV20]. Worth [VN90a]. write [YG02]. Writing [YTY12]. written [BFMX03]. WSN [JAC19]. WSQ [Fed93]. Wuppertal [AFL96]. WVSN [LGM21]. Wyndham [IEE88a]. Wyner [ARZG03, CX04, DVS^{+}14, EC09, JA03, JFC12, LCLX04, RMG05, SA03, SWW12, WWY^{+}07]. WZT [KLAH99].

X [R^{+}07, Fet96, Gai98, KN11, LKRR91]. X-Masking [KN11]. X-ray [LKRR91]. X10 [CMK^{+}16]. X9.32 [Ano06c]. X9.32-2006 [Ano06c]. Xampling [ME10]. XBM [Mia99]. XLIII [Gol06]. XMill [LS99, XMi03]. XML [Che01, Che06, HS06, HAY06, Kro00, LE07, LDM05, LS99, MPC03, NHHS02, XMi03, YYO04]. XPAND [MK06]. XPath [ACM^{+}15]. XPRESS [MPC03]. XR [Ric10, Ric11, TCKM10]. XR/HD [TCKM10]. XT [RAE16]. ’XX [SCxx]. Xz [Dia16].

Year [MF98c]. years [Ver98]. yield [GL03a]. YK [BY00]. Yokohama [Ano03]. Yuval [Hoc95].

Z [TDG^{+}19]. Z-checker [TDG^{+}19]. z13 [SLJ^{+}15]. z15 [BWB^{+}20, KMA^{+}20]. Zador [GL01]. Zdelta [TMS02]. Zero [ABP09, Cha89, Sha11, TKRR02, WB91, CT95, GLL04, KLAH99, LMO97]. Zero-Crossing [Cha89]. zerodelay [GLL04]. Zero-Error [Sha11, TKRR02]. Zero-Frequency [WB91]. zero-trees [LMO97]. zeroblocks [Hsi01]. Zerotree [KA99, BW95, Cre96, FJ96, HB96b, IA00, RC98]. Zerotrees [Sha93a, Fow98, KF96, Sha93b]. ZFP [DFH^{+}19, FDH^{+}20, TDL^{+}19]. Ziegler [CBK96]. Zip [NN97]. Zipf [Ano07i, Bak07]. Ziv [AHCI^{+}12, Apo05, Sav97, ARZG03, AJ96a, ANdlF04, ANdlF07, ACP04, AN10, ANS12, ACF^{+}21, AF18, BK93a, BFG09, BEGV18, BES05, CPS08, CX04, CI07, CIS08, De 11, DVS^{+}14, DJSS14, EC09, FT95, FT98, Fen93, Fen95, FOF09, FIKS18, GKPR96, GVYZ97, HPZ11, Hor95, II02, JS95, JST01, JA03, JMnRS19, Jen99, JFC12, KS96a, KS98, KKP16a, KKP16b, KY99, KW01, KS00b, Köp21, KM97a, KM99a, KVP20, Lan83a, LCLX04, Log04, LS95b, LS96, LST99, MW85, Mun91, NR99, Nav01, Nav02, NT05a, Per21, PFP99, PR98, RMG05, RS00, RSV^{+}00, RL99, Ryt02, Ryt03, SA03, SWW12, Tho91b, WWY^{+}07, WO00, Wil91b, WZ91b]. Ziv-78 [AF18]. Ziv-based [AN10]. Ziv-Like [KVP20]. ZLIB [DG96, KGG^{+}14, lGAR03a, lGAR03b]. Zonal [CHB86, SS00]. Zone [DSG12]. Zopfli [AKSV15]. Zstandard [Ano16]. Zur [Haa10, Haa12].

References

Anderson:1987:BIM

[A^{+}87] K. L. Anderson et al. Binary-image-manipulation algorithm in the image view facility. IBM Journal of Research and Development, 31(1):16–31, January 1987. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Amidon:1971:ASB

[AA71] E. L. Amidon and G. S. Akin. Algorithmic selection of the best method for compressing map data strings. Communications of the ACM, 14(12):769–774, December 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Alistair:2001:U

[AA01] ?. Alistair and ?. Abbott. [unknown], 2001. URL ftp://ftp.

pd.uwa.edu.au/pub/Wavelets/.

Agarwal:2004:AGA

[AAJ04] R. C. Agarwal, S. Amalapurapu, and S. Jain. An approximation to the greedy algorithm for differential compression of very large files. In Storer and Cohn [SC04], page ?? ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1281499. IEEE Computer Society

Order Number: P2082.

Akrour:2021:FHI

[AAL^{+}21] Leila Akrour, Soltane Ameur, Mourad Lahdir, Régis Fournier, and Amine Nait Ali. Fast hyperspectral image encoder based on supervised multimodal scheme. International Journal of Image and Graphics (IJIG), 21(01):??, January 2021. ISSN

0219-4678. URL https://www.worldscientific.com/doi/
10.1142/S0219467821500078.

Amir:2012:QDP

[AALR12] Amihood Amir, Yonatan Aumann, Avivit Levy, and Yuri Roshko. Quasi-distinct parsing and optimal compression methods. Theoretical Computer Science, 422(1):1–14, March 9, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-

2294 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304397511009418.

Amelio:2011:ICM

[AAR11] A. Amelio, A. Apostolico, and S. E. Rombo. Image compression by 2D Motif basis. In Storer and Marcellin [SM11b], pages 153–162. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=5749473. IEEE

Computer Society Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Al-Araj:1995:RIV

[AAS95] A. G. Al-Araj and K. Sayood. Recursively indexed vector quantization of non-stationary sources. In Storer and Cohn [SC95], page 450. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

515560. IEEE catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Amir:1992:ETD

[AB92] A. Amir and C. Benson. Efficient two-dimensional compressed matching. In Storer and Cohn [SC92], pages 279– 288. ISBN 0-8186-2717-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=227453. IEEE catalog

number 91TH0436-6.

Arnold:1997:CEL

[AB97] R. Arnold and T. Bell. A corpus for the evaluation of lossless compression algorithms. In Storer and Cohn [SC97], pages 201–210. ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1997. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=582019. IEEE

Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Al-Bahadili:2008:NLD

[AB08] Hussein Al-Bahadili. A novel lossless data compression scheme based on the error correcting Hamming codes. Computers and Mathematics with Applications, 56(1):143–150, July 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-

7668 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0898122107008334.

AlZahir:2010:LCM

[AB10] S. AlZahir and A. Borici. Lossless compression of maps, charts, and graphs via color separation. In Storer and Marcellin [SM10b], page 518. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=
5453472.

Abbas:2014:FDC

[ABA14] Ahmed Mohammed Abbas, Azuraliza Abu Bakar, and Mohd Zakree Ahmad. Fast dynamic clustering SOAP messages based compression and aggregation model for enhanced performance of Web services. Journal of Network and Computer Applications, 41(??):80–88, May 2014. CO-DEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804513002154.

Abbas:1993:ANN

[Abb93] Hazem M. Abbas. Artificial neural networks models for image data compression. Ph.D. thesis, Queen’s University at Kingston, Kingston, ON, Canada, 1993. 162 pp. URL http:

//search.proquest.com/docview/304058068.

Aslam:2010:MQS

[ABB10] S. Aslam, A. Bobick, and C. Barnes. Modeling the quantization staircase function. In Storer and Marcellin [SM10b], page 520. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=5453474.

Aberg:1999:TAS

[Abe99] J. Aberg. On taking advantage of similarities between parameters in lossless sequential coding. In Storer and Cohn [SC99], page ?? ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1999. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=785670. IEEE

Computer Society Order Number PR00096. IEEE Order Plan Catalog Number PR00096.

Abel:2004:RPD

[Abe04] J. Abel. Record preprocessing for data compression. In Storer and Cohn [SC04], page ?? ISBN 0-7695-2082-

0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1281497. IEEE Computer Society Order Number: P2082.

Abel:2005:FEP

[Abe05] J. Abel. A fast and efficient post BWT-stage for the Burrows– Wheeler compression algorithm. In Storer and Cohn [SC05], page ?? ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=1402206.

Abel:2007:IFC

[Abe07a] Jürgen Abel. Incremental frequency count — a post BWTstage for the Burrows–Wheeler compression algorithm. Software — Practice and Experience, 37(3):247–265, March 2007. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Abel:2007:U

[Abe07b] Jürgen Abel. [unknown], 2007. URL http://www.data-

compression.info/JuergenAbel/Preprints/;   http://
www.data-compression.info/JuergenAbel/Preprints/Preprint_
After_BWT_Stages.pdf.

Abel:2010:PBS

[Abe10] Jürgen Abel. Post BWT stages of the Burrows–Wheeler compression algorithm. Software — Practice and Experience, 40 (9):751–777, August 2010. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Anderson:1991:VHS

[ABF^{+}91] R. Anderson, J. Bowers, W.-C. Fang, D. Johnson, J.-J. Lee, and R. Nixon. A very high speed noiseless data compression chip for space imaging applications. In Storer and Reif [SR91b], page ?? ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1990. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=213299. IEEE

Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Amir:1994:OTD

[ABF94] Amihood Amir, Gary Benson, and Martin Farach. Optimal two-dimensional compressed matching. In Automata, languages and programming (Jerusalem, 1994), volume 820 of Lecture Notes in Comput. Sci., pages 215–226. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994.

Ajtai:2002:CEU

[ABF^{+}02] Miklos Ajtai, Randal Burns, Ronald Fagin, Darrell D. E. Long, and Larry Stockmeyer. Compactly encoding unstructured inputs with differential compression. Journal of the ACM, 49 (3):318–367, May 2002. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Azhar:1994:DCT

[ABG^{+}94] S. Azhar, G. J. Badros, A. Glodjo, M.-Y. Kao, and J. H. Reif. Data compression techniques for stock market prediction. In Storer and Cohn [SC94], pages 72–82. ISBN 0-8186-5636-0, 0-8186-5637-9 (case). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1994. URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

305914. IEEE Catalog Number 93TH0626-2. IEEE Computer Society Press Order Number 5637-02.

Al-Bahadili:2008:ACW

[ABH08] Hussein Al-Bahadili and Shakir M. Hussain. An adaptive character wordlength algorithm for data compression. Computers and Mathematics with Applications, 55(6):1250–1256, March 2008. CODEN CMAPDK. ISSN 0898-1221 (print),

1873-7668 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0898122107005925.

Apostolico:1999:LGD

[ABL99] A. Apostolico, M. E. Bock, and S. Lonardi. Linear global detectors of redundant and rare substrings. In Storer and Cohn [SC99], pages 168–177. ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1999. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=755666. IEEE

Computer Society Order Number PR00096. IEEE Order Plan Catalog Number PR00096.

Adjeroh:2008:BWT

[ABM08] Donald Adjeroh, Tim Bell, and Amar Mukherjee. The Burrows–Wheeler Transform: Data Compression, Suffix Arrays, and Pattern Matching. Springer Science+Business Media, LLC, Boston, MA, USA, 2008. ISBN 0-387-78908-1, 0-387-78909-X (e-book). 300 pp. LCCN ????

Arandiga:2010:PVW

[ABM10] F. Aràndiga, A. M. Belda, and P. Mulet. Point-value WENO multiresolution applications to stable image compression. Journal of Scientific Computing, 43(2):158–182, May 2010. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL http://link.springer.com/article/
10.1007/s10915-010-9351-8;   http://link.springer.
com/content/pdf/10.1007/s10915-010-9351-8;   http:/
/www.springerlink.com/openurl.asp?genre=article&issn=
0885-7474&volume=43&issue=2&spage=158-182.

Antonini:1990:ICU

[ABMD90] Marc Antonini, Michel Barlaud, Pierre Mathieu, and Ingrid Daubechies. Image coding using vector quantization in the wavelet transform domain. In International Conference on Acoustics, Speech, and Signal Processing, volume 4, pages 2297–2300. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, April 1990. ISSN 1520-6149 (print), 2379-190X (electronic).

Antonini:1992:ICU

[ABMD92] Marc Antonini, Michel Barlaud, Pierre Mathieu, and Ingrid Daubechies. Image coding using wavelet transform. IEEE Transactions on Image Processing, 1(2):205–220, April 1992. CODEN IIPRE4. ISSN 1057-7149 (print), 1941-0042

(electronic). URL http://ieeexplore.ieee.org/abstract/
document/136597/.

Abousleman:1995:CHI

[Abo95] G. P. Abousleman. Compression of hyperspectral imagery using hybrid DPCM/DCT and entropy-constrained trellis coded quantization. In Storer and Cohn [SC95], pages 322–331.

ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ????  URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=515522. IEEE catalog

number 95TH8037. IEEE Computer Society Press order number PR07010.

Abousleman:2006:CHI

[Abo06] Glen P. Abousleman. Coding of hyperspectral imagery with trellis-coded quantization. In G. Motta, F. Rizzo, and J. A. Storer, editors, Hyperspectral Data Compression, page ?? Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006.

Agulhari:2009:ZPS

[ABP09] C. M. Agulhari, I. S. Bonatti, and P. L. D. Peres. A zero padding SVD encoder to compress electrocardiogram. In Storer and Marcellin [SM09], page 434. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=4976488. IEEE Computer Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

Abramson:1963:ITC

[Abr63] N. Abramson. Information Theory and Coding. McGraw-Hill, New York, NY, USA, 1963. ?? pp.

Abrahamson:1989:ADS

[Abr89] David M. Abrahamson. An adaptive dependency source model for data compression. Communications of the ACM, 32(1): 77–83, January 1989. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.org/

pubs/toc/Abstracts/0001-0782/63243.html.

Abrahams:1993:CMC

[Abr93] J. Abrahams. Codes with monotonic codeword lengths. In Storer and Cohn [SC93], pages 52–59. ISBN 0-8186-3391-3 (microfiche), 0-8186-3392-1 (casebound). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1993. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=253145. IEEE Computer Society Press order number 3392-02. IEEE catalog number 93TH0536-3.

Abrahams:1994:HTC

[Abr94] J. Abrahams. Huffman-type codes for infinite source distributions. In Storer and Cohn [SC94], pages 83–89. ISBN 0-8186-5636-0, 0-8186-5637-9 (case). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1994.

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=305915. IEEE Catalog Number 93TH0626-2. IEEE Computer Society Press Order Number 5637-02.

Al-Bahadili:2010:BLT

[ABR10] H. Al-Bahadili and A. Rababa’a. A bit-level text compression scheme based on the HCDC algorithm. International Journal of Computer Applications, 32(3):355–361, 2010. ISSN 1206-212X (print), 1925-7074 (electronic). URL https:

//www.tandfonline.com/doi/full/10.2316/Journal.202.
2010.3.202-2914.

Atya:2016:PAE

[ABS^{+}16] Ahmed Osama Fathy Atya, Ioannis Broustis, Shailendra Singh, Dimitris Syrivelis, Srikanth V. Krishnamurthy, and Thomas F. La Porta. A policy-aware enforcement logic for appropriately invoking network coding. IEEE/ACM Transactions on Networking, 24(4):2005–2018, August 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Arps:1998:PMF

[AC98] R. Arps and C. Constantinescu. PRECIS: a method for fast compression of periodic halftones. In Storer and Cohn [SC98], page ?? ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1998. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=672234. IEEE

Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Apostolico:2006:TCC

[AC06] A. Apostolico and Yong Wook Choi. Textual compression by collapsible tries. In Storer and Cohn [SC06], page 437. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1607280.

Andriani:2007:LCC

[AC07] S. Andriani and G. Calvagno. Lossless compression of colour video sequence using optimal prediction theory — Octopus. In Storer and Cohn [SC07], page 375. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=4148776. IEEE Computer Society Order Number P2791.

Apostolico:2009:PRP

[AC09] A. Apostolico and F. Cunial. Probing the randomness of proteins by their subsequence composition. In Storer and Marcellin [SM09], pages 173–182. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=4976461. IEEE Computer Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

Apostolico:2011:SSG

[AC11] A. Apostolico and F. Cunial. Sequence similarity by gapped LZW. In Storer and Marcellin [SM11b], pages 343–352. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5749492. IEEE Computer Society

Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Arvinti:2014:PDM

[AC14] B. Arvinti and M. Costache. The performance of the Daubechies mother wavelets on ECG compression. In 2014 11th International Symposium on Electronics and Telecommunications (ISETC). IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, November 2014.

Arandiga:2002:AHF

[ACD02] F. Arandiga, G. Chiavassa, and R. Donat. Applications of Harten’s framework for multiresolution: From conservation laws to image compression. In Barth et al. [BCH02], pages 281–296. CODEN LNCSA6. ISBN 3-540-42420-2 (print), 3-642-56205-1 (e-book). ISSN 1439-7358. LCCN QA403.3 .M85

2002. URL http://link.springer.com/content/pdf/10.

1007/978-3-642-56205-1_6. Papers from the Yosemite Educational Symposium (YES), Yosemite valley, California, Fall 2000.

Arroyuelo:2021:EPL

[ACF^{+}21] Diego Arroyuelo, Rodrigo Cánovas, Johannes Fischer, Dominik Köppl, Marvin Löbel, Gonzalo Navarro, and Rajeev Raman. Engineering practical Lempel–Ziv tries. ACM Journal of Experimental Algorithmics, 26(1):14:1–14:47, December 2021. CODEN ???? ISSN 1084-6654. URL https://dl.acm.org/

doi/10.1145/3481638.

Achterberg:1976:BCD

[Ach76] Donald Dean Achterberg. Block Code Data Compression: a Rate Distortion Theoretic Approach. Ph.D. thesis, University of Southern California, Los Angeles, CA, USA, 1976. ???? pp.

URL http://search.proquest.com/docview/302800779.

Acharya:1994:VAA

[Ach94] Tinku Acharya. VLSI Algorithms and Architectures for Data Compression. Ph.D. thesis, University of Central Florida, Orlando, FL, USA, 1994. 145 pp. URL http://search.

proquest.com/docview/304145977.

Apostolico:2008:TCR

[ACK08] A. Apostolico, F. Cunial, and V. Kaul. Table compression by record intersections. In Storer and Marcellin [SM08], pages 13–22. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33

INTERNET. URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4483279.

ACM:1978:CRT

[ACM78] ACM, editor. Conference record of the tenth annual ACM Symposium on Theory of Computing: papers presented at the Symposium, San Diego, California, May 1–3, 1978. ACM Press, New York, NY 10036, USA, 1978. LCCN QA76.6.A13 1978.

ACM:1979:CRE

[ACM79] ACM, editor. Conference record of the eleventh annual ACM Symposium on Theory of Computing: papers presented at the Symposium, Atlanta, Georgia, April 30–May 2, 1979. ACM Press, New York, NY 10036, USA, 1979. LCCN QA75.5.A14

ACM:1995:PTS

[ACM95] ACM, editor. Proceedings of the twenty-seventh annual ACM Symposium on Theory of Computing: Las Vegas, Nevada, May 29–June 1, 1995. ACM Press, New York, NY 10036, USA, 1995. ISBN 0-89791-718-9. LCCN QA 76.6 A13 1995. ACM order no. 508950.

ACM:1997:PEA

[ACM97] ACM, editor. Proceedings of the Eighth Annual ACM-SIAM Symposium on Discrete Algorithms, New Orleans, Louisiana, January 5–7, 1997. ACM Press, New York, NY 10036, USA, 1997. CODEN PAAAF2. ISBN 0-89871-390-

0. LCCN ???? URL http://www.acm.org/pubs/contents/
proceedings/soda/314161/.

ACM:2003:PAS

[ACM03] ACM, editor. Proceedings of the 2003 ACM SIGMOD International Conference on Management of Data 2003, San Diego, California, June 09–12, 2003. ACM Press, New York, NY 10036, USA, 2003. ISBN 1-58113-634-X. LCCN ????

ACM:2013:SPF

[ACM13] ACM, editor. STOC ’13: Proceedings of the Forty-fifth Annual ACM Symposium on Theory of Computing: June 1–4, 2013, Palo Alto, California, USA. ACM Press, New York, NY 10036, USA, 2013. ISBN 1-4503-2029-5.

Arroyuelo:2015:FMX

[ACM^{+}15] Diego Arroyuelo, Francisco Claude, Sebastian Maneth, Veli Mäkinen, Gonzalo Navarro, Kim Nguy˜ ên, Jouni Sirén, and Niko Välimäki. Fast in-memory XPath search using compressed indexes. Software — Practice and Experience, 45(3): 399–434, March 2015. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Apostolico:2004:MZL

[ACP04] A. Apostolico, M. Comin, and L. Parida. Motifs in Ziv– Lempel-Welch Clef. In Storer and Cohn [SC04], pages 72–81. ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1281452. IEEE Computer

Society Order Number: P2082.

Apostolico:2005:LCE

[ACP05] A. Apostolico, M. Comin, and L. Parida. Off-line compression by extensible motifs. In Storer and Cohn [SC05], page ?? ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1402207.

Apostolico:2009:GCB

[AD09] Alberto Apostolico and Guido Drovandi. Graph compression by BFS. Algorithms (Basel), 2(3):1031–1044, September 2009. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https:

//www.mdpi.com/1999-4893/2/3/1031.

Altingovde:2008:ICB

[ADCU08] Ismail Sengor Altingovde, Engin Demir, Fazli Can, and Özgür Ulusoy. Incremental cluster-based retrieval using compressed cluster-skipping inverted files. ACM Transactions on Information Systems, 26(3):15:1–15:??, June 2008. CODEN ATISET. ISSN 1046-8188.

Adjeroh:2006:OCR

[Adj06] D. A. Adjeroh. Optimal coding rate selection for 3D video using RCPC codes. In Storer and Cohn [SC06], page 436. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1607279.

Adiego:2006:UWS

[AdlF06] J. Adiego and P. de la Fuente. On the use of words as source alphabet symbols in PPM. In Storer and Cohn [SC06], page 435. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=1607278.

Adiego:2004:MPP

[AdlPN04] J. Adiego, P. de la Puente, and G. Navarro. Merging prediction by partial matching with structural contexts model. In Storer and Cohn [SC04], page ?? ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1281498. IEEE Computer Society Order Number: P2082.

Systems:2001:APD

[Ado01] Adobe Systems. Adobe Portable Document Format Version 1.4. Addison-Wesley, Reading, MA, USA, third edition, December 2001. ?? pp.

Adobe:2006:AAF

[Ado06] Adobe Systems Incorporated. Adobe Acrobat family / about

Adobe PDF, 2006. URL http://www.adobe.com/products/
acrobat/adobepdf.html.

Andrews:1993:MRV

[AECG93] B. D. Andrews, M. Effros, P. A. Chou, and R. M. Gray. A mean-removed variation of weighted universal vector quantization for image coding. In Storer and Cohn [SC93], pages 302– 309. ISBN 0-8186-3391-3 (microfiche), 0-8186-3392-1 (casebound). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 1993. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=253119. IEEE Computer

Society Press order number 3392-02. IEEE catalog number 93TH0536-3.

Apostolico:1987:RTU

[AF87] Alberto Apostolico and A. S. Fraenkel. Robust transmission of unbounded strings Using Fibonacci representations. IEEE Transactions on Information Theory, 33(2):238–245, March 1987. CODEN IETTAW. ISSN 0018-9448 (print), 1557-9654 (electronic).

Anedda:1988:CQI

[AF88a] C. Anedda and L. Felician. P -compressed quadtrees for image storing. The Computer Journal, 31(4):353–357, 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Anedda:1988:PCQ

[AF88b]  C. Anedda and L. Felician.  P-compressed quadtrees

for image storing. The Computer Journal, 31(4):353– 357, August 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.

co.uk/computer_journal/hdb/Volume_31/Issue_04/tiff/
353.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_04/tiff/354.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_31/Issue_04/

tiff/355.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_31/Issue_04/tiff/356.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_31/Issue_04/
tiff/357.tif.

Ageenko:1999:FAM

[AF99] Eugene I. Ageenko and Pasi Fränti. Forward-adaptive method for context-based compression of large binary images. Software — Practice and Experience, 29(11): 943–952, September 1999. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http:

//www3.interscience.wiley.com/cgi-bin/abstract?ID=
63501201;   http://www3.interscience.wiley.com/cgi-
bin/fulltext?ID=63501201&PLACEBO=IE.pdf.

Aghito:2004:CBC

[AF04] S. M. Aghito and S. Forchhammer. Context based coding of binary shapes by object boundary straightness analysis. In Storer and Cohn [SC04], pages 399–408. ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1281485. IEEE Computer Society Order Number: P2082.

Arz:2018:LZC

[AF18] Julian Arz and Johannes Fischer. Lempel–Ziv-78 compressed string dictionaries. Algorithmica, 80(7):2012–2047, July 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Alakuijala:2019:BGP

[AFF^{+}19] Jyrki Alakuijala, Andrea Farruggia, Paolo Ferragina, Eugene Kliuchnikov, Robert Obryk, Zoltan Szabadka, and Lode Vandevenne. Brotli: a general-purpose data compressor. ACM Transactions on Information Systems, 37(1):4:1–4:??, January 2019. CODEN ATISET. ISSN 1046-8188.

Albert:1997:VCW

[AFHU97] J. Albert, S. Frank, U. Hafner, and M. Unger. Video compression with weighted finite automata. In Storer and Cohn [SC97], page ?? ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1997. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=582071. IEEE

Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Alefeld:1996:SCV

[AFL96] Götz Alefeld, Andreas Frommer, and Bruno Lang, editors. Scientific computing and validated numerics: proceedings of the International Symposium on Scientific Computing, Computer Arithmetic and Validated Numerics SCAN-95 held in Wuppertal, Germany, September 26–29, 1995, volume 90 of Mathematical Research. Akademie Verlag, Berlin, Germany, 1996. ISBN 3-05-501737-4. ISSN 0138-3019. LCCN QA76.95 .I575 1995.

Anshel:1986:MIP

[AG86] Michael Anshel and William Gewirtz, editors. Mathematics of information processing, volume 34 of Proceedings of symposia in applied mathematics, 0160-7634; AMS short course lecture notes. American Mathematical Society, Providence, RI, USA, 1986. ISBN 0-8218-0086-8. xi + 233 pp. LCCN QA76.9.M35 M39 1986.

Aaron:2002:CSI

[AG02] A. Aaron and B. Girod. Compression with side information using turbo codes. In Storer and Cohn [SC02], pages 252– 261. ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2002. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=999963. IEEE

Computer Society Order Number PR01477.

Agrawal:2009:DDM

[AG09] N. Agrawal and A. Gupta. DCT domain message embedding in spread-spectrum steganography system. In Storer and Marcellin [SM09], page 433. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ????

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=4976487. IEEE Computer Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

Agarwal:2006:AGA

[AGJA06] R. C. Agarwal, K. Gupta, S. Jain, and S. Amalapurapu. An approximation to the greedy algorithm for differential compression. IBM Journal of Research and Development, 50(1): 149–166, January 2006. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http://www.research.

ibm.com/journal/rd/501/agarwal.html.

Andoni:2013:HFU

[AGMP13] Alexandr Andoni, Assaf Goldberger, Andrew McGregor, and Ely Porat. Homomorphic fingerprints under misalignments: sketching edit and shift distances. In ACM [ACM13], pages 931–940. ISBN 1-4503-2029-5.

Atallah:1996:PMI

[AGS96] M. Atallah, Y. Genin, and W. Szpankowski. Pattern matching image compression. In Storer and Cohn [SC96], page ?? ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1996. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=488349. IEEE Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Akansu:1992:MSD

[AH92] Ali Akansu and R. Haddad. Multiresolution Signal Decomposition. Academic Press, New York, USA, 1992. ?? pp.

Ahn:2002:GCD

[AH02] Jeong-Hwan Ahn and Yo-Sung Ho. Geometry compression of 3-D mesh models using a joint prediction. In Storer and Cohn [SC02], page ?? ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2002. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=999989. IEEE Computer Society Order Number PR01477.

Anagnostopoulos:2014:APC

[AH14] Christos Anagnostopoulos and Stathes Hadjiefthymiades. Advanced principal component-based compression schemes for wireless sensor networks. ACM Transactions on Sensor Networks, 11(1):7:1–7:??, August 2014. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Al-Hafeedh:2012:CIB

[AHCI^{+}12] Anisa Al-Hafeedh, Maxime Crochemore, Lucian Ilie, Evguenia Kopylova, W. F. Smyth, German Tischler, and Munina Yusufu. A comparison of index-based Lempel–Ziv LZ77 factorization algorithms. ACM Computing Surveys, 45(1):5:1–5:??, November 2012. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

Ahlswede:1997:UCI

[AHK97] Rudolf Ahlswede, Te Sun Han, and Kingo Kobayashi. Universal coding of integers and unbounded search trees. IEEE Transactions on Information Theory, 43(2):669–682, 1997. CODEN IETTAW. ISSN 0018-9448 (print), 1557-9654 (electronic).

Aitken:1986:PMD

[Ait86] Robert Campbell Aitken. Performance measures of data compression techniques in fault detection. Thesis (M.Sc.), University of Victoria, Victoria, BC, Canada, 1986. ???? pp.

Acharya:1995:ELC

[AJ95] Tinku Acharya and Joseph F. JáJá. Enhancing LZW coding using a variable-length binary encoding. Technical Report TR 1995-70, Institute for Systems Research, University of Maryland, College Park, MD, USA, 1995. ?? pp.

Acharya:1996:ELZ

[AJ96a] T. Acharya and J. F. JaJa. Enhancing Lempel–Ziv codes using an on-line variable length binary encoding. In Storer and Cohn [SC96], page ?? ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1996. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=488347. IEEE

Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Acharya:1996:LVL

[AJ96b] Tinku Acharya and Joseph F. JáJá. An on-line variable length binary encoding of text. Information Sciences, 94:1–22, 1996. CODEN ISIJBC. ISSN 0020-0255 (print), 1872-6291 (electronic).

Agnihotri:2008:IDS

[AJN08] S. Agnihotri, H. S. Jamadagni, and P. Nuggehalli. Interactive distributed source coding in asymmetric communication scenarios. In Storer and Marcellin [SM08], page 503.

ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33 INTER-

NET. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=4483330.

Appadwedula:1998:JSC

[AJRK98] S. Appadwedula, D. L. Jones, K. Ramchandran, and I. Konzintsev. Joint source channel matching for a wireless communications link. In Storer and Cohn [SC98], page ?? ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1998. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=672230. IEEE Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Adams:1999:PER

[AK99] M. D. Adams and F. Kossentini. Performance evaluation of reversible integer-to-integer wavelet transforms for image compression. In Storer and Cohn [SC99], page ?? ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1999. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=785671. IEEE Computer Society Or-

der Number PR00096. IEEE Order Plan Catalog Number PR00096.

Ahuja:2008:SPM

[AK08] S. Ahuja and M. Krunz. Server placement in multipledescription-based media streaming. In Storer and Marcellin [SM08], pages 372–381. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008;

QA76.9.D33 INTERNET. URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=4483315.

Akimov:2006:LCC

[AKF06] A. Akimov, A. Kolesnikov, and P. Franti. Lossless compression of color map images by context tree modeling. In Storer and Cohn [SC06], pages 10 pp.–421. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=1607276.

Almurib:2018:ADI

[AKL18] Haider A. F. Almurib, Thulasiraman Nandha Kumar, and Fabrizio Lombardi. Approximate DCT image compression using inexact computing. IEEE Transactions on Computers, 67(2): 149–159, ???? 2018. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://ieeexplore.

ieee.org/document/7990539/.

Abedini:2010:SBS

[AKS10] N. Abedini, S. P. Khatri, and S. A. Savari. A SATbased scheme to determine optimal fix-free codes. In Storer and Marcellin [SM10b], pages 169–178. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=5453461.

Alakuijala:2015:CBD

[AKSV15] Jyrki Alakuijala, Evgenii Kliuchnikov, Zoltan Szabadka, and Lode Vandevenne. Comparison of Brotli, Deflate, Zopfli, LZMA, LZHAM and Bzip2 compression algorithms. Report, Google, Inc., September 22, 2015. 6 pp. URL https:

//github.com/google/brotli/blob/master/docs/brotli-
comparison-study-2015-09-22.pdf.

Aleh:1981:DUB

[AL81] Avner Aleh and K. Dan Levin. The determination of upper bounds for economically effective compression in packet switching networks. ACM SIGMETRICS Performance Evaluation Review, 11(1):64–72, Spring 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Apostolico:1998:STP

[AL98] A. Apostolico and S. Lonardi. Some theory and practice of greedy off-line textual substitution. In Storer and Cohn [SC98], pages 119–128. ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1998. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=672138. IEEE

Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Apostolico:2000:CBS

[AL00] A. Apostolico and S. Lonardi. Compression of biological sequences by greedy off-line textual substitution. In Storer and Cohn [SC00], pages 143–152. ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2000. URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

838154. This millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Archer:2001:COT

[AL01] C. Archer and T. K. Leen. The coding-optimal transform. In Storer and Cohn [SC01c], pages 381–390. ISBN 0-7695-1031-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 2001. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=917169. IEEE CSP num-

ber 01PR1031.

Auli-Llinas:2010:LAB

[AL10] F. Aulí-Llinás. Local average-based model of probabilities for JPEG2000 bitplane coder. In Storer and Marcellin [SM10b], pages 59–68. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=
5453434.

Arnavut:1998:BST

[ALA98] Z. Arnavut, D. Leavitt, and M. Abdulazizoglu. Block sorting transformations. In Storer and Cohn [SC98], page ?? ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1998. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=672232. IEEE Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Auli-Llinas:2011:RDO

[ALM11] F. Aulí-Llinás and M. W. Marcellin. Rate-distortion optimized adaptive scanning order for bitplane image coding engines. In Storer and Marcellin [SM11b], pages 163–172. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5749474. IEEE Computer Society

Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Aronica:2018:OPL

[ALM^{+}18] Salvatore Aronica, Alessio Langiu, Francesca Marzi, Salvatore Mazzola, and Filippo Mignosi. On optimal parsing for LZ78-like compressors. Theoretical Computer Science, 710(??):19–28, February 1, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304397517301524.

Auli-Llinas:2012:EQD

[ALMJR^{+}12] F. Aulí-Llinás, M. W. Marcellin, L. Jimenez-Rodriguez, I. Blanes, and J. Serra-Sagristà. Embedded quantizer design for low rate lossy image coding. In Storer and Marcellin [SM12], pages 89–98. ISBN 1-4673-0715-7 (paperback). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=6189240. IEEE Computer Society order number P4656.

Auli-Llinas:2009:HAD

[ALMSS09] F. Aulí-Llinás, M. W. Marcellin, and J. Serra-Sagristà. Highly accurate distortion estimation for JPEG2000 through PDFbased estimators. In Storer and Marcellin [SM09], pages 391– 400. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=4976483. IEEE

Computer Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

Amat:2016:NMA

[ALRT16] S. Amat, J. Liandrat, J. Ruiz, and J. C. Trillo. On a nonlinear mean and its application to image compression using multiresolution schemes. Numerical Algorithms, 71(4):729–752, April 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-015-0019-1.

Auli-Llinas:2006:ERC

[ALSSMPBR06] F. Aulí-Llinás, J. Serra-Sagristà, J. L. Monteagudo-Pereira, and J. Bartrina-R. Efficient rate control for JPEG2000 coder and decoder. In Storer and Cohn [SC06], pages 282–291. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1607263.

An:2012:EDP

[ALYK12] Baik Song An, Manhee Lee, Ki Hwan Yum, and Eun Jung Kim. Efficient data packet compression for cache coherent multiprocessor systems. In Storer and Marcellin [SM12], pages 129–138. ISBN 1-4673-0715-7 (paperback). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ????

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=6189244. IEEE Computer Society order number P4656.

Arya:1993:AFV

[AM93] S. Arya and D. M. Mount. Algorithms for fast vector quantization. In Storer and Cohn [SC93], pages 381–390. ISBN 0-8186-3391-3 (microfiche), 0-8186-3392-1 (casebound). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1993. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=253111. IEEE Computer Society Press order number 3392-02. IEEE catalog number 93TH0536-3.

Acharya:1995:TBB

[AM95] T. Acharya and A. Mukherjee. A tree based binary encoding of text using LZW algorithm. In Storer and Cohn [SC95], page 463. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=515573. IEEE catalog

number 95TH8037. IEEE Computer Society Press order number PR07010.

Albers:1996:ACA

[AM96] S. Albers and M. Mitzenmacher. Average case analyses of list update algorithms, with applications to data compression. Lecture Notes in Computer Science, 1099:514–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arnavut:1997:BSC

[AM97] Z. Arnavut and S. S. Magliveras. Block sorting and compression. In Storer and Cohn [SC97], pages 181–190. ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1997. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=582009. IEEE Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Albers:1998:ACA

[AM98] Susanne Albers and Michael Mitzenmacher. Average case analyses of list update algorithms, with applications to data compression. Algorithmica, 21(3):312–329, July 1998. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=21&issue=3&spage=312.

Adler:2001:TCW

[AM01] M. Adler and M. Mitzenmacher. Towards compressing Web graphs. In Storer and Cohn [SC01c], pages 203–212. ISBN 0-7695-1031-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2001. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=917151. IEEE

CSP number 01PR1031.

Anh:2005:IIC

[AM05] Vo Ngoc Anh and Alistair Moffat. Inverted index compression using word-aligned binary codes. Information Retrieval, 8:151–166, 2005. CODEN IFRTFY. ISSN 1386-4564 (print),

1573-7659 (electronic). URL http://link.springer.com/
article/10.1023/B%3AINRT.0000048490.99518.5c.

Anh:2010:LMW

[AM10] V. N. Anh and A. Moffat. Local modeling for WebGraph compression. In Storer and Marcellin [SM10b], page 519. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=5453475.

Andersson:2013:DDS

[AMAM13] M. Andersson, J. Munkberg, and T. Akenine-Möller. Depth and distance: Stochastic depth buffer compression using generalized plane encoding. Computer Graphics Forum, 32(2pt1): 103–112, May 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Adjeroh:2002:PMB

[AMB^{+}02] D. Adjeroh, A. Mukherjee, T. Bell, M. Powell, and N. Zhang. Pattern matching in BWT-transformed text. In Storer and Cohn [SC02], page ?? ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2002. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=999988. IEEE Computer Society Order Number PR01477.

Azar:2019:EEI

[AMBC19] Joseph Azar, Abdallah Makhoul, Mahmoud Barhamgi, and Raphaël Couturier. An energy efficient IoT data compression approach for edge machine learning. Future Generation Computer Systems, 96(??):168–175, July 2019. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167739X18331716.

AFB:2006:DBT

[Ame06] American Foundation for the Blind. Duxbury Braille Trans-

lator, 2006. URL http://www.afb.org/prodProfile.asp?
ProdID=42.

Amjad:2008:DCP

[Amj08] Hasan Amjad. Data compression for proof replay. Journal of Automated Reasoning, 41(3–4):193–218, November 2008. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670

(electronic). URL http://link.springer.com/article/10.
1007/s10817-008-9109-2.

Abu-Mostafa:2000:MCL

[AMM00] Y. S. Abu-Mostafa and R. J. McEliece. Maximal codeword lengths in Huffman codes. Computers and Mathematics with Applications, 39(11):129–134, June 2000. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S089812210000119X.

Adiego:2009:HPW

[AMPF09] J. Adiego, M. A. Martinez-Prieto, and P. Fuente. High performance word-codeword mapping algorithm on PPM. In Storer and Marcellin [SM09], pages 23–32. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=4976446. IEEE Computer Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

Adiego:2010:MPT

[AMPHTSM10] J. Adiego, M. A. Martinez-Prieto, J. E. Hoyos-Torio, and F. Sanchez-Martinez. Modelling parallel texts for boosting compression. In Storer and Marcellin [SM10b], page 517. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=5453473.

Amsterdam:1986:DCH

[Ams86] Jonathan Amsterdam. Data compression with Huffman coding. Byte Magazine, 11(5):99–108, May 1986. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Alecu:2003:OED

[AMS^{+}03] A. Alecu, A. Munteanu, P. Schelkens, J. Cornelis, and S. Dewitte. On the optimality of embedded deadzone scalarquantizers for wavelet-based L-infinite-constrained image coding. In Storer and Cohn [SC03], page ?? ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 2003. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1194032. IEEE Computer

Society Order number PR01896.

amu:2006:U

[amu06] amu. [unknown], 2006. URL http://ifa.amu.edu.pl/~

kprzemek.

Adjeroh:2006:CPS

[AN06] D. Adjeroh and Fei Nan. On compressibility of protein sequences. In Storer and Cohn [SC06], page 434. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1607277.

Adjeroh:2008:SSS

[AN08] D. Adjeroh and Fei Nan. Suffix sorting via Shannon–Fano– Elias codes. In Storer and Marcellin [SM08], page 502.

ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33 INTER-

NET. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=4483329.

Arroyuelo:2010:PAR

[AN10] Diego Arroyuelo and Gonzalo Navarro. Practical approaches to reduce the space requirement of Lempel–Ziv-based compressed text indices. ACM Journal of Experimental Algorithmics, 15 (1):15:1–15:??, March 2010. ISSN 1084-6654.

Anderson:2007:RVJ

[And07] M. Anderson. Resources: Virtual jamming. IEEE Spectrum, 44(7):53–56, July 2007. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Andujar:2012:ACT

[And12] C. Andujar. Adaptive compression of texture pyramids. Computer Graphics Forum, 31(6):1973–1983, September 2012. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Adiego:2004:LZC

[ANdlF04] J. Adiego, G. Navarro, and P. de la Fuente. Lempel–Ziv compression of structured text. In Storer and Cohn [SC04], pages 112–121. ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=1281456. IEEE

Computer Society Order Number: P2082.

Adiego:2007:LZC

[ANdlF07] Joaquín Adiego, Gonzalo Navarro, and Pablo de la Fuente. Lempel–Ziv compression of highly structured documents. Journal of the American Society for Information Science and Technology: JASIST, 58(4):461–478, February 15, 2007. CO-DEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Anonymous:1972:G

[Ano72] Anonymous. G.711, 1972. URL http://en.wikipedia.org/

wiki/G.711.

Anonymous:1982:DCH

[Ano82] Anonymous. Data compression hardware and software market: Data compression products in the U.S. Frost and Sullivan, 106 Fulton Strett, New York 10038, NY, USA, May 1982. vii + 184 pp. LCCN HD9696.D383 U6316 1982.

Anonymous:1986:CDF

[Ano86a] Anonymous. Compression/decompression of font patterns. IBM Technical Disclosure Bulletin, 28(8):3563–3564, January 1986. CODEN IBMTAA. ISSN 0018-8689.

Anonymous:1986:DOD

[Ano86b] Anonymous. Don’t overlook data compression. Computers & Security, 5(1):8–9, March 1986. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
0167404886901100.

Anonymous:1988:DCTb

[Ano88a] Anonymous. Data compression: Techniques and applications. October 1983–September 1986. Computers & Security, 7(5):518, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
0167404888902611.

Anonymous:1988:DCTc

[Ano88b] Anonymous. Data compression: Techniques and applications. October 1983–September 1986. Computers & Security, 7(5):519, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
0167404888902684.

Anonymous:1988:DCTd

[Ano88c] Anonymous. Data compression: Techniques and applications. October 1986–September 1987. Computers & Security, 7(5):519, October 1988. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
0167404888902696.

Anonymous:1991:DDC

[Ano91] Anonymous. The DDJ data compression contest. Dr. Dobb’s Journal of Software Tools, 16(2):52–??, February 1991. CO-DEN DDJOEB. ISSN 1044-789X.

Anonymous:1992:CCG

[Ano92a] Anonymous. Can compression get you out of a squeeze? Datamation, 38(20):65–??, October 01, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:DYDb

[Ano92b] Anonymous. Double your drives with data compression. Datamation, 38(24):49–??, December 01, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:FDC

[Ano92c] Anonymous. On-the-Fly disk compression: Diet disk compression utility, automated Apple menus, and organizing Usenet news. Byte Magazine, 17(6):357–??, June 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:DCM

[Ano94] Anonymous. Data compression on the Macintosh. Byte Magazine, 19(2):62–??, February 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:WCM

[Ano95] Anonymous. Wavelets challenge MPEG, fractals — developers claim 300-to-1 compression. Byte Magazine, 20(12):34–??, December 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1997:LGH

[Ano97] Anonymous. Linux Gazette: History of the portable network graphics (PNG) format. Linux Journal, 36:??, April 1997. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Anonymous:1998:BRIc

[Ano98] Anonymous. Book review: Information systems and data compression: By Jerzy A. Seidler. Kluwer Academic Publishers, Boston. (1997). 472 pages. $160.00, NLG 335.00, GBP 108.80. Computers and Mathematics with Applications, 35(5):

135, March 1998. CODEN CMAPDK. ISSN 0898-1221 (print),

1873-7668 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0898122198907532.

Anonymous:2000:BRPa

[Ano00] Anonymous. Book reviews: PNG: The Definitive Guide by Michael J. Hammel; The Cathedral and the Bazaar by Peter Salus. Linux Journal, 69:??, January 2000. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http:

//noframes.linuxjournal.com/lj-issues/issue69/3628.
html;   http://noframes.linuxjournal.com/lj-issues/
issue69/3763.html.

Anonymous:2002:SHS

[Ano02] Anonymous. Secure Hash Standard. Technical report, National Institute for Standards and Technology, Gaithersburg, MD, USA, August 1, 2002. iv + 71 pp. URL http://csrc.

nist.gov/publications/fips/fips180-2/fips180-2.pdf.

FIPS Publication 180-2. Officially replaced by [Ano12b], but that newer version is much shorter.

Anonymous:2003:CCV

[Ano03] Anonymous, editor. Cool Chips VI:An International Symposium on Low-Power and High-Speed Chips, Yokohama Joho Bunka Center, Yokohama, Japan (Yokohama Media & Communications Center, Yokohama, Japan) April 16–18, 2003. ????, ????, 2003. ISBN ???? LCCN ????

Anonymous:2004:CPa

[Ano04a] Anonymous. The Capocelli Prize. In Storer and Cohn [SC04], page iv. ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=1281443. IEEE

Computer Society Order Number: P2082.

Anonymous:2004:WSF

[Ano04b] Anonymous. Was that slip and fall for real? [video surveillance and VideoSave]. IEEE Spectrum, 41(5):18, May 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2006:DCAa

[Ano06a] Anonymous. Data compression: audio, 2006. URL http://

en.wikipedia.org/wiki/Audio_data_compression.

Anonymous:2006:DCAb

[Ano06b] Anonymous. Data compression: audio, 2006. URL

http://www.uiciechi.it/vecchio/cnt/schede/windots-
eng.html.

Anonymous:2006:DCF

[Ano06c] Anonymous, editor. Data compression in financial telecommunications: American National Standard for Financial Services: ANSI X9.32-2006. Accredited Standards Committee X9, Inc., Financial Industry Standards, Annapolis, MD, USA, 2006. x + 30 leaves pp. LCCN HG4515.5 .D38 2006. Revision of X9.32-1998. Date approved: July 6, 2006.

Anonymous:2006:DHP

[Ano06d] Anonymous. DonationCoder home page, 2006. URL http://

www.donationcoder.com/Reviews/Archive/ArchiveTools/
index.html.

Anonymous:2006:FDT

[Ano06e] Anonymous. Filter design toolbox: Examples of adaptive filters that use LMS algorithms, 2006. URL http://www.

weizmann.ac.il/matlab/toolbox/filterdesign/adaptiv9.

html. Weizmann Institute of Science, 234 Herzl Street, Rehovot 76100, Israel.

Anonymous:2006:HDS

[Ano06f] Anonymous. H.264 draft standard, 2006. URL ftp:/

/standards.polycom.com/JVT_Site/draft_standard/JVT-
G050r1.zip.

Anonymous:2006:HIT

[Ano06g] Anonymous. The H.264 integer transform, 2006. URL ftp:/

/standards.polycom.com/JVT_Site/2002_01_Geneva/JVT-
B038r2.doc;   ftp://standards.polycom.com/JVT_Site/
2002_01_Geneva/JVT-B039r2.doc.

Anonymous:2006:HSR

[Ano06h] Anonymous. The H.264 standards repository, 2006. URL ftp:

//standards.polycom.com/JVT_Site/.

Anonymous:2006:IBG

[Ano06i] Anonymous. IA-32 (32-bit generation of Intel microprocessor

architecture), 2006. URL http://en.wikipedia.org/wiki/
IA-32.

Anonymous:2006:KBS

[Ano06j] Anonymous. Karlheinz Brandenburg and the secret his-

tory of MP3, 2006. URL http://www.bbc.co.uk/dna/h2g2/
A406973.

Anonymous:2006:LCK

[Ano06k] Anonymous. Levenstein coding [also known as Leven-

shtein coding], 2006. URL http://en.wikipedia.org/wiki/
Levenstein_coding.

Anonymous:2006:MMG

[Ano06l] Anonymous. Mathematical mysteries: the Goldbach conjec-

ture, 2006. URL http://pass.maths.org/issue2/xfile/.

Anonymous:2006:MAF

[Ano06m] Anonymous. Monkey’s Audio: a fast and powerful lossless au-

dio compressor, 2006. URL http://www.monkeysaudio.com/
index.html.

Anonymous:2006:MT

[Ano06n] Anonymous. Morse tape, 2006. URL http://memory.loc.

gov/mss/mmorse/071/071009/0001d.jpg.

Anonymous:2006:OS

[Ano06o] Anonymous. Ogg squish. Web page, 2006. URL http://www.

xiph.org/ogg/flac.html.

Anonymous:2006:PAa

[Ano06p] Anonymous. Proceedings available. In Storer and Cohn [SC06], page ?? ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=1607232.

Anonymous:2006:RSA

[Ano06q] Anonymous. Rarissimo software: Automatically compress/uncompress files preserving streams and Macintosh

files integrity, 2006. URL http://peccatte.karefil.com/
software/Rarissimo/RarissimoEN.htm.

Anonymous:2006:JCI

[Ano06r] Anonymous. JBIG2.com: An introduction to JBIG2: JBIG2

primer, 2006. URL http://jbig2.com/.

Anonymous:2006:Ma

[Ano06s]  Anonymous. mpeg-4.als, 2006. URL http://www.nue.tu-
berlin.de/forschung/projekte/lossless/mp4als.html.

Anonymous:2006:Ua

[Ano06t] Anonymous. [unknown], 2006. URL http://f-cpu.seul.

org/whygee/ddj-3r/ddj-3R.html.

Anonymous:2006:Ub

[Ano06u] Anonymous. [unknown], 2006. URL http://f-cpu.seul.

org/whygee/phasing-in_codes/PhasingInCodes.nb.

Anonymous:2006:USC

[Ano06v] Anonymous. [UNRAR source code], 2006. URL http://www.

rarlab.com/rar/unrarsrc-3.5.4.tar.gz.

Anonymous:2006:UBU

[Ano06w] Anonymous. UTF-16 (16-bit Unicode Transformation For-

mat), 2006. URL http://en.wikipedia.org/wiki/UTF-16.

Anonymous:2007:CP

[Ano07a] Anonymous. The Capocelli Prize. In Storer and Cohn [SC07], page xiii. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=4148738. IEEE

Computer Society Order Number P2791.

Anonymous:2007:DI

[Ano07b] Anonymous. Datacompression.info, 2007. URL http://

datacompression.info/IncredibleClaims.shtml.

Anonymous:2007:FSH

[Ano07c] Anonymous. Facsimile & SSTV history, 2007. URL http://

www.hffax.de/html/hauptteil_faxhistory.htm.

Anonymous:2007:PC

[Ano07d] Anonymous. Program committee. In Storer and Cohn [SC07], page xii. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=4148737. IEEE

Computer Society Order Number P2791.

Anonymous:2007:T

[Ano07e] Anonymous. Telegraphy, 2007. URL http://www.technikum29.

de/en/communication/fax.shtm.

Anonymous:2007:Ua

[Ano07f] Anonymous. [unknown], 2007. URL http://www.codethatword.

com/.

Anonymous:2007:Ub

[Ano07g] Anonymous. [unknown], 2007. URL http://www.datacompression.

info/patents.shtml.

Anonymous:2007:Uc

[Ano07h] Anonymous. [unknown], 2007. URL http://www.gemstartvguide.

com.

Anonymous:2007:ZL

[Ano07i] Anonymous. Zipf’s law, 2007. URL http://en.wikipedia.

org/wiki/Zipf's_law.

Anonymous:2008:CP

[Ano08a] Anonymous. The Capocelli Prize. In Storer and Marcellin [SM08], page vi. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008;

QA76.9.D33 INTERNET. URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=4483276.

Anonymous:2008:CC

[Ano08b] Anonymous. Compression curiosities, 2008. URL http://

www.maximumcompression.com/compression_fun.php.

Anonymous:2008:HP

[Ano08c] Anonymous. Hutter prize, 2008. URL http://en.wikipedia.

org/wiki/Hutter_Prize.

Anonymous:2008:MCL

[Ano08d] Anonymous. Maximum compression: Lossless data compression software benchmarks / comparisons, 2008. URL http://

www.maximumcompression.com.

Anonymous:2008:PSL

[Ano08e] Anonymous. PAQ [a series of lossless data compression

archivers], 2008. URL http://en.wikipedia.org/wiki/PAQ.

Anonymous:2008:KNH

[Ano08f] Anonymous. kgbarchiver.net home page, 2008. URL http:

//kgbarchiver.net/.

Anonymous:2008:SR

[Ano08g] Anonymous. strange.rar, 2008. URL http://www.

maximumcompression.com/strange.rar.

Anonymous:2008:TR

[Ano08h]  Anonymous. test.rar, 2008. URL http://www.maximumcompression.
com/test.rar.

Anonymous:2008:VSV

[Ano08i] Anonymous. VC-1 [SMPTE 421M video codec standard], 2008.

URL http://en.wikipedia.org/wiki/VC-1.

Anonymous:2008:WMC

[Ano08j] Anonymous. Windows media codecs, 2008. URL http://www.

microsoft.com/windows/windowsmedia/forpros/codecs/
video.aspx.

Anonymous:2009:AAS

[Ano09a] Anonymous. ASCII [American Standard Code for Information

Interchange], 2009. URL http://en.wikipedia.org/wiki/
ASCII.

Anonymous:2009:CP

[Ano09b] Anonymous. The Capocelli Prize. In Storer and Marcellin [SM09], page vi. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

4976441. IEEE Computer Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

Anonymous:2009:HSA

[Ano09c] Anonymous. The Haiku Society of America, 2009. URL http:

//www.hsa-haiku.org/index.htm.

Anonymous:2009:LEW

[Ano09d] Anonymous. List of English words containing Q not followed

by U, 2009. URL http://en.wikipedia.org/wiki/List_of_
English_words_containing_Q_not_followed_by_U.

Anonymous:2009:LMP

[Ano09e] Anonymous. LZX: The most powerful archiver available for

the Amiga, 2009. URL http://xavprods.free.fr/lzx/.

Anonymous:2009:SFC

[Ano09f] Anonymous. StuffIt [a family of computer software utilities for archiving and compressing files on the Macintosh and Microsoft Windows platforms], 2009. URL http://en.

wikipedia.org/wiki/Stuffit.

Anonymous:2010:CP

[Ano10] Anonymous. The Capocelli Prize. In Storer and Marcellin [SM10b], page vi. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=
5453519.

Anonymous:2011:CP

[Ano11a] Anonymous. The Capocelli Prize. In Storer and Marcellin [SM11b], page vi. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

5749553. IEEE Computer Society Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Anonymous:2011:PA

[Ano11b] Anonymous. Proceedings available. In Storer and Marcellin [SM11b], page ?? ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:/

/ieeexplore.ieee.org/servlet/opac?punumber=5749456.

IEEE Computer Society Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Anonymous:2012:CP

[Ano12a] Anonymous. The Capocelli Prize. In Storer and Marcellin [SM12], page vi. ISBN 1-4673-0715-7 (paperback). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ????

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=6189302. IEEE Computer Society order number P4656.

Anonymous:2012:SHS

[Ano12b] Anonymous. Secure Hash Standard (SHS). Federal Information Processing Standards Publication FIPS Pub 180-4, National Institute for Standards and Technology, Gaithersburg, MD, USA, March 2012. v + 30 pp. URL http://

csrc.nist.gov/publications/fips/fips180-4/fips-180-
4.pdf;   http://csrc.nist.gov/publications/PubsFIPS.
html#fips180-4.

Anonymous:2016:Z

[Ano16] Anonymous. Zstandard. Web site., 2016. URL http://

facebook.github.io/zstd/;   https://github.com/facebook/
zstd.

Anonymous:2018:HDC

[Ano18] Anonymous. HBO’s data compression stars — [back story]. IEEE Spectrum, 55(9):4, September 2018. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:20xx:IDM

[Anoxx] Anonymous. Ingrid Daubechies, mother of the wavelet. Web

story, 20xx. URL http://focusonbelgium.be/en/Do%20you%
20know%20these%20Belgians/ingrid-daubechies-mother-
wavelet.

Ahmed:1974:DCT

[ANR74] N. Ahmed, T. Natarajan, and K. R. Rao. Discrete Cosine Transfom. IEEE Transactions on Computers, C-23(1):90–93, January 1974. CODEN ITCOB4. ISSN 0018-9340 (print),

1557-9956 (electronic). URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1672377. See comments

[Sha75].

Arroyuelo:2012:SLZ

[ANS12] Diego Arroyuelo, Gonzalo Navarro, and Kunihiko Sadakane. Stronger Lempel–Ziv based compressed text indexing. Algorithmica, 62(1–2):54–101, February 2012. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=62&issue=1&spage=54.

Aizu:1995:ISM

[ANT95a] M. Aizu, O. Nakagawa, and M. Takagi. An image segmentation method based on a color space distribution model. In Storer and Cohn [SC95], page 439. ISBN 0-8186-7012-

6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=515549. IEEE catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Antoshenkov:1995:BAB

[Ant95b] G. Antoshenkov. Byte-aligned bitmap compression. In Storer and Cohn [SC95], page 476. ISBN 0-8186-7012-

6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=515586. IEEE catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Antoshenkov:1997:DBO

[Ant97] Gennady Antoshenkov. Dictionary-based order-preserving string compression (*). VLDB Journal: Very Large Data Bases, 6(1):26–39, February 1997. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http://ftp.

informatik.rwth-aachen.de/dblp/db/indices/a-tree/a/
Antoshenkov:Gennady.html;   http://link.springer.de/
link/service/journals/00778/bibs/7006001/70060026.
htm;   http://link.springer.de/link/service/journals/
00778/papers/7006001/70060026.pdf;   http://link.springer.
de/link/service/journals/00778/papers/7006001/70060026.

ps.gz. Electronic edition.

Al-Okaily:2017:TBC

[AOAAH17] Anas Al-Okaily, Badar Almarri, Sultan Al Yami, and Chun-Hsi Huang. Toward a better compression for DNA sequences using Huffman encoding. Journal of Computational Biology, 24(4):280–288, April 2017. CODEN JCOBEM. ISSN 1066-5277 (print), 1557-8666 (electronic). URL https://

www.liebertpub.com/doi/abs/10.1089/cmb.2016.0151;
https://www.liebertpub.com/doi/pdf/10.1089/cmb.2016.
0151.

Andriotis:2013:JSD

[AOT13] Panagiotis Andriotis, George Oikonomou, and Theo Tryfonas. JPEG steganography detection with Benford’s Law. Digital Investigation, 9(3–4):246–257, 2013. ISSN 1742-2876 (print),

1873-202X (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1742287613000066.

Apostolico:2003:CWF

[AP03] A. Apostolico and L. Parida. Compression and the wheel of fortune. In Storer and Cohn [SC03], pages 143–152. ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 2003. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1194005. IEEE Computer

Society Order number PR01896.

Ali:1992:UFT

[APC92] M. Ali, C. Papadopoulos, and T. Clarkson. The use of fractal theory in a video compression system. In Storer and Cohn [SC92], pages 259–268. ISBN 0-8186-2717-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ????

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=227455. IEEE catalog number 91TH0436-6.

Apiki:1991:LDC

[Api91] Steve Apiki. Lossless data compression: An explanation of two compression algorithms. Byte Magazine, 16(3):309–312, 314, 386–387, March 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Apostolico:2005:LZW

[Apo05] A. Apostolico. Of Lempel–Ziv–Welch parses with refillable gaps. In Storer and Cohn [SC05], pages 338–347. ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1402195.

Apple:2006:U

[App06] Apple Computer. [unknown], 2006. URL http://www.apple.

com/quicktime/technologies/aac/.

Arpe:2006:COG

[AR06] J. Arpe and R. Reischuk. On the complexity of optimal grammar-based compression. In Storer and Cohn [SC06], pages 173–182. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=1607252.

Acevedo:2008:LWB

[AR08] D. Acevedo and A. Ruedin. A lossless wavelet-based predictive multispectral image compressor. In Storer and Marcellin [SM08], page 501. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008;

QA76.9.D33 INTERNET. URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=4483328.

Akyol:2009:NDQ

[AR09a] E. Akyol and K. Rose. Nonuniform dithered quantization. In Storer and Marcellin [SM09], page 435. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=4976489. IEEE Computer Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

Akyol:2009:TCD

[AR09b] E. Akyol and K. Rose. On transform coding with dithered quantizers. In Storer and Marcellin [SM09], pages 243–251. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=4976468. IEEE Computer

Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

Akyol:2012:CRQ

[AR12a] E. Akyol and K. Rose. On constrained randomized quantization. In Storer and Marcellin [SM12], pages 72–81. ISBN 1-4673-0715-7 (paperback). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=6189238. IEEE Com-

puter Society order number P4656.

Akyol:2012:TOM

[AR12b] E. Akyol and K. Rose. Towards optimality in multiterminal transform coding. In Storer and Marcellin [SM12], pages 247–256. ISBN 1-4673-0715-7 (paperback). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

6189256. IEEE Computer Society order number P4656.

Atoofian:2018:DTS

[AR18] Ehsan Atoofian and Sean Rea. Data-type specific cache compression in GPGPUs. The Journal of Supercomputing, 74(4): 1609–1635, April 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Ang:1991:VCM

[ARA91] P. H. Ang, P. A. Ruetz, and D. Auld. Video compression makes big gains. IEEE Spectrum, 28(10):16–19, October 1991. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Arai:2013:MDH

[Ara13] Kohei Arai. Method for data hiding based on LeGall 5/3 (Cohen–Daubechies–Feauveau: CDF 5/3) wavelet with data compression and random scanning of secret imagery data. International Journal of Wavelets, Multiresolution and Information Processing, 11(4):1360006, 18, 2013. CODEN IJWMIP. ISSN 0219-6913 (print), 1793-690X (electronic).

Awan:2003:SRA

[ARH03] A. M. Awan, N. M. Rajpoot, and S. A. Husain. Stackrun adaptive wavelet image coding. In Storer and Cohn [SC03], page ?? ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2003. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1194033. IEEE Computer Society Order number PR01896.

Arnavut:1996:LCU

[Arn96] Z. Arnavut. Lossless compression using inversions on multiset permutations. In Storer and Cohn [SC96], page ?? ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1996. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=488348. IEEE Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Arnavut:1997:RTB

[Arn97] Z. Arnavut. A remapping technique based on permutations for lossless compression of multispectral images. In Storer and Cohn [SC97], pages 407–416. ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1997. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=582067. IEEE Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Arnavut:1999:MFP

[Arn99] Z. Arnavut. Move-to-front and permutation based inversion coding. In Storer and Cohn [SC99], page ?? ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1999. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=785672. IEEE Computer Society Or-

der Number PR00096. IEEE Order Plan Catalog Number PR00096.

Arnavut:2000:MFI

[Arn00] Z. Arnavut. Move-to-front and inversion coding. In Storer and Cohn [SC00], pages 193–202. ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2000. URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

838159. This millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Arnavut:2002:GBT

[Arn02] Z. Arnavut. Generalization of the BWT transformation and inversion ranks. In Storer and Cohn [SC02], page ?? ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2002. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=999990. IEEE

Computer Society Order Number PR01477.

Aronson:1977:CST

[Aro77] Jules Aronson. Computer science and technology: data compression — a comparison of methods. NBS special publication 500-12, Department of Commerce, National Bureau of Standards, Institute for Computer Sciences and Technology, Washington, DC, USA, June 1977. iv + 31 + 1 pp. URL

http://ucblibraries.colorado.edu/circulation/forms/
ericrequest.htm;   http://www.eric.ed.gov/contentdelivery/
servlet/ERICServlet?accno=ED149732. ERIC Document

Number: ED149732.

Aggarwal:2001:AOS

[ARR01] A. Aggarwal, S. L. Regunathan, and K. Rose. Asymptotically optimal scalable coding for minimum weighted mean square error. In Storer and Cohn [SC01c], pages 43–52. ISBN 0-7695-1031-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 2001. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=917135. IEEE CSP num-

ber 01PR1031.

Akyol:2010:OAM

[ARR10] E. Akyol, K. Rose, and T. Ramstad. Optimized analog mappings for distributed source-channel coding. In Storer and Marcellin [SM10b], pages 159–168. ISBN 0-7695-3994-

7. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=5453458.

Aaron:2003:WZC

[ARZG03] A. Aaron, S. Rane, Rui Zhang, and B. Girod. Wyner–Ziv coding for video: applications to compression and error resilience. In Storer and Cohn [SC03], pages 93–102. ISBN 0-7695-1896-

6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 2003. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1194000. IEEE Computer

Society Order number PR01896.

Aberg:1997:TCC

[AS97] J. Aberg and Yu. M. Shtarkov. Text compression by context tree weighting. In Storer and Cohn [SC97], pages 377–386. ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33

D232 1997. URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=582062. IEEE Computer Society

Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Arelakis:2014:CVA

[AS14a] Angelos Arelakis and Per Stenstrom. A case for a value-aware cache. IEEE Computer Architecture Letters, 13(1):1–4, January/June 2014. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Arelakis:2014:SSC

[AS14b] Angelos Arelakis and Per Stenstrom. SC2: a statistical compression cache scheme. ACM SIGARCH Computer Architecture News, 42(3):145–156, June 2014. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Alzina:1999:PMI

[ASG99] M. Alzina, W. Szpankowski, and A. Grama. 2D-pattern matching image and video compression. In Storer and Cohn [SC99], pages 424–433. ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1999. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=755692. IEEE

Computer Society Order Number PR00096. IEEE Order Plan Catalog Number PR00096.

Adelson:1987:OPT

[ASH87] E. H. Adelson, E. Simoncelli, and R. Hingorani. Orthogonal pyramid transforms for image coding. In Proceedings SPIE, pages 50–58. vol. 845, Cambridge, MA, USA, October 1987.

Araujo:2020:ECQ

[ASJ20] Leonardo C. Araujo, Joao P. H. Sansao, and Mario C. S. Junior. Effects of color quantization on JPEG compression. International Journal of Image and Graphics (IJIG), 20(03):??, July 2020. ISSN 0219-4678. URL https://www.

worldscientific.com/doi/10.1142/S0219467820500266.

Al-Shammary:2012:RAS

[ASK12] Dhiah Al-Shammary and Ibrahim Khalil. Redundancyaware SOAP messages compression and aggregation for enhanced performance. Journal of Network and Computer Applications, 35(1):365–381, January 2012. CO-DEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S108480451100172X.

Anand:2020:CTE

[ASLB20] Ashima Anand, Amit Kumar Singh, Zhihan Lv, and Guarav Bhatnagar. Compression-then-encryption-based secure watermarking technique for smart healthcare system. IEEE Multi-Media, 27(4):133–143, 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Al-Shaykh:1996:LCN

[ASM96] O. K. Al-Shaykh and R. M. Mersereau. Lossy compression of noisy cardiac image sequences. In Storer and Cohn [SC96], pages 43–52. ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1996. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=488309. IEEE

Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Arildsen:2010:FLS

[AsM^{+}10] T. Arildsen, J. stergaard, M. N. Murthi, S. V. Andersen, and S. H. Jensen. Fixed-lag smoothing for low-delay predictive coding with noise shaping for lossy networks. In Storer and Marcellin [SM10b], pages 279–287. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=5453446.

Aljifri:2003:ITU

[ASP03] Hassan Aljifri, Marcel Smets, and Alexander Pons. IP traceback using header compression. Computers & Security, 22(2):136–151, February 2003. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

https://www.sciencedirect.com/science/article/pii/
S0167404803002128.

Aberg:1997:TUI

[ASS97] J. Aberg, Yu. M. Shtarkov, and B. J. M. Smeets. Towards understanding and improving escape probabilities in PPM. In Storer and Cohn [SC97], pages 22–31. ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1997. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=581954. IEEE Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Aberg:1998:NUP

[ASS98] J. Aberg, Yu. M. Shtarkov, and B. J. M. Smeets. Non-uniform PPM and context tree models. In Storer and Cohn [SC98], pages 279–288. ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1998. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=672156. IEEE

Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

AT-T:1996:U

[AT-96] AT-T. [unknown], 1996. URL http://www.djvu.att.com/.

Abel:2005:UTP

[AT05] J. Abel and W. Teahan. Universal text preprocessing for data compression. IEEE Transactions on Computers, 54(5):497– 507, May 2005. CODEN ITCOB4. ISSN 0018-9340 (print),

1557-9956 (electronic). URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1407841.

Arps:1988:MPV

[ATL^{+}88] Ronald B. Arps, Thomas K. Truong, David J. Lu, Richard C. Pasco, and Theodore David Friedman. A multi-purpose VLSI chip for adaptive data compression of bilevel images. IBM Journal of Research and Development, 32(6):775–795 (or 775– 794??), November 1988. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

ATSC:2006:U

[ATS06] ATSC. [unknown], 2006. URL http://www.atsc.org/

standards/a_52b.pdf.

Atteson:1995:FDO

[Att95] Kevin Scott Atteson. A formalism for the design of optimal adaptive text data compression rules. Ph.D. thesis, University of Pennsylvania, Philadelphia, PA, USA, 1995. 148 pp. URL

http://search.proquest.com/docview/304212670.

Aucsmith:1998:IHI

[Auc98] David Aucsmith, editor. Information hiding II: Second International Workshop, IH’98: Portland, Oregon, USA, April 1998: Proceedings, volume 1525 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. CODEN LNCSD9. ISBN 3-540-65386-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25I48 1998. URL http://

link.springer-ny.com/link/service/series/0558/tocs/
t1525.htm;   http://www.springerlink.com/content/978-
3-540-65386-8;   http://www.springerlink.com/openurl.
asp?genre=issue&issn=0302-9743&volume=1525.

Ausbeck:1998:CMP

[Aus98] P. J. Ausbeck, Jr. Context models for palette images. In Storer and Cohn [SC98], pages 309–318. ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1998. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=672159. IEEE Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Ausbeck:1999:SPC

[Aus99] P. J. Ausbeck, Jr. A streaming piecewise-constant model. In Storer and Cohn [SC99], pages 208–217. ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1999. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=755670. IEEE Computer Society Or-

der Number PR00096. IEEE Order Plan Catalog Number PR00096.

Ausbeck:2000:SIM

[Aus00] P. J. Ausbeck, Jr. The skip-innovation model for sparse images. In Storer and Cohn [SC00], pages 43–52. ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2000. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=838144. This millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Aubert:2018:PDC

[AVE18] Pierre Aubert, Thomas Vuillaume, and Nahid Emad. Polynomial data compression for large-scale physics experiments.

Computing and Software for Big Science, 2(1):??, November 2018. CODEN ???? ISSN 2510-2036 (print), 2510-2044

(electronic). URL https://link.springer.com/article/
10.1007/s41781-018-0010-3.

Alameldeen:2004:ACC

[AW04] Alaa R. Alameldeen and David A. Wood. Adaptive cache compression for high-performance processors. ACM SIGARCH Computer Architecture News, 32(2):212, March 2004. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Altaf:2015:LPM

[AW15] Muhammad Shoaib Bin Altaf and David A. Wood. LogCA: A performance model for hardware accelerators. IEEE Computer Architecture Letters, 14(2):132–135, July/December 2015. CO-DEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Burrows:1992:LDC

[B^{+}92] Michael Burrows et al. On-line data compression in a logstructured file system. Technical report, Digital Systems Research Center, Palo Alto, CA, USA, 1992. ?? pp.

Borgelt:2013:TAD

[B^{+}13] Christian Borgelt et al., editors. Towards advanced data analysis by combining soft computing and statistics, volume 285 of Studies in fuzziness and soft computing. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 016194883, 3-642-30278-5 (e-book). LCCN ????

Burt:1983:LPC

[BA83] Peter J. Burt and Edward H. Adelson. The Laplacian pyramid as a compact image code. IEEE Transactions on Communications, COM-31(4):532–540, April 1983. CODEN IECMBT. ISSN 0090-6778 (print), 1558-0857 (electronic).

Boussakta:2004:FAD

[BA04] Said Boussakta and Hamoud O. Alshibami. Fast algorithm for the 3-D DCT-II. IEEE Transactions on Signal Processing, 52(4):??, 2004. CODEN ITPRED. ISSN 1053-587X (print), 1941-0476 (electronic).

Barr:2006:EAL

[BA06] Kenneth C. Barr and Krste Asanović. Energy-aware lossless data compression. ACM Transactions on Computer Systems, 24(3):250–291, August 2006. CODEN ACSYEC. ISSN 0734-2071 (print), 1557-7333 (electronic).

Batista:2008:TPP

[BA08] L. Batista and L. A. Alexandre. Text pre-processing for lossless compression. In Storer and Marcellin [SM08], page 506. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33 INTER-

NET. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=4483333.

Beal:2016:CPP

[BA16] Richard Beal and Donald Adjeroh. Compressed parameterized pattern matching. Theoretical Computer Science, 609 (part 1)(??):129–142, January 4, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304397515008440.

Bras-Amoros:2006:SIC

[BAGCGC^{+}06] M. Bras-Amoros, J. Gonzalez-Conjero, P. Guitart-Colom, J. Serra-Sagristà, and F. Garcia-Vilchez. Still image compression through exhaustive two-valued shape-adaptive searches. In Storer and Cohn [SC06], page 441. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=1607284.

Bakulina:2007:AZL

[Bak07] M. P. Bakulina. Application of Zipf’s law for text compression. Diskretn. Anal. Issled. Oper. Ser. 2, 14(2):3–13, 2007. CODEN ???? ISSN 1560-9901.

Ball:1997:GIF

[Bal97] Jimmy Ball. GIF images on the fly: New images from old, in real time. Linux Journal, 43:??, November 1997. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL

ftp://ftp.ssc.com/pub/lj/listings/issue43/2411.tgz.

Behnamfar:2002:PIC

[BAL02] F. Behnamfar, F. Alajaji, and T. Linder. Progressive image communication over binary channels with additive bursty noise. In Storer and Cohn [SC02], pages 272–281. ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2002. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=999965. IEEE

Computer Society Order Number PR01477.

Banikazemi:2009:LDC

[Ban09] M. Banikazemi. LZB: Data compression with bounded references. In Storer and Marcellin [SM09], page 436. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4976490. IEEE Computer Society

Order Number P3592. BMS Part Number CFP09DCC-PRT. Barba:1981:MAD

[Bar81] Joseph Barba. A Modified Adaptive Delta Modulator for Video Data Compression. Ph.D. thesis, City University of New York, New York, NY, USA, 1981. 219 pp. URL http://search.

proquest.com/docview/303098954.

Barnsley:1988:FE

[Bar88] Michael Barnsley. Fractals Everywhere. Academic Press, New York, USA, 1988. ISBN 0-12-079062-9. xii + 394 pp. LCCN QA614.86 .B37 1988.

Baran:1990:PSG

[Bar90a] N. Baran. Putting the squeeze on graphics: Compression technologies for full-color graphics and full-motion video. Byte Magazine, 15(13):289–290, 292–294, December 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Barsimantov:1990:DC

[Bar90b] Elie Oscar Barsimantov. Data compression. Thesis (M.S.), University of Colorado, Boulder, CO, USA, 1990. 132 leaves pp.

Barnes:1994:NMP

[Bar94] C. F. Barnes. A new multiple path search technique for residual vector quantizers. In Storer and Cohn [SC94], pages 42– 51. ISBN 0-8186-5636-0, 0-8186-5637-9 (case). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1994.

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=305911. IEEE Catalog Number 93TH0626-2. IEEE Computer Society Press Order Number 5637-02.

Baran:2000:NVNb

[Bar00a] Nicholas Baran. News and views: Neural net crowned king; smart shirts monitor vital signs; McMaster team wins college design contest; Java applet tracks space station; Sandia Red Team batting 1000; new digital geometry compression algorithm. Dr. Dobb’s Journal of Software Tools, 25(10):18, October 2000. CODEN DDJOEB. ISSN 1044-789X.

Baran:2000:NVV

[Bar00b] Nicholas Baran. News and views: Volunteers needed for weather simulation model; Bill Gates meets Bill Blass — not!; computer lemon law may debut in Pennsylvania; inventor of Huffman encoding dies; green cards, not guest visas; online forums on electronic recycling. Dr. Dobb’s Journal of Software Tools, 25(1):18, January 2000. CODEN DDJOEB. ISSN 1044-789X.

Barnes:2004:SAS

[Bar04] C. F. Barnes. Successive approximation source coding and image enabled data mining. In Storer and Cohn [SC04], page ?? ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1281501. IEEE Computer

Society Order Number: P2082.

Barron:2005:DWP

[Bar05] David W. Barron. David Wheeler: a personal memoir. The Computer Journal, 48(6):650–651, November 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/cgi/content/
full/48/6/650;   http://comjnl.oxfordjournals.org/
cgi/reprint/48/6/650.

Barbero:2011:LRS

[Bar11] J. M. Barbero. Lossy RAID storage architecture for JPEG 2000 images. In Storer and Marcellin [SM11b], page 445. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5749502. IEEE Computer Society

Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Barbero:2012:SRS

[Bar12] J. M. Barbero. Scalable RAID storage based on the structure of multimedia file. In Storer and Marcellin [SM12], page 390. ISBN 1-4673-0715-7 (paperback). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

6189271. IEEE Computer Society order number P4656.

Bassiouni:1985:DCS

[Bas85] M. A. Bassiouni. Data compression in scientific and statistical databases. IEEE Transactions on Software Engineering, SE-11(10):1047–1058, October 1985. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?arnumber=1701920.

Basu:1991:TCU

[Bas91] D. Basu. Text compression using several Huffman trees. In Storer and Reif [SR91b], page ?? ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1990. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=213309. IEEE Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Bass:1992:EP

[Bas92] Thomas A. Bass. Eudaemonic Pie. Penguin Books, Harmondsworth, UK, 1992. ?? pp.

Bunton:1992:PDM

[BB92] Suzanne Bunton and Gaetano Borriello. Practical dictionary management for hardware data compression. Communications of the ACM, 35(1):95–104, January 1992. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org/pubs/toc/Abstracts/
0001-0782/129622.html.

Bradley:1993:WTV

[BB93] J. N. Bradley and C. M. Brislawn. Wavelet transform-vector quantization compression of supercomputer ocean models. In Storer and Cohn [SC93], pages 224–233. ISBN 0-8186-3391-3 (microfiche), 0-8186-3392-1 (casebound). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1993. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=253127. IEEE Computer Society Press order number 3392-02. IEEE catalog number 93TH0536-3.

Brown:1995:PRB

[BB95] R. Brown and A. F. Boden. A posteriori restoration of block transform-compressed data. In Storer and Cohn [SC95], page 426. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=515536. IEEE catalog

number 95TH8037. IEEE Computer Society Press order number PR07010.

Blandford:2002:ICT

[BB02] D. Blandford and G. Blelloch. Index compression through document reordering. In Storer and Cohn [SC02], pages 342– 351. ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2002. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=999972. IEEE

Computer Society Order Number PR01477.

Baron:2005:ASS

[BB05] D. Baron and Y. Bresler. Antisequential suffix sorting for BWT-based data compression. IEEE Transactions on Computers, 54(4):385–397, April 2005. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=
1401858.

Boufounos:2007:QSR

[BB07] P. Boufounos and R. Baraniuk. Quantization of sparse representations. In Storer and Cohn [SC07], page 378. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4148779. IEEE Computer Society

Order Number P2791.

Bouttefroy:2010:MRM

[BBBP10] P. L. M. Bouttefroy, A. Bouzerdoum, A. Beghdadi, and S. L. Phung. Multi-resolution mean-shift algorithm for vector quantization. In Storer and Marcellin [SM10b], page 523. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5453500.

Bhattacharjee:1995:EDC

[BBC95] S. Bhattacharjee, J. Bhattacharya, and P. P. Chaudhuri. An efficient data compression hardware based on cellular automata. In Storer and Cohn [SC95], page 472. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=515582. IEEE catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Banister:2002:RVT

[BBF02] B. A. Banister, B. Belzer, and T. R. Fischer. Robust video transmission over binary symmetric channels with packet erasures. In Storer and Cohn [SC02], pages 162–171. ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 2002. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=999954. IEEE Computer

Society Order Number PR01477.

Bjorner:2010:CDC

[BBG10] Nikolaj Bjørner, Andreas Blass, and Yuri Gurevich. Contentdependent chunking for differential compression, the local maximum approach. Journal of Computer and System Sciences, 76(3–4):154–203, May/June 2010. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0022000009000580.

Bradley:1993:FWS

[BBH93] Jonathan N. Bradley, Christopher M. Brislawn, and Tom Hopper. The FBI wavelet/scalar quantization standard for grayscale fingerprint image compression. In Proceedings of Visual Information Processing II, Orlando, FL, volume 1961, pages 293–304. SPIE Optical Engineering Press, Bellingham, WA, USA, April 1993.

Bhatt:1997:DTM

[BBH97] B. Bhatt, D. Birks, and D. Hermreck. Digital television: making it work. IEEE Spectrum, 34(10):19–28, October 1997. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bottcher:2011:SMC

[BBH11] S. Böttcher, A. Bültmann, and R. Hartel. Search and modification in compressed texts. In Storer and Marcellin [SM11b], pages 403–412. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

5749498. IEEE Computer Society Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Byrnes:1994:WTA

[BBHB94] James S. Byrnes, Jennifer L. Byrnes, Kathryn A. Hargreaves, and Karl Berry, editors. Wavelets and Their Applications, volume 442 of NATO ASI Series C: Mathematical and Physical Sciences. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1994. ISBN 0-7923-3078-1. Proceedings of the NATO Advanced Study Institute at Il Ciocco, Barga, Italy in August, 1992.

Bremler-Barr:2014:CSP

[BBHK14] Anat Bremler-Barr, David Hay, and Yaron Koral. CompactDFA: Scalable pattern matching using longest prefix match solutions. IEEE/ACM Transactions on Networking, 22 (2):415–428, April 2014. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Bottcher:2012:FID

[BBHS12] S. Bottcher, A. Bultmann, R. Hartel, and J. Schlussler. Fast insertion and deletion in compressed texts. In Storer and Marcellin [SM12], page 393. ISBN 1-4673-0715-7 (paperback). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=6189274. IEEE Computer Society order number P4656.

Brown:1995:PSI

[BBKF95] Allen Brown, Anne Brüggemann-Klein, and An Feng, editors. Proceedings of the Sixth International Conference on Electronic Publishing, Document Manipulation and Typography, 24–26 September, 1996, Palo Alto, CA, USA, volume 8(2) of Electronic Publishing—Origination, Dissemination, and Design. John Wiley, New York, NY, USA, 1995. ISSN 0894-3982.

Brislawn:1996:FCS

[BBOH96] Christopher Brislawn, Jonathan Bradley, R. Onyshczak, and Tom Hopper. The FBI compression standard for digitized fingerprint images. In Proceedings SPIE, Denver, CO, volume 2847, pages 344–355. SPIE Optical Engineering Press, Bellingham, WA, USA, August 1996.

Bradley:1995:WSC

[BBQ^{+}95] J. N. Bradley, C. M. Brislawn, D. J. Quinlan, H. D. Zhang, and V. Nuri. Wavelet subband coding of computer simulation output using the A++ array class library. In Storer and Cohn [SC95], page 454. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=515564. IEEE catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Bjorklund:2014:CFS

[BBZ14] Henrik Björklund, Johanna Björklund, and Niklas Zechner. Compression of finite-state automata through failure transitions. Theoretical Computer Science, 557(??):87–100, November 6, 2014. CODEN TCSCDI. ISSN 0304-3975 (print),

1879-2294 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304397514006720.

Bai:2005:IML

[BC05] Yuhua Bai and T. Cooklev. An improved method for lossless data compression. In Storer and Cohn [SC05], page ?? ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1402208.

Balachandran:2008:SHC

[BC08] S. Balachandran and C. Constantinescu. Sequence of hashes compression in data de-duplication. In Storer and Marcellin [SM08], page 505. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008;

QA76.9.D33 INTERNET. URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=4483332.

Bursalioglu:2010:FMD

[BC10] O. Y. Bursalioglu and G. Caire. A flexible multiple description coding scheme based on rateless codes. In Storer and Marcellin [SM10b], pages 269–278. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=5453447.

Buttcher:2010:IRI

[BCC10] Stefan Büttcher, Charles L. A. Clarke, and Gordon V. Cormack. Information Retrieval: Implementing and Evaluating Search Engines. MIT Press, Cambridge, MA, USA, 2010. ISBN 0-262-02651-1 (hardcover). xxiv + 606 pp. LCCN TK5105.884 .B98 2010.

Bille:2018:DRC

[BCC^{+}18] Philip Bille, Anders Roy Christiansen, Patrick Hagge Cording, Inge Li Gørtz, Frederik Rye Skjoldjensen, Hjalte Wedel Vildhøj, and Søren Vind. Dynamic relative compression, dynamic partial sums, and substring concatenation. Algorithmica, 80(11):3207–3224, November 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Basile:1995:UHS

[BCD^{+}95] C. Basile, A. P. Cavallerano, M. S. Deiss, R. Keeler, J. S. Lim, W. C. Luplow, W. H. Paik, E. Petajan, R. Rast, G. Reitmeier, T. R. Smith, and C. Todd. The US HDTV standard. IEEE Spectrum, 32(4):36–45, April 1995. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Barcaccia:1998:PMT

[BCD98] P. Barcaccia, A. Cresti, and S. De Agostino. Pattern matching in text compressed with the ID heuristic. In Storer and Cohn [SC98], pages 113–118. ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1998. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=672137. IEEE Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Barbay:2014:EFC

[BCG^{+}14] Jérémy Barbay, Francisco Claude, Travis Gagie, Gonzalo Navarro, and Yakov Nekrich. Efficient fully-compressed sequence representations. Algorithmica, 69(1):232–268, May 2014. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL http://link.springer.com/article/10.
1007/s00453-012-9726-3.

Barth:2002:MMM

[BCH02] Timothy J. Barth, Tony Chan, and Robert Haimes, editors. Multiscale and Multiresolution Methods: Theory and Applications, volume 20 of Lecture Notes in Computational Science and Engineering. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. CODEN LNCSA6. ISBN 3-540-42420-2 (print), 3-642-56205-1 (e-book). ISSN 1439-7358. LCCN QA403.3

.M85 2002. URL http://link.springer.com/book/10.
1007/978-3-642-56205-1;   http://www.springerlink.
com/content/978-3-642-56205-1. Papers from the Yosemite

Educational Symposium (YES), Yosemite valley, California, Fall 2000.

Becker:2004:FRI

[BCP04] A. Becker, W. Chan, and D. Poulouin. Flicker reduction in intraframe codecs. In Storer and Cohn [SC04], pages 252–261. ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1281470. IEEE Computer

Society Order Number: P2082.

Bruni:2020:FLD

[BCP20] Vittoria Bruni, Mariantonia Cotronei, and Francesca Pitolli. A family of level-dependent biorthogonal wavelet filters for image compression. Journal of Computational and Applied Mathematics, 367(??):Article 112467, March 15, 2020. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0377042719304704.

Bergadano:1998:HDC

[BCR98] Francesco Bergadano, Bruno Crispo, and Giancarlo Ruffo. High dictionary compression for proactive password checking. ACM Transactions on Information and System Security, 1(1):3–25, November 1998. CODEN ATISBQ. ISSN 1094-9224 (print), 1557-7406 (electronic). URL http://www.

acm.org:80/pubs/citations/journals/tissec/1998-1-1/
p3-bergadano/.

Bassino:2006:OPC

[BCSV06] F. Bassino, J. Clement, G. Seroussi, and A. Viola. Optimal prefix codes for some families of two-dimensional geometric distributions. In Storer and Cohn [SC06], pages 113–122. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1607246.

Bell:1990:TC

[BCW90] Timothy C. Bell, John G. Cleary, and Ian H. Witten. Text Compression. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1990. ?? pp.

Bell:1982:KSC

[BD82] D. A. Bell and S. M. Deen. Key space compression and hashing in PRECI. The Computer Journal, 25(4):486–492, November 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/
content/25/4/486.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_25/Issue_04/tiff/
486.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_04/tiff/487.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_25/Issue_04/
tiff/488.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_04/tiff/489.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_25/Issue_04/
tiff/490.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_04/tiff/491.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_25/Issue_04/
tiff/492.tif.

Bhushan:2018:CCU

[BD18] S. N. Bharath Bhushan and Ajit Danti. Classification of compressed and uncompressed text documents. Future Generation Computer Systems, 88(??):614–623, November 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167739X17320708.

Biswas:2006:QTP

[BdB06] A. Biswas and A. C. den Brinker. Quantization of transmission parameters in stereo linear predictive systems. In Storer and Cohn [SC06], pages 262–271. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=1607261.

Brisaboa:2012:CDB

[BdBN12] N. R. Brisaboa, G. de Bernardo, and G. Navarro. Compressed dynamic binary relations. In Storer and Marcellin [SM12], pages 52–61. ISBN 1-4673-0715-7 (paperback). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=6189236. IEEE Computer Society order number P4656.

Bhattacharjee:1997:PAA

[BDCC97] S. Bhattacharjee, S. Das, Y. Chowdhury, and P. P. Chaudhuri. A pipelined architecture algorithm for image compression. In Storer and Cohn [SC97], page ?? ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1997. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=582080. IEEE Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Bertram:2004:GBS

[BDHJ04] Martin Bertram, Mark A. Duchaineau, Bernd Hamann, and Kenneth I. Joy. Generalized B-spline subdivision-surface wavelets for geometry compression. IEEE Transactions on Visualization and Computer Graphics, 10(3):326–338, May/June 2004. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL http://csdl.computer.org/

comp/trans/tg/2004/03/v0326abs.htm;   http://csdl.
computer.org/dl/trans/tg/2004/03/v0326.htm;   http:
//csdl.computer.org/dl/trans/tg/2004/03/v0326.pdf.

Bode:1996:PVM

[BDLS96] Arndt Bode, Jack Dongarra, T. Ludwig, and V. Sunderam, editors. Parallel virtual machine, EuroPVM ’96: third European PVM conference, Munich, Germany, October 7–9, 1996: proceedings, volume 1156 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-540-61779-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.E975 1996.

Belinskaya:1995:NOC

[BDS95] D. Belinskaya, S. DeAgostino, and J. A. Storer. Near optimal compression with respect to a static dictionary on a practical massively parallel architecture. In Storer and Cohn [SC95], pages 172–181. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ????

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=515507. IEEE catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Balan:2000:ADW

[BDV00] Radu Balan, Ingrid Daubechies, and Vinay Vaishampayan. The analysis and design of windowed Fourier frame based multiple description source coding schemes. IEEE Transactions on Information Theory, 46(7):2491–2536, November 2000. CO-DEN IETTAW. ISSN 0018-9448 (print), 1557-9654 (electronic).

Bani-Eqbal:1995:ESF

[BE95a] Behnam Bani-Eqbal. Enhancing the speed of fractal image compression. Optical Engineering, 34(6):1705–1710, June 1995. CODEN OPEGAR. ISSN 0091-3286. URL ftp:

//links.uwaterloo.ca/pub/Fractals/Papers/bani95.ps.
gz.

Bani-Eqbal:1995:SFI

[BE95b] Behnam Bani-Eqbal. Speeding up fractal image compression. Proceedings of the SPIE — The International Society for Optical Engineering, 2418(??):67–74, February 1995. CODEN PSISDG. ISSN ????

Bani-Eqbal:1996:CTF

[BE96] B. Bani-Eqbal. Combining tree and feature classification in fractal encoding of images. In Storer and Cohn [SC96], page ?? ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN QA76.9.D33 D37 1996. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=488350. IEEE

Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Becker:1993:IVC

[Bec93] Jeffrey Chaim Becker. Increasing VLSI Chip Address and Data I/O Pin Bandwidth Through Compression. Ph.D. thesis, University of California, Davis, Davis, CA, USA, 1993. 104 pp.

URL http://search.proquest.com/docview/304049364.

Beebe:1979:IFD

[Bee79] Nelson H. F. Beebe. Integral file data compression. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, May 18, 1979. 39 pp.

Bille:2018:TST

[BEGV18] Philip Bille, Mikko Berggren Ettienne, Inge Li Gørtz, and Hjalte Wedel Vildhøj. Time-space trade-offs for Lempel–Ziv compressed indexing. Theoretical Computer Science, 713(??):66–77, February 22, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0304397517309167.

Beheshti:2008:DCL

[Beh08] S. Beheshti. Data compression and linear modeling. In Storer and Marcellin [SM08], page 507. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33 INTERNET.

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=
&arnumber=4483334.

Bell:1986:BOT

[Bel86] Timothy C. Bell. Better OPM/L text compression. IEEE Transactions on Communications, 34(12):1176–1182, December 1986. CODEN IECMBT. ISSN 0090-6778 (print), 1558-0857 (electronic).

Bell:1987:UTI

[Bel87] Timothy C. Bell. A Unifying Theory and Improvements for Existing Approaches to Text Compression. Ph.D. thesis, Department of Computer Science, University of Canterbury, Christchurch, New Zealand, ????, 1987. ?? pp.

Bellis:2006:HMF

[Bel06] Mary Bellis. The history of MP3: Fraunhofer Gesellschaft

and MP3, 2006. URL http://inventors.about.com/od/
mstartinventions/a/;   http://inventors.about.com/
od/mstartinventions/a/MPThree.htm.

Belkasim:2011:MRA

[Bel11] S. Belkasim. Multi-resolution analysis using symmetrized odd and even DCT transforms. In Storer and Marcellin [SM11b], page 447. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=5749504. IEEE

Computer Society Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Benschop:1996:MPS

[Ben96] L. C. Benschop. 100 Mbit per second VLSI implementation of sliding window coding. In Storer and Cohn [SC96], page ?? ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1996. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=488352. IEEE

Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Benschop:1997:LDC

[Ben97] Leonard Cornelius Benschop. Lossless data compression in VLSI. Masterarbeit, Technische Universiteit Eindhoven, Eindhoven, The Netherlands, 1997. 242 pp. URL http://search.

proquest.com/docview/304401746.

Benatti:2010:QIC

[Ben10] Fabio Benatti, editor. Quantum information, computation and cryptography: an introductory survey of theory, technology and experiments, volume 808 of Lecture notes in physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 3-642-11914-X, 3-642-11913-1. ???? pp. LCCN Q360 Q36 2010.

Berkovich:1998:REC

[BEQ98] S. Berkovich and E. El-Qawasmeh. Reversing the errorcorrection scheme for a fault-tolerant indexing. In Storer and Cohn [SC98], page ?? ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1998. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=672237. IEEE

Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Berger:1971:RDT

[Ber71] Toby Berger. Rate distortion theory; a mathematical basis for data compression. Prentice-Hall series in information and system sciences; Prentice-Hall electrical engineering series. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1971. ISBN 0-13-753103-6. xiii + 311 pp. LCCN Q385 .B47.

Berke:1999:CAP

[Ber99] J. Berke. Comparison and application possibilities of JPEG and fractal-based image compressing methods in the development of multimedia-based material. In Storer and Cohn [SC99], page ?? ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1999. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=785674. IEEE

Computer Society Order Number PR00096. IEEE Order Plan Catalog Number PR00096.

Brittain:2005:GTD

[BES05] Nathanael J. Brittain and Mahmoud R. El-Sakka. Grayscale two-dimensional Lempel–Ziv encoding. In International Conference on Image Analysis and Recognition, ICIAR’2005, volume 3656 of Lecture Notes in Computer Science, pages 328– 334. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005.

Brittain:2007:GTT

[BES07] Nathanael Brittain and Mahmoud R. El-Sakka. Grayscale true two-dimensional dictionary-based image compression. Journal of Visual Communication and Image Representation, 18(1): 35–44, February 2007. CODEN JVCRE7. ISSN 1047-3203 (print), 1095-9076 (electronic).

Betzos:1994:AFC

[Bet94] George Anastasios Betzos. Algorithms for fast compression of quadtree and octree data using content addressable memories. Ph.D. thesis, Syracuse University, Syracuse, NY, USA,

1994. 236 pp. URL http://search.proquest.com/docview/
304139463.

Barrett:1995:LBD

[BF95a] C. W. Barrett and R. L. Frost. Lattice-based designs of direct sum codebooks for vector quantization. In Storer and Cohn [SC95], page 436. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

515546. IEEE catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Buttigieg:1995:MPM

[BF95b] Victor Buttigieg and P. G. Farrell. A Maximum A-Posteriori (MAP) decoding algorithm for variable-length error-correcting codes. In Codes and Cyphers: Cryptography and Coding IV, pages 103–119. Institute of Mathematics and Its Applications, Southend-on-sea, Essex, England, 1995.

Brunk:1996:FRS

[BF96] H. Brunk and N. Farvardin. Fixed-rate successively refinable scalar quantizers. In Storer and Cohn [SC96], pages 250–259. ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1996. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=488330. IEEE

Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Brunk:1998:ETC

[BF98] H. Brunk and N. Farvardin. Embedded trellis coded quantization. In Storer and Cohn [SC98], pages 93–102. ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1998. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=672135. IEEE Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Banister:1999:QCT

[BF99a] Brian A. Banister and Thomas R. Fischer. Quadtree classification and TCQ image coding. In Storer and Cohn [SC99], pages 149–157. ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1999. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=755664. IEEE

Computer Society Order Number PR00096. IEEE Order Plan Catalog Number PR00096.

Barequet:1999:SSI

[BF99b] R. Barequet and M. Feder. SICLIC: a simple inter-color lossless image coder. In Storer and Cohn [SC99], pages 501–510. ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1999. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=755700. IEEE Computer Society Or-

der Number PR00096. IEEE Order Plan Catalog Number PR00096.

Brunk:1999:JDP

[BF99c] H. Brunk and N. Farvardin. Joint design of progressive fixedrate source-channel codes. In Storer and Cohn [SC99], page ?? ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1999. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=785675. IEEE Computer Society Or-

der Number PR00096. IEEE Order Plan Catalog Number PR00096.

Boettcher:2007:MBA

[BF07] J. B. Boettcher and J. E. Fowler. A modified BISK algorithm for 3D dual-treewavelet transform coding. In Storer and Cohn [SC07], page 377. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=4148778. IEEE Computer Society Order Number P2791.

Beirami:2011:RTP

[BF11] A. Beirami and F. Fekri. The redundancy of two-part codes for finite-length parametric sources. In Storer and Marcellin [SM11b], page 446. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

5749503. IEEE Computer Society Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Beirami:2012:MAU

[BF12] A. Beirami and F. Fekri. Memory-assisted universal source coding. In Storer and Marcellin [SM12], page 392. ISBN 1-4673-0715-7 (paperback). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=6189273. IEEE Com-

puter Society order number P4656.

Bursalioglu:2009:JSC

[BFCP09] O. Y. Bursalioglu, M. Fresia, G. Caire, and H. V. Poor. Joint source-channel coding at the application layer. In Storer and Marcellin [SM09], pages 93–102. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=4976453. IEEE Computer Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

Bille:2009:IAS

[BFG09] Philip Bille, Rolf Fagerberg, and Inge Li Gørtz. Improved approximate string matching and regular expression matching on Ziv–Lempel compressed texts. ACM Transactions on Algorithms, 6(1):3:1–3:??, December 2009. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Brisaboa:2010:NSV

[BFL^{+}10] N. R. Brisaboa, A. Faria, J. R. Lopez, G. Navarro, and E. R. Lopez. A new searchable variable-to-variable compressor. In Storer and Marcellin [SM10b], pages 199–208. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=5453462.

Behr:2003:ECE

[BFMX03] F. Behr, V. Fossum, M. Mitzenmacher, and D. Xiao. Estimating and comparing entropies across written natural languages using PPM compression. In Storer and Cohn [SC03], page ?? ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2003. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1194035. IEEE Computer Society Order number PR01896.

Brisaboa:2010:DLT

[BFNP10] Nieves Brisaboa, Antonio Fariña, Gonzalo Navarro, and José Paramá. Dynamic lightweight text compression. ACM Transactions on Information Systems, 28(3):10:1–10:??, June 2010. CODEN ATISET. ISSN 1046-8188.

Balster:2011:PCR

[BFT11] Eric J. Balster, Benjamin T. Fortener, and William F. Turri. Post-compression rate-distortion development for embedded block coding with optimal truncation in JPEG2000 imagery. International Journal of Image and Graphics (IJIG), 11(4): 611–627, October 2011. CODEN ???? ISSN 0219-4678.

Bruce:1996:AWA

[BG96] Andrew Bruce and Hong-Ye Gao. Applied wavelet analysis with S-Plus. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 0-387-94714-0 (softcover). xxi + 338 pp. LCCN TK5102.9.B78 1996.

URL http://www.springer-ny.com/catalog/np/jul96np/
DATA/0-387-94714-0.html.

Banerjee:2003:TRU

[BG03a] S. Banerjee and M. J. Gormish. The transport of reversible and unreversible embedded wavelets (TRUEW). In Storer and Cohn [SC03], page ?? ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2003. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1194034. IEEE Computer Society Order number PR01896.

Bosi:2003:IDA

[BG03b] Marina Bosi and Richard E. Goldberg. Introduction To Digital Audio Coding and Standards. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2003. ?? pp.

Ba:2006:NTC

[BG06] D. E. Ba and V. K. Goyal. Nonlinear transform coding: polar coordinates revisited. In Storer and Cohn [SC06], page 438. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=1607281.

Barret:2009:LHI

[BGBV09] M. Barret, J.-L. Gutzwiller, I. P. A. Bita, and F. D. Vedova. Lossy hyperspectral images coding with exogenous quasi optimal transforms. In Storer and Marcellin [SM09], pages 411–419. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=4976485. IEEE

Computer Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

Bodlaender:1991:CAM

[BGK91] H. Bodlaender, T. Gonzalez, and T. Kloks. Complexity aspects of map compression. In Storer and Reif [SR91b], pages 287–296. ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1990. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=213352. IEEE

Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Bodlaender:1995:CAT

[BGK95] Hans L. Bodlaender, Teofilo F. Gonzalez, and Ton Kloks. Complexity aspects of two-dimensional data compression. Nordic Journal of Computing, 2(4):462–495, Winter 1995. CODEN NJCOFR. ISSN 1236-6064. URL http://www.cs.helsinki.

fi/njc/References/bodlaendergk1995:462.html.

B:1999:DIF

[BGM99] Praveen B., Deepak Gupta, and Rajat Moona. Design and implementation of a file system with on-the-fly data compression for GNU/Linux. Software — Practice and Experience, 29(10):863–874, August 1999. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http:

//www3.interscience.wiley.com/cgi-bin/abstract?ID=
63001367;   http://www3.interscience.wiley.com/cgi-
bin/fulltext?ID=63001367&PLACEBO=IE.pdf.

Bekesi:1996:WCA

[BGPW96a] J. Bekesi, G. Galambos, U. Pferschy, and G. J. Woeginger. Worst-case analysis for on-line data compression. Lecture Notes in Computer Science, 1120:288–300, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bekesi:1996:FGA

[BGPW96b] J. Békési, G. Galambos, U. Pferschy, and G. J. Wöginger. The fractional greedy algorithm for data compression. Computing: Archiv für informatik und numerik, 56(1):29–46, March 1996. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.springer.at/springer.py?Page=
10&Key=362&cat=300607/tocs/springer.py?Page=47&Key=
340&cat=3&id_abstract=284&id_volume=25&id_journal=
8.

Bekesi:1997:GAL

[BGPW97] József Békési, Gábor Galambos, Ulrich Pferschy, and Gerhard J. Woeginger. Greedy algorithms for on-line data compression. Journal of Algorithms, 25(2):274–289, November 1997. CODEN JOALDV. ISSN 0196-6774 (print), 1090-

2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677497908853.

Babu:2001:SMBa

[BGR01a] Shivnath Babu, Minos Garofalakis, and Rajeev Rastogi. SPARTAN: a model-based semantic compression system for massive data tables. In Sellis and Mehrotra [SM01a], pages 283–294. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Babu:2001:SMBb

[BGR01b] Shivnath Babu, Minos Garofalakis, and Rajeev Rastogi. SPARTAN: a model-based semantic compression system for massive data tables. ACM SIGMOD Record, 30(2):283–294, June 2001. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bahl:1996:SOC

[BGU96] Paramvir Bahl, Paul S. Gauthier, and Robert A. Ulichney. Software-only compression, rendering, and playback of digital video. Digital Technical Journal, 7(4):52–75, March 1996. CODEN DTJOEL. ISSN 0898-901X. URL ftp://ftp.

digital.com/pub/Digital/info/DTJ/v7n4/Softwareonly_
Compression_Ren_20apr1996DTJK04P8.ps;   http://www.
digital.com:80/info/DTJK04;   http://www.digital.com:

80/info/DTJK04/DTJK04AH.HTM;   http://www.digital.
com:80/info/DTJK04/DTJK04P8.PS;   http://www.digital.
com:80/info/DTJK04/DTJK04PF.PDF;   http://www.digital.
com:80/info/DTJK04/DTJK04SC.TXT.

Bergen:1992:DSF

[BH92] Helen A. Bergen and James M. Hogan. Data security in a fixed-model arithmetic coding compression algorithm. Computers & Security, 11(5):445–461, September 1992. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489290011F.

Barnes:1993:CVR

[BH93a] C. F. Barnes and E. J. Holder. Classified variable rate residual vector quantization applied to image subband coding. In Storer and Cohn [SC93], pages 272–281. ISBN 0-8186-3391-3 (microfiche), 0-8186-3392-1 (casebound). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1993. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=253122. IEEE Computer Society Press order number 3392-02. IEEE catalog number 93TH0536-3.

Bergen:1993:CPA

[BH93b] Helen A. Bergen and James M. Hogan. A chosen plaintext attack on an adaptive arithmetic coding compression algorithm. Computers & Security, 12(2):157–167, March 1993. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/016740489390099Q.

Bauer:2000:ISC

[BH00] R. Bauer and J. Hagenauer. Iterative source/channel-decoding using reversible variable length codes. In Storer and Cohn [SC00], pages 93–102. ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2000. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=838149. This

millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Bauer:2001:VLC

[BH01] Rainer Bauer and Joachim Hagenauer. On variable-length codes for iterative source/channel-decoding. In Storer and Cohn [SC01c], pages 273–282. ISBN 0-7695-1031-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2001. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=917158. IEEE CSP number 01PR1031.

Banihashemi:2007:DOR

[BH07] A. H. Banihashemi and A. Hatam. A distortion optimal rate allocation algorithm for transmission of embedded bitstreams over noisy channels. In Storer and Cohn [SC07], page 376. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=4148777. IEEE Computer

Society Order Number P2791.

Bonny:2010:HBC

[BH10] Talal Bonny and Jörg Henkel. Huffman-based code compression techniques for embedded processors. ACM Transactions on Design Automation of Electronic Systems (TODAES), 15 (4):31:1–31:??, September 2010. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Bredies:2012:TVB

[BH12] K. Bredies and M. Holler. A total variation-based JPEG decompression model. SIAM Journal on Imaging Sciences, 5(1): 366–393, ???? 2012. CODEN SJISBI. ISSN 1936-4954.

Bottou:1998:CAB

[BHB98] L. Bottou, P. G. Howard, and Y. Bengio. The Z-coder adaptive binary coder. In Storer and Cohn [SC98], pages 13–22. ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1998. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=672124. IEEE Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Bok:2020:PCS

[BHLY20] Kyoungsoo Bok, Jieun Han, Jongtae Lim, and Jaesoo Yoo. Provenance compression scheme based on graph patterns for large RDF documents. The Journal of Supercomputing, 76(8): 6376–6398, August 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Bi:2006:SMI

[BHS06] Dongsheng Bi, M. W. Hoffman, and K. Sayood. State machine interpretation of arithmetic codes for joint source and channel coding. In Storer and Cohn [SC06], pages 143–152. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1607249.

Baker:2019:EIQ

[BHT19] A. H. Baker, D. M. Hammerling, and T. L. Turton. Evaluating image quality measures to assess the impact of lossy data compression applied to climate simulation data. Computer Graphics Forum, 38(3):517–528, June 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bainbridge:1998:MIC

[BI98] D. Bainbridge and S. Inglis. Musical image compression. In Storer and Cohn [SC98], pages 209–218. ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1998. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=672149. IEEE Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Bi:2008:RDC

[Bi08] Qiu Yan Bi. Research on Data Compression Algorithm and Communication Network Simulation in Power System. Ph.D. thesis, Shandong University, Shandong, Peoples Republic of

China, 2008. URL http://search.proquest.com/docview/
1026704409.

Biggs:2008:CII

[Big08] Norman Biggs. Codes: an Introduction to Information Communication and Cryptography. Springer undergraduate mathematics series. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. ISBN 1-84800-272-6 (paperback). x + 273 pp. LCCN QA268 .B54 2008.

Bille:2020:ESI

[Bil20] Philip Bille. Editorial: Special issue on data compression algorithms and their applications. Algorithms (Basel), 13(1), January 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/13/1/28.

Bajaj:2001:RIC

[BIP01] Chandrajit Bajaj, Insung Ihm, and Sanghun Park. 3D RGB image compression for interactive applications. ACM Transactions on Graphics, 20(1):10–38, January 2001. CODEN AT-GRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/citations/journals/tog/2001-
20-1/p10-bajaj/.

Bist:1994:DSQ

[Bis94] A. Bist. Differential state quantization of high order Gauss– Markov process. In Storer and Cohn [SC94], pages 62–71. ISBN 0-8186-5636-0, 0-8186-5637-9 (case). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1994.

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=305913. IEEE Catalog Number 93TH0626-2. IEEE Computer Society Press Order Number 5637-02.

Bist:1995:NPQ

[Bis95] A. Bist. Narrowband process quantization. In Storer and Cohn [SC95], page 459. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

515569. IEEE catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Biskup:2008:GSH

[Bis08] M. T. Biskup. Guaranteed synchronization of Huffman codes. In Storer and Marcellin [SM08], pages 462–471. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33 INTER-

NET. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=4483324.

Boss:1995:ACI

[BJ95] R. D. Boss and E. W. Jacobs. Archetype classification in an iterated transformation image compression algorithm. In Yuval Fisher, editor, Fractal Image Compression: Theory and Application, chapter 4, pages 79–90. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 0-387-94211-4.

Breslow:2018:MFF

[BJ18] Alex D. Breslow and Nuwan S. Jayasena. Morton filters: faster, space-efficient cuckoo filters via biasing, compression, and decoupled logical sparsity. Proceedings of the VLDB Endowment, 11(9):1041–1055, May 2018. CODEN ???? ISSN 2150-8097.

Bar-Joseph:2003:CHC

[BJCO03] Ziv Bar-Joseph and Daniel Cohen-Or. Compression: Hierarchical context-based pixel ordering. Computer Graphics Forum, 22(3):349–358, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bocca:1994:ICV

[BJZ94] Jorge B. Bocca, Matthias Jarke, and Carlo Zaniolo, editors. 20th International Conference on Very Large Data Bases, September 12–15, 1994, Santiago, Chile proceedings. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1994. ISBN 1-55860-153-8. LCCN QA76.9.D3 I559 1994.

Bahl:1974:IDC

[BK74] L. R. Bahl and H. Kobayashi. Image data compression by predictive coding. II. encoding algorithms. IBM Journal of Research and Development, 18(2):172–179, March 1974. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Bookstein:1990:CIT

[BK90] Abraham Bookstein and Shmuel T. Klein. Compression, information theory, and grammars: a unified approach. ACM Transactions on Information Systems, 8(1):27–49, January 1990. CODEN ATISET. ISSN 1046-8188. URL http://www.

acm.org:80.

Bookstein:1991:FCB

[BK91] A. Bookstein and S. T. Klein. Flexible compression for bitmap sets. In Storer and Reif [SR91b], pages 402–410. ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1990. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=213340. IEEE Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Buhmann:1992:COV

[BK92] J. Buhmann and H. Kuhnel. Complexity optimized vector quantization: a neural network approach. In Storer and Cohn [SC92], pages 12–21. ISBN 0-8186-2717-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

227480. IEEE catalog number 91TH0436-6.

Bell:1993:LMS

[BK93a] Timothy C. Bell and David Kulp. Longest-match string searching for Ziv–Lempel compression. Software — Practice and Experience, 23(7):757–771, July 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bookstein:1993:HCDc

[BK93b] A. Bookstein and S. T. Klein. Is Huffman coding dead. Computing: Archiv für informatik und numerik, 50(4):279–296, December 1993. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). Also published in [BK93c, BK93d].

Bookstein:1993:HCDa

[BK93c] Abraham Bookstein and S. T. Klein. Is Huffman coding dead? In Korfhage et al. [KRW^{+}93], pages 80–87. ISBN 0-89791-605-0. LCCN ???? Also published in [BK93d, BK93b].

Bookstein:1993:HCDb

[BK93d] Abraham Bookstein and S. T. Klein. Is Huffman coding dead? In Storer and Cohn [SC93], pages 464–?? ISBN 0-8186-3391-3 (microfiche), 0-8186-3392-1 (casebound). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1993. URL http://ieeexplore.ieee.org/servlet/opac?

punumber=452. Also published in [BK93c, BK93b].

Bhaskaran:1994:CHS

[BK94] R. Bhaskaran and S. C. Kwatra. Compression of HDTV signals for low bit-rate transmission using motion compensated subband transform coding and a self-organization neural network. In Storer and Cohn [SC94], pages 195–204. ISBN 0-8186-5636-0, 0-8186-5637-9 (case). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1994.

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=305927. IEEE Catalog Number 93TH0626-2. IEEE Computer Society Press Order Number 5637-02.

Bernardini:1996:ATT

[BK96] R. Bernardini and J. Kovacevic. Arbitrary tilings of the timefrequency plane using local bases. In Storer and Cohn [SC96], pages 369–378. ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1996. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=488342. IEEE

Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Bernardini:1997:OSF

[BK97] R. Bernardini and J. Kovacevic. Orthonormal sets of filters obtained by modulations and rotations of a prototype. In Storer and Cohn [SC97], page ?? ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1997. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=582079. IEEE Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Balkenhol:2000:UDC

[BK00] B. Balkenhol and S. Kurtz. Universal data compression based on the Burrows–Wheeler transformation: theory and practice. IEEE Transactions on Computers, 49(10):1043–1053, October 2000. CODEN ITCOB4. ISSN 0018-9340 (print),

1557-9956 (electronic). URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=888040.

Bergman:2005:FDP

[BK05] E. Bergman and S. T. Klein. Fast decoding of prefix encoded texts. In Storer and Cohn [SC05], pages 143–152. ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1402175.

Ballard:2020:TPC

[BKK20] Grey Ballard, Alicia Klinvex, and Tamara G. Kolda. TuckerMPI: a parallel C++/MPI software package for largescale data compression via the Tucker tensor decomposition. ACM Transactions on Mathematical Software, 46(2): 13:1–13:31, June 2020. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL https://dl.acm.org/

doi/abs/10.1145/3378445.

Bookstein:1992:MBC

[BKR92] A. Bookstein, S. T. Klein, and T. Raita. Model based concordance compression. In Storer and Cohn [SC92], pages 82– 91. ISBN 0-8186-2717-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=227473. IEEE catalog

number 91TH0436-6.

Bookstein:1993:CRF

[BKR^{+}93] A. Bookstein, S. T. Klein, T. Raita, I. K. Ravichandra Rao, and M. D. Patil. Can random fluctuation be exploited in data compression? In Storer and Cohn [SC93], pages 70– 78. ISBN 0-8186-3391-3 (microfiche), 0-8186-3392-1 (casebound). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 1993. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=253143. IEEE Computer

Society Press order number 3392-02. IEEE catalog number 93TH0536-3.

Bookstein:1994:MMC

[BKR94] A. Bookstein, S. T. Klein, and T. Raita. Markov models for clusters in concordance compression. In Storer and Cohn [SC94], pages 116–125. ISBN 0-8186-5636-0, 0-8186-5637-9 (case). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1994. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=305919. IEEE

Catalog Number 93TH0626-2. IEEE Computer Society Press Order Number 5637-02.

Bookstein:1995:MWO

[BKR^{+}95] A. Bookstein, S. T. Klein, T. Raita, P. Lankinen, and C. W. Sze. Modeling word occurrences for the compression of concordances. In Storer and Cohn [SC95], page 462. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=515572. IEEE catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Bookstein:1997:MWO

[BKR97a] A. Bookstein, S. T. Klein, and T. Raita. Modeling word occurrences for the compression of concordances. ACM Transactions on Information Systems, 15(3):254–290, July 1997. CODEN ATISET. ISSN 1046-8188. URL http://www.acm.org:80/

tois/abstracts/bookstein.html.

Bookstein:1997:ORT

[BKR97b] A. Bookstein, S. T. Klein, and T. Raita. An overhead reduction technique for mega-state compression schemes. In Storer and Cohn [SC97], pages 367–376. ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1997. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=582061. IEEE Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Benoit:2009:MCS

[BKRSR09] Anne Benoit, Harald Kosch, Veronika Rehn-Sonigo, and Yves Robert. Multi-criteria scheduling of pipeline workflows (and application to the JPEG encoder). The International Journal of High Performance Computing Applications, 23(2):171–187, May 2009. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-

2846 (electronic). URL http://hpc.sagepub.com/content/
23/2/171.full.pdf+html.

Balkenhol:1999:MBW

[BKS99] B. Balkenhol, S. Kurtz, and Y. M. Shtarkov. Modifications of the Burrows and Wheeler data compression algorithm. In Storer and Cohn [SC99], pages 188–197. ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1999. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=755668. IEEE Computer Society Or-

der Number PR00096. IEEE Order Plan Catalog Number PR00096.

Betzel:2018:ACT

[BKS^{+}18] Filipe Betzel, Karen Khatamifard, Harini Suresh, David J. Lilja, John Sartori, and Ulya Karpuzcu. Approximate communication: Techniques for reducing communication bottlenecks in large-scale parallel systems. ACM Computing Surveys, 51 (1):1:1–1:??, April 2018. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

Bianchi:1989:DCH

[BKV89] Mark J. Bianchi, Jeffery J. Kato, and David J. Van Maren. Data compression in a half-inch reel-to-reel tape drive. Hewlett–Packard Journal, 40(3):26–31, June 1989. CODEN HPJOAX. ISSN 0018-1153.

Baby:2005:ULA

[BKV05] Thomas Baby, Youngmin Kim, and Amitabh Varshney. Unsupervised learning applied to progressive compression of timedependent geometry. Computers and Graphics, 29(3):451–461, June 2005. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Bocharova:1994:TCL

[BKYK94] I. E. Bocharova, V. D. Kolesnik, V. Yu Krachkovsky, and B. D. Kudryashov. On trellis codes for linear predictive speech compression. Lecture Notes in Computer Science, 829:115–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bao:1997:CCB

[BL97] P. G. Bao and S. W.-C. Lam. Calligraphic character boundary coding with rational B-spline based on energy minimization using genetic algorithm. In Storer and Cohn [SC97], page ?? ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1997. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=582076. IEEE

Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Burroughs:1998:DCT

[BL98] S. H. Burroughs and T. R. Lattrell. Data compression technology in ASIC cores. IBM Journal of Research and Development, 42(6):725–731, ???? 1998. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http://www.almaden.

ibm.com/journal/rd/426/burroughs.html.

Bartkowiak:2001:SAO

[BŁ01] Maciej Bartkowiak and Adam Łuczak. A simple algorithm for ordering and compression of vector codebooks. Lecture Notes in Computer Science, 2124:102–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer-ny.com/link/
service/series/0558/bibs/2124/21240102.htm;   http:
//link.springer-ny.com/link/service/series/0558/papers/
2124/21240102.pdf.

Blackstock:1987:LGE

[Bla87] Steve Blackstock. LZW and GIF explained, 1987. URL

http://www.cis.udel.edu/~amer/CISC651/lzw.and.gif.
explained.html.

Blanchette:2009:PPM

[Bla09] Jasmin Christian Blanchette. Proof pearl: Mechanizing the textbook proof of Huffman’s algorithm. Journal of Automated Reasoning, 43(1):1–18, June 2009. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL http://link.

springer.com/article/10.1007/s10817-009-9116-y.

Bornholt:2016:DBA

[BLC^{+}16] James Bornholt, Randolph Lopez, Douglas M. Carmean, Luis Ceze, Georg Seelig, and Karin Strauss. A DNA-based archival storage system. ACM SIGARCH Computer Architecture News, 44(2):637–649, May 2016. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Berners-Lee:1996:RHT

[BLFF96]  T. Berners-Lee, R. Fielding, and H. Frystyk.  RFC

1945: Hypertext Transfer Protocol — HTTP/1.0, May

1996. URL ftp://ftp.internic.net/rfc/rfc1945.txt;
 http://www.faqs.org/rfcs/rfc1945.html;   https://
www.math.utah.edu/pub/rfc/rfc1945.txt. Status: INFOR-

MATIONAL.

Blitstein:1991:DIM

[Bli91] R. Blitstein. A design to increase media-independent data integrity and availability through the use of robotic media management systems. In Storer and Reif [SR91b], page ?? ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1990. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=213332. IEEE Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Blinn:1993:WDD

[Bli93] J. F. Blinn. What’s the deal with the DCT. IEEE Computer Graphics and Applications, 13(4):78–83, July 1993. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Baek:2014:DHD

[BLNK14] Seungcheol Baek, Hyung Gyu Lee, Chrysostomos Nicopoulos, and Jongman Kim. Designing hybrid DRAM/PCM main memory systems utilizing dual-phase compression. ACM Transactions on Design Automation of Electronic Systems (TODAES), 20(1):11:1–11:??, November 2014. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Bloom:1996:NTC

[Blo96a] C. Bloom. New techniques in context modeling and arithmetic encoding. In Storer and Cohn [SC96], page ?? ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1996. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=488354. IEEE Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Bloom:1996:LND

[Blo96b] Charles R. Bloom. LZP: a new data compression algorithm. In Storer and Cohn [SC96], page 425. ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1996. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=488353. IEEE Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Bloom:1998:SPC

[Blo98] Charles R. Bloom. Solving the problems of context modeling,

1998. URL http://www.cbloom.com/papers/ppmz.zip.

Beferull-Lozano:2001:CLC

[BLO01] B. Beferull-Lozano and A. Ortega. Construction of low complexity regular quantizers for overcomplete expansions in R^{N}.

In Storer and Cohn [SC01c], pages 193–202. ISBN 0-7695-1031-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 2001. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=917150. IEEE CSP num-

ber 01PR1031.

Boucheron:2011:HSV

[BLS11] L. E. Boucheron, P. L. D. Leon, and S. Sandoval. Hybrid scalar/vector quantization of mel-frequency cepstral coefficients for low bit-rate coding of speech. In Storer and Marcellin [SM11b], pages 103–112. ISBN 1-61284-279-

8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=5749468. IEEE Computer Society Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Bhardwaj:2019:MCA

[BLSM19] Kartikeya Bhardwaj, Ching-Yi Lin, Anderson Sartor, and Radu Marculescu. Memory- and communication-aware model compression for distributed deep learning inference on IoT. ACM Transactions on Embedded Computing Systems, 18(5s): 82:1–82:??, October 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://dl.acm.org/

ft_gateway.cfm?id=3358205.

Bell:1989:NDD

[BM89] T. Bell and A. Moffat. A note on the DMC data compression scheme. The Computer Journal, 32(1):16– 20, February 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/32/1/16.full.pdf+html;
http://www3.oup.co.uk/computer_journal/hdb/Volume_
32/Issue_01/tiff/16.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_32/Issue_01/tiff/17.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_
01/tiff/18.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_01/tiff/19.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_32/Issue_01/
tiff/20.tif.

Bailey:1990:PDC

[BM90] R. L. Bailey and R. Mukkamala. Pipelining data compression algorithms. The Computer Journal, 33(4):308–313, August 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/
content/33/4/308.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_04/tiff/
308.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_04/tiff/309.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_33/Issue_04/
tiff/310.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_04/tiff/311.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_33/Issue_04/
tiff/312.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_04/tiff/313.tif.

Broder:1996:PBC

[BM96] A. Broder and M. Mitzenmacher. Pattern-based compression of text images. In Storer and Cohn [SC96], pages 300–309. ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1996. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=488335. IEEE

Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Bayrakeri:1997:TSV

[BM97] S. Bayrakeri and R. M. Mersereau. Temporally scalable video coding using nonlinear deinterlacing. In Storer and Cohn [SC97], page ?? ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1997. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=582078. IEEE

Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Bell:1998:CSM

[BM98] T. Bell and B. McKenzie. Compression of sparse matrices by arithmetic coding. In Storer and Cohn [SC98], pages 23–32. ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33

D232 1998. URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=672126. IEEE Computer Society

Press order number PR08406. IEEE order plan catalog number 98TB100225.

Bentley:1999:DCU

[BM99] J. Bentley and D. McIlroy. Data compression using long common strings. In Storer and Cohn [SC99], pages 287–295. ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1999. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=755678. IEEE Computer Society Or-

der Number PR00096. IEEE Order Plan Catalog Number PR00096.

Butterman:2003:ERB

[BM03] L. Butterman and N. Memon. An error-resilient blocksorting compression algorithm. In Storer and Cohn [SC03], page ?? ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2003. URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

1194036. IEEE Computer Society Order number PR01896.

Bocherer:2011:MDD

[BM11a] G. Böcherer and R. Mathar. Matching dyadic distributions to channels. In Storer and Marcellin [SM11b], pages 23–32. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5749460. IEEE Computer Society

Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Bonyadi:2011:NHQ

[BM11b] Mohammad Reza Bonyadi and Mohsen Ebrahimi Moghaddam. A nonuniform high-quality image compression method to preserve user-specified compression ratio. International Journal of Image and Graphics (IJIG), 11(3):355–375, July 2011. CO-DEN ???? ISSN 0219-4678.

Bilgin:2004:WCE

[BMA04] A. Bilgin, M. W. Marcellin, and M. I. Altbach. Wavelet compression of ECG signals by JPEG2000. In Storer and Cohn [SC04], page ?? ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

1281503. IEEE Computer Society Order Number: P2082.

Blalock:2018:STS

[BMG18] Davis Blalock, Samuel Madden, and John Guttag. Sprintz: Time series compression for the Internet of Things. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 2(3):1–23, September 2018. CODEN ???? ISSN 2474-9567 (electronic). URL https://dl.acm.

org/doi/abs/10.1145/3264903.

Bowery:2008:HPC

[BMH08] Jim Bowery, Matt Mahoney, and Marcus Hutter. Hutter prize for compressing human knowledge, 2008. URL http://prize.

hutter1.net.

Baker:1999:CDE

[BMM99] Brenda Baker, Udi Manber, and Robert Muth. Compressing differences of executable code. In ACM SIGPLAN Workshop on Compiler Support for System Software (WCSSS ’99), page ?? ACM Press, New York, NY 10036, USA, 1999.

Benini:2002:ODP

[BMM02] Luca Benini, Alberto Macii, and Enrico Macii. Offline data profiling techniques to enhance memory compression in embedded systems. Lecture Notes in Computer Science, 2451:314–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer-ny.com/link/
service/series/0558/bibs/2451/24510314.htm;   http:
//link.springer-ny.com/link/service/series/0558/papers/
2451/24510314.pdf.

Bell:1993:RDC

[BMNM^{+}93] Timothy C. Bell, Alistair Moffat, Craig G. Nevill-Manning, Ian H. Witten, and Justin Zobel. Research: Data compression in full-text retrieval systems. Journal of the American Society for Information Science, 44(9):508–531, October 1993. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bassiouni:1991:EIC

[BMT91] M. Bassiouni, A. Mukherjee, and N. Tzannes. Experiments on improving the compression of special data types. In Storer and Reif [SR91b], page ?? ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1990. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=213328. IEEE

Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Belazzougui:2014:AIC

[BN14] Djamal Belazzougui and Gonzalo Navarro. Alphabetindependent compressed text indexing. ACM Transactions on Algorithms, 10(4):23:1–23:??, August 2014. ISSN 1549-6325 (print), 1549-6333 (electronic).

Bhaskaran:1993:OPL

[BNK93] V. Bhaskaran, B. K. Natarajan, and K. Konstantinides. Optimal piecewise-linear compression of images. In Storer and Cohn [SC93], pages 168–177. ISBN 0-8186-3391-3 (microfiche), 0-8186-3392-1 (casebound). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1993. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=253133. IEEE Computer Society Press order number 3392-02. IEEE catalog number 93TH0536-3.

BOCU:2001:U

[BOC01] BOCU. [unknown], 2001. URL http://oss.software.ibm.

com/icu/docs/papers/binary_ordered_compression_for_
unicode.html.

Boden:1995:QDB

[Bod95] A. F. Boden. Quantization distortion in block transformcompressed data. In Storer and Cohn [SC95], page 427. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ????  URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=515537. IEEE catalog

number 95TH8037. IEEE Computer Society Press order number PR07010.

Bonneau:1995:UCO

[Bon95] Robert Bonneau. Using the centroid operator for faster multiresolution image compression and pattern recognition. Proceedings of the SPIE — The International Society for Optical Engineering, 2569:813–824, 1995. CODEN PSISDG. ISSN 0277-786X (print), 1996-756X (electronic).

Booth:1999:DTU

[Boo99] S. A. Booth. Digital TV in the United States. IEEE Spectrum, 36(3):39–41, March 1999. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Borjesson:1980:DCS

[Bör80] Per Ola Börjesson. Data compression and signal estimation with application to digitized electrocardiograms. Doctoral thesis, Lunds universitet, Lund, Sweden, 1980. 44 pp.

Born:1995:FFH

[Bor95] Günter Born. The File Formats Handbook. International Thomson Computer Press, 20 Park Plaza Suite 1001, Boston, MA 02116 USA, 1995. ISBN 1-85032-117-5, 1-85032-128-0. xvii + 1274 pp. LCCN QA76.9.F5 B67 1995. US$59.95.

Borm:2018:ACL

[Bör18] Steffen Börm. Adaptive compression of large vectors. Mathematics of Computation, 87(87):209–235, 2018. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL

http://www.ams.org/journals/mcom/2018-87-309/S0025-
5718-2017-03203-8;   http://www.ams.org/journals/mcom/
2018-87-309/S0025-5718-2017-03203-8/S0025-5718-2017-
03203-8.pdf;   https://www.ams.org/mathscinet/search/
authors.html?mrauthid=678579.

Bowyer:1970:DCT

[Bow70] Duane Ernest Bowyer. Data Compression and Transmission. Ph.D. thesis, University of Florida, Gainesville, FL, USA,

1970. 146 pp. URL http://search.proquest.com/docview/
302406194.

Boyd:1991:ESD

[Boy91] Colin Boyd. Enhancing secrecy by data compression: Theoretical and practical aspects. Lecture Notes in Computer Science, 547:266–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer-

ny.com/link/service/series/0558/bibs/0547/05470266.
htm;   http://link.springer-ny.com/link/service/series/
0558/papers/0547/05470266.pdf.

Berstel:1985:TC

[BP85] Jean Berstel and Dominique Perrin. Theory of Codes. Academic Press, New York, USA, 1985. ?? pp.

Bottou:1998:LCP

[BP98] L. Bottou and S. Pigeon. Lossy compression of partially masked still images. In Storer and Cohn [SC98], page ?? ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1998. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=672238. IEEE Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Barros:2000:VAC

[BP00] João Paulo Barros and Rui Pais. A versatile assignment in CS 2 (poster session): a file compression utility based on the Huffman code. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):185, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Biskup:2009:GSH

[BP09] M. T. Biskup and W. Plandowski. Guaranteed synchronization of Huffman codes with known position of decoder. In Storer and Marcellin [SM09], pages 33–42. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=4976447. IEEE Computer Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

Bell:2006:CC

[BPHA06] Timothy Bell, Matt Powell, Joffre Horlor, and Ross Arnold. The Canterbury Corpus, 2006. URL http://corpus.

canterbury.ac.nz.

Bell:2002:SBC

[BPMA02] T. Bell, M. Powell, A. Mukherjee, and D. Adjeroh. Searching BWT compressed text with the Boyer–Moore algorithm and binary search. In Storer and Cohn [SC02], pages 112– 121. ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2002. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=999949. IEEE

Computer Society Order Number PR01477.

Bouras:2006:IMS

[BPT06] Christos J. Bouras, Alexandros Panagopoulos, and Thrasyvoulos Tsiatsos. Implementation of 3D mesh streaming and compression techniques in NVEs. Computer Animation and Virtual Worlds, 17(2):127–140, May 2006. CODEN ???? ISSN 1546-427X.

Bajaj:1999:SRC

[BPZ99] C. L. Bajaj, V. Pascucci, and G. Zhuang. Single-resolution compression of arbitrary triangular meshes with properties. In Storer and Cohn [SC99], pages 247–256. ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1999. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=755674. IEEE Computer Society Or-

der Number PR00096. IEEE Order Plan Catalog Number PR00096.

Bellaachia:2004:FSC

[BR04] A. Bellaachia and I. A. L. Rassan. Fast searching over compressed text using a new coding technique: tagged sub-optimal code (TSC). In Storer and Cohn [SC04], page ?? ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1281502. IEEE Computer Society

Order Number: P2082.

Bischoff:2005:TCB

[BR05] U. Bischoff and J. Rossignac. TetStreamer: compressed backto-front transmission of Delaunay tetrahedra meshes. In Storer and Cohn [SC05], pages 93–102. ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=1402170.

Bernardini:2006:DCF

[BR06] R. Bernardini and R. Rinaldo. Distributed coding via folding functions. In Storer and Cohn [SC06], page 440. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1607283.

Burtscher:2007:HTC

[BR07] M. Burtscher and P. Ratanaworabhan. High throughput compression of double-precision floating-point data. In Storer and Cohn [SC07], pages 293–302. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=4148768. IEEE Computer Society Order Number P2791.

Burtscher:2009:FHS

[BR09a] M. Burtscher and P. Ratanaworabhan. FPC: a high-speed compressor for double-precision floating-point data. IEEE Transactions on Computers, 58(1):18–31, January 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=4589203.

Burtscher:2009:PPC

[BR09b] M. Burtscher and P. Ratanaworabhan. pFPC: a parallel compressor for floating-point data. In Storer and Marcellin [SM09], pages 43–52. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

4976448. IEEE Computer Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

Burtscher:2010:GST

[BR10] M. Burtscher and P. Ratanaworabhan. gFPC: a self-tuning compression algorithm. In Storer and Marcellin [SM10b], pages 396–405. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=5453485.

Brandenburg:1999:MAE

[Bra99] Karlheinz Brandenburg. MP3 and AAC explained. In ????, editor, The AES 17th International Conference, Florence, Italy, Sept. 2–5, page ?? ????, ????, 1999. URL http://www.cselt.

it/mpeg/tutorials.htm.

Bartrina-Rapesta:2008:JAR

[BRALSSMP08] J. Bartrina-Rapesta, F. Aulí-Llinás, J. Serra-Sagristà, and J. L. Monteagudo-Pereira. JPEG2000 arbitrary ROI coding through rate-distortion optimization techniques. In Storer and Marcellin [SM08], pages 292–301. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33 INTERNET.

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=
&arnumber=4483307.

Bhattar:2010:ALC

[BRD10] R. K. Bhattar, K. R. Ramakrishnan, and K. S. Dasgupta. Analysis of LDPC codes for compression of nonuniform sources with side information using density evolution. In Storer and Marcellin [SM10b], page 522. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=5453501.

Brent:1987:LAD

[Bre87] R. P. Brent. A linear algorithm for data compression. Australian Computer Journal, 19(2):64–68, May 1987. CODEN ACMJB2. ISSN 0004-8917.

Bretz:2004:SDB

[Bre04] E. A. Bretz. Speedy data backups. IEEE Spectrum, 41(1):94– 96, January 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bridges:1991:DIC

[Bri91] John Bridges. Differential image compression. Dr. Dobb’s Journal of Software Tools, 16(2):38, 40–42, 44, 46, 48, 51, February 1991. CODEN DDJOEB. ISSN 1044-789X.

Bassiouni:1988:SDC

[BRM88] M. A. Bassiouni, N. Ranganathan, and A. Mukherjee. A scheme for data compression in supercomputers. In IEEE [IEE88b], pages 272–278. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

Bartrina-Rapesta:2012:MIC

[BRNMG^{+}12] J. Bartrina-Rapesta, M. Navarro, J. Munoz-Gomez, M. W. Marcellin, J. Ruberte, and J. Serra-Sagristà. MicroCT image coding based on air filtering. In Storer and Marcellin [SM12], page 391. ISBN 1-4673-0715-7 (paperback). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ????

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=6189272. IEEE Computer Society order number P4656.

Broukhis:2008:CCC

[Bro08] Leonid A. Broukhis. The Calgary corpus Compression & SHA-1 crack Challenge, 2008. URL http://mailcom.com/ challenge. On July 2, 2010, the challenge was accepted by Alexander Rhatushnyak who sent me an entry of size 580170 (payout of $121.17 received).

Ballester-Ripoll:2019:TDI

[BRP19] Rafael Ballester-Ripoll and Renato Pajarola. Tensor decompositions for integral histogram compression and look-up. IEEE Transactions on Visualization and Computer Graphics, 25(2): 1435–1446, February 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL https:

//www.computer.org/csdl/trans/tg/2019/02/08281540-
abs.html.

Bruckheim:1969:OSO

[Bru69] Arthur Jay Bruckheim. Optimum and Sub-Optimum Compression of Second Order Markov Data. Ph.D. thesis, University of Maryland, College Park, MD, USA, 1969. 155 pp. URL

http://search.proquest.com/docview/302460376.

brucelindbloom:2007:I

[bru07] brucelindbloom. info, 2007. URL http://www.brucelindbloom.

com/.

Bernardini:2008:RCP

[BRV08] R. Bernardini, R. Rinaldo, and A. Vitali. A reliable chunkless peer-to-peer architecture for multimedia streaming. In Storer and Marcellin [SM08], pages 242–251. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33 INTER-

NET. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=4483302.

Bryan:1995:CSC

[Bry95] John Bryan. Compression scorecard: Compression’s more efficient than ever if you pick the right code. Our scorecards compare compression quality and speed of seven leading choices. Byte Magazine, 20(5):107–??, May 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Barnsley:1988:BWC

[BS88a] Michael F. Barnsley and Alan D. Sloan. A better way to compress images. Byte Magazine, 13(??):215–222, January 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Bhasker:1988:CMM

[BS88b] J. Bhasker and T. Samad. Compacting MIMOLA microcode. ACM SIGMICRO Newsletter, 19(1–2):40–44, June 1988. URL

https://dl.acm.org/doi/10.1145/62197.62206.

Brandenburg:1994:IMA

[BS94] Karlheinz Brandenburg and Gerhard Stoll. ISO-MPEG-1 audio: a generic standard for coding of high-quality digital audio. Journal of the Audio Engineering Society, 42(10):780–792, October 1994. CODEN ADIOA3. ISSN 1549-4950.

Bauer:1995:VCU

[BS95] M. Bauer and K. Sayood. Video coding using 3-dimensional DCT and dynamic code selection. In Storer and Cohn [SC95], page 451. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=515561. IEEE

catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Breinholt:1998:AGH

[BS98] Greg Breinholt and Christoph Schierz. Algorithm 781: generating Hilbert’s space-filling curve by recursion. ACM Transactions on Mathematical Software, 24(2):184–189, June 1998. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (elec-

tronic). URL http://www.acm.org:80/pubs/citations/
journals/toms/1998-24-2/p184-breinholt/.

Barros:2002:SMD

[BS02] J. Barros and S. D. Servetto. Sequencing multiple descriptions. In Storer and Cohn [SC02], page ?? ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2002. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=999991. IEEE Computer Society Order Number PR01477.

Babacan:2004:PIC

[BS04] S. D. Babacan and K. Sayood. Predictive image compression using conditional averages. In Storer and Cohn [SC04], page ?? ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1281500. IEEE Computer

Society Order Number: P2082.

Behroozi:2006:SIA

[BS06] H. Behroozi and M. R. Soleymani. Side information aware coding strategy in the quadratic Gaussian CEO problem. In Storer and Cohn [SC06], page 439. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=1607282.

Basheer:2017:CBQ

[BS17] Aseel Basheer and Kewei Sha. Cluster-based quality-aware adaptive data compression for streaming data. Journal of Data and Information Quality (JDIQ), 9(1):2:1–2:??, October 2017. CODEN ???? ISSN 1936-1955.

Basavarajaiah:2020:SCD

[BS20] Madhushree Basavarajaiah and Priyanka Sharma. Survey of compressed domain video summarization techniques. ACM Computing Surveys, 52(6):116:1–116:29, January 2020. CO-DEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (elec-

tronic). URL https://dl.acm.org/ft_gateway.cfm?id=
3355398.

Beirami:2016:PLN

[BSF16] Ahmad Beirami, Mohsen Sardari, and Faramarz Fekri. Packetlevel network compression: realization and scaling of the network-wide benefits. IEEE/ACM Transactions on Networking, 24(3):1588–1604, June 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Barbero:2010:DCJ

[BSG10] J. M. Barbero, E. Santos, and A. Gutierrez. Dual contribution of JPEG 2000 images for unidirectional links. In Storer and Marcellin [SM10b], page 521. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=5453477.

Bauer:2005:UAM

[BSS05] M. Bauer, S. M. Schuster, and K. Sayood. The use of average mutual information profile as a species signature. In Storer and Cohn [SC05], page ?? ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=1402209.

Blanes:2009:CRK

[BSS09] I. Blanes and J. Serra-Sagristà. Clustered reversible-KLT for progressive lossy-to-lossless 3 d image coding. In Storer and Marcellin [SM09], pages 233–242. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=4976467. IEEE Computer Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

Bienkowski:2018:OAF

[BSSU18] Marcin Bienkowski, Nadi Sarrar, Stefan Schmid, and Steve Uhlig. Online aggregation of the forwarding information base: Accounting for locality and churn. IEEE/ACM Transactions on Networking, 26(1):591–604, February 2018. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Baxter:1996:LCE

[BST96] J. Baxter and J. Shawe-Taylor. Learning to compress ergodic sources. In Storer and Cohn [SC96], page ?? ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1996. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=488351. IEEE Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Bentley:1986:LAD

[BSTW86] Jon Louis Bentley, Daniel D. Sleator, Robert E. Tarjan, and Victor K. Wei. A locally adaptive data compression scheme. Communications of the ACM, 29(4):320–330, April 1986. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org/pubs/toc/Abstracts/
0001-0782/5688.html.

Benedetto:1993:NMA

[BT93] J. Benedetto and A. Teolis. Nonlinear method and apparatus for coding and decoding acoustic signals with data compression and noise suppression using cochlear filters, wavelet analysis, and irregular sampling reconstruction. U.S. Patent pending, 1993.

Baig:2017:MHC

[BT17] Mohammad Haris Baig and Lorenzo Torresani. Multiple hypothesis colorization and its application to image compression. Computer Vision and Image Understanding: CVIU, 164(??):111–123, November 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1077314217300267.

Boukaram:2018:BQS

[BTLK18] Wajih Halim Boukaram, George Turkiyyah, Hatem Ltaief, and David E. Keyes. Batched QR and SVD algorithms on GPUs with applications in hierarchical matrix compression. Parallel Computing, 74(??):19–33, ???? 2018. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167819117301461.

Budihardjo:1976:DCA

[Bud76] Peter Setiadi Budihardjo. Data Compression Algorithms for Encoding of Digital Sources. [unknown], Computer Communications Network Group, University of Waterloo, Waterloo, ON, Canada, 1976. xii + 152 pp. URL http://search.

proquest.com/docview/302833917. Report E-69.

Bunton:1995:SDD

[Bun95] S. Bunton. The structure of DMC [dynamic Markov compression]. In Storer and Cohn [SC95], pages 72–81. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=515497. IEEE catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Bunton:1997:ETL

[Bun97a] S. Bunton. An executable taxonomy of on-line modeling algorithms. In Storer and Cohn [SC97], pages 42–51. ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1997. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=581959. IEEE Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Bunton:1997:GIP

[Bun97b] S. Bunton. Generalization and improvement to PPM’s ‘blending’. In Storer and Cohn [SC97], page ?? ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1997. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=582082. IEEE Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Bunton:1997:PSS

[Bun97c] S. Bunton. A percolating state selector for suffix-tree context models. In Storer and Cohn [SC97], pages 32–41. ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1997. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=581957. IEEE Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Bunton:1998:BSC

[Bun98] S. Bunton. Bayesian state combining for context models. In Storer and Cohn [SC98], pages 329–338. ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1998. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=672161. IEEE Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Busch:2009:SCH

[Bus09] Stephan Busch. Squeeze Chart 2013 home page, 2009. URL

http://www.squeezechart.com/.

Buttigieg:1995:VLE

[But95] Victor Buttigieg. Variable-Length Error-Correcting Codes. Ph.D. thesis, University of Manchester, ????, 1995. ?? pp.

Buyanovsky:1994:AC

[Buy94] George Buyanovsky. Associative coding. Monitor, 8:10–19, August 1994. Hard copies of the Russian source and English translation are available from the authors of this book. Send requests to the authors’ email address found in the Preface.

Buyanovsky:2002:PC

[Buy02] George Buyanovsky. Private communications, 2002.

Belzer:1996:STC

[BV96] B. Belzer and J. D. Villasenor. Symmetric trellis coded vector quantization. In Storer and Cohn [SC96], pages 13–22. ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1996. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=488306. IEEE Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Boldi:2004:WFC

[BV04a] Paolo Boldi and Sebastiano Vigna. The WebGraph Framework. I: Compression techniques. In Proceedings of the 13th International World Wide Web Conference (WWW 2004), pages 595–601. ACM Press, New York, NY 10036, USA, 2004.

Boldi:2004:WFI

[BV04b] Paolo Boldi and Sebastiano Vigna. The WebGraph Framework. II: Codes for the World-Wide Web. In Storer and Cohn [SC04], page ?? ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

1281504. IEEE Computer Society Order Number: P2082.

Boldi:2005:CWW

[BV05] Paolo Boldi and Sebastiano Vigna. Codes for the World–Wide Web. Internet Mathematics, 2(4):405–427, 2005. ISSN 1542-7951 (print), 1944-9488 (electronic).

Bell:1994:RBG

[BW94a] Timothy C. Bell and Ian H. Witten. The relationship between greedy parsing and symbolwise text compression. Journal of the ACM, 41(4):708–724, July 1994. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0004-5411/179892.html.

Burrows:1994:BSL

[BW94b] Michael Burrows and D. J. Wheeler. A block-sorting lossless data compression algorithm. Report 124, Digital Systems Research Center, Palo Alto, CA, USA, May 10, 1994. ?? pp.

Barnes:1995:EWZ

[BW95] C. F. Barnes and J. P. Watkins. Embedded wavelet zerotree coding with direct sum quantization structures. In Storer and Cohn [SC95], pages 252–261. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=515515. IEEE catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Bao:1998:HIC

[BW98] Paul Bao and Xiaolin Wu. Hybrid image compression scheme based on wavelet transform and adaptive context modeling. In Storer and Cohn [SC98], page ?? ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1998. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=672235. IEEE Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Bach:1999:LAT

[BW99] J. Bach and I. H. Witten. Lexical attraction for text compression. In Storer and Cohn [SC99], page ?? ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1999. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=785673. IEEE Computer Society Or-

der Number PR00096. IEEE Order Plan Catalog Number PR00096.

Baron:2010:MAL

[BW10] D. Baron and T. Weissman. An MCMC approach to lossy compression of continuous sources. In Storer and Marcellin [SM10b], pages 40–48. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=
5453443.

Bartlett:2011:DSD

[BW11] N. Bartlett and F. Wood. Deplump for streaming data. In Storer and Marcellin [SM11b], pages 363–372. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=5749494. IEEE Computer Society Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Berry:2020:IZP

[BWB^{+}20] C. J. Berry, D. Wolpert, B. Bell, A. Jatkowski, J. Surprise, G. Strevig, J. Isakson, O. Geva, B. Deskin, M. Cichanowski, G. Biran, D. Hamid, C. Cavitt, G. Fredeman, D. Chidambarrao, B. Bruen, M. Wood, S. Carey, D. Turner, and L. Sigal. IBM z15: Physical design improvements to significantly increase content in the same technology. IBM Journal of Research and Development, 64(5/6):8:1–8:12, 2020. CODEN IB-MJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Bell:1989:MTC

[BWC89] Timothy Bell, Ian H. Witten, and John G. Cleary. Modeling for text compression. ACM Computing Surveys, 21(4):557– 591, December 1989. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic). URL http://www.acm.org/

pubs/toc/Abstracts/0360-0300/76896.html.

Bilgin:2003:DCJ

[BWM03] A. Bilgin, Zhenyu Wu, and M. W. Marcellin. Decompression of corrupt JPEG2000 codestreams. In Storer and Cohn [SC03], pages 123–132. ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2003. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1194003. IEEE Computer Society Order number PR01896.

Bowen-Wright:1995:LCS

[BWS95] R. Bowen-Wright and K. Sayood. Lossless compression by simulated annealing. In Storer and Cohn [SC95], page 452. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ????  URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=515562. IEEE catalog

number 95TH8037. IEEE Computer Society Press order number PR07010.

Bentley:1976:AOA

[BY76] J. L. Bentley and A. Chi-Chih Yao. An almost optimal algorithm for unbounded searching. Information Processing Letters, 5(3):82–87, August ??, 1976. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Berger:1990:OEB

[BY90] T. Berger and R. Yeung. Optimum ‘1’-ended binary prefix codes. IEEE Transactions on Information Theory, 36(6):1435– 1441, November 1990. CODEN IETTAW. ISSN 0018-9448 (print), 1557-9654 (electronic).

Banerji:2000:AYA

[BY00] A. Banerji and En-Hui Yang. Applications of YK algorithm to the Internet transmission of Web-data: implementation issues and modifications. In Storer and Cohn [SC00], page ?? ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2000. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=838193. This millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Byun:2020:CSM

[BYJ20] Seokhyun Byun, Seunghyun Yoon, and Kyomin Jung. Comparative studies on machine learning for paralinguistic signal compression and classification. The Journal of Supercomputing, 76(10):8357–8371, October 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https://

link.springer.com/article/10.1007/s11227-020-03346-
3.

Britanak:2006:DCS

[BYR06] Vladimir Britanak, Patrick C. Yip, and Kamisetty Ramamohan Rao. Discrete Cosine and Sine Transforms: General Properties, Fast Algorithms and Integer Approximations. Academic Press, New York, USA, 2006. ?? pp.

Bajaj:1997:PIT

[BZ97] C. L. Bajaj and Guozhong Zhuang. Progressive image transmission: an adaptive quadtree-pruning approach. In Storer and Cohn [SC97], page ?? ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1997.

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=582075. IEEE Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Belloulata:2008:NOB

[BZ08] K. Belloulata and Shiping Zhu. A new object-based system for fractal video sequences compression. In Storer and Marcellin [SM08], page 508. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008;

QA76.9.D33 INTERNET. URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=4483335.

Beery:2009:MLM

[BZ09] T. A. Beery and R. Zamir. Multi level multiple descriptions. In Storer and Marcellin [SM09], pages 63–72. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4976450. IEEE Computer Society

Order Number P3592. BMS Part Number CFP09DCC-PRT.

Bai:2012:MDV

[BZL^{+}12] Huihui Bai, Mengmeng Zhang, Meiqin Liu, Anhong Wang, and Yao Zhao. Multiple description video coding using macro block level correlation of inter-/intra-descriptions. In Storer and Marcellin [SM12], page 389. ISBN 1-4673-0715-7 (paperback). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=6189270. IEEE Computer Society order number P4656.

Bilgin:1998:ELC

[BZM98] A. Bilgin, G. Zweig, and M. W. Marcellin. Efficient lossless coding of medical image volumes using reversible integer wavelet transforms. In Storer and Cohn [SC98], pages 428–437. ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN QA76.9.D33 D232 1998. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=672188. IEEE

Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Bai:2008:PET

[BZZ08] Huihui Bai, Yao Zhao, and Ce Zhu. Priority encoding transmission based multiple description video coding over packet loss network. In Storer and Marcellin [SM08], page 504. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33 INTER-

NET. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=4483331.

Connor:1993:CTR

[CA93] J. T. Connor and L. E. Atlas. Coding theory and regularization. In Storer and Cohn [SC93], pages 158–167. ISBN 0-8186-3391-3 (microfiche), 0-8186-3392-1 (casebound). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1993. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=253134. IEEE Computer Society Press order number 3392-02. IEEE catalog number 93TH0536-3.

Constantinescu:1997:FRC

[CA97] C. Constantinescu and R. Arps. Fast residue coding for lossless textual image compression. In Storer and Cohn [SC97], pages 397–406. ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1997. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=582065. IEEE

Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Constantinescu:2002:FPA

[CA02] C. Constantinescu and R. Arps. Fast peak autocorrelation finding for periodicity-exploiting compression methods. In Storer and Cohn [SC02], page ?? ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2002. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=999994. IEEE Computer Society Order Number PR01477.

Cai:2006:DIW

[CA06] Weiting Cai and Malek Adjouadi. Design and implementation of wavelet-domain video compression using multiresolution motion estimation and compensation. International Journal of Image and Graphics (IJIG), 6(4):533–549, October 2006. CODEN ???? ISSN 0219-4678.

Chu:1996:HBB

[CAC96] Chung-Tao Chu, D. Anastassiou, and Shih-Fu Chang. Hybrid block-based/segment-based video compression at very low bit rate. In Storer and Cohn [SC96], page ?? ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1996. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=488363. IEEE Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Chu:1997:MAC

[CAC97] Chung-Tao Chu, D. Anastassiou, and Shih-Fu Chang. Motionadapted content-based temporal scalability in very low bitrate video coding. In Storer and Cohn [SC97], page ?? ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1997. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=582087. IEEE Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Cachin:1998:ITM

[Cac98] Christian Cachin. An information-theoretic model for steganography. In Aucsmith [Auc98], pages 306–318. CO-DEN LNCSD9. ISBN 3-540-65386-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.A25I48

1998. URL http://link.springer-ny.com/link/service/
series/0558/bibs/1525/15250306.htm;   http://link.
springer-ny.com/link/service/series/0558/papers/1525/
15250306.pdf.

Cinque:2003:WOP

[CAL03] L. Cinque, S. De Agostino, and F. Liberati. A work-optimal parallel implementation of lossless image compression by block matching. Nordic Journal of Computing, 10(1):13–??, Spring 2003. CODEN NJCOFR. ISSN 1236-6064. Selected papers of the Prague Stringology Conference (PSC’02), September 23– 24, 2002.

Caldwell:2006:PCO

[Cal06a] Chris K. Caldwell. Prime conjectures and open questions,

2006. URL http://primes.utm.edu/notes/conjectures/.

Calgary:2006:U

[Cal06b] Calgary. [unknown], 2006. URL ftp://ftp.cpsc.ucalgary.

ca/pub/projects/text.compression.corpus.

Camana:1979:VBC

[Cam79] P. Camana. Video-bandwidth compression: a study in tradeoffs. IEEE Spectrum, 16(6):24–29, June 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Campos:2006:RC

[Cam06] Arturo San Emeterio Campos. Range coder, 2006. URL http:

//www.arturocampos.com/ac_range.html.

Capon:1959:PMR

[Cap59] J. Capon. A probabilistic model for run-length coding of pictures. IEEE Transactions on Information Theory, 5(4):157– 163, December 1959. CODEN IETTAW. ISSN 0018-9448 (print), 1557-9654 (electronic).

CapitantdeVillebonne:1978:DID

[Cap78] Patrice Jean-Marie Capitant de Villebonne. Design and Implementation of Distortion-Free Compression Techniques for LANDSAT Data and Television Images. Ph.D. thesis, Stanford University, Stanford, CA, USA, 1978. 79 pp. URL http:

//search.proquest.com/docview/302940557.

Cappellini:1985:DCE

[Cap85] Vito Cappellini, editor. Data compression and error control techniques with applications. Academic Press, New York, USA, 1985. ISBN 0-12-159260-X. xiv + 304 pp. LCCN TK5102.5 .D336 1985.

Capocelli:1989:CAR

[Cap89] Renato Capocelli. Comments and additions to ‘robust transmission of unbounded strings using Fibonacci representations’.

IEEE Transactions on Information Theory, 35(1):191–193, January 1989. CODEN IETTAW. ISSN 0018-9448 (print), 1557-9654 (electronic).

Carpentieri:1997:NLI

[Car97a] B. Carpentieri. A new lossless image compression algorithm based on arithmetic coding. Lecture Notes in Computer Science, 1311:54–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carson:1997:SPP

[Car97b] George S. Carson. Standards pipeline: PNG, VRML 97, BIIF, imaging standards. Computer Graphics, 31(3):18–20, August 1997. CODEN CGRADI, CPGPBZ. ISSN 0097-8930 (print), 1558-4569 (electronic).

Cardinal:2000:TBS

[Car00] J. Cardinal. Tree-based search for ECVQ. In Storer and Cohn [SC00], page ?? ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2000. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=838195. This

millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Cardinal:2001:DTS

[Car01] J. Cardinal. Design of tree-structured multiple description vector quantizers. In Storer and Cohn [SC01c], pages 23–32. ISBN 0-7695-1031-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2001. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=917133. IEEE

CSP number 01PR1031.

Carpentieri:2020:CNG

[Car20] Bruno Carpentieri. Compression of next-generation sequencing data and of DNA digital files. Algorithms (Basel), 13(6), June 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/13/6/151.

Cathey:1984:DCN

[Cat84] Silvia Gilbert Cathey. Data Compression for Noiseless Channels (Huffman Codes, Lattice Theory). Ph.D. thesis, Clemson University, Clemson, SC, USA, 1984. 137 pp. URL http://

search.proquest.com/docview/303289543.

Chiu:1996:VCU

[CB96] Yi-Jen Chiu and T. Berger. Video compression using fax techniques. In Storer and Cohn [SC96], page ?? ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1996. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=488357. IEEE Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Chin:1997:PRC

[CB97] Yi-Jen Chin and T. Berger. Perceptual rate control algorithms for fax-based video compression. In Storer and Cohn [SC97], page ?? ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1997. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=582086. IEEE

Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Chan:2004:ERC

[CB04] W. Chan and A. Becker. Efficient rate control for motion JPEG2000. In Storer and Cohn [SC04], page ?? ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1281505. IEEE Computer Society

Order Number: P2082.

Cortes:2000:SCR

[CBC00] Toni Cortes, Yolanda Becerra, and Raúl Cervera. Swap compression: resurrecting old ideas. Software — Practice and Experience, 30(5):567–587, April 25, 2000. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (elec-

tronic). URL http://www3.interscience.wiley.com/cgi-
bin/abstract/71004434/START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?ID=71004434&PLACEBO=IE.
pdf.

Channappayya:2008:RBS

[CBHC08] S. S. Channappayya, A. C. Bovik, R. W. Heath, and C. Caramanis. Rate bounds on SSIM index of quantized image DCT coefficients. In Storer and Marcellin [SM08], pages 352–361. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33 INTER-

NET. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=4483313.

Chang:1996:IZS

[CBK96] C. C. Chang, D. J. Buehrer, and H. C. Kowng. An improvement to Ziegler’s sparse matrix compression algorithm. The Journal of Systems and Software, 35(1):67–71, October 1996. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chandra:2014:HPM

[CBR14] Surendar Chandra, John Boreczky, and Lawrence A. Rowe. High performance many-to-many intranet screen sharing with DisplayCast. ACM Transactions on Multimedia Computing, Communications, and Applications, 10(2):19:1–19:??, February 2014. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Aldus:1988:TIF

[CC88a] Aldus Corporation and Microsoft Corporation. Tag image file format (TIFF) specification revision 5.0. Technical report, Aldus Corporation, 411 First Avenue South, Suite 200, Seattle, WA 98104, Tel: (206) 622-5500, and Microsoft Corporation, 16011 NE 36th Way, Box 97017, Redmond, WA 98073-9717, Tel: (206) 882-8080, August 8, 1988.

Aldus:tiff

[CC88b] Aldus Corporation and Microsoft Corporation. Tag image file format (TIFF) specification revision 5.0. Technical report, Aldus Corporation, 411 First Avenue South, Suite 200, Seattle, WA 98104, Tel: (206) 622-5500, and Microsoft Corporation, 16011 NE 36th Way, Box 97017, Redmond, WA 98073-9717, Tel: (206) 882-8080, August 8 1988.

Chang:1993:NLA

[CC93] Henry Ker-Chang Chang and Shing Hong Chen. A new locally adaptive data compression scheme using multilist structure. The Computer Journal, 36(6):570–578, December 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/
content/36/6/570.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/Volume_36/Issue_06/Vol36_06.
body.html#AbstractChang.

Chu:1998:AFC

[CC98] Hsueh-Ting Chu and Chaur-Chin Chen. On accelerating fractal compression. In Storer and Cohn [SC98], page ?? ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1998. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=672256. IEEE Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Cao:2000:PCR

[CC00] L. Cao and Chang Wen Chen. Product code and recurrent alternative decoding for wireless image transmission. In Storer and Cohn [SC00], page ?? ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2000. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=838194. This

millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Chang:2001:BLL

[CC01] Chin-Chen Chang and Guu-In Chen. Bi-level lossless calligraphy compression using location propagation and width inheritance. International Journal of Computer Processing of Oriental Languages (IJCPOL), 14(4):309–??, 2001. CODEN ???? ISSN 0219-4279.

Chandra:2003:TDC

[CC03] A. Chandra and K. Chakrabarty. Test data compression and test resource partitioning for system-on-a-chip using frequency-directed run-length (FDR) codes. IEEE Transactions on Computers, 52(8):1076–1088, August 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=1223641.

Chen:2006:GCD

[CC06a] Chih-Chun Chen and Jung-Hong Chuang. Geometry compression and decompression: Texture adaptation for progressive meshes. Computer Graphics Forum, 25(3):343–350, September 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chuang:2006:USF

[CC06b] J.-C. Chuang and C.-C. Chang. Using a simple and fast image compression algorithm to hide secret information. International Journal of Computer Applications, 28 (4):329–333, 2006. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.tandfonline.com/doi/full/10.
1080/1206212X.2006.11441818.

Chang:2019:FPC

[CC19] Yeim-Kuan Chang and Han-Chen Chen. Fast packet classification using recursive endpoint-cutting and bucket compression on FPGA. The Computer Journal, 62(2):198–214, February 1, 2019. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://academic.oup.com/comjnl/
article/62/2/198/5026282.

Cen:1999:DTE

[CCA99] S. Cen, P. Cosman, and F. Azadegan. Decision trees for error concealment in video decoding. In Storer and Cohn [SC99], pages 384–393. ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1999. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=755688. IEEE

Computer Society Order Number PR00096. IEEE Order Plan Catalog Number PR00096.

Chakareski:2002:CRD

[CCA02] J. Chakareski, P. A. Chou, and B. Aazhang. Computing rate-distortion optimized policies for streaming media to wireless clients. In Storer and Cohn [SC02], pages 53–62. ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2002. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=999943. IEEE

Computer Society Order Number PR01477.

Carpentieri:2019:OPL

[CCD^{+}19] Bruno Carpentieri, Arcangelo Castiglione, Alfredo De Santis, Francesco Palmieri, and Raffaele Pizzolante. One-pass lossless data hiding and compression of remote sensing data. Future Generation Computer Systems, 90(??):222–239, January 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-

7115 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167739X18306629.

Canteaut:2018:SCP

[CCF^{+}18] Anne Canteaut, Sergiu Carpov, Caroline Fontaine, Tancrède Lepoint, María Naya-Plasencia, Pascal Paillier, and Renaud Sirdey. Stream ciphers: A practical solution for efficient homomorphic-ciphertext compression. Journal of Cryptology: the journal of the International Association for Cryptologic Research, 31(3):885–916, July 2018. CO-DEN JOCREQ. ISSN 0933-2790 (print), 1432-1378 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s00145-017-9273-9.

Chaddha:1996:CRH

[CCG96] N. Chaddha, P. A. Chou, and R. M. Gray. Constrained and recursive hierarchical table-lookup vector quantization. In Storer and Cohn [SC96], pages 220–229. ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1996. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=488327. IEEE Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Chu:1994:ICC

[CCH94] Chyi-Hwa Chu, Ming-Guey Chern, and Yuang-Cheh Hsueh. Image compression by cardinality distribution. Computers and Graphics, 18(5):715–722, September–October 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

ccitt:2001:U

[cci01] ccitt. [unknown], 2001. URL http://src.doc.ic.ac.uk/

computing/ccitt/ccitt-standards/1988/.

Chang:2021:TLR

[CCKH21] Chin-Chen Chang, Jui-Feng Chang, Wei-Jiun Kao, and Ji-Hwei Horng. Two-layer reversible data hiding for VQcompressed images based on De-clustering and indicator-free search-order coding. Future Internet, 13(8):215, August 20, 2021. CODEN ???? ISSN 1999-5903. URL https://www.

mdpi.com/1999-5903/13/8/215.

Chen:1996:ABPa

[CCM96a] I-Cheng K. Chen, John T. Coffey, and Trevor N. Mudge. Analysis of branch prediction via data compression. Operating Systems Review, 30(5):128–137, December 1996. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chen:1996:ABPb

[CCM96b] I-Cheng K. Chen, John T. Coffey, and Trevor N. Mudge. Analysis of branch prediction via data compression. ACM SIGPLAN Notices, 31(9):128–137, September 1996. CO-DEN SINODQ. ISBN 0-89791-767-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//www.acm.org:80/pubs/citations/proceedings/asplos/

237090/p128-chen/. Co-published as SIGOPS Operating Systems Review 30(5), December 1996, and as SIGARCH Computer Architecture News, 24(special issue), October 1996.

Chen:2003:OAP

[CCMW03] D. Chen, Yi-Jen Chiang, N. Memon, and Xiaolin Wu. Optimal alphabet partitioning for semi-adaptive coding of sources of unknown sparse distributions. In Storer and Cohn [SC03], pages 372–381. ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2003. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1194028. IEEE Computer Society Order number PR01896.

Chen:2005:OPG

[CCMW05] Dan Chen, Yi-Jen Chiang, N. Memon, and Xiaolin Wu. Optimized prediction for geometry compression of triangle meshes. In Storer and Cohn [SC05], pages 83–92. ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=1402169.

Calhoun:2019:EFL

[CCO^{+}19] Jon Calhoun, Franck Cappello, Luke N. Olson, Marc Snir, and William D. Gropp. Exploring the feasibility of lossy compression for PDE simulations. The International Journal of High Performance Computing Applications, 33(2):397–410, March 1, 2019. CODEN IHPCFL. ISSN 1094-3420 (print),

1741-2846 (electronic). URL https://journals.sagepub.
com/doi/full/10.1177/1094342018762036.

Cappellari:2009:LLS

[CCRCK09] L. Cappellari, C. Cruz-Reyes, G. Calvagno, and J. Kari. Lossy to lossless spatially scalable depth map coding with cellular automata. In Storer and Marcellin [SM09], pages 332–341. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=4976477. IEEE Computer

Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

Chellappa:2004:DFM

[CCV04] Vijay Chellappa, P. C. Cosman, and G. M. Voelker. Dual frame motion compensation with uneven quality assignment. In Storer and Cohn [SC04], pages 262–271. ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1281471. IEEE Computer Society Order Number: P2082.

Chellappa:2005:ECD

[CCV05] V. Chellappa, P. C. Cosman, and G. M. Voelker. Error concealment for dual frame video coding with uneven quality. In Storer and Cohn [SC05], pages 319–328. ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=1402193.

Culik:1991:UFG

[CD91] K. Culik II and S. Dube. Using fractal geometry for image compression. In Storer and Reif [SR91b], page ?? ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1990. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=213302. IEEE Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Capocelli:1992:CSS

[CD92] R. Capocelli and A. De Santis. On the construction of statistically synchronizable codes. IEEE Transactions on Information Theory, 38(2):407–414, March 1992. CODEN IETTAW. ISSN 0018-9448 (print), 1557-9654 (electronic).

Capocelli:1994:BPC

[CD94] R. Capocelli and A. De Santis. Binary prefix codes ending in a ‘1’. IEEE Transactions on Information Theory, 40(4):1296– 1302, July 1994. CODEN IETTAW. ISSN 0018-9448 (print), 1557-9654 (electronic).

Cook:1995:ISS

[CD95] Gregory W. Cook and Edward J. Delp. An investigation of scalable SIMD I/O techniques with application to parallel JPEG compression. Journal of Parallel and Distributed Computing, 30(2):111–128, November 1, 1995. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

http://www.idealibrary.com/links/doi/10.1006/jpdc.
1995.1132/production;   http://www.idealibrary.com/
links/doi/10.1006/jpdc.1995.1132/production/pdf.

Cvetkovic:2000:SBO

[CD00] Z. Cvetkovic and Ingrid Daubechies. Single-bit oversampled A/D conversion with exponential accuracy in the bitrate. In Storer and Cohn [SC00], pages 343–352. ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2000. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=838174. This millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Cinque:2007:PDL

[CD07] L. Cinque and S. De Agostino. A parallel decoder for lossless image compression by block matching. In Storer and Cohn [SC07], pages 183–192. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ????

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=4148757. IEEE Computer Society Order Number P2791.

Cerra:2008:MCD

[CD08] D. Cerra and M. Datcu. A model conditioned data compression based similarity measure. In Storer and Marcellin [SM08], page 509. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008;

QA76.9.D33 INTERNET. URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=4483336.

Cerra:2009:ACC

[CD09] D. Cerra and M. Datcu. Algorithmic cross-complexity and relative complexity. In Storer and Marcellin [SM09], pages 342–351. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=4976478. IEEE

Computer Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

Cerra:2010:SMU

[CD10] D. Cerra and M. Datcu. A similarity measure using smallest context-free grammars. In Storer and Marcellin [SM10b], pages 346–355. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=5453478.

Cao:2007:SSA

[CDAM07] Minh Duc Cao, T. I. Dix, L. Allison, and C. Mears. A simple statistical algorithm for biological sequence compression. In Storer and Cohn [SC07], pages 43–52. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=4148743. IEEE Computer Society Order Number P2791.

Cheng:1996:FHR

[CDC96] J.-M. Cheng, L. M. Duyanovich, and D. J. Craft. A fast, highly reliable data compression chip and algorithm for storage systems. IBM Journal of Research and Development, 40(6): 603–613, November 1996. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http://www.almaden.

ibm.com/journal/rd40-6.html#two.

Chen:2005:ESM

[CDDM05] S. Chen, S. Diggavi, S. Dusad, and S. Muthukrishnan. Efficient string matching algorithms for combinatorial universal denoising. In Storer and Cohn [SC05], pages 153–162. ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1402176.

Cohen:2002:ICW

[CDGO02] Albert Cohen, Ingrid Daubechies, Onur G. Guleryuz, and Michael T. Orchard. On the importance of combining waveletbased nonlinear approximation with coding strategies. IEEE Transactions on Information Theory, 48(7):1895–1921, July 2002. CODEN IETTAW. ISSN 0018-9448 (print), 1557-9654 (electronic).

Cvetkovic:2002:IBF

[CDL02] Z. Cvetkovic, Ingrid Daubechies, and B. F. Logan. Interpolation of bandlimited functions from quantized irregular samples. In Storer and Cohn [SC02], pages 412–421. ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2002. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=999981. IEEE

Computer Society Order Number PR01477.

Cvetkovic:2007:SBO

[CDL07] Zoran Cvetković, Ingrid Daubechies, and Benjamin F. Logan, Jr. Single-bit oversampled A/D conversion with exponential accuracy in the bit rate. IEEE Transactions on Information Theory, 53(11):3979–3989, November 2007. CODEN IET-TAW. ISSN 0018-9448 (print), 1557-9654 (electronic). Data Compression Conference (DCC 2002), Snowbird, UT, April 02–04, 2002.

Cinque:2009:PPA

[CDL09] L. Cinque, S. De Agostino, and L. Lombardi. Practical parallel algorithms for dictionary data compression. In Storer and Marcellin [SM09], page 441. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=4976495. IEEE Computer Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

Cinque:2010:SCP

[CDL10] Luigi Cinque, Sergio De Agostino, and Luca Lombardi. Scalability and communication in parallel low-complexity lossless compression. Mathematics in Computer Science, 3(4):391–406, June 2010. CODEN ???? ISSN 1661-8270 (print), 1661-8289

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1661-8270&volume=3&issue=4&
spage=391.

Cinque:2013:BIC

[CDL13] Luigi Cinque, Sergio De Agostino, and Luca Lombardi. Binary image compression via monochromatic pattern substitution: Sequential and Parallel implementations. Mathematics in Computer Science, 7(2):155–166, June 2013. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:/

/link.springer.com/article/10.1007/s11786-013-0153-
x.

Cappello:2019:UCL

[CDL^{+}19] Franck Cappello, Sheng Di, Sihuan Li, Xin Liang, Ali Murat Gok, Dingwen Tao, Chun Hong Yoon, Xin-Chuan Wu, Yuri Alexeev, and Frederic T. Chong. Use cases of lossy compression for floating-point data in scientific data sets. The International Journal of High Performance Computing Applications, 33(6): 1201–1220, November 1, 2019. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/1094342019853336.

Cinque:2005:SLC

[CDLW05] Luigi Cinque, Sergio De Agostino, Franco Liberati, and Bart Westgeest. A simple lossless compression heuristic for grey scale images. International Journal of Foundations of Computer Science (IJFCS), 16(6):1111–??, December 2005. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Cohen:2002:SRS

[CDMW02] A. S. Cohen, S. C. Draper, E. Martinian, and G. W. Wornell. Source requantization: successive degradation and bit stealing. In Storer and Cohn [SC02], pages 102–111. ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2002. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=999948. IEEE

Computer Society Order Number PR01477.

Culik:1993:ECW

[CDR93] K. Culik II, S. Dube, and P. Rajcani. Efficient compression of wavelet coefficients for smooth and fractal-like data. In Storer and Cohn [SC93], pages 234–243. ISBN 0-8186-3391-3 (microfiche), 0-8186-3392-1 (casebound). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1993. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=253126. IEEE Computer Society Press order number 3392-02. IEEE catalog number 93TH0536-3.

Claypoole:1998:AWT

[CDSB98] R. Claypoole, G. Davis, W. Sweldens, and R. Baraniuk. Adaptive wavelet transforms for image coding using lifting. In Storer and Cohn [SC98], page ?? ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1998. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=672259. IEEE Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Chen:2011:IPB

[CDW11] Zhifeng Chen, S. Doken, and Dapeng Wu. An improved parametric bit rate model for frame-level rate control in video coding. In Storer and Marcellin [SM11b], page 451. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5749508. IEEE Computer Society

Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Charrier:2003:NVM

[CE03] C. Charrier and T. Eude. A new visual masking tool for JPEG2000. In Storer and Cohn [SC03], page ?? ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 2003. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1194037. IEEE Computer

Society Order number PR01896.

Choukse:2018:CEM

[CEA18] Esha Choukse, Mattan Erez, and Alaa Alameldeen. Compress-Points: An evaluation methodology for compressed memory systems. IEEE Computer Architecture Letters, 17(2):126–129, July/December 2018. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Cen:2009:DPL

[Cen09] Feng Cen. Design of punctured LDPC codes for ratecompatible asymmetric Slepian–Wolf coding. In Storer and Marcellin [SM09], page 438. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ????

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=4976492. IEEE Computer Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

Cespedes:1991:ACM

[Ces91] Ernesto Raul Cespedes. Adaptive compression of multisensor image data. Ph.D. thesis, Mississippi State University, Mississippi State, MS 39762, USA, 1991. 204 pp. URL http://

search.proquest.com/docview/303923204.

Chen:1996:RQI

[CF96] Qing Chen and T. R. Fischer. Robust quantization for image coding and noisy digital transmission. In Storer and Cohn [SC96], pages 3–12. ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1996. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=488305. IEEE

Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Crump:1997:ILB

[CF97] V. J. Crump and T. R. Fischer. Intraframe low bit rate video coding robust to packet erasure. In Storer and Cohn [SC97], page ?? ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1997. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=582088. IEEE

Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Chande:1998:JSC

[CF98] V. Chande and N. Farvardin. A joint source-channel coding scheme for robust image transmission. In Storer and Cohn [SC98], page ?? ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1998. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=672244. IEEE

Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Chande:1999:JSC

[CF99] V. Chande and N. Farvardin. Joint source-channel coding for progressive transmission of embedded source coders. In Storer and Cohn [SC99], pages 52–61. ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1999. URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

755654. IEEE Computer Society Order Number PR00096. IEEE Order Plan Catalog Number PR00096.

Chen:2005:TSC

[CFC05] Zhe Chen, David Dagan Feng, and Weidong Cai. Temporal and spatial compression of dynamic positron emission tomography in sinogram domain. International Journal of Image and Graphics (IJIG), 5(4):839–??, October 2005. CODEN ???? ISSN 0219-4678.

Cantone:2012:ABM

[CFG12] Domenico Cantone, Simone Faro, and Emanuele Giaquinta. Adapting Boyer–Moore-like algorithms for searching Huffman encoded texts. International Journal of Foundations of Computer Science (IJFCS), 23(2):343–356, February 2012. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Cozzolino:2012:EEM

[CFGL12] Angelo Cozzolino, Francesco Flammini, Valentina Galli, and Mariangela Lamberti. Evaluating the effects of MJPEG compression on motion tracking in metro railway surveillance. Lecture Notes in Computer Science, 7517:142–154, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33140-4_13/.

Chang:2010:PRN

[CFY^{+}10] Weiling Chang, Binxing Fang, Xiaochun Yun, Shupeng Wang, and Xiangzhan Yu. A pseudo-random number generator based on LZSS. In Storer and Marcellin [SM10b], page 524. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5453503.

Cate:1991:CCCa

[CG91a] Vincent Cate and Thomas Gross. Combining the concepts of compression and caching for a two-level filesystem. Operating Systems Review, 25(3S):200–211, April 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Cate:1991:CCCb

[CG91b] Vincent Cate and Thomas Gross. Combining the concepts of compression and caching for a two-level filesystem. ACM SIGARCH Computer Architecture News, 19(2):200–211, April 1991. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Chakareski:2003:RDO

[CG03] J. Chakareski and B. Girod. Rate-distortion optimized packet scheduling and routing for media streaming with path diversity. In Storer and Cohn [SC03], pages 203–212. ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 2003. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1194011. IEEE Computer

Society Order number PR01896.

Chakareski:2004:CRD

[CG04] J. Chakareski and B. Girod. Computing rate-distortion optimized policies for streaming media with rich acknowledgments. In Storer and Cohn [SC04], pages 202–211. ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1281465. IEEE Computer Society

Order Number: P2082.

Chow:1997:PIF

[CGC97] S. K. Chow, M. Gillies, and S. L. Chan. Parallel implementation of fractal image compression using multiple digital signal processors. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Constantinescu:2011:MDC

[CGC11] C. Constantinescu, J. Glider, and D. Chambliss. Mixing deduplication and compression on active data sets. In Storer and Marcellin [SM11b], pages 393–402. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=5749497. IEEE Computer Society Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Cosman:1991:CVQ

[CGR91] P. C. Cosman, R. M. Gray, and E. A. Riskin. Combining vector quantization and histogram equalization. In Storer and Reif [SR91b], pages 113–118. ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1990. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=213378. IEEE Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Chan:1993:JCD

[CGS93] W.-Y. Chan, A. Gersho, and S.-W. Soong. Joint codebook design for summation product-code vector quantizers. In Storer and Cohn [SC93], pages 42–51. ISBN 0-8186-3391-3 (microfiche), 0-8186-3392-1 (casebound). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1993. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=253146. IEEE Computer Society Press order number 3392-02. IEEE catalog number 93TH0536-3.

Campobello:2015:CLL

[CGSS15] G. Campobello, O. Giordano, A. Segreto, and S. Serrano. Comparison of local lossless compression algorithms for Wireless Sensor Networks. Journal of Network and Computer Applications, 47(??):23–31, January 2015. CO-DEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804514002203.

Cormack:1987:DCU

[CH87] G. V. Cormack and R. N. S. Horspool. Data compression using dynamic Markov Modelling. The Computer Journal, 30(6):541–550, December 1987. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/30/6/541.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_06/tiff/541.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_30/Issue_06/
tiff/542.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_06/tiff/543.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_30/Issue_06/
tiff/544.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_06/tiff/545.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_30/Issue_06/
tiff/546.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_06/tiff/547.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_30/Issue_06/
tiff/548.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_06/tiff/549.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_30/Issue_06/
tiff/550.tif.

Cheng:2006:NLU

[CH06] J. Cheng and Tien-Ke Huang. New lower and upper bounds on the expected length of optimal one-to-one codes. In Storer and Cohn [SC06], pages 43–52. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=1607239.

Courbet:2009:HCR

[CH09] Clement Courbet and Celine Hudelot. Hierarchies and compression: Random accessible hierarchical mesh compression for interactive visualization. Computer Graphics Forum, 28(5): 1311–1318, July 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Courbet:2011:TPM

[CH11] Clément Courbet and Céline Hudelot. Taylor prediction for mesh geometry compression. Computer Graphics Forum, 30 (1):139–151, March 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chaitin:1966:LPC

[Cha66] Gregory J. Chaitin. On the length of programs for computing finite binary sequences. Journal of the ACM, 13(4):547–569, October 1966. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Chaitin:1977:AIT

[Cha77] Gregory J. Chaitin. Algorithmic information theory. IBM Journal of Research and Development, 21(4):350–359, July 1977. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Chan:1987:NST

[Cha87] Barry Wai Kwong Chan. A new subword-tree based data compression algorithm. Thesis (M.Sc.)(C.S.), School of Computer Science, The University of New Brunswick, Fredericton, NB,

Canada, 1987. ???? pp. URL http://search.proquest.com/
docview/303653948.

Chakreyavanich:1989:RGD

[Cha89] Udomsak Chakreyavanich. Regular Grid DEM Data Compression By Using Zero-Crossing: the Automatic Breakline Detection Method. Ph.D. thesis, The Ohio State University, Columbus, OH, USA, 1989. 149 pp. URL http://search.proquest.

com/docview/303812819.

Chang:1991:EDC

[Cha91a] D. K. Chang. Exact data compression using hierarchical dictionaries. In Storer and Reif [SR91b], page ?? ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1990. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=213330. IEEE Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Chang:1991:DCU

[Cha91b] Daniel K. Chang. Data compression using hierarchical dictionaries. The Journal of Systems and Software, 15(3):233–238, July 1991. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chang:1991:UDCb

[Cha91c] Soo-Young Chang. Universal data compression by using tree models. Ph.D. thesis, The Pennsylvania State University, University Park, PA, USA, 1991. 166 pp. URL http://search.

proquest.com/docview/303950372.

Chang:1993:SPM

[Cha93] Daniel Kuo-Yee Chang. String pattern matching and lossless data compression. Ph.D. thesis, City University of New York, New York, NY, USA, 1993. 110 pp. URL http://search.

proquest.com/docview/304028781.

Charlap:1995:BFFa

[Cha95a] David Charlap. The BMP file format. part 1. Dr. Dobb’s Journal of Software Tools, 20(3):44, 46–48, 50, March 1995. CODEN DDJOEB. ISSN 1044-789X.

Charlap:1995:BFFb

[Cha95b] David Charlap. The BMP file format. part 2. Dr. Dobb’s Journal of Software Tools, 20(4):34, 36–38, 40, 42, 92–93, April 1995. CODEN DDJOEB. ISSN 1044-789X.

Chaddha:1996:MBM

[Cha96] N. Chaddha. Model-based motion compensated compression for synthetic animations. In Storer and Cohn [SC96], page ?? ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1996. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=488355. IEEE

Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Chaitin:1997:LM

[Cha97] Gregory J. Chaitin. The Limits of Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. ?? pp.

Chapin:2000:SBT

[Cha00] B. Chapin. Switching between two on-line list update algorithms for higher compression of Burrows–Wheeler transformed data. In Storer and Cohn [SC00], pages 183–192. ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2000. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=838158. This millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Chang:2004:COE

[Cha04] S. Chang. Classification oriented embedded image coding. In Storer and Cohn [SC04], page ?? ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1281506. IEEE Computer Society Order Number: P2082.

Chaitin:2006:LR

[Cha06] Gregory Chaitin. The limits of reason. Scientific American, 294(3):74–81, March 2006. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.cs.auckland.ac.nz/CDMTCS/chaitin/sciamer3.html;
 http://www.cs.auckland.ac.nz/CDMTCS/chaitin/sciamer3.
pdf;   http://www.sciam.com/sciammag/?contents=2006-
03;   http://www.scientificamerican.com/article.cfm?
id=the-limits-of-reason.

Chaitin:2007:LR

[Cha07] Gregory Chaitin. The limits of reason, 2007. URL http://

www.cs.auckland.ac.nz/CDMTCS/chaitin/sciamer3.html.

Cheng:1986:AAZ

[CHB86] L. M. Cheng, A. S. Ho, and R. E. Burge. An adaptive asteroid zonal filter for data compression. Computer Vision, Graphics, and Image Processing, 34(3):292–301, June 1986. CODEN CVGPDB. ISSN 0734-189X.

Chen:1977:DCB

[Che77] Po Hsiin Chen. Data Compression for Binary and Multi-Level Satellite Imagery. Ph.D. thesis, Purdue University, West Lafayette, IN, USA, 1977. 176 pp. URL http://search.

proquest.com/docview/302819991.

Chen:1991:EDC

[Che91] Ji Chen. The effect of data compression on image quality in high-resolution digital chest radiography. Ph.D. thesis, University of Michigan, Ann Arbot, MI, USA, 1991. 216 pp. URL

http://search.proquest.com/docview/303978189.

Chen:1999:SDC

[Che99] Xiao Chen. Study on Data Compression Coding in Multimedia Communication. D.Eng. thesis, Huazhong (Central China)

University of Science and Technology, Huazhong, Peoples Re-

public of China, 1999. URL http://search.proquest.com/
docview/1027904819.

Cheney:2000:SMT

[Che00] J. Cheney. Statistical models for term compression. In Storer and Cohn [SC00], page ?? ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2000. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=838197. This

millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Cheney:2001:CXM

[Che01] J. Cheney. Compressing XML with multiplexed hierarchical PPM models. In Storer and Cohn [SC01c], pages 163– 172. ISBN 0-7695-1031-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2001. URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

917147. IEEE CSP number 01PR1031.

Chen:2005:RDC

[Che05] Long Yan Chen. Research on data compression algorithm for isoline. M.S., Dalian University of Technology, Dalian, Peoples Republic of China, 2005. URL http://search.proquest.

com/docview/1024701902.

Cheney:2006:TXD

[Che06] J. Cheney. Tradeoffs in XML database compression. In Storer and Cohn [SC06], pages 392–401. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=1607274.

Chen:2015:RBD

[Che15] Yanlai Chen. Reduced basis decomposition: a certified and fast lossy data compression algorithm. Computers and Mathematics with Applications, 70(10):2566–2574, November 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-

7668 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0898122115004630.

Chi:1998:SML

[Chi98a] Chi-Hung Chi. Study on mult-lingual LZ77 and LZ78 text compression. In Storer and Cohn [SC98], page ?? ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1998. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=672254. IEEE Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Chiu:1998:PPS

[Chi98b] Y.-J. Chiu. A perceptual preprocessor to segment video for motion estimation. In Storer and Cohn [SC98], page ?? ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1998. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=672255. IEEE Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Chiu:1999:PBV

[Chi99] Y.-J. Chiu. A perceptual-based video coder for error resilience. In Storer and Cohn [SC99], page ?? ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1999. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=785678. IEEE Computer Society Or-

der Number PR00096. IEEE Order Plan Catalog Number PR00096.

Chen:2010:NFE

[CHMW10] Hao Chen, Ruimin Hu, Dan Mao, and Zhongyuan Wang. A novel frame error concealment algorithm based on dynamic texture synthesis. In Storer and Marcellin [SM10b], page 526. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=5453505.

Chomsky:1956:TMD

[Cho56] N. Chomsky. Three models for the description of language. IRE Transactions on Information Theory, 2(3):113–124, 1956. CODEN IRITAY. ISSN 0096-1000.

Chow:1996:SRC

[Cho96] James G. Chow. Successive refinement and the compression of multi-dimensional data. Ph.D. thesis, Cornell University, Ithaca, NY, USA, 1996. 100 pp. URL http://search.

proquest.com/docview/304303621.

Choe:2005:CET

[Cho05] Geon Ho Choe. Computational ergodic theory, volume 13 of Algorithms and computation in mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. ISBN 3-540-23121-8 (hardcover). ISSN 1431-1550. xix + 453 pp. LCCN QA313 .C485 2005.

Chien:2008:GBW

[CHSS08] Yu-Feng Chien, Wing-Kai Hon, R. Shah, and J. Scott Vitter. Geometric Burrows–Wheeler transform: Linking range searching and text indexing. In Storer and Marcellin [SM08], pages 252–261. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33

INTERNET. URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4483303.

Chiu:2010:ECT

[CHSV10] Sheng-Yuan Chiu, Wing-Kai Hon, R. Shah, and J. S. Vitter. I/O-efficient compressed text indexes: From theory to practice. In Storer and Marcellin [SM10b], pages 426–434. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=5453486.

Chu:1992:LDC

[Chu92] K. C. Chu. Lossless data compression circuit and method. U.S. Patent No. 5,150,430, September 1992.

Chung:1997:EHD

[Chu97] Kuo-Liang Chung. Efficient Huffman decoding. Information Processing Letters, 61(2):97–99, February 28, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See comment [TM98].

Chu:2002:LLD

[Chu02] A. Chu. LZAC lossless data compression. In Storer and Cohn [SC02], page ?? ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2002. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=999992. IEEE Computer Society Order Number PR01477.

Cho:2004:TSA

[CHW04] Chuan-Yu Cho, Shih-Yu Huang, and Jia-Shung Wang. A-TDB: a self adaptive block-matching algorithm for varying motion contents. In Storer and Cohn [SC04], page ?? ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1281508. IEEE Computer Society

Order Number: P2082.

Constantinescu:2007:LZC

[CI07] Sorin Constantinescu and Lucian Ilie. The Lempel–Ziv complexity of fixed points of morphisms. SIAM Journal on Discrete Mathematics, 21(2):466–481, ???? 2007. CODEN SJD-MEC. ISSN 0895-4801 (print), 1095-7146 (electronic).

Cieplinski:1998:ECM

[Cie98] Leszek Cieplinski. Entropy-constrained multiresolution vector quantisation for image coding. Fundamenta Informaticae, 34 (4):389–396, July 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Cleary:1995:IAC

[CIRM95] John G. Cleary, Sean A. Irvine, and Ingrid Rinsma-Melchert. On the insecurity of arithmetic coding. Computers & Security, 14(2):167–180, ???? 1995. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/016740489597050K.

Crochemore:2008:SAC

[CIS08] M. Crochemore, L. Ilie, and W. F. Smyth. A simple algorithm for computing the Lempel–Ziv factorization. In Storer and Marcellin [SM08], pages 482–488. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33 INTERNET.

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=
&arnumber=4483326.

Chen:2009:SWC

[CJDL09] Chao Chen, Xiangyang Ji, Qionghai Dai, and Xiaodong Liu. Slepian–Wolf coding of binary finite memory source using Burrows–Wheeler transform. In Storer and Marcellin [SM09], page 440. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=4976494. IEEE

Computer Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

Culik:1993:ICUa

[CK93a] Karel Culik II and Jarkko Kari. Image compression using weighted finite automata. Computers and Graphics, 17(3): 305–313, May–June 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Culik:1993:ICUb

[CK93b] Karel Culik II and Jarkko Kari. Image compression using weighted finite automata. Lecture Notes in Computer Science, 711:392–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Culik:1994:IDC

[CK94a] Karel Culik II and J. Kari. Image-data compression using edge-optimizing algorithm for WFA inference. Information Processing and Management, 30(6):829–838, 1994. CODEN IPMADK. ISSN 0306-4573 (print), 1873-5371 (electronic).

Culik:1994:IAW

[CK94b] Karel Culik II and Jarkko Kari. Inference algorithm for weighted finite automata and image compression. In Y. Fisher, editor, Fractal Image Encoding and Compression, page ?? Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994.

Cheung:1995:EVL

[CK95a] Kar-Ming Cheung and A. Kiely. An efficient variable length coding scheme for an IID source. In Storer and Cohn [SC95], pages 182–191. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

515508. IEEE catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Culik:1995:FSM

[CK95b] Karel Culik II and Jarkko Kari. Finite state methods for compression and manipulation of images. In Storer and Cohn [SC95], pages 142–151. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ????

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=515504. IEEE catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Cohn:1996:PSP

[CK96] M. Cohn and R. Khazan. Parsing with suffix and prefix dictionaries. In Storer and Cohn [SC96], pages 180–189. ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1996. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=488323. IEEE Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Conley:2006:CMA

[CK06] E. S. Conley and S. T. Klein. Compression of multilingual aligned texts. In Storer and Cohn [SC06], page 442. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1607285.

Conley:2013:IAB

[CK13] Ehud S. Conley and Shmuel T. Klein. Improved alignmentbased algorithm for multilingual text compression. Mathematics in Computer Science, 7(2):137–153, June 2013. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11786-012-0138-1.

Chang:2009:TIC

[CKC09] Chin-Chen Chang, Ching-Lin Kuo, and Chang-Chu Chen. Three improved codebook searching algorithms for image compression using vector quantizer. International Journal of Computer Applications, 31(1):16–22, 2009. ISSN 1206-212X (print), 1925-7074 (electronic). URL https://www.tandfonline.

com/doi/full/10.1080/1206212X.2009.11441919.

Chi:1995:EHC

[CKCW95] Chi-Hung Chi, Chi-Kwun Kan, Kwok-Shing Cheng, and Ling Wong. Extending Huffman coding for multilingual text compression. In Storer and Cohn [SC95], page 437. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=515547. IEEE catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Choueka:1985:EVH

[CKP85] Y. Choueka, Shmuel T. Klein, and Y. Perl. Efficient variants of Huffman codes in high level languages. In Proceedings of the 8th ACM-SIGIR Conference, Montréal, QC, Canada, pages 122–130. ACM Press, New York, NY 10036, USA, 1985.

Chung:1995:NAS

[CKS95] W. C. Chung, F. Kossentini, and M. J. T. Smith. A new approach to scalable video coding. In Storer and Cohn [SC95], pages 381–390. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

515528. IEEE catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Coleman:2008:RDF

[CKS08] T. P. Coleman, N. Kiyavash, and V. G. Subramanian. The rate-distortion function of a Poisson process with a queueing distortion measure. In Storer and Marcellin [SM08], pages 63–72. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33

INTERNET. URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4483284.

Curewitz:1993:PPD

[CKV93] Kenneth M. Curewitz, P. Krishnan, and Jeffrey Scott Vitter. Practical prefetching via data compression. ACM SIGMOD Record, 22(2):257–266, June 1993. CODEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Culik:1997:CSLa

[CKV97] K. Culik II, J. Kari, and V. Valenta. Compression of silhouettelike images based on WFA. In Storer and Cohn [SC97], page ?? ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1997. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=582089. IEEE

Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Chevion:1991:HEM

[CKW91] D. Chevion, E. D. Karnin, and E. Walach. High efficiency, multiplication free approximation of arithmetic coding. In Storer and Reif [SR91b], pages 43–52. ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1990. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=213372. IEEE Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Chang:1993:RCO

[CKW93] Chin-Chen Chang, Huey-Cheue Kowng, and Tzong-Chen Wu. A refinement of a compression-oriented addressing scheme. BIT (Nordisk tidskrift for informationsbehandling), 33(4):529–535, December 1993. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.mai.liu.se/BIT/contents/
bit33.html;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&volume=33&issue=4&
spage=529.

Chang:1991:ECC

[CL91] M. Chang and G. G. Langdon, Jr. Effect of coefficient coding on JPEG baseline image compression. In Storer and Reif [SR91b], page ?? ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1990. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=213331. IEEE

Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Cocurullo:1995:NAV

[CL95] F. Cocurullo and F. Lavagetto. A new algorithm for vector quantization. In Storer and Cohn [SC95], page 456. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ????  URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=515566. IEEE catalog

number 95TH8037. IEEE Computer Society Press order number PR07010.

Crochemore:1996:PMT

[CL96] Maxime Crochemore and Thierry Lecroq. Pattern-matching and text-compression algorithms. ACM Computing Surveys, 28(1):39–41, March 1996. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic). URL http://www.acm.

org/pubs/citations/journals/surveys/1996-28-1/p39-
crochemore/;   http://www.acm.org/pubs/toc/Abstracts/
surveys/234331.html.

Chan:1997:FTD

[CL97] R. K. W. Chan and M. C. Lee. A fast three dimensional Discrete Cosine Transform. In Storer and Cohn [SC97], page ?? ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1997. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=582085. IEEE

Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Cherniavsky:2006:PLD

[CL06] N. Cherniavsky and R. E. Ladner. Practical low delay broadcast of compressed variable bit rate movies. In Storer and Cohn [SC06], pages 362–371. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=1607271.

Colin:2016:CTC

[CL16] Alexei Colin and Brandon Lucia. Chain: tasks and channels for reliable intermittent programs. ACM SIGPLAN Notices, 51(10):514–530, October 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chunfeng:2012:ICA

[CLA12] Liu Chunfeng, Wang Leizhen, and Yang Aimin, editors. Information computing and applications: Third International Conference, ICICA 2012, Chengde, China, September 14–16, 2012. Proceedings Part II, volume 308 of Communications in computer and information science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 3-642-34040-7, 3-642-34041-5 (e-book). LCCN ????

Cinque:2002:PAL

[CLD02] L. Cinque, F. Liberati, and S. De Agostino. A parallel algorithm for lossless image compression by block matching. In Storer and Cohn [SC02], page ?? ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2002. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=999993. IEEE Computer Society Order Number PR01477.

Cinque:2004:SLC

[CLD04] L. Cinque, F. Liberati, and S. De Agostino. A simple lossless compression heuristic for RGB images. In Storer and Cohn [SC04], page ?? ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1281509. IEEE Computer Society Order Number: P2082.

Crochemore:2014:NGP

[CLM14] Maxime Crochemore, Alessio Langiu, and Filippo Mignosi. Note on the greedy parsing optimality for dictionary-based text compression. Theoretical Computer Science, 525(??):55–59, March 13, 2014. CODEN TCSCDI. ISSN 0304-3975 (print),

1879-2294 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304397514000309.

Coleman:2004:SNA

[CLME04] T. P. Coleman, A. H. Lee, M. Medard, and M. Effros. On some new approaches to practical Slepian–Wolf compression inspired by channel coding. In Storer and Cohn [SC04], pages 282–291. ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=1281473. IEEE

Computer Society Order Number: P2082.

Cormen:2001:IA

[CLR01] Thomas H. Cormen, Charles E. (Eric) Leiserson, and Ronald L. Rivest. Introduction to Algorithms. MIT Press, Cambridge, MA, USA, second edition, 2001. ISBN 0-262-53196-8 (paperback), 0-262-03293-7 (hardcover), 0-07-013151-1 (McGraw-Hill), 0-07-297054-5 (McGraw-Hill with CD-ROM). xxi + 1180 pp. LCCN QA76.6 .I5858 2001.

Corliss:2003:DID

[CLR03] Marc L. Corliss, E. Christopher Lewis, and Amir Roth. A DISE implementation of dynamic code decompression. ACM SIGPLAN Notices, 38(7):232–243, July 2003. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Corliss:2005:IED

[CLR05] Marc L. Corliss, E. Christopher Lewis, and Amir Roth. The implementation and evaluation of dynamic code decompression using DISE. ACM Transactions on Embedded Computing Systems, 4(1):38–72, February 2005. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Cheng:2019:CBC

[CLS19] Kin-On Cheng, Ngai-Fong Law, and Wan-Chi Siu. Clusteringbased compression for population DNA sequences. IEEE/ ACM Transactions on Computational Biology and Bioinformatics, 16(1):208–221, January 2019. CODEN ITCBCY. ISSN 1545-5963 (print), 1557-9964 (electronic).

Chang:1991:UDCa

[CM91] S.-Y. Chang and J. J. Metzner. Universal data compression algorithms by using full tree models. In Storer and Reif [SR91b], page ?? ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1990. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=213303. IEEE

Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Carpentieri:1997:NTV

[CM97] B. Carpentieri and G. Motta. A new trellis vector residual quantizer with applications to speech and image coding. In Storer and Cohn [SC97], page ?? ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1997. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=582083. IEEE Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Cabrera-Mercader:1999:RCM

[CM99] Carlos R. Cabrera-Mercader. Robust compression of multispectral remote sensing data. Ph.D. thesis, Massachusetts Institute of Technology, Cambridge, MA, USA, 1999. ???? pp. URL

http://search.proquest.com/docview/304562114.

Chou:2002:VDC

[CM02] Peter H. Chou and Teresa H. Meng. Vertex data compression through vector quantization. IEEE Transactions on Visualization and Computer Graphics, 8(4):373–382, October/December 2002. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL http://csdl.

computer.org/comp/trans/tg/2002/04/v0373abs.htm;
http://csdl.computer.org/dl/trans/tg/2002/04/v0373.
htm;   http://csdl.computer.org/dl/trans/tg/2002/04/
v0373.pdf.

Costa:2004:ERL

[CM04] M. H. M. Costa and H. S. Malvar. Efficient run-length encoding of binary sources with unknown statistics. In Storer and Cohn [SC04], page ?? ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1281510. IEEE Computer Society Order Number: P2082.

Cheng:2005:MEF

[CM05] Jimming Cheng and M. Mitzenmacher. The MARKOV EX-PERT for finding episodes in time series. In Storer and Cohn [SC05], page ?? ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=
1402211.

Cheng:2010:NEL

[Cm10] S. Cheng and R. ma. The non-existence of length-5 perfect Slepian–Wolf codes of three sources. In Storer and Marcellin [SM10b], page 528. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ????

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=
&arnumber=5453507.

Coleman:2005:TPM

[CME05] T. P. Coleman, M. Medard, and M. Effros. Towards practical minimum-entropy universal decoding. In Storer and Cohn [SC05], pages 33–42. ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=
1402164.

Coleman:2006:TSV

[CME06] T. P. Coleman, M. Medard, and M. Effros. Time-sharing vs. source-splitting in the Slepian–Wolf problem: error exponents analysis. In Storer and Cohn [SC06], pages 53–62. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1607240.

Cheng:2016:FCL

[CMK^{+}16] Long Cheng, Avinash Malik, Spyros Kotoulas, Tomas E. Ward, and Georgios Theodoropoulos. Fast compression of large Semantic Web data using X10. IEEE Transactions on Parallel and Distributed Systems, 27(9):2603–2617, September 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL https://www.computer.org/csdl/trans/td/
2016/09/07313002-abs.html.

Chen:2017:VME

[CMMS17] Jiageng Chen, Rashed Mazumder, Atsuko Miyaji, and Chunhua Su. Variable message encryption through blockcipher compression function. Concurrency and Computation: Practice and Experience, 29(7):??, April 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chui:1994:WTA

[CMP94] Charles K. Chui, Laura Montefusco, and Luigia Puccio, editors. Wavelets: Theory, Algorithms, and Applications, Proceedings of the International Conference in Chui:1994:WTA, Sicily, 14–20 October 1993. University of Messina, Academic Press, New York, USA, 1994. ISBN 0-12-174575-9.

Cheung:1999:SSV

[CMP99] G. Cheung, S. McCanne, and C. Papadimitriou. Software synthesis of variable-length code decoder using a mixture of programmed logic and table lookups. In Storer and Cohn [SC99], pages 121–130. ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1999. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=755661. IEEE

Computer Society Order Number PR00096. IEEE Order Plan Catalog Number PR00096.

Coifman:1990:SPC

[CMQW90] Ronald R. Coifman, Yves Meyer, Stephen R. Quake, and Mladen Victor Wickerhauser. Signal processing and compression with wavelet packets. Preprint, Yale University,

New Haven, April 1990. URL ftp://math.yale.edu/pub/
wavelets/cmqw.tex.

Coifman:1993:SPC

[CMQW93] Ronald R. Coifman, Yves Meyer, Stephen R. Quake, and Mladen Victor Wickerhauser. Signal processing and compression with wavelet packets. In Meyer and Roques [MR93b], pages 77–93. ISBN 2-86332-130-7. URL ftp://math.yale.

edu/pub/wavelets/cmqw.tex.

Coifman:1994:SPC

[CMQW94] Ronald R. Coifman, Yves Meyer, Stephen R. Quake, and Mladen Victor Wickerhauser. Signal processing and compression with wavelet packets. In Byrnes et al. [BBHB94], pages 363–379. ISBN 0-7923-3078-1. URL ftp://math.yale.

edu/pub/wavelets/cmqw.tex. Proceedings of the NATO Ad-

vanced Study Institute at Il Ciocco, Barga, Italy in August, 1992.

Crochemore:2000:DCU

[CMRS00] Maxime Crochemore, Filippo Mignosi, Antonio Restivo, and Sergio Salemi. Data compression using antidictionaries. Proceedings of the IEEE, 88(11):1756–1768, 2000. CODEN IEEPAD. ISSN 0018-9219 (print), 1558-2256 (electronic). A special issue on lossless data compression, edited by J. Storer.

Chou:1999:MDD

[CMW99] P. A. Chou, S. Mehrotra, and A. Wang. Multiple description decoding of overcomplete expansions using projections onto convex sets. In Storer and Cohn [SC99], pages 72–81.

ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1999. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=755656. IEEE Computer Society Or-

der Number PR00096. IEEE Order Plan Catalog Number PR00096.

Chou:2000:FPA

[CMWM00] P. A. Chou, A. Mohr, A. Wang, and S. Mehrotra. FEC and pseudo-ARQ for receiver-driven layered multicast of audio and video. In Storer and Cohn [SC00], pages 440–449. ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2000. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=838184. This millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Choi:1996:VQD

[CN96] Sunghyun Choi and Wee-Keong Ng. Vector quantizer design using genetic algorithms. In Storer and Cohn [SC96], page ?? ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1996. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=488358. IEEE

Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Crochemore:2002:IAB

[CN02] Maxime Crochemore and G. Navarro. Improved antidictionary based compression. In XII International Conference of the Chilean Computer Science Society (SCCC’02), pages 7–13. ????, ????, November 2002.

Chen:2007:NCD

[CN07] S. Chen and A. Nucci. Nonuniform compression in databases with Haar wavelet. In Storer and Cohn [SC07], pages 223–232. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=4148761. IEEE Computer

Society Order Number P2791.

Carretero:1998:CPRb

[CNCC98] J. Carretero, J. No, A. Choudhary, and P. Chen. COMPAS-SION: a parallel I/O runtime system including chunking and compression for irregular applications. Lecture Notes in Computer Science, 1457:262–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carretero:1998:CPRa

[CNPC98] J. Carretero, J. No, S.-S. Park, and A. Choudhary. COMPAS-SION: a parallel I/O runtime system including chunking and compression for irregular applications. Lecture Notes in Computer Science, 1401:668–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Claude:2012:DES

[CNS12] F. Claude, P. K. Nicholson, and D. Seco. Differentially encoded search trees. In Storer and Marcellin [SM12], pages 357–366. ISBN 1-4673-0715-7 (paperback). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

6189267. IEEE Computer Society order number P4656.

Cao:2017:DEC

[CNW^{+}17] Yi Cao, Javad Nejati, Muhammad Wajahat, Aruna Balasubramanian, and Anshul Gandhi. Deconstructing the energy consumption of the mobile page load. ACM SIGMETRICS Performance Evaluation Review, 45(1):68, June 2017. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chrysafis:1997:ECB

[CO97] C. Chrysafis and A. Ortega. Efficient context-based entropy coding for lossy wavelet image compression. In Storer and Cohn [SC97], pages 241–250. ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1997. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=582047. IEEE Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Chrysafis:1998:LBR

[CO98] C. Chrysafis and A. Ortega. Line based reduced memory, wavelet image compression. In Storer and Cohn [SC98], pages 398–407. ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1998. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=672177. IEEE

Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Coalson:2006:FFLb

[Coa06a] Josh Coalson. FLAC-free lossless audio codec, 2006. URL

http://sourceforge.net/projects/flac.

Coalson:2006:FFLa

[Coa06b] Josh Coalson. flac: free lossless audio codec, 2006. URL

http://flac.sourceforge.net/id.html.

Cody:1992:FWT

[Cod92] Mac A. Cody. The fast wavelet transform. Dr. Dobb’s Journal of Software Tools, 17(4):16–18, 20, 24, 26, 28, 100–101, April 1992. CODEN DDJOEB. ISSN 1044-789X.

Cole:1985:NPP

[Col85] A. J. Cole. A note on Peano polygons and gray codes. International Journal of Computer Mathematics, 18:3–13, 1985. CODEN IJCMAT. ISSN 0020-7160.

Cole:1986:DTB

[Col86] A. J. Cole. Direct transformations between sets of integers and Hilbert polygons. International Journal of Computer Mathematics, 20:115–122, 1986. CODEN IJCMAT. ISSN 0020-7160.

Cole:1987:CTR

[Col87] A. J. Cole. Compaction techniques for raster scan graphics using space-filling curves. The Computer Journal, 30 (1):87–92, February 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/30/1/87.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_30/Issue_01/tiff/87.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_01/tiff/
88.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_30/Issue_01/tiff/89.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_01/tiff/
90.tif;   http://www3.oup.co.uk/computer_journal/hdb/

Volume_30/Issue_01/tiff/91.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_01/tiff/
92.tif.

Cole:1993:MTF

[Col93] B. Cole. Multimedia-the technology framework. IEEE Spectrum, 30(3):32–39, March 1993. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cohen-Or:1999:DCS

[COMF99] Daniel Cohen-Or, Yair Mann, and Shachar Fleishman. Deep compression for streaming texture intensive animations. Computer Graphics, 33(Annual Conference Series):261–268, 1999. CODEN CGRADI, CPGPBZ. ISSN 0097-8930 (print), 1558-4569 (electronic). URL http://www.acm.org/pubs/

citations/proceedings/graph/311535/p261-cohen-or/.

Cook:2005:HCE

[Coo05] Robert P. Cook. Heuristic compression of an English word list. Software — Practice and Experience, 35(6):577–581, May 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cormack:1985:DCD

[Cor85] Gordon V. Cormack. Data compression on a database system. Communications of the ACM, 28(12):1336–1342, December 1985. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.org/pubs/

toc/Abstracts/0001-0782/214963.html.

Cormack:1986:DCU

[Cor86] Gordon V. Cormack. Data compression using dynamic Markov modelling. Technical report, Faculty of Mathematics, University of Waterloo, Waterloo, ON, Canada, 1986. 19 pp.

Cottingham:1985:CDS

[Cot85] Marion S. Cottingham. A compressed data structure for surface representation. Computer Graphics Forum, 4(3):217–228, September 1985. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cover:1996:UDC

[Cov96] T. M. Cover. Universal data compression and portfolio selection. In IEEE [IEE96], pages 534–538. CODEN ASF-PDV. ISBN 0-7803-3762-X (casebound), 0-8186-7594-2 (softbound), 0-8186-7596-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1996. IEEE catalog number 96CH35973. IEEE Computer Society Press order number PR07594.

Cho:2003:IER

[CP03] Sungdae Cho and W. A. Pearlman. Improved error resilient embedded video coding. In Storer and Cohn [SC03], pages 83–92. ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2003. URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

1193999. IEEE Computer Society Order number PR01896.

Castiglione:2015:CBA

[CPD^{+}15] Arcangelo Castiglione, Raffaele Pizzolante, Alfredo De Santis, Bruno Carpentieri, Aniello Castiglione, and Francesco Palmieri. Cloud-based adaptive compression and secure management services for 3D healthcare data. Future Generation Computer Systems, 43–44(??):120–134, February 2015. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167739X14001332.

Chaddha:1996:JIC

[CPG96] N. Chaddha, K. Perlmutter, and R. M. Gray. Joint image classification and compression using hierarchical tablelookup vector quantization. In Storer and Cohn [SC96], pages 23–32. ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1996. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=488307. IEEE

Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Constantinescu:2009:BSO

[CPL09] C. Constantinescu, J. Pieper, and Tiancheng Li. Block size optimization in deduplication systems. In Storer and Marcellin [SM09], page 442. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

4976496. IEEE Computer Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

Cosman:1995:TSV

[CPP95] P. C. Cosman, S. M. Perlmutter, and K. O. Perlmutter. Tree-structured vector quantization with significance map for wavelet image coding. In Storer and Cohn [SC95], pages 33– 41. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=515493. IEEE catalog

number 95TH8037. IEEE Computer Society Press order number PR07010.

Chou:2003:TTB

[CPR03] J. Chou, S. S. Pradhan, and K. Ramchandran. Turbo and trellis-based constructions for source coding with side information. In Storer and Cohn [SC03], pages 33–42. ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 2003. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1193994. IEEE Computer

Society Order number PR01896.

Chen:2008:LZF

[CPS08] Gang Chen, Simon J. Puglisi, and W. F. Smyth. Lempel– Ziv factorization using less time & space. Mathematics in Computer Science, 1(4):605–623, June 2008. CO-DEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1661-8270&volume=1&issue=4&spage=605.

Chen:2020:SDC

[CPT^{+}20] Hsing-Chung Chen, Karisma Trinanda Putra, Shian-Shyong Tseng, Chin-Ling Chen, and Jerry Chun-Wei Lin. A spatiotemporal data compression approach with low transmission cost and high data fidelity for an air quality monitoring system. Future Generation Computer Systems, 108(??):488– 500, July 2020. CODEN FGSEVI. ISSN 0167-739X (print),

1872-7115 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167739X19324197.

Chen:1991:IHI

[CR91] K. Chen and T. V. Ramabadran. An improved hierarchical interpolation (HINT) method for the reversible compression of grayscale images. In Storer and Reif [SR91b], page ?? ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1990. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=213325. IEEE Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Chen:1995:FPM

[CR95a] Shenfeng Chen and J. H. Reif. Fast pattern matching for entropy bounded text. In Storer and Cohn [SC95], pages 282– 291. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=515518. IEEE catalog

number 95TH8037. IEEE Computer Society Press order number PR07010.

Crouse:1995:JOU

[CR95b] M. Crouse and Kannan Ramchandran. JPEG optimization using an entropy-constrained quantization framework. In Storer and Cohn [SC95], pages 342–351. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=515524. IEEE catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Chen:1996:ELC

[CR96] Shenfeng Chen and J. H. Reif. Efficient lossless compression of trees and graphs. In Storer and Cohn [SC96], page ?? ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1996. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=488356. IEEE Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Chen:1997:FCV

[CR97] Shenfeng Chen and J. H. Reif. Fast and compact volume rendering in the compressed transform domain. In Storer and Cohn [SC97], pages 271–280. ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1997. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=582050. IEEE Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Candes:2006:EBL

[CR06] E. Candes and J. Romberg. Encoding the l_{p} ball from limited measurements. In Storer and Cohn [SC06], pages 33–42. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1607238.

Cooprider:2007:OCC

[CR07] Nathan Dean Cooprider and John David Regehr. Offline compression for on-chip RAM. ACM SIGPLAN Notices, 42(6): 363–372, June 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Craig:1996:CDC

[Cra96] Richard Craig. Considerations of data compression and symmetric encryption techniques for a personal communications service. Thesis (M.S.), University of Mississippi, Oxford, MS, USA, 1996. various pp.

Craft:1998:FHD

[Cra98a] D. J. Craft. A fast hardware data compression algorithm and some algorithmic extensions. IBM Journal of Research and Development, 42(6):733–745, ???? 1998. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http:/

/www.almaden.ibm.com/journal/rd/426/craft.html.

Craft:1998:SPP

[Cra98b] D. J. Craft. Simple pre-processors significantly improve LZ 1 compression. In Storer and Cohn [SC98], page ?? ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1998. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=672261. IEEE Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Craft:1998:SPS

[Cra98c] David J. Craft. Simple pre-processors significantly improve LZ 1 compression. In Storer and Cohn [SC98], pages 538–?? ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1998. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=672118. IEEE Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Corvetto:2010:RDL

[CRA10] A. Corvetto, A. Ruedin, and D. Acevedo. Robust detection and lossless compression of the foreground in magnetic resonance images. In Storer and Marcellin [SM10b], page 529. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=5453506.

Cressman:1994:ADC

[Cre94] David C. Cressman. Analysis of data compression in the DLT2000 tape drive. Digital Technical Journal, 6 (2):62–71, Spring 1994. CODEN DTJOEL. ISSN 0898-

901X. URL ftp://ftp.digital.com/pub/Digital/info/
DTJ/v6n2/Analysis_of_Data_Compression_i_01oct1994DTJE05P8.
ps;   http://www.digital.com:80/info/DTJE05/DTJE05SC.
TXT.

Creusere:1996:RIC

[Cre96] C. D. Creusere. Robust image coding using the embedded zerotree wavelet algorithm. In Storer and Cohn [SC96], page ?? ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1996. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=488364. IEEE

Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Creusere:1997:LMC

[Cre97] C. D. Creusere. Out-of-loop motion compensation for reduced complexity video encoding. In Storer and Cohn [SC97], page ?? ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1997. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=582084. IEEE

Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Creusere:1998:SCR

[Cre98] C. D. Creusere. Successive coefficient refinement for embedded lossless image compression. In Storer and Cohn [SC98], page ?? ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1998. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=672262. IEEE

Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Creusere:2002:APA

[Cre02] C. D. Creusere. An analysis of perceptual artifacts in MPEG scalable audio coding. In Storer and Cohn [SC02], pages 152– 161. ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2002. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=999953. IEEE

Computer Society Order Number PR01477.

Chen:2008:ISA

[CRN08] Su Chen, S. Ranjan, and A. Nucci. IPzip: a stream-aware IP compression algorithm. In Storer and Marcellin [SM08], pages 182–191. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33

INTERNET. URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4483296.

Crocker:1995:PPN

[Cro95] Lee Daniel Crocker. PNG: The Portable Network Graphic format. Dr. Dobb’s Journal of Software Tools, 20(7):36, 38, 40, 42, 44, July 1995. CODEN DDJOEB. ISSN 1044-789X.

Chandramouli:1998:ECM

[CRR98] R. Chandramouli, N. Ranganathan, and S. J. Ramadoss. Empirical channel matched quantizer design and UEP for robust image transmission. In Storer and Cohn [SC98], page ?? ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1998. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=672245. IEEE Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Carpentieri:1992:SMP

[CS92] B. Carpentieri and J. A. Storer. A split-merge parallel blockmatching algorithm for video displacement estimation. In Storer and Cohn [SC92], pages 239–248. ISBN 0-8186-2717-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=227457. IEEE catalog number 91TH0436-6.

Constantinescu:1993:LAV

[CS93] C. Constantinescu and J. A. Storer. On-line adaptive vector quantization with variable size codebook entries. In Storer and Cohn [SC93], pages 32–41. ISBN 0-8186-3391-3 (microfiche), 0-8186-3392-1 (casebound). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1993. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=253147. IEEE Computer Society Press order number 3392-02. IEEE catalog number 93TH0536-3.

Constantinescu:1994:ITS

[CS94a] C. Constantinescu and J. A. Storer. Improved techniques for single-pass adaptive vector quantization. Proceedings of the IEEE, 82(6):933–939, June 1994. CODEN IEEPAD. ISSN 0018-9219 (print), 1558-2256 (electronic).

Constantinescu:1994:OAV

[CS94b] C. Constantinescu and J. A. Storer. Online adaptive vector quantization with variable size codebook entries. Information Processing and Management, 30(6):745–758, 1994. CODEN IPMADK. ISSN 0306-4573 (print), 1873-5371 (electronic).

Constantinescu:1995:ASP

[CS95] C. Constantinescu and J. A. Storer. Application of single-pass adaptive VQ to bilevel images. In Storer and Cohn [SC95], page 423. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=515533. IEEE

catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Christiansen:2002:FMD

[CS02] B. O. Christiansen and K. E. Schauser. Fast motion detection for thin client compression. In Storer and Cohn [SC02], pages 332–341. ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2002. URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

999971. IEEE Computer Society Order Number PR01477.

Chen:2005:DJS

[CS05] Qingyu Chen and K. P. Subbalakshmi. Distributed joint source-channel decoding for correlated Markov sources. In Storer and Cohn [SC05], page ?? ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=1402210.

Cuzzocrea:2013:SIP

[CS13] Alfredo Cuzzocrea and Domenico Saccà. Special issue papers: Exploiting compression and approximation paradigms for effective and efficient online analytical processing over sensor network readings in data grid environments. Concurrency and Computation: Practice and Experience, 25(14):2016–2035, September 25, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chang:2000:APP

[CSCW00] Yung-Ching Chang, Bin-Kai Shyu, Chuan-Yu Cho, and Jia-Shung Wang. Adaptive post-processing for region-based fractal image compression. In Storer and Cohn [SC00], page ?? ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2000. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=838196. This millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Chen:1977:FCA

[CSF77] Wen-Hsiung Chen, C. Harrison Smith, and S. C. Fralick. A fast computational algorithm for the discrete cosine transform. IEEE Transactions on Communications, 25(9):1004– 1009, September 1977. CODEN IECMBT. ISSN 0090-6778 (print), 1558-0857 (electronic).

Christiansen:2000:NCT

[CSM00] B. O. Christiansen, K. E. Schauser, and M. Munke. A novel codec for thin client computing. In Storer and Cohn [SC00], pages 13–22. ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2000. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=838141. This

millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Christiansen:2001:STC

[CSM01] B. O. Christiansen, K. E. Schauser, and M. Munke. Streaming thin client compression. In Storer and Cohn [SC01c], pages 223–232. ISBN 0-7695-1031-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2001. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=917153. IEEE CSP number 01PR1031.

Cho:2005:CWS

[CSP05] Yushin Cho, A. Said, and W. A. Pearlman. Coding the wavelet spatial orientation tree with low computational complexity. In Storer and Cohn [SC05], page ?? ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=1402212.

Chobey:2011:NPM

[CST11] D. K. Chobey, A. Sharma, and A. K. Tiwari. A novel prediction model for lossless video compression. In Storer and Marcellin [SM11b], page 452. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ????

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=5749509. IEEE Computer Society Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Cleary:1995:EZF

[CT95] J. G. Cleary and W. J. Teahan. Experiments on the zero frequency problem. In Storer and Cohn [SC95], page 480. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ????  URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=515590. IEEE catalog

number 95TH8037. IEEE Computer Society Press order number PR07010.

Cleary:1997:ULC

[CT97] John G. Cleary and W. J. Teahan. Unbounded length contexts for PPM. The Computer Journal, 40(2–3):67–75, ???? 1997. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/40/2_and_3/67.full.pdf+html;   http://
www.oup.co.uk/computer_journal/Volume_40/Issue_02/
Vol40_02.body.html#AbstractCleary;   http://www3.oup.
co.uk/computer_journal/Volume_40/Issue_02/Vol40_02.
body.html#AbstractCleary.

Chapin:1998:HCB

[CT98a] B. Chapin and S. R. Tate. Higher compression from the Burrows–Wheeler transform by modified sorting. In Storer and Cohn [SC98], page ?? ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1998. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=672253. IEEE

Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Chung:1998:JSC

[CT98b] S.-Y. Chung and M. D. Trott. Joint source-channel coding using space-filling curves for bandwidth compression. In Storer and Cohn [SC98], page ?? ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1998. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=672258. IEEE Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Cleary:1999:OIP

[CT99] J. G. Cleary and W. J. Teahan. An open interface for probabilistic models of text. In Storer and Cohn [SC99], page ?? ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1999. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=785679. IEEE Computer Society Or-

der Number PR00096. IEEE Order Plan Catalog Number PR00096.

Cover:2006:EIT

[CT06] Thomas M. Cover and Joy A. Thomas. Elements of Information Theory. Wiley-Interscience, New York, NY, USA, second edition, 2006. ISBN 0-471-74881-1 (e-book), 0-471-24195-4 (print), 0-471-74882-X. xxiii + 748 pp. LCCN Q360 .C68 2006.

URL http://www.loc.gov/catdir/enhancements/fy0624/
2005047799-d.html;   http://www.loc.gov/catdir/enhancements/
fy0624/2005047799-t.html;   http://www.loc.gov/catdir/
enhancements/fy0826/2005047799-b.html.

Celikcan:2009:OSC

[CT09] U. Celikcan and E. Tuncel. Optimized source-channel coding of video signals in packet loss environments. In Storer and Marcellin [SM09], page 437. ISBN 0-7695-3592-

5. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=4976491. IEEE Computer Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

Chen:2009:THC

[CTC^{+}09] D. M. Chen, S. S. Tsai, V. Chandrasekhar, G. Takacs, J. Singh, and B. Girod. Tree histogram coding for mobile image matching. In Storer and Marcellin [SM09], pages 143–152. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4976458. IEEE Computer Society

Order Number P3592. BMS Part Number CFP09DCC-PRT.

Chen:2010:IIC

[CTC^{+}10] D. M. Chen, S. S. Tsai, V. Chandrasekhar, G. Takacs, R. Vedantham, R. Grzeszczuk, and B. Girod. Inverted index compression for scalable image matching. In Storer and Marcellin [SM10b], page 525. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ????

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=
&arnumber=5453502.

Chen:2007:MDC

[CTD07] Jun Chen, Chao Tian, and S. Diggavi. Multiple description coding for stationary and ergodic sources. In Storer and Cohn [SC07], pages 73–82. ISBN 0-7695-2791-

4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=4148746. IEEE Computer Society Order Number P2791.

Cleary:1995:ULC

[CTW95] John G. Cleary, W. J. Teahan, and Ian H. Witten. Unbounded length contexts for PPM. In Storer and Cohn [SC95], pages 52– 61. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=515495. IEEE catalog

number 95TH8037. IEEE Computer Society Press order number PR07010.

Cushing:1990:DCA

[Cus90] John Aikin Cushing. Data compression algorithms for English text based on word frequency. Working paper, Institute of Economic Research, Kobe University of Commerce, Kobe, Japan, 1990. 57 pp.

Culik:1996:FAB

[CV96] Karel Culik II and V. Valenta. Finite automata based compression of bi-level images. In Storer and Cohn [SC96], pages 280–289. ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1996. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=488333. IEEE

Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Culik:1997:FAB

[CV97] Karel Culik II and Vladimir Valenta. Finite automata based compression of bi-level and simple color images. Computers and Graphics, 21(1):61–68, January–February 1997. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.elsevier.com/cgi-bin/cas/
tree/store/cag/cas_sub/browse/browse.cgi?year=1997&
volume=21&issue=1&aid=9600070.

Chaddha:1995:HVQ

[CVC95] Navin Chaddha, Mohan Vishwanath, and P. A. Chou. Hierarchical vector quantization of perceptually weighted block transforms. In Storer and Cohn [SC95], pages 3–12. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ????  URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=515490. IEEE catalog

number 95TH8037. IEEE Computer Society Press order number PR07010.

Caillaud:2016:GPC

[CVDL16] F. Caillaud, V. Vidal, F. Dupont, and G. Lavoué. Geometry: Progressive compression of arbitrary textured meshes. Computer Graphics Forum, 35(7):475–484, October 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cvetkovic:1998:ASC

[Cve98] Z. Cvetkovic. Accurate subband coding with low resolution quantization. In Storer and Cohn [SC98], pages 448–457. ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1998. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=672192. IEEE Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Cvetkovic:1999:SCQ

[Cve99] Z. Cvetkovic. Source coding with quantized redundant expansions: accuracy and reconstruction. In Storer and Cohn [SC99], pages 344–353. ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1999. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=755684. IEEE

Computer Society Order Number PR00096. IEEE Order Plan Catalog Number PR00096.

Culik:1997:CSLb

[CVK97] Karel Culik II, Vladimir Valenta, and Jarkko Kari. Compression of silhouette-like images based on WFA. J.UCS: The Journal of Universal Computer Science, 3(10):1100–1113, October 28, 1997. CODEN ???? ISSN 0948-695X (print), 0948-6968

(electronic). URL http://medoc.springer.de:8000/jucs/
jucs_3_10/compression_of_silhouette_like.

Cleary:1984:DCU

[CW84] John G. Cleary and I. H. Witten. Data compression using adaptive coding and partial string matching. IEEE Transactions on Communications, COM-32(4):396–402, April 1984. CODEN IECMBT. ISSN 0090-6778 (print), 1558-0857 (electronic).

Chang:1991:LOP

[CW91] Chin-Chen Chang and Tzong-Chen Wu. A letter-oriented perfect hashing scheme based upon Sparse table compression. Software — Practice and Experience, 21(1):35–49, January 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chang:1997:LAD

[CW97] Chin-Chen Chang and Chih-Hung Wang. A locally adaptive data compression strategy for Chinese–English characters. The Journal of Systems and Software, 36(2):167–179, February 1997. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Cannane:2001:GPC

[CW01] Adam Cannane and Hugh E. Williams. General-purpose compression for efficient retrieval. Journal of the American Society for Information Science and Technology: JASIST, 52(5):430– 437, ???? 2001. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Cannane:2002:GPC

[CW02] Adam Cannane and Hugh E. Williams. A general-purpose compression scheme for large collections. ACM Transactions on Information Systems, 20(3):329–355, 2002. CODEN ATISET. ISSN 1046-8188.

Cumming:2003:CRD

[CW03] I. Cumming and Jing Wang. Compression of RADARSAT data with block adaptive wavelets. In Storer and Cohn [SC03], page ?? ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2003. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1194038. IEEE Computer Society Order number PR01896.

Cameron:2007:CCS

[CW07] Michael Cameron and Hugh Williams. Comparing compressed sequences for faster nucleotide BLAST searches. IEEE/ACM Transactions on Computational Biology and Bioinformatics, 4 (3):349–364, July 2007. CODEN ITCBCY. ISSN 1545-5963 (print), 1557-9964 (electronic).

Chlopkowski:2015:GPL

[CW15] Marek Chlopkowski and Rafal Walkowiak. A general purpose lossless data compression method for GPU. Journal of Parallel and Distributed Computing, 75(??):40–52, January 2015. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731514001956.

Chen:1999:MEF

[CWL99] Hong-Chung Chen, Yue-Li Wang, and Yu-Feng Lan. A memory-efficient and fast Huffman decoding algorithm. Information Processing Letters, 69(3):119–122, February 12, 1999. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic). URL http://www.elsevier.nl/gej-ng/10/23/20/
47/19/20/abstract.html;   http://www.elsevier.nl/gej-
ng/10/23/20/47/19/20/article.pdf.

Cheng:2015:CMD

[CWLS15] Kin-On Cheng, Paula Wu, Ngai-Fong Law, and Wan-Chi Siu. Compression of multiple DNA sequences using intra-sequence and inter-sequence similarities. IEEE/ACM Transactions on Computational Biology and Bioinformatics, 12(6):1322–1332, November 2015. CODEN ITCBCY. ISSN 1545-5963 (print), 1557-9964 (electronic).

Chon:2011:ICV

[CWRL11] Jaehong Chon, S. Whittle, E. A. Riskin, and R. E. Ladner. Improving compressed video sign language conversations in the presence of data loss. In Storer and Marcellin [SM11b], pages 383–392. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=5749496. IEEE

Computer Society Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Carpentieri:2000:LCC

[CWS00] B. Carpentieri, M. J. Weinberger, and G. Seroussi. Lossless compression of continuous-tone images. Proceedings of the IEEE, 88(11):1797–1809, November 2000. CODEN IEEPAD. ISSN 0018-9219 (print), 1558-2256 (electronic).

Cui:2020:DFR

[CWW20] Tingting Cui, Wei Wang, and Meiqin Wang. Distinguisher on full-round compression function of GOST R. Information Processing Letters, 156(??):Article 105902, April 2020. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0020019019301851.

Chen:2010:ICU

[CWY^{+}10] Zhuoyuan Chen, Jiangtao Wen, Shiqiang Yang, Yuxing Han, and J. D. Villasenor. Image compression using the DCT and noiselets: a new algorithm and its rate distortion performance. In Storer and Marcellin [SM10b], page 527. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5453504.

Chau:1991:SCB

[CWYW91] W. K. Chau, S. K. M. Wong, X. D. Yang, and S. J. Wan. On the selection of color basis for image compression. In Storer and Reif [SR91b], page ?? ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1990. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=213320. IEEE

Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Cannane:1999:GPC

[CWZ99] A. Cannane, H. E. Williams, and J. Zobel. A generalpurpose compression scheme for databases. In Storer and Cohn [SC99], page ?? ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1999. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=785676. IEEE

Computer Society Order Number PR00096. IEEE Order Plan Catalog Number PR00096.

Cheng:2004:SRW

[CX04] S. Cheng and Zixiang Xiong. Successive refinement for the Wyner–Ziv problem and layered code design. In Storer and Cohn [SC04], page ?? ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

1281507. IEEE Computer Society Order Number: P2082.

Chen:2012:CGT

[CXF12] Minjie Chen, Mantao Xu, and P. Franti. Compression of GPS trajectories. In Storer and Marcellin [SM12], pages 62–71. ISBN 1-4673-0715-7 (paperback). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

6189237. IEEE Computer Society order number P4656.

Changjun:2012:SFS

[CXYQ12] Fu Changjun, Ji Xiangyang, Zhang Yongbing, and Dai Qionghai. A single frame super-resolution method based on matrix completion. In Storer and Marcellin [SM12], pages 297–306. ISBN 1-4673-0715-7 (paperback). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

6189261. IEEE Computer Society order number P4656.

Changjun:2012:PVE

[CXZD12] Fu Changjun, Ji Xiangyang, Yongbing Zhang, and Qionghai Dai. Packet video error concealment based on compressed sensing and regularized least squares. In Storer and Marcellin [SM12], page 394. ISBN 1-4673-0715-7 (paperback). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=6189275. IEEE Computer Society order number P4656.

Chou:1991:UMP

[CY91] W.-K. Chou and D. Y. Y. Yun. A uniform model for parallel Fast Fourier Transform (FFT) and Fast Discrete Cosine Transform (FDCT). In Storer and Reif [SR91b], page ?? ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1990. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=213304. IEEE Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Chang:1992:CDC

[CY92] Kuang-Yao Chang and Gregory T. Yang. CCIGS: a data compression system for Chinese fonts and binary images using classification techniques. Software — Practice and Experience, 22 (12):1027–1047, December 1992. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chen:1994:CCI

[CYH94] Yung-Sheng S. Chen, Hung-Tien T. Yen, and Wen-Hsing H. Hsu. Compression of color image via the technique of surface fitting. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 56(3): 272–279, May 1994. CODEN CGMPE5. ISSN 1049-

9652. URL http://www.idealibrary.com/links/artid/
cgip.1994.1024/production;   http://www.idealibrary.
com/links/artid/cgip.1994.1024/production/pdf.

Chen:2008:DIH

[CYL^{+}08] Xi Chen, Lei Yang, H. Lekatsas, R. P. Dick, and Li Shang. Design and implementation of a high-performance microprocessor cache compression algorithm. In Storer and Marcellin [SM08], pages 43–52. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008;

QA76.9.D33 INTERNET. URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=4483282.

Calude:1998:TNL

[CZ98a] Cristian Calude and Tudor Zamfirescu. The typical number is a lexicon. New Zealand Journal of Mathematics, 27:7–13, 1998. ISSN 1171-6096 (print), 1179-4984 (electronic).

Chan:1998:FPL

[CZ98b] T. F. Chan and H. M. Zhou. Feature preserving lossy image compression using nonlinear PDEs. In Storer and Cohn [SC98], page ?? ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1998. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=672240. IEEE

Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Chi:1999:DCM

[CZ99] C.-H. Chi and Y. Zhang. Design consideration for multilingual cascading text compressors. In Storer and Cohn [SC99], page ?? ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN QA76.9.D33 D37 1999. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=785677. IEEE

Computer Society Order Number PR00096. IEEE Order Plan Catalog Number PR00096.

Chen:2020:VLC

[CZAR20] Tian Chen, Yongsheng Zuo, Xin An, and Fuji Ren. A variablelength compatible compression scheme based on tristate signals. The Journal of Supercomputing, 76(2):1020–1033, February 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Chen:1992:AHS

[CZS92] O. T.-C. Chen, Z. Zhang, and B. J. Sheu. An adaptive high-speed lossy data compression. In Storer and Cohn [SC92], pages 349–358. ISBN 0-8186-2717-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

227446. IEEE catalog number 91TH0436-6.

Chen:2011:ITF

[CZT^{+}11] Hao Chen, Ye Zhang, Yu Tao, Bin Zou, and Wenyan Tang. An improved temporal frame interpolation algorithm for H.264 video compression. In Storer and Marcellin [SM11b], page 449. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=5749506. IEEE

Computer Society Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Chen:2019:LAC

[CZZR19] Siguang Chen, Shujun Zhang, Xiaoyao Zheng, and Xiukai Ruan. Layered adaptive compression design for efficient data collection in industrial wireless sensor networks. Journal of Network and Computer Applications, 129(??):37–45, March 1, 2019. CODEN JNCAF3. ISSN 1084-8045 (print), 1095-

8592 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1084804519300025.

Chen:2011:FVQ

[CZZZ11] Yushi Chen, Yuhhang Zhang, Ye Zhang, and Zhixin Zhou. Fast vector quantization algorithm for hyperspectral image compression. In Storer and Marcellin [SM11b], page 450. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5749507. IEEE Computer Society

Order Number P4352; BMS Part Number: CFP11DCC-PRT.

DeVore:1992:ICT

[D^{+}92] R. DeVore et al. Image compression through wavelet transform coding. IEEE Transactions on Information Theory, 38(2):719– 746, March 1992. CODEN IETTAW. ISSN 0018-9448 (print), 1557-9654 (electronic).

Bresolin:2008:EAA

[dABdSFNA08] A. de Andrade Bresolin, D. da Silva Freitas, A. D. D. Neto, and P. J. Alsina. European and American audio-visual speech recognition, using SVM in Portuguese language. In Storer and Marcellin [SM08], page 511. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33 INTERNET. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=4483338.

Daciuk:2001:EAC

[Dac01] Jan Daciuk. Experiments with automata compression. Lecture Notes in Computer Science, 2088:105–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer-ny.com/link/
service/series/0558/bibs/2088/20880105.htm;   http:
//link.springer-ny.com/link/service/series/0558/papers/
2088/20880105.pdf.

Dasarathy:1995:IDC

[Das95] Belur V. Dasarathy, editor. Image Data Compression: Block Truncation Coding (BTC) Techniques. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-8186-6847-4 (paper). ???? pp. LCCN TA1637 .I423 1995.

Daubechies:1988:OBC

[Dau88] Ingrid Daubechies. Orthonormal bases of compactly supported wavelets. Communications on Pure and Applied Mathematics (New York), 41(7):909–996, October 1988. CODEN CPAMAT, CPMAMV. ISSN 0010-3640 (print), 1097-0312 (electronic). Reprinted in [?].

Davisson:1964:TDC

[Dav64] Lee David Davisson. Theory of Data Compression. Ph.D. thesis, University of California, Los Angeles, Los Angeles, CA,

USA, 1964. 46 pp. URL http://search.proquest.com/
docview/302143966.

Davisson:1966:CST

[Dav66] I. D. Davisson. Comments on ‘sequence time coding for data compression’. Proceedings of the IEEE, 54:2010, December 1966. CODEN IEEPAD. ISSN 0018-9219 (print), 1558-2256 (electronic).

Davison:1993:VTC

[Dav93] Andrew Davison. Vague text compression. ACM SIGACT News, 24(1):68–74, Winter 1993. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Davis:1995:SQW

[Dav95a] G. Davis. Self-quantized wavelet subtrees: a waveletbased theory for fractal image compression. In Storer and Cohn [SC95], pages 232–241. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ????

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=515513. IEEE catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Davis:1995:ASQ

[Dav95b] Geoffrey Davis. Adaptive self-quantization of wavelet subtrees: a wavelet-based theory of fractal image compression. Proceedings of the SPIE — The International Society for Optical Engineering, 2569:294–307, July 1995. CODEN PSISDG. ISSN 0277-786X (print), 1996-756X (electronic). URL http://www.

cs.dartmouth.edu/~gdavis/papers/spie_sandiego.ps.g.

Davis:1996:IIM

[Dav96] Geoffrey Davis. Implicit image models in image fractal compression. Proceedings of the SPIE — The International Society for Optical Engineering, 2569(??):88–97, ???? 1996. CODEN PSISDG. ISBN 0-8194-2213-4. ISSN ????

Davis:2012:LAP

[Dav12] Ernest Davis. Linear algebra and probability for computer science applications. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2012. ISBN 1-4665-0155-3 (hardcover). xviii + 413 pp. LCCN QA76.9.M35 D38 2012.

Das:2002:WBL

[DB02] B. Das and S. Banerjee. A wavelet based low complexity embedded block coding algorithm. In Storer and Cohn [SC02], page ?? ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2002. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=999995. IEEE Computer Society Order Number PR01477.

Das:2003:NRA

[DB03] Bipul Das and S. Banerjee. A novel RAM architecture for bitplane based coding. In Storer and Cohn [SC03], page ?? ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2003. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=1194040. IEEE

Computer Society Order number PR01896.

Dube:2007:BRP

[DB07] D. Dube and V. Beaudoin. Bit recycling with prefix codes. In Storer and Cohn [SC07], page 379. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=4148780. IEEE Computer Society Order Number P2791.

Dube:2008:AML

[DB08] D. Dube and V. Beaudoin. All-match LZ77 bit recycling. In Storer and Marcellin [SM08], page 513. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33 INTER-

NET. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=4483340.

Dube:2010:LDC

[DB10] D. Dube and V. Beaudoin. Lossless data compression via substring enumeration. In Storer and Marcellin [SM10b], pages 229–238. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=5453467.

Duval:2000:SDC

[DBTNT00] L. C. Duval, Van Bui-Tran, T. Q. Nguyen, and T. D. Tran. Seismic data compression using GenLOT: towards “optimality”? In Storer and Cohn [SC00], page ?? ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2000. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=838199. This millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Davis:1997:ICU

[DC97] G. M. Davis and S. Chawla. Image coding using optimized significance tree quantization. In Storer and Cohn [SC97], pages 387–396. ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1997. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=582064. IEEE

Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Dong:2003:RDB

[DC03] Yanting Dong and L. Carin. Rate-distortion bound for joint compression and classification. In Storer and Cohn [SC03], page ?? ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2003. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1194042. IEEE Computer Society Order number PR01896.

Di:2018:OEB

[DC18] Sheng Di and Franck Cappello. Optimization of error-bounded lossy compression for hard-to-compress HPC data. IEEE Transactions on Parallel and Distributed Systems, 29(1):129– 143, January 2018. CODEN ITDSEO. ISSN 1045-9219 (print),

1558-2183 (electronic). URL https://www.computer.org/
csdl/trans/td/2018/01/08031063-abs.html.

daCunha:2007:SMV

[dCDV07] A. L. da Cunha, Minh Do, and M. Vetterli. A stochastic model for video and its information rates. In Storer and Cohn [SC07], pages 3–12. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

4148739. IEEE Computer Society Order Number P2791.

DeAgostino:2017:CVE

[DCP17] Sergio De Agostino, Bruno Carpentieri, and Raffaele Pizzolante. Concurrent vs. exclusive reading in parallel decoding of LZ-compressed files. Algorithms (Basel), 10(1), March 2017. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/10/1/21.

DePrisco:1998:DEH

[DD98] Roberto De Prisco and Alfredo De Santis. On the data expansion of the Huffman compression algorithm. The Computer Journal, 41(3):137–144, ???? 1998. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/41/3/137.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/
Volume_41/Issue_03/Vol41_03.body.html#AbstractDePrisco.

Daubechies:2006:DCI

[DDGV06] Ingrid Daubechies, Ronald A. DeVore, C. Sinan Güntürk, and Vinay A. Vaishampayan. A/D conversion with imperfect quantizers. IEEE Transactions on Information Theory, 52(3):874– 885, March 2006. CODEN IETTAW. ISSN 0018-9448 (print), 1557-9654 (electronic). IEEE International Symposium on Circuits and Systems, Phoenix, AZ, May 26–29, 2002.

Denecker:1998:ILH

[DDL98] K. Denecker, P. De Neve, and I. Lemahieu. Improved lossless halftone image coding using a fast adaptive context template selection scheme. In Storer and Cohn [SC98], page ?? ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1998. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=672265. IEEE Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

DiLillo:2012:CBT

[DDT^{+}12] Antonella Di Lillo, Ajay Daptardar, Kevin Thomas, James A. Storer, and Giovanni Motta. Compression-based tools for navigation with an image database. Algorithms (Basel), 5(1):1–17, March 2012. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/5/1/1.

DeAgostino:1994:PCP

[De 94a] Sergio De Agostino. P-complete problems in data compression. Theoretical Computer Science, 127(1):181–186, May 09, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294

(electronic). URL http://www.elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/browse.cgi?year=1994&
volume=127&issue=1&aid=1568. See erratum [De 00b].

DeCanne:1994:ERC

[De 94b] B. De Canne. Error-recovery codes. Dr. Dobb’s Journal of Software Tools, 19(15):44, 46, 48, 50, 92–93, December 1994. CODEN DDJOEB. ISSN 1044-789X.

DeAgostino:2000:WOP

[De 00a] S. De Agostino. Work-optimal parallel decoders for LZ2 data compression. In Storer and Cohn [SC00], pages 393–399. ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2000. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=838179. This millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

DeAgostino:2000:EPC

[De 00b]  Sergio De Agostino.  Erratum to “P-complete Prob-

lems in Data Compression” [Theoret. Comput. Sci. 127 (1994) 181–186]. Theoretical Computer Science, 234(1– 2):325–326, March 6, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://

www.elsevier.nl/gej-ng/10/41/16/170/21/37/abstract.
html;   http://www.elsevier.nl/gej-ng/10/41/16/170/
21/37/article.pdf. See [De 94a].

DeAgostino:2003:AWO

[De 03] S. De Agostino. Almost work-optimal PRAM EREW decoders of LZ compressed text. In Storer and Cohn [SC03], page ?? ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2003. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1194041. IEEE Computer Society Order number PR01896.

DeAgostino:2005:BSD

[De 05] S. De Agostino. Bounded size dictionary compression: relaxing the LRU deletion heuristic. In Storer and Cohn [SC05], page ?? ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=1402213.

DeAgostino:2006:LIC

[De 06] S. De Agostino. Lossless image compression by block matching on a mesh of trees. In Storer and Cohn [SC06], page 443. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1607286.

DeAgostino:2011:LZD

[De 11] Sergio De Agostino. Lempel–Ziv data compression on parallel and distributed systems. Algorithms (Basel), 4(3):183–199, September 2011. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/4/3/183.

Dionysian:1992:VPR

[DE92] R. Dionysian and M. D. Ercegovac. Variable precision representation for efficient VQ codebook storage. In Storer and Cohn [SC92], pages 319–328. ISBN 0-8186-2717-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ????

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=227449. IEEE catalog number 91TH0436-6.

Degener:1994:DSC

[Deg94] Jutta Degener. Digital speech compression. Dr. Dobb’s Journal of Software Tools, 19(15):30, 32, 34, 88–89, December 1994. CODEN DDJOEB. ISSN 1044-789X.

Delfino:1993:PMI

[Del93] Erik Delfino. PC monitor — the Internet toolkit: File compression and archive utilities. Online, 17(6):90–??, November 1, 1993. CODEN ONLIDN. ISSN 0146-5422.

Delport:1995:AMC

[Del95] V. Delport. Alternative methods for codebook design in vector quantization. In Storer and Cohn [SC95], page 485. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ????  URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=515595. IEEE catalog

number 95TH8037. IEEE Computer Society Press order number PR07010.

Denisowski:2001:HDI

[Den01] P. Denisowski. How does it sound? [voice clarity analysis]. IEEE Spectrum, 38(2):60–64, February 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Deorowicz:2000:IBW

[Deo00] Sebastian Deorowicz. Improvements to Burrows–Wheeler compression algorithm. Software — Practice and Experience, 30(13):1465–1483, November 10, 2000. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (elec-

tronic). URL http://www3.interscience.wiley.com/cgi-
bin/abstract/73001202/START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?ID=73001202&PLACEBO=IE.
pdf.

Deorowicz:2002:SSA

[Deo02]  Sebastian Deorowicz.  Second step algorithms in the

Burrows–Wheeler compression algorithm. Software — Practice and Experience, 32(2):99–111, February 2002. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X

(electronic).  URL http://www3.interscience.wiley.
com/cgi-bin/abstract/88013462/START;   http://www3.
interscience.wiley.com/cgi-bin/fulltext?ID=88013462&
PLACEBO=IE.pdf.

Destasio:2008:EU

[Des08] B. Destasio. Emilcont ultracompression, 2008. URL http://

www.freewebs.com/emilcont/.

Deutsch:1996:RDC

[Deu96a] L. P. Deutsch. RFC 1951: DEFLATE compressed data format specification version 1.3, May 1996. URL ftp://ftp.

internic.net/rfc/rfc1951.txt;   http://www.ietf.org/

rfc/rfc1951;   https://www.math.utah.edu/pub/rfc/rfc1951.

txt. Status: INFORMATIONAL.

Deutsch:1996:RGF

[Deu96b] L. Peter Deutsch. RFC 1952: GZIP file format specification

version 4.3, May 1996. URL ftp://ftp.internic.net/rfc/
rfc1952.txt;   http://www.gzip.org/zlib/rfc-gzip.html;
 https://www.math.utah.edu/pub/rfc/rfc1952.txt. Sta-

tus: INFORMATIONAL.

Daehlen:1993:IPI

[DF93] Morten Dæhlen and Michael Floater. Iterative polynomial interpolation and data compression. Numerical Algorithms, 5 (1–4):165–177, November 1993. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Algorithms for approximation, III (Oxford, 1992).

Du:2008:LCP

[DF08] Qian Du and James E. Fowler. Low-complexity principal component analysis for hyperspectral image compression. The International Journal of High Performance Computing Applications, 22(4):438–448, November 2008. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL http://hpc.

sagepub.com/content/22/4/438.full.pdf+html.

Danieli:2010:MMI

[DFA^{+}10] M. Danieli, S. Forchhammer, J. D. Andersen, L. P. B. Christensen, and S. S. Christensen. Maximum mutual information vector quantization of log-likelihood ratios for memory efficient HARQ implementations. In Storer and Marcellin [SM10b], pages 30–39. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=
5453440.

Diffenderfer:2019:EAZ

[DFH^{+}19] James Diffenderfer, Alyson L. Fox, Jeffrey A. Hittinger, Geoffrey Sanders, and Peter G. Lindstrom. Error analysis of ZFP compression for floating-point data. SIAM Journal on Scientific Computing, 41(3):A1867–A1898, ???? 2019. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

Davisson:1976:DC

[DG76] Lee D. Davisson and Robert M. Gray, editors. Data compression, volume 14 of Benchmark papers in electrical engineering and computer science. Dowden, Hutchinson and Ross, Stroudsburg, PA, USA, 1976. ISBN 0-87933-089-9. xv + 407 pp. LCCN TK5102.5 .D335 1976.

Deutsch:1996:RZC

[DG96] L. P. Deutsch and J-L. Gailly. RFC 1950: ZLIB compressed data format specification version 3.3, May 1996. URL ftp://

ftp.internic.net/rfc/rfc1950.txt;   http://www.ietf.
org/rfc/rfc1950;   https://www.math.utah.edu/pub/rfc/

rfc1950.txt. Status: INFORMATIONAL.

Dutta:2017:EFC

[DG17] Tanima Dutta and Hari Prabhat Gupta. An efficient framework for compressed domain watermarking in P frames of high-efficiency video coding (HEVC)-encoded video. ACM Transactions on Multimedia Computing, Communications, and Applications, 13(1):12:1–12:??, January 2017. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Dechevsky:2009:WCD

[DGG09] Lubomir T. Dechevsky, Niklas Grip, and Joakim Gundersen. Wavelet compression, data fitting and approximation based on adaptive composition of Lorentz-type tresholding and Besovtype non-treshold shrinkage. R&D report 5/2009, Narvik University College, Narvik, Norway, 2009. 9 pp.

Dechevsky:2010:WCD

[DGG10] Lubomir T. Dechevsky, Niklas Grip, and Joakim Gundersen. Wavelet Compression, Data Fitting and Approximation Based on Adaptive Composition of Lorentz-Type Thresholding and Besov-Type Non-threshold Shrinkage, pages 738–746. ????, ????, 2010. LCCN ????

Diaz-Gutierrez:2005:GIH

[DGGP05] Pablo Diaz-Gutierrez, M. Gopi, and Renato Pajarola. Geometry II: Hierarchyless simplification, stripification and compression of triangulated two-manifolds. Computer Graphics Forum, 24(3):457–467, September 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Danciu:1997:FCC

[DH97] I. M. Danciu and J. C. Hart. Fractal color compression in the L^{∗} a^{∗} b^{∗} uniform color space. In Storer and Cohn [SC97], page ?? ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1997. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=582090. IEEE

Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Danciu:1998:FCC

[DH98] I. M. Danciu and J. C. Hart. Fractal color compression in the L^{∗} a^{∗} b^{∗} uniform color space. In Storer and Cohn [SC98], page ?? ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1998. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=672264. IEEE

Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Deever:2000:WYS

[DH00] A. Deever and S. S. Hemami. What’s your sign?: efficient sign coding for embedded wavelet image coding. In Storer and Cohn [SC00], pages 273–282. ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2000. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=838167. This

millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Dong:2018:CAP

[DH18] D. Dong and J. Herbert. Content-aware partial compression for textual big data analysis in Hadoop. IEEE Transactions on Big Data, 4(4):459–472, December 2018. ISSN 2332-7790.

Dawy:2004:ICT

[DHH04] Z. Dawy, J. Hagenauer, and A. Hoffmann. Implementing the context tree weighting method for content recognition. In Storer and Cohn [SC04], page ?? ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1281512. IEEE Computer Society Order Number: P2082.

Demko:1985:CFO

[DHN85] Stephen Demko, Laurie Hodges, and Bruce Naylor. Construction of fractal objects with iterated function systems. Computers and Graphics, 19(3):271–278, July 1985. CODEN CGRADI, CPGPBZ. ISSN 0097-8930 (print), 1558-4569 (electronic).

Demiroglu:2001:JSC

[DHS01a] C. Demiroglu, M. W. Hoffman, and K. Sayood. Joint source channel coding using arithmetic codes and trellis coded modulation. In Storer and Cohn [SC01c], pages 302–311. ISBN 0-7695-1031-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2001. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=917161. IEEE

CSP number 01PR1031.

Dorward:2001:LDP

[DHS^{+}01b] S. Dorward, D. Huang, S. A. Savari, G. Schuller, and B. Yu. Low delay perceptually lossless coding of audio signals. In Storer and Cohn [SC01c], pages 312–320. ISBN 0-7695-1031-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 2001. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=917162. IEEE CSP num-

ber 01PR1031.

Drmota:2002:PAR

[DHS02] M. Drmota, Hsien-Kuei Hwang, and W. Szpankowski. Precise average redundancy of an idealized arithmetic coding. In Storer and Cohn [SC02], pages 222–231. ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2002. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=999960. IEEE Computer Society Order Number PR01477.

Deng:2017:EOD

[DHW^{+}17] Ze Deng, Wei Han, Lizhe Wang, Rajiv Ranjan, Albert Y. Zomaya, and Wei Jie. An efficient online direction-preserving compression approach for trajectory streaming data. Future Generation Computer Systems, 68(??):150–162, March 2017. CODEN FGSEVI. ISSN 0167-739X (print), 1872-

7115 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167739X16303508.

Davidson:2004:FDC

[DI04] Michael Davidson and Lucian Ilie. Fast data compression with antidictionaries. Fundamenta Informaticae, 64(1–4):119–134, July 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Davidson:2005:FDC

[DI05] Michael Davidson and Lucian Ilie. Fast data compression with antidictionaries. Fundamenta Informaticae, 64(1–4):119–134, July 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

DiazDiaz:2016:XFI

[Dia16] Antonio Diaz Diaz. Xz format inadequate for long-term archiving. Web document., June 11, 2016. URL https:/

/www.nongnu.org/lzip/xz_inadequate.html. Updated 20-

Jul-2018.

Dickman:1992:PCS

[Dic92] Chris Dickman. The President and CEO of Storm Technology discusses file compression issues. Electronic Composition and Imaging, 6(??):52–??, March 1992. ISSN 0838-9535.

Dion:1993:CPC

[Dio93] Pierre J. Dion. Cross-platform compression. Dr. Dobb’s Journal of Software Tools, 18(12):32, 34, 36–37, December 1993. CODEN DDJOEB. ISSN 1044-789X.

Disque:1988:SCT

[Dis88] Tom Disque. Simple compression technique works with text and graphics. C Users Journal, 6(8):23–??, August 1988. ISSN 0898-9788.

Du:2009:CPL

[DJCM09] Zhiyan Du, P. Jaromersky, Yi-Jen Chiang, and N. Memon. Out-of-core progressive lossless compression and selective decompression of large triangle meshes. In Storer and Marcellin [SM09], pages 420–429. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ????

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=4976486. IEEE Computer Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

Dunton:2020:PEM

[DJID20] Alec M. Dunton, Lluís Jofre, Gianluca Iaccarino, and Alireza Doostan. Pass-efficient methods for compression of highdimensional turbulent flow data. Journal of Computational Physics, 423(??):Article 109704, December 15, 2020. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999120304782.

DeVore:1991:DCU

[DJL91] R. A. DeVore, B. Jawerth, and B. J. Lucier. Data compression using wavelets: error, smoothness and quantization. In Storer and Reif [SR91b], pages 186–195. ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1990. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=213386. IEEE Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Do:2014:FRL

[DJSS14] Huy Hoang Do, Jesper Jansson, Kunihiko Sadakane, and Wing-Kin Sung. Fast relative Lempel–Ziv self-index for similar sequences. Theoretical Computer Science, 532(??):14–30, May 1, 2014. CODEN TCSCDI. ISSN 0304-3975 (print),

1879-2294 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304397513005409.

Du:1995:CCE

[DK95] Junchen Du and S. P. Kim. Classified conditional entropy coding of LSP parameters. In Storer and Cohn [SC95], page 435. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=515545. IEEE catalog

number 95TH8037. IEEE Computer Society Press order number PR07010.

Drinic:2002:PPC

[DK02] M. Drinic and D. Kirovski. PPMexe: PPM for compressing software. In Storer and Cohn [SC02], pages 192–201.

ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2002. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=999957. IEEE

Computer Society Order Number PR01477.

Dado:2016:TCG

[DKB^{+}16] Bas Dado, Timothy R. Kol, Pablo Bauszat, Jean-Marc Thiery, and Elmar Eisemann. Texturing & compression: Geometry and attribute compression for voxel scenes. Computer Graphics Forum, 35(2):397–407, May 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Docef:1995:MFS

[DKCS95] A. Docef, F. Kossentini, W. C. Chung, and M. J. T. Smith. Multiplication-free subband coding of color images. In Storer and Cohn [SC95], pages 352–361. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=515525. IEEE catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Dragotti:2001:QOF

[DKG01] P. L. Dragotti, J. Kovacevic, and V. K. Goyal. Quantized oversampled filter banks with erasures. In Storer and Cohn [SC01c], pages 173–182. ISBN 0-7695-1031-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2001. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=917148. IEEE CSP number 01PR1031.

Drinic:2003:PMC

[DKP03] M. Drinic, D. Kirovski, and M. Potkonjak. PPM model cleaning. In Storer and Cohn [SC03], pages 163–172. ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 2003. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1194007. IEEE Computer

Society Order number PR01896.

Dua:2021:PLH

[DKS21] Yaman Dua, Vinod Kumar, and Ravi Shankar Singh. Parallel lossless HSI compression based on RLS filter. Journal of Parallel and Distributed Computing, 150(??):60–68, April 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731520304202.

Drinic:2007:PPC

[DKV07] Milenko Drinić, Darko Kirovski, and Hoi Vo. PPMexe: Program compression. ACM Transactions on Programming Languages and Systems, 29(1):3:1–3:31, January 2007. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Duan:2001:CLD

[DL01] Jiangang Duan and Jin Li. Compression of the layered depth image. In Storer and Cohn [SC01c], pages 331–340. ISBN 0-7695-1031-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2001. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=917164. IEEE

CSP number 01PR1031.

Nogueiras:1995:HPB

[dlCNLA^{+}95] A. de la Cruz Nogueiras, M. Gamez Lau, A. Cerdeira Altuzarra, M. Estrada del Cueto, and P. Goga. A high performance block compression algorithm for small systems-software and hardware implementations. In Storer and Cohn [SC95], page 422. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=515532. IEEE

catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Dorrigiv:2008:LUA

[DLOM08] R. Dorrigiv, A. Lopez-Ortiz, and J. I. Munro. List update algorithms for data compression. In Storer and Marcellin [SM08], page 512. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008;

QA76.9.D33 INTERNET. URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=4483339.

Dai:2005:ABB

[DLT05] Wei Dai, Lijie Liu, and T. D. Tran. Adaptive block-based image coding with pre-/post-filtering. In Storer and Cohn [SC05], pages 73–82. ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=
1402168.

Daly:2016:DRA

[DLT16] James Daly, Alex X. Liu, and Eric Torng. A difference resolution approach to compressing access control lists. IEEE/ ACM Transactions on Networking, 24(1):610–623, February 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Davis:1998:PRL

[DLY^{+}98] G. Davis, L. Lau, R. Young, F. Duncalfe, and L. Brebber. Parallel run length encoding compression: Reducing I/O in dynamic environmental simulations. The International Journal of High Performance Computing Applications, 12(4):396– 410, Winter 1998. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic).

Deligiannis:2009:MCN

[DMC^{+}09] N. Deligiannis, A. Munteanu, T. Clerckx, P. Schelkens, and J. Cornelis. Modeling the correlation noise in spatial domain distributed video coding. In Storer and Marcellin [SM09], page 443. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=4976497. IEEE

Computer Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

deMedeiros:2010:LAC

[dMGdC10] R. J. V. de Medeiros, E. C. Gurjao, and J. M. de Carvalho. Lossy audio compression via compressed sensing. In Storer and Marcellin [SM10b], page 545. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=5453498.

DiLillo:2008:MRI

[DMS08] A. Di Lillo, G. Motta, and J. A. Storer. Multiresolution rotation-invariant texture classification using feature extraction in the frequency domain and vector quantization. In Storer and Marcellin [SM08], pages 452–461. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33 INTERNET.

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=
&arnumber=4483323.

DiLillo:2010:SRU

[DMS10] A. Di Lillo, G. Motta, and J. A. Storer. Shape recognition using vector quantization. In Storer and Marcellin [SM10b], pages 484–493. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=
5453468.

Derpich:2008:QGR

[DØG08] M. S. Derpich, J. Østergaard, and G. C. Goodwin. The quadratic Gaussian rate-distortion function for source uncorrelated distortions. In Storer and Marcellin [SM08], pages 73–82. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33

INTERNET. URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4483285.

Dahl:2009:CIS

[DØJJ09] J. Dahl, J. Østergaard, T. L. Jensen, and S. H. Jensen. l_{1} compression of image sequences using the structural similarity index measure. In Storer and Marcellin [SM09], pages 133–142. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=4976457. IEEE Computer

Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

Desoky:1988:CID

[DOK88] Ahmed Desoky, Carol O’Connor, and Thomas Klein. Compression of image data using arithmetic coding. In Wegman et al. [WGM88], pages 812–814. URL http://www.dtic.mil/

dtic/tr/fulltext/u2/a208838.pdf.

Dolby:2006:DLH

[Dol06a] Dolby. Dolby Laboratories home page, 2006. URL http://

www.dolby.com/.

Dolby:2006:TDH

[Dol06b] Thomas Dolby. Thomas Dolby home page, 2006. URL http:

//www.thomasdolby.com/.

Dosik:1972:OBC

[Dos72] Paul H. Dosik. An Optimal Buffer Controlled Data Compression System. Ph.D. thesis, Polytechnic University, New York,

NY, USA, 1972. 190 pp. URL http://search.proquest.
com/docview/302676863.

Douglis:1993:CCU

[Dou93] Fred Douglis. The compression cache: Using on-line compression to extend physical memory. In USENIX [USE93], pages 519–529. ISBN 1-880446-48-0. LCCN QA 76.76 O63

U84 1993. URL http://www.usenix.org/publications/
library/proceedings/sd93/.

Dozier:1991:OED

[Doz91] J. Dozier. Overview of the EOS data and information system. In Storer and Reif [SR91b], page ?? ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1990. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=213295. IEEE Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

dellaPorta:1558:MN

[dP58] Giambattista della Porta. Magia Naturalis. ????, Naples[, Italy], 1558. ?? pp. First edition, four volumes 1558, second edition, 20 volumes 1589. Translated by Thomas Young and Samuel Speed, Natural Magick by John Baptista Porta, a Neopolitane, London 1658.

Dance:1976:ALD

[DP76] D. L. Dance and U. W. Pooch. An adaptive on line data compression system. The Computer Journal, 19 (3):216–224, August 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/19/3/216.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_03/tiff/216.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_19/Issue_03/
tiff/217.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_03/tiff/218.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_19/Issue_03/
tiff/219.tif;   http://www3.oup.co.uk/computer_journal/

hdb/Volume_19/Issue_03/tiff/220.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_19/Issue_03/
tiff/221.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_03/tiff/222.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_19/Issue_03/
tiff/223.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_03/tiff/224.tif.

Santis:1991:OAC

[DP91] A. De Santis and G. Persiano. An optimal algorithm for the construction of optimal prefix codes with given fringe. In Storer and Reif [SR91b], pages 297–306. ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1990. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=213351. IEEE Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Diamantini:2001:EDC

[DP01] Claudia Diamantini and Maurizio Panti. An efficient data compression approach to the classification task. Lecture Notes in Computer Science, 2035:444–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/series/0558/bibs/
2035/20350444.htm;   http://link.springer-ny.com/link/
service/series/0558/papers/2035/20350444.pdf.

Dutra:2008:CHI

[DPdS08] A. J. S. Dutra, W. A. Pearlman, and E. A. B. da Silva. Compression of hyperspectral images with LVQ-SPECK. In Storer and Marcellin [SM08], pages 93–102. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33 INTERNET.

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=
&arnumber=4483287.

Dahmen:1993:WAM

[DPS93] W. Dahmen, S. Prössdorf, and R. Schneider. Wavelet approximation methods for pseudodifferential equations. II. Matrix compression and fast solution. Advances in Computational Mathematics, 1(3):259–335, October 1993. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF02072014.

Dvorsky:1999:WBC

[DPS99] J. Dvorsky, J. Pokorny, and V. Snasel. Word-based compression methods for large text documents. In Storer and Cohn [SC99], page ?? ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1999. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=785680. IEEE

Computer Society Order Number PR00096. IEEE Order Plan Catalog Number PR00096.

Dhem:1997:LCA

[DQ97] J. F. Dhem and J. J. Quisquater. Lossless compression algorithms for smart cards: A progress report. Future Generation Computer Systems, 13(1):27–38, June 1997. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167739X97891091.

Dewitte:1983:OBC

[DR83] J. Dewitte and J. Ronson. Original block coding scheme for low bit rate image transmission. In H. W. Schussler, editor, Signal Processing II: Theories and Applications—Proceedings of EUSIPCO 83, pages 143–146. Elsevier Science, Inc., Amsterdam, The Netherlands, 1983.

Duni:2006:HRT

[DR06] E. R. Duni and B. D. Rao. High-rate training of Gaussian mixture vector quantizers. In Storer and Cohn [SC06], page 445. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=1607288.

Das:1994:VDV

[DRG94] A. Das, A. V. Rao, and A. Gersho. Variable-dimension vector quantization of speech spectra for low-rate vocoders. In Storer and Cohn [SC94], pages 420–429. ISBN 0-8186-5636-0, 0-8186-5637-9 (case). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1994. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=305949. IEEE

Catalog Number 93TH0626-2. IEEE Computer Society Press Order Number 5637-02.

Demay:2010:RDB

[DRR10] G. Demay, V. Rathi, and L. K. Rasmussen. Rate distortion bounds for binary erasure source using sparse graph codes. In Storer and Marcellin [SM10b], pages 49–58. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=5453435.

Davis:1981:DCL

[DRT81] R. H. Davis, C. Rinaldi, and C. J. Trebilcock. Data compression in limited capacity microcomputer systems. Information Processing Letters, 13(4–5):138–141, End ??, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

deRooij:2012:ARD

[dRV12] Steven de Rooij and Paul M. B. Vitanyi. Approximating ratedistortion graphs of individual data: Experiments in lossy compression and denoising. IEEE Transactions on Computers, 61 (3):395–407, March 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Agostino:1992:PAO

[DS92] S. De Agostino and J. A. Storer. Parallel algorithms for optimal compression using dictionaries with the prefix property. In Storer and Cohn [SC92], pages 52–61. ISBN 0-8186-2717-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=227476. IEEE catalog number 91TH0436-6.

DeAgostino:1996:LVL

[DS96] Sergio De Agostino and James A. Storer. On-line versus offline computation in dynamic text compression. Information Processing Letters, 59(3):169–174, August 12, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Demir:1998:JSC

[DS98] N. Demir and K. Sayood. Joint source/channel coding for variable length codes. In Storer and Cohn [SC98], pages 139–148. ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1998. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=672140. IEEE

Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Daptardar:2004:AKM

[DS04] A. Daptardar and D. Shapira. Adapting the Knuth–Morris– Pratt algorithm for pattern matching in Huffman encoded texts. In Storer and Cohn [SC04], page ?? ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1281511. IEEE Computer Society Order Number: P2082.

Daptardar:2006:RCC

[DS06] A. H. Daptardar and J. A. Storer. Reduced complexity content-based image retrieval using vector quantization. In Storer and Cohn [SC06], pages 342–351. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=1607269.

Daptardar:2008:VBI

[DS08] A. H. Daptardar and J. A. Storer. VQ based image retrieval using color and position features. In Storer and Marcellin [SM08], pages 432–441. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008;

QA76.9.D33 INTERNET. URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=4483321.

Duan:2012:RDA

[DSG12] Yizhou Duan, Jun Sun, and Zongming Guo. Rate-distortion analysis and modeling of dead-zone plus uniform threshold scalar quantization for generalized Gaussian random variables. In Storer and Marcellin [SM12], page 395. ISBN 1-4673-0715-7 (paperback). ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=6189276. IEEE Computer Society

order number P4656.

DiLillo:2007:TCU

[DSM07] A. Di Lillo, J. A. Storer, and G. Motta. Texture classification using VQ with feature extraction based on transforms motivated by the human visual system. In Storer and Cohn [SC07], page 392. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=4148793. IEEE

Computer Society Order Number P2791.

Doshi:2007:DFC

[DSMJ07] V. Doshi, D. Shah, M. Medard, and S. Jaggi. Distributed functional compression through graph coloring. In Storer and Cohn [SC07], pages 93–102. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=4148748. IEEE Computer Society Order Number P2791.

Datcu:1995:CFP

[DSSR95a] M. Datcu, G. Schwarz, K. Schmidt, and C. Reck. Correction of fixed pattern background and restoration of JPEG compressed CCD images. In Storer and Cohn [SC95], page 424. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=515534. IEEE catalog

number 95TH8037. IEEE Computer Society Press order number PR07010.

Datcu:1995:HAJ

[DSSR95b] M. Datcu, G. Schwarz, K. Schmidt, and C. Reck. Histogram analysis of JPEG compressed images as an aid in image deblocking. In Storer and Cohn [SC95], page 425. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=515535. IEEE catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Diggavi:2000:DAM

[DSV00a] S. N. Diggavi, N. J. A. Sloane, and V. A. Vaishampayan. Design of asymmetric multiple description lattice vector quantizers. In Storer and Cohn [SC00], pages 490–499. ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2000. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=838189. This millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Dragotti:2000:AOF

[DSV00b] P. L. Dragotti, S. D. Servetto, and M. Vetterli. Analysis of optimal filter banks for multiple description coding. In Storer and Cohn [SC00], pages 323–332. ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2000. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=838172. This millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Deng:2012:VIA

[DTZZ12] Aidong Deng, Jianeng Tang, Li Zhao, and Cairong Zou. The variable-interval arithmetic coding using asymptotic deterministic randomness for data compression and encryption. Journal of Statistical Computation and Simulation, 82(10):1545–1555, 2012. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Du:2000:DCP

[Du00] Ying Du. Data compression — the process of encoding data to reduce their storage requirements. M.Eng., Donghua University, Donghua, Peoples Republic of China, 2000. URL http:/

/search.proquest.com/docview/1027130485.

Dube:2011:USS

[Dub11] D. Dubé. On the use of stronger synchronization to boost compression by substring enumeration. In Storer and Marcellin [SM11b], page 454. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ????

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=5749511. IEEE Computer Society Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Damme:2019:CES

[DUH^{+}19] Patrick Damme, Annett Ungethüm, Juliana Hildebrandt, Dirk Habich, and Wolfgang Lehner. From a comprehensive experimental survey to a cost-based selection strategy for lightweight integer compression algorithms. ACM Transactions on Database Systems, 44(3):9:1–9:??, June 2019. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

URL https://dl.acm.org/ft_gateway.cfm?id=3323991.

Dumitrescu:2006:FAD

[Dum06] S. Dumitrescu. Faster algorithm for designing optimal prefixfree codes with unequal letter costs. In Storer and Cohn [SC06], page 444. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=1607287.

Dumitrescu:2008:SEO

[Dum08] S. Dumitrescu. Speed-up of encoder optimization step in multiple description scalar quantizer design. In Storer and Marcellin [SM08], pages 382–391. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33 INTERNET. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=4483316.

Dunham:1980:AAS

[Dun80] James G. Dunham. Abstract alphabet sliding-block entropy compression coding with a fidelity criterion. Annals of Probability, 8(6):1085–1092, December 1980. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http:/

/projecteuclid.org/euclid.aop/1176994570.

Dunlavey:1981:LES

[Dun81] Michael R. Dunlavey. Letter to the Editor: On spelling correction and beyond. Communications of the ACM, 24(9):608, September 1981. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Pet80, Mil81, Nix81, Pet81].

Durbin:1960:FTS

[Dur60] J. Durbin. The fitting of time-series models. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 28:233–344, 1960. ISSN 0373-1138.

Duvanenko:1988:ICC

[Duv88] Victor J. Duvanenko. Image compression via compilation. Dr. Dobb’s Journal of Software Tools, 13(11):42–47, 82–90, November 1988. CODEN DDJOEB. ISSN 0888-3076.

DeMarchi:2001:AAN

[DV01] Stefano De Marchi and Marco Vianello. Approximating the approximant: a numerical code for polynomial compression of discrete integral operators. Numerical Algorithms, 28(1–4):101–116, December 2001. CODEN NUALEG.

ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
37/7/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/37/7/fulltext.pdf;   http:
//ipsapp007.lwwonline.com/content/getfile/5058/37/
7/abstract.htm;   http://ipsapp007.lwwonline.com/content/
getfile/5058/37/7/fulltext.pdf.

DVB:2006:DVB

[DVB06] DVB. Digital Video Broadcasting Project (DVB) home page,

2006. URL http://www.dvb.org/.

Donoho:1998:DCH

[DVDD98] David L. Donoho, Martin Vetterli, R. A. DeVore, and Ingrid Daubechies. Data compression and harmonic analysis. IEEE Transactions on Information Theory, 44(6):2435–2476, October 1998. CODEN IETTAW. ISSN 0018-9448 (print), 1557-9654 (electronic). Information theory: 1948–1998.

Denecker:1997:ECS

[DvOL97] K. Denecker, J. van Overloop, and I. Lemahieu. An experimental comparison of several lossless image coders for medical images. In Storer and Cohn [SC97], page ?? ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1997. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=582091. IEEE Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Deligiannis:2014:PRW

[DVS^{+}14] Nikos Deligiannis, Frederik Verbist, Jürgen Slowack, Rik van de Walle, Peter Schelkens, and Adrian Munteanu. Progressively refined Wyner–Ziv video coding for visual sensors. ACM Transactions on Sensor Networks, 10(2):21:1–21:??, January 2014. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Dumitrescu:2003:OVR

[DW03] S. Dumitrescu and Xiaolin Wu. Optimal variable rate multiplexing of scalable code streams. In Storer and Cohn [SC03], page ?? ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2003. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1194043. IEEE Computer Society Order number PR01896.

Dumitrescu:2005:GOG

[DW05] S. Dumitrescu and Xiaolin Wu. On global optimality of gradient descent algorithms for fixed-rate scalar multiple description quantizer design. In Storer and Cohn [SC05], pages 388–397. ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1402200.

Dumitrescu:2004:FAO

[DWB04] S. Dumitrescu, X. Wu, and G. Bahl. Fast algorithms for optimal two-description scalar quantizer design. In Storer and Cohn [SC04], pages 42–51. ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1281449. IEEE Computer Society Order Number: P2082.

Dumitrescu:2002:GOU

[DWW02] S. Dumitrescu, Xiaolin Wu, and Zhe Wang. Globally optimal uneven error-protected packetization of scalable code streams. In Storer and Cohn [SC02], pages 73–82. ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2002. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=999945. IEEE Computer Society Order Number PR01477.

Dai:2011:BSP

[DX11] Wenrui Dai and Hongkai Xiong. Bitwise structured prediction model for lossless image coding. In Storer and Marcellin [SM11b], page 453. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

5749510. IEEE Computer Society Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Dai:2008:NSC

[DXS08] Wenrui Dai, Hongkai Xiong, and Li Song. On non-sequential context modeling with application to executable data compression. In Storer and Marcellin [SM08], pages 172–181.

ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33 INTER-

NET. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=4483295.

Desoky:1991:PAV

[DY91] A. Desoky and Y. You. Performance analysis of a vector quantization algorithm of image data. In Storer and Reif [SR91b], page ?? ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1990. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=213318. IEEE

Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Deng:2012:OAC

[DY12] Xi Deng and Yuanyuan Yang. Online adaptive compression in delay sensitive wireless sensor networks. IEEE Transactions on Computers, 61(10):1429–1442, October 2012. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Dobson:1996:LCW

[DYW^{+}96] K. Dobson, J. Yang, N. Whitney, K. Smart, and P. Rigstad. A low complexity wavelet based audio compression method. In Storer and Cohn [SC96], page ?? ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1996. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=488365. IEEE Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Dai:2003:BCC

[DZ03] Vito Dai and A. Zakhor. Binary combinatorial coding. In Storer and Cohn [SC03], page ?? ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 2003. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1194039. IEEE Computer

Society Order number PR01896.

Dumitrescu:2008:IMD

[DZ08] S. Dumitrescu and Ting Zheng. Improved multiple description framework based on successively refinable quantization and uneven erasure protection. In Storer and Marcellin [SM08], page 514. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33

INTERNET. URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4483341.

Dai:2008:TDC

[DZZ08] Lu Dai, Li Zhang, and Xiaolin Zhao. A three dimensional combinative lifting algorithm for wavelet transform using 9/7 filter. In Storer and Marcellin [SM08], page 510. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33 INTER-

NET. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=4483337.

Estes:1995:EEF

[EA95] R. R. Estes, Jr. and V. R. Algazi. Efficient error free chain coding of binary documents. In Storer and Cohn [SC95], pages 122–131. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=515502. IEEE

catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Edgoose:1998:MML

[EA98] T. Edgoose and L. Allison. Minimum message length hidden Markov modelling. In Storer and Cohn [SC98], pages 169–178. ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1998. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=672145. IEEE

Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Eriksson:2004:NMT

[EANG04] T. Eriksson, J. B. Anderson, M. Novak, and N. Goertz. New methods for trellis source coding at rates above and below one. In Storer and Cohn [SC04], page ?? ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1281513. IEEE Computer Society Order Number: P2082.

Elgohary:2016:CLA

[EBH^{+}16] Ahmed Elgohary, Matthias Boehm, Peter J. Haas, Frederick R. Reiss, and Berthold Reinwald. Compressed linear algebra for large-scale machine learning. Proceedings of the VLDB Endowment, 9(12):960–971, August 2016. CODEN ???? ISSN 2150-8097.

Elgohary:2017:SML

[EBH^{+}17] Ahmed Elgohary, Matthias Boehm, Peter J. Haas, Frederick R. Reiss, and Berthold Reinwald. Scaling machine learning via compressed linear algebra. ACM SIGMOD Record, 46(1):42– 49, March 2017. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Elgohary:2018:CLA

[EBH^{+}18] Ahmed Elgohary, Matthias Boehm, Peter J. Haas, Frederick R. Reiss, and Berthold Reinwald. Compressed linear algebra for large-scale machine learning. VLDB Journal: Very Large Data Bases, 27(5):719–744, October 2018. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Elgohary:2019:CLA

[EBH^{+}19] Ahmed Elgohary, Matthias Boehm, Peter J. Haas, Frederick R. Reiss, and Berthold Reinwald. Compressed linear algebra for declarative large-scale machine learning. Communications of the ACM, 62(5):83–91, May 2019. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:

//cacm.acm.org/magazines/2019/5/236413/fulltext.

Esmaili:2009:LCS

[EC09] G. Esmaili and P. Cosman. Low complexity spatio-temporal key frame encoding for Wyner–Ziv video coding. In Storer and Marcellin [SM09], pages 382–390. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=4976482. IEEE Computer Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

Ebrahim:2016:MIB

[EC16] Mansoor Ebrahim and Wai Chong Chia. Multiview image block compressive sensing with joint multiphase decoding for visual sensor network. ACM Transactions on Multimedia Computing, Communications, and Applications, 12(2):30:1–30:??, March 2016. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Erdozain:2012:ODE

[ECBL12] A. Erdozain, P. M. Crespo, and B. Beferull-Lozano. Optimum distortion exponent in parallel fading channels by using analog joint source-channel coding schemes. In Storer and Marcellin [SM12], pages 277–286. ISBN 1-4673-0715-7 (paperback). ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=6189259. IEEE Computer Society

order number P4656.

Effros:1994:VDW

[ECG94] M. Effros, P. A. Chou, and R. M. Gray. Variable dimension weighted universal vector quantization and noiseless coding. In Storer and Cohn [SC94], pages 2–11. ISBN 0-8186-5636-0, 0-8186-5637-9 (case). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1994.

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=305907. IEEE Catalog Number 93TH0626-2. IEEE Computer Society Press Order Number 5637-02.

ECMA-151

[ECM91a] ECMA. ECMA-151: Data Compression for Information Interchange — Adaptive Coding with Embedded Dictionary — DCLZ Algorithm. ECMA (European Association for Standardizing Information and Communication Systems), Geneva,

Switzerland, June 1991. URL http://www.ecma.ch/ecma1/
STAND/ECMA-151.HTM.

ECMA-159

[ECM91b] ECMA. ECMA-159: Data Compression for Information Interchange — Binary Arithmetic Coding Algorithm. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzerland, December 1991.

URL http://www.ecma.ch/ecma1/STAND/ECMA-159.HTM.

ECMA-222

[ECM95] ECMA. ECMA-222: Adaptive Lossless Data Compression Algorithm. ECMA (European Association for Standardizing Information and Communication Systems), Geneva, Switzer-

land, June 1995. URL http://www.ecma.ch/ecma1/STAND/
ECMA-222.HTM.

Eck:1998:NCS

[ECM98] P. Eck, X. Changsong, and R. Matzner. A new compression scheme for syntactically structured messages (programs) and its application to Java and the Internet. In Storer and Cohn [SC98], page ?? ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1998. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=672266. IEEE

Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

ECMA-321

[ECM01] ECMA. ECMA-321: Streaming Lossless Data Compression Algorithm — (SLDC). ECMA (European Association for Standardizing Information and Communication Systems), Geneva,

Switzerland, June 2001. URL http://www.ecma.ch/ecma1/
STAND/ecma-321.htm.

Eberhard:2017:ACF

[EEH17] Sebastian Eberhard, Gabriel Ebner, and Stefan Hetzl. Algorithmic compression of finite tree languages by rigid acyclic grammars. ACM Transactions on Computational Logic, 18(4): 26:1–26:??, December 2017. CODEN ???? ISSN 1529-3785 (print), 1557-945X (electronic).

Ekbatani:2007:TSF

[EEJ07] S. Ekbatani, F. Etemadi, and H. Jafarkhani. Transmission over slowly fading channels using unreliable quantized feedback. In Storer and Cohn [SC07], pages 353–362. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4148774. IEEE Computer Society

Order Number P2791.

Eichinger:2015:TSC

[EEKB15] Frank Eichinger, Pavel Efros, Stamatis Karnouskos, and Klemens Böhm. A time-series compression technique and its application to the smart grid. VLDB Journal: Very Large Data Bases, 24(2):193–218, April 2015. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Effros:1997:FWU

[Eff97] M. Effros. Fast weighted universal transform coding: toward optimal, low complexity bases for image compression. In Storer and Cohn [SC97], pages 211–220. ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1997. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=582021. IEEE Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Effros:1998:PMR

[Eff98] M. Effros. Practical multi-resolution source coding: TSVQ revisited. In Storer and Cohn [SC98], pages 53–62. ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1998. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=672131. IEEE Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Effros:1999:ULS

[Eff99] M. Effros. Universal lossless source coding with the Burrows– Wheeler transform. In Storer and Cohn [SC99], pages 178–187. ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1999. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=755667. IEEE Computer Society Or-

der Number PR00096. IEEE Order Plan Catalog Number PR00096.

Effros:2000:PPB

[Eff00a] M. Effros. PPM performance with BWT complexity: a new method for lossless data compression. In Storer and Cohn [SC00], pages 203–212. ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2000. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=838160. This

millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Engelson:2000:LCH

[EFF00b] V. Engelson, D. Fritzson, and P. Fritzson. Lossless compression of high-volume numerical data from simulations. In Storer and Cohn [SC00], page ?? ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2000. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=838221. This

millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Ebrahimi:2016:JPT

[EFPS16] Touradj Ebrahimi, Siegfried Foessel, Fernando Pereira, and Peter Schelkens. JPEG Pleno: Toward an efficient representation of visual reality. IEEE MultiMedia, 23(4):14–20, October/December 2016. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL https://www.computer.

org/csdl/mags/mu/2016/04/mmu2016040014-abs.html.

Effros:2003:SKL

[EFZ03] M. Effros, Hanying Feng, and K. Zeger. Suboptimality of the Karhunen–Loeve transform for transform coding. In Storer and Cohn [SC03], pages 293–302. ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 2003. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1194020. IEEE Computer

Society Order number PR01896.

Erol:1998:HVC

[EGCK98] B. Erol, M. Gallant, G. Cote, and F. Kossentini. The H.263+ video coding standard: complexity and performance. In Storer and Cohn [SC98], pages 259–268. ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1998. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=672154. IEEE Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Esnaola:2008:DCC

[EGF08] I. Esnaola and J. Garcia-Frias. Distributed compression of correlated signals using random projections. In Storer and Marcellin [SM08], pages 322–331. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33 INTERNET. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=4483310.

Ehrlacher:2021:AHS

[EGLS21] Virginie Ehrlacher, Laura Grigori, Damiano Lombardi, and Hao Song. Adaptive hierarchical subtensor partitioning for tensor compression. SIAM Journal on Scientific Computing, 43(1):A139–A163, ???? 2021. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

Eriksson:2006:TBV

[EGNA06] T. Eriksson, N. Goertz, M. Novak, and J. B. Anderson. Trellis based variable rate residual image coding over noisy channels. In Storer and Cohn [SC06], pages 252–261. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1607260.

Elliott:1991:CTD

[EGS91] J. A. Elliott, P. M. Grant, and G. G. Sexton. Concurrent techniques for developing motion video compression algorithms. In Storer and Reif [SR91b], page ?? ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1990. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=213335. IEEE Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Eriksson:2005:BST

[EHAN05] T. Eriksson, S. Hellerbrand, J. B. Anderson, and M. Novak. On the block size of trellis quantizers. In Storer and Cohn [SC05], page ?? ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=1402214.

Epstein:1992:MKW

[EHSC92] B. R. Epstein, R. Hingorani, J. M. Shapiro, and M. Czigler. Multispectral KLT-wavelet data compression for Landsat thematic mapper images. In Storer and Cohn [SC92], pages 200– 208. ISBN 0-8186-2717-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=227461. IEEE catalog

number 91TH0436-6.

Etemadi:2007:ORP

[EJ07] F. Etemadi and H. Jafarkhani. Optimal rate and power allocation for layered transmission with superposition coding. In Storer and Cohn [SC07], page 380. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=4148781. IEEE Computer Society Order Number P2791.

Ewetz:2016:CRC

[EK16] Rickard Ewetz and Cheng-Kok Koh. Construction of reconfigurable clock trees for MCMM designs using mode separation and scenario compression. ACM Transactions on Design Automation of Electronic Systems (TODAES), 21(4):57:1–57:??, September 2016. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Ekstrand:1996:LCG

[Eks96] Nicklas Ekstrand. Lossless compression of grayscale images via context tree weighting. In Storer and Cohn [SC96], pages 132–139. ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1996. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=488318. IEEE

Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Etezadi:2012:PRT

[EKT12] F. Etezadi, A. Khisti, and M. D. Trott. Prospicient real-time coding of Markov sources over burst erasure channels: Lossless case. In Storer and Marcellin [SM12], pages 267–276. ISBN 1-4673-0715-7 (paperback). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=6189258. IEEE Com-

puter Society order number P4656.

Earl:2003:ESF

[EL03] E. Earl and R. E. Ladner. Enhanced Sequitur for finding structure in data. In Storer and Cohn [SC03], page ?? ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 2003. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1194044. IEEE Computer

Society Order number PR01896.

Elias:1975:UCS

[Eli75] P. Elias. Universal codeword sets and representations of the integers. IEEE Transactions on Information Theory, IT-21 (2):194–203, March 1975. CODEN IETTAW. ISSN 0018-9448 (print), 1557-9654 (electronic).

Elnahas:1984:DCA

[Eln84] Sharaf Eldin Elnahas. Data Compression with Applications to Digital Radiology. D.Sc. thesis, Washington University, St. Louis, MO, USA, 1984. 162 pp. URL http://search.

proquest.com/docview/303316909.

Eastman:1984:AMC

[ELZC84] Willard L. Eastman, Abraham Lempel, Jacob Ziv, and Martin Cohn. Apparatus and method for compressing data signals and restoring the signal. US Patent 4,558,302., August 7,

1984. URL http://www.google.com/patents/US4464650;
https://en.wikipedia.org/wiki/Graphics_Interchange_

Format. US Patent Application 291,870 filed 10 August 1981. See also the related patent [Wel84a] issued on the same day. This patent was held by Sperry and its successor, Unisys, and vigorously enforced until its expiry on 7 August 2001. It led to multimillion-dollar lawsuit settlements, and the move away from LZW compression to several other methods that are patent free.

Effros:2002:CCO

[EM02] M. Effros and D. Muresan. Codecell contiguity in optimal fixed-rate and entropy-constrained network scalar quantizers. In Storer and Cohn [SC02], pages 312–321. ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 2002. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=999969. IEEE Computer

Society Order Number PR01477.

Ebrahimi-Moghadam:2003:IFB

[EMS03] A. Ebrahimi-Moghadam and S. Shirani. Image foveation based on vector quantization. In Storer and Cohn [SC03], page ?? ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2003. URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

1194045. IEEE Computer Society Order number PR01896.

Ernvall:1984:ELM

[EN84] J. Ernvall and O. Nevalainen. Estimating the length of minimal spanning trees in compression of files. BIT (Nordisk tidskrift for informationsbehandling), 24(1):19–32, March 1984. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&volume=24&issue=1&
spage=19.

Eriksson:2002:ICM

[ENA02] T. Eriksson, M. Novak, and J. B. Anderson. Image coding with the MAP criterion. In Storer and Cohn [SC02], page ?? ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2002. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=999996. IEEE

Computer Society Order Number PR01477.

Eriksson:2003:MCT

[ENA03] T. Eriksson, M. Novak, and J. B. Anderson. MAP criterion trellis source coding for short data sequences. In Storer and Cohn [SC03], pages 43–52. ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 2003. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1193995. IEEE Computer

Society Order number PR01896.

Eriksson:2005:SBV

[ENA05] T. Eriksson, M. Novak, and J. B. Anderson. Short-block variable-rate trellis quantization. In Storer and Cohn [SC05], pages 251–260. ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=
1402186.

Engbert:1993:FHC

[Eng93] Ludger Engbert. On-the-fly Huffman coding. C Users Journal, 11(12):29–??, December 1993. ISSN 0898-9788.

Eisenberger:1965:SSU

[EP65] Isidore Eisenberger and Edward C. Posner. Systematic statistics used for data compression in space telemetry. Journal of the American Statistical Association, 60(309):97–133, March 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2283140.

See corrigenda [EP67].

Eisenberger:1967:CSS

[EP67] Isidore Eisenberger and Edward C. Posner. Corrigenda: Systematic statistics used for data compression in space telemetry. Journal of the American Statistical Association, 62(320): 1518, December 1967. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2283812. See [EP65].

Even:1978:EEC

[ER78] Shimon Even and Michael Rodeh. Economical encoding of commas between strings. Communications of the ACM, 21 (4):315–317, April 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Einarsson:1987:DCD

[ER87] Goran Einarsson and Goran Roth. Data compression of digital color pictures. Computers and Graphics, 11(4):409–426, 1987. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Ekstrand:1998:WDE

[ES98] N. Ekstrand and B. Smeets. Weighting of double exponential distributed data in lossless image compression. In Storer and Cohn [SC98], page ?? ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1998. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=672268. IEEE

Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Ekstrand:2000:SNC

[ES00] N. Ekstrand and B. Smeets. Some notes on the context mapping function in lossless data compression. In Storer and Cohn [SC00], page ?? ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2000. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=838200. This

millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Eskicioglu:1995:MDM

[Esk95] A. M. Eskicioglu. A multi-dimensional measure for image quality. In Storer and Cohn [SC95], page 469. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=515579. IEEE catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Elmoataz:2015:LLG

[ETT15] Abderrahim Elmoataz, Matthieu Toutain, and Daniel Tenbrinck. On the p-Laplacian and ∞-Laplacian on graphs with applications in image and data processing. SIAM Journal on Imaging Sciences, 8(4):2412–2451, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Edwards:2014:PAB

[EV14] James A. Edwards and Uzi Vishkin. Parallel algorithms for Burrows–Wheeler compression and decompression. Theoretical Computer Science, 525(??):10–22, March 13, 2014. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304397513007615.

Evans:1979:SRE

[Eva79] Arnold Kerry Evans. Statistical Representation of Electrocardiographic Body Surface Maps for Feature Identification, Data Compression and Classification. Ph.D. thesis, The University of Utah, Salt Lake City, UT 84112, USA, 1979. 96 pp. URL

http://search.proquest.com/docview/302988278.

Evans:1998:CGP

[Eva98] W. S. Evans. Compression via guided parsing. In Storer and Cohn [SC98], page ?? ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1998. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=672269. IEEE

Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Even:1998:FVA

[Eve98] S. Even. Four value-adding algorithms. IEEE Spectrum, 35(5): 33–38, May 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ewing:1996:CJF

[EW96] Gary J. Ewing and Christopher J. Woodruff. Comparison of JPEG and fractal-based image compression on target acquisition by human observers. Optical Engineering, 35(1):284–288, January 1996. CODEN OPEGAR. ISSN 0091-3286.

Eng:1981:SPB

[EY81] K. Y. Eng and O. C. Yue. Spectral properties and bandlimiting effects of time-compressed TV signals in a timecompression multiplexing system. The Bell System Technical Journal, 60(9):2167–2185, November 1981. CODEN BST-

JAN. ISSN 0005-8580. URL http://bstj.bell-labs.com/
BSTJ/images/Vol60/bstj60-9-2167.pdf;   http://www.
alcatel-lucent.com/bstj/vol60-1981/articles/bstj60-
9-2167.pdf.

Fowler:1993:RVB

[FA93] J. E. Fowler and S. C. Ahalt. Robust, variable bit-rate coding using entropy-biased codebooks. In Storer and Cohn [SC93], pages 361–370. ISBN 0-8186-3391-3 (microfiche), 0-8186-3392-1 (casebound). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1993. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=253113. IEEE

Computer Society Press order number 3392-02. IEEE catalog number 93TH0536-3.

Fowler:1994:DVQ

[FA94] J. E. Fowler and S. C. Ahalt. Differential vector quantization of real-time video. In Storer and Cohn [SC94], pages 205–214. ISBN 0-8186-5636-0, 0-8186-5637-9 (case). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1994.

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=305928. IEEE Catalog Number 93TH0626-2. IEEE Computer Society Press Order Number 5637-02.

Fowler:1997:AVQa

[FA97] J. E. Fowler and S. C. Ahalt. Adaptive vector quantization using generalized threshold replenishment. In Storer and Cohn [SC97], pages 317–326. ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1997.

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=582055. IEEE Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Fowler:1995:QBJ

[FAEY95] B. Fowler, R. Arps, A. El Gamal, and D. Yang. Quadtree based JBIG compression. In Storer and Cohn [SC95], pages 102– 111. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=515500. IEEE catalog

number 95TH8037. IEEE Computer Society Press order number PR07010.

Faller:1973:ASD

[Fal73] N. Faller. An adaptive system for data compression. In Record of the 7th Asilomar Conference on Circuits, Systems, and Computers, pages 593–597. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1973.

Fano:1949:TI

[Fan49] R. M. Fano. The transmission of information. Technical Report 65, Research Laboratory for Electronics, MIT, Cambridge, MA, USA, 1949. ?? pp.

Fang:1966:IIE

[Fan66] I. Fang. It isn’t ETAOIN SHRDLU; It’s ETAONI RSHDLC. Journalism Quarterly, 43:761–762, 1966. CODEN JOQUAX. ISSN 0196-3031.

Fawcett:1996:CCN

[Faw96] R. Fawcett. Combination coding: a new entropy coding technique. In Storer and Cohn [SC96], page ?? ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1996. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=488366. IEEE Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Fenwick:1998:CUF

[FB98] P. Fenwick and S. Brierley. Compression of Unicode files. In Storer and Cohn [SC98], page ?? ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1998. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=672274. IEEE Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Fu:2016:FPP

[FB16] Yongquan Fu and Ernst Biersack. False-positive probability and compression optimization for tree-structured Bloom filters. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 1(4):19:1–19:39, September 2016. CODEN ???? ISSN 2376-3639 (print), 2376-

3647 (electronic). URL http://dl.acm.org/citation.cfm?
id=2940324.

Finamore:2004:VPE

[FBS04] W. A. Finamore, S. V. B. Bruno, and D. Silva. Vector permutation encoding for the uniform sources. In Storer and Cohn [SC04], page ?? ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1281515. IEEE Computer Society Order Number: P2082.

Frost:1991:DPR

[FBX91] R. L. Frost, C. F. Barnes, and F. Xu. Design and performance of residual quantizers. In Storer and Reif [SR91b], pages 129–138. ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1990. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=213380. IEEE

Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Fagin:1994:RPF

[FC94] B. Fagin and P. Chintrakulchai. A reprogrammable processor for fractal image compression. Lecture Notes in Computer Science, 849:129–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fan:1998:PPL

[FC98] G. L. Fan and W. K. Cham. Post-processing for low bitrate wavelet-based image coding using multiscale edge characterization. In Storer and Cohn [SC98], page ?? ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1998. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=672270. IEEE Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Fox:1991:OPM

[FCDH91] Edward A. Fox, Qi Fan Chen, Amjad M. Daoud, and Lenwood S. Heath. Order preserving minimal perfect hash functions and information retrieval. ACM Transactions on Information Systems, 9(3):281–308, July 1991. CODEN ATISET. ISSN 1046-8188. URL http://www.acm.org:80. Special Issue on Research and Development in Information Retrieval.

Fiorini:2013:DTDb

[FCL13] Rodolfo A. Fiorini, Andrea Condorelli, and Giuseppe Laguteta. Discrete tomography data footprint reduction via natural compression. Fundamenta Informaticae, 125(3–4):273– 284, July 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Frank:2000:TCU

[FCW00] E. Frank, Chang Chui, and I. H. Witten. Text categorization using compression models. In Storer and Cohn [SC00], page ?? ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2000. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=838202. This millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

devices:2006:FFH

[fd06] flac. devices. FLAC format hardware, 2006. URL http://

flac.sourceforge.net/links.html#hardware.

Fox:2020:SAI

[FDH^{+}20] Alyson Fox, James Diffenderfer, Jeffrey Hittinger, Geoffrey Sanders, and Peter Lindstrom. Stability analysis of inline ZFP compression for floating-point data in iterative methods. SIAM Journal on Scientific Computing, 42(5):A2701–A2730, ???? 2020. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

Feng:2011:SRI

[FDKB11] Wu-Chi Feng, Thanh Dang, John Kassebaum, and Tim Bauman. Supporting region-of-interest cropping through constrained compression. ACM Transactions on Multimedia Computing, Communications, and Applications, 7(3):17:1–17:??, August 2011. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Frank:1980:PIT

[FDU80] Amalie J. Frank, J. D. Daniels, and Diane R. Unangst. Progressive image transmission using a growth-geometry coding. Proceedings of the IEEE, 68(7):897–909, July 1980. CODEN IEEPAD. ISSN 0018-9219 (print), 1558-2256 (electronic).

Frank-Dayan:2000:ULB

[FDZ00] Y. Frank-Dayan and R. Zamir. Universal lattice-based quantizers for multiple descriptions. In Storer and Cohn [SC00], pages 500–509. ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2000. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=838190. This

millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Fleming:1999:GMD

[FE99] M. Fleming and M. Effros. Generalized multiple description vector quantization. In Storer and Cohn [SC99], pages 3–12. ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1999. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=755648. IEEE Computer Society Or-

der Number PR00096. IEEE Order Plan Catalog Number PR00096.

Fleming:2001:NVQ

[FE01] M. Fleming and M. Effros. Network vector quantization. In Storer and Cohn [SC01c], pages 13–22. ISBN 0-7695-1031-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 2001. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=917132. IEEE CSP num-

ber 01PR1031.

Faridi:2006:DCQ

[FE06] A. Faridi and A. Ephremides. Distortion control for queues with deadlines. In Storer and Cohn [SC06], pages 312–321. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1607266.

Feder:1988:F

[Fed88] Jens Feder. Fractals. Plenum Press, New York, NY, USA; London, UK, 1988. ?? pp.

FBI:1993:WGF

[Fed93] Federal Bureau of Investigation. WSQ grayscale fingerprint image compression specification, ver. 2.0. Document IAFIS-IC-0110v2, Criminal Justice Information Services, Washington, DC, USA (??), February 1993. ?? pp.

Felt:1974:DCT

[Fel74] Douglas Charles Felt. Data-Compression Techniques Using Adaptive Translation-Permutation Codes. Ph.D. thesis, Washington State University, Pullman, WA, USA, 1974. 113 pp.

URL http://search.proquest.com/docview/302713192.

Feldspar:2003:EDA

[Fel03] Antaeus Feldspar. An explanation of the Deflate algorithm,

2003. URL http://www.zlib.org/feldspar.html. This ar-

ticle was originally posted to comp.compression on 23 August 1997.

Fenwick:1993:ZLE

[Fen93] P. M. Fenwick. Ziv–Lempel encoding with multi-bit flags. In Storer and Cohn [SC93], pages 138–147. ISBN 0-8186-3391-3 (microfiche), 0-8186-3392-1 (casebound). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1993. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=253136. IEEE Computer Society Press order number 3392-02. IEEE catalog number 93TH0536-3.

Fenwick:1995:DZL

[Fen95] P. Fenwick. Differential Ziv–Lempel text compression. J.UCS: The Journal of Universal Computer Science, 1(8): 591–602, August 28, 1995. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://www.jucs.org/

differential_ziv_lempel_text_compression.

Fenwick:1996:SRT

[Fen96a] P. Fenwick. Symbol ranking text compression. Technical Report 132, Department of Computer Science, University of Auckland, Auckland, NZ, June 1996. ?? pp.

Fenwick:1996:PEC

[Fen96b] Peter Fenwick. Punctured Elias codes for variable-length coding of the integers. Technical Report 137, Department of Computer Science, The University of Auckland, Auckland, NZ,

December 1996. 15 pp. URL http://www.firstpr.com.au/
audiocomp/lossless/TechRep137.pdf.

Fenwick:1997:SRT

[Fen97] P. Fenwick. Symbol ranking text compressors. In Storer and Cohn [SC97], page ?? ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1997. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=582093. IEEE

Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Fenwick:2002:VLI

[Fen02] P. Fenwick. Variable-length integer codes based on the Goldbach Conjecture, and other additive codes. IEEE Transactions on Information Theory, 48(8):2412–2417, August 2002. CODEN IETTAW. ISSN 0018-9448 (print), 1557-9654 (electronic).

Fenwick:2012:PCE

[Fen12] P. M. Fenwick. PPM compression without escapes. Software — Practice and Experience, 42(2):255–260, February 2012.

CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Fetik:1996:MXG

[Fet96] Richard Fetik. Multimedia X: Generating GIFs on the fly. The X Journal: Computing Technology with the X Window System, 5(5):86–??, May 1996. CODEN XJOUEA. ISSN 1056-7003.

Fazzinga:2013:RDC

[FFFM13] Bettina Fazzinga, Sergio Flesca, Filippo Furfaro, and Elio Masciari. RFID-data compression for supporting aggregate queries. ACM Transactions on Database Systems, 38(2):11:1– 11:??, June 2013. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Farruggia:2019:BDC

[FFFV19] Andrea Farruggia, Paolo Ferragina, Antonio Frangioni, and Rossano Venturini. Bicriteria data compression. SIAM Journal on Computing, 48(5):1603–1642, ???? 2019. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Federovsky:1998:BPB

[FFW98] Eitan Federovsky, Meir Feder, and Sholomo Weiss. Branch prediction based on universal data compression algorithms. ACM SIGARCH Computer Architecture News, 26(3):62–72, June 1998. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Fiala:1989:DCF

[FG89] E. R. Fiala and D. H. Greene. Data compression with finite Windows. Communications of the ACM, 32(4):490–505, April 1989. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.org/pubs/

toc/Abstracts/0001-0782/63341.html.

Fox:1997:CAT

[FG97] A. Fox and T. Grun. Compressing address trace data for cache simulations. In Storer and Cohn [SC97], page ?? ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1997. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=582096. IEEE Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Flierl:2001:MME

[FG01] M. Flierl and B. Girod. Multihypothesis motion estimation for video coding. In Storer and Cohn [SC01c], pages 341–350. ISBN 0-7695-1031-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2001. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=917165. IEEE CSP number 01PR1031.

Flierl:2007:HPA

[FG07] M. Flierl and B. Girod. Half-pel accurate motion-compensated orthogonal video transforms. In Storer and Cohn [SC07], pages 13–22. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=4148740. IEEE

Computer Society Order Number P2791.

Feygin:1993:MEV

[FGC93] G. Feygin, P. G. Gulak, and P. Chow. Minimizing error and VLSI complexity in the multiplication free approximation of arithmetic coding. In Storer and Cohn [SC93], pages 118– 127. ISBN 0-8186-3391-3 (microfiche), 0-8186-3392-1 (casebound). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 1993. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=253138. IEEE Computer

Society Press order number 3392-02. IEEE catalog number 93TH0536-3.

Feygin:1994:AAV

[FGC94] G. Feygin, P. G. Gulak, and P. Chow. Architectural advances in the VLSI implementation of arithmetic coding for binary image compression. In Storer and Cohn [SC94], pages 254–263. ISBN 0-8186-5636-0, 0-8186-5637-9 (case). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1994.

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=305933. IEEE Catalog Number 93TH0626-2. IEEE Computer Society Press Order Number 5637-02.

Ferragina:2008:SCS

[FGG^{+}08] Paolo Ferragina, Roberto Grossi, Ankur Gupta, Rahul Shah, and Jeffrey Scott Vitter. On searching compressed string collections cache-obliviously. In Lenzerini and Lembo [LL08], pages 181–190. ISBN 1-605-60932-3. LCCN ????

Foschini:2004:FCS

[FGGV04] L. Foschini, R. Grossi, A. Gupta, and J. S. Vitter. Fast compression with a static model in high-order entropy. In Storer and Cohn [SC04], pages 62–71. ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1281451. IEEE Computer Society Order Number: P2082.

Fielding:1999:RHT

[FGM^{+}99] R. Fielding, J. Gettys, J. Mogul, H. Frystyk, L. Masinter, P. Leach, and T. Berners-Lee. RFC 2616: Hypertext Transfer Protocol — HTTP/1.1, 1999. URL http://www.faqs.org/

rfcs/rfc2616.html.

Fulber-Garcia:2021:LDF

[FGM21] Vinicius Fulber-Garcia and Sérgio Luis Sardi Mergen. LUISA: Decoupling the frequency model from the context model in prediction-based compression. The Computer Journal, 64(9): 1437–1450, September 2021. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://academic.

oup.com/comjnl/article/64/9/1437/5868170.

Ferragina:2005:BTC

[FGMS05] Paolo Ferragina, Raffaele Giancarlo, Giovanni Manzini, and Marinella Sciortino. Boosting textual compression in optimal linear time. Journal of the ACM, 52(4):688–713, July 2005. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Frey:1996:FEC

[FH96] B. J. Frey and G. E. Hinton. Free energy coding. In Storer and Cohn [SC96], pages 73–81. ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1996. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=488312. IEEE Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Fojtik:1998:IMP

[FH98] J. Fojtik and V. Hlavac. Invisible modification of the palette color image enhancing lossless compression. Lecture Notes in Computer Science, 1451:1029–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fowler:2002:OBM

[FH02] J. E. Fowler and Li Hua. Omnidirectionally balanced multiwavelets for vector wavelet transforms. In Storer and Cohn [SC02], pages 422–431. ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2002. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=999982. IEEE Computer Society Order Number PR01477.

Fiala:2008:DUS

[FH08] M. Fiala and J. Holub. DCA using suffix arrays. In Storer and Marcellin [SM08], page 516. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33 INTERNET. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=4483343.

Fischer:2008:PEB

[FHS08] J. Fischer, V. Heun, and H. M. Stiihler. Practical entropybounded schemes for O(1)-range minimum queries. In Storer and Marcellin [SM08], pages 272–281. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33 INTERNET.

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=
&arnumber=4483305.

Fischer:2018:LZF

[FIKS18] Johannes Fischer, Tomohiro I, Dominik Köppl, and Kunihiko Sadakane. Lempel–Ziv factorization powered by space efficient suffix trees. Algorithmica, 80(7):2048–2081, July 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

FIPS:2001:AAE

[FIP01] FIPS. Announcing the Advanced Encryption Standard (AES). National Institute for Standards and Technology, Gaithersburg, MD, USA, November 26, 2001. 51 pp. URL http:

//csrc.nist.gov/publications/fips/fips197/fips-197.

pdf. Federal Information Processing Standards Publication 197.

Fisher:1995:FIC

[Fis95a] Yuval Fisher, editor. Fractal Image Compression: Theory and Application. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 0-387-94211-4. ?? pp.

Fisher:1995:NAF

[Fis95b] Yuval Fisher, editor. NATO ASI on Fractal Image Encoding and Analysis: July 8–17, 1995, Trondheim, Norway. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 3-540-63196-8. LCCN TA1637 .N38 1998.

Franques:1996:EWB

[FJ96] V. T. Franques and V. K. Jain. Enhanced wavelet-based zerotree coding of images. In Storer and Cohn [SC96], page ?? ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1996. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=488368. IEEE

Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Forchhammer:2000:CLP

[FJ00] S. Forchhammer and O. R. Jensen. Content layer progressive coding of digital maps. In Storer and Cohn [SC00], pages 233–242. ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2000. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=838163. This

millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Fang:2007:VBO

[FJ07] Yong Fang and Jechang Jeong. VLR-based optimal positioning of resynchronization markers. In Storer and Cohn [SC07], page 381. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=4148782. IEEE

Computer Society Order Number P2791.

Fu:2011:CMV

[FJD11] Changjun Fu, Xiangyang Ji, and Qionghai Dai. Compressed multi-view imaging with joint reconstruction. In Storer and Marcellin [SM11b], page 448. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=5749505. IEEE Computer Society Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Fraenkel:1985:NCS

[FK85a] A. S. Fraenkel and S. T. Klein. Novel compression of sparse bit-strings—preliminary report. In A. Apostolico and Z. Galil, editors, Combinatorial Algorithms on Words, volume 12 of NATO ASI Series F, pages 169–183. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985.

Fraenkel:1985:RUC

[FK85b] A. S. Fraenkel and Shmuel T. Klein. Robust universal complete codes as alternatives to Huffman codes. Technical Report CS85-16, Department of Applied Mathematics, Weizmann Institute of Science, Rehovot, Israel, October 1985. ?? pp.

Fraenkel:1990:BHC

[FK90] Aviezri S. Fraenkel and Shmuel T. Klein. Bidirectional Huffman coding. The Computer Journal, 33(4):296–307, August 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/
content/33/4/296.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_33/Issue_04/tiff/
296.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_04/tiff/297.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_33/Issue_04/
tiff/298.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_04/tiff/299.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_33/Issue_04/
tiff/300.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_04/tiff/301.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_33/Issue_04/

tiff/302.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_04/tiff/303.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_33/Issue_04/
tiff/304.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_04/tiff/305.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_33/Issue_04/
tiff/306.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_33/Issue_04/tiff/307.tif.

Ferens:1995:AWS

[FK95] K. Ferens and W. Kinsner. Adaptive wavelet subband coding for music compression. In Storer and Cohn [SC95], page 460. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=515570. IEEE catalog

number 95TH8037. IEEE Computer Society Press order number PR07010.

Fraenkel:1996:RUC

[FK96] Aviezri S. Fraenkel and Shmuel T. Klein. Robust universal complete codes for transmission and compression. Discrete Applied Mathematics, 64(1):31–55, January 1996. CODEN DAMADU. ISSN 0166-218X (print), 1872-6771 (electronic).

Freedman:2009:FDR

[FK09] D. Freedman and P. Kisilev. Fast data reduction via KDE approximation. In Storer and Marcellin [SM09], page 445. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=4976499. IEEE Computer

Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

Feig:1990:DCT

[FL90] Ephraim N. Feig and Elliot Linzer. Discrete cosine transform algorithms for image data compression. In Proceedings Electronic Imaging ’90 East, Boston, MA, USA, pages 84–87. ????, ????, 1990.

Fortnow:1998:NOL

[FL98] L. Fortnow and S. Laplante. Nearly optimal language compression using extractors. Lecture Notes in Computer Science, 1373:84–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Faro:2013:EOS

[FL13] Simone Faro and Thierry Lecroq. The exact online string matching problem: a review of the most recent results. ACM Computing Surveys, 45(2):13:1–13:??, February 2013. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

Freytag:2003:VPI

[FLA^{+}03] Johann Christoph Freytag, Peter C. Lockemann, Serge Abiteboul, Michael J. Carey, Patricia G. Selinger, and Andreas Heuer, editors. VLDB 2003: Proceedings of 29th International Conference on Very Large Data Bases, September 9–12, 2003, Berlin, Germany. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2003. ISBN 0-12-722442-4. LCCN ????

Ferragina:2009:CIL

[FLMM09] Paolo Ferragina, Fabrizio Luccio, Giovanni Manzini, and S. Muthukrishnan. Compressing and indexing labeled trees, with applications. Journal of the ACM, 57(1):4:1–4:33, November 2009. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Franaszek:2006:DCR

[FLMPR06] P. A. Franaszek, L. A. Lastras-Montano, Song Peng, and J. T. Robinson. Data compression with restricted parsings. In Storer and Cohn [SC06], pages 203–212. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=1607255.

Fan:2020:ALC

[FLS20] Q. Fan, D. J. Lilja, and S. S. Sapatnekar. Adaptive-length coding of image data for low-cost approximate storage. IEEE Transactions on Computers, 69(2):239–252, February 2020. CODEN ITCOB4. ISSN 2326-3814.

Fu:2010:TSB

[FLZW10] Jingjing Fu, Zhouchen Lin, Bing Zeng, and Feng Wu. Tree structure based analyses on compressive sensing for binary sparse sources. In Storer and Marcellin [SM10b], page 530. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=5453509.

Fowell:1989:FPF

[FM89] Richard A. Fowell and David D. McNeil. Faster plots by fan data-compression. IEEE Computer Graphics and Applications, 9(2):58–66, March 1989. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Feder:1995:UCA

[FM95a] M. Feder and N. Merhav. Universal coding for arbitrarily varying sources and for hierarchies of model classes. In Storer and Cohn [SC95], pages 82–91. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=515498. IEEE catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Feder:1995:CPG

[FM95b] T. Feder and R. Motwani. Clique partitions, graph compression and speeding-up algorithms. Journal of Computer and System Sciences, 51(2):261–272, October 1995. CO-DEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022000085710653.

Franceschini:1996:DCU

[FM96] R. W. Franceschini and A. Mukherjee. Data compression using encrypted text. In Storer and Cohn [SC96], page ?? ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1996. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=488369. IEEE Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Fout:2012:APB

[FM12] Nathaniel Fout and Kwan-Liu Ma. An adaptive predictionbased approach to lossless compression of floating-point volume data. IEEE Transactions on Visualization and Computer Graphics, 18(12):2295–2304, December 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ferragina:2007:CRS

[FMMN07] Paolo Ferragina, Giovanni Manzini, Veli Mäkinen, and Gonzalo Navarro. Compressed representations of sequences and full-text indexes. ACM Transactions on Algorithms, 3(2):20:1– 20:??, May 2007. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Fraenkel:1983:TCP

[FMP83] Aviezri S. Fraenkel, Moshe Mor, and Y. Perl. Is text compression by prefixes and suffixes practical? Acta Informatica, 20 (4):371–389, December 1983. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Franti:1993:TSM

[FN93] Pasi Fränti and Olli Nevalainen. A two-stage modelling method for compressing binary images by arithmetic coding. The Computer Journal, 36(7):615–622, ???? 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/
content/36/7/615.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/Volume_36/Issue_07/Vol36_07.
body.html#AbstractFranti.

Fredriksson:2009:SRA

[FN09] Kimmo Fredriksson and Fedor Nikitin. Simple random access compression. Fundamenta Informaticae, 92(1–2):63–81, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Funasaka:2017:ALL

[FNI17] Shunji Funasaka, Koji Nakano, and Yasuaki Ito. Adaptive loss-less data compression method optimized for GPU decompression. Concurrency and Computation: Practice and Experience, 29(24):??, December 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Franti:1994:CDI

[FNK94] Pasi Fränti, Olli Nevalainen, and Timo Kaukoranta. Compression of digital images by block truncation coding: a survey. The Computer Journal, 37(4):308–332, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/
content/37/4/308.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/Volume_37/Issue_04/Vol37_04.
body.html#AbstractFranti.

Farina:2008:WBS

[FNP08] A. Farina, G. Navarro, and J. R. Parama. Word-based statistical compressors as natural language compression boosters. In Storer and Marcellin [SM08], pages 162–171. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33 INTER-

NET. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=4483294.

Ferragina:2011:OPT

[FNV11] Paolo Ferragina, Igor Nitto, and Rossano Venturini. On optimally partitioning a text to improve its compression. Algorithmica, 61(1):51–74, September 2011. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=61&issue=1&spage=51.

Fan:2010:OOT

[FO10] Zihong Fan and A. Ortega. Optimization of overlapped tiling for efficient 3D image retrieval. In Storer and Marcellin [SM10b], pages 494–503. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ????

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=
&arnumber=5453471.

Ferreira:2009:USA

[FOF09] A. Ferreira, A. Oliveira, and M. Figueiredo. On the use of suffix arrays for memory-efficient Lempel–Ziv data compression. In Storer and Marcellin [SM09], page 444. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4976498. IEEE Computer Society

Order Number P3592. BMS Part Number CFP09DCC-PRT.

Ferreira:2011:SWU

[FOF11] A. Ferreira, A. Oliveira, and M. Figueiredo. Sliding window update using suffix arrays. In Storer and Marcellin [SM11b], page 456. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=5749513. IEEE

Computer Society Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Fan:2009:OTF

[FOhC09] Zihong Fan, A. Ortega, and Cheng hao Chien. Overlapped tiling for fast random oblique plane access of 3D object datasets. In Storer and Marcellin [SM09], pages 163–172. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=4976460. IEEE Computer

Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

Fong:1981:DOA

[Fon81] Chung-Bin Fong. The Decision Oriented Approach to Information-Preserving Data Compression of Digitized Images. Ph.D. thesis, University of Pittsburgh, Pittsburgh, PA,

USA, 1981. 147 pp. URL http://search.proquest.com/
docview/303140921.

Farina:2012:ISI

[FOP12] A. Farina, A. Ordonez, and J. R. Parama. Indexing sequences of IEEE 754 double precision numbers. In Storer and Marcellin [SM12], pages 367–376. ISBN 1-4673-0715-7 (paperback). ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=6189268. IEEE Computer Society

order number P4656.

Forkert:1996:EAW

[For96] R. D. Forkert. Elements of adaptive wavelet image compression. In Storer and Cohn [SC96], page ?? ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1996. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=488367. IEEE Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Forchhammer:1999:ICU

[For99] S. Forchhammer. Image coding using Markov models with hidden states. In Storer and Cohn [SC99], page ?? ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1999. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=785681. IEEE Computer Society Or-

der Number PR00096. IEEE Order Plan Catalog Number PR00096.

Fan:2017:MPC

[FOSS17] Xinxin Fan, Adilet Otemissov, Francesco Sica, and Andrey Sidorenko. Multiple point compression on elliptic curves. Designs, Codes, and Cryptography, 83(3):565–588, June 2017. CODEN DCCREC. ISSN 0925-1022 (print), 1573-7586 (electronic).

Foulds:1985:DRT

[Fou85] Richard Arthur Foulds. A Data Reduction Technique for Minimum Bandwidth Transmission of Intelligible American Sign Language for the Deaf (Compression, Telecommunications, Communication, Rehabilitation). Ph.D. thesis, Tufts University, Medford, MA, USA, 1985. 215 pp. URL http://search.

proquest.com/docview/303423885.

Fowler:1997:AVQb

[Fow97a] J. E. Fowler. Adaptive vector quantization — Part I: a unifying structure. In Storer and Cohn [SC97], page ?? ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1997. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=582094. IEEE Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Fowler:1997:AVQc

[Fow97b] J. E. Fowler. Adaptive vector quantization. II. Classification and comparison of algorithms. In Storer and Cohn [SC97], page ?? ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1997. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=582095. IEEE

Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Fowler:1998:VCU

[Fow98] J. E. Fowler. Video coding using vector zerotrees and adaptive vector quantization. In Storer and Cohn [SC98], page ?? ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1998. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=672275. IEEE Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Fowler:2000:QOS

[Fow00] J. E. Fowler. QccPack: an open-source software library for quantization, compression, and coding. In Storer and Cohn [SC00], page ?? ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2000. URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=
838201;   http://www.ece.msstate.edu/~fowler/QccPack.

This millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Fowler:2006:ARW

[Fow06] J. E. Fowler. Analysis of redundant-wavelet multihypothesis for motion compensation. In Storer and Cohn [SC06], pages 352–361. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=1607270.

Fowler:2008:CPP

[Fow08] J. E. Fowler. Compressive-projection principal component analysis for the compression of hyperspectral signatures. In Storer and Marcellin [SM08], pages 83–92. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33 INTERNET.

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=
&arnumber=4483286.

Fowler:2009:CPP

[Fow09] J. E. Fowler. Compressive-projection principal component analysis and the first eigenvector. In Storer and Marcellin [SM09], pages 223–232. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ????

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=4976466. IEEE Computer Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

Fellner:2019:GPP

[FP19] Andreas Fellner and Bruno Woltzenlogel Paleo. Greedy pebbling for proof space compression. International Journal on Software Tools for Technology Transfer (STTT), 21(1):71– 86, February 2019. CODEN ???? ISSN 1433-2779 (print),

1433-2787 (electronic). URL https://link.springer.com/
article/10.1007/s10009-017-0459-0;   https://link.
springer.com/content/pdf/10.1007/s10009-017-0459-0.
pdf.

Feischl:2020:SCE

[FP20] Michael Feischl and Daniel Peterseim. Sparse compression of expected solution operators. SIAM Journal on Numerical Analysis, 58(6):3144–3164, ???? 2020. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Feig:1995:ICU

[FPR95] Ephraim Feig, Heidi Peterson, and Viresh Ratnakar. Image compression using spatial prediction. In IEEE [IEE95b], pages 2339–2342. CODEN IPRODJ. ISBN 0-7803-2431-5. ISSN 0736-7791. LCCN TK7882.S65 .I16 1995. Five volumes. IEEE Catalog No. 95CH35732.

Ferguson:1984:SSH

[FR84] T. J. Ferguson and J. H. Rabinowitz. Self-synchronizing Huffman codes. IEEE Transactions on Information Theory, 30 (4):687–693, July 1984. CODEN IETTAW. ISSN 0018-9448 (print), 1557-9654 (electronic).

Forchhammer:1995:CPH

[FR95] S. Forchhammer and J. Rissanen. Coding with partially hidden Markov models. In Storer and Cohn [SC95], pages 92– 101. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=515499. IEEE catalog

number 95TH8037. IEEE Computer Society Press order number PR07010.

Floor:2006:NIM

[FR06] P. A. Floor and T. A. Ramstad. Noise immunity for 1:N and M:1 nonlinear mappings for source-channel coding. In Storer and Cohn [SC06], page 446. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=1607289.

Fraser:1967:DCA

[Fra67] A. G. Fraser. Data compression and automatic programming. The Computer Journal, 10(2):165–167, August 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/10/2/165.full.pdf+html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_10/Issue_02/
100165.sgm.abs.html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_02/tiff/165.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_10/Issue_
02/tiff/166.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_02/tiff/167.tif.

Franz:1997:ACS

[Fra97] M. Franz. Adaptive compression of syntax trees and iterative dynamic code optimization: Two basic technologies for mobile object systems. Lecture Notes in Computer Science, 1222:263– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fraser:2006:IDI

[Fra06] Christopher W. Fraser. An instruction for direct interpretation of LZ77-compressed programs. Software — Practice and Experience, 36(4):397–411, ???? 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Freeman:1961:EAG

[Fre61] Herbert Freeman. On the encoding of arbitrary geometric configurations. IRE Transactions on Electronic Computers, EC-10(2):260–268, June 1961. CODEN IRELAO.

ISSN 0367-9950. URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5219197.

Freeman:1991:ACD

[Fre91] G. H. Freeman. Asymptotic convergence of dual-tree entropy codes. In Storer and Reif [SR91b], pages 208–217. ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1990. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=213360. IEEE Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Freeman:1993:DCV

[Fre93] G. H. Freeman. Divergence and the construction of variableto-variable-length lossless codes by source-word extensions. In Storer and Cohn [SC93], pages 79–88. ISBN 0-8186-3391-3 (microfiche), 0-8186-3392-1 (casebound). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1993. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=253142. IEEE Computer Society Press order number 3392-02. IEEE catalog number 93TH0536-3.

Frey:1998:GMM

[Fre98] Brendan J. Frey. Graphical models for machine learning and digital communication. Adaptive computation and machine learning. MIT Press, Cambridge, MA, USA, 1998. ISBN 0-262-06202-X. xiii + 195 pp. LCCN Q325.5 .F74 1998. URL http:/

/cognet.mit.edu/library/books/view?isbn=026206202X.

Franaszek:1996:PCC

[FRT96] P. Franaszek, J. Robinson, and J. Thomas. Parallel compression with cooperative dictionary construction. In Storer and Cohn [SC96], pages 200–209. ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1996. URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

488325. IEEE Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Fryer:1985:SDC

[Fry85] Jeffrey Keefe Fryer. A survey of data compression techniques and their application to VSPC data sets. M.Sc.C.S., The University of New Brunswick, Fredericton, NB, Canada, 1985.

???? pp. URL http://search.proquest.com/docview/
303433066.

Feder:1998:UDC

[FS98] M. Feder and A. C. Singer. Universal data compression and linear prediction. In Storer and Cohn [SC98], pages 511–520. ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1998. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=672225. IEEE

Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Forchhammer:2002:PCP

[FS02] S. Forchhammer and J. M. Salinas. Progressive coding of palette images and digital maps. In Storer and Cohn [SC02], pages 362–371. ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2002. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=999974. IEEE Computer Society Order Number PR01477.

Florencio:2003:CSB

[FS03] D. A. Florencio and P. Simard. Can the sample being transmitted be used to refine its own PDF estimate? In Storer and Cohn [SC03], pages 233–242. ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 2003. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1194014. IEEE Computer

Society Order number PR01896.

Faramarzpour:2004:LLC

[FS04] N. Faramarzpour and S. Shirani. Lossless and lossy compression of DNA microarray images. In Storer and Cohn [SC04], page ?? ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=1281514. IEEE

Computer Society Order Number: P2082.

Fang:1991:NNB

[FSC91] W. C. Fang, B. J. Sheu, and O. T.-C. Chen. A neural network based VLSI vector quantizer for real-time image compression. In Storer and Reif [SR91b], pages 342–351. ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1990. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=213346. IEEE Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Filgueira:2011:ACE

[FSC^{+}11] Rosa Filgueira, David E. Singh, Jesús Carretero, Alejandro Calderón, and Félix García. Adaptive-CoMPI: Enhancing MPI-based applications’ performance and scalability by using adaptive compression. The International Journal of High Performance Computing Applications, 25(1):93–114, February 2011. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846

(electronic). URL http://hpc.sagepub.com/content/25/1/
93.full.pdf+html.

Forchhammer:2012:RAB

[FSL^{+}12] S. Forchhammer, M. Salmistraro, K. J. Larsen, Xin Huang, and Huynh Van Luong. Rate-adaptive BCH coding for Slepian–Wolf coding of highly correlated sources. In Storer and Marcellin [SM12], pages 237–246. ISBN 1-4673-0715-7 (paperback). ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=6189255. IEEE Computer Society

order number P4656.

Factor:2001:SCC

[FSY01] M. Factor, D. Sheinwald, and B.-A. Yassour. Software compression in the client/server environment. In Storer and Cohn [SC01c], pages 233–242. ISBN 0-7695-1031-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2001. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=917154. IEEE CSP number 01PR1031.

Farach:1995:SML

[FT95] Martin Farach and Mikkel Thorup. String matching in Lempel–Ziv compressed strings. In ACM [ACM95], pages 703–712. ISBN 0-89791-718-9. LCCN QA 76.6 A13 1995.

URL http://www.acm.org/pubs/articles/proceedings/
stoc/225058/p703-farach/p703-farach.pdf;   http://
www.acm.org/pubs/citations/proceedings/stoc/225058/

p703-farach/. ACM order no. 508950.

Farach:1998:SML

[FT98] Martin Farach and Mikkel Thorup. String matching in Lempel–Ziv compressed strings. Algorithmica, 20(4):388–404, April 1998. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-

0541 (electronic). URL http://www.springerlink.com/
openurl.asp?genre=article&issn=0178-4617&volume=20&
issue=4&spage=388.

Forst:2000:MHT

[FT00] Gunnar Forst and Anders Thorup. Minimal Huffman trees. Acta Informatica, 36(9/10):721–734, April 2000. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/0036009/00360721.htm;
http://link.springer-ny.com/link/service/journals/
00236/papers/0036009/00360721.pdf.

Fredriksson:2004:ESM

[FT04] Kimmo Fredriksson and Jorma Tarhio. Efficient string matching in Huffman compressed texts. Fundamenta Informaticae, 63(1):1–16, January 2004. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Fenwick:2003:BWA

[FTL03] P. Fenwick, M. Titchener, and M. Lorenz. Burrows Wheeler — alternatives to move to front. In Storer and Cohn [SC03], page ?? ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2003. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1194047. IEEE Computer Society Order number PR01896.

Freina:1998:PFI

[FU98] T. Freina and A. Uhl. Predictive fractal image coding: hybrid algorithms and compression of residuals. In Storer and Cohn [SC98], page ?? ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1998. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=672277. IEEE

Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Fu:2003:RND

[Fu03] Min Fu. The research on nonlinear data compression and reconstruction based on MRA. M.S., Huazhong (Central China) University of Science and Technology, Huazhong, Peoples Re-

public of China, 2003. URL http://search.proquest.com/
docview/1024720586.

Funet:2001:TI

[Fun07] Funet. Test images, 2001, 2007. URL ftp://nic.funet.fi/

pub/graphics/misc/test-images/.

Ferragina:2016:CCO

[FV16] Paolo Ferragina and Rossano Venturini. Compressed cacheoblivious string B-tree. ACM Transactions on Algorithms, 12 (4):52:1–52:??, September 2016. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Ferragina:2020:PIF

[FV20] Paolo Ferragina and Giorgio Vinciguerra. The PGMindex: a fully-dynamic compressed learned index with provable worst-case bounds. Proceedings of the VLDB Endowment, 13(8):1162–1175, April 2020. CODEN ???? ISSN

2150-8097. URL https://dl.acm.org/doi/abs/10.14778/
3389133.3389135.

Fresia:2008:DSC

[FVP08] M. Fresia, L. Vandendorpe, and H. V. Poor. Distributed source coding using Raptor codes for hidden Markov sources. In Storer and Marcellin [SM08], page 517. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33 INTERNET.

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=
&arnumber=4483344.

Fernandez:2010:EAC

[FVW10] F. Fernandez, A. Viola, and M. J. Weinberger. Efficient algorithms for constructing optimal bi-directional context sets. In Storer and Marcellin [SM10b], pages 179–188. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5453460.

Fischer:1991:ECT

[FW91] T. R. Fischer and M. Wang. Entropy-constrained trellis coded quantization. In Storer and Reif [SR91b], pages 103–112. ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1990. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=213377. IEEE Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Fulscher:1993:ERI

[FW93] Markus P. Fülscher and Per-Olof Widmark. An electron repulsion integral compression algorithm. Journal of Computational Chemistry, 14(1):8–12, January 1993. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Fenwick:1995:VQW

[FW95] P. M. Fenwick and S. A. Woolford. Vector quantisation for wavelet based image compression. In Storer and Cohn [SC95], page 465. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=515575. IEEE

catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Fowler:2002:JEC

[FW02] J. E. Fowler and Y. Wang. Joint embedded coding of data and grid using first-generation wavelet transforms. In Storer and Cohn [SC02], pages 432–441. ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2002. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=999983. IEEE Computer Society Order Number PR01477.

Forchhammer:2001:LID

[FWA01] S. Forchhammer, Xiaolin Wu, and J. D. Andersen. Lossless image data sequence compression using optimal context quantization. In Storer and Cohn [SC01c], pages 53–62. ISBN 0-7695-1031-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 2001. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=917136. IEEE CSP num-

ber 01PR1031.

Flierl:1998:LOD

[FWG98] M. Flierl, T. Wiegand, and B. Girod. A locally optimal design algorithm for block-based multi-hypothesis motioncompensated prediction. In Storer and Cohn [SC98], pages 239–248. ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1998. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=672152. IEEE

Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Furht:1999:PRM

[FWI99] Borko Furht, Raymond Westwater, and Jeffrey Ice. Project reports: Multimedia broadcasting over the Internet. II. Video compression. IEEE MultiMedia, 6(1):85–89, January–March 1999. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166

(electronic). URL http://dlib.computer.org/mu/books/
mu1999/pdf/u1085.pdf.

Fan:2008:NMD

[FWS^{+}08] Yuhua Fan, Jia Wang, Jun Sun, Peng Wang, and Songyu Yu. A novel multiple description video codec based on Slepian– Wolf coding. In Storer and Marcellin [SM08], page 515. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33 INTER-

NET. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=4483342.

Fan:2011:NTS

[FWS11] Yuhua Fan, Jia Wang, and Jun Sun. On natural type selection in universal multiple description coding. In Storer and Marcellin [SM11b], pages 273–282. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=5749485. IEEE Computer Society Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Feng:1998:LSE

[FWY^{+}98] Wu Feng, Gao Wen, Xiang YangZhao, Gao Peng, and Chen DaTong. On-line sprite encoding with large global motion estimation. In Storer and Cohn [SC98], page ?? ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1998. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=672272. IEEE Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Fu:2008:SIR

[FWZ08] Jingjing Fu, Feng Wu, and Bing Zeng. Spectral information recovery for compressed image restoration. In Storer and Marcellin [SM08], page 518. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008;

QA76.9.D33 INTERNET. URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=4483345.

Fan:2012:DSV

[FWZ^{+}12] Xiaopeng Fan, Feng Wu, Debin Zhao, O. C. Au, and Wen Gao. Distributed soft video broadcast (DCAST) with explicit motion. In Storer and Marcellin [SM12], pages 199–208. ISBN 1-4673-0715-7 (paperback). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=6189251. IEEE Com-

puter Society order number P4656.

Fowler:1995:OLP

[FY95] J. E. Fowler and R. Yagel. Optimal linear prediction for the lossless compression of volume data. In Storer and Cohn [SC95], page 458. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

515568. IEEE catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Farber:2002:QUE

[FZ02] B. Farber and K. Zeger. Quantizers with uniform encoders and channel optimized decoders. In Storer and Cohn [SC02], pages 292–301. ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2002. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=999967. IEEE Computer Society Order Number PR01477.

Farber:2005:QMS

[FZ05] B. Farber and K. Zeger. Quantization of multiple sources using integer bit allocation. In Storer and Cohn [SC05], pages 368– 377. ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=1402198.

Feng:2003:NSC

[FZE03] Hanying Feng, Qian Zhao, and M. Effros. Network source coding using entropy constrained dithered quantization. In Storer and Cohn [SC03], page ?? ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2003. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1194046. IEEE Computer Society Order number PR01896.

Gottlieb:1975:CCM

[G^{+}75] D. Gottlieb et al. A classification of compression methods and their usefulness for a large data processing center. In Proceedings of the National Computer Conference, May 19– 22, 1975, Anaheim, California., volume 44, pages 453–458. AFIPS Press, Montvale, NJ, USA, 1975.

Gilbert:2008:ACG

[GA08] J. Gilbert and D. M. Abrahamson. Adaptive compression of graph structured text. In Storer and Marcellin [SM08], page 519. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33

INTERNET. URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4483346.

Gadallah:1991:DCT

[Gad91] M. E. Gadallah. Data Compression Techniques for Isolated and Connected Word Recognition (Speech Signals, Template Matching). Ph.D. thesis, Cranfield University, Cranfield, Bedfordshire MK43 0AL, UK, 1991. 301 pp. URL http://search.

proquest.com/docview/303965529.

Gage:1994:NAD

[Gag94] Philip Gage. A new algorithm for data compression. C Users Journal, 12(2):23–28, February 1994. ISSN

0898-9788. URL http://www.ddj.com/cpp/184402829;
jsessionid=LGSEIODZNDHKIQSNDLRSKHSCJUNN2JVN?_requestid=
927467.

Gage:1997:RAD

[Gag97] Philip Gage. Random access data compression. C/C++ Users Journal, 15(9):23–??, September 1997. CODEN CCUJEX. ISSN 1075-2838.

Gagie:2004:DSC

[Gag04] T. Gagie. Dynamic Shannon coding. In Storer and Cohn [SC04], page ?? ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

1281516. IEEE Computer Society Order Number: P2082.

Gagie:2006:DAC

[Gag06] T. Gagie. Dynamic asymmetric communication. In Storer and Cohn [SC06], page 447. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=
1607290.

Gaither:1998:TDD

[Gai98] Justin Gaither. A tale of DXPC: Differential X protocol compression. Linux Journal, 53:??, September 1998. CODEN LI-JOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Gallager:1978:VTH

[Gal78] Robert G. Gallager. Variations on a theme by Huffman. IEEE Transactions on Information Theory, IT-24(6):668–674, November 1978. CODEN IETTAW. ISSN 0018-9448 (print), 1557-9654 (electronic).

Ganssle:1994:DC

[Gan94] Jack G. Ganssle. Data compression. Web document, 1994.

URL http://www.avocetsystems.com/company/articles/
magazine/acompres.htm.

Gardner:1972:MG

[Gar72] Martin Gardner. Mathematical games. Scientific American, 227(2):106, August 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:

//www.gemstartvguide.com;   http://www.macrovision.
com/products/ce_manufacturers/ipg_ce/vcr_plus.htm.

Gaur:2016:BVC

[GAS16] Jayesh Gaur, Alaa R. Alameldeen, and Sreenivas Subramoney. Base-victim compression: an opportunistic cache compression architecture. ACM SIGARCH Computer Architecture News, 44(3):317–328, June 2016. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Gammerman:1992:EUM

[GB92] A. Gammerman and A. Bellotti. Experiments using minimallength encoding to solve machine learning problems. In Storer and Cohn [SC92], pages 359–367. ISBN 0-8186-2717-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=227445. IEEE catalog number 91TH0436-6.

Gutmann:1994:HAT

[GB94] P. C. Gutmann and T. C. Bell. A hybrid approach to text compression. In Storer and Cohn [SC94], pages 225–233. ISBN 0-8186-5636-0, 0-8186-5637-9 (case). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1994.

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=305930. IEEE Catalog Number 93TH0626-2. IEEE Computer Society Press Order Number 5637-02.

Gilbert:1998:LDI

[GB98] Jeffrey M. Gilbert and Robert W. Brodersen. A lossless 2-D image compression technique for synthetic discretetone images. In Storer and Cohn [SC98], pages 359– 368. ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D232 1998. URL http://bwrc.eecs.berkeley.
edu/Publications/1999/A_lossless_2-D_image_compression_
technique/JMG_DCC98.pdf;   http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=672166. IEEE Com-

puter Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Gillibert:2006:NLI

[GB06] L. Gillibert and A. Bretto. Near-lossless 3D-image compression using hypergraphs. In Storer and Cohn [SC06], page 450. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1607293.

Gillibert:2009:HGL

[GB09] Luc Gillibert and Alain Bretto. Hypergraphs for generic lossless image compression. Fundamenta Informaticae, 91(3–4): 533–546, August 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gauravaram:2016:BIC

[GBK16] Praveen Gauravaram, Nasour Bagheri, and Lars R. Knudsen. Building indifferentiable compression functions from the PGV compression functions. Designs, Codes, and Cryptography, 78(2):547–581, February 2016. CODEN DCCREC. ISSN 0925-1022 (print), 1573-7586 (electronic). URL http://link.

springer.com/article/10.1007/s10623-014-0020-z.

Gibson:1998:DCM

[GBL^{+}98] Jerry D. Gibson, Toby Berger, Tom Lookabaugh, David Lindbergh, and Richard L. Baker. Digital Compression for Multimedia: Principles and Standards. Morgan Kaufmann series in multimedia information and systems. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1998. ISBN 1-55860-369-7. xvii + 476 pp. LCCN QA76.575 .D535 1998. US$74.95, CAN$104.95, UK£49.95.

Guzun:2016:HQO

[GC16] Gheorghi Guzun and Guadalupe Canahuate. Hybrid query optimization for hard-to-compress bit-vectors. VLDB Journal: Very Large Data Bases, 25(3):339–354, June 2016. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Gao:1995:CLL

[GCBV95] Zheng Gao, Feng Chen, B. Belzer, and J. Villasenor. A comparison of the Z, E_{8} and Leech lattices for image subband quantization. In Storer and Cohn [SC95], pages 312– 321. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=515521. IEEE catalog

number 95TH8037. IEEE Computer Society Press order number PR07010.

Gentile:1996:RTI

[GCK^{+}96] A. Gentile, H. Cat, F. Kossentini, F. Sorbello, and D. S. Wills. Real-time implementation of full-search vector quantization on a low memory SIMD architecture. In Storer and Cohn [SC96], page ?? ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1996. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=488370. IEEE

Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Gergel:2006:UFL

[GCL06] B. Gergel, H. Cheng, and Xiaobo Li. A unified framework for lossless image set compression. In Storer and Cohn [SC06], page 448. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=1607291.

Guivarch:2000:JSC

[GCS00] L. Guivarch, J.-C. Carlach, and P. Siohan. Joint sourcechannel soft decoding of Huffman codes with turbo-codes. In Storer and Cohn [SC00], pages 83–92. ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2000. URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

838148. This millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Gandoin:2002:PLC

[GD02] Pierre-Marie Gandoin and Olivier Devillers. Progressive lossless compression of arbitrary simplicial complexes. ACM Transactions on Graphics, 21(3):372–379, July 2002. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Gehrig:2006:DSC

[GD06] N. Gehrig and P. L. Dragotti. Distributed sampling and compression of scenes with finite rate of innovation in camera sensor networks. In Storer and Cohn [SC06], pages 83–92. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1607243.

Gueguen:2007:MBS

[GD07] L. Gueguen and M. Datcu. The model based similarity metric. In Storer and Cohn [SC07], page 382. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=4148783. IEEE Computer Society Order Number P2791.

Gastpar:2003:DPC

[GDV03] M. Gastpar, P.-L. Dragotti, and M. Vetterli. The distributed, partial, and conditional Karhunen–Loeve transforms. In Storer and Cohn [SC03], pages 283–292. ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 2003. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1194019. IEEE Computer

Society Order number PR01896.

Gu:2005:SCM

[GE05] Wei-Hsin Gu and M. Effros. Source coding for a multihop network. In Storer and Cohn [SC05], page ?? ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1402218.

Gu:2006:MRC

[GE06] Wei-Hsin Gu and M. Effros. On multi-resolution coding and a two-hop network. In Storer and Cohn [SC06], page 451. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1607294.

Gunduz:2007:LTC

[GE07] D. Gunduz and E. Erkip. Lossless transmission of correlated sources over a multiple access channel with side information. In Storer and Cohn [SC07], pages 83–92. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4148747. IEEE Computer Society

Order Number P2791.

Gedik:2014:DFG

[Ged14] Bugra Gedik. Discriminative fine-grained mixing for adaptive compression of data streams. IEEE Transactions on Computers, 63(9):2228–2244, September 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Gerencser:1991:APS

[Ger91] L. Gerencser. Asymptotics of predictive stochastic complexity. In Storer and Reif [SR91b], pages 228–238. ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1990. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=213358. IEEE Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Garcia-Frias:2001:JSC

[GF01] J. Garcia-Frias. Joint source-channel decoding of correlated sources over noisy channels. In Storer and Cohn [SC01c], pages 283–292. ISBN 0-7695-1031-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2001. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=917159. IEEE CSP number 01PR1031.

Gog:2010:ASD

[GF10] S. Gog and J. Fischer. Advantages of shared data structures for sequences of balanced parentheses. In Storer and Marcellin [SM10b], pages 406–415. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=5453484.

Geelen:2009:SLE

[GFC^{+}09] Bert Geelen, Vissarion Ferentinos, Francky Catthoor, Gauthier Lafruit, Diederik Verkest, Rudy Lauwereins, and Thanos Stouraitis. Spatial locality exploitation for runtime reordering of JPEG2000 wavelet data layouts. ACM Transactions on Design Automation of Electronic Systems (TODAES), 15(1): 8:1–8:??, December 2009. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Garcia-Frias:2007:EPK

[GFE07] J. Garcia-Frias and I. Esnaola. Exploiting prior knowledge in the recovery of signals from noisy random projections. In Storer and Cohn [SC07], pages 333–342. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=4148772. IEEE Computer Society Order Number P2791.

Gong:1999:ECR

[GFH99] Y. Gong, M. K. H. Fan, and C.-M. Huang. On entropyconstrained residual vector quantization design. In Storer and Cohn [SC99], page ?? ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1999. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=785683. IEEE

Computer Society Order Number PR00096. IEEE Order Plan Catalog Number PR00096.

Gong:2000:ICU

[GFH00] Yun Gong, M. K. H. Fan, and Chien-Min Huang. Image compression using lossless coding on VQ indexes. In Storer and Cohn [SC00], page ?? ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2000. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=838230. This

millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Garcia-Frias:1997:ATC

[GFV97] J. Garcia-Frias and J. D. Villasenor. An analytical treatment of channel-induced distortion in run length coded subbands. In Storer and Cohn [SC97], pages 52–61. ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1997. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=581965. IEEE Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Garcia-Frias:1998:TDH

[GFV98] J. Garcia-Frias and J. D. Villasenor. Turbo decoding of hidden Markov sources with unknown parameters. In Storer and Cohn [SC98], pages 159–168. ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1998. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=672143. IEEE Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Gersho:1992:VQS

[GG92] Allen Gersho and Robert M. Gray. Vector Quantization and Signal Compression. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1992. ?? pp.

Guionnet:2004:JSC

[GG04] T. Guionnet and C. Guillemot. Joint source-channel decoding of quasiarithmetic codes. In Storer and Cohn [SC04], pages 272–281. ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=1281472. IEEE

Computer Society Order Number: P2082.

Gray:2006:QJE

[GG06] R. M. Gray and J. T. Gill III. Quantization with joint entropy/ memory constraints. In Storer and Cohn [SC06], pages 223– 235. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=1607257.

Grossberg:2009:CSL

[GGG^{+}09] M. Grossberg, I. Gladkova, S. Gottipati, M. Rabinowitz, P. Alabi, T. George, and A. Pacheco. A comparative study of lossless compression algorithms on multi-spectral imager data. In Storer and Marcellin [SM09], page 447. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=4976501. IEEE Computer Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

Grassi:2012:KSP

[GGM12] Elena Grassi, Federico Di Gregorio, and Ivan Molineris. KungFQ: a simple and powerful approach to compress fastq files. IEEE/ACM Transactions on Computational Biology and Bioinformatics, 9(6):1837–1842, November 2012. CODEN ITCBCY. ISSN 1545-5963 (print), 1557-9964 (electronic).

Gray:2008:RDF

[GH08] R. M. Gray and T. Hashimoto. Rate-distortion functions for nonstationary Gaussian autoregressive processes. In Storer and Marcellin [SM08], pages 53–62. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33 INTERNET.

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=
&arnumber=4483283.

Gharavi:1984:CVL

[Gha84] H. Gharavi. Conditional variable-length coding for gray-level pictures. AT&T Bell Laboratories Technical Journal, 63(2): 249–260, 1984. CODEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Gharavi:1987:CRL

[Gha87] H. Gharavi. Conditional run-length and variable-length coding of digital pictures. IEEE Transactions on Communications, COM-35(6):671–677, June 1987. CODEN IECMBT. ISSN 0090-6778 (print), 1558-0857 (electronic).

Guan:2014:EBG

[GHD^{+}14] Tao Guan, Yunfeng He, Liya Duan, Jianzhong Yang, Juan Gao, and Junqing Yu. Efficient BOF generation and compression for on-device mobile visual location recognition. IEEE MultiMedia, 21(2):32–41, April/June 2014. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Ghido:2003:AOP

[Ghi03] F. Ghido. An asymptotically optimal predictor for stereo lossless audio compression. In Storer and Cohn [SC03], page ?? ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2003. URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

1194048. IEEE Computer Society Order number PR01896.

Ghido:2004:EAL

[Ghi04] F. Ghido. An efficient algorithm for lossless compression of IEEE float audio. In Storer and Cohn [SC04], pages 429–438. ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1281488. IEEE Computer

Society Order Number: P2082.

Ghido:2005:QCA

[Ghi05] F. Ghido. QLFC — a compression algorithm using the Burrows–Wheeler transform. In Storer and Cohn [SC05], page ?? ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=1402216.

Ghido:2006:CPR

[Ghi06] F. Ghido. Combined prediction and residual coding for lossless audio compression. In Storer and Cohn [SC06], page 449. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1607292.

Ganardi:2018:TCU

[GHLN18] Moses Ganardi, Danny Hucke, Markus Lohrey, and Eric Noeth. Tree compression using string grammars. Algorithmica, 80(3):885–917, March 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Gupta:2006:CDS

[GHSV06] A. Gupta, Wing-Kai Hon, R. Shah, and J. S. Vitter. Compressed data structures: dictionaries and data-aware measures. In Storer and Cohn [SC06], pages 213–222. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1607256.

Glusker:2005:TCM

[GHV05] Mark Glusker, David M. Hogan, and Pamela Vass. The ternary calculating machine of Thomas Fowler. IEEE Annals of the History of Computing, 27(3):4–22, July/September 2005. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Gillman:1992:DHT

[Gil92] P. Gillman. Data handling and text compression. Journal of Information Science, Principles and Practice, 18(2):105–110, ???? 1992. CODEN JISCDI. ISSN 0165-5515 (print), 1741-6485 (electronic).

Gilchrist:2008:JGC

[Gil08] Jeff Gilchrist. Jeff Gilchrist’s compression archive comparison

test, 2008. URL http://compression.ca/act/.

Ginesta:1995:DHO

[Gin95] Xavier Ginesta. Design of high-order statistical models of an information source and their application to data compression. Ph.D. thesis, Polytechnic University, New York, NY, USA,

1995. 126 pp. URL http://search.proquest.com/docview/
304209689.

Girod:1999:BDS

[Gir99] Bernd Girod. Bidirectionally decodable streams of prefix codewords. IEEE Communications Letters, 3(8):245–247, August 1999. CODEN ICLEF6. ISSN 1089-7798 (print), 1558-2558 (electronic).

Givens:1958:CPU

[Giv58] Wallace Givens. Computation of plane unitary rotations transforming a general matrix to triangular form. Journal of the Society for Industrial and Applied Mathematics, 6(1):26–50, March 1958. CODEN JSIMAV. ISSN 0368-4245 (print), 1095-712X (electronic).

Girishwaingankar:2020:PNB

[GJ20] Poorva Girishwaingankar and Sangeeta Milind Joshi. The PHY-NGSC-based ORT run length encoding scheme for video compression. International Journal of Image and Graphics (IJIG), 20(02):??, April 2020. CODEN ???? ISSN

0219-4678. URL https://www.worldscientific.com/doi/
10.1142/S0219467820500072.

Goel:1988:DCA

[GK88a] B. Goel and S. Kwatra. A data compression algorithm for color images based on run-length coding and fractal geometry. In IEEE [IEE88a], pages 1253–1256. ISBN ???? LCCN TK5101.A1 I34 1988. IEEE catalog number 88CH2538-7.

Goel:1988:IIC

[GK88b] B. Goel and S. Kwatra. A data compression algorithm for color images based on run-length coding and fractal geometry. In IEEE [IEE88a], pages 1253–1256. ISBN ???? LCCN TK5101.A1 I34 1988. IEEE catalog number 88CH2538-7.

Ginesta:1994:VQC

[GK94] X. Ginesta and S. P. Kim. Vector quantization of contextual information for lossless image compression. In Storer and Cohn [SC94], pages 390–399. ISBN 0-8186-5636-0, 0-8186-5637-9 (case). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1994. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=305947. IEEE

Catalog Number 93TH0626-2. IEEE Computer Society Press Order Number 5637-02.

Ginesta:1995:DFF

[GK95] X. Ginesta and S. P. Kim. A derailment-free finite-state vector quantizer with optimized state codebooks. In Storer and Cohn [SC95], pages 152–161. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ????

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=515505. IEEE catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Goyal:1998:OMD

[GK98] V. K. Goyal and J. Kovacevic. Optimal multiple description transform coding of Gaussian vectors. In Storer and Cohn [SC98], pages 388–397. ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1998.

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=672173. IEEE Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Gallant:1999:RDO

[GK99] M. Gallant and F. Kossentini. Rate-distortion optimized spatial scalability for DCT-based video coding. In Storer and Cohn [SC99], page ?? ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1999. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=785682. IEEE

Computer Society Order Number PR00096. IEEE Order Plan Catalog Number PR00096.

Gagie:2009:LMA

[GKN09] T. Gagie, M. Karpinski, and Y. Nekrich. Low-memory adaptive prefix coding. In Storer and Marcellin [SM09], pages 13–22. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=4976445. IEEE

Computer Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

Gasieniec:1996:EAL

[GKPR96] L. Gasieniec, M. Karpinski, W. Plandowski, and W. Rytter. Efficient algorithms for Lempel–Ziv encoding. Lecture Notes in Computer Science, 1097:392–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gasieniec:2005:RTT

[GKPS05] L. Gasieniec, R. Kolpakov, I. Potapov, and P. Sant. Realtime traversal in grammar-based compressed files. In Storer and Cohn [SC05], page ?? ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=
1402215.

Gaj:2017:DCR

[GKSB17] Sibaji Gaj, Aditya Kanetkar, Arijit Sur, and Prabin Kumar Bora. Drift-compensated robust watermarking algorithm for H.265/HEVC video stream. ACM Transactions on Multimedia Computing, Communications, and Applications, 13(1):11:1– 11:??, January 2017. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Goyal:1999:QFE

[GKV99] V. K. Goyal, J. Kovacevic, and M. Vetterli. Quantized frame expansions as source-channel codes for erasure channels. In Storer and Cohn [SC99], pages 326–335. ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1999. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=755682. IEEE Computer Society Or-

der Number PR00096. IEEE Order Plan Catalog Number PR00096.

Gray:2001:ZEC

[GL01] R. M. Gray and J. Li. On Zador’s entropy-constrained quantization theorem. In Storer and Cohn [SC01c], pages 3–12. ISBN 0-7695-1031-0. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN QA76.9.D33 D37 2001. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=917131. IEEE

CSP number 01PR1031.

Golin:2003:RRT

[GL03a] M. J. Golin and Yiu-Cho Leung. Recurrence relations on transfer matrices yield good lower and upper bounds on the channel capacity of some 2-dimensional constrained systems. In Storer and Cohn [SC03], page ?? ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 2003. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1194049. IEEE Computer

Society Order number PR01896.

Gray:2003:HRM

[GL03b] R. M. Gray and T. Linder. High rate mismatch in entropy constrained quantization. In Storer and Cohn [SC03], pages 173–182. ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2003. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1194008. IEEE Computer Society Order number PR01896.

Gray:2004:RCH

[GL04] R. M. Gray and T. Linder. Results and conjectures on high rate quantization. In Storer and Cohn [SC04], pages 3–12. ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1281445. IEEE Computer

Society Order Number: P2082.

Gray:2009:BAO

[GL09] R. M. Gray and T. Linder. Bits in asymptotically optimal lossy source codes are asymptotically Bernoulli. In Storer and Marcellin [SM09], pages 272–281. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=4976471. IEEE Computer Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

Guerra:2016:PCS

[GLI16] Aníbal Guerra, Jaime Lotero, and Sebastián Isaza. Performance comparison of sequential and parallel compression applications for DNA raw data. The Journal of Supercomputing, 72(12):4696–4717, December 2016. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Gyorgy:2004:FPL

[GLL04] A. Gyorgy, T. Linder, and G. Lugosi. A “follow the perturbed leader”-type algorithm for zero-delay quantization of individual sequences. In Storer and Cohn [SC04], pages 342–351. ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1281479. IEEE Computer

Society Order Number: P2082.

Goshi:2003:ULP

[GLM^{+}03] J. Goshi, R. Ladner, A. Mohr, E. Riskin, and A. Lippman. Unequal loss protection for H.263 compressed video. In Storer and Cohn [SC03], pages 73–82. ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 2003. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1193998. IEEE Computer

Society Order number PR01896.

Grotschel:2010:PMP

[GLM10] Martin Grötschel, Klaus Lucas, and V. L. (Volker Ludwig) Mehrmann, editors. Produktionsfaktor Mathematik = Production factor mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 3-642-11247-1 (paperback). xiv + 402 pp. LCCN HB241 .P76513

2010. URL http://d-nb.info/1000034348/04;   http://
deposit.d-nb.de/cgi-bin/dokserv?id=3423837\%26prov=
M\%26dok\_var=1\%26dok\_ext=htm.

Guan:2021:CER

[GLM^{+}21] Hui Guan, Shaoshan Liu, Xiaolong Ma, Wei Niu, Bin Ren, Xipeng Shen, Yanzhi Wang, and Pu Zhao. CoCoPIE: enabling real-time AI on off-the-shelf mobile devices via compressioncompilation co-design. Communications of the ACM, 64(6):62– 68, June 2021. CODEN CACMA2. ISSN 0001-0782 (print),

1557-7317 (electronic). URL https://dl.acm.org/doi/10.
1145/3418297.

Gross:1999:CMV

[GLS99] Markus H. Gross, Lars Lippert, and Oliver G. Staadt. Compression methods for visualization. Future Generation Computer Systems, 15(1):11–29, February 12, 1999. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

URL http://visinfo.zib.de/EVlib/Show?EVL-1999-101;
 http://www.elsevier.com/gej-ng/10/19/19/30/18/18/
abstract.html.

Gilbert:1959:VLB

[GM59] E. N. Gilbert and E. F. Moore. Variable length binary encodings. The Bell System Technical Journal, 38(4):933–967, July 1959. CODEN BSTJAN. ISSN 0005-8580. URL http://bstj.

bell-labs.com/BSTJ/images/Vol38/bstj38-4-933.pdf;
http://www.alcatel-lucent.com/bstj/vol38-1959/articles/
bstj38-4-933.pdf. Monograph 3515.

Guan:1999:CAC

[GM99] Y. Guan and R. J. Moorhead. Compression of arbitrary cutting planes. In Storer and Cohn [SC99], page ?? ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1999. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=785685. IEEE Computer Society Or-

der Number PR00096. IEEE Order Plan Catalog Number PR00096.

Gobbetti:2006:GCD

[GMC^{+}06] E. Gobbetti, F. Marton, P. Cignoni, M. Di Benedetto, and F. Ganovelli. Geometry compression and decompression: C-BDAM — compressed batched dynamic adaptive meshes for terrain rendering. Computer Graphics Forum, 25(3):333–342, September 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Guo:2006:FLC

[GMNK06] Jiangling Guo, S. Mitra, B. Nutter, and T. Karp. A fast and low complexity image codec based on backward coding of wavelet trees. In Storer and Cohn [SC06], pages 292–301. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1607264.

Galvagno:1995:CGI

[GMR95] G. Galvagno, G. A. Mian, and R. Rinaldo. Coding gain of intra/inter-frame subband systems. In Storer and Cohn [SC95], page 449. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=515559. IEEE

catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Guthe:2009:BCP

[GMSK09] Michael Guthe, Gero Müller, Martin Schneider, and Reinhard Klein. BTF-CIELab: a perceptual difference measure for quality assessment and compression of BTFs. Computer Graphics Forum, 28(1):101–113, March 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gagie:2017:CSS

[GMV17] Travis Gagie, Giovanni Manzini, and Daniel Valenzuela. Compressed spaced suffix arrays. Mathematics in Computer Science, 11(2):151–157, June 2017. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Galil:1979:NFG

[GN79] Zvi Galil and Amnon Naamad. Network flow and generalized path compression. In ACM [ACM79], pages 13–26. LCCN QA75.5.A14 1979.

Golin:2001:OPF

[GN01] M. J. Golin and HyeonSuk Na. Optimal prefix-free codes that end in a specified pattern and similar problems: the uniform probability case. In Storer and Cohn [SC01c], pages 143–152. ISBN 0-7695-1031-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2001. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=917145. IEEE CSP number 01PR1031.

Golin:2005:GKM

[GN05] M. J. Golin and Hyeon-Suk Na. Generalizing the Kraft– McMillan inequality to restricted languages. In Storer and Cohn [SC05], pages 163–172. ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ????

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=
&arnumber=1402177.

Gonzalez:2015:LCS

[GNF15] Rodrigo González, Gonzalo Navarro, and Héctor Ferrada. Locally compressed suffix arrays. ACM Journal of Experimental Algorithmics, 19(??):1.1:1–1.1:??, February 2015. CODEN ???? ISSN 1084-6654.

Gamal:1984:IDC

[GO84] A. El Gamal and A. Orlitsky. Interactive data compression. In IEEE [IEE84], pages 100–108. CODEN ASFPDV. ISBN 0-8186-8591-3, 0-8186-0591-X (paperback), 0-8186-4591-1 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1984. IEEE catalog no. 84CH2085-9.

Gisladottir:1995:ORM

[GO95] J. V. Gisladottir and M. T. Orchard. Optimal representation of motion fields for video compression. In Storer and Cohn [SC95], pages 401–410. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ????

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=515530. IEEE catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Guleryuz:1996:RDB

[GO96] O. G. Guleryuz and M. T. Orchard. Rate-distortion based temporal filtering for video compression. In Storer and Cohn [SC96], pages 122–131. ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1996. URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

488317. IEEE Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Grossi:2015:FCT

[GO15] Roberto Grossi and Giuseppe Ottaviano. Fast compressed tries through path decompositions. ACM Journal of Experimental Algorithmics, 19(??):1.8:1–1.8:??, February 2015. CODEN ???? ISSN 1084-6654.

Golomb:1966:RLE

[Gol66] Solomon W. Golomb. Run-length encodings. IEEE Transactions on Information Theory, IT-12(3):399–401, 1966. CODEN IETTAW. ISSN 0018-9448 (print), 1557-9654 (electronic).

Golomb:1972:RLE

[Gol72] S. W. Golomb. Run-length encodings. IEEE Transactions on Information Theory, IT-12(3):399–401, 1972. CODEN IET-TAW. ISSN 0018-9448 (print), 1557-9654 (electronic).

Golin:1992:DIC

[Gol92] S. Golin. DVI image compression, second generation. Proceedings of the SPIE — The International Society for Optical Engineering, 1657:106–113, February 1992. CODEN PSISDG. ISSN 0277-786X (print), 1996-756X (electronic).

Goldbach:2006:LXG

[Gol06] Ch. Goldbach. Lettre XLIII: Goldbach à Euler: Moscau d.

7 Juni 1742, 2006. URL http://www.math.dartmouth.edu/~
euler/correspondence/letters/OO0765.pdf.

Goodman:1991:TPI

[Goo91] Stephen Darrow Goodman. Temporal pattern identification in a self-organizing neural network with an application to data compression. Ph.D. thesis, Georgia Institute of Technology, Atlanta, GA, USA, 1991. 209 pp. URL http://search.

proquest.com/docview/303954063.

Gooda:2003:ATD

[Goo03] Abdel-Hamid Mohamed Gooda. Application of the techniques of data compression and nearest neighbor classification to information retrieval. D.Sc. thesis, The George Washington University, Washington, DC, USA, 2003. ???? pp. URL http://

search.proquest.com/docview/305328703.

Gove:1994:MHI

[Gov94] R. J. Gove. The MVP: a highly-integrated video compression chip. In Storer and Cohn [SC94], pages 215–224. ISBN 0-8186-5636-0, 0-8186-5637-9 (case). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1994.

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=305929. IEEE Catalog Number 93TH0626-2. IEEE Computer Society Press Order Number 5637-02.

Guha:1995:PIC

[GP95] R. Guha and R. K. Pollock. Parallel image compression using vector quantization. In Storer and Cohn [SC95], page ?? ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ????  URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=515599. IEEE catalog

number 95TH8037. IEEE Computer Society Press order number PR07010.

Gibbons:2002:EAS

[GPM02] Alan Gibbons, Ida Pu, and Muthu Muthukrishnan. Exact analyses of a simple heuristic employed in array compression. The Computer Journal, 45(2):213–220, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_45/Issue_02/450213.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_
45/Issue_02/pdf/450213.pdf.

Gray:1998:QCD

[GPO98] R. M. Gray, K. O. Perlmutter, and R. A. Olshen. Quantization, classification, and density estimation for Kohonen’s Gaussian mixture. In Storer and Cohn [SC98], pages 63–72. ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1998. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=672132. IEEE Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Gaston:2011:QCM

[GPV11] B. Gastón, J. Pujol, and M. Villanueva. Quasi-cyclic minimum storage regenerating codes for distributed data compression. In Storer and Marcellin [SM11b], pages 33–42. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5749461. IEEE Computer Society

Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Gasieniec:1999:AOF

[GR99] L. Gasieniec and W. Rytter. Almost-optimal fully LZWcompressed pattern matching. In Storer and Cohn [SC99], pages 316–325. ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1999. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=755681. IEEE

Computer Society Order Number PR00096. IEEE Order Plan Catalog Number PR00096.

Gray:1953:PCC

[Gra53] Frank Gray. Pulse code communication, March 17, 1953. United States Patent 2,632,058.

Gray:1991:IC

[Gra91] R. M. Gray. Image compression. In Storer and Reif [SR91b], pages 474–475. ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1990. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=213293. IEEE

Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Graf:1997:EIC

[Gra97] U. Graf. Encoding of intervals with conditional coding. In Storer and Cohn [SC97], page ?? ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1997. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=582097. IEEE Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Graf:1999:SSW

[Gra99] U. Graf. Sorted sliding window compression. In Storer and Cohn [SC99], page ?? ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1999. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=785684. IEEE

Computer Society Order Number PR00096. IEEE Order Plan Catalog Number PR00096.

Graves:2003:RID

[Gra03] C. A. Graves. Recasting of the image deblocking problem into the domain of image denoising. In Storer and Cohn [SC03], page ?? ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2003. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1194050. IEEE Computer Society Order number PR01896.

Gray:2005:LFF

[Gra05] R. M. Gray. A Lagrangian formulation of fixed-rate quantization. In Storer and Cohn [SC05], pages 261–269. ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1402187.

Gladkova:2005:CAI

[GRG05] I. Gladkova, L. Roytman, and M. Goldberg. Compression algorithm for infrared hyperspectral sounder data. In Storer and Cohn [SC05], page ?? ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=1402217.

Gorjiara:2008:MDC

[GRG08] Bita Gorjiara, Mehrdad Reshadi, and Daniel Gajski. Merged dictionary code compression for FPGA implementation of custom microcoded PEs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 1(2):11:1–11:??, June 2008. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Grimm:1973:ALP

[Gri73] Richard E. Grimm. The autobiography of Leonardo Pisano. Fibonacci Quarterly, 11(1):99–104, February 1973. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/11-1/grimm.pdf.

GonzalezSmith:1985:PAD

[GS85] M. E. Gonzalez Smith and J. A. Storer. Parallel algorithms for data compression. Journal of the ACM, 32(2):344–373, April 1985. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.org/pubs/

toc/Abstracts/0004-5411/3152.html.

Ginesta:1996:LCF

[GS96] X. Ginesta and S. Sambhwani. Low complexity finite-state scalar quantization of image subbands. In Storer and Cohn [SC96], page ?? ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1996. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=488371. IEEE

Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Gupta:2005:SWC

[GS05] M. R. Gupta and A. Stroilov. Segmenting for wavelet compression. In Storer and Cohn [SC05], page ?? ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1402219.

Georgiadis:2012:SAV

[GS12] G. Georgiadis and S. Soatto. Scene-aware video modeling and compression. In Storer and Marcellin [SM12], pages 149–158. ISBN 1-4673-0715-7 (paperback). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

6189246. IEEE Computer Society order number P4656.

Goncalves:2019:NRR

[GSPR19] Mateus A. Gonçalves, Lizandro S. Santos, Fernando C. Peixoto, and Teodorico C. Ramalho. NMR relaxation and relaxivity parameters of MRI probes revealed by optimal wavelet signal compression of molecular dynamics simulations. International Journal of Quantum Chemistry, 119(10):e25896:1– e25896:??, May 15, 2019. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Galli:2001:THO

[GSS01] N. Galli, B. Seybold, and K. Simon. Tetris-hashing or optimal table compression. Discrete Applied Mathematics, 110(1):41– 58, June 1, 2001. CODEN DAMADU. ISSN 0166-218X (print), 1872-6771 (electronic).

Gibson:2002:DLW

[GSW02] D. Gibson, M. Spann, and S. I. Woolley. Diagnostically lossless 3D wavelet compression for digital angiogram video. In Storer and Cohn [SC02], page ?? ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2002. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=999997. IEEE Computer Society Order Number PR01477.

Grumbach:1993:CDS

[GT93] S. Grumbach and F. Tahi. Compression of DNA sequences. In Storer and Cohn [SC93], pages 340–350. ISBN 0-8186-3391-3 (microfiche), 0-8186-3392-1 (casebound). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1993. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=253115. IEEE Computer Society Press order number 3392-02. IEEE catalog number 93TH0536-3.

Gu:2005:LAO

[Gu05] Hung-Yan Gu. A large-alphabet-oriented scheme for Chinese and English text compression. Software — Practice and Experience, 35(11):1027–1039, September 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Guleryuz:2002:IDI

[Gul02] O. G. Guleryuz. Iterated denoising for image recovery. In Storer and Cohn [SC02], pages 3–12. ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2002. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=999938. IEEE Computer Society Order Number PR01477.

Guleryuz:2004:PWC

[Gul04] O. G. Guleryuz. Predicting wavelet coefficients over edges using estimates based on nonlinear approximants. In Storer and Cohn [SC04], pages 162–171. ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ????

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=1281461. IEEE Computer Society Order Number: P2082.

Gunther:1996:DCS

[Gün96] U. Günther. Data compression and serial communication with generalized T-codes. J.UCS: The Journal of Universal Computer Science, 2(11):769–795, November 28, 1996. CODEN ???? ISSN 0948-695X (print), 0948-6968 (elec-

tronic). URL http://www.jucs.org/data_compression_
and_serial_communiction.

Grbovic:2009:DEU

[GV09] M. Grbovic and S. Vucetic. Decentralized estimation using learning vector quantization. In Storer and Marcellin [SM09], page 446. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=4976500. IEEE

Computer Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

Garcia-Vilchez:2008:HIC

[GVSSBRAL08] F. Garcia-Vilchez, J. Serra-Sagristà, J. Bartrina-Rapesta, and F. Aulí-Llinás. Hyperspectral image coding using 3D transform and the recommendation CCSDS-122-B-1. In Storer and Marcellin [SM08], pages 103–112. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33 INTERNET. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=4483288.

Goyal:1995:QOE

[GVT95] V. K. Goyal, M. Vetterli, and N. T. Thao. Quantization of overcomplete expansions. In Storer and Cohn [SC95], pages 13–22. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=515491. IEEE catalog

number 95TH8037. IEEE Computer Society Press order number PR07010.

Gallager:1975:OSC

[GvV75] Robert G. Gallager and David C. van Voorhis. Optimal source codes for geometrically distributed integer alphabets. IEEE Transactions on Information Theory, IT-21(3):228–230, March 1975. CODEN IETTAW. ISSN 0018-9448 (print), 1557-9654 (electronic).

Greene:1997:PZL

[GVYZ97] D. Greene, M. Vishwanath, F. Yao, and Tong Zhang. Progressive Ziv–Lempel encoding of synthetic images. In Storer and Cohn [SC97], page ?? ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1997. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=582099. IEEE

Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Gibbons:1973:CFC

[GW73] Charles J. Gibbons and M. Wells. Correspondence: On “File Compression Using Variable Length Encodings”. The Computer Journal, 16(3):287, August 1973. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://www3.oup.co.uk/computer_journal/hdb/Volume_
16/Issue_03/tiff/287.tif. See [Wel72].

Gonzalez:1992:DIP

[GW92] Rafael C. Gonzalez and Richard E. Woods. Digital Image Processing. Addison-Wesley, Reading, MA, USA, 1992. ?? pp.

Gemelos:2005:ERP

[GW05] G. M. Gemelos and T. Weissman. On the entropy rate of pattern processes. In Storer and Cohn [SC05], pages 233– 242. ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=1402184.

Gera:2012:FCF

[GWSR12] Y. Gera, Zhe Wang, S. Simon, and T. Richter. Fast and context-free lossless image compression algorithm based on JPEG-LS. In Storer and Marcellin [SM12], page 396. ISBN 1-4673-0715-7 (paperback). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=6189277. IEEE Com-

puter Society order number P4656.

Gasthaus:2010:LCB

[GWT10] J. Gasthaus, F. Wood, and Yee Whye Teh. Lossless compression based on the sequence memoizer. In Storer and Marcellin [SM10b], pages 337–345. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=5453479.

Gu:2009:HBJ

[GYLC09] Xiaozhuo Gu, Jianzu Yang, Julong Lan, and Zhenhuan Cao. Huffman-based join-exit-tree scheme for contributory key management. Computers & Security, 28(1–2):29–39, February/ March 2009. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0167404808000680.

Guenter:1993:MCC

[GYM93] Brian K. Guenter, Hee Cheol Yun, and Russell M. Mersereau. Motion compensated compression of computer animation frames. Computer Graphics, 27(Annual Conference Series): 297–304, 1993. CODEN CGRADI, CPGPBZ. ISSN 0097-8930 (print), 1558-4569 (electronic). URL http://www.

acm.org:80/pubs/citations/proceedings/graph/166117/
p297-guenter/.

Gertner:1991:GSM

[GZ91] I. Gertner and Y. Y. Zeevi. Generalized scanning and multiresolution image compression. In Storer and Reif [SR91b], page ?? ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1990. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=213327. IEEE

Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Gormish:1996:DSQ

[GZS^{+}96] M. J. Gormish, A. Zandi, E. L. Schwartz, A. Keith, and M. Boliek. Device selective quantization for reversible wavelets. In Storer and Cohn [SC96], page ?? ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1996. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=488372. IEEE Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Goyal:1997:UTC

[GZV97] V. K. Goyal, J. Zhuang, and M. Vetterli. Universal transform coding based on backward adaptation. In Storer and Cohn [SC97], pages 231–240. ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1997.

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=582046. IEEE Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Haffner:1998:HQD

[H^{+}98] Patrick Haffner et al. High-quality document image compression with DjVu. Journal of Electronic Imaging, 7(3): 410–425, 1998. CODEN JEIME5. ISSN 1017-9909 (print), 1560-229X (electronic). URL http://citeseer.ist.psu.

edu/old/523172.html.

Huang:2010:SDC

[H^{+}10] Bormin Huang et al., editors. Satellite data compression, communications, and processing VI: 3–5 August 2010, San Diego, California, United States, volume 7810. SPIE Optical Engineering Press, Bellingham, WA, USA, 2010. ISBN 0-8194-8306-0. LCCN TK5102.92 .S3744 2010.

Haar:1910:TOF

[Haa10] A. Haar. Zur Theorie der Orthogonalen Funktionensysteme. (German) [On the theory of orthogonal function systems]. Mathematische Annalen, 69:331–371, 1910. CODEN MAANA3. ISSN 0025-5831 (print), 1432-1807 (electronic).

Haar:1912:TOF

[Haa12] A. Haar. Zur Theorie der Orthogonalen Funktionensysteme. (German) [On the theory of orthogonal function systems]. Mathematische Annalen, 71:38–53, 1912. CODEN MAANA3. ISSN 0025-5831 (print), 1432-1807 (electronic).

Haeberli:1996:GOC

[Hae96] Paul Haeberli. Grafica Obscura: Collected computer graphics

hacks, 1996. URL http://www.sgi.com/grafica/huffman/.

Hafner:1995:ACW

[Haf95] Ullrich Hafner. Asymmetric coding in (m)-WFA image compression. Report 132, Department of Computer Science, University of Würzburg, Würzburg, Germany, December 1995. ?? pp.

Hafner:1996:RIC

[Haf96] U. Hafner. Refining image compression with weighted finite automata. In Storer and Cohn [SC96], pages 359–368. ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1996. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=488341. IEEE Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Hafner:2001:FOS

[Haf01] Ullrich Hafner. FIASCO — an open-source fractal image and sequence codec. Linux Journal, 81:152–155, January 2001. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Haghighat:2007:DCU

[Hag07] Javad Haghighat. Data compression using error correcting codes. Ph.D. thesis, Department of Electrical and Computer Engineering, Concordia University, Montréal, QC, Canada, March 2007. xiii + 86 pp. URL http:/

/www.collectionscanada.gc.ca/obj/thesescanada/vol2/
002/NR30135.PDF. 2 microfiches.

Hahn:1974:NTC

[Hah74] Bruce Hahn. A new technique for compression and storage of data. Communications of the ACM, 17(8):434–436, August 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Huang:2005:FPV

[HAH05] Bormin Huang, A. Ahuja, and Hung-Lung Huang. Fast precomputed VQ with optimal bit allocation for lossless compression of ultraspectral sounder data. In Storer and Cohn [SC05], pages 408–417. ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=
1402202.

Halfhill:1994:HSD

[Hal94] Tom Halfhill. How safe is data compression: On the fly data compression can double your storage space — but not without some hassles and headaches. Byte Magazine, 19(2):56–??, February 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Halbach:2004:UEP

[Hal04] T. Halbach. Unequal error protection of SNR-scalable DPCMcoded video. In Storer and Cohn [SC04], page ?? ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1281517. IEEE Computer Society

Order Number: P2082.

Haley:2011:PCC

[Hal11] Brent Kreh Haley. A Pipeline for the Creation, Compression, and Display of Streamable 3D Motion Capture Based Skeletal Animation Data. Ph.D. thesis, The Ohio State University, Columbus, OH, USA, 2011. 158 pp. URL http://search.

proquest.com/docview/863205140.

Hammerstrom:1993:NNW

[Ham93] D. Hammerstrom. Neural networks at work. IEEE Spectrum, 30(6):26–32, June 1993. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hamidi:1994:WTI

[Ham94] Seyed Mostafa Hamidi. Wavelet transform for image data compression. Ph.D. thesis, The University of Alabama in Huntsville, Huntsville, AL, USA, 1994. 150 pp. URL http:/

/search.proquest.com/docview/304121118.

Hamalainen:1995:VRG

[Ham95a] J. Hamalainen. Video recording goes digital. IEEE Spectrum, 32(4):76–79, April 1995. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hamzaoui:1995:CCS

[Ham95b] Raouf Hamzaoui. Codebook clustering by self-organizing maps for fractal image compression. In Fisher [Fis95b], pages xiii + 368. ISBN 3-540-63196-8. LCCN TA1637

.N38 1998. URL ftp://ftp.informatik.uni-freiburg.de/
papers/fractal/Hamz95.ps.gz.

Ham:2017:DDS

[HAM17] Tae Jun Ham, Juan L. Aragón, and Margaret Martonosi. Decoupling data supply from computation for latency-tolerant communication in heterogeneous architectures. ACM Transactions on Architecture and Code Optimization, 14(2):16:1–16:??, July 2017. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Hart:1996:FIC

[Har96] John C. Hart. Fractal image compression and recurrent iterated function systems. IEEE Computer Graphics and Applications, 16(4):25–33, July 1996. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Haralambous:2004:FCG

[Har04] Yannis Haralambous. Fontes & codages: Glyphes et caractères à l’ère du numérique. (French) [Fonts and Encodings: Glyphs and characters in the numerical era]. O’Reilly, Paris, France, 2004. ISBN 2-84177-273-X. xx + 990 pp. LCCN ???? URL

http://www.oreilly.com/catalog/9782841772735.

Haralambous:2007:FE

[Har07] Yannis Haralambous. Fonts and Encodings. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-10242-9.

xx + 1016 pp. LCCN Z250. URL http://www.oreilly.com/

catalog/9780596102425. Translation of [Har04] from French by P. Scott Horne.

Harrusi:2010:FCP

[HAR10] S. Harrusi, A. Averbuch, and N. Rabin. A fast compact prefix encoding for pattern matching in limited resources devices. In Storer and Marcellin [SM10b], page 533. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5453510.

Hassanieh:2018:SFT

[Has18] Haitham Hassanieh. The Sparse Fourier Transform: Theory and Practice, volume 19 of ACM book series. ACM Press, New York, NY 10036, USA, 2018. ISBN 1-947487-07-8 (hardcover), 1-947487-04-3 (paperback), 1-947487-05-1 (e-book). xvii + 260 pp. LCCN QC20.7.F67 H37 2018.

Hatton:1995:ODC

[Hat95] Edward Douglas Hatton. Optimized data compression in the absence of a priori probabilities. Ph.D. thesis, University of Waterloo, Waterloo, ON, Canada, 1995. 141 pp. URL http:

//search.proquest.com/docview/304299385.

Haukeli:2012:LDC

[Hau12a] Martin Haukeli. Lossless data compression using cellular automata. Masteroppgave i informatikk, Institutt for informatikk, Universitetet i Oslo, Oslo, Norway, 2012. 97 pp.

Hautamaki:2012:RBC

[Hau12b] Ville Hautamäki. Review of A Concise Introduction to Data Compression by David Salomon. ACM SIGACT News, 43(2):

9–10, June 2012. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hawley:1990:POAa

[Haw90a] Michael Hawley. The personal orchestra, or audio data compression by 10,000:1. In USENIX [USE90], pages 289–330.

Hawley:1990:POAb

[Haw90b] Michael Hawley. The personal orchestra, or audio data compression by 10,000:1. Computing systems: the journal of the USENIX Association, 3(2):289–329, Spring 1990. CODEN CMSYE2. ISSN 0895-6340.

Harrusi:2006:XSC

[HAY06] S. Harrusi, A. Averbuch, and A. Yehudai. XML syntax conscious compression. In Storer and Cohn [SC06], pages 10 pp.– 411. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=1607275.

Haley:1996:IVR

[HB96a] Michael B. Haley and Edwin H. Blake. Incremental volume rendering using hierarchical compression. Computer Graphics Forum, 15(3):C45–C55, September 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Huber:1996:GLD

[HB96b] A. K. Huber and S. E. Budge. Global and local distortion inference during embedded zerotree wavelet decompression. In Storer and Cohn [SC96], page ?? ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1996. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=488373. IEEE Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Hong:1999:FAR

[HB99] S.-W. Hong and P. Bao. Finite automata and regularized edge-preserving wavelet transform scheme. In Storer and Cohn [SC99], page ?? ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN QA76.9.D33 D37 1999. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=785687. IEEE

Computer Society Order Number PR00096. IEEE Order Plan Catalog Number PR00096.

Hirschberg:2008:ECM

[HB08] D. S. Hirschberg and P. Baldi. Effective compression of monotone and quasi-monotone sequences of integers. In Storer and Marcellin [SM08], page 520. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33 INTERNET. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=4483347.

Hatam:2009:ARA

[HB09] A. Hatam and A. H. Banihashemi. Adaptive rate allocation algorithm for transmission of multiple embedded bit streams over time-varying noisy channels. In Storer and Marcellin [SM09], page 448. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ????

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=4976502. IEEE Computer Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

Holliday:2017:SDL

[HBL^{+}17] Andrew Holliday, Mohammadamin Barekatain, Johannes Laurmaa, Chetak Kandaswamy, and Helmut Prendinger. Speedup of deep learning ensembles for semantic segmentation using a model compression technique. Computer Vision and Image Understanding: CVIU, 164(??):16–26, November 2017. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1077314217300826.

Haroune-Belkacem:2019:SES

[HBSASA19] Nassima Haroune-Belkacem, Fouzi Semchedine, Ahmed Al-Shammari, and Djamil Aissani. SMCA: an efficient SOAP messages compression and aggregation technique for improving web services performance. Journal of Parallel and Distributed Computing, 133(??):149–158, November 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519304940.

Horspool:1992:CWB

[HC92] R. Nigel Horspool and Gordon V. Cormack. Constructing word-based text compression algorithms. In Storer and Cohn [SC92], pages 62–71. ISBN 0-8186-2717-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

227475. IEEE catalog number 91TH0436-6.

Helfgott:1997:MPS

[HC97] H. Helfgott and M. Cohn. On maximal parsings of strings. In Storer and Cohn [SC97], pages 291–299. ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1997. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=582052. IEEE Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Helfgott:1998:LTC

[HC98] H. Helfgott and M. Cohn. Linear-time construction of optimal context trees. In Storer and Cohn [SC98], pages 369–377. ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1998. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=672167. IEEE Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Hu:2006:ITS

[HC06] Yu-Chen Hu and Chin-Chen Chang. An improved treestructured codebook search algorithm for grayscale image compression. Fundamenta Informaticae, 70(3):251–260, May 2006. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hernandez-Cabronero:2012:SSC

[HCBMSS12] Miguel Hernández-Cabronero, Ian Blanes, Michael W. Marcellin, and Joan Serra-Sagristà. Standard and specific compression techniques for DNA microarray images. Algorithms (Basel), 5(1):30–49, March 2012. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-
4893/5/1/30.

Huo:2004:JCL

[HCD04] X. Huo, J. Chen, and D. L. Donoho. JBEAM: coding lines and curves via digital beamlets. In Storer and Cohn [SC04], pages 449–458. ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=1281490. IEEE

Computer Society Order Number: P2082.

Hemami:1999:RVL

[HCL99] S. S. Hemami, T. Chang, and R. Lau. Resynchronizing variable-length codes for robust image transmission. In Storer and Cohn [SC99], page ?? ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1999. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=785686. IEEE

Computer Society Order Number PR00096. IEEE Order Plan Catalog Number PR00096.

Hernandez-Cabronero:2012:DMI

[HCMGB^{+}12] M. Hernandez-Cabronero, J. Munoz-Gomez, I. Blanes, M. W. Marcellin, and J. Serra-Sagristà. DNA microarray image coding. In Storer and Marcellin [SM12], pages 32–41. ISBN 1-4673-0715-7 (paperback). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=6189234. IEEE Com-

puter Society order number P4656.

Hurst:1988:IDC

[HD88] G. J. B. Hurst and M. Dupuis. Integral data compression for FPS 64-bit processors: Improved I/O and reduced storage. Journal of Computational Chemistry, 9(2):148–157, March 1988. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Hanus:2009:SCS

[HDCH09] P. Hanus, J. Dingel, G. Chalkidis, and J. Hagenauer. Source coding scheme for multiple sequence alignments. In Storer and Marcellin [SM09], pages 183–192. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=4976462. IEEE Computer Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

Hijazi:2012:MAD

[HDK^{+}12] H. Hijazi, A. Diallo, M. Kieffer, L. Liberti, and C. Weidmann. A MILP approach for designing robust variable-length codes based on exact free distance computation. In Storer and Marcellin [SM12], pages 257–266. ISBN 1-4673-0715-7 (paperback). ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=6189257. IEEE Computer Society

order number P4656.

Heath:1972:OBC

[Hea72] F. G. Heath. Origins of the binary code. Scientific American, 227(2):76, August 1972. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Heshmati:2008:DCT

[Hes08] Ashkan Heshmati. Data compression and transmission in wireless sensor networks. Master of Applied Science (Electrical Engineering), Department of Electrical and Computer Engineering, Concordia University, Montréal, QC, Canada, 2008.

xii + 84 pp. URL http://www.collectionscanada.gc.ca/
obj/thesescanada/vol2/002/MR28915.PDF. 1 microfiche.

Haggag:2011:MEF

[HESF11] M. N. Haggag, M. El-Sharkawy, and G. Fahmy. Modified efficient fast multiplication-free integer transformation for the 2-D DCT H.265 standard. In Storer and Marcellin [SM11b], page 455. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=5749512. IEEE

Computer Society Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Hlavac:1997:ANL

[HF97] V. Hlavac and J. Fojtik. Adaptive non-linear predictor for lossless image compression. Lecture Notes in Computer Science, 1296:279–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hlavac:1999:APL

[HF99] V. Hlaváč and J. Fojtík. Adaptive predictor for lossless image compression. Computing: Archiv für informatik und numerik, 62(4):339–354, December 1999. CODEN CMPTA2.

ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer-ny.com/link/service/journals/00607/
bibs/9062004/90620339.htm;   http://link.springer-
ny.com/link/service/journals/00607/papers/9062004/
90620339.pdf.

Hua:2007:SSI

[HG07] Gang Hua and O. G. Guleryuz. Spatial sparsity induced temporal prediction for hybrid video compression. In Storer and Cohn [SC07], pages 23–32. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=4148741. IEEE Computer Society Order Number P2791.

Hu:2009:AJS

[HGFL09] Yichuan Hu, J. Garcia-Frias, and M. Lamarca. Analog joint source channel coding using space-filling curves and MMSE decoding. In Storer and Marcellin [SM09], pages 103–112. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=4976454. IEEE Computer

Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

Huo:2005:BAQ

[HGH05] Longshe Huo, Wen Gao, and Qingming Huang. Bandwidth adaptive quality smoothing for unequal error protected scalable video streaming. In Storer and Cohn [SC05], page ?? ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1402220.

Hou:2011:CIC

[HGR11] Xin Hou, K. S. Gurumoorthy, and A. Rajwade. Color image compression using a learned dictionary of pairs of orthonormal bases. In Storer and Marcellin [SM11b], page 458. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5749515. IEEE Computer Society

Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Hamzaoui:1997:QBV

[HGS97] R. Hamzaoui, B. Ganz, and D. Saupe. Quadtree based variable rate oriented mean shape-gain vector quantization. In Storer and Cohn [SC97], pages 327–336. ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1997. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=582056. IEEE Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Hankerson:1999:TMI

[HH99] Darrel Hankerson and Greg A. Harris. Transform methods and image compression. Linux Journal, 57:??, January 1999. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Huang:2008:TBJ

[HH08] Yuh-Ming Huang and Y. S. Han. Trellis-based joint Huffman and convolutional soft-decision priority-first decoding. In Storer and Marcellin [SM08], page 521. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33 INTERNET.

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=
&arnumber=4483348.

Hsu:2011:DET

[HHCH11] Chun-Lung Hsu, Yu-Sheng Huang, Ming-Da Chang, and Hung-Yen Huang. Design of an error-tolerance scheme for discrete wavelet transform in JPEG 2000 encoder. IEEE Transactions on Computers, 60(5):628–638, May 2011. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Heymann:2007:SDC

[HHG07] Roger W. Heymann, Bormin Huang, and Irina Gladkova, editors. Satellite data compression, communications, and archiving III: 29–30 August 2007, San Diego, California, USA, volume 6683. SPIE Optical Engineering Press, Bellingham, WA, USA, 2007. ISBN 0-8194-6831-2. LCCN TK5102.92 .S372

Hsieh:2020:WBQ

[HHLW20] Jui-Hung Hsieh, King-Chu Hung, Je-Hung Liu, and Tsung-Ching Wu. Wavelet-based quality-constrained ECG data compression system without decoding process. IEEE MultiMedia, 27(2):33–45, 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Heymann:2008:SDC

[HHSS08] Roger W. Heymann, Bormin Huang, and Joan Serra-Sagristà, editors. Satellite data compression, communications, and processing IV: 10–11 August 2008, San Diego, California, USA, volume 7084. SPIE Optical Engineering Press, Bellingham, WA, USA, 2008. ISBN 0-8194-7304-9 (paperback). LCCN TK5102.92 .S3739 2008.

Huang:2005:SDC

[HHW05] Bormin Huang, Roger W. Heymann, and Charles C. Wang, editors. Satellite data compression, communications, and archiving: 31 July–1 August, 2005, San Diego, California, USA, volume 5889 of Proceedings of SPIE. SPIE Optical Engineering Press, Bellingham, WA, USA, 2005. ISBN 0-8194-5894-5 (paperback). ISSN 0277-786X (print), 1996-756X (electronic). ???? pp. LCCN TK5102.92 .S37 2005.

Herrero:2012:NPS

[HI12] R. Herrero and V. Ingle. A new preprocessing stage for compression of ultraspectral images. In Storer and Marcellin [SM12], page 397. ISBN 1-4673-0715-7 (paperback). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ????

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=6189278. IEEE Computer Society order number P4656.

Hilbert:1891:SAL

[Hil91] D. Hilbert. Über stetige Abbildung einer Linie auf ein Flächenstück. (German) [On a continuous map of a line on a surface element]. Mathematische Annalen, 38:459–460, 1891. CODEN MAANA3. ISSN 0025-5831 (print), 1432-1807 (electronic).

Hilbert:1975:JCS

[Hil75] Edward Erwin Hilbert. A Joint Clustering /Data Compression Concept. Ph.D. thesis, University of Southern California, Los Angeles, CA, USA, 1975. ???? pp. URL http://search.

proquest.com/docview/302793284.

Hameed:2020:LCE

[HIMM20] Mustafa Emad Hameed, Masrullizam Mat Ibrahim, Nurulfajar Abd Manap, and Ali A. Mohammed. A lossless compression and encryption mechanism for remote monitoring of ECG data using Huffman coding and CBC–AES. Future Generation Computer Systems, 111(??):829–840, October 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-

7115 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167739X19313950.

Hayden:2002:TQC

[HJW02] Patrick Hayden, Richard Jozsa, and Andreas Winter. Trading quantum for classical resources in quantum data compression. Journal of Mathematical Physics, 43(9):4404–4444, September 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hoffman:2005:GHS

[HJW05] Dean Hoffman, Peter Johnson, and Nadine Wilson. Generating Huffman sequences. Journal of Algorithms, 54(1):115–121, January 2005. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0196677404000653.

Hassan:2006:EVB

[HKA06] M. Hassan, M. Krunz, and S. Ahuja. Efficient video broadcast over wireless channels using adaptive playback. In Storer and Cohn [SC06], page 454. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=1607297.

Hovestadt:2013:AOC

[HKKW13] Matthias Hovestadt, Odej Kao, Andreas Kliem, and Daniel Warneke. Adaptive online compression in clouds — making informed decisions in virtual machine environments. Journal of Grid Computing, 11(2):167–186, June 2013. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:/

/link.springer.com/article/10.1007/s10723-013-9249-

Harada:2010:LCM

[HKM10] N. Harada, Y. Kamamoto, and T. Moriya. Lossless compression of mapped domain linear prediction residual for ITU-T recommendation G.711.0. In Storer and Marcellin [SM10b], page 532. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=5453511.

Herzog:2008:PRA

[HKMS08] Robert Herzog, Shinichi Kinuwaki, Karol Myszkowski, and Hans-Peter Seidel. Perception for rendering and animation: Render2MPEG: a perception-based framework towards integrating rendering and video compression. Computer Graphics Forum, 27(2):183–192, April 2008. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hon:2011:CDM

[HKS^{+}11] Wing-Kai Hon, Tsung-Han Ku, R. Shah, S. V. Thankachan, and J. S. Vitter. Compressed dictionary matching with one error. In Storer and Marcellin [SM11b], pages 113–122. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5749469. IEEE Computer Society

Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Hon:2013:FCD

[HKS^{+}13] Wing-Kai Hon, Tsung-Han Ku, Rahul Shah, Sharma V. Thankachan, and Jeffrey Scott Vitter. Faster compressed dictionary matching. Theoretical Computer Science, 475 (??):113–119, March 4, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304397512009826.

Huang:2010:CTC

[HKW10] Ying-Zong Huang, Y. Kochman, and G. W. Wornell. Causal transmission of colored source frames over a packet erasure channel. In Storer and Marcellin [SM10b], pages 129–138. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=5453457.

Hirschberg:1990:EDP

[HL90] Daniel S. Hirschberg and Debra A. Lelewer. Efficient decoding of prefix codes. Communications of the ACM, 33(4): 449–459, April 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.org/

pubs/toc/Abstracts/0001-0782/77566.html. This is the

first of four key papers behind the bzip2 compression tools. The others are [BW94b, Whe97, SB97a].

Han:1997:EDE

[HL97] Wei-Yu Han and Ja-Chen Lin. Edge detection and edgepreserved compression for error- diffused images. Computers and Graphics, 21(6):757–767, November–December 1997. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684

(electronic). URL http://www.elsevier.com/cgi-bin/cas/
tree/store/cag/cas_sub/browse/browse.cgi?year=1997&
volume=21&issue=6&aid=9700055.

Hong:2000:GTI

[HL00] E. S. Hong and R. E. Ladner. Group testing for image compression. In Storer and Cohn [SC00], pages 3–12. ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2000. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=838140. This millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Hong:2002:EGC

[HL02] E. S. Hong and R. E. Ladner. Extended Golomb codes for binary Markov sources. In Storer and Cohn [SC02], page ?? ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2002. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=999999. IEEE

Computer Society Order Number PR01477.

Hashempour:2005:AAC

[HL05] H. Hashempour and F. Lombardi. Application of arithmetic coding to compression of VLSI test data. IEEE Transactions on Computers, 54(9):1166–1177, September 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=1471676.

He:2017:ISA

[HL17] Bin He and Gang Li. Intelligent self-adaptation data behavior control inspired by speech acts. ACM Transactions on Sensor Networks, 13(2):13:1–13:??, June 2017. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Hu:2007:BPV

[HLC07] Yu-Chen Hu, Chia-Chen Lin, and Kang-Liang Chi. Block prediction vector quantization for grayscale image compression. Fundamenta Informaticae, 78(2):257–270, October 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Han:2020:ADL

[HLL^{+}20] R. Han, C. H. Liu, S. Li, S. Wen, and X. Liu. Accelerating deep learning systems via critical set identification and model compression. IEEE Transactions on Computers, 69(7):1059– 1070, 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hudson:2017:JSG

[HLNS17] Graham Hudson, Alain Leger, Birger Niss, and Istvan Sebestyen. JPEG at 25: Still going strong. IEEE MultiMedia, 24(2):96–103, April/June 2017. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

URL https://www.computer.org/csdl/mags/mu/2017/02/
mmu2017020096-abs.html.

Hong:2001:GTW

[HLR01] E. S. Hong, R. E. Ladner, and E. A. Riskin. Group testing for wavelet packet image compression. In Storer and Cohn [SC01c], pages 73–82. ISBN 0-7695-1031-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2001. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=917138. IEEE CSP number 01PR1031.

Hon:2004:CID

[HLS^{+}04] Wing-Kai Hon, Tak-Wah Lam, Kunihiko Sadakane, Wing king Sung, and Siu-Ming Yiu. Compressed index for dynamic text. In Storer and Cohn [SC04], pages 102–111. ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1281455. IEEE Computer Society

Order Number: P2082.

Hoang:1994:EBM

[HLV94] D. T. Hoang, P. M. Long, and J. S. Vitter. Explicit bit minimization for motion-compensated video coding. In Storer and Cohn [SC94], pages 175–184. ISBN 0-8186-5636-0, 0-8186-5637-9 (case). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1994. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=305925. IEEE

Catalog Number 93TH0626-2. IEEE Computer Society Press Order Number 5637-02.

Hoang:1995:MDC

[HLV95] D. T. Hoang, P. M. Long, and J. S. Vitter. Multipledictionary compression using partial matching. In Storer and Cohn [SC95], pages 272–281. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ????

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=515517. IEEE catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Hoang:1996:ECM

[HLV96] D. T. Hoang, P. M. Long, and J. S. Vitter. Efficient cost measures for motion compensation at low bit rates. In Storer and Cohn [SC96], pages 102–111. ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1996. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=488315. IEEE Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Hoang:1997:LFM

[HLV97] D. T. Hoang, E. L. Linzer, and J. S. Vitter. A lexicographic framework for MPEG rate control. In Storer and Cohn [SC97], pages 101–110. ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1997. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=581982. IEEE

Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Hunt:1976:ADF

[HM76] James W. Hunt and M. Douglas McIlroy. An algorithm for differential file comparison. Computing Science Technical Report 41, Bell Labs, Murray Hill, NJ, USA, June 1976. ?? pp.

Held:1987:DCT

[HM87] Gilbert Held and Thomas (Thomas R.) Marshall. Data compression: techniques and applications: hardware and software considerations. John Wiley, New York, NY, USA, second edition, 1987. ISBN 0-471-91280-8. xii + 206 pp. LCCN QA76.9.D33 H44 1987.

Held:1991:DCT

[HM91] Gilbert Held and Thomas (Thomas R.) Marshall. Data compression: techniques and applications: hardware and software considerations. John Wiley, New York, NY, USA, third edition, 1991. ISBN 0-471-92941-7. xi + 301 pp. LCCN QA76.9.D33 H44 1991.

Hung:1993:ACO

[HM93] A. C. Hung and H.-Y. Meng. Adaptive channel optimization of vector quantized data. In Storer and Cohn [SC93], pages 282– 291. ISBN 0-8186-3391-3 (microfiche), 0-8186-3392-1 (casebound). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 1993. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=253121. IEEE Computer

Society Press order number 3392-02. IEEE catalog number 93TH0536-3.

Hung:1994:MRQ

[HM94] A. C. Hung and T. H.-Y. Meng. Multidimensional rotations for quantization. In Storer and Cohn [SC94], pages 32–41. ISBN 0-8186-5636-0, 0-8186-5637-9 (case). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1994.

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=305910. IEEE Catalog Number 93TH0626-2. IEEE Computer Society Press Order Number 5637-02.

Hahn:1996:DLV

[HM96] P. J. Hahn and V. J. Mathews. Distortion-limited vector quantization. In Storer and Cohn [SC96], pages 340–348. ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1996. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=488339. IEEE Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Hahn:1997:PLI

[HM97] P. J. Hahn and V. J. Mathews. Perceptually lossless image compression. In Storer and Cohn [SC97], page ?? ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1997. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=582100. IEEE Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Hoang:2007:CBC

[HM07] Anh Tuan Hoang and Mehul Motani. Collaborative broadcasting and compression in cluster-based wireless sensor networks. ACM Transactions on Sensor Networks, 3(3):17:1–17:??, August 2007. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Han:2010:ETD

[HMR10] Jingning Han, V. Melkote, and K. Rose. Estimation-theoretic delayed decoding of predictively encoded video sequences. In Storer and Marcellin [SM10b], pages 119–128. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=5453454.

Haddadin:1992:SVQ

[HMS92] O. S. Haddadin, V. J. Mathews, and T. G. Stockham, Jr. Subband vector quantization of images using hexagonal filter banks. In Storer and Cohn [SC92], pages 2–11. ISBN 0-8186-2717-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=227481. IEEE catalog number 91TH0436-6.

Havas:1994:GHH

[HMWC94] G. Havas, B. S. Majewski, N. C. Wormald, and Z. J. Czech. Graphs, hypergraphs and hashing. In van Leeuwen [vL94], pages 153–165. CODEN LNCSD9. ISBN 0-86636-292-4 (New York), 3-540-57899-4 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .I647 1993. URL http://

link.springer-ny.com/link/service/series/0558/tocs/
t0790.htm;   http://www.springerlink.com/content/978-
0-86636-292-4;   http://www.springerlink.com/openurl.
asp?genre=issue&issn=0302-9743&volume=790.

Hong:2002:OCR

[HN02] B. Hong and A. Nosratinia. Overhead-constrained rateallocation for scalable video transmission over networks. In Storer and Cohn [SC02], page ?? ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2002. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=999998. IEEE Computer Society Order Number PR01477.

Hong:2007:ATC

[HN07] E. S. Hong and S.-F. Newman. Applying Tunstall coding in the existing SEED format for seismographic data. In Storer and Cohn [SC07], page 384. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=4148785. IEEE Computer Society Order Number P2791.

Huang:2016:CFL

[HNC^{+}16] Xiaomeng Huang, Yufang Ni, Dexun Chen, Songbin Liu, Haohuan Fu, and Guangwen Yang. Czip: A fast lossless compression algorithm for climate data. International Journal of Parallel Programming, 44(6):1248–1267, December 2016. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640

(electronic). URL http://link.springer.com/article/10.
1007/s10766-016-0403-z.

Huybrechs:2019:HFA

[HO19] Daan Huybrechs and Peter Opsomer. High-frequency asymptotic compression of dense BEM matrices for general geometries without ray tracing. Journal of Scientific Computing, 78(2):710–745, February 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:/

/link.springer.com/article/10.1007/s10915-018-0786-

7;   https://link.springer.com/content/pdf/10.1007/
s10915-018-0786-7.pdf.

Hoang:1999:RTV

[Hoa99] D. T. Hoang. Real-time VBR rate control of MPEG video based upon lexicographic bit allocation. In Storer and Cohn [SC99], pages 374–383. ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1999. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=755687. IEEE

Computer Society Order Number PR00096. IEEE Order Plan Catalog Number PR00096.

Hoces:1995:MRF

[Hoc95] Carlos Hoces. Media reviews: Fractal compression theory: Image Compression Using Fractals, edited by Yuval Fisher (Springer-Verlag, 1995, ISBN 0-387-94211-4, 324 pp., $49.50). IEEE MultiMedia, 2(3):76, Fall 1995. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Hoces:1996:MRB

[Hoc96] Carlos Hoces. Media reviews: a broad guide to data compression: Introduction to Data Compression by Khalid Sayood (Morgan Kaufmann Publishers, 1995, 600 pp., $64.95, ISBN 1-55860-346-8). IEEE MultiMedia, 3(3):91–92, Fall 1996. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166

(electronic). URL http://dlib.computer.org/mu/books/
mu1996/pdf/u3090.pdf.

Hoffman:1997:DCD

[Hof97] Roy (Roy L.) Hoffman. Data compression in digital systems. Digital multimedia standards series. Chapman and Hall, Ltd., London, UK, 1997. ISBN 0-412-08551-8. xvi + 415 pp. LCCN QA76.9.D33 H64 1997.

Huang:2004:CFI

[HOG04] Y.-Z. Huang, D. O’Brien, and R. M. Gray. Classification of features and images using Gauss mixtures with VQ clustering. In Storer and Cohn [SC04], pages 13–21. ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1281446. IEEE Computer Society

Order Number: P2082.

Hooker:1978:OSQ

[Hoo78] Lawrence Eugene Hooker. Optimum Sampling and Quantizing for Multiple Channel Data Compression Systems. Ph.D. thesis, North Carolina State University, Raleigh, NC, USA,

1978. 306 pp. URL http://search.proquest.com/docview/
302897622.

Hopkins:2009:RTD

[Hop09] Brian Hopkins, editor. Resources for teaching discrete mathematics: classroom projects, history modules, and articles, volume 74 of MAA notes. Mathematical Association of America, Washington, DC, USA, 2009. ISBN 0-88385-184-9. xiv + 323 pp. LCCN QA13 .R4675 2009.

Horspool:1991:IL

[Hor91a] N. R. Horspool. Improving LZW. In James A. Storer, editor, Proceedings of the 1991 Data Compression Conference, pages 332–341. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1991.

Horspool:1991:ILD

[Hor91b] R. N. Horspool. Improving LZW [data compression algorithm]. In Storer and Reif [SR91b], pages 332–341. ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1990. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=213347. IEEE Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Horspool:1995:ENG

[Hor95] R. N. Horspool. The effect of non-greedy parsing in Ziv– Lempel compression methods. In Storer and Cohn [SC95], pages 302–311. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

515520. IEEE catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Hekland:2005:USM

[HOR05] F. Hekland, G. E. Oien, and T. A. Ramstad. Using 2: 1 Shannon mapping for joint source-channel coding. In Storer and Cohn [SC05], pages 223–232. ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=1402183.

Horstmann:2006:WST

[Hor06] Cay S. Horstmann. Windows shell tutorial, 2006. URL http:/

/www.horstmann.com/bigj/help/windows/tutorial.html.

Hosang:2002:CEA

[Hos02] M. Hosang. A character elimination algorithm for lossless data compression. In Storer and Cohn [SC02], page ?? ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2002. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=1000000. IEEE

Computer Society Order Number PR01477.

Houston:1979:AFM

[Hou79] George Branston Houston. The Application of Formal Models to a Fully Adaptive General Purpose Digital Data Compression System. Ph.D. thesis, University of Washington, Seattle, WA,

USA, 1979. 359 pp. URL http://search.proquest.com/
docview/302947120.

Houghton:1997:EEC

[Hou97] A. D. Houghton. The Engineer’s Error Coding Handbook. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. ISBN 1-4613-0447-4. 272 pp. LCCN TK7888.4.

Howard:1993:DAE

[How93] Paul Glor Howard. The design and analysis of efficient lossless data compression systems. Ph.D. thesis, Brown University, Providence, RI, USA, 1993. 135 pp. URL http://search.

proquest.com/docview/304072223.

Howard:1996:LLC

[How96] P. G. Howard. Lossless and lossy compression of text images by soft pattern matching. In Storer and Cohn [SC96], pages 210–219. ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1996. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=488326. IEEE

Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Howard:1998:IEC

[How98] Paul G. Howard. Interleaving entropy codes. In Proceedings Compression and Complexity of Sequences 1997, Salerno, Italy, pages 45–55. ????, ????, June 1998.

Hopper:1992:CGS

[HP92] T. Hopper and F. Preston. Compression of grey-scale fingerprint images. In Storer and Cohn [SC92], pages 309– 318. ISBN 0-8186-2717-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=227450. IEEE catalog

number 91TH0436-6.

Holzmann:1995:EHD

[HP95] Gerard J. Holzmann and Björn Pehrson. The Early History of Data Networks. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ??

pp. URL http://labit501.upct.es/ips/libros/TEHODN/
ch-2-5.3.html.

Hon:2011:CPS

[HPST11] Wing-Kai Hon, M. Patil, R. Shah, and S. V. Thankachan. Compressed property suffix trees. In Storer and Marcellin [SM11b], pages 123–132. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

5749470. IEEE Computer Society Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Huang:2011:SDCb

[HPT11] Bormin Huang, Antonio J. Plaza, and Carole Thiébaut, editors. Satellite data compression, communications, and processing VII: 23–24 August 2011, San Diego, California, United States, volume 8157. SPIE Optical Engineering Press, Bellingham, WA, USA, 2011. ISBN 0-8194-8767-8. ISSN 0277-786X (print), 1996-756X (electronic). LCCN QC371 .S67 v. 8157; TK5102.92 .S3745 2011.

Huang:2009:SDC

[HPV09] Bormin Huang, Antonio J. Plaza, and Raffaele Vitulli, editors. Satellite data compression, communication, and processing V: 4–5 August 2009, San Diego, California, United States, volume 7455. SPIE Optical Engineering Press, Bellingham, WA, USA, 2009. ISBN 0-8194-7745-1. LCCN TK5102.92 .S374 2009.

Hoobin:2011:RLZ

[HPZ11] Christopher Hoobin, Simon J. Puglisi, and Justin Zobel. Relative Lempel–Ziv factorization for efficient storage and retrieval of Web collections. Proceedings of the VLDB Endowment, 5 (3):265–273, November 2011. CODEN ???? ISSN 2150-8097.

Hunter:1980:IDF

[HR80] R. Hunter and A. H. Robinson. International digital facsimile coding standards. Proceedings of the IEEE, 68(7):854–867, July 1980. CODEN IEEPAD. ISSN 0018-9219 (print), 1558-2256 (electronic).

Hakonsen:2006:ITF

[HR06a] G. Hakonsen and T. A. Ramstad. Image transmission over flat fading channels using joint source channel coding. In Storer and Cohn [SC06], page 452. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=1607295.

Hekland:2006:DSM

[HR06b] F. Hekland and T. A. Ramstad. Digitising the 2:1 Shannon mappings for transport over heterogeneous networks. In Storer and Cohn [SC06], page 455. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=1607298.

Holub:2011:LDC

[HŘŠ11] J. Holub, J. Řezníček, and F. Šimek. Lossless data compression testbed: ExCom and Prague Corpus. In Storer and Marcellin [SM11b], page 457. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ????

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=5749514. IEEE Computer Society Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Hirschberg:1994:PAD

[HS94] D. S. Hirschberg and L. M. Stauffer. Parsing algorithms for dictionary compression on the PRAM. In Storer and Cohn [SC94], pages 136–145. ISBN 0-8186-5636-0, 0-8186-5637-9 (case). ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN QA76.9.D33 D37 1994. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=305921. IEEE

Catalog Number 93TH0626-2. IEEE Computer Society Press Order Number 5637-02.

Helfgott:1997:GNS

[HS97] H. A. Helfgott and J. A. Storer. Generalized node splitting and bilevel image compression. In Storer and Cohn [SC97], page ?? ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1997. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=582102. IEEE

Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Hans:1998:AIA

[HS98] M. Hans and R. Schafer. AudioPaK — an integer arithmetic lossless audio codec. In Storer and Cohn [SC98], page ?? ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1998. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=672286. IEEE Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Hans:2001:LCD

[HS01] Mat Hans and R. W. Schafer. Lossless compression of digital audio. IEEE signal processing magazine, 18(4):21–32, July 2001. CODEN ISPRE6. ISSN 1053-5888 (print), 1558-0792 (electronic).

Hariharan:2006:ERC

[HS06] S. Hariharan and P. Shankar. Evaluating the role of context in syntax directed compression of XML documents. In Storer and Cohn [SC06], page 453. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=1607296.

Haleem:2007:OSC

[HS07] M. A. Haleem and K. P. Subbalakshmi. Optimal sourcechannel decoder for correlated Markov sources over additive Markov channels. In Storer and Cohn [SC07], page 383. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4148784. IEEE Computer Society

Order Number P2791.

Hsiang:2001:EIC

[Hsi01] Shih-Ta Hsiang. Embedded image coding using zeroblocks of subband/wavelet coefficients and context modeling. In Storer and Cohn [SC01c], pages 83–92. ISBN 0-7695-1031-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 2001. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=917139. IEEE CSP num-

ber 01PR1031.

Hu:2020:ATL

[HSL^{+}20] Xinjue Hu, Jingming Shan, Yu Liu, Lin Zhang, and Shervin Shirmohammadi. An adaptive two-layer light field compression scheme using GNN-based reconstruction. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(2s):72:1–72:23, July 2020. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/3395620.

Hecht:1942:EQV

[HSP42] S. Hecht, S. Schlaer, and M. H. Pirenne. Energy, quanta and vision. Journal of the Optical Society of America, 38:196–208, 1942. CODEN JOSAAH. ISSN 0030-3941.

Holub:2013:GAD

[HSP^{+}13] Petr Holub, Martin Srom, Martin Pulec, Jirí Matela, and Martin Jirman. GPU-accelerated DXT and JPEG compression schemes for low-latency network transmissions of HD, 2K, and 4K video. Future Generation Computer Systems, 29(8):1991–2006, October 2013. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167739X13001209.

Huybrechs:2004:NWN

[HSV04] Daan Huybrechs, Jo Simoens, and Stefan Vandewalle. A note on wave number dependence of wavelet matrix compression for integral equations with oscillatory kernel. Journal of Computational and Applied Mathematics, 172(2):233–246, December 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print),

1879-1778 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0377042704001177.

Hon:2008:CID

[HSV^{+}08] Wing-Kai Hon, R. Shah, J. S. Vitter, Tak-Wah Lam, and Siu-Lung Tarn. Compressed index for dictionary matching. In Storer and Marcellin [SM08], pages 23–32. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33 INTER-

NET. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=4483280.

Hamzaoui:2002:RBV

[HSX02] R. Hamzaoui, V. Stankovic, and Zixiang Xiong. Rate-based versus distortion-based optimal joint source-channel coding. In Storer and Cohn [SC02], pages 63–72. ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2002. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=999944. IEEE Computer Society Order Number PR01477.

Han:2017:CCF

[HSZ17] Yunheng Han, Weiwei Sun, and Baihua Zheng. COMPRESS: a comprehensive framework of trajectory compression in road networks. ACM Transactions on Database Systems, 42(2): 11:1–11:??, June 2017. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Heroy:2018:RGC

[HTS^{+}18] Samuel Heroy, Dane Taylor, F. Bill Shi, M. Gregory Forest, and Peter J. Mucha. Rigid graph compression: Motif-based rigidity analysis for disordered fiber networks. Multiscale Modeling & Simulation, 16(3):1283–1304, ???? 2018. CODEN MMSUBT. ISSN 1540-3459 (print), 1540-3467 (electronic).

Haemmerle:1999:CBS

[HU99a] J. Haemmerle and A. Uhl. Classification based speed-up methods for fractal image compression on multicomputers. Lecture Notes in Computer Science, 1557:276–285, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hammerle:1999:CBS

[HU99b] Jutta Hämmerle and Andreas Uhl. Classification based speed-up methods for fractal image compression on multicomputers. Lecture Notes in Computer Science, 1557:276–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer-ny.com/link/
service/series/0558/bibs/1557/15570276.htm;   http:
//link.springer-ny.com/link/service/series/0558/papers/
1557/15570276.pdf.

Huang:1991:MPI

[Hua91] Shih-Chi Huang. Multilayer perceptrons for image data compression and speech recognition. Ph.D. thesis, University of Notre Dame, Notre Dame, IN, USA, 1991. 200 pp. URL

http://search.proquest.com/docview/303940741.

Huang:2001:DHI

[Hua01] Cai Wei Huang. Design of the down hole instruments of multiplexed logging system and research on the data compression of DHV logging system. M.Eng., Beijing University of Aeronautics and Astronautics, Beijing, Peoples Republic of

China, 2001. URL http://search.proquest.com/docview/
1024707067.

Huang:2003:DCM

[Hua03] Zheng Huang. Data compression methods for multisensor Kalman filter fusion. M.S., Sichuan University, Chengdu, Peoples Republic of China, 2003. URL http://search.

proquest.com/docview/1024712628.

Huang:2011:SDCa

[Hua11] Bormin Huang, editor. Satellite Data Compression. Springer Science+Business Media, LLC, New York, NY, 2011. ISBN 1-4614-1182-3, 1-4614-1183-1. ix + 309 pp. LCCN QA76.9.D33 .S35 2011; QA76.9.D33 INTERNET. URL http://www.loc.

gov/catdir/enhancements/fy1208/2011939205-d.html;
http://www.loc.gov/catdir/enhancements/fy1208/2011939205-
t.html.

Huffman:1952:MCM

[Huf52] David Huffman. A method for the construction of minimum redundancy codes. Proceedings of the Institute of Radio Engineers, 40(9):1098–1101, 1952. CODEN PIREAE. ISSN 0096-8390.

Husson:1970:DCS

[Hus70] Georges Eugene Husson. Data compression systems. Dip.Eng. (Elec.), M.Eng., McGill University, Montréal, QC, Canada,

1970. iii + 125 pp. URL http://digitool.library.mcgill.
ca/thesisfile49103.pdf;   http://www.collectionscanada.
ca/obj/thesescanada/vol1/QMM/TC-QMM-49103.pdf.

Huynh:1994:MDF

[Huy94] L. Huynh. Multiplication and division free adaptive arithmetic coding techniques for bi-level images. In Storer and Cohn [SC94], pages 264–273. ISBN 0-8186-5636-0, 0-8186-5637-9 (case). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1994. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=305934. IEEE

Catalog Number 93TH0626-2. IEEE Computer Society Press Order Number 5637-02.

Howard:1991:AAC

[HV91a] P. G. Howard and J. S. Vitter. Analysis of arithmetic coding for data compression. In Storer and Reif [SR91b], pages 3–12. ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1990. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=213368. IEEE

Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Howard:1991:NML

[HV91b] P. G. Howard and J. S. Vitter. New methods for lossless image compression using arithmetic coding. In Storer and Reif [SR91b], pages 257–266. ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1990. URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

213355. IEEE Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Howard:1992:PLI

[HV92a] P. G. Howard and J. S. Vitter. Parallel lossless image compression using Huffman and arithmetic coding. In Storer and Cohn [SC92], pages 299–308. ISBN 0-8186-2717-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=227451. IEEE catalog number 91TH0436-6.

Howard:1992:EMH

[HV92b] Paul G. Howard and J. S. Vitter. Error modeling for hierarchical lossless image compression. In Storer and Cohn [SC92], pages 269–278. ISBN 0-8186-2717-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

227454. IEEE catalog number 91TH0436-6.

Howard:1992:NML

[HV92c] Paul G. Howard and J. S. Vitter. New methods for lossless image compression using arithmetic coding. Information Processing and Management, 28(6):765–779, 1992. CODEN IP-MADK. ISSN 0306-4573 (print), 1873-5371 (electronic).

Howard:1992:PIA

[HV92d] Paul G. Howard and J. S. Vitter. Practical implementations of arithmetic coding. In James A. Storer, editor, Image and Text Compression, pages 85–112. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Nether-

lands, 1992. URL http://www.cs.duke.edu/~jsv/Papers/
catalog/node85.html.

Howard:1993:DAF

[HV93a] P. G. Howard and J. S. Vitter. Design and analysis of fast text compression based on quasi-arithmetic coding. In Storer and Cohn [SC93], pages 98–107. ISBN 0-8186-3391-3 (microfiche), 0-8186-3392-1 (casebound). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1993. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=253140. IEEE Computer Society Press order number 3392-02. IEEE catalog number 93TH0536-3.

Howard:1993:FEL

[HV93b] Paul G. Howard and J. S. Vitter. Fast and efficient lossless image compression. In Storer and Cohn [SC93], pages 351– 360. ISBN 0-8186-3391-3 (microfiche), 0-8186-3392-1 (casebound). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 1993. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=253114. IEEE Computer

Society Press order number 3392-02. IEEE catalog number 93TH0536-3.

Howard:1994:DAF

[HV94a] Paul G. Howard and J. S. Vitter. Design and analysis of fast text compression based on quasi-arithmetic coding. Information Processing and Management, 30(6):777–790, 1994. CODEN IPMADK. ISSN 0306-4573 (print), 1873-5371

(electronic). URL http://www.cs.duke.edu/~jsv/Papers/
catalog/node91.html.

Howard:1994:FPL

[HV94b] Paul G. Howard and J. S. Vitter. Fast progressive lossless image compression. In Proceedings of the Image and Video Compression Conference: IS&T/SPIE 1994 Symposium on Electronic Imaging: Science & Technology, February 1994, San Jose, CA, volume 2186, pages 98–109. SPIE Optical Engineering Press, Bellingham, WA, USA, February 1994.

Howard:1996:PLI

[HV96] Paul G. Howard and Jeffrey Scott Vitter. Parallel lossless image compression using Huffman and arithmetic coding. Information Processing Letters, 59(2):65–73, July 22, 1996. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hitchcock:2006:DER

[HV06a] John M. Hitchcock and N. V. Vinodchandran. Dimension, entropy rates, and compression. Journal of Computer and System Sciences, 72(4):760–782, June 2006. CO-DEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022000005001248.

Huybrechs:2006:TDW

[HV06b] Daan Huybrechs and Stefan Vandewalle. A two-dimensional wavelet-packet transform for matrix compression of integral equations with highly oscillatory kernel. Journal of Computational and Applied Mathematics, 197(1):218–232, December 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print),

1879-1778 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0377042705006552.

Han:2008:OBR

[HV08] Sunhyoung Han and N. Vasconcelos. Object-based regions of interest for image compression. In Storer and Marcellin [SM08], pages 132–141. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008;

QA76.9.D33 INTERNET. URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=4483291.

Hartenstein:1998:ATQ

[HW98] H. Hartenstein and Xiaolin Wu. Analysis of trellis quantization for near-lossless image coding. In Storer and Cohn [SC98], page ?? ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1998. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=672288. IEEE

Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Huang:2006:OIA

[HW06] Xiang Huang and Xiaolin Wu. Optimal index assignment for multiple description lattice vector quantization. In Storer and Cohn [SC06], pages 272–281. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=1607262.

Heymann:2006:SDC

[HWS06] Roger W. Heymann, Charles Chien-Yao Wang, and Timothy J. Schmit, editors. Satellite data compression, communications, and archiving II: 13–14 August, 2006, San Diego, California, USA, volume 6300 of Proceedings of SPIE. SPIE Optical Engineering Press, Bellingham, WA, USA, 2006. ISBN 0-8194-6379-5. ISSN 0277-786X (print), 1996-756X (electronic). LCCN TK5102.92 .S37 2006. URL http://www.loc.gov/

catdir/toc/fy0708/2007270591.html.

Hon:2010:CIA

[HWY10] Wing-Kai Hon, Winson Wu, and Ting-Shuo Yang. Compressed indexes for approximate library management. In Storer and Marcellin [SM10b], page 534. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=5453513.

Hamano:2010:DCB

[HY10] K. Hamano and H. Yamamoto. Data compression based on a dictionary method using recursive construction of T-codes. In Storer and Marcellin [SM10b], page 531. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=5453508.

hydrogenaudio:2006:HAT

[hyd06] hydrogenaudio. hydrogenaudio: the audio technology enthusi-

ast’s resource, 2006. URL http://www.hydrogenaudio.org/
forums/.

Hsu:1995:ASC

[HZ95] William H. Hsu and Amy E. Zwarico. Automatic synthesis of compression techniques for heterogeneous files. Software — Practice and Experience, 25(10):1097–1116, October 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hu:2011:IPA

[HZKL11] Yichuan Hu, Jianzhong Zhang, Farooq Khan, and Ying Li. Improving PPM algorithm using dictionaries. In Storer and Marcellin [SM11b], page 459. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=5749516. IEEE Computer Society Order Number P4352; BMS Part Number: CFP11DCC-PRT.

He:2002:IFG

[HZZY02] Yuwen He, Xuejun Zhao, Yuzhuo Zhong, and Shiqiang Yang. Improved fine granular scalable coding with interlayer prediction. In Storer and Cohn [SC02], pages 172– 181. ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2002. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=999955. IEEE

Computer Society Order Number PR01477.

Iren:2000:SNN

[IA00] S. Iren and P. D. Amer. SPIHT-NC: network-conscious zerotree encoding. In Storer and Cohn [SC00], pages 313–322.

ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2000. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=838171. This millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

IBM:1988:IJR

[IBM88] IBM. IBM Journal of Research and Development, volume 32(6). IBM Corporation, San Jose, CA, USA, November 1988. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). ?? pp. Special issue on encoding.

Iyengar:1998:EFS

[IC98] Vikram Iyengar and Krishnendu Chakrabarty. An efficient finite-state machine implementation of Huffman decoders. Information Processing Letters, 64(6):271–275, January 15, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [TM98].

Ismaeil:1999:MEU

[IDKW99] I. R. Ismaeil, A. Docef, F. Kossentini, and R. Ward. Motion estimation using long-term motion vector prediction. In Storer and Cohn [SC99], page ?? ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1999. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=785688. IEEE

Computer Society Order Number PR00096. IEEE Order Plan Catalog Number PR00096.

Ivie:1991:TIC

[IE91] E. L. Ivie and J. J. Ekstrom. Techniques for index compression. In Storer and Reif [SR91b], page ?? ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1990. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=213310. IEEE Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

IEEE:1984:ASF

[IEE84] IEEE, editor. 25th annual Symposium on Foundations of Computer Science, October 24–26, 1984, Singer Island, Florida.

IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1984. CODEN ASFPDV. ISBN 0-8186-8591-3, 0-8186-0591-X (paperback), 0-8186-4591-1 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1984. IEEE catalog no. 84CH2085-9.

IEEE754:1985:ISB

[IEE85] IEEE754. ANSI/IEEE Standard 754-1985: IEEE standard for binary floating-point arithmetic., 1985.

IEEE:1988:DTS

[IEE88a] IEEE, editor. Digital technology — spanning the universe: IEEE International Conference on Communications ’88: Philadelphia, PA, June 12–15, 1988, Wyndham Franklin Plaza Hotel: conference record. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1988. ISBN ???? LCCN TK5101.A1 I34 1988. IEEE catalog number 88CH2538-7.

IEEE:1988:PSN

[IEE88b] IEEE, editor. Proceedings, Supercomputing ’88: November 14– 18, 1988, Orlando, Florida, volume 1. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1988. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

IEEE:1991:PAS

[IEE91] IEEE, editor. Proceedings: 32nd annual Symposium on Foundations of Computer Science, San Juan, Puerto Rico, October 1–4, 1991. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1991. CO-DEN ASFPDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

IEEE:1995:HCV

[IEE95a] IEEE, editor. Hot chips VII: symposium record: Stanford University, Stanford, California, August 1995. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN ???? LCCN ????

IEEE:1995:IPIa

[IEE95b] IEEE, editor. ICASSP ’95: Proceedings of the 1995 IEEE International Conference on Acoustics, Speech, and Signal Processing, May 9–12, 1995, Detroit, Michigan. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. CODEN IPRODJ. ISBN 0-7803-2431-5. ISSN 0736-7791. LCCN TK7882.S65 .I16 1995. Five volumes. IEEE Catalog No. 95CH35732.

IEEE:1995:IPIb

[IEE95c] IEEE, editor. ICIP’95: Proceedings, International Conference on Image Processing: October 23–26, 1995, Washington, DC, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-7803-3122-2 (casebound), 0-8186-7310-9 (softbound), 0-7803-3123-0 (microfiche), 0-7803-2749-7 (CD-ROM). LCCN TK8315.I222 1995. Three volumes. IEEE catalog number 95CB35819.

IEEE:1996:ASF

[IEE96] IEEE, editor. 37th Annual Symposium on Foundations of Computer Science: October 14–16, 1996, Burlington, Vermont. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. CODEN ASF-PDV. ISBN 0-7803-3762-X (casebound), 0-8186-7594-2 (softbound), 0-8186-7596-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1996. IEEE catalog number 96CH35973. IEEE Computer Society Press order number PR07594.

IEEE:2004:IIC

[IEE04] IEEE, editor. 2004 IEEE International Conference on Acoustics, Speech, and Signal Processing: proceedings: May 17– 21, 2004, Fairmont Queen Elizabeth Hotel, Montreal, Quebec, Canada (ICASSP ’04). IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2004. ISBN 0-7803-8484-9. LCCN TK7882.S65 I61

2004. URL http://ieeexplore.ieee.org/servlet/opac?

punumber=9248. IEEE Catalog Number: 04CH37568.

IEEE:2005:ICI

[IEE05] IEEE, editor. 2005 International Conference on Image Processing (ICIP): September 11–14, 2005, Genova, Italy. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2005. ISBN 0-7803-9134-9. LCCN ???? Five volumes. IEEE catalog number 05CH37687.

IEEE:2009:DCC

[IEE09] IEEE, editor. 2009 Data Compression Conference, January 2009. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2009. ISBN 1-4244-3753-9 (paperback). LCCN ????

IEEE:2010:DCCa

[IEE10] IEEE, editor. 2010 Data Compression Conference, January 2010. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2010. ISBN 1-4244-6425-0 (paperback). LCCN ????

IEEE:2011:DCC

[IEE11a] IEEE, editor. 2011 Data Compression Conference, March 2011. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2011. ISBN 1-61284-279-8 (paperback). LCCN ????

IEEE:2011:FIC

[IEE11b] IEEE, editor. 2011 First International Conference on Data Compression, Communications and Processing, September 2011. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2011. ISBN 1-4577-1458-2 (paperback). LCCN ????

Ingber:2006:NAD

[IF06] A. Ingber and M. Feder. Non-asymptotic design of finite state universal predictors for individual sequences. In Storer and Cohn [SC06], pages 3–12. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=1607235.

Ingber:2007:PPB

[IF07] A. Ingber and M. Feder. Power preserving 2: 1 bandwidth reduction mappings. In Storer and Cohn [SC07], page 385. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=4148786. IEEE Computer

Society Order Number P2791.

Imaizumi:2012:KDS

[IFK12] Shoko Imaizumi, Masaaki Fujiyoshi, and Hitoshi Kiya. A key derivation scheme for hierarchical access control to JPEG 2000 coded images. Lecture Notes in Computer Science, 7088: 180–191, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/
chapter/10.1007/978-3-642-25346-1_17/.

Ignatchenko:1998:AOD

[Ign98] Sergey Ignatchenko. An algorithm for online data compression. C/C++ Users Journal, 16(10):63–??, October 1998. CODEN CCUJEX. ISSN 1075-2838.

Iwata:2002:LZC

[II02] Kazunori Iwata and Naohiro Ishii. Lempel–Ziv coding in reinforcement learning. Lecture Notes in Computer Science, 2412:531–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer-

ny.com/link/service/series/0558/bibs/2412/24120531.
htm;   http://link.springer-ny.com/link/service/series/
0558/papers/2412/24120531.pdf.

Ivanov:2000:CDN

[IK00] Denis V. Ivanov and Yevgeniy P. Kuzmin. Color distribution — a new approach to texture compression. Computer Graphics Forum, 19(3):??, August 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=420&amp;

 http://www.blackwellpublishers.co.uk/asp/journal.
asp?ref=0167-7055&amp;   iid=3&amp; src=ard&amp; vid=
19.

Ingber:2011:DLS

[IK11] A. Ingber and Y. Kochman. The dispersion of lossy source coding. In Storer and Marcellin [SM11b], pages 53–62. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5749463. IEEE Computer Society

Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Ishii:1982:HED

[IKM82] Atsushi Ishii, Kenji Kouno, and Yasunori Maezawa. Highly efficient data compression method for newspaper image data.

Fujitsu scientific and technical journal, V 18(N 2):199–225, June 1982. CODEN FUSTA4. ISSN 0016-2523.

Ibarria:2007:SP

[ILR07] L. Ibarria, P. Lindstrom, and J. Rossignac. Spectral predictors. In Storer and Cohn [SC07], pages 163–172. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4148755. IEEE Computer Society

Order Number P2791.

Ibarria:2003:CCC

[ILRS03] Lawrence Ibarria, Peter Lindstrom, Jarek Rossignac, and Andrzej Szymczak. Compression: Out-of-core compression and decompression of large n-dimensional scalar fields. Computer Graphics Forum, 22(3):343–348, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Isal:2001:PSB

[IM01] R. Y. K. Isal and A. Moffat. Parsing strategies for BWT compression. In Storer and Cohn [SC01c], pages 429–438. ISBN 0-7695-1031-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2001. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=917174. IEEE

CSP number 01PR1031.

IMA:2006:IIM

[IMA06] IMA. IMA: The Institute for the Musical Arts home pages,

2006. URL http://www.ima.org/.

Islam:2013:MSC

[IMB^{+}13] Tanzima Zerin Islam, Kathryn Mohror, Saurabh Bagchi, Adam Moody, Bronis R. de Supinski, and Rudolf Eigenmann. McrEngine: a scalable checkpointing system using data-aware aggregation and compression. Scientific Programming, 21(3– 4):149–163, ???? 2013. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Indjic:1991:RPS

[Ind91] D. Indjic. Reduction in power system load data training sets size using fractal approximation theory. In Storer and Reif [SR91b], page ?? ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN QA76.9.D33 D37 1990. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=213315. IEEE

Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Ingber:2008:SJS

[Ing08] A. Ingber. Simple joint source-channel coding schemes for colored Gaussian sources. In Storer and Marcellin [SM08], page 522. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33

INTERNET. URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4483349.

I:2016:FLF

[INI^{+}16] Tomohiro I., Yuto Nakashima, Shunsuke Inenaga, Hideo Bannai, and Masayuki Takeda. Faster Lyndon factorization algorithms for SLP and LZ78 compressed text. Theoretical Computer Science, 656 (Part B)(??):215–224, December 20, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-

2294 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030439751600195X.

Ilie:2006:SCS

[IP06] Lucian Ilie and Cristian Popescu. The shortest common superstring problem and viral genome compression. Fundamenta Informaticae, 73(1–2):153–164, October 2006. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

ITU-R:1997:MSA

[IRB97] ITU-R and BS1116. Methods for the subjective assessment of small impairments in audio systems including multichannel sound systems. UTR-R document BS 1116 (Rev. 1), ITU, Geneva, Switzerland, 1997. ?? pp.

Isenburg:2001:TSC

[Ise01] Martin Isenburg. Triangle strip compression. Computer Graphics Forum, 20(2):??, June 2001. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ishihara:1989:UHT

[Ish89] Fumiko Ishihara. The use of Hadamard Transform as a data compression technique in the development of a threedimensional fluorescence spectral library for qualitative analysis. Ph.D. thesis, Virginia Polytechnic Institute and State University, Blacksburg, VA, USA, 1989. 154 pp. URL http:/

/search.proquest.com/docview/303819442.

ISO:1984:IPS

[ISO84] ISO. Information processing systems — data communication high-level data link control procedure-frame structure: IS 3309, October 1984.

ISO:1993:ITC

[ISO93] ISO/IEC. International Standard IS 11172-3: Information technology—coding of moving pictures and associated audio for digital storage media at up to about 1.5 mbits/s—part 3: Audio., 1993.

ISO:1995:IIIl

[ISO95] ISO. ISO/IEC 10918-2:1995: Information technology — Digital compression and coding of continuous-tone still images: Compliance testing. International Organization for Standardization, Geneva, Switzerland, 1995. ISBN ???? 181 pp. LCCN

???? ???? URL http://www.iso.ch/cate/d20689.html.

ISO:1997:IIIm

[ISO97] ISO. ISO/IEC 10918-3:1997: Information technology — Digital compression and coding of continuous-tone still images: Extensions. International Organization for Standardization, Geneva, Switzerland, 1997. ISBN ???? 80 pp. LCCN ????

CHF 164. URL http://www.iso.ch/cate/d25037.html.

ISO:1998:IGT

[ISO98] ISO. ISO 12639:1998: Graphic technology — Prepress digital data exchange — Tag image file format for image technology (TIFF/IT). International Organization for Standardization, Geneva, Switzerland, 1998. ISBN ???? 43 pp. LCCN ????

CHF 128. URL http://www.iso.ch/cate/d2181.html.

ISO:1999:IIA

[ISO99a] ISO. ISO/IEC 10918-3:1997/Amd 1:1999: Provisions to allow registration of new compression types and versions in the SPIFF header. International Organization for Standardization, Geneva, Switzerland, 1999. ISBN ???? 1 pp. LCCN ????

CHF 13. URL http://www.iso.ch/cate/d30961.html.

ISO:1999:IIIw

[ISO99b] ISO. ISO/IEC 10918-4:1999: Information technology — Digital compression and coding of continuous-tone still images: Registration of JPEG profiles, SPIFF profiles, SPIFF tags, SPIFF colour spaces, APPn markers, SPIFF compression types and Registration Authorities (REGAUT). International Organization for Standardization, Geneva, Switzerland, 1999. ISBN ???? 29 pp. LCCN ???? CHF 104. URL http://www. iso.ch/cate/d25431.html. Available in English only.

ISO:2000:ITJ

[ISO00] ISO/IEC. International Standard IS 15444-1: Information technology—JPEG 2000 image coding system., March 16, 2000. This is the FDC (final committee draft) version 1.0.

ISO:2003:IOS

[ISO03a] ISO. International Organization for Standardization, 2003.

URL http://www.iso.ch/.

ISO:2003:ITH

[ISO03b] ISO/IEC. International Standard ISO/IEC 13818-7, Information technology, Generic coding of moving pictures and associated audio information, Part 7: Advanced Audio Coding (AAC). International Organization for Standardization, Geneva, Switzerland, second edition, 2003. ?? pp. Issued 2003-08-01.

ISO:2009:IID

[ISO09] ISO. ISO 28500:2009: Information and documentation — WARC file format. International Organization for Standardization, Geneva, Switzerland, May 2009. ?? pp. URL https:

//www.iso.org/standard/68004.html.

ISO:2017:IID

[ISO17] ISO. ISO 28500:2017: Information and documentation — WARC file format. International Organization for Standardization, Geneva, Switzerland, second edition, August 2017. 26

pp. URL https://www.iso.org/standard/68004.html.

Israelsen:1991:VIV

[Isr91] P. Israelsen. VLSI implementation of a vector quantization processor. In Storer and Reif [SR91b], page ?? ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1990. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=213298. IEEE Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

ITU-T:1989:CRG

[ITU89] ITU-T. CCITT recommendation G.711: Pulse code modulation (PCM) of voice frequencies, 1989.

ITU-T:1990:RGK

[ITU90] ITU-T. Recommendation G.726 (12/90): 40, 32, 24, 16 kbit/s Adaptive Differential Pulse Code Modulation (AD-PCM), 1990.

ITU:1991:BAQ

[ITU91] ITU/TG10. Basic audio quality requirements for digital audio bit-rate reduction systems for broadcast emission and primary distribution. ITU-R document TG-10-2/3-E, ITU, ????, October 1991. ?? pp.

ITU-T:1994:ECP

[ITU94] ITU-T. ITU-T Recommendation V.42, Revision 1: Errorcorrecting procedures for DCEs using asynchronous-to-synchronous conversion, 1994.

ITU-T264:2002:AVC

[ITU02] ITU-T264. ITU-T Recommendation H.264, ISO/IEC 11496-10: Advanced Video Coding, September 2002. Final Committee Draft, Document JVT-E022.

ITU:2006:AMS

[ITU06a] ITU. Audiovisual and multimedia systems [list of standards],

2006. URL http://www.itu.int/rec/T-REC-h.

G131:2006:TEC

[ITU06b] ITU. ITU-T recommendation G.131: Talker echo and its con-

trol, 2006. URL http://www.itu.int/rec/T-REC-G.131/en.

Inglis:1994:CBT

[IW94a] S. Inglis and I. H. Witten. Compression-based template matching. In Storer and Cohn [SC94], pages 106–115.

ISBN 0-8186-5636-0, 0-8186-5637-9 (case). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1994.

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=305918. IEEE Catalog Number 93TH0626-2. IEEE Computer Society Press Order Number 5637-02.

Iyer:1994:DCS

[IW94b] Balakrishna R. Iyer and David Wilhite. Data compression support in databases. In Bocca et al. [BJZ94], pages 695–704. ISBN 1-55860-153-8. LCCN QA76.9.D3 I559 1994. URL http:

//www.vldb.org/dblp/db/conf/vldb/vldb94-695.html.

Inglis:1996:BLD

[IW96] S. J. Inglis and I. H. Witten. Bi-level document image compression using layout information. In Storer and Cohn [SC96], page ?? ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1996. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=488374. IEEE

Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Itagaki:2000:PSC

[IY00] S. Itagaki and H. Yokoo. PPM*-style context sorting compression method using a prefix list. In Storer and Cohn [SC00], page ?? ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2000. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=838203. This

millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Inoue:2011:TCD

[IYY^{+}11] Hiroaki Inoue, Junya Yamada, Hideyuki Yoneda, Katsumi Togawa, Masato Motomura, and Koichiro Furuta. Test compression for dynamically reconfigurable processors. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 4 (4):40:1–40:??, December 2011. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Jagmohan:2003:WZE

[JA03] A. Jagmohan and N. Ahuja. Wyner–Ziv encoded predictive multiple descriptions. In Storer and Cohn [SC03], pages 213–222. ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2003. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1194012. IEEE Computer Society Order number PR01896.

Jain:2007:EBP

[JA07] S. K. Jain and D. A. Adjeroh. Edge-based prediction for lossless compression of hyperspectral images. In Storer and Cohn [SC07], pages 153–162. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ????

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=4148754. IEEE Computer Society Order Number P2791.

Jacobson:1992:RAH

[Jac92] G. Jacobson. Random access in Huffman-coded files. In Storer and Cohn [SC92], pages 368–377. ISBN 0-8186-2717-

4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=227444. IEEE catalog number 91TH0436-6.

Jellali:2019:BDS

[JAC19] Zakia Jellali, Leïla Najjar Atallah, and Sofiane Cherif. Bidimensional signal compression based on linear prediction coding: Application to WSN. ACM Transactions on Sensor Networks, 15(3):29:1–29:??, August 2019. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic). URL https://dl.

acm.org/ft_gateway.cfm?id=3317688.

Jagadish:1990:CTM

[Jag90] H. V. Jagadish. A compression technique to materialize transitive closure. ACM Transactions on Database Systems, 15(4):558–598, December 1990. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL http:/

/www.acm.org/pubs/articles/journals/tods/1990-15-4/
p558-jagadish/p558-jagadish.pdf;   http://www.acm.
org/pubs/citations/journals/tods/1990-15-4/p558-jagadish/

;   http://www.acm.org/pubs/toc/Abstracts/tods/99944.
html.

Jakobsson:1978:HCB

[Jak78] M. Jakobsson. Huffman coding in bit-vector compression. Information Processing Letters, 7(6):304–307, October ??, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jakobsson:1985:CCS

[Jak85] Matti Jakobsson. Compression of character strings by an adaptive dictionary. BIT (Nordisk tidskrift for informationsbehandling), 25(4):593–603, December 1985. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=25&issue=4&spage=593.

Jakobsson:1988:ROP

[Jak88] Matti Jakobsson. Research: One-pass text compression with a subword dictionary. Journal of the American Society for Information Science, 39(4):262–269, July 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Jenkins:2012:ADL

[JAL^{+}12] John Jenkins, Isha Arkatkar, Sriram Lakshminarasimhan, Neil Shah, and Eric R. Schendel. Analytics-driven lossless data compression for rapid in-situ indexing, storing, and querying. Lecture Notes in Computer Science, 7447:16–30, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32597-7_2/.

Jayant:1997:SCC

[Jay97] N. Jayant, editor. Signal Compression: Coding of Speech, Audio, Text, Image and Video. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1997. ?? pp.

Jordan:1974:COR

[JB74] B. W. Jordan, Jr. and R. C. Barrett. A cell organized raster display for line drawings. Communications of the ACM, 17(2): 70–77, February 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Jung:1995:RTV

[JB95] Bongjin Jung and W. P. Burleson. Real-time VLSI compression for high-speed wireless local area networks. In Storer and Cohn [SC95], page 431. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=515541. IEEE catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Jiang:1999:LCI

[JB99] W. Jiang and L. Bruton. Lossless color image compression using chromatic correlation. In Storer and Cohn [SC99], page ?? ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1999. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=785690. IEEE Computer Society Or-

der Number PR00096. IEEE Order Plan Catalog Number PR00096.

Jafarkhani:1997:ECS

[JBF97] H. Jafarkhani, H. Brunk, and N. Farvardin. Entropyconstrained successively refinable scalar quantization. In Storer and Cohn [SC97], pages 337–346. ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1997. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=582057. IEEE Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Jakob:2017:PPP

[JBG17] Johannes Jakob, Christoph Buchenau, and Michael Guthe. Parallel processing: A parallel approach to compression and decompression of triangle meshes using the GPU. Computer Graphics Forum, 36(5):71–80, August 2017. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

JBIG:2003:JH

[JBI03a] JBIG (Joint Bi-level Image experts Group). JBIG homepage,

2003. URL http://www.jpeg.org/jbig/.

JBIG2:2003:JD

[JBI03b] JBIG2. JBIG2 development, 2003. URL http://www.jpeg.

org/jbig/jbigpt2.html.

Jayant:1995:ACT

[JC95] Nikil S. Jayant and Edward Y. Chen. Audio compression: technology and applications. AT&T Technical Journal, 74(2): 23–34, 1995. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Joshi:1993:ISC

[JCF93] R. L. Joshi, V. J. Crump, and T. R. Fischer. Image subband coding using arithmetic and trellis coded quantization. IEEE Transactions on Circuits and Systems for Video Technology, 5 (6):515–523, December 1993. CODEN ITCTEM. ISSN 1051-8215 (print), 1558-2205 (electronic).

Joo:2008:ECP

[JCS^{+}08] Yongsoo Joo, Youngjin Cho, Donghwa Shin, Jaehyun Park, and Naehyuck Chang. An energy characterization platform for memory devices and energy-aware data compression for multilevel-cell flash memory. ACM Transactions on Design Automation of Electronic Systems (TODAES), 13(3):43:1– 43:??, July 2008. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Ji:2017:LDC

[JCS^{+}17] Cheng Ji, Li-Pin Chang, Liang Shi, Congming Gao, Chao Wu, Yuangang Wang, and Chun Jason Xue. Lightweight data compression for mobile flash storage. ACM Transactions on Embedded Computing Systems, 16(5s):183:1–183:??, October 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Jiang:2003:UEW

[JdVL03] Zhuhan Jiang, Olivier de Vel, and Bruce Litow. Unification and extension of weighted finite automata applicable to image compression. Theoretical Computer Science, 302(1–3):275– 294, June 13, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jagemar:2016:AMC

[JEEL16] Marcus Jägemar, Sigrid Eldh, Andreas Ermedahl, and Björn Lisper. Automatic message compression with overload protection. The Journal of Systems and Software, 121(??):209–222, November 2016. CODEN JSSODM. ISSN 0164-1212 (print),

1873-1228 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0164121216300267.

Jordan:1998:PCB

[JEK98] Corinne Le Buhan Jordan, Touradj Ebrahimi, and Murat Kunt. Progressive content-based shape compression for retrieval of binary images. Computer Vision and Image Understanding: CVIU, 71(2):198–212, August 1998. CO-DEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (elec-

tronic). URL http://www.idealibrary.com/links/artid/
cviu.1998.0707/production;   http://www.idealibrary.
com/links/artid/cviu.1998.0707/production/pdf;   http:
//www.idealibrary.com/links/artid/cviu.1998.0707/production/
ref.

Jena:1999:ILZ

[Jen99] S. K. Jena. An improved Lempel–Ziv algorithm for sequential data compression. International Journal of High Speed Computing (IJHSC), 10(3):275–284, September 1999. CODEN IH-SCEZ. ISSN 0129-0533.

Jez:2015:FFC

[Jez15] Artur Jez. Faster fully compressed pattern matching by recompression. ACM Transactions on Algorithms, 11(3):20:1–20:??, January 2015. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Jez:2016:RSP

[Jez16] Artur Jez. Recompression: a simple and powerful technique for word equations. Journal of the ACM, 63(1):4:1–4:??, March 2016. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Jafarkhani:1996:FRS

[JF96] H. Jafarkhani and N. Farvardin. Fast reconstruction of subband decomposed signals for progressive transmission. In Storer and Cohn [SC96], pages 230–239. ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1996. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=488328. IEEE Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Ji:2012:ECB

[JFC12] Wen Ji, P. Frossard, and Yiqiang Chen. EXIT chart-based side information refinement for Wyner–Ziv video coding. In Storer and Marcellin [SM12], pages 209–218. ISBN 1-4673-0715-7 (paperback). ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=6189252. IEEE Computer

Society order number P4656.

Jeong:2005:MDC

[JG05] Sangoh Jeong and R. M. Gray. Minimum distortion color image retrieval based on Lloyd-clustered Gauss mixtures. In Storer and Cohn [SC05], pages 279–288. ISBN 0-7695-2309-

9. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=1402189.

Jiang:2002:ESF

[JGL02] Jianmin Jiang, Baofeng Guo, and Pengjie Li. Extracting shape features in JPEG-2000 compressed images. Lecture Notes in Computer Science, 2457:123–??, 2002. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.de/link/service/
series/0558/bibs/2457/24570123.htm;   http://link.
springer.de/link/service/series/0558/papers/2457/24570123.
pdf.

Joseph:2020:PAN

[JGM^{+}20] V. Joseph, G. L. Gopalakrishnan, S. Muralidharan, M. Garland, and A. Garg. A programmable approach to neural network compression. IEEE Micro, 40(5):17–25, 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Jiang:1995:ILD

[Jia95] Jianmin Jiang. Improving LZFG data compression algorithm. In Storer and Cohn [SC95], page 475. ISBN 0-8186-7012-

6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=515585. IEEE catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Jahedi:2018:CDB

[JJM18] S. Jahedi, F. Javadi, and M. J. Mehdipour. Compression and decompression based on discrete weighted transform. Applied Mathematics and Computation, 335(??):133–145, October 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print),

1873-5649 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0096300318303667.

Jang:2013:PSP

[JKL13] Hakbeom Jang, Channoh Kim, and Jae W. Lee. Practical speculative parallelization of variable-length decompression algorithms. ACM SIGPLAN Notices, 48(5):55–64, May 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jung:2015:CCD

[JKV15] Eun-Sung Jung, Rajkumar Kettimuthu, and Venkatram Vishwanath. Cluster-to-cluster data transfer with data compression over wide-area networks. Journal of Parallel and Distributed Computing, 79–80(??):90–103, May 2015. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731514001713.

Jin:2008:FPD

[JLJ08] Soonjong Jin, Hyuk Lee, and Jechang Jeong. Fast partial distortion elimination algorithm for lossless and lossy motion estimation using Hadamard transform and probability model. In Storer and Marcellin [SM08], page 523. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33 INTER-

NET. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=4483350.

Jiang:2020:PAD

[JLJ^{+}20] Hao Jiang, Chunwei Liu, Qi Jin, John Paparrizos, and Aaron J. Elmore. PIDS: attribute decomposition for improved compression and query performance in columnar storage. Proceedings of the VLDB Endowment, 13(6):925–938, February 2020. CO-DEN ???? ISSN 2150-8097. URL https://dl.acm.org/doi/

abs/10.14778/3380750.3380761.

Jiang:2010:TSC

[JLL^{+}10] Wenfei Jiang, Wenyu Liu, L. J. Latecki, Hui Liang, Changqing Wang, and Bing Feng. Two-step coding for high definition video compression. In Storer and Marcellin [SM10b], page 535. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=5453512.

Janson:2004:ASC

[JLS04] S. Janson, S. Lonardi, and W. Szpankowski. On the average sequence complexity. In Storer and Cohn [SC04], page ?? ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1281518. IEEE Computer

Society Order Number: P2082.

Jackson:1996:PPF

[JM96] David Jeff Jackson and Wagdy Mahmoud. Parallel pipelined fractal image compression using quadtree recomposition. The Computer Journal, 39(1):1–13, ???? 1996. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/39/1/
1.full.pdf+html;   http://www.oup.co.uk/jnls/list/
comjnl/hdb/Volume_39/Issue_01/390001.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/Volume_39/
Issue_01/Vol39_01.body.html#AbstractJackson.

Jones:1997:CCM

[JM97] D. K. Jones and M. W. Maier. Compression comparisons for multiview stereo. In Storer and Cohn [SC97], page ?? ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1997. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=582103. IEEE Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Jana:2003:OTC

[JM03] Soumya Jana and P. Moulin. Optimal transform coding of Gaussian mixtures for joint classification/reconstruction. In Storer and Cohn [SC03], pages 313–322. ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 2003. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1194022. IEEE Computer

Society Order number PR01896.

Jahannia:2019:PZL

[JMnRS19] Marieh Jahannia, Morteza Mohammad-noori, Narad Rampersad, and Manon Stipulanti. Palindromic Ziv–Lempel and Crochemore factorizations of m-Bonacci infinite words. Theoretical Computer Science, 790(??):16–40, October 22, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304397519303184.

Jalali:2009:ISU

[JMW09] S. Jalali, A. Montanari, and T. Weissman. An implementable scheme for universal lossy compression of discrete Markov sources. In Storer and Marcellin [SM09], pages 292–301. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4976473. IEEE Computer Society

Order Number P3592. BMS Part Number CFP09DCC-PRT.

Jacquin:1997:CAP

[JOC97] A. Jacquin, H. Okada, and P. Crouch. Content-adaptive postfiltering for very low bit rate video. In Storer and Cohn [SC97], pages 111–120. ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1997. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=581986. IEEE

Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Joglekar:1973:DCR

[Jog73] Anil Narayan Joglekar. Data Compression In Recursive Estimation with Applications to Navigation Systems. Ph.D. thesis, Stanford University, Stanford, CA, USA, 1973. 149 pp. URL

http://search.proquest.com/docview/302696388.

Johnson:1989:BDC

[Joh89] Michael Paul Johnson. Beyond DES: Data compression and the MPJ encryption algorithm. M.S. thesis, University of Colorado at Colorado Springs, Colorado Springs, CO, USA,

1989. 135 pp. URL http://search.proquest.com/docview/
303808583.

Jones:1988:AST

[Jon88] Douglas W. Jones. Application of splay trees to data compression. Communications of the ACM, 31(8):996–1007, August 1988. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.org/pubs/

toc/Abstracts/0001-0782/63036.html.

Jones:1991:PED

[Jon91] D. W. Jones. Practical evaluation of a data compression algorithm. In Storer and Reif [SR91b], pages 372–381. ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1990. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=213343. IEEE Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Janowska:2007:PCT

[JP07] Agata Janowska and Wojciech Penczek. Path compression in timed automata. Fundamenta Informaticae, 79(3–4):379–399, February 2007. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

JPEG:2000:JON

[JPE00] JPEG 2000 Organization. JPEG 2000 our new standard!, 2000.

URL http://www.jpeg.org/JPEG2000.htm.

Jagmohan:2002:MDC

[JR02] A. Jagmohan and K. Ratakonda. Multiple description coding of predictively encoded sequences. In Storer and Cohn [SC02], pages 13–22. ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2002. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=999939. IEEE Computer Society Order Number PR01477.

Joseph:2012:EQP

[JR12] A. Benjamin Joseph and Baskaran Ramachandran. Enhanced quality preserved image compression technique using edge assisted wavelet based interpolation. Lecture Notes in Computer Science, 7135:146–153, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-29280-4_16/.

Jimenez-Rodriguez:2011:PEJ

[JRALMSS11] L. Jiménez-Rodríguez, F. Aulí-Llinás, M. W. Marcellin, and J. Serra-Sagristà. Pre-encoded JPEG2000 video transmission in a video-on-demand scenario. In Storer and Marcellin [SM11b], page 460. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

5749517. IEEE Computer Society Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Jacquet:1995:ABL

[JS95] Philippe Jacquet and Wojciech Szpankowski. Asymptotic behavior of the Lempel–Ziv parsing scheme and digital search trees. Theoretical Computer Science, 144(1–2):161–197, June 26, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-

2294 (electronic). URL http://www.elsevier.com/cgi-bin/
cas/tree/store/tcs/cas_sub/browse/browse.cgi?year=
1995&volume=144&issue=1-2&aid=1932.

Ji:1999:MVA

[JS99a] T. Ji and W. E. Stark. Modified Viterbi algorithm for predictive TCQ. In Storer and Cohn [SC99], page ?? ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1999. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=785689. IEEE Computer Society Or-

der Number PR00096. IEEE Order Plan Catalog Number PR00096.

Jin:1999:FPD

[JS99b] Jionghua Jin and Jianjun Shi. Feature-preserving data compression of stamping tonnage information using wavelets. Technometrics, 41(4):327–339, November 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1271349.

Jia:2020:SED

[JSC20] Yichen Jia, Zili Shao, and Feng Chen. SlimCache: an efficient data compression scheme for flash-based key-value caching.

ACM Transactions on Storage, 16(2):14:1–14:34, June 2020. CODEN ???? ISSN 1553-3077 (print), 1553-3093 (electronic).

URL https://dl.acm.org/doi/abs/10.1145/3383124.

Jevdjic:2017:ASC

[JSCM17] Djordje Jevdjic, Karin Strauss, Luis Ceze, and Henrique S. Malvar. Approximate storage of compressed and encrypted videos. Operating Systems Review, 51(2):361–373, June 2017. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Jacquet:2004:EHM

[JSS04] P. Jacquet, G. Seroussi, and W. Szpankowski. On the entropy of a hidden Markov process. In Storer and Cohn [SC04], pages 362–371. ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=1281481. IEEE

Computer Society Order Number: P2082.

Jansson:2015:LDT

[JSS15] Jesper Jansson, Kunihiko Sadakane, and Wing-Kin Sung. Linked dynamic tries with applications to LZ-compression in sublinear time and space. Algorithmica, 71(4):969–988, April 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541

(electronic). URL http://link.springer.com/article/10.
1007/s00453-013-9836-6.

Jacquet:2001:APL

[JST01] P. Jacquet, W. Szpankowski, and J. Tang. Average profile of the Lempel–Ziv parsing scheme for a Markovian source. Algorithmica, 31(3):318–360, July 2001. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=31&issue=3&spage=318.

Mathematical analysis of algorithms.

Jackson:1996:PAD

[JT96] David Jeff Jackson and Greg Scott Tinney. Performance analysis of distributed implementations of a fractal image compression algorithm. Concurrency: Practice and Experience, 8(5):357–386, June 1996. CODEN CPEXEI. ISSN

1040-3108. URL http://www3.interscience.wiley.com/
cgi-bin/abstract?ID=23284.

Jafarkhani:1998:SRT

[JT98] H. Jafarkhani and V. Tarokh. Successively refinable trellis coded quantization. In Storer and Cohn [SC98], pages 83–92. ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33

D232 1998. URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=672134. IEEE Computer Society

Press order number PR08406. IEEE order plan catalog number 98TB100225.

Jiang:2012:EPC

[JTC^{+}12] Wenfei Jiang, Jiang Tian, Kangying Cai, Fan Zhang, and Tao Luo. Efficient progressive compression of 3D points by maximizing tangent-plane continuity. In Storer and Marcellin [SM12], page 398. ISBN 1-4673-0715-7 (paperback). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=6189279. IEEE Computer Society order number P4656.

Jahnke:2010:SCM

[JU10] Tobias Jahnke and Tudor Udrescu. Solving chemical master equations by adaptive wavelet compression. Journal of Computational Physics, 229(16):5724–5741, August 10, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110001919.

Jurgen:1988:TSE

[Jur88] R. K. Jurgen. Technology ’88: the specialities [electrical engineering]. IEEE Spectrum, 25(1):70, January 1988. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1991:CDH

[Jur91] R. K. Jurgen. The challenges of digital HDTV. IEEE Spectrum, 28(4):28–30, April 1991. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1992:TCE

[Jur92] R. K. Jurgen. Technology 1992 — consumer electronics. IEEE Spectrum, 29(1):52–54, January 1992. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jin:2004:ESI

[JV04] T. Jin and J. Vaisey. Efficient side-information context description for context-based adaptive entropy coders. In Storer and Cohn [SC04], page ?? ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1281519. IEEE Computer Society Order Number: P2082.

Joung:1998:BCM

[JWK98] Hwayong Joung, E. K. Wong, and S. P. Kim. Block context modeling approach for binary image coding. In Storer and Cohn [SC98], page ?? ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1998. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=672289. IEEE

Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Jia:2003:SAC

[JYHC03] Yunwei Jia, En-Hui Yang, Da-Ke He, and S. Chan. Speeding up arithmetic coding using greedy re-normalization. In Storer and Cohn [SC03], page ?? ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2003. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1194051. IEEE Computer Society Order number PR01896.

Jiang:2012:CTI

[JZL^{+}12] Wenfei Jiang, Fan Zhang, L. J. Latecki, Zhibo Chen, and Yi Hu. Coefficient thresholding with image restoration. In Storer and Marcellin [SM12], page 399. ISBN 1-4673-0715-7 (paperback). ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=6189280. IEEE Computer Society

order number P4656.

Kedem:1995:ASH

[KA95] G. Kedem and T. Alexander. Application specific hardware compression of ray-casting data. In Storer and Cohn [SC95], page 484. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=515594. IEEE

catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Kleider:1998:ARC

[KA98] J. E. Kleider and G. P. Abousleman. Adaptive-rate coding modulation system for digital image transmission. In Storer and Cohn [SC98], page ?? ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1998.

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=672294. IEEE Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Kawanaka:1999:ZCW

[KA99] A. Kawanaka and V. R. Algazi. Zerotree coding of wavelet coefficients for image data on arbitrarily shaped support. In Storer and Cohn [SC99], page ?? ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1999. URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

785691. IEEE Computer Society Order Number PR00096. IEEE Order Plan Catalog Number PR00096.

Koc:2012:MPD

[KA12] B. Koc and Z. Arnavut. A modified pseudo-distance technique for lossless compression on color-mapped images. In Storer and Marcellin [SM12], page 401. ISBN 1-4673-0715-7 (paperback). ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=6189282. IEEE Computer Society

order number P4656.

Khalid:2011:WAH

[KAB11] Azra Khalid, Uzma Afsheen, and Saad Umer Baig. Wavelet analysis and high quality JPEG2000 compression using Daubechies wavelet. In International Conference on Graphic and Image Processing (ICGIP 2011). SPIE Optical Engineering Press, Bellingham, WA, USA, October 2011.

Kampf:1998:PFC

[Kam98] F. A. Kampf. Performance as a function of compression. IBM Journal of Research and Development, 42(6):759–766, ???? 1998. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556

(electronic). URL http://www.almaden.ibm.com/journal/
rd/426/kampf.html.

Karp:1961:MRC

[Kar61] R. S. Karp. Minimum-redundancy coding for the discrete noiseless channel. Transactions of the IRE, 7:27–38, 1961.

Karwowski:2012:IAA

[Kar12] Damian Karwowski. Improved adaptive arithmetic coding for HEVC video compression technology. Lecture Notes in Computer Science, 7594:121–128, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-33564-8_15/.

Kobayashi:1974:IDC

[KB74] H. Kobayashi and L. R. Bahl. Image data compression by predictive coding. I. prediction algorithms. IBM Journal of Research and Development, 18(2):164–171, March 1974. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Konstantinides:1992:MAI

[KB92] Konstantinos Konstantinides and Vasudev Bhaskaran. Monolithic architectures for image processing and compression. IEEE Computer Graphics and Applications, 12(6):75–86, November 1992. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kasperovich:1993:FDC

[KB93] L. V. Kasperovich and V. F. Babkin. Fast discrete cosine transform approximation for JPEG image compression. Lecture Notes in Computer Science, 719:98–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Knutsson:1999:AEE

[KB99] Björn Knutsson and Mats Björkman. Adaptive end-to-end compression for variable-bandwidth communication. Computer Networks (Amsterdam, Netherlands: 1999), 31(7):767– 779, April 8, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://www.elsevier.com/

cas/tree/store/comnet/sub/1999/31/7/2105.pdf.

Kopansky:2003:ESC

[KB03] A. Kopansky and M. Bystrom. Extending source codes for error resilience. In Storer and Cohn [SC03], pages 23–32. ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2003. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=1193993. IEEE

Computer Society Order number PR01896.

Kelly:2006:BCI

[KB06] B. Kelly and D. Brailsford. The B-coder: an improved binary arithmetic coder and probability estimator. In Storer and Cohn [SC06], page 456. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=1607299.

Klein:1989:STR

[KBD89] Shmuel T. Klein, Abraham Bookstein, and Scott Deerwester. Storing text retrieval systems on CD-ROM. compression and encryption considerations. ACM Transactions on Information Systems, 7(3):230–245, July 1989. CODEN ATISET. ISSN 1046-8188. URL http://www.acm.org:80. Special Issue on Research and Development in Information Retrieval.

Klein:2008:UFC

[KBN08] S. T. Klein and M. K. Ben-Nissan. Using Fibonacci compression codes as alternatives to dense codes. In Storer and Marcellin [SM08], pages 472–481. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33 INTERNET. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=4483325.

Ko:1991:DDC

[KC91] C. Ko and W. Chung. 2-D discrete cosine transform array processor using non-planar connections. In Storer and Reif [SR91b], page ?? ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1990. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=213305. IEEE

Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Kandadai:2006:PWA

[KC06] S. Kandadai and C. D. Creusere. Perceptually-weighted audio coding that scales to extremely low bitrates. In Storer and Cohn [SC06], pages 382–391. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=1607273.

Kawahara:1998:HSS

[KCB98] M. Kawahara, Yi-Jen Chiu, and T. Berger. High-speed software implementation of Huffman coding. In Storer and Cohn [SC98], page ?? ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1998. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=672291. IEEE

Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Ko:2021:LCL

[KCBM21] Yousun Ko, Alex Chadwick, Daniel Bates, and Robert Mullins. Lane compression: a lightweight lossless compression method for machine learning on embedded systems. ACM Transactions on Embedded Computing Systems, 20(2):16:1–16:26, March 2021. CODEN ???? ISSN 1539-9087 (print), 1558-

3465 (electronic). URL https://dl.acm.org/doi/10.1145/
3431815.

Kung:2017:CPD

[KCCW17] Sun-Yuan Kung, Thee Chanyaswad, J. Morris Chang, and Peiyuan Wu. Collaborative PCA/DCA learning methods for compressive privacy. ACM Transactions on Embedded Computing Systems, 16(3):76:1–76:??, July 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Kim:2006:GCD

[KCL06] Junho Kim, Sungyul Choe, and Seungyong Lee. Geometry compression and decompression: Multiresolution random accessible mesh compression. Computer Graphics Forum, 25(3): 323–331, September 2006. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kim:2018:EMS

[KCL18] Yong Hwi Kim, Junho Choi, and Kwan H. Lee. An efficient method for specular-enhanced BTF compression. Computers and Graphics, 75(??):A4, October 2018. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0097849318300918.

Kozintsev:1998:ITU

[KCR98] I. Kozintsev, J. Chou, and K. Ramchandran. Image transmission using arithmetic coding based continuous error detection. In Storer and Cohn [SC98], pages 339–348. ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1998. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=672162. IEEE Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Kossentini:1993:LBR

[KCS93] F. Kossentini, W. C. Chung, and M. J. T. Smith. Low bit rate coding of Earth science images. In Storer and Cohn [SC93], pages 371–380. ISBN 0-8186-3391-3 (microfiche), 0-8186-3392-1 (casebound). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1993. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=253112. IEEE

Computer Society Press order number 3392-02. IEEE catalog number 93TH0536-3.

Kong:2018:EVE

[KD18] Lingchao Kong and Rui Dai. Efficient video encoding for automatic video analysis in distributed wireless surveillance systems. ACM Transactions on Multimedia Computing, Communications, and Applications, 14(3):72:1–72:??, August 2018. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Knoll:2012:MLP

[KdF12] B. Knoll and N. de Freitas. A machine learning perspective on predictive coding with PAQ8. In Storer and Marcellin [SM12], pages 377–386. ISBN 1-4673-0715-7 (paperback). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=6189269. IEEE Computer Society order number P4656.

Kitakami:2011:ERM

[KE11] M. Kitakami and T. Ebihara. Error recovery method for PPM compressed data. In Storer and Marcellin [SM11b], page 462. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=5749519. IEEE

Computer Society Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Kelsey:2002:CIL

[Kel02] John Kelsey. Compression and information leakage of plaintext. Lecture Notes in Computer Science, 2365:263–276, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer-ny.com/link/
service/series/0558/bibs/2365/23650263.htm;   http:
//link.springer-ny.com/link/service/series/0558/papers/
2365/23650263.pdf.

Kendall:1961:CGD

[Ken61] Maurice G. Kendall. A Course in the Geometry of n-Dimensions. Hafner, New York, NY, USA, 1961. ?? pp.

Keshet:1991:NAT

[Kes91] Yehoshua Keshet. A new approach towards seismic data compression and automated classification of seismic attributes, with application to East Africa rift lakes data. Ph.D. thesis, Duke University, Durham, NC 27708, USA, 1991. 185 pp. URL

http://search.proquest.com/docview/303948606.

Kiselyov:1994:SSM

[KF94] O. Kiselyov and P. Fisher. Self-similarity of the multiresolutional image/video decomposition: smart expansion as compression of still and moving pictures. In Storer and Cohn [SC94], pages 331–340. ISBN 0-8186-5636-0, 0-8186-5637-9 (case). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1994. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=305941. IEEE

Catalog Number 93TH0626-2. IEEE Computer Society Press Order Number 5637-02.

Kiselyov:1995:MPL

[KF95] O. Kiselyov and P. Fisher. Multiresolutional piecewise-linear image decompositions: quantization error propagation and design of “stable” compression schemes. In Storer and Cohn [SC95], page 470. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

515580. IEEE catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Kiselyov:1996:ICI

[KF96] O. Kiselyov and P. Fisher. Image compression with iterated function systems, finite automata and zerotrees: grand unification. In Storer and Cohn [SC96], page ?? ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1996. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=488375. IEEE Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Karadimitriou:1997:ICM

[KF97] K. Karadimitriou and M. Fenstermacher. Image compression in medical image databases using set redundancy. In Storer and Cohn [SC97], page ?? ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1997. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=582104. IEEE

Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Kopylov:2002:CTC

[KF02] P. Kopylov and P. Franti. Context tree compression of multicomponent map images. In Storer and Cohn [SC02], pages 212– 221. ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2002. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=999959. IEEE

Computer Society Order Number PR01477.

Kopylov:2003:OLO

[KF03] P. Kopylov and P. Franti. Optimal layer ordering in the compression of map images. In Storer and Cohn [SC03], pages 323–332. ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2003. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1194023. IEEE Computer Society Order number PR01896.

Kaukoranta:1999:RCS

[KFN99] T. Kaukoranta, P. Franti, and O. Nevalainen. Reduced comparison search for the exact GLA. In Storer and Cohn [SC99], pages 33–41. ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1999. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=755651. IEEE

Computer Society Order Number PR00096. IEEE Order Plan Catalog Number PR00096.

KumaraSwamy:2004:RBF

[KFSM04] K. KumaraSwamy, C. Faloutsos, G. Shan, and V. Megalooikonomou. Relation between fractal dimension and performance of vector quantization. In Storer and Cohn [SC04], page ?? ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=1281523. IEEE

Computer Society Order Number: P2082.

Kim:1995:VBM

[KG95] S. P. Kim and X. Ginesta. VQ-based model design algorithms for text compression. In Storer and Cohn [SC95], page 434. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=515544. IEEE catalog

number 95TH8037. IEEE Computer Society Press order number PR07010.

Kari:1998:ICM

[KG98] J. Kari and M. Gavrilescu. Intensity controlled motion compensation. In Storer and Cohn [SC98], pages 249–258. ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1998. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=672153. IEEE Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Kirsanov:2003:ALC

[KG03] D. Kirsanov and S. J. Gortler. Arc-length compression. In Storer and Cohn [SC03], page ?? ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 2003. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1194055. IEEE Computer

Society Order number PR01896.

Kukunas:2014:HPZ

[KGG^{+}14] James T. Kukunas, Vinodh Gopal, Jim Guilford, Sean Gulley, Arjan van de Ven, and Wajdi Feghali. High performance ZLIB compression on Intel architecture processors. White paper 330343-002, Intel Corporation, ????, April 2014. 19 pp.

Kelner:2000:MDL

[KGK00] J. A. Kelner, V. K. Goyal, and J. Kovacevic. Multiple description lattice vector quantization: variations and extensions. In Storer and Cohn [SC00], pages 480–489. ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2000. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=838188. This millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Koyuturk:2006:NDB

[KGR06] Mehmet Koyutürk, Ananth Grama, and Naren Ramakrishnan. Nonorthogonal decomposition of binary matrices for boundederror data compression and analysis. ACM Transactions on Mathematical Software, 32(1):33–69, March 2006. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Khuri:2000:IPL

[KH00] Sami Khuri and Hsiu-Chin Hsu. Interactive packages for learning image compression algorithms. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32 (3):73–76, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Krunz:2004:ARC

[KH04] M. M. Krunz and M. Hassan. Adaptive rate control scheme for video streaming over wireless channels. In Storer and Cohn [SC04], pages 242–251. ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1281469. IEEE Computer Society Order Number: P2082.

Kim:2008:DRE

[KH08] A. N. Kim and F. Hekland. Dimension reduction and expansion: Distributed source coding in a noisy environment. In Storer and Marcellin [SM08], pages 332–341. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33 INTER-

NET. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=4483311.

Kammerl:2008:TAD

[KHCS08] J. Kammerl, P. Hinterseer, S. Chaudhuri, and E. Steinbach. A theoretical analysis of data reduction using the Weber quantizer. In Storer and Marcellin [SM08], page 524. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33 INTER-

NET. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=4483351.

Kim:2021:CCA

[KHHK21] Jinkwon Kim, Seokin Hong, Jeongkyu Hong, and Soontae Kim. CID: Co-architecting instruction cache and decompression system for embedded systems. IEEE Transactions on Computers, 70(7):1132–1145, July 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Klinc:2009:CDE

[KHJ^{+}09] D. Klinc, C. Hazay, A. Jagmohan, H. Krawczyk, and T. Rabin. On compression of data encrypted with block ciphers. In Storer and Marcellin [SM09], pages 213–222. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4976465. IEEE Computer Society

Order Number P3592. BMS Part Number CFP09DCC-PRT.

Kida:2009:STB

[Kid09] T. Kida. Suffix tree based VF-coding for compressed pattern matching. In Storer and Marcellin [SM09], page 449.

ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=4976503. IEEE Computer

Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

Kientzle:1999:AAI

[Kie99] Tim Kientzle. Algorithm alley: Implementing fast DCTs. Dr. Dobb’s Journal of Software Tools, 24(3):115–119, March 1999. CODEN DDJOEB. ISSN 1044-789X. URL

http://www.ddj.com/1999/9902/9902toc.htm;   http:/
/www.ddj.com/ftp/1999/1999_03/aa399.txt;   http://
www.ddj.com/ftp/1999/1999_03/aa399.zip.

Kiely:2004:SGP

[Kie04] A. Kiely. Selecting the Golomb parameter in Rice coding. Ipn (interplanetary network) progress report, ????, ????, November 15, 2004. 1–8 pp.

Kawashima:1995:UCD

[KIHO95] T. Kawashima, T. Igarashi, R. Hines, and M. Ogawa. A universal compressed data format for foreign file systems. In Storer and Cohn [SC95], page 429. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=515539. IEEE catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Kirovski:2004:PCC

[Kir04] D. Kirovski. Point compression for certificates of authenticity. In Storer and Cohn [SC04], page ?? ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1281521. IEEE Computer Society Order Number: P2082.

Kirovski:2005:PSC

[Kir05] D. Kirovski. A point-set compression heuristic for fiber-based certificates of authenticity. In Storer and Cohn [SC05], pages 103–112. ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=1402171.

Kjelsoe:1995:MMF

[KJ95] M. Kjelsoe and S. Jones. Memory management in flashmemory disks with data compression. Lecture Notes in Computer Science, 986:399–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kirovski:2003:F

[KJ03] D. Kirovski and N. Jojic. FaceCerts. In Storer and Cohn [SC03], page ?? ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2003. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1194054. IEEE Computer Society Order number PR01896.

Koyuncu:2010:SDQ

[KJ10] E. Koyuncu and H. Jafarkhani. A systematic distributed quantizer design method with an application to MIMO broadcast channels. In Storer and Marcellin [SM10b], pages 297–306. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=5453448.

Khalil:1999:CWP

[KJP99] H. Khalil, A. Jacquin, and C. Podilchuk. Constrained wavelet packets for tree-structured video coding algorithms. In Storer and Cohn [SC99], pages 354–363. ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1999. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=755685. IEEE Computer Society Or-

der Number PR00096. IEEE Order Plan Catalog Number PR00096.

Kiely:2001:ABE

[KK01] A. Kiely and M. Klimesh. An adaptable binary entropy coder. In Storer and Cohn [SC01c], pages 391–400. ISBN 0-7695-1031-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 2001. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=917170. IEEE CSP num-

ber 01PR1031.

Kaufman:2003:ECS

[KK03] Y. Kaufman and S. T. Klein. An efficient compression scheme supporting partial decoding. In Storer and Cohn [SC03], page ?? ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2003. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1194052. IEEE Computer Society Order number PR01896.

Kaufman:2004:STC

[KK04] Y. Kaufman and S. T. Klein. Semilossless text compression. In Storer and Cohn [SC04], page ?? ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1281520. IEEE Computer Society Order Number: P2082.

Kim:2007:NAD

[KK07] Il Ho Kim and H. J. Ko. A new approach to decoding of BCH and Reed–Solomon codes using syzygy. In Storer and Cohn [SC07], page 387. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=4148788. IEEE Computer Society Order Number P2791.

Komogortsev:2008:PRT

[KK08] Oleg V. Komogortsev and Javed I. Khan. Predictive real-time perceptual compression based on eye-gaze-position analysis. ACM Transactions on Multimedia Computing, Communications, and Applications, 4(3):23:1–23:??, August 2008. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Kim:2002:RCU

[KKA02] Hyun Mun Kim, Hyung-Suk Kim, and T. Acharya. Rate control using conditional mean estimator. In Storer and Cohn [SC02], page ?? ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2002. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1000001. IEEE Computer Society Order Number PR01477.

Keller:2014:GSC

[KKFL14] Orgad Keller, Tsvi Kopelowitz, Shir Landau Feibish, and Moshe Lewenstein. Generalized substring compression. Theoretical Computer Science, 525(??):42–54, March 13, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304397513007627.

Kim:2011:QLM

[KKJ11] Jong-Nam Kim, Won-Hee Kim, and Tae-Il Jung. Quassi lossless motion estimation algorithm using fast elimination of checking points. In Storer and Marcellin [SM11b], page 461. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=5749518. IEEE

Computer Society Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Kuhn:2016:DCC

[KKL16] Michael Kuhn, Julian Kunkel, and Thomas Ludwig. Data compression for climate data. Supercomputing Frontiers and Innovations, 3(1):75–94, ???? 2016. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (electronic). URL http://superfri.org/

superfri/article/view/101.

Kirovski:1999:PBP

[KKMS99] Darko Kirovski, Johnson Kin, and William H. Mangione-Smith. Procedure based program compression. International Journal of Parallel Programming, 27(6):457–475, December 1999. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=0885-7458&volume=27&issue=6&
spage=457.

Kavousianos:2007:OSH

[KKN07] X. Kavousianos, E. Kalligeros, and D. Nikolos. Optimal selective Huffman coding for test-data compression. IEEE Transactions on Computers, 56(8):1146–1152, August 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=4264327.

Karkkainen:2012:STB

[KKP12] J. Karkkainen, D. Kempa, and S. J. Puglisi. Slashing the time for BWT inversion. In Storer and Marcellin [SM12], pages 99–108. ISBN 1-4673-0715-7 (paperback). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ????

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=6189241. IEEE Computer Society order number P4656.

Karkkainen:2016:LLZ

[KKP16a] Juha Kärkkäinen, Dominik Kempa, and Simon J. Puglisi. Lazy Lempel–Ziv factorization algorithms. ACM Journal of Experimental Algorithmics, 21(1):2.4:1–2.4:??, November 2016. CO-DEN ???? ISSN 1084-6654.

Karkkainen:2017:LLZ

[KKP16b] Juha Kärkkäinen, Dominik Kempa, and Simon J. Puglisi. Lazy Lempel–Ziv factorization algorithms. ACM Journal of Experimental Algorithmics, 21(1):2.4:1–2.4:??, 2016. CODEN ???? ISSN 1084-6654.

Kaklamanis:2000:HAS

[KKS00] C. Kaklamanis, C. Konstantopoulos, and A. Svolos. A hypercube algorithm for sliding window compression. Parallel Processing Letters, 10(4):315–??, December 2000. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (elec-

tronic). URL http://ejournals.wspc.com.sg/ppl/10/
1004/S0129626400000305.html.

Kounchev:2010:PDTb

[KKT10] Ognyan Kounchev, Damyan Kalaglarsky, and Milcho Tsvetkov. Polyharmonic Daubechies type wavelets in image processing and astronomy, II. arXiv.org, ??(??):1–9, June 29, 2010. CO-DEN ???? ISSN 2331-8422. URL https://arxiv.org/abs/

1006.5739.

Kay:1995:GFF

[KL95] David C. Kay and John R. Levine. Graphics File Formats. Windcrest/McGraw-Hill, Blue Ridge Summit, PA, USA, second edition, 1995. ISBN 0-07-034025-0. xx + 476 pp. LCCN T385.K376 1995. US$29.95.

Kolarav:1997:CFD

[KL97] K. Kolarav and W. Lynch. Compression of functions defined on surfaces of 3D objects. In Storer and Cohn [SC97], pages 281–290. ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1997. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=582051. IEEE

Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Kim:2001:SPM

[KL01] Wonkoo Kim and Ching-Chung Li. A study on preconditioning multiwavelet systems for image compression. Lecture Notes in Computer Science, 2251:22–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer-ny.com/link/
service/series/0558/bibs/2251/22510022.htm;   http:
//link.springer-ny.com/link/service/series/0558/papers/
2251/22510022.pdf.

Kim:2003:RTP

[KL03] Junki Kim and Ho Suk Lee. Real-time preprocessing and video object segmentation for high compression and content-based MPEG-4 coding. In Storer and Cohn [SC03], page ?? ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2003. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=1194053. IEEE

Computer Society Order number PR01896.

Kirovski:2005:PAG

[KL05] D. Kirovski and Z. Landau. Parameter analysis for the generalized LZ compression of audio. In Storer and Cohn [SC05], page ?? ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=1402222.

Kato:2007:CLP

[KL07] M. Kato and Chia-Tien Dan Lo. Compression for low power consumption in battery-powered handsets. In Storer and Cohn [SC07], page 386. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

4148787. IEEE Computer Society Order Number P2791.

Kaser:2016:CBI

[KL16] Owen Kaser and Daniel Lemire. Compressed bitmap indexes: beyond unions and intersections. Software — Practice and Experience, 46(2):167–198, February 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Klasing:1998:ICCa

[Kla98a] R. Klasing. Improved compressions of cube-connected cycles networks. Lecture Notes in Computer Science, 1517:242–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klasing:1998:ICCb

[Kla98b] R. Klasing. Improved compressions of cube-connected cycles networks. IEEE Transactions on Parallel and Distributed Systems, 9(8):803–812, August 1998. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://dlib.

computer.org/td/books/td1998/pdf/l0803.pdf;   http:
//www.computer.org/tpds/td1998/l0803abs.htm.

Kolarov:1999:FCW

[KLAH99] K. Kolarov, W. Lynch, B. Arrighi, and B. Hoover. A fractional chip wavelet zero tree codec (WZT) for video compression. In Storer and Cohn [SC99], page ?? ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1999. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=785692. IEEE Computer Society Or-

der Number PR00096. IEEE Order Plan Catalog Number PR00096.

Klein:1989:ACM

[Kle89] Thomas Joseph Klein. Arithmetic coding and the move-to-front method: a study of two data compression schemes. M.S. thesis, University of Louisville, Louisville, KY, USA,

1989. 167 pp. URL http://search.proquest.com/docview/
303725202.

Klein:1997:EOR

[Kle97] S. T. Klein. Efficient optimal recompression. The Computer Journal, 40(2–3):117–126, ???? 1997. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/40/2_and_3/117.
full.pdf+html;   http://www.oup.co.uk/computer_journal/
Volume_40/Issue_02/Vol40_02.body.html#AbstractKlein;
 http://www3.oup.co.uk/computer_journal/Volume_40/
Issue_02/Vol40_03.body.html#AbstractKlein.

Klein:2007:FDF

[Kle07] S. T. Klein. Fast decoding of Fibonacci encoded texts. In Storer and Cohn [SC07], page 388. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=4148789. IEEE Computer Society Order Number P2791.

Klimesh:2000:OSD

[Kli00] M. Klimesh. Optimal subtractive dither for near-lossless compression. In Storer and Cohn [SC00], pages 223–232. ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2000. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=838162. This millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Kasirajan:2012:NDA

[KLJ12] Priya Kasirajan, Carl Larsen, and S. Jagannathan. A new data aggregation scheme via adaptive compression for wireless sensor networks. ACM Transactions on Sensor Networks, 9 (1):5:1–5:??, November 2012. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Kashyap:2005:DSC

[KLMXL05] A. Kashyap, L. A. Lastras-Montano, C. Xia, and Zhen Liu. Distributed source coding in dense sensor networks. In Storer and Cohn [SC05], pages 13–22. ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=1402162.

Kanai:1995:OCI

[KLR^{+}95] J. Kanai, S. Latifi, G. Rajarathinam, G. Nagy, and H. Bunke. Operations on compressed image data. In Storer and Cohn [SC95], page 432. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

515542. IEEE catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Kochanek:2008:MC

[KLUZ08] J. Kochanek, J. Lansky, P. Uzel, and M. Zemlicka. Multistream compression. In Storer and Marcellin [SM08], page 527. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33

INTERNET. URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4483354.

Kaplan:2007:SAB

[KLV07] Haim Kaplan, Shir Landau, and Elad Verbin. A simpler analysis of Burrows–Wheeler-based compression. Theoretical Computer Science, 387(3):220–235, November 22, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kwong:1995:SCA

[KM95] S. Kwong and K. F. Man. A speech coding algorithm based on predictive coding. In Storer and Cohn [SC95], page 455. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=515565. IEEE catalog

number 95TH8037. IEEE Computer Society Press order number PR07010.

Kosaraju:1997:SEB

[KM97a] S. Rao Kosaraju and G. Manzini. Some entropic bounds for Lempel–Ziv algorithms. In Storer and Cohn [SC97], page ?? ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1997. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=582106. IEEE

Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Kruse:1997:DCU

[KM97b] H. Kruse and A. Mukherjee. Data compression using text encryption. In Storer and Cohn [SC97], page ?? ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1997. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=582107. IEEE Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Kruse:1998:PTI

[KM98] H. Kruse and A. Mukherjee. Preprocessing text to improve compression ratios. In Storer and Cohn [SC98], page ?? ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1998. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=672295. IEEE Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Kosaraju:1999:CLE

[KM99a] S. Rao Kosaraju and Giovanni Manzini. Compression of low entropy strings with Lempel–Ziv algorithms. SIAM Journal on Computing, 29(3):893–911, June 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:/

/epubs.siam.org/sam-bin/dbq/article/33110.

Kruse:1999:ITC

[KM99b] H. Kruse and A. Mukherjee. Improving text compression ratios with the Burrows–Wheeler transform. In Storer and Cohn [SC99], page ?? ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1999. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=785693. IEEE

Computer Society Order Number PR00096. IEEE Order Plan Catalog Number PR00096.

Kieffer:2011:AOS

[KM11] J. C. Kieffer and J. Marcos. Advances in optimal structured source code design. In Storer and Marcellin [SM11b], pages 13–22. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=5749459. IEEE

Computer Society Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Kalpathy:2016:DPA

[KM16] Ravi Kalpathy and Hosam Mahmoud. Degree profile of m-ary search trees: a vehicle for data structure compression. Probability in the Engineering and Informational Sciences, 30(1): 113–123, January 2016. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/790B4495DD2902C2E6B5A66DAA03F5E7.

Klein:2020:DVD

[KMA^{+}20] M. Klein, A. Misra, B. Abali, P. Sethia, S. Weishaupt, B. Giamei, M. Farrell, and T. J. Slegel. Design and verification of DEFLATE acceleration as an architected instruction in z15. IBM Journal of Research and Development, 64(5/6):9:1–9:10, 2020. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Kamamoto:2010:LCP

[KMH10] Y. Kamamoto, T. Moriya, and N. Harada. Low-complexity PARCOR coefficient quantizer and prediction order estimator for G.711.0 (lossless speech coding). In Storer and Marcellin [SM10b], pages 475–483. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ????

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=
&arnumber=5453469.

Klonatos:2012:TOS

[KMM^{+}12] Yannis Klonatos, Thanos Makatos, Manolis Marazakis, Michail D. Flouris, and Angelos Bilas. Transparent online storage compression at the block-level. ACM Transactions on Storage, 8(2):5:1–5:??, May 2012. CODEN ???? ISSN 1553-3077 (print), 1553-3093 (electronic).

Korn:2002:RVG

[KMMV02] D. Korn, J. MacDonald, J. Mogul, and K. Vo. RFC 3284: The VCDIFF generic differencing and compression data format,

2002. URL ftp://ftp.internic.net/rfc/rfc3284.txt.

Kum:2005:RTM

[KMP05] Sang-Uok Kum and Ketan Mayer-Patel. Real-time multidepth stream compression. ACM Transactions on Multimedia Computing, Communications, and Applications, 1(2):128–150, May 2005. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Kossentini:1996:ACU

[KMS96] F. Kossentini, M. W. Macon, and M. J. T. Smith. Audio coding using variable-depth multistage quantizers. In Storer and Cohn [SC96], page ?? ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1996. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=488377. IEEE

Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Kozen:1998:EAO

[KMS98] D. Kozen, Y. Minsky, and B. Smith. Efficient algorithms for optimal video transmission. In Storer and Cohn [SC98], pages 229–238. ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1998. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=672151. IEEE

Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Kirchhoffer:2018:PDV

[KMSW18] Heiner Kirchhoffer, Detlev Marpe, Heiko Schwarz, and Thomas Wiegand. Properties and design of variable-to-variable length codes. ACM Transactions on Multimedia Computing, Communications, and Applications, 14(3):75:1–75:??, August 2018. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Karpovsky:1990:ORC

[KN90] M. G. Karpovsky and P. Nagvajara. Optimal robust compression of test responses. IEEE Transactions on Computers, 39 (1):138–141, January 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=46290.

Kashyap:1999:CDS

[KN99] N. Kashyap and D. L. Neuhoff. Codes for data synchronization with timing. In Storer and Cohn [SC99], pages 443–452. ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1999. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=755694. IEEE Computer Society Or-

der Number PR00096. IEEE Order Plan Catalog Number PR00096.

Karpinski:2005:ACO

[KN05] M. Karpinski and Y. Nekricht. Algorithms for construction of optimal and almost-optimal length-restricted codes. In Storer and Cohn [SC05], page ?? ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=1402221.

Kreft:2010:LLC

[KN10] S. Kreft and G. Navarro. LZ77-like compression with fast random access. In Storer and Marcellin [SM10b], pages 239–248. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=5453442.

Kinsman:2011:TOT

[KN11] Adam B. Kinsman and Nicola Nicolici. Trade-offs in test data compression and deterministic X-masking of responses. IEEE Transactions on Computers, 60(4):498–507, April 2011. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kim:2010:EWD

[KNB10] Y. Kim, M. S. Nadar, and A. Bilgin. Exploiting waveletdomain dependencies in compressed sensing. In Storer and Marcellin [SM10b], page 536. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ????

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=
&arnumber=5453515.

Knowlton:1980:PTG

[Kno80] Kenneth Knowlton. Progressive transmission of grey-scale and binary pictures by simple, efficient, and lossless encoding schemes. Proceedings of the IEEE, 68(7):885–896, July 1980. CODEN IEEPAD. ISSN 0018-9219 (print), 1558-2256 (electronic).

Knott:1998:AAA

[Kno98] Gary D. Knott. Algorithm alley: Adaptive delta modulation: Algorithms for audio compression. Dr. Dobb’s Journal of Software Tools, 23(4):124, 126, 128, 139, April 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://www.ddj.com/ftp/

1998/1998_04/aa498.txt.

Kumar:2008:IMC

[KNTG08] V. Santhosh Kumar, R. Nanjundiah, M. J. Thazhuthaveetil, and R. Govindarajan. Impact of message compression on the scalability of an atmospheric modeling application on clusters. Parallel Computing, 34(1):1–16, January 2008. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Knuth:1984:TB

[Knu84] Donald E. Knuth. The TEXbook, volume A of Computers and Typesetting. Addison-Wesley, Reading, MA, USA, 1984. ISBN 0-201-13447-0 (hardcover), 0-201-13448-9 (paperback). ix + 483 pp. LCCN Z253.4.T47 K58 1986. US$15.95 (paperback), US$32.95 (hardcover). Second printing, revised, October 1984. Sixth printing, revised, January 1986; also published as Computers & Typesetting, Vol. A. Twenty-fourth printing, summer 1993, contains final revisions. Twenty-first printing, June 1992.

Knuth:1985:DHC

[Knu85] Donald E. Knuth. Dynamic Huffman coding. Journal of Algorithms, 6(2):163–180, June 1985. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0196677485900367.

Knuth:1986:MP

[Knu86a]  Donald E. Knuth.  : The Program, volume D of

Computers and Typesetting. Addison-Wesley, Reading, MA, USA, 1986. ISBN 0-201-13438-1. xv + 560 pp. LCCN Z250.8.M46 K578 1986. Third printing, xviii + 566 pp., 1991.

Knuth:1986:TP

[Knu86b] Donald E. Knuth. TEX: The Program, volume B of Computers and Typesetting. Addison-Wesley, Reading, MA, USA, 1986. ISBN 0-201-13437-3. xv + 594 pp. LCCN Z253.4.T47 K578 1986. Reprinted with corrections May 1988. Fourth printing, xviii + 600 pp., 1991.

Knuth:1986:MB

[Knu86c]  Donald E. Knuth. The  book, volume C of Com-

puters and Typesetting. Addison-Wesley, Reading, MA, USA, 1986. ISBN 0-201-13445-4 (hardcover), 0-201-13444-6 (paperback). xi + 361 pp. LCCN Z250.8.M46 K58 1986.

Knuth:1986:CMT

[Knu86d] Donald E. Knuth. Computer Modern Typefaces, volume E of Computers and Typesetting. Addison-Wesley, Reading, MA, USA, 1986. ISBN 0-201-13446-2. xv + 588 pp. LCCN Z250.8.M46 K574 1986.

Knuth:1986:CT

[Knu86e] Donald Ervin Knuth. Computers and Typesetting. Addison-Wesley, Reading, MA, USA, 1986. ISBN 0-201-13447-0 (v.A), 0-201-13437-3 (v.B), 0-201-13445-4 (v.C), 0-201-13438-1 (v.D), 0-201-13446-2 (v.E). ?? pp. LCCN Z253.4.T47K58 1986.

Knutsson:1998:OGW

[Knu98] Lukas Knutsson. Optimized GIFs from Windows DIBs. C/C++ Users Journal, 16(12):??, December 1998. CODEN CCUJEX. ISSN 1075-2838.

Knuth:1973:ACP

[Knuta] Donald E. Knuth. The Art of Computer Programming, Volume 1, Fundamental Algorithms. Addison-Wesley, Reading, MA, USA, second edition, 1973. ISBN 0-201-03821-8. xxi + 634 pp.

Kieffer:1996:TQM

[KNY96] J. Kieffer, G. Nelson, and E-H. Yang. Tutorial on the quadrisection method and related methods for lossless data compres-

sion., 1996. URL http://www.ece.umn.edu/users/kieffer/
index.html.

Kohlhaas:1972:CFC

[Koh72] Charles Albert Kohlhaas. Calculation of Formation Compression From Core and Well-Tested Data. Ph.D. thesis, Colorado School of Mines, Golden, CO, USA, 1972. ???? pp. URL

http://search.proquest.com/docview/302554471.

Koloda:2012:SEC

[KØJ^{+}12] J. Koloda, J. Østergaard, S. H. Jensen, A. M. Peinado, and V. Sanchez. Sequential error concealment for video/ images by weighted template matching. In Storer and Marcellin [SM12], pages 159–168. ISBN 1-4673-0715-7 (paperback). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=6189247. IEEE Computer Society order number P4656.

Kominek:1995:CFE

[Kom95] J. Kominek. Convergence of fractal encoded images. In Storer and Cohn [SC95], pages 242–251. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=515514. IEEE catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Koman:1996:GAS

[Kom96] Richard Koman. GIF Animation Studio: Animating Your Web Site. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, October 1996. ISBN 1-56592-230-1. xi + 159 pp. LCCN TK5105.888 .K65 1996. US$39.95.

URL http://www.oreilly.com/catalog/9781565922303;
 http://www.oreilly.com/catalog/gif.

Kim:2009:FIP

[KOO09] Chanyul Kim, N. E. O’Connor, and YunJe Oh. Fast intra prediction in the transform domain. In Storer and Marcellin [SM09], page 450. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

4976504. IEEE Computer Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

Kopooshian:1997:DCE

[Kop97] Hrag H. Kopooshian. Data compression and encryption. Thesis (M.S.), California State University, Northridge, Northridge, CA, USA, 1997. vi + 127 pp.

Koppl:2021:RLZ

[Köp21] Dominik Köppl. Reversed Lempel–Ziv factorization with suffix trees. Algorithms (Basel), 14(6), June 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/14/6/161.

Kortman:1967:DCR

[Kor67] C. M. Kortman. Data compression by redundancy reduction. IEEE Spectrum, 4(3):133–139, March 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kochman:2008:NSP

[KØZ08] Y. Kochman, J. Østergaard, and R. Zamir. Noise-shaped predictive coding for multiple descriptions of a colored Gaussian source. In Storer and Marcellin [SM08], pages 362–371. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33 INTER-

NET. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=4483314.

Kiely:1995:SCM

[KP95a] A. Kiely and F. Pollara. Subband coding methods for seismic data compression. In Storer and Cohn [SC95], page 447. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=515557. IEEE catalog

number 95TH8037. IEEE Computer Society Press order number PR07010.

Kleijn:1995:SCS

[KP95b] W. B. Kleijn and K. K. Paliwal. Speech Coding and Synthesis. Elsevier, Amsterdam, The Netherlands, 1995. ?? pp.

Kim:1997:EWV

[KP97] Beong-Jo Kim and W. A. Pearlman. An embedded wavelet video coder using three-dimensional set partitioning in hierarchical trees (SPIHT). In Storer and Cohn [SC97], pages 251–260. ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1997. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=582048. IEEE

Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Kerkyacharian:2003:RCC

[KP03] Gérard Kerkyacharian and Dominique Picard. Replicant compression coding in Besov spaces. ESAIM: Probability and Statistics, 7(??):239–??, March 2003. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Krajcevski:2016:GGD

[KPM16] Pavel Krajcevski, Srihari Pratapa, and Dinesh Manocha. GST: GPU-decodable supercompressed textures. ACM Transactions on Graphics, 35(6):230:1–230:??, November 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kanakagiri:2017:MMD

[KPM17] Raghavendra Kanakagiri, Biswabandan Panda, and Madhu Mutyam. MBZip: Multiblock data compression. ACM Transactions on Architecture and Code Optimization, 14(4):42:1– 42:??, December 2017. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Kuroda:1996:FBM

[KPY96] Hideo Kuroda, Dan C. Popescu, and Hong Yan. Fast block matching method for image data compression based on fractal models. Proceedings of the SPIE — The International Society for Optical Engineering, 2501(??):1257–1266, ???? 1996. CODEN PSISDG. ISSN ????

Kess:1997:CGR

[KR97] B. L. Kess and S. E. Reichenbach. Capturing global redundancy to improve compression of large images. In Storer and Cohn [SC97], pages 62–71. ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1997. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=581967. IEEE Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Khalil:2001:RPV

[KR01] H. Khalil and K. Rose. Robust predictive vector quantizer design. In Storer and Cohn [SC01c], pages 33–42. ISBN 0-7695-1031-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 2001. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=917134. IEEE CSP num-

ber 01PR1031.

Kim:2010:BES

[KR10] A. N. Kim and T. A. Ramstad. Bandwidth expansion in a simple Gaussian sensor network using feedback. In Storer and Marcellin [SM10b], pages 259–268. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=5453444.

Kraft:1949:DQG

[Kra49] L. G. Kraft. A device for quantizing, grouping, and coding amplitude modulated pulses. Master’s thesis, Department of Electrical Engineering, MIT, Cambridge, MA, USA, 1949. ?? pp.

Kratzer:1991:NAD

[Kra91] K. P. Kratzer. A neural approach to data compression and classification. Lecture Notes in Computer Science, 541:250–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krause:2010:FFP

[Kra10] Philipp Klaus Krause. ftc — floating precision texture compression. Computers and Graphics, 34(5):594–601, October 2010. CODEN COGRD2. ISSN 0097-8493 (print), 1873-

7684 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0097849310000890.

Krivoulets:2002:CST

[Kri02] A. Krivoulets. On coding of sources with two-sided geometric distribution using binary decomposition. In Storer and Cohn [SC02], page ?? ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2002. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1000002. IEEE Computer Society Order Number PR01477.

Kim:2009:FPD

[KRK09] Jong-Nam Kim, Tae-Kyung Ryu, and Won-Hee Kim. A fast partial distortion elimination algorithm using dithering matching pattern. In Storer and Marcellin [SM09], page 451. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4976505. IEEE Computer Society

Order Number P3592. BMS Part Number CFP09DCC-PRT.

Kroeker:2000:PIM

[Kro00] Kirk L. Kroeker. Products: iCompression’s MPEG-2 encoder reference design kit for USB; free project management support tools; Force Computers’ protocol software for telecom integration; Data Junction’s free XML tool; Fookes Software’s text editor for programmers; Research Systems’ IDL wavelet toolkit; Be opens source code to BeOS 5; JafSoft’s utility for converting text files into RTF. Computer, 33(5):98–100, May 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814

(electronic). URL http://dlib.computer.org/co/books/
co2000/pdf/r5098.pdf.

Klein:2021:OSR

[KRSS21] Shmuel T. Klein, Jakub Radoszewski, Tamar C. Serebro, and Dana Shapira. Optimal skeleton and reduced Huffman trees. Theoretical Computer Science, 852(??):157–171, January 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print),

1879-2294 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304397520306605.

Korodi:2005:LDC

[KRT05] G. Korodi, J. Rissanen, and I. Tabus. Lossless data compression using optimal tree machines. In Storer and Cohn [SC05], pages 348–357. ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=
1402196.

Kruger:1992:BTC

[Kru92] Anton Kruger. Block truncation compression. Dr. Dobb’s Journal of Software Tools, 17(4):48, 50, 53–55, 104, 106, April 1992. CODEN DDJOEB. ISSN 1044-789X.

Korfhage:1993:PSA

[KRW^{+}93] Robert R. Korfhage, Edie M. Rasmussen, Peter Willett, et al., editors. Proceedings of the sixteenth Annual International ACM SIGIR Conference on Research and Development in Information Retrieval: Pittsburgh, PA, USA, June 27–July 1, 1993. ACM Press, New York, NY 10036, USA, 1993. ISBN 0-89791-605-0. LCCN ????

Kordes:1989:UMC

[KS89] F. L. G. Kordes and J. J. Schuurman. The use of MEBAS in creating a simulation environment for compression and encryption. Report NLR TP 89130 U, National Lucht-en Ruimtevaartlaboratorium, Amsterdam, The Netherlands, 1989. 69 pp.

Koch:1991:DCV

[KS91] E. Koch and M. Sommer. Data compression in view of data sealing. In Storer and Reif [SR91b], page ?? ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1990. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=213334. IEEE Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Kaerkkaeinen:1996:LZI

[KS96a] J. Kaerkkaeinen and E. Sutinen. Lempel–Ziv index for qgrams. Lecture Notes in Computer Science, 1136:378–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kontoyiannis:1996:SEE

[KS96b] I. Kontoyiannis and Y. M. Suhov. Stationary entrophy estimation via string matching. In Storer and Cohn [SC96], page ?? ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1996. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=488376. IEEE

Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Karkkainen:1998:LZI

[KS98] Juha Kärkkäinen and Erkki Sutinen. Lempel–Ziv index for qgrams. Algorithmica, 21(1):137–154, 1998. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). Fourth European Symposium on Algorithms (Barcelona, 1996).

Klein:2000:NCM

[KS00a] S. T. Klein and D. Shapira. A new compression method for compressed matching. In Storer and Cohn [SC00], pages 400–409. ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche).

ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 2000. URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=838180. This millen-

nial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Knessl:2000:ABH

[KS00b] Charles Knessl and Wojciech Szpankowski. Asymptotic behavior of the height in a digital search tree and the longest phrase of the Lempel–Ziv scheme. SIAM Journal on Computing, 30 (3):923–964, June 2000. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http://epubs.siam.

org/sam-bin/dbq/article/35681.

Klein:2001:PMH

[KS01] S. T. Klein and D. Shapira. Pattern matching in Huffman encoded texts. In Storer and Cohn [SC01c], pages 449–458. ISBN 0-7695-1031-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2001. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=917176. IEEE

CSP number 01PR1031.

Klein:2002:SCD

[KS02] S. T. Klein and D. Shapira. Searching in compressed dictionaries. In Storer and Cohn [SC02], pages 142–151. ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2002. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=999952. IEEE

Computer Society Order Number PR01477.

Klein:2005:CPM

[KS05] S. T. Klein and D. Shapira. Compressed pattern matching in JPEG images. In Storer and Cohn [SC05], page ?? ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1402223.

Klein:2006:CPM

[KS06a] Shmuel T. Klein and Dana Shapira. Compressed pattern matching in JPEG images. International Journal of Foundations of Computer Science (IJFCS), 17(6):1297–1306, December 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Krintz:2006:AFC

[KS06b] Chandra Krintz and Sezgin Sucu. Adaptive on-the-fly compression. IEEE Transactions on Parallel and Distributed Systems, 17(1):15–24, January 2006. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Klein:2007:CDE

[KS07] S. T. Klein and D. Shapira. Compressed delta encoding for LZSS encoded files. In Storer and Cohn [SC07], pages 113– 122. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=4148750. IEEE Com-

puter Society Order Number P2791.

Klein:2008:HCN

[KS08] S. T. Klein and D. Shapira. Huffman coding with nonsorted frequencies. In Storer and Marcellin [SM08], page 526. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33

INTERNET. URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4483353.

Klein:2011:SDM

[KS11a] S. T. Klein and D. Shapira. The string-to-dictionary matching problem. In Storer and Marcellin [SM11b], pages 143– 152. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=5749472. IEEE

Computer Society Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Klein:2011:CMD

[KS11b] Shmuel T. Klein and Dana Shapira. Compressed matching in dictionaries. Algorithms (Basel), 4(1):61–74, March 2011. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https:

//www.mdpi.com/1999-4893/4/1/61.

Klein:2016:CMF

[KS16] Shmuel T. Klein and Dana Shapira. Compressed matching for feature vectors. Theoretical Computer Science, 638(??):52–62, July 25, 2016. CODEN TCSCDI. ISSN 0304-3975 (print),

1879-2294 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304397515011822.

Kim:2016:BPC

[KSCE16] Jungrae Kim, Michael Sullivan, Esha Choukse, and Mattan Erez. Bit-plane compression: transforming data for better compression in many-core architectures. ACM SIGARCH Computer Architecture News, 44(3):329–340, June 2016. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Klein:2006:MDE

[KSS06] S. T. Klein, T. C. Serebro, and D. Shapira. Modeling delta encoding of compressed files. In Storer and Cohn [SC06], page 457. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=1607300.

Kovacevic:1995:ABT

[KSY95] J. Kovacevic, R. J. Safranek, and E. M. Yeh. Adaptive bidirectional time-recursive interpolation for deinterlacing. In Storer and Cohn [SC95], page 446. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=515556. IEEE catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Krichevsky:1981:PUC

[KT81] R. E. Krichevsky and V. K. Trofimov. The performance of universal coding. IEEE Transactions on Information Theory, IT-27:199–207, March 1981. CODEN IETTAW. ISSN 0018-9448 (print), 1557-9654 (electronic).

Kliewer:2002:CFO

[KT02] J. Kliewer and R. Thobaben. Combining FEC and optimal soft-input source decoding for the reliable transmission of correlated variable-length encoded signals. In Storer and Cohn [SC02], pages 83–91. ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2002. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=999946. IEEE Computer Society Order Number PR01477.

Korodi:2007:NML

[KT07a] G. Korodi and I. Tabus. Normalized maximum likelihood model of order-1 for the compression of DNA sequences. In Storer and Cohn [SC07], pages 33–42. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=4148742. IEEE Computer Society Order Number P2791.

Korodi:2007:CAN

[KT07b] Gergely Korodi and Ioan Tabus. Compression of annotated nucleotide sequences. IEEE/ACM Transactions on Computational Biology and Bioinformatics, 4(3):447–457, July 2007. CODEN ITCBCY. ISSN 1545-5963 (print), 1557-9964 (electronic).

Kitakami:2009:LIC

[KT09] M. Kitakami and K. Tai. Lossless image compression by PPMbased prediction coding. In Storer and Marcellin [SM09], page 452. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=4976506. IEEE

Computer Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

Khaled:2013:IAC

[KT13] Elleithy Khaled and Sobh Tarek, editors. Innovations and advances in computer, information, systems sciences, and engineering, volume 152 of Lecture notes in electrical engineering. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 1-4614-3534-X, 1-4614-3535-8 (e-book). LCCN ????

Kasai:2010:RCS

[KTMS10] K. Kasai, T. Tsujimoto, R. Matsumoto, and K. Sakaniwa. Rate-compatible Slepian–Wolf coding with short non-binary LDPC codes. In Storer and Marcellin [SM10b], pages 288– 296. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=5453449.

Kida:1998:MPM

[KTS^{+}98] T. Kida, M. Takeda, A. Shinohara, M. Miyazaki, and S. Arikawa. Multiple pattern matching in LZW compressed text. In Storer and Cohn [SC98], pages 103–112. ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1998. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=672136. IEEE Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Kida:1999:SAP

[KTSA99] Takuya Kida, Masayuki Takeda, Ayumi Shinohara, and Setsuo Arikawa. Shift-and approach to pattern matching in LZW compressed text. Lecture Notes in Computer Science, 1645: 1–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer-ny.com/
link/service/series/0558/bibs/1645/16450001.htm;
http://link.springer-ny.com/link/service/series/0558/
papers/1645/16450001.pdf.

Kumaran:1995:DWB

[KU95] Muthu Kumaran and Scott E. Umbaugh. A dynamic windowbased runlength coding algorithm applied to gray-level images. Graphical models and image processing: GMIP, 57 (4):267–282, July 1995. CODEN GMIPF4. ISSN 1077-

3169. URL http://www.idealibrary.com/links/artid/
gmip.1995.1025/production;   http://www.idealibrary.
com/links/artid/gmip.1995.1025/production/pdf.

Koshelevat:2004:MOB

[KUCV04] O. Koshelevat, B. Usevitch, S. Cabrera, and E. Vidal. MSE optimal bit rate allocation in the application of JPEG2000 part 2 to meteorological data. In Storer and Cohn [SC04], page ?? ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=1281522. IEEE

Computer Society Order Number: P2082.

Kulekci:2011:CCM

[Kül11] M. Oğuzhan Külekci. Compressed context modeling for text compression. In Storer and Marcellin [SM11b], pages 373– 382. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=5749495. IEEE

Computer Society Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Kulekci:2012:SBW

[Kül12] M. Oguzhan Külekci. On scrambling the Burrows–Wheeler transform to provide privacy in lossless compression. Computers & Security, 31(1):26–32, February 2012. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0167404811001416.

Kundu:1982:CLM

[Kun82] S. Kundu. On a class of linear maps for data compression. IEEE Transactions on Software Engineering, SE-8(5):530–532, September/October 1982. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?arnumber=1702981.

Kurien:1983:PDC

[Kur83] Thomas Kurien. A Partitioned Data Compression Algorithm. Ph.D. thesis, The University of Connecticut, Storrs, CT, USA,

1983. 167 pp. URL http://search.proquest.com/docview/
303265149.

Kutil:2002:OBA

[Kut02] R. Kutil. Optimization of bitstream assembly in parallel multimedia compression. PDCP: Parallel and Distributed Computing Practices, 5(3):??, September 2002. CODEN ???? ISSN 1097-2803.

Kim:2019:IED

[KV19] Jungwon Kim and Jeffrey S. Vetter. Implementing efficient data compression and encryption in a persistent key–value store for HPC. The International Journal of High Performance Computing Applications, 33(6):1098–1112, November 1, 2019. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (elec-

tronic). URL https://journals.sagepub.com/doi/full/
10.1177/1094342019847264.

Kosolobov:2020:LZL

[KVP20] Dmitry Kosolobov, Daniel Valenzuela, and Simon J. Puglisi. Lempel–Ziv-like parsing in small space. Algorithmica, 82(11): 3195–3215, November 2020. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL https://link.

springer.com/article/10.1007/s00453-020-00722-6.

Kinsner:1995:RSF

[KW95] W. Kinsner and L. Wall. Reduced-search fractal block coding of images. In Storer and Cohn [SC95], page 461. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=515571. IEEE catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Klein:2000:PHD

[KW00] S. T. Klein and Y. Wiseman. Parallel Huffman decoding. In Storer and Cohn [SC00], pages 383–392. ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2000. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=838178. This millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Klein:2001:PLZ

[KW01] Shmuel Tomi Klein and Yair Wiseman. Parallel Lempel–Ziv coding (extended abstract). Lecture Notes in Computer Science, 2089:18–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer-

ny.com/link/service/series/0558/bibs/2089/20890018.
htm;   http://link.springer-ny.com/link/service/series/
0558/papers/2089/20890018.pdf.

Klein:2003:PHD

[KW03] S. T. Klein and Y. Wiseman. Parallel Huffman decoding with applications to JPEG files. The Computer Journal, 46(5):487–497, September 2003. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://www3.oup.co.uk/computer_journal/hdb/Volume_
46/Issue_05/460487.sgm.abs.html;   http://www3.oup.

co.uk/computer_journal/hdb/Volume_46/Issue_05/pdf/

460487.pdf. This is an extended version of a paper that appeared earlier in Klein, S. T. and Wiseman, Y. (2000) Proc. Data Compression Conf. DCC’00, Snowbird, Utah, pp. 383– 392. IEEE Computer Society Press, Los Alamitos, CA.

Kim:2016:ECI

[KWC^{+}16] Dongwook Kim, Youjip Won, Jaehyuk Cha, Sungroh Yoon, Jongmoo Choi, and Sooyong Kang. Exploiting compressioninduced internal fragmentation for power-off recovery in SSD. IEEE Transactions on Computers, 65(6):1720–1733, ???? 2016. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Khansari:1992:CIC

[KWLG92] M. R. K. Khansari, I. Widjaja, and A. Leon-Garcia. Convolutional interpolative coding algorithms. In Storer and Cohn [SC92], pages 209–218. ISBN 0-8186-2717-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

227460. IEEE catalog number 91TH0436-6.

Kang:2012:SLC

[KXPZ12] Xiangui Kang, Xianyu Xu, Anjie Peng, and Wenjun Zeng. Scalable lossy compression for pixel-value encrypted images. In Storer and Marcellin [SM12], page 400. ISBN 1-4673-0715-7 (paperback). ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=6189281. IEEE Computer

Society order number P4656.

Kieffer:1999:STB

[KY99] J. C. Kieffer and E.-H. Yang. A simple technique for bounding the pointwise redundancy of the 1978 Lempel–Ziv algorithm. In Storer and Cohn [SC99], pages 434–442. ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1999. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=755693. IEEE Computer Society Or-

der Number PR00096. IEEE Order Plan Catalog Number PR00096.

Khalifeh:2008:OAT

[KY08] A. Khalifeh and H. Yousefi’zadeh. Optimal audio transmission over wireless tandem channels. In Storer and Marcellin [SM08], page 525. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008;

QA76.9.D33 INTERNET. URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=4483352.

Khalifeh:2010:HMT

[KY10] A. Khalifeh and H. Yousefi’zadeh. A hybrid media transmission scheme for wireless VoIP. In Storer and Marcellin [SM10b], pages 465–474. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=
5453466.

Kyle:1990:EIC

[Kyl90] Jim Kyle. ECOMP image compression. Computer Language Magazine, 7(12):78–??, December 1990. CODEN COMLEF. ISSN 0749-2839.

Kieffer:1996:LCB

[KYNC96] J. Kieffer, E-H. Yang, G. Nelson, and P. Cosman. Lossless compression via bisection trees, 1996. URL http://www.ece.

umn.edu/users/kieffer/index.html.

Kieffer:1998:CPM

[KYPY98] J. Kieffer, En-Hui Yang, T. Park, and S. Yakowitz. Complexity of preprocessor in MPM data compression system. In Storer and Cohn [SC98], page ?? ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1998. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=672292. IEEE Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Kochman:2002:APV

[KZ02] Y. Kochman and R. Zamir. Adaptive parametric vector quantization by natural type selection. In Storer and Cohn [SC02], pages 392–401. ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2002. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=999979. IEEE Computer Society Order Number PR01477.

Liebchen:2005:MAL

[L^{+}05] Tilman Liebchen et al. The MPEG-4 Audio Lossless Coding (ALS) Standard — technology and applications. In AES 119th Convention, New York, October 7–10, 2005, page ?? ACM Press, New York, NY 10036, USA, 2005. URL http:

//www.nue.tu-berlin.de/forschung/projekte/lossless/
mp4als.html. Available at http://www.nue.tu-berlin.de/
forschung/projekte/lossless/mp4als.html.

Lindsay:1991:RBV

[LA91] R. A. Lindsay and D. E. Abercrombie. Restricted boundary vector quantization. In Storer and Reif [SR91b], pages 159–165. ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1990. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=213383. IEEE

Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Ling:1996:HSM

[LA96] Nam Ling and R. Advani. A high speed motion estimator using 2-D log search algorithm. In Storer and Cohn [SC96], page ?? ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1996. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=488380. IEEE

Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Lin:2005:JSC

[LA05] Zihuai Lin and Tor Aulin. Joint source and channel coding using trellis coded CPM: soft decoding. In Storer and Cohn [SC05], pages 213–222. ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ????

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=
&arnumber=1402182.

Lenhardt:2012:GHS

[LA12] R. Lenhardt and J. Alakuijala. Gipfeli — high speed compression algorithm. In Storer and Marcellin [SM12], pages 109–118. ISBN 1-4673-0715-7 (paperback). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

6189242. IEEE Computer Society order number P4656.

Lee:2014:JEP

[LAB^{+}14] Jae-Gil Lee, Gopi Attaluri, Ronald Barber, Naresh Chainani, Oliver Draese, Frederick Ho, Stratos Idreos, Min-Soo Kim, Sam Lightstone, Guy Lohman, Konstantinos Morfonios, Keshava Murthy, Ippokratis Pandis, Lin Qiao, Vijayshankar Raman, Vincent Kulandai Samy, Richard Sidle, Knut Stolze, and Liping Zhang. Joins on encoded and partitioned data. Proceedings of the VLDB Endowment, 7(13):1355–1366, August 2014. CODEN ???? ISSN 2150-8097.

Lafe:2000:CAT

[Laf00] Olu Lafe. Cellular automata transforms: theory and applications in multimedia compression, encryption and modeling, volume MMSA16 of Multimedia systems and applications series. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2000. ISBN 0-7923-7857-1. xii + 177 pp. LCCN QA267.5.C45 L34 2000.

Langdon:1981:CBW

[LaJR81] Glen G. Langdon and and J. Rissanen. Compression of black white images with arithmetic coding. IEEE Transactions on Communications, COM-29(6):858–867, June 1981. CODEN IECMBT. ISSN 0090-6778 (print), 1558-0857 (electronic).

Lakhani:1997:EJI

[Lak97] G. Lakhani. Enhancements to the JPEG implementation of block smoothing method. In Storer and Cohn [SC97], page ?? ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1997. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=582108. IEEE

Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Lakhani:1998:AJQ

[Lak98] G. Lakhani. Adjustments for JPEG de-quantization coefficients. In Storer and Cohn [SC98], page ?? ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1998. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=672297. IEEE Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Lake:2000:PGM

[Lak00a] J. M. Lake. Prediction by grammatical match. In Storer and Cohn [SC00], pages 153–162. ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2000. URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

838155. This millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Lakhani:2000:MHC

[Lak00b] G. Lakhani. A modification to the Huffman coding of JPEG’s baseline compression algorithm. In Storer and Cohn [SC00], page ?? ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2000. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=838204. This

millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Lakhani:2002:UMH

[Lak02] G. Lakhani. Using multiple Huffman code tables for optimal coding of DCT blocks. In Storer and Cohn [SC02], page ?? ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2002. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=1000003. IEEE

Computer Society Order Number PR01477.

Lambert:1999:IAC

[Lam99] Sean M. Lambert. Implementing Associative Coder of Buyanovsky (ACB) data compression. M.S. thesis, Montana State University, Bozeman, MT, USA, 1999. ?? pp. URL

sum1els@mindless.com.

Langdon:1983:NZL

[Lan83a] Glen G. Langdon. A note on the Ziv–Lempel model for compressing individual sequences. IEEE Transactions on Information Theory, IT-29(2):284–287, March 1983. CODEN IET-TAW. ISSN 0018-9448 (print), 1557-9654 (electronic).

Langdon:1983:ARL

[Lan83b] Glenn G. Langdon. An adaptive run length coding algorithm. IBM Technical Disclosure Bulletin, 26(7B):3783–3785, December 1983. CODEN IBMTAA. ISSN 0018-8689.

Langdon:1984:PVM

[Lan84] Glen G. Langdon. On parsing vs. mixed-order model structures for data compression. IBM research report RJ-4163 (46091), IBM Corporation, San Jose, CA, USA, January 18, 1984. ?? pp.

Langdon:1991:PQC

[Lan91] G. G. Langdon, Jr. Probabilistic and Q-coder algorithms for binary source adaptation. In Storer and Reif [SR91b], pages 13–22. ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1990. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=213369. IEEE

Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Lanza:1998:EDC

[Lan98] P. Lanza. Entropy and data compression performances. In Storer and Cohn [SC98], page ?? ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1998. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=672298. IEEE Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Lansky:2009:U

[Lan09] Lansky. [unknown], 2009. URL http://www.ksi.mff.cuni.

cz/~lansky/SC/.

Lange:2010:AP

[Lan10] Kenneth Lange. Applied probability. Springer texts in statistics. Springer Science+Business Media, LLC, New York, NY, USA, second edition, 2010. ISBN 1-4419-7165-3, 1-4419-7164-5. xv + 436 pp. LCCN QA273 .L36 2010.

Li:2007:DHB

[LAP07] Hui Li, P. Agarwal, and B. Prabhakaran. Data hiding based compression mechanism for 3D models. In Storer and Cohn [SC07], page 391. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=4148792. IEEE Computer Society Order Number P2791.

Lima:2007:NFS

[LAPL07] L. Lima, D. Alfonso, L. Pezzoni, and R. Leonardi. New fast search algorithm for base layer of H.264 scalable video coding extension. In Storer and Cohn [SC07], page 393. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4148794. IEEE Computer Society

Order Number P2791.

Larsson:1996:EAS

[Lar96] N. J. Larsson. Extended application of suffix trees to data compression. In Storer and Cohn [SC96], pages 190–199. ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1996. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=488324. IEEE Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Larsson:1998:CTB

[Lar98] N. J. Larsson. The context trees of block sorting compression. In Storer and Cohn [SC98], pages 189–198. ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1998. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=672147. IEEE Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Larsson:1999:SSM

[Lar99] N. Jesper Larsson. Structures of string matching and data compression. Ph.D. thesis, Lunds Universitet, Lund, Sweden,

1999. 130 pp. URL http://search.proquest.com/docview/
304568808.

Lau:2009:RLH

[Lau09] Raymond Lau. Raymond Lau home page, 2009. URL http:/

/www.raylau.com/biography.html.

Lawrence:1977:NUC

[Law77] John C. Lawrence. A new universal coding scheme for the binary memoryless source. IEEE Transactions on Information Theory, 23(4):466–472, July 1977. CODEN IETTAW. ISSN 0018-9448 (print), 1557-9654 (electronic).

Lynch:1981:ADC

[LB81] Clifford A. Lynch and E. B. Brownrigg. Application of data compression to a large bibliographic data base. In Zaniolo and Delobel [ZD81], pages 435–447. LCCN QA76.9.D3

I559 1981. URL http://www.vldb.org/dblp/db/conf/vldb/

LynchB81.html. IEEE catalog no. 81CH1701-2.

Lamb:1990:CRT

[LB90] D. A. Lamb and D. T. Barnard. Compact representation of text files. Technical report, Queen’s University, Kingston, Ontario, March 1990. 1–7 pp.

LBE:2007:U

[LBE07] LBE. [unknown], 2007. URL http://in.geocities.com/

iamthebiggestone/how_lbe_works.htm.

Linde:1980:AVQ

[LBG80] Y. Linde, A. Buzo, and R. M. Gray. An algorithm for vector quantization design. IEEE Transactions on Communications, COM-28:84–95, January 1980. CODEN IECMBT. ISSN 0090-6778 (print), 1558-0857 (electronic).

Liang:1991:RDP

[LBH91] K. M. Liang, S. E. Budge, and R. W. Harris. Rate distortion performance of VQ and PVQ compression algorithms. In Storer and Reif [SR91b], page ?? ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1990. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=213316. IEEE Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Lemire:2016:SCI

[LBK16] Daniel Lemire, Leonid Boytsov, and Nathan Kurz. SIMD compression and the intersection of sorted integers. Software — Practice and Experience, 46(6):723–749, June 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lalgudi:2008:MDC

[LBMN08] H. G. Lalgudi, A. Bilgin, M. W. Marcellin, and M. S. Nadar. Multi-dimensional compression using JPEG2000. In Storer and Marcellin [SM08], page 528. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33 INTERNET.

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=
&arnumber=4483355.

Llados-Bernaus:1998:CAF

[LBS98] R. Llados-Bernaus and R. L. Stevenson. Codeword assignment for fixed-length entropy coded video streams. In Storer and Cohn [SC98], pages 269–278. ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1998. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=672155. IEEE Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Lu:1991:ASM

[LC91] Cheng-Chang Lu and Ngee-Boo Choong. Adaptive source modeling using conditioning tree with maximum entropy reduction criterions. In Storer and Reif [SR91b], page ?? ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1990. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=213338. IEEE Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Leontaris:2003:VCI

[LC03] A. Leontaris and P. C. Cosman. Video compression with intra/inter mode switching and a dual frame buffer. In Storer and Cohn [SC03], pages 63–72. ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 2003. URL http://ieeexplore.ieee.org/

stamp/stamp.jsp?tp=&arnumber=1193997. IEEE Computer

Society Order number PR01896.

Lavu:2003:EQG

[LCB03] S. Lavu, Hyeokio Choi, and R. Baraniuk. Estimationquantization geometry coding using normal meshes. In Storer and Cohn [SC03], pages 362–371. ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 2003. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1194027. IEEE Computer

Society Order number PR01896.

Chang:2009:BLC

[lCcYxFpW09] Wei ling Chang, Xiao chun Yun, Bin xing Fang, and Shu peng Wang. The block LZSS compression algorithm. In Storer and Marcellin [SM09], page 439. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=4976493. IEEE Computer Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

Liu:2011:JST

[LCG11] Jiaying Liu, Yongjin Cho, and Zongming Guo. Joint spatialtemporal layer bit allocation with S-domain dependent R-D modeling. In Storer and Marcellin [SM11b], page 468. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5749525. IEEE Computer Society

Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Li:2021:TRT

[LCJP21] Tianyi Li, Lu Chen, Christian S. Jensen, and Torben Bach Pedersen. TRACE: real-time compression of streaming trajectories in road networks. Proceedings of the VLDB Endowment, 14(7):1175–1187, March 2021. CODEN ???? ISSN 2150-

8097. URL https://dl.acm.org/doi/10.14778/3450980.
3450987.

Lin:2003:EDC

[LCL03] Chun-Yuan Lin, Yeh-Ching Chung, and Jen-Shiuh Liu. Efficient data compression methods for multidimensional sparse array operations based on the EKMR scheme. IEEE Transactions on Computers, 52(12):1640–1646, December 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=1252859.

Liu:2004:SWC

[LCLX04] Z. Liu, S. Cheng, A. D. Liveris, and Zixiang Xiong. Slepian– Wolf coded nested quantization (SWC-NQ) for Wyner–Ziv coding: performance analysis and code design. In Storer and Cohn [SC04], pages 322–331. ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ????

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=1281477. IEEE Computer Society Order Number: P2082.

Li:2003:TDC

[LCT03] Lei Li, Krishnendu Chakrabarty, and Nur A. Touba. Test data compression using dictionaries with selective entries and fixedlength indices. ACM Transactions on Design Automation of Electronic Systems (TODAES), 8(4):470–490, October 2003. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Lansky:2007:CTM

[LCV07] J. Lansky, K. Chernik, and Z. Vlckova. Comparison of text models for BWT. In Storer and Cohn [SC07], page 389. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=4148790. IEEE Computer

Society Order Number P2791.

Lin:2007:TED

[LCY07] Chin-Feng Lin, Yeh-Ching Chung, and Don-Lin Yang. TRLE — an efficient data compression scheme for image composition of volume rendering on distributed memory multicomputers. The Journal of Supercomputing, 39(3):321–345, March 2007. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=0920-8542&volume=39&issue=3&
spage=321.

Little:2010:OSM

[LD10] G. Little and J. Diamond. Optimum string match choices in LZSS. In Storer and Marcellin [SM10b], page 538. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5453493.

Liao:1999:TCB

[LDK99] Stan Liao, Srinivas Devadas, and Kurt Keutzer. A text-compression-based method for code size minimization in embedded systems. ACM Transactions on Design Automation of Electronic Systems (TODAES), 4(1): 12–38, January 1999. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL http://www.

acm.org/pubs/articles/journals/todaes/1999-4-1/p12-
liao/p12-liao.pdf;   http://www.acm.org/pubs/citations/
journals/todaes/1999-4-1/p12-liao/.

Leighton:2005:AGB

[LDM05] G. Leighton, J. Diamond, and T. Muldner. AXECHOP: a grammar-based compressor for XML. In Storer and Cohn [SC05], page ?? ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=
1402224.

Litow:1995:WFA

[LdV95] Bruce Litow and Olivier de Val. The weighted finite automaton inference problem. Technical Report 95-1, James Cook University, Brisbane, Queensland, 1995. ?? pp.

Luo:2020:STS

[LDZ^{+}20] Guoliang Luo, Zhigang Deng, Xin Zhao, Xiaogang Jin, Wei Zeng, Wenqiang Xie, and Hyewon Seo. Spatio-temporal segmentation based adaptive compression of dynamic mesh sequences. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(1):14:1–14:24, April 2020. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/doi/abs/10.1145/3377475.

LeGall:1991:MVC

[Le 91] Didier Le Gall. MPEG: a video compression standard for multimedia applications. Communications of the ACM, 34 (4):46–58, April 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.org/

pubs/toc/Abstracts/0001-0782/103090.html.

Lamparter:1994:ECC

[LE94] B. Lamparter and W. Effelsberg. eXtended color cell compression — a runtime-efficient compression scheme for software video. Lecture Notes in Computer Science, 868:181–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Logeswaran:2001:PSS

[LE01] R. Logeswaran and C. Eswaran. Performance survey of several lossless compression algorithms for telemetry applications. International Journal of Computer Applications, 23(1):1–9, 2001. CODEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.tandfonline.com/doi/full/10.
1080/1206212X.2001.11441623.

Logeswaran:2002:RBN

[LE02] R. Logeswaran and C. Eswaran. Radial basis neural network for lossless data compression. International Journal of Computer Applications, 24(1):14–19, 2002. ISSN 1206-212X (print), 1925-7074 (electronic). URL https://www.tandfonline.

com/doi/full/10.1080/1206212X.2002.11441655.

League:2007:TBC

[LE07] C. League and K. Eng. Type-based compression of XML data. In Storer and Cohn [SC07], pages 273–282. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4148766. IEEE Computer Society

Order Number P2791.

Lea:1978:TCA

[Lea78] R. M. Lea. Text compression with an associative parallel processor. The Computer Journal, 21(1):45–56, February 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/
content/21/1/45.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_21/Issue_01/tiff/45.
tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_21/Issue_01/tiff/46.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_21/Issue_01/tiff/
47.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_21/Issue_01/tiff/48.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_21/Issue_01/tiff/

49.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_21/Issue_01/tiff/50.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_21/Issue_01/tiff/
51.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_21/Issue_01/tiff/52.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_21/Issue_01/tiff/
53.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_21/Issue_01/tiff/54.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_21/Issue_01/tiff/
55.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_21/Issue_01/tiff/56.tif.

Lee:1981:DCS

[Lee81] Dong Hee Lee. Data Compression with Source Coding for Unknown Source Statistics. Ph.D. thesis, University of Maryland, College Park, MD, USA, 1981. 86 pp. URL http://search.

proquest.com/docview/303164012.

Lee:1991:PDC

[Lee91] E. T. Lee. Pictorial data compression using array grammars. In Storer and Reif [SR91b], page ?? ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1990. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=213317. IEEE Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Lee:1999:EAP

[Lee99] W. S. Lee. Edge-adaptive prediction for lossless image coding. In Storer and Cohn [SC99], pages 483–490. ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1999. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=755698. IEEE Computer Society Or-

der Number PR00096. IEEE Order Plan Catalog Number PR00096.

Lee:2000:TWT

[Lee00] Wee Sun Lee. Trees, windows and tiles for wavelet image compression. In Storer and Cohn [SC00], pages 283–292. ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2000. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=838168. This millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Lehman:2002:AAG

[Leh02] Eric Allen Lehman. Approximation algorithms for grammarbased data compression. Ph.D. thesis, Massachusetts Institute of Technology, Cambridge, MA, USA, 2002. ???? pp. URL

http://search.proquest.com/docview/305445788.

Lelewer:1991:DCM

[Lel91] Debra Ann Lelewer. Data compression on machines with limited memory. Ph.D. thesis, University of California, Irvine, Irvine, CA, USA, 1991. 99 pp. URL http://search.

proquest.com/docview/303921041.

Lesk:1988:GII

[Les88] Michael Lesk. GRAB — inverted indexes with low storage overhead. Computing systems: the journal of the USENIX Association, 1(3):207–220, Summer 1988. CODEN CMSYE2. ISSN 0895-6340.

Levinson:1947:WRE

[Lev47] N. Levinson. The Weiner RMS error criterion in filter design and prediction. Journal of Mathematical Physics, 25:261–278, 1947. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Levoy:1995:PAJ

[Lev95] Marc Levoy. Polygon-assisted JPEG and MPEG compression of synthetic images. Computer Graphics, 29 (Annual Conference Series):21–28, November 1995. CODEN CGRADI, CPGPBZ. ISSN 0097-8930 (print), 1558-4569 (elec-

tronic). URL http://www.acm.org:80/pubs/citations/
proceedings/graph/218380/p21-levoy/.

Levine:1996:SVQ

[Lev96] E. Levine. Stochastic vector quantization, and stochastic VQ with state feedback using neural networks. In Storer and Cohn [SC96], pages 330–339. ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1996. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=488338. IEEE Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Lewalle:1995:TCW

[Lew95] Jacques Lewalle. Tutorial on continuous wavelet analysis of experimental data, 1995. URL ftp.mame.syr.edu/

pub/jlewalle/tutor.ps.Z;   http://www.ecs.syr.edu/
faculty/lewalle/tutor/tutor.html.

Larsen:1995:CSI

[LF95] M. F. Larsen and R. L. Frost. Constraining the size of the instantaneous alphabet in trellis quantizers. In Storer and Cohn [SC95], pages 23–32. ISBN 0-8186-7012-

6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=515492. IEEE catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Lutai:2001:MSD

[LF01] Guan Lütai and Lu Feng. A method with scattered data spline and wavelets for image compression. Lecture Notes in Computer Science, 2251:49–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/0558/bibs/2251/
22510049.htm;   http://link.springer-ny.com/link/service/
series/0558/papers/2251/22510049.pdf.

Li:2003:AIC

[LF03] Xiuqi Li and Borko Furht. An approach to image compression using three-dimensional DCT. In Proceeding of The Sixth International Conference on Visual Information System 2003 (VIS2003), September 24–26, page ?? ????, ????, 2003.

Leu:2004:CDM

[LF04] Fang-Yie Leu and Yao-Chung Fan. Compressing a directed massive graph using small world model. In Storer and Cohn [SC04], page ?? ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

1281526. IEEE Computer Society Order Number: P2082.

Latendresse:2005:GFI

[LF05] Mario Latendresse and Marc Feeley. Generation of fast interpreters for Huffman compressed bytecode. Science of Computer Programming, 57(3):295–317, September 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Louis-Gavet:1978:DAI

[LG78] Guy Louis-Gavet. Diverses applications issues d’une fonction f de compactage basée sur une étude mathématique du langage naturel (compactage de données, comparaison de textes, hash-coding). [various applications issued from a compression function f based on a mathematical study of the natural language (data compression, comparison of texts, hashcoding)]. RAIRO Informatique/Computer Science (Revue Française d’Automatique, d’Informatique et de Recherche Operationnelle), 12(1):47–71, ???? 1978. CODEN RSINDN.

Lin:1999:VQV

[LG99] K. K. Lin and R. M. Gray. Vector quantization of video with two codebooks. In Storer and Cohn [SC99], page ?? ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1999. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=785694. IEEE Computer Society Or-

der Number PR00096. IEEE Order Plan Catalog Number PR00096.

Lin:2001:RDO

[LG01a] K. K. Lin and R. M. Gray. Rate-distortion optimization for the SPIHT encoder. In Storer and Cohn [SC01c], pages 123–132. ISBN 0-7695-1031-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2001. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=917143. IEEE CSP number 01PR1031.

Lin:2001:VRC

[LG01b] K. K. Lin and R. M. Gray. Video residual coding using SPIHT and dependent optimization. In Storer and Cohn [SC01c], pages 113–122. ISBN 0-7695-1031-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2001. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=917142. IEEE CSP number 01PR1031.

Gailly:2003:ZMS

[lGAR03a] Jean loup Gailly, Mark Adler, and Greg Roelofs. zlib: a massively spiffy yet delicately unobtrusive compression library (also free, not to mention unencumbered by patents), 2003.

URL http://www.gzip.org/zlib/.

Gailly:2003:ZTD

[lGAR03b] Jean loup Gailly, Mark Adler, and Greg Roelofs. zlib tech-

nical details, 2003. URL http://www.zlib.org/zlib_tech.
html.

Lippert:1997:CDV

[LGK97] L. Lippert, M. H. Gross, and C. Kurmann. Compression domain volume rendering for distributed environments. Computer Graphics Forum, 16(3):C95–C107, September 4– 8, 1997. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=146&amp; http://www.

blackwellpublishers.co.uk/asp/journal.asp?ref=0167-
7055&amp;   iid=3&amp; src=ard&amp; vid=16.

Lin:2005:ALT

[LGLK05] Song Lin, D. Gunopulos, S. Lonardi, and V. Kalogeraki. Applying LVQ techniques to compress historical information in sensor networks. In Storer and Cohn [SC05], page ?? ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1402225.

Lakshmi:2021:ECT

[LGM21] T. C. Subbu Lakshmi, D. Gnanadurai, and I. Muthulakshmi. Energy conserving texture-based adaptable compressive sensing scheme for WVSN. Concurrency and Computation: Practice and Experience, 33(3):e5178:1–e5178:??, February 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:1999:JIC

[LGO99] J. Li, R. M. Gray, and R. Olshen. Joint image compression and classification with vector quantization and a two dimensional hidden Markov model. In Storer and Cohn [SC99], pages 23–32. ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN QA76.9.D33 D37 1999. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=755650. IEEE

Computer Society Order Number PR00096. IEEE Order Plan Catalog Number PR00096.

leGuen:2000:BRL

[lGPL^{+}00] D. le Guen, S. Pateux, C. Labit, G. Moury, and D. Lebedeff. Bit rate and local quality control for on-board satellite image compression. In Storer and Cohn [SC00], page ?? ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2000. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=838205. This millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Langdon:1992:JML

[LGS92] G. Langdon, A. Gulati, and E. Seiler. On the JPEG model for lossless image compression. In Storer and Cohn [SC92], pages 172–180. ISBN 0-8186-2717-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=227464. IEEE

catalog number 91TH0436-6.

Lelewer:1987:DC

[LH87] Debra A. Lelewer and Daniel S. Hirschberg. Data compression. ACM Computing Surveys, 19(3):261–296, September 1987. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (elec-

tronic). URL http://www.acm.org/pubs/toc/Abstracts/
0360-0300/45074.html;   http://www.ics.uci.edu/~dan/
pubs/DataCompression.html. Reprinted in Japanese BIT

Special issue in Computer Science (1989), 16–195.

Lelewer:1991:SCM

[LH91] D. A. Lelewer and D. S. Hirschberg. Streamlining context models for data compression. In Storer and Reif [SR91b], pages 313–322. ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1990. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=213349. IEEE

Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Lin:2020:KPS

[LHD^{+}20] Weiyao Lin, Xiaoyi He, Wenrui Dai, John See, Tushar Shinde, Hongkai Xiong, and Lingyu Duan. Key-point sequence lossless compression for intelligent video analysis. IEEE MultiMedia, 27(3):12–22, 2020. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Lee:2003:AMM

[LHK03] J.-S. Lee, W.-K. Hong, and S.-D. Kim. Adaptive methods to minimize decompression overhead for compressed on-chip caches. International Journal of Computer Applications, 25 (2):98–105, 2003. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.tandfonline.com/doi/full/10.
1080/1206212X.2003.11441690.

Liu:2005:DBV

[LHS05] Hongmei Liu, Jiwu Huang, and Yun Q. Shi. DWT-based video data hiding robust to MPEG compression and frame loss. International Journal of Image and Graphics (IJIG), 5(1):111– ??, January 2005. CODEN ???? ISSN 0219-4678.

Lekatsas:2000:ACL

[LHW00] H. Lekatsas, J. Henkel, and W. Wolf. Arithmetic coding for low power embedded system design. In Storer and Cohn [SC00], pages 430–439. ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2000. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=838183. This

millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Lin:2012:FAH

[LHY12] Yih-Kai Lin, Shu-Chien Huang, and Cheng-Hsing Yang. A fast algorithm for Huffman decoding based on a recursion Huffman tree. The Journal of Systems and Software, 85(4):974– 980, April 2012. CODEN JSSODM. ISSN 0164-1212 (print),

1873-1228 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0164121211002925.

Li:1986:SIS

[Li86] Yuk Lun Li. A study of invariant speech characteristics as applied to phoneme extrapolation (speech recognition, data compression). M.E. thesis, The Cooper Union for the Advancement of Science and Art, New York, NY, USA, 1986. 176 pp. URL

http://search.proquest.com/docview/303473611.

Li:2002:SCA

[Li02a] J. Li. A source coding approach to classification by vector quantization and the principle of minimum description length. In Storer and Cohn [SC02], pages 382–391. ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2002. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=999978. IEEE

Computer Society Order Number PR01477.

Li:2002:ECP

[Li02b] Xin Li. Embedded coding of palette images in the topological space. In Storer and Cohn [SC02], page ?? ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 2002. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1000005. IEEE Computer

Society Order Number PR01477.

Li:2002:LBR

[Li02c] Xin Li. Low bit rate image coding in the scale space. In Storer and Cohn [SC02], pages 33–42. ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2002. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=999941. IEEE Computer Society Order Number PR01477.

Lindstrom:2008:LCH

[LI08] P. Lindstrom and M. Isenburg. Lossless compression of hexahedral meshes. In Storer and Marcellin [SM08], pages 192– 201. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33

INTERNET. URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4483297.

Li:2010:PAP

[Li10] Mengdong Li. Preimage awareness proofs of two compression functions. In Yang [Yan10], pages 660–664. ISBN 1-4244-6942-

2. LCCN QA76.9.A25. URL http://ieeexplore.ieee.org/
servlet/opac?punumber=5680738.

Liang:1984:WHP

[Lia84] Franklin Mark Liang. Word Hy-phen-a-tion by Com-pu-ter. Ph.D. dissertation, Computer Science Department, Stanford University, Stanford, CA, USA, March 1984. 92

pp. URL http://www.tug.org/docs/liang/;   http://
wwwlib.umi.com/dissertations/fullcit/8329742.

Liaw:1995:RGF

[Lia95] Wilson MacGyver Liaw. Reading GIF files. Dr. Dobb’s Journal of Software Tools, 20(2):56, 58, 60, 103, 106–107, February 1995. CODEN DDJOEB. ISSN 1044-789X.

Linde:1978:DTT

[Lin78] Yoseph Linde. The Design of Tree and Trellis Data Compression Systems. Ph.D. thesis, Stanford University, Stanford,

CA, USA, 1978. 142 pp. URL http://search.proquest.
com/docview/302929849.

Lin:1995:AIQ

[Lin95a] Jianhua Lin. Adaptive image quantization based on learning classifier systems. In Storer and Cohn [SC95], page 471. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=515587. IEEE catalog

number 95TH8037. IEEE Computer Society Press order number PR07010.

Lindley:1995:JLIa

[Lin95b] Craig A. Lindley. JPEG-like image compression. part 1. Dr. Dobb’s Journal of Software Tools, 20(7):50, 52, 54–58, July 1995. CODEN DDJOEB. ISSN 1044-789X.

Lindley:1995:JLIb

[Lin95c] Craig A. Lindley. JPEG-Like image compression. part 2. Dr. Dobb’s Journal of Software Tools, 20(8):62, 64–66, 102–105, August 1995. CODEN DDJOEB. ISSN 1044-789X.

Lin:1997:SAT

[Lin97] J. Lin. The search accuracy of tree-structured VQ. In Storer and Cohn [SC97], page ?? ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1997.

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=582112. IEEE Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Lindarto:1998:MCB

[Lin98] T. Lindarto. A multimode context-based lossless wavelet image coder. In Storer and Cohn [SC98], page ?? ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1998. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=672300. IEEE Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Lindstrom:2014:FRC

[Lin14] Peter Lindstrom. Fixed-rate compressed floating-point arrays. IEEE Transactions on Visualization and Computer Graphics, 20(12):2674–2683, December 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic),

2160-9306. URL http://csdl.computer.org/csdl/trans/
tg/2014/12/06876024-abs.html.

Liou:1991:OPK

[Lio91] Ming Liou. Overview of the p× 64 kbits/s video coding standard. Communications of the ACM, 34(4):59–63, April 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Little:1995:IEC

[Lit95] R. R. Little. An investigation of effective compression ratios for the proposed synchronous data compression proto. In Storer and Cohn [SC95], page 490. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=515597. IEEE catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Liu:1991:IUD

[Liu91] Chuang-Chun Liu. Identification and universal data compression of hidden Markov processes. Ph.D. thesis, University of Maryland, College Park, MD, USA, 1991. 76 pp. URL http:

//search.proquest.com/docview/303967206.

Liu:1997:NTS

[Liu97] C. P. Liu. A new two successive process image compression technique using subband coding and JPEG discrete cosine transform coding. Graphical models and image processing: GMIP, 59(3):179–191, May 1997. CODEN GMIPF4. ISSN 1077-3169. URL http://www.idealibrary.

com/links/artid/gmip.1997.0430/production;   http:/
/www.idealibrary.com/links/artid/gmip.1997.0430/production/
pdf.

Liu:2007:GQR

[Liu07] Xiteng Liu. Guided quaternary reaching method for waveletbased image compression. In Storer and Cohn [SC07], page

394. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=4148795. IEEE Com-

puter Society Order Number P2791.

Liu:2008:CDW

[Liu08a] Xiteng Liu. Composition of DCT and wavelet transform for image compression. In Storer and Marcellin [SM08], page

532. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33

INTERNET. URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4483359.

Liu:2008:MRR

[Liu08b] Xiteng Liu. Maximally robust redundant system with minimal coherence. In Storer and Marcellin [SM08], page 533. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33 INTER-

NET. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=4483360.

Liu:2010:NAT

[Liu10] Xiteng Liu. A new approach to time-frequency analysis. In Storer and Marcellin [SM10b], page 540. ISBN 0-7695-3994-

7. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=5453495.

Lee:2005:IEC

[LJ05] B. K. Lee and L. K. John. Implications of executing compression and encryption applications on general purpose processors. IEEE Transactions on Computers, 54(7):917–922, July 2005. CODEN ITCOB4. ISSN 0018-9340 (print),

1557-9956 (electronic). URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1432674.

Lexa:2007:JOD

[LJ07] M. A. Lexa and D. H. Johnson. Joint optimization of distributed broadcast quantization systems for classification. In Storer and Cohn [SC07], pages 363–374. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=4148775. IEEE Computer Society Order Number P2791.

Li:2019:IID

[LJF19] Rongjia Li, Chenhui Jin, and Ruya Fan. Improved integral distinguishers on compression function of GOST R hash function. The Computer Journal, 62(4):535–544, April 2019. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://academic.oup.com/comjnl/article/
62/4/535/5224765.

Liu:2021:DBF

[LJPE21] Chunwei Liu, Hao Jiang, John Paparrizos, and Aaron J. Elmore. Decomposed bounded floats for fast compression and queries. Proceedings of the VLDB Endowment, 14(11):2586– 2598, July 2021. CODEN ???? ISSN 2150-8097. URL https:

//dl.acm.org/doi/10.14778/3476249.3476305.

Liu:2008:PAD

[LJZ^{+}08] Da Liu, Xiangyang Ji, Debin Zhao, Xiaobin Zhu, Zhi Bian, and Wen Gao. Performance analysis of dual frame motion compensation. In Storer and Marcellin [SM08], page 531. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33 INTER-

NET. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=4483358.

Lewis:1991:KVC

[LK91] A. S. Lewis and G. Knowles. A 64 Kb/s video codec using the 2-D wavelet transform. In Storer and Reif [SR91b], pages 196–201. ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1990. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=213362. IEEE

Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Lam:1996:ESC

[LK96a] Wai-Man Lam and Sanjeev R. Kulkarni. Extended synchronizing codewords for binary prefix codes. IEEE Transactions on Information Theory, 42(3):984–987, May 1996. CODEN IETTAW. ISSN 0018-9448 (print), 1557-9654 (electronic).

Langi:1996:CAO

[LK96b] A. Langi and W. Kinsner. Compression of aerial ortho images based on image denoising. In Storer and Cohn [SC96], page ?? ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1996. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=488378. IEEE

Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Lee:2008:VHV

[LK08] J. B. Lee and H. Kalva. The VC-1 and H.264 Video Compression Standards for Broadband Video Services. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. ?? pp.

Lynch:1999:FMC

[LKA99] W. Lynch, K. Kolarov, and B. Arrighi. Fast, modified Z-coding of wavelet pyramids. In Storer and Cohn [SC99], page ?? ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1999. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=785695. IEEE Computer Society Or-

der Number PR00096. IEEE Order Plan Catalog Number PR00096.

Lozano:2002:QFE

[LKA02] A. C. Lozano, J. Kovacevic, and M. Andrews. Quantized frame expansions in a wireless environment. In Storer and Cohn [SC02], pages 232–241. ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2002. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=999961. IEEE Computer Society Order Number PR01477.

Lemire:2012:RRB

[LKG12] Daniel Lemire, Owen Kaser, and Eduardo Gutarra. Reordering rows for better compression: Beyond the lexicographic order. ACM Transactions on Database Systems, 37(3):20:1– 20:??, August 2012. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Li:2000:DPR

[LKH^{+}00] A. H. Li, S. Kittitornkun, Yu-Hen Hu, Dong-Seek Park, and J. Villasenor. Data partitioning and reversible variable length codes for robust video communications. In Storer and Cohn [SC00], pages 460–469. ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2000. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=838186. This

millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Lee:2017:IEE

[LKK^{+}17] S. Lee, K. Kim, G. Koo, H. Jeon, M. Annavaram, and W. W. Ro. Improving energy efficiency of GPUs through data compression and compressed execution. IEEE Transactions on Computers, 66(5):834–847, May 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lu:2012:SBM

[LKR12] Weizhi Lu, K. Kpalma, and J. Ronsin. Sparse binary matrices of LDPC codes for compressed sensing. In Storer and Marcellin [SM12], page 405. ISBN 1-4673-0715-7 (paperback). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=6189286. IEEE Computer Society order number P4656.

Lemire:2017:SVF

[LKR17] Daniel Lemire, Nathan Kurz, and Christoph Rupp. Stream VByte: Faster byte-oriented integer compression. Information Processing Letters, 130(??):1–6, February 2017. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0020019017301679.

Lee:1991:DIC

[LKRR91] H. Lee, Y. Kim, A. H. Rowberg, and E. A. Riskin. 3-D image compression for X-ray CT images using displacement estimation. In Storer and Reif [SR91b], page ?? ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1990. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=213308. IEEE Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Ling:1996:CDT

[LL96] Nam Ling and Jui-Rua Li. A codebook design technique for better image quality in vector quantization. In Storer and Cohn [SC96], page ?? ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1996. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=488379. IEEE

Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Lefebvre:2002:CLL

[LL02] Arnaud Lefebvre and Thierry Lecroq. Compror: On-line lossless data compression with a factor oracle. Information Processing Letters, 83(1):1–6, July 16, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lenzerini:2008:PTS

[LL08] Maurizio Lenzerini and Domenico Lembo, editors. Proceedings of the Twenty-Seventh ACM SIGMOD-SIGACT-SIGART Symposium on Principles of Database Systems: PODS’08, Vancouver, BC, Canada, June 9–11, 2008. ACM Press, New York, NY 10036, USA, 2008. ISBN 1-605-60932-3. LCCN ????

Li:2016:TLL

[LL16] Yimei Li and Yao Liang. Temporal lossless and lossy compression in wireless sensor networks. ACM Transactions on Sensor Networks, 12(4):37:1–37:??, November 2016. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Lee:2008:IWB

[LLA08] Jungwon Lee, Teahyung Lee, and D. V. Anderson. Improved wavelet-based embedded image coding using a dynamic index reordering vector quantizer. In Storer and Marcellin [SM08], page 530. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33

INTERNET. URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4483357.

Lee:1995:PVQ

[LLCC95] H. J. Lee, J. C. Liu, A. K. Chan, and C. K. Chui. A parallel vector quantization algorithm for SIMD multiprocessor systems. In Storer and Cohn [SC95], page 479. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=515589. IEEE catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Liu:2011:VEI

[LLCC11] Shaoli Liu, Ling Li, Yunji Chen, and Tianshi Chen. Video encoding without integer-pel motion estimation. In Storer and Marcellin [SM11b], page 469. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ????

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=5749526. IEEE Computer Society Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Llewellyn:1987:DCS

[Lle87] J. A. Llewellyn. Data compression for a source with Markov characteristics. The Computer Journal, 30(2):149–156, April 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/
content/30/2/149.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_02/tiff/
149.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_02/tiff/150.tif;   http://www3.

oup.co.uk/computer_journal/hdb/Volume_30/Issue_02/
tiff/151.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_02/tiff/152.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_30/Issue_02/
tiff/153.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_02/tiff/154.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_30/Issue_02/
tiff/155.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_02/tiff/156.tif.

Luneau:2009:CWM

[LLJ09] J.-M. Luneau, J. Lebrun, and S. H. Jensen. Complex wavelet modulation subbands for speech compression. In Storer and Marcellin [SM09], page 457. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=4976511. IEEE Computer Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

Lai:2012:IVS

[LLLZ12] Yi Lai, Xuguang Lan, Yuehu Liu, and Nanning Zheng. Improved view synthesis with depth reliability maps. In Storer and Marcellin [SM12], page 402. ISBN 1-4673-0715-7 (paperback). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=6189283. IEEE Computer Society order number P4656.

Loeffler:1989:PFD

[LLM89] C. Loeffler, A. Ligtenberg, and G. Moschytz. Practical fast 1-D DCT algorithms with 11 multiplications. In Proceedings of the International Conference on Acoustics, Speech, and Signal Processing (ICASSP ’89), pages 988–991. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1989.

Lan:2004:SWC

[LLN^{+}04] C.-F. Lan, A. D. Liveris, K. Narayanan, Zixiang Xiong, and C. Georghiades. Slepian–Wolf coding of multiple M-ary sources using LDPC codes. In Storer and Cohn [SC04], page ?? ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1281525. IEEE Computer Society

Order Number: P2082.

Lloyd:1982:LSQ

[Llo82] S. P. Lloyd. Least squares quantization in PCM. IEEE Transactions on Information Theory, IT-28:129–137, March 1982. CODEN IETTAW. ISSN 0018-9448 (print), 1557-9654 (electronic).

Loew:1996:APM

[LLP96] Murray H. Loew, Dunling Li, and Raymond L. Pickholtz. Adaptive PIFS model in fractal image compression. Proceedings of the SPIE — The International Society for Optical Engineering, 2707(??):284–293, ???? 1996. CODEN PSISDG. ISSN ????

Lavoue:2019:PET

[LLPP19] Guillaume Lavoue, Michael Langer, Adrien Peytavie, and Pierre Poulin. A psychophysical evaluation of texture compression masking effects. IEEE Transactions on Visualization and Computer Graphics, 25(2):1336–1346, February 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (elec-

tronic), 2160-9306. URL https://www.computer.org/csdl/
trans/tg/2019/02/08292880-abs.html.

Li:2021:TAE

[LLS^{+}21] Yang Li, Guangcan Liu, Yubao Sun, Qingshan Liu, and Shengyong Chen. 3D tensor auto-encoder with application to video compression. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(2):48:1–48:18, June 2021. CODEN ???? ISSN 1551-6857 (print), 1551-

6865 (electronic). URL https://dl.acm.org/doi/10.1145/
3431768.

Liao:2016:LCG

[LLT^{+}16] Xiaofei Liao, Li Lin, Guang Tan, Hai Jin, Xiaobin Yang, Wei Zhang, and Bo Li. LiveRender: a cloud gaming system based on compressed graphics streaming. IEEE/ACM Transactions on Networking, 24(4):2128–2139, August 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Liu:2011:FBC

[LLW11] Xiao Lin Liu, Chong Luo, and Feng Wu. Formulating binary compressive sensing decoding with asymmetrical property. In Storer and Marcellin [SM11b], pages 213–222. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5749479. IEEE Computer Society

Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Laney:2014:AED

[LLW^{+}14] Daniel Laney, Steven Langer, Christopher Weber, Peter Lindstrom, and Al Wegener. Assessing the effects of data compression in simulations using physically motivated metrics. Scientific Programming, 22(2):141–155, ???? 2014. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Lu:2008:LCE

[LLWC08] Qin Lu, Wusheng Luo, Jidong Wang, and Bo Chen. Lowcomplexity and energy efficient image compression scheme for wireless sensor networks. Computer Networks (Amsterdam, Netherlands: 1999), 52(13):2594–2603, September 17, 2008. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Lin:2014:SSE

[LLWD14] Xing Lin, Yebin Liu, Jiamin Wu, and Qionghai Dai. Spatialspectral encoded compressive hyperspectral imaging. ACM Transactions on Graphics, 33(6):233:1–233:??, November 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lu:2000:PLI

[LLY00] Tianyu Lu, Zisheng Le, and D. Y. Y. Yun. Piecewise linear image coding using surface triangulation and geometric compression. In Storer and Cohn [SC00], pages 410–419. ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2000. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=838181. This millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Luo:2014:ICH

[LLYH14] Da Luo, Weiqi Luo, Rui Yang, and Jiwu Huang. Identifying compression history of wave audio and its applications. ACM Transactions on Multimedia Computing, Communications, and Applications, 10(3):30:1–30:??, April 2014. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Linder:1993:URC

[LLZ93] T. Linder, G. Lugosi, and K. Zeger. Universality and rates of convergence in lossy source coding. In Storer and Cohn [SC93], pages 89–97. ISBN 0-8186-3391-3 (microfiche), 0-8186-3392-1 (casebound). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1993. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=253141. IEEE

Computer Society Press order number 3392-02. IEEE catalog number 93TH0536-3.

Linder:1996:DVQ

[LLZ96] T. Linder, G. Lugosi, and K. Zeger. Designing vector quantizers in the presence of source noise or channel noise. In Storer and Cohn [SC96], pages 33–42. ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1996. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=488308. IEEE Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Lai:2009:PRM

[LLZ09] Zhongyuan Lai, Wenyu Liu, and Yuan Zhang. Perceptual relevance measure for generic shape coding. In Storer and Marcellin [SM09], page 453. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ????

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=4976507. IEEE Computer Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

Liu:1991:OFA

[LM91] Y. Liu and H. Ma. ω-orbit finite automata for data compression. In Storer and Reif [SR91b], pages 166–175. ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1990. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=213384. IEEE Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Larsson:1999:ODB

[LM99] N. J. Larsson and A. Moffat. Offline dictionary-based compression. In Storer and Cohn [SC99], pages 296–305. ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1999. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=755679. IEEE Computer Society Or-

der Number PR00096. IEEE Order Plan Catalog Number PR00096.

Larsson:2000:LDB

[LM00] Larsson and Moffat. Off-line dictionary-based compression. Proceedings of the IEEE, 88(11):1722–1732, 2000. CO-DEN IEEPAD. ISSN 0018-9219 (print), 1558-2256 (elec-

tronic). URL http://www.bic.kyoto-u.ac.jp/pathway/
rwan/software/restore.html. An earlier, shorter version

was published Proceedings of the Conference on Data Compression 1999, pages 296–305.

Liddell:2001:LRC

[LM01] M. Liddell and A. Moffat. Length-restricted coding in static and dynamic frameworks. In Storer and Cohn [SC01c], pages 133–142. ISBN 0-7695-1031-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2001. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=917144. IEEE CSP number 01PR1031.

Liddell:2002:ICM

[LM02] M. Liddell and A. Moffat. Incremental calculation of minimum-redundancy length-restricted codes. In Storer and Cohn [SC02], pages 182–191. ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2002. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=999956. IEEE Computer Society Order Number PR01477.

Liddell:2003:HPC

[LM03] M. Liddell and A. Moffat. Hybrid prefix codes for practical use. In Storer and Cohn [SC03], pages 392–401. ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 2003. URL http://ieeexplore.ieee.org/

stamp/stamp.jsp?tp=&arnumber=1194030. IEEE Computer

Society Order number PR01896.

Liu:2008:RDO

[LM08] Hui Liu and Siliang Ma. R–D optimized tree-structured compression algorithms with discrete directional wavelet transform. Journal of Computational and Applied Mathematics, 219(1):302–311, September 15, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0377042707004116.

Lalgudi:2008:LBV

[LMBN08] H. G. Lalgudi, M. W. Marcellin, A. Bilgin, and M. S. Nadar. Lifting-based view compensated compression of volume rendered images for efficient remote visualization. In Storer and Marcellin [SM08], pages 442–451. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33 INTERNET.

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=
&arnumber=4483322.

Lastras-Montano:2005:NTG

[LMC05] L. A. Lastras-Montano and V. Castelli. Near tightness of the El Gamal and Cover region for two descriptions. In Storer and Cohn [SC05], pages 3–12. ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=1402161.

Liu:2009:CIR

[LMH^{+}09] Yanwei Liu, Siwei Ma, Qingming Huang, Debin Zhao, Wen Gao, and Nan Zhang. Compression-induced rendering distortion analysis for texture/depth rate allocation in 3D video compression. In Storer and Marcellin [SM09], pages 352–361. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=4976479. IEEE Computer

Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

Lin:2021:EBL

[LMJ^{+}21] Xuelian Lin, Shuai Ma, Jiahao Jiang, Yanchen Hou, and Tianyu Wo. Error bounded line simplification algorithms for trajectory compression: an experimental evaluation. ACM Transactions on Database Systems, 46(3):11:1–11:44, September 2021. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-

4644 (electronic). URL https://dl.acm.org/doi/10.1145/
3474373.

Lohrey:2011:TSC

[LMM11] M. Lohrey, S. Maneth, and R. Mennicke. Tree structure compression with RePair. In Storer and Marcellin [SM11b], pages 353–362. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=5749493. IEEE

Computer Society Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Liang:1997:RBV

[LMO97] J. Liang, I. Moccagatta, and K. Oehler. Region-based video coding with embedded zero-trees. In Storer and Cohn [SC97], page ?? ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1997. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=582110. IEEE

Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Linares:1994:EBT

[LMS94] I. Linares, R. M. Mersereau, and M. J. T. Smith. Enhancement of block transform coded images using residual spectra adaptive postfiltering. In Storer and Cohn [SC94], pages 321–330. ISBN 0-8186-5636-0, 0-8186-5637-9 (case). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1994. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=305940. IEEE

Catalog Number 93TH0626-2. IEEE Computer Society Press Order Number 5637-02.

Lee:2018:IEE

[LMSE18] Hochan Lee, Mansureh S. Moghaddam, Dongkwan Suh, and Bernhard Egger. Improving energy efficiency of coarse-grain reconfigurable arrays through modulo schedule compression/ decompression. ACM Transactions on Architecture and Code Optimization, 15(1):1:1–1:??, April 2018. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Lemmerling:2003:EIS

[LMV03] Philippe Lemmerling, Nicola Mastronardi, and Sabine Van Huffel. Efficient implementation of a structured total least squares based speech compression method. Linear Algebra and its Applications, 366(C):295–315, June 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-

1856 (electronic). URL http://www.elsevier.nl/gej-
ng/10/30/19/215/25/42/abstract.html;   http://www.
sciencedirect.com/science?_ob=GatewayURL&_origin=SOCJLA&
_urlversion=4&_method=citationSearch&_version=1&_piikey=
S0024379502004652&_volkey=00243795%23366%23295&_refkey=
Lemmerling%232003%23295%23315&md5=dba27dbbdd79bc42830964db1b0f2c94.

Special issue on Structured Matrices: Analysis, Algorithms and Applications.

Lelewer:1991:HCD

[LN91] Debra A. Lelewer and Cheng Ng. An honors course in data compression. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23(1):146–150, March 1991. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Li:2002:LCI

[LN02] Fei Li and J. Nieh. Low-complexity interpolation coding for server-based computing. In Storer and Cohn [SC02], page ?? ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2002. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=1000004. IEEE

Computer Society Order Number PR01477.

L:2017:BDW

[LN17] Biji C. L. and Achuthsankar S. Nair. Benchmark dataset for whole genome sequence compression. IEEE/ACM Transactions on Computational Biology and Bioinformatics, 14(6): 1228–1236, November 2017. CODEN ITCBCY. ISSN 1545-5963 (print), 1557-9964 (electronic).

Lu:1999:PJS

[LNA99] J. Lu, A. Nosratinia, and B. Aazhang. Progressive joint sourcechannel coding in feedback channels. In Storer and Cohn [SC99], pages 140–148. ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1999. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=755663. IEEE

Computer Society Order Number PR00096. IEEE Order Plan Catalog Number PR00096.

Liu:2016:ILL

[LNhCW16] Wei Jun Liu, Ge Nong, Wai hong Chan, and Yi Wu. Improving a lightweight LZ77 computation algorithm for running faster. Software — Practice and Experience, 46(9):1201–1217, September 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Long:1997:TCA

[LNV97] P. M. Long, A. I. Natsev, and J. S. Vitter. Text compression via alphabet re-representation. In Storer and Cohn [SC97], pages 161–170. ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1997. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=582003. IEEE

Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Lengwehasatit:1999:CDT

[LO99] K. Lengwehasatit and A. Ortega. Complexity-distortion tradeoffs in vector matching based on probabilistic partial distance techniques. In Storer and Cohn [SC99], pages 394–403. ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1999. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=755689. IEEE Computer Society Or-

der Number PR00096. IEEE Order Plan Catalog Number PR00096.

Logeswaran:2004:FTS

[Log04] R. Logeswaran. Fast two-stage Lempel–Ziv lossless numeric telemetry data compression using a neural network predictor. J.UCS: The Journal of Universal Computer Science, 10(9): 1199–1211, September 28, 2004. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http://www.jucs.

org/jucs_10_9/fast_two_stage_lempel.

Logeswaran:2006:FTN

[Log06] R. Logeswaran. Fault tolerant neural predictors for compression of sensor telemetry data. J.UCS: The Journal of Universal Computer Science, 12(10):1439–1454, ???? 2006. CODEN ???? ISSN 0948-695X (print), 0948-6968 (elec-

tronic). URL http://www.jucs.org/jucs_12_10/fault_
tolerant_neural_predictors.

Lopez-Ortiz:2005:FSS

[LOMSS05] Alejandro López-Ortiz, Mehdi Mirzazadeh, Mohammad Ali Safari, and Hossein Sheikhattar. Fast string sorting using order-preserving compression. ACM Journal of Experimental Algorithmics, 10:1.4:1–1.4:??, ???? 2005. CODEN ???? ISSN 1084-6654.

Lopes:2012:ICS

[Lop12] Dorsey Lopes. Important concepts in signal processing, image processing and data compression. University Publications, Delhi, India, 2012. ISBN 81-323-3604-6. ???? pp. LCCN ????

Lasch:2020:FSS

[LOS20] Robert Lasch, Ismail Oukid, and Kai-Uwe Sattler. Faster & strong: string dictionary compression using sampling and fast vectorized decompression. VLDB Journal: Very Large Data Bases, 29(6):1263–1285, November 2020. CO-DEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

tronic). URL https://link.springer.com/article/10.
1007/s00778-020-00620-x.

Ling:1975:BOI

[LP75] H. Ling and F. P. Palermo. Block-oriented information compression. IBM Journal of Research and Development, 19(2): 141–145, March 1975. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Larmore:1995:CHT

[LP95] Lawrence L. Larmore and Teresa M. Przytycka. Constructing Huffman trees in parallel. SIAM Journal on Computing, 24(6): 1163–1169, December 1995. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http://epubs.

siam.org/sam-bin/dbq/article/23324.

Liu:2007:FDW

[LP07] Ying Liu and W. A. Pearlman. Four-dimensional wavelet compression of 4-D medical images using scalable 4-D SBHP. In Storer and Cohn [SC07], pages 233–242. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=4148762. IEEE Computer Society Order Number P2791.

Liu:2010:GDH

[LPG10] Yang Liu, Balakrishnan Prabhakaran, and Xiaohu Guo. Geometry I: Dirichlet harmonic shape compression with feature preservation for parameterized surfaces. Computer Graphics Forum, 29(7):2039–2048, September 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lin:1997:OJC

[LPR97] C. C. Lin, D. J. Pease, and R. R. Raje. An optimal-joint-coordinate block matching algorithm for motion-compensated coding. In Storer and Cohn [SC97], page ?? ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1997. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=582111. IEEE Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Loia:2002:FRC

[LPS02] Vincenzo Loia, Witold Pedrycz, and Salvatore Sessa. Fuzzy relation calculus in the compression and decompression of fuzzy relations. International Journal of Image and Graphics (IJIG), 2(4):617–??, October 2002. CODEN ???? ISSN 0219-4678.

Liu:2015:GTB

[LPZJ15] Shaohui Liu, Anand Paul, Guochao Zhang, and Gwanggil Jeon. A game theory-based block image compression method in encryption domain. The Journal of Supercomputing, 71 (9):3353–3372, September 2015. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http://link.

springer.com/article/10.1007/s11227-015-1413-0.

Liebchen:2004:MES

[LR04] T. Liebchen and Y. A. Reznik. MPEG-4 ALS: an emerging standard for lossless audio coding. In Storer and Cohn [SC04], pages 439–448. ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

1281489. IEEE Computer Society Order Number: P2082.

Liu:2021:LRD

[LRH^{+}21] Xinqian Liu, Jiadong Ren, Haitao He, Qian Wang, and Chen Song. Low-rate DDoS attacks detection method using data compression and behavior divergence measurement. Computers & Security, 100(??):Article 102107, January 2021. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167404820303801.

Lopez-Rubio:2001:DTN

[LRMPGR01] Ezequiel López-Rubio, José Muñoz-Pérez, and José Antonio Gómez-Ruiz. Dynamic topology networks for colour image compression. Lecture Notes in Computer Science, 2085:168–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer-ny.com/link/
service/series/0558/bibs/2085/20850168.htm;   http:
//link.springer-ny.com/link/service/series/0558/papers/
2085/20850168.pdf.

LoPresto:1997:ICB

[LRO97] S. M. LoPresto, K. Ramchandran, and M. T. Orchard. Image coding based on mixture modeling of wavelet coefficients and a fast estimation-quantization framework. In Storer and Cohn [SC97], pages 221–230. ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1997. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=582045. IEEE Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Leung:1991:ICE

[LS91] Wai-Hong Leung and S. S. Skiena. Inducing codes from examples. In Storer and Reif [SR91b], pages 267–276. ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1990. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=213354. IEEE Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Lin:1992:IST

[LS92] J. Lin and J. A. Storer. Improving search for tree-structured vector quantization. In Storer and Cohn [SC92], pages 339– 348. ISBN 0-8186-2717-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=227447. IEEE catalog

number 91TH0436-6.

Lin:1993:DPT

[LS93] J. Lin and J. A. Storer. Design and performance of treestructured vector quantizers. In Storer and Cohn [SC93], pages 292–301. ISBN 0-8186-3391-3 (microfiche), 0-8186-3392-1 (casebound). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1993. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=253120. IEEE

Computer Society Press order number 3392-02. IEEE catalog number 93TH0536-3.

Luczak:1994:LDC

[LS94] T. Luczak and W. Szpankowski. A lossy data compression based on string matching: Preliminary analysis and suboptimal algorithms. Lecture Notes in Computer Science, 807: 102–112, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lawler:1995:AEC

[LS95a] E. L. Lawler and S. Sarkissian. Adaptive error correcting codes based on cooperative play of the game of “twenty questions with a liar”. In Storer and Cohn [SC95], page 464. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=515574. IEEE catalog

number 95TH8037. IEEE Computer Society Press order number PR07010.

Louchard:1995:GLZ

[LS95b] G. Louchard and W. Szpankowski. Generalized Lempel–Ziv parsing scheme and its preliminary analysis of the average profile. In Storer and Cohn [SC95], pages 262–271. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=515516. IEEE catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Louchard:1996:ARR

[LS96] G. Louchard and W. Szpankowski. Average redundancy rate of the Lempel–Ziv code. In Storer and Cohn [SC96], pages 92–101. ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1996. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=488314. IEEE

Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Liefke:1999:XEC

[LS99] Hartmut Liefke and Dan Suciu. XMill: an efficient compressor for XML data. In Proceedings of the ACM SIGMOD Symposium on the Management of Data, 2000, pages 153–164. ACM Press, New York, NY 10036, USA, 1999. URL http:

//citeseer.ist.psu.edu/327443.html.

Lonardi:2003:JSC

[LS03] S. Lonardi and W. Szpankowski. Joint source-channel LZ’77 coding. In Storer and Cohn [SC03], pages 273–282. ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2003. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=1194018. IEEE

Computer Society Order number PR01896.

Lakhani:2004:UPM

[LS04] Gopal Lakhani and Radhakrishnan Sethurama. Using partialmatching approach with Sequitur for contact-based coding. In Storer and Cohn [SC04], page ?? ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1281524. IEEE Computer Society Order Number: P2082.

Lin:1991:COT

[LSC91] J. Lin, J. A. Storer, and M. Cohn. On the complexity of optimal tree pruning for source coding. In Storer and Reif [SR91b], pages 63–72. ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1990. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=213363. IEEE

Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Lin:2010:PDW

[LSC10] Ting-Lan Lin, Jihyun Shin, and P. Cosman. Packet dropping for widely varying bit reduction rates using a networkbased packet loss visibility model. In Storer and Marcellin [SM10b], pages 445–454. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=
5453488.

Li:1997:VDC

[LSD97] Hua Harry Li, Shan Sun, and Haluk Derin, editors. Video data compression for multimedia computing: statistically based and biologically inspired techniques, volume SECS 378 of The Kluwer international series in engineering and computer science; Multimedia systems and applications. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1997. ISBN 0-7923-9790-8. xvi + 421 pp. LCCN

QA76.575 .V52 1997. URL http://www.loc.gov/catdir/
enhancements/fy0820/96047147-d.html;   http://www.
loc.gov/catdir/enhancements/fy0820/96047147-t.html.

Lee:2012:CSR

[LSH12] S. S. Lee, M. Shishibori, and C. Y. Han. Compression of search range of VP-tree for multimedia data retrieval applications. In Storer and Marcellin [SM12], page 403. ISBN 1-4673-0715-7 (paperback). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=6189284. IEEE Com-

puter Society order number P4656.

Louchard:1999:APG

[LST99] Guy Louchard, Wojciech Szpankowski, and Jing Tang. Average profile of the generalized digital search tree and the generalized Lempel–Ziv algorithm. SIAM Journal on Computing, 28(3):904–934, June 1999. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http://epubs.siam.

org/sam-bin/dbq/article/30181.

Liu:2008:IPE

[LSW08] Dong Liu, Xiaoyan Sun, and Feng Wu. Intra prediction via edge-based inpainting. In Storer and Marcellin [SM08], pages 282–291. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33

INTERNET. URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4483306.

Liu:2009:IIW

[LSW09] Dong Liu, Xiaoyan Sun, and Feng Wu. Improving inverse wavelet transform by compressive sensing decoding with deconvolution. In Storer and Marcellin [SM09], page 455. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4976509. IEEE Computer Society

Order Number P3592. BMS Part Number CFP09DCC-PRT. Lemire:2016:CFS

[LSYKK16] Daniel Lemire, Gregory Ssi-Yan-Kai, and Owen Kaser. Consistently faster and smaller compressed bitmaps with Roaring. Software — Practice and Experience, 46(11):1547–1569, November 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lu:2003:KSC

[LT03] Yue Lu and Chew Lim Tan. Keyword searching in compressed document images. In Storer and Cohn [SC03], page ?? ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2003. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=1194056. IEEE

Computer Society Order number PR01896.

Leontaris:2008:DCI

[LT08a] A. Leontaris and A. M. Tourapis. Drift characterization of intra prediction and quantization in H.264. In Storer and Marcellin [SM08], pages 212–221. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33 INTERNET. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=4483299.

Li:2008:ETI

[LT08b] Zhen Li and A. M. Tourapis. An estimation-theoretic interpretation of video rate distortion optimization with Lagrangian formulation. In Storer and Marcellin [SM08], pages 222–231. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33

INTERNET. URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4483300.

Lowell:2009:NFN

[LT09] D. Lowell and D. E. Tamir. New families and new members of integer sequence based coding methods. In Storer and Marcellin [SM09], page 456. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=4976510. IEEE Computer Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

Li:2004:SWC

[LTB04] J. Li, Z. Tu, and R. S. Blum. Slepian–Wolf coding for nonuniform sources using turbo codes. In Storer and Cohn [SC04], pages 312–321. ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

1281476. IEEE Computer Society Order Number: P2082.

Lechner:2011:SGC

[LTO11] G. Lechner, R. Timo, and L. Ong. Sparse graph codes for the two-way relay network with correlated sources. In Storer and Marcellin [SM11b], page 466. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=5749523. IEEE Computer Society Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Lagana:1991:ECD

[LTZ91] M. R. Lagana, G. Turrini, and G. Zanchi. Experiments on the compression of dictionary entries. In Storer and Reif [SR91b], page ?? ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1990. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=213329. IEEE

Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Lu:1992:DCS

[Lu92] Wei-Wei Lu. Data compression for space applications. D.Phil. thesis, University of Sussex, Brighton, UK, 1992. 158 pp. URL

http://search.proquest.com/docview/304001463.

Lucente:1996:CHB

[Luc96] M. Lucente. Computational holographic bandwidth compression. IBM Systems Journal, 35(3&4):349–366, 1996. CODEN IBMSA7. ISSN 0018-8670.

Lucco:2000:SSD

[Luc00] Steven Lucco. Split-stream dictionary program compression. ACM SIGPLAN Notices, 35(5):27–34, May 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://delivery.acm.

org/10.1145/350000/349307/p27-lucco.pdf;   http://
www.acm.org/pubs/articles/proceedings/pldi/349299/
p27-lucco/p27-lucco.pdf;   http://www.acm.org/pubs/
citations/proceedings/pldi/349299/p27-lucco/.

Lumban:2013:RPD

[Lum13] Gaol Ford Lumban, editor. Recent progress in data engineering and Internet technology. Volume 1, volume 156 of Lecture notes in electrical engineering. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 3-642-28806-5, 3-642-28807-3 (e-book). LCCN ????

Luse:1994:BFF

[Lus94] Marv Luse. The BMP file format. Dr. Dobb’s Journal of Software Tools, 19(10):18–20, 22, 82, 84, 85, September 1994. CODEN DDJOEB. ISSN 1044-789X.

Luther:1988:YTC

[Lut88a] A. C. Luther. You are there … and in control (digital interactive video). IEEE Spectrum, 25(9):45–50, September 1988. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Luther:1988:YTL

[Lut88b] A. C. Luther. You are there … and in control (digital interactive video). IEEE Spectrum, 25(9):45–50, September 1988. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lin:1992:NOV

[LV92] Jyh-Han Lin and J. S. Vitter. Nearly optimal vector quantization via linear programming. In Storer and Cohn [SC92], pages 22–31. ISBN 0-8186-2717-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=227479. IEEE catalog

number 91TH0436-6.

Lin:1995:PAF

[LV95] H. Lin and A. N. Venetsanopoulos. A pyramid algorithm for fast fractal image compression. In Proceedings ICIP-95 (IEEE International Conference on Image Processing), volume III, pages 596–599. ????, ????, October 1995.

Lin:1996:PLF

[LV96] Huawu Lin and Anastasios N. Venetsanopoulos. Perceptually lossless fractal image compression. Proceedings of the SPIE — The International Society for Optical Engineering, 2727(??): 1394–1399, ???? 1996. CODEN PSISDG. ISSN ????

Lakovic:2002:DEC

[LV02] Ksenija Laković and John Villasenor. On design of errorcorrecting reversible variable length codes. IEEE Communications Letters, 6(8):337–339, August 2002. CODEN ICLEF6. ISSN 1089-7798 (print), 1558-2558 (electronic).

Lakovic:2003:ACE

[LV03] Ksenija Laković and John Villasenor. An algorithm for construction of efficient fix-free codes. IEEE Communications Letters, 7(8):391–393, August 2003. CODEN ICLEF6. ISSN 1089-7798 (print), 1558-2558 (electronic).

Li:2008:IKC

[LV08] Ming Li and P. M. B. (Paul Michael Béla) Vitányi. An Introduction to Kolmogorov Complexity and Its Applications. Texts in computer science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 2008. ISBN 0-387-33998-1 (hardcover), 0-387-49820-6 (e-book). xxiii + 790 pp. LCCN QA267.7 .L5 2008. URL http://link.

springer.com/10.1007/978-0-387-49820-1.

Lobaz:2014:HLB

[LV14] P. Lobaz and L. Vása. Hierarchical Laplacian-based compression of triangle meshes. Graphical Models, 76(6):682–690, November 2014. CODEN GRMOFM. ISSN 1524-0703 (print),

1524-0711 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S1524070314000502.

Lekatsas:1999:RAD

[LW99] H. Lekatsas and W. Wolf. Random access decompression using binary arithmetic coding. In Storer and Cohn [SC99], pages 306–315. ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1999. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=755680. IEEE

Computer Society Order Number PR00096. IEEE Order Plan Catalog Number PR00096.

Lin:2000:LPC

[LW00] Kun-Jin Lin and Cheng-Wen Wu. A low-power CAM design for LZ data compression. IEEE Transactions on Computers, 49(10):1139–1145, October 2000. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=
&arnumber=888055.

Liu:2004:HSB

[LW04] Yali Liu and J. Wen. Heuristic search based soft-input softoutput decoding of arithmetic codes. In Storer and Cohn [SC04], page ?? ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

1281527. IEEE Computer Society Order Number: P2082.

Loomis:2008:VTO

[LW08] Jay Loomis and Mike Wasson. VC-1 technical overview, 2008.

URL http://www.microsoft.com/windows/windowsmedia/
howto/articles/vc1techoverview.aspx. VC-1 is a video

codec specification that has been standardized by the Society of Motion Picture and Television Engineers (SMPTE) and implemented by Microsoft as Microsoft Windows Media Video (WMV) 9.

Lin:2011:EPB

[LW11a] Yih-Lon Lin and Ming-Sheng Wu. An edge property-based neighborhood region search strategy for fractal image compression. Computers and Mathematics with Applications, 62 (1):310–318, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0898122111003993.

Lin:2011:RLC

[LW11b] Yu-Hsun Lin and Ja-Ling Wu. Rendering lossless compression of depth image. In Storer and Marcellin [SM11b], page

467. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=5749524. IEEE

Computer Society Order Number P4352; BMS Part Number: CFP11DCC-PRT.

LeGia:2014:DCS

[LW14] Q. T. Le Gia and H. Wendland. Data compression on the sphere using multiscale radial basis function approximation. Advances in Computational Mathematics, 40(4):923–943, August 2014. CODEN ACMHEX. ISSN 1019-7168 (print),

1572-9044 (electronic). URL http://link.springer.com/
article/10.1007/s10444-013-9334-z.

Li:2021:RGL

[LW21] Ke Li and Gang Wu. A randomized generalized low rank approximations of matrices algorithm for high dimensionality reduction and image compression. Numerical Linear Algebra with Applications, 28(1):e2338:1–e2338:??, January 2021. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Li:2010:SMM

[LWW10] Zhi Li, Feng Wu, and J. Wright. On the systematic measurement matrix for compressed sensing in the presence of gross errors. In Storer and Marcellin [SM10b], pages 356–365. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=5453481.

Liu:2012:LMO

[LWZI12] Zhenyu Liu, Dongsheng Wang, Junwei Zhou, and T. Ikenaga. Lagrangian multiplier optimization using Markov chain based rate and piecewise approximated distortion models. In Storer and Marcellin [SM12], page 404. ISBN 1-4673-0715-7 (paperback). ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=6189285. IEEE Computer Society

order number P4656.

Liveris:2003:DCB

[LXG03] A. D. Liveris, Zixiang Xiong, and C. N. Georghiades. Distributed compression of binary sources using conventional parallel and serial concatenated convolutional codes. In Storer and Cohn [SC03], pages 193–202. ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2003. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1194010. IEEE Computer Society Order number PR01896.

Luo:2018:ACS

[LXG^{+}18] Wenchang Luo, Yao Xu, Boyuan Gu, Weitian Tong, Randy Goebel, and Guohui Lin. Algorithms for communication scheduling in data gathering network with data compression. Algorithmica, 80(11):3158–3176, November 2018. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Liu:1991:RDC

[LY91] Chengwen Liu and Clement Yu. Research: Data compression using word encoding with Huffman code. Journal of the American Society for Information Science, 42(9):685–698, October 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Loewenstern:1997:SLE

[LY97] D. Loewenstern and P. N. Yianilos. Significantly lower entropy estimates for natural DNA sequences. In Storer and Cohn [SC97], pages 151–160. ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1997.

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=581998. IEEE Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Li:2016:RTA

[LYC16] Shing-Han Li, David C. Yen, and Ying-Ping Chuang. A real-time audit mechanism based on the compression technique. ACM Transactions on Management Information Systems (TMIS), 7(2):4:1–4:??, August 2016. CODEN ???? ISSN 2158-656X (print), 2158-6578 (electronic).

Lynch:1966:STC

[Lyn66a] T. J. Lynch. Sequence time coding for data compression. Proceedings of the IEEE, 54:1490–1491, October 1966. CODEN IEEPAD. ISSN 0018-9219 (print), 1558-2256 (electronic).

Lynch:1966:DCE

[Lyn66b] Thomas Joseph Lynch. Data Compression with Error-Control Coding for Space Telemetry. Ph.D. thesis, University of Maryland, College Park, MD, USA, 1966. 130 pp. URL http://

search.proquest.com/docview/302201464.

Lynn:1969:QCC

[Lyn69] Kenneth C. Lynn. A quantitative comparison of conventional information compression techniques in dental literature. American Documentation, 20(2):149–151, April 1969. CODEN AM-DOA7. ISSN 0096-946X.

Lyon:2009:BHP

[Lyo09] Richard F. Lyon. A brief history of pixel, 2009. URL http:/

/www.foveon.com/files/ABriefHistoryofPixel2.pdf.

Lempel:1976:CFS

[LZ76] A. Lempel and J. Ziv. On the complexity of finite sequences. IEEE Transactions on Information Theory, 22(1):75–81, 1976. CODEN IETTAW. ISSN 0018-9448 (print), 1557-9654 (electronic).

Lin:2005:PCP

[LZ05] Yongjing Lin and Youtao Zhang. Performance comparison of path matching algorithms over compressed control flow traces. In Storer and Cohn [SC05], pages 113–122. ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1402172.

Lansky:2006:CST

[LZ06] J. Lansky and M. Zemlicka. Compression of small text files using syllables. In Storer and Cohn [SC06], page 458. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1607301.

Lansky:2007:CSS

[LZ07] J. Lansky and M. Zemlicka. Compression of a set of strings. In Storer and Cohn [SC07], page 390. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=4148791. IEEE Computer Society Order Number P2791.

Lu:2010:IAF

[LZCF10] Jiyuan Lu, Peizhao Zhang, Hongyang Chao, and P. Fisher. An integrated algorithm for fractional pixel interpolation and motion estimation of H.264. In Storer and Marcellin [SM10b], page 541. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=5453494.

Lan:2010:ARB

[LZM^{+}10a] Xuguang Lan, Nanning Zheng, Wen Ma, Miao Hui, and Jianru Xue. Arbitrary ROI-based wavelet video coding. In Storer and Marcellin [SM10b], page 537. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=5453490.

Liu:2010:ERD

[LZM10b] Da Liu, Debin Zhao, and Siwei Ma. Error resilient dual frame motion compensation with uneven quality protection. In Storer and Marcellin [SM10b], page 539. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=5453492.

Lahsini:2011:DVC

[LZPB11] C. Lahsini, S. Zaibi, R. Pyndiah, and A. Bouallegue. Distributed video coding in pixel domain using spatial correlation at the decoder. In Storer and Marcellin [SM11b], page 463. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=5749520. IEEE

Computer Society Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Lai:2010:ADE

[LZR^{+}10] Zhongyuan Lai, Junhuan Zhu, Zhou Ren, Wenyu Liu, and Baolan Yan. Arbitrary directional edge encoding schemes for the operational rate-distortion optimal shape coding framework. In Storer and Marcellin [SM10b], pages 20–29. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5453441.

Lai:2011:ADM

[LZWL11a] Zhongyuan Lai, Zhen Zuo, Zhe Wang, and Wenyu Liu. Accurate distortion measurement using analytical model for the B-spline-based shape coding. In Storer and Marcellin [SM11b], page 465. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=5749522. IEEE

Computer Society Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Lai:2011:HAD

[LZWL11b] Zhongyuan Lai, Zhen Zuo, Zhe Wang, and Wenyu Liu. A hybrid admissible distortion checking algorithm for the Bspline-based operational rate-distortion optimal shape coding. In Storer and Marcellin [SM11b], page 464. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5749521. IEEE Computer Society

Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Lan:2008:PPA

[LZX^{+}08] Xuguang Lan, Nanning Zheng, Jianru Xue, Weike Chen, Bin Wang, Wen Ma, and Songlin Zhao. A peer-to-peer architecture based on scalable video coding. In Storer and Marcellin [SM08], page 529. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008;

QA76.9.D33 INTERNET. URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=4483356.

Lan:2009:JNS

[LZX^{+}09] Xuguang Lan, Nanning Zheng, Jianru Xue, Ce Li, and Songlin Zhao. Joint network-source video coding based on Lagrangian rate allocation. In Storer and Marcellin [SM09], page 454. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=4976508. IEEE Computer

Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

Liu:2011:TRL

[LZX^{+}11] Xianming Liu, Debin Zhao, Ruiqin Xiong, Siwei Ma, Wen Gao, and Huifang Sun. Transductive regression with local and global consistency for image super-resolution. In Storer and Marcellin [SM11b], pages 173–182. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=5749475. IEEE Computer Society Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Linder:1998:MDR

[LZZ98] T. Linder, R. Zamir, and K. Zeger. The multiple description rate region for high resolution source coding. In Storer and Cohn [SC98], pages 149–158. ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1998. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=672141. IEEE Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Liu:2012:MSS

[LZZ^{+}12] Xianming Liu, Deming Zhai, Guangtao Zhai, Debin Zhao, Ruiqin Xiong, and Wen Gao. Multi-scale spatial error concealment via hybrid Bayesian regression. In Storer and Marcellin [SM12], pages 169–178. ISBN 1-4673-0715-7 (paperback). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=6189248. IEEE Computer Society order number P4656.

Matthews:1998:CRT

[M^{+}98] Michael B. Matthews et al., editors. Conference record of the Thirty-Second Asilomar Conference on Signals, Systems and Computers: November 1–4, 1998, Pacific Grove, California. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1998. ISBN 0-7803-5148-7, 0-7803-5149-5, 0-7803-5150-9. LCCN TK5101.A1 A85 1998; TK454.2 .A8 1998.

Martinez:2009:DAC

[M^{+}09] V. J. (Vicent J.) Martinez et al. Data analysis in cosmology, volume 665 of Lecture notes in physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 3-540-23972-3, 3-540-44767-9 (e-). ISSN 0075-8450 (print), 1616-6361 (electronic). xii + 636 pp. LCCN QB981 .D38 2009.

Ma:1978:DC

[Ma78] Joong Soo Ma. Data Compression. Ph.D. thesis, University of Massachusetts Amherst, Amherst, MA, USA, 1978. 148 pp.

URL http://search.proquest.com/docview/302919063.

Ma:1996:MRB

[Ma96] Kai-Kuang Ma. Media reviews: Block truncation coding in perspective: Image Data Compression: Block Truncation Coding, by Belur V. Dasarathy (IEEE Computer Society Press, 1995, 284 pp., $50 list, $40 members, ISBN 0-89186-6876-4). IEEE MultiMedia, 3(2):83–84, Summer 1996. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166

(electronic). URL http://dlib.computer.org/mu/books/
mu1996/pdf/u2083.pdf.

Moffat:2005:BCN

[MA05] A. Moffat and V. N. Anh. Binary codes for non-uniform sources. In Storer and Cohn [SC05], pages 133–142. ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1402174.

Moffat:2006:BCL

[MA06] Alistair Moffat and Vo Ngoc Anh. Binary codes for locally homogeneous sequences. Information Processing Letters, 99 (5):175–180, September 15, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Maa:1994:IEB

[Maa94] Chia-Yiu Y. Maa. Identifying the existence of bar codes in compressed images. Computer Vision, Graphics, and Image Processing. Graphical Models and Image Processing, 56 (4):352–356, July 1994. CODEN CGMPE5. ISSN 1049-

9652. URL http://www.idealibrary.com/links/artid/
cgip.1994.1032/production;   http://www.idealibrary.
com/links/artid/cgip.1994.1032/production/pdf.

Miguel:2004:RCW

[MAC^{+}04] A. C. Miguel, A. R. Askew, A. Chang, S. Hauck, R. E. Ladner, and E. A. Riskin. Reduced complexity wavelet-based predictive coding of hyperspectral images for FPGA implementation. In Storer and Cohn [SC04], pages 469–478. ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1281492. IEEE Computer Society

Order Number: P2082.

MacCormick:2012:NAC

[Mac12] John MacCormick. Nine algorithms that changed the future: the ingenious ideas that drive today’s computers. Princeton University Press, Princeton, NJ, USA, 2012. ISBN 0-691-14714-0 (hardcover), 0-691-15819-3 (paperback). x + 2 + 219 pp. LCCN QA76 .M21453 2012. URL http:

//press.princeton.edu/chapters/s9528.pdf;   http://
www.jstor.org/stable/10.2307/j.ctt7t71s. With a fore-

word by Christopher M. Bishop.

Mahoney:1990:GB

[Mah90] Michael S. Mahoney. Goldbach’s biography. In Charles Coulston Gillispie, editor, Dictionary of Scientific Biography, page ?? Scribner’s, New York, NY, USA, 1990. 14 volumes, 1970–1990.

Mahoney:2008:BAR

[Mah08a] Matt Mahoney. Better Archiver with Recursive Functionality

(BARF), 2008. URL http://www.cs.fit.edu/~mmahoney/
compression/barf.html.

Mahoney:2008:DCP

[Mah08b] Matt Mahoney. Data compression programs, 2008. URL http:

//www.cs.fit.edu/~mmahoney/compression. As of July 23,

2009, this page is no longer maintained. The newest version

can be found at http://mattmahoney.net/dc/.

Maier:1995:AES

[Mai95] M. W. Maier. Algorithm evaluation for the synchronous data compression standards. In Storer and Cohn [SC95], page

444. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=515554;   http://
ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=515596.

IEEE catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Makinen:1989:ITA

[Mäk89] E. Mäkinen. On implementing two adaptive data-compression schemes. The Computer Journal, 32(3):238–240, June 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/
content/32/3/238.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_32/Issue_03/tiff/
238.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_03/tiff/239.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_32/Issue_03/
tiff/240.tif.

Mallat:1989:TMS

[Mal89] Stephane Mallat. A theory for multiresolution signal decomposition: The wavelet representation. IEEE Transactions on Pattern Analysis and Machine Intelligence, 11(7):674–693, July 1989. CODEN ITPIDJ. ISSN 0162-8828.

Malvar:1999:FPW

[Mal99] H. S. Malvar. Fast progressive wavelet coding. In Storer and Cohn [SC99], pages 336–343. ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1999. URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

755683. IEEE Computer Society Order Number PR00096. IEEE Order Plan Catalog Number PR00096.

Malvar:2000:FPI

[Mal00] H. S. Malvar. Fast progressive image coding without wavelets. In Storer and Cohn [SC00], pages 243–252. ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2000. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=838164. This millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Malvar:2001:FAE

[Mal01] H. S. Malvar. Fast adaptive encoder for bi-level images. In Storer and Cohn [SC01c], pages 253–262. ISBN 0-7695-1031-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 2001. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=917156. IEEE CSP num-

ber 01PR1031.

Malvar:2006:ARL

[Mal06] H. S. Malvar. Adaptive run-length/Golomb–Rice encoding of quantized generalized Gaussian sources with unknown statistics. In Storer and Cohn [SC06], pages 23–32. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1607237.

Malvar:2007:LNL

[Mal07] H. S. Malvar. Lossless and near-lossless audio compression using integer-reversible modulated lapped transforms. In Storer and Cohn [SC07], pages 323–332. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=4148771. IEEE Computer Society Order Number P2791.

Moskowitz:2010:ITE

[MAL10] I. S. Moskowitz, F. Ahmed, and P. A. Lafferty. Information theoretic effects of JPEG compression on image steganography. International Journal of Computer Applications, 32 (3):318–327, 2010. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.tandfonline.com/doi/full/10.
2316/Journal.202.2010.3.202-2736.

Mandelbrot:1977:FFC

[Man77] Benoît B. Mandelbrot. Fractals: form, chance, and dimension. W. H. Freeman, New York, NY, USA, revised edition, 1977. ISBN 0-7167-0473-0, 0-7167-0474-9 (paperback). xvi + 365 pp. LCCN QA447 .M3613 1977a. Translated from the French.

Mandelbrot:1982:FGN

[Man82] Benoît B. Mandelbrot. The Fractal Geometry of Nature. W. H. Freeman, New York, NY, USA, 1982. ISBN 0-7167-1186-9. v + 460 pp. LCCN QA447 .M25 1982; QA447 .M357 1982. Schriftenreihe für den Referenten. [Series for the Referee].

Manber:1997:TCS

[Man97] Udi Manber. A text compression scheme that allows fast searching directly in the compressed file. ACM Transactions on Information Systems, 15(2):124–136, April 1997. CODEN ATISET. ISSN 1046-8188. URL http://www.acm.org:80/

tois/abstracts/manber.html.

Manning:1998:MCV

[Man98] Manning. Motion compensated video compression overview,

1998. URL http://www.newmediarepublic.com/dvideo/
compression/adv08.html.

Manzini:2001:ABW

[Man01] Giovanni Manzini. An analysis of the Burrows — Wheeler transform. Journal of the ACM, 48(3):407–430, May 2001. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/citations/
journals/jacm/2001-48-3/p407-manzini/.

Marr:1980:TDP

[Mar80] James Douglas Marr. Two-Dimensional Prediction for Data Rate Compression of Lpc Parameters. Ph.D. thesis, Georgia Institute of Technology, Atlanta, GA, USA, 1980. 154 pp. URL

http://search.proquest.com/docview/303008916.

Marking:1990:DGI

[Mar90] Michael P. Marking. Decoding group 3 images. C Users Journal, 8(6):45–54, June 1990. ISSN 0898-9788.

Markas:1993:DCA

[Mar93] Tassos Markas. Data compression: Algorithms and architectures. Ph.D. thesis, Duke University, Durham, NC 27708, USA,

1993. 170 pp. URL http://search.proquest.com/docview/
304040162.

Marquant:2002:PPT

[Mar02] G. Marquant. Perceptual preprocessing techniques applied to video compression: some result elements and analysis. In Storer and Cohn [SC02], page ?? ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2002. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1000006. IEEE Computer Society Order Number PR01477.

Masui:1991:KDC

[Mas91] T. Masui. Keyword dictionary compression using efficient trie implementation. In Storer and Reif [SR91b], page ?? ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1990. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=213323. IEEE Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Meyer:2000:FAW

[MAS00] F. G. Meyer, A. Averbuch, and J. O. Strömberg. Fast adaptive wavelet packet image compression. IEEE Transactions on Image Processing, 9(5):792–800, May 2000. CODEN IIPRE4. ISSN 1057-7149 (print), 1941-0042 (electronic).

Meyer:1998:FWP

[MASC98] F. G. Meyer, A. Z. Averbuch, J. O. Stromberg, and R. R. Coifman. Fast wavelet packet image compression. In Storer and Cohn [SC98], page ?? ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1998. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=672305. IEEE

Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Massberg:2015:GHC

[Maß15] Jens Maßberg. Generalized Huffman coding for binary trees with choosable edge lengths. Information Processing Letters, 115(4):502–506, April 2015. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0020019014002634.

Mathews:1996:EDC

[Mat96] G. Jason Mathews. Evaluating data-compression algorithms. Dr. Dobb’s Journal of Software Tools, 21(1):50–53, January 1996. CODEN DDJOEB. ISSN 1044-789X.

Mattern:2012:MSD

[Mat12] C. Mattern. Mixing strategies in data compression. In Storer and Marcellin [SM12], pages 337–346. ISBN 1-4673-0715-7 (paperback). ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=6189265. IEEE Computer Society

order number P4656.

Matthews:2015:HCL

[Mat15] Suzanne J. Matthews. Heterogeneous compression of large collections of evolutionary trees. IEEE/ACM Transactions on Computational Biology and Bioinformatics, 12(4):807–814, July 2015. CODEN ITCBCY. ISSN 1545-5963 (print), 1557-9964 (electronic).

Max:1960:QMD

[Max60] Joel Max. Quantizing for minimum distortion. IRE Transactions on Information Theory, IT-6:7–12, March 1960. CODEN IRITAY. ISSN 0096-1000.

Mayer:1999:BME

[May99] J. Mayer. A blending model for efficient compression of smooth images. In Storer and Cohn [SC99], pages 228–237. ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1999. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=755672. IEEE Computer Society Or-

der Number PR00096. IEEE Order Plan Catalog Number PR00096.

Malak:1991:IDL

[MB91] M. Malak and J. Baker. An image database for low bandwidth communication links. In Storer and Reif [SR91b], pages 53–62. ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1990. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=213373. IEEE

Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Mahmoud:2000:ELB

[MB00a] H. A. Mahmoud and M. Bayoumi. An efficient low-bit rate motion compensation technique based on quadtree. In Storer and Cohn [SC00], page ?? ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2000. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=838207. This

millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Mahmoud:2000:ESE

[MB00b] H. A. Mahmoud and M. Bayoumi. An efficient successive elimination algorithm for block-matching motion estimation. In Storer and Cohn [SC00], page ?? ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2000. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=838206. This millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Mitran:2002:TSC

[MB02] P. Mitran and J. Bajcsy. Turbo source coding: a noiserobust approach to data compression. In Storer and Cohn [SC02], page ?? ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2002. URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

1000008. IEEE Computer Society Order Number PR01477.

Menconi:2008:DCG

[MBB08] Giulia Menconi, Vieri Benci, and Marcello Buiatti. Data compression and genomes: a two-dimensional life domain map. Journal of Theoretical Biology, 253(2):281–288, July 21, 2008. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-

8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519308001252.

Mukherjee:1991:MDE

[MBBA91] A. Mukherjee, H. Bheda, M. A. Bassiouni, and T. Acharya. Multibit decoding/encoding of binary codes using memory based architectures. In Storer and Reif [SR91b], pages 352–361. ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1990. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=213345. IEEE

Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Mitin:2003:PSI

[MBWP03] Alexander V. Mitin, Jon Baker, Krzysztof Wolinski, and Peter Pulay. Parallel stored-integral and semidirect Hartree–Fock and DFT methods with data compression. Journal of Computational Chemistry, 24(2):154–160, January 30, 2003. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Mazza:2013:MSC

[MC13] Tommaso Mazza and Stefano Castellana. Multi-sided compression performance assessment of ABI SOLiD WES data. Algorithms (Basel), 6(2):309–318, June 2013. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/6/2/309.

McCreight:1976:SES

[McC76] Edward M. McCreight. A space-economical suffix tree construction algorithm. Journal of the ACM, 23(2):262–272, April 1976. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

McConnell:1992:FDF

[McC92] Kenneth R. McConnell. FAX: Digital Facsimile Technology and Applications. Artech House Inc., Norwood, MA, USA,

1992. ?? pp. URL http://www.mathworks.com/.

Mitra:2003:CDE

[McC03] Tulika Mitra and Tzi cker Chiueh. Compression-domain editing of 3D models. In Storer and Cohn [SC03], pages 343– 352. ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2003. URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

1194025. IEEE Computer Society Order number PR01896.

McGeoch:1993:DC

[McG93] Catherine C. McGeoch. Data compression. American Mathematical Monthly, 100(5):493–497, May 1993. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Munkberg:2008:PHT

[MCHAM08] J. Munkberg, P. Clarberg, J. Hasselgren, and T. Akenine-Möller. Practical HDR texture compression. Computer Graphics Forum, 27(6):1664–1676, September 2008. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mo:2012:TBD

[MCL12] Shangfeng Mo, Hong Chen, and Yinglong Li. Topology-based data compression in wireless sensor networks. Lecture Notes in Computer Science, 7418:22–34, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-32281-5_4/.

McMillan:1956:TII

[McM56] Brockway McMillan. Two inequalities implied by unique decipherability. IEEE Transactions on Information Theory, 2(4): 115–116, December 1956. CODEN IETTAW. ISSN 0018-9448 (print), 1557-9654 (electronic).

Mera:2017:ATP

[MCM^{+}17] Maria Isabel Mera, Jonah Caplan, Seyyed Hasan Mozafari, Brett H. Meyer, and Peter Milder. Area, throughput, and power trade-offs for FPGA- and ASIC-based execution stream compression. ACM Transactions on Embedded Computing Systems, 16(4):96:1–96:??, August 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Milani:2009:BIS

[MCRKC09] S. Milani, C. Cruz-Reyes, J. Kari, and G. Calvagno. A binary image scalable coder based on reversible cellular automata transform and arithmetic coding. In Storer and Marcellin [SM09], page 460. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

4976514. IEEE Computer Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

Majani:1995:MBJ

[MD95] E. E. Majani and W. C. Dias. Making the best of JPEG at very high compression ratios: Rectangular pixel averaging for Mars PathFinder. In Storer and Cohn [SC95], page ?? ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=515600. IEEE catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Mondal:2006:QCP

[MDH06] B. Mondal, S. Dutta, and R. W. Heath, Jr. Quantization on the complex projective space. In Storer and Cohn [SC06], pages 242–251. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=1607259.

Marron:1967:ADC

[MdM67] B. A. Marron and P. A. D. de Maine. Automatic data compression. Communications of the ACM, 10(11):711–715, November 1967. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Also published in/as: File Organization, Siwerts and Zeitlinge (Amsterdam), 1969.

Manduchi:2000:SPT

[MDPM00] R. Manduchi, S. Dolinar, F. Pollara, and A. Matache. Semantic progressive transmission for deep space communications. In Storer and Cohn [SC00], page ?? ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2000. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=838208. This millennial edition of the DCC proceedings is dedicated to the memory of David A.

Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Muresan:2002:QHS

[ME02] D. Muresan and M. Effros. Quantization as histogram segmentation: globally optimal scalar quantizer design in network systems. In Storer and Cohn [SC02], pages 302–311. ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2002. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=999968. IEEE

Computer Society Order Number PR01477.

Mishali:2010:XAD

[ME10] M. Mishali and Y. C. Eldar. Xampling: Analog data compression. In Storer and Marcellin [SM10b], pages 366–375. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=5453480.

Markham:2009:IIM

[MEIS09] T. S. Markham, S. C. Evans, J. Impson, and E. Steinbrecher. Implementation of an incremental MDL-based two part compression algorithm for model inference. In Storer and Marcellin [SM09], pages 322–331. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=4976476. IEEE Computer Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

Mantiuk:2006:BCH

[MEMS06] Rafał Mantiuk, Alexander Efremov, Karol Myszkowski, and Hans-Peter Seidel. Backward compatible high dynamic range MPEG video compression. ACM Transactions on Graphics, 25 (3):713–723, July 2006. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Martinez-Enriquez:2011:LTG

[MEO11] E. Martínez-Enríquez and A. Ortega. Lifting transforms on graphs for video coding. In Storer and Marcellin [SM11b], pages 73–82. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

5749465. IEEE Computer Society Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Mermelstein:1991:TAS

[Mer91] P. Mermelstein. Trends in audio and speech compression for storage and real-time communication. In Storer and Reif [SR91b], page ?? ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1990. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=213306. IEEE

Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Meridian:2003:MHP

[Mer03] Meridian. Meridian home page, 2003. URL http://www.

meridian-audio.com/.

Merhav:1994:MRL

[MF94] N. Merhav and M. Feder. The minimax redundancy is a lower bound for most sources. In Storer and Cohn [SC94], pages 52– 61. ISBN 0-8186-5636-0, 0-8186-5637-9 (case). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1994.

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=305912. IEEE Catalog Number 93TH0626-2. IEEE Computer Society Press Order Number 5637-02.

Martins:1996:BLI

[MF96] B. Martins and S. Forchhammer. Bi-level image compression with tree coding. In Storer and Cohn [SC96], pages 270–279. ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1996. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=488332. IEEE

Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Manor:1997:ITU

[MF97a] D. Manor and M. Feder. An iterative technique for universal lossy compression of individual sequences. In Storer and Cohn [SC97], pages 141–150. ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1997.

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=581995. IEEE Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Martins:1997:LLC

[MF97b] B. Martins and S. Forchhammer. Lossy/lossless coding of bilevel images. In Storer and Cohn [SC97], page ?? ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1997. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=582116. IEEE Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Martins:1998:LCV

[MF98a] B. Martins and S. Forchhammer. Lossless compression of video using motion compensation. In Storer and Cohn [SC98], page ?? ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1998. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=672302. IEEE

Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Mello:1998:LBR

[MF98b] M. A. Mello and T. R. Fischer. Low bit-rate video coding with trellis source codes. In Storer and Cohn [SC98], page ?? ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1998. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=672304. IEEE

Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Moore:1998:DCY

[MF98c] Robert L. Moore and D. Gregory Foley. Date compression and Year 2000 challenges. Dr. Dobb’s Journal of Software Tools, 23(5):20–22, 24, 109, May 1998. CODEN DDJOEB. ISSN

1044-789X. URL http://www.ddj.com/ftp/1998/1998_05/
y2kcomp.txt;   http://www.ddj.com/ftp/1998/1998_05/
y2kcomp.zip.

Meron:2004:OFS

[MF04] E. Meron and M. Feder. Optimal finite state universal coding of individual sequences. In Storer and Cohn [SC04], pages 332–341. ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=1281478. IEEE

Computer Society Order Number: P2082.

Mun:2010:BCS

[MF10] Sungkwang Mun and J. E. Fowler. Block compressed sensing of images using directional transforms. In Storer and Marcellin [SM10b], page 547. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ????

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=
&arnumber=5453522.

Mun:2011:RRB

[MF11] Sungkwang Mun and J. E. Fowler. Residual reconstruction for block-based compressed sensing of video. In Storer and Marcellin [SM11b], pages 183–192. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=5749476. IEEE Computer Society Order Number P4352; BMS Part Number: CFP11DCC-PRT.

McGregor:1996:FFC

[MFCM96] D. R. McGregor, R. J. Fryer, P. Cockshott, and P. Murray. Faster fractal compression. Dr. Dobb’s Journal of Software Tools, 21(1):34, 36, 38–40, January 1996. CODEN DDJOEB. ISSN 1044-789X.

Muller:2009:JVL

[MFD^{+}09] Daniel Muller, Bernhard Fleck, George Dimitoglou, Benjamin W. Caplins, Desmond E. Amadigwe, Juan Pablo Garcia Ortiz, Benjamin Wamsler, Alen Alexanderian, V. Keith Hughitt, and Jack Ireland. JHelioviewer: Visualizing large sets of solar images using JPEG 2000. Computing in Science and Engineering, 11(5):38–47, September/October 2009. CODEN CSENFA. ISSN 0740-7475 (print), 1558-1918 (electronic).

Mohammed:2007:BAB

[MFT07] H. M. Mohammed, N. Farber, and H. Thoma. Bit allocation based on motion vector analysis for H.264/AVC. In Storer and Cohn [SC07], pages 313–322. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=4148770. IEEE Computer Society Order Number P2791.

Maleh:2008:SRS

[MG08] R. Maleh and A. C. Gilbert. Sublinear recovery of sparse wavelet signals. In Storer and Marcellin [SM08], pages 342– 351. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33

INTERNET. URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4483312.

Mao:2010:STE

[MG10] M. Z. Mao and R. M. Gray. Stationary and Trellis encoding for IID sources and simulation. In Storer and Marcellin [SM10b], pages 3–12. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=
5453439.

Marcellin:2000:OJ

[MGBB00] M. W. Marcellin, M. J. Gormish, A. Bilgin, and M. P. Boliek. An overview of JPEG-2000. In Storer and Cohn [SC00], pages 523–541. ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2000. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=838192. This

millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Munoz-Gomez:2011:INF

[MGBRMSS11] Juan Muñoz-Gómez, Joan Bartrina-Rapesta, Michael W. Marcellin, and Joan Serra-Sagristà. Influence of noise filtering in coding computed tomography with JPEG2000. In Storer and Marcellin [SM11b], pages 413–422. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=5749499. IEEE Computer Society Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Motta:2007:DCE

[MGC07] G. Motta, J. Gustafson, and S. Chen. Differential compression of executable code. In Storer and Cohn [SC07], pages 103–112. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=4148749. IEEE Computer

Society Order Number P2791.

Mallet:2008:CBI

[MGD08] A. Mallet, L. Gueguen, and M. Datcu. Complexity based image artifact detection. In Storer and Marcellin [SM08], page 534. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33

INTERNET. URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4483361.

Magnani:2005:OOB

[MGG05] A. Magnani, A. Ghosh, and R. M. Gray. Optimal one-bit quantization. In Storer and Cohn [SC05], pages 270–278. ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1402188.

Mao:2011:AOS

[MGL11] M. Z. Mao, R. M. Gray, and T. Linder. On asymptotically optimal stationary source codes for IID sources. In Storer and Marcellin [SM11b], pages 3–12. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=5749458. IEEE Computer Society Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Merry:2013:FPB

[MGM13] Bruce Merry, James Gain, and Patrick Marais. Fast in-place binning of laser range-scanned point sets. Journal on Computing and Cultural Heritage (JOCCH), 6(3):14:1–14:??, July 2013. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Misra:2008:HRF

[MGV08] V. Misra, V. K. Goyal, and L. R. Varshney. High-resolution functional quantization. In Storer and Marcellin [SM08], pages 113–122. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33

INTERNET. URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4483289.

Malvar:2003:LCT

[MHKK03] H. S. Malvar, A. Hallapuro, M. Karczewicz, and L. Kerofsky. Low-complexity transform and quantization in H.264/ AVC. IEEE Transactions on Circuits and Systems for Video Technology, 13(7):598–603, July 2003. CODEN ITCTEM. ISSN 1051-8215 (print), 1558-2205 (electronic). URL http://

research.microsoft.com/pubs/77882/MalvarCSVTJuly03.
pdf.

Malvar:2006:LCT

[MHKK06] H. S. Malvar, A. Hallapuro, M. Karczewicz, and L. Kerofsky. Low-complexity transform and quantization in H.264/AVC,

2006. URL http://research.microsoft.com/apps/pubs/
default.aspx?id=77882;   http://research.microsoft.
com/pubs/77882/MalvarCSVTJuly03.pdf. See original paper

[MHKK03].

Mitarai:2001:CPM

[MHM^{+}01] S. Mitarai, M. Hirao, T. Matsumoto, A. Shinohara, M. Takeda, and S. Arikawa. Compressed pattern matching for SE-QUITUR. In Storer and Cohn [SC01c], pages 469–478. ISBN 0-7695-1031-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2001. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=917178. IEEE

CSP number 01PR1031.

Millett:1991:ICM

[MI91] R. P. Millett and E. L. Ivie. Index compression method with compressed mode Boolean operators. In Storer and Reif [SR91b], page ?? ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1990. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=213324. IEEE

Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Miastkowski:1990:MFG

[Mia90] S. Miastkowski. A ‘more filling’ generation of tape backup. Byte Magazine, 15(13):235–237, December 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Miano:1999:CIF

[Mia99] John Miano. Compressed Image File Formats: JPEG, PNG, GIF, XBM, BMP: Your guide to graphics files on the Web. ACM Press and Addison-Wesley, New York, NY, USA, 1999. ISBN 0-201-60443-4, 0-201-61657-2. xi + 264 pp.

Miller:1981:LES

[Mil81] Lance A. Miller. Letter to the Editor: On spelling correction and beyond. Communications of the ACM, 24(9):608–609, September 1981. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Pet80, Dun81, Nix81, Pet81].

Miller:1995:MCD

[Mil95] Mark Christopher Miller. Multiscale compression of digital terrain data to meet real-time rendering rate constraints. Ph.D. thesis, University of California, Davis, Davis, CA, USA,

1995. 216 pp. URL http://search.proquest.com/docview/
304175054.

Mishra:1991:AID

[Mis91] Sanjay Mishra. An algorithm for image data compression based on vector quantization. M.S. thesis, Texas A&I University, Kingsville, TX, USA, 1991. 128 pp. URL http://

search.proquest.com/docview/303967682.

Mistereggen:2011:CSR

[Mis11] Eirik Mistereggen. Compression of Short Read data sets. Masteroppgave i informatikk, Institutt for informatikk, Universitetet i Oslo, Oslo, Norway, 2011.

Mitzenmacher:2001:HFO

[Mit01] M. Mitzenmacher. On the hardness of finding optimal multiple preset dictionaries. In Storer and Cohn [SC01c], pages 411–418. ISBN 0-7695-1031-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2001. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=917172. IEEE CSP number 01PR1031.

Miu:2007:RDD

[Miu07] Hong Zhi Miu. The research and design on data compression algorithm. [unknown], Zhejiang University, Zhejiang, Peoples Republic of China, 2007. URL http://search.proquest.

com/docview/1024942932.

Mayne:1975:ICF

[MJ75] A. Mayne and E. B. James. Information compression by factorising common strings. The Computer Journal, 18(2):157–160, May 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/18/2/157.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_02/tiff/157.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_18/Issue_02/
tiff/158.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_02/tiff/159.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_18/Issue_02/
tiff/160.tif.

Madisetti:1992:RTI

[MJB92] A. Madisetti, R. Jain, and R. L. Baker. Real time implementation of pruned tree search vector quantization. In Storer and Cohn [SC92], pages 152–161. ISBN 0-8186-2717-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=227466. IEEE catalog number 91TH0436-6.

Munoz-Jimenez:2008:NBM

[MJZMA08] V. Munoz-Jimenez, A. Zergainoh-Mokraoui, and J.-P. Astruc. New bidirectional motion estimation using mesh-based frame interpolation for videoconferencing applications. In Storer and Marcellin [SM08], page 536. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33 INTERNET. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=4483363.

Mannan:2003:BHC

[MK03] M. Abdul Mannan and M. Kaykobad. Block Huffman coding. Computers and Mathematics with Applications, 46(10–11):1581–1587, November/December 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0898122103901933.

Mitra:2006:XET

[MK06] S. Mitra and K. S. Kim. XPAND: an efficient test stimulus compression technique. IEEE Transactions on Computers, 55(2):163–173, February 2006. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=
1566577.

Moriya:2010:ELC

[MKH10] T. Moriya, Y. Kamamoto, and N. Harada. Enhanced lossless coding tools of LPC residual for ITU-T G.711.0. In Storer and Marcellin [SM10b], page 546. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=5453523.

Min:2012:DCT

[MKK12] Joonki Min, Jookyoung Kim, and Youngmi Kwon. Data compression technique for wireless sensor networks. Lecture Notes in Computer Science, 7425:9–16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-32645-
5_2/.

Mitra:2008:NPA

[MKM^{+}08] Suman K. Mitra, Malay K. Kundu, C. A. Murthy, Bhargab B. Bhattacharya, and Tinku Acharya. A new probabilistic approach for fractal based image compression. Fundamenta Informaticae, 87(3–4):417–433, November 2008. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mahdian:2006:WGC

[MKNG06] A. Mahdian, H. Khalili, E. Nourbakhsh, and M. Ghodsi. Web graph compression by edge elimination. In Storer and Cohn [SC06], page 459. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=1607302.

Meyer:2012:DPD

[MKSS12] Quirin Meyer, Benjamin Keinert, Gerd Sußner, and Marc Stamminger. Data-parallel decompression of triangle mesh topology. Computer Graphics Forum, 31(8):2541–2553, December 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Miguel:2006:CLM

[MKW^{+}06] A. C. Miguel, J. F. Keane, J. Whiteaker, H. Zhang, and A. Paulovich. Compression of LC/MS proteomic data. In Storer and Cohn [SC06], page 460. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=1607303.

Mayer:1998:PPE

[ML98] J. Mayer and G. G. Langdon. Post-processing enhancement of decompressed images using variable order Bézier polynomials and distance transform. In Storer and Cohn [SC98], page ?? ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1998. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=672303. IEEE

Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Maglo:2015:MCS

[MLDH15] Adrien Maglo, Guillaume Lavoué, Florent Dupont, and Céline Hudelot. 3D mesh compression: Survey, comparisons, and emerging trends. ACM Computing Surveys, 47(3):44:1–44:??, April 2015. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

Milidiu:2003:FDM

[MLMD03] R. L. Milidiu, E. S. Laber, L. O. Moreno, and J. C. Duarte. A fast decoding method for prefix codes. In Storer and Cohn [SC03], page ?? ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2003. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1194057. IEEE Computer Society Order number PR01896.

Milidiu:1999:BCL

[MLP99a] R. L. Milidiu, E. S. Laber, and A. A. Pessoa. Bounding the compression loss of the FGK algorithm. In Storer and Cohn [SC99], page ?? ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1999. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=785696. IEEE

Computer Society Order Number PR00096. IEEE Order Plan Catalog Number PR00096.

Milidiu:1999:WEP

[MLP99b] R. L. Milidiu, E. S. Laber, and A. A. Pessoa. A workefficient parallel algorithm for constructing Huffman codes. In Storer and Cohn [SC99], pages 277–286. ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1999. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=755677. IEEE Computer Society Or-

der Number PR00096. IEEE Order Plan Catalog Number PR00096.

Mahapatra:2003:PAA

[MLS03] Nihar R. Mahapatra, Jiangjiang Liu, and Krishnan Sundaresan. The performance advantage of applying compression to the memory system. ACM SIGPLAN Notices, 38(2s):86–96, February 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Manber:1993:SAN

[MM93] Udi Manber and Gene Myers. Suffix arrays: a new method for on-line string searches. SIAM Journal on Computing, 22(5): 935–948, October 1993. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Moellenhoff:1995:QMS

[MM95] M. S. Moellenhoff and M. W. Maier. Queuing models of synchronous compressors. In Storer and Cohn [SC95], page 445. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=515555. IEEE catalog

number 95TH8037. IEEE Computer Society Press order number PR07010.

Meadows:1997:LBR

[MM97] S. Meadows and S. Mitra. Low bit rate color image coding with adaptive encoding of wavelet coefficients. In Storer and Cohn [SC97], page ?? ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1997. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=582118. IEEE

Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Mukherjee:1999:JOF

[MM99] K. Mukherjee and A. Mukherjee. Joint optical flow motion compensation and video compression using hybrid vector quantization. In Storer and Cohn [SC99], page ?? ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1999. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=785698. IEEE Computer Society Or-

der Number PR00096. IEEE Order Plan Catalog Number PR00096.

Mukherjee:2000:SCD

[MM00] K. Mukherjee and A. Mukherjee. A spatially coherent discrete wavelet transform — accessing the localization property for data compression. In Storer and Cohn [SC00], page ?? ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2000. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=838211. This millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Milenkovic:2003:SBT

[MM03] A. Milenkovic and M. Milenkovic. Stream-based trace compression. IEEE Computer Architecture Letters, 2(1):4, January 2003. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Morhac:2005:MNS

[MM05] M. Morhác and V. Matousek. Multidimensional nuclear spectra compression using fast adaptive Fourier-based transforms. Computer Physics Communications, 165(2):127–138, January 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465504004965.

Milidiu:2006:CCP

[MM06] R. L. Milidiu and C. G. Mello. Crypto-compression prefix coding. In Storer and Cohn [SC06], page 461. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1607304.

Mukherjee:1999:BRD

[MMA99] K. Mukherjee, A. Mukherjee, and T. Acharya. ‘bit rate on demand’ using pruned tree-structured hierarchical lookup vector quantization. In Storer and Cohn [SC99], pages 42–51. ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1999. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=755652. IEEE Computer Society Or-

der Number PR00096. IEEE Order Plan Catalog Number PR00096.

Mukherjee:2000:DIA

[MMA00] K. Mukherjee, A. Mukherjee, and T. Acharya. Distributed Internet-adaptive image compression. In Storer and Cohn [SC00], page ?? ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2000. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=838210. This

millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Milenkovic:2007:AHS

[MMB07] M. Milenkovic, A. Milenkovic, and M. Burtscher. Algorithms and hardware structures for unobtrusive real-time compression of instruction and data address traces. In Storer and Cohn [SC07], pages 283–292. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=4148767. IEEE Computer Society Order Number P2791.

Mitra:1998:SPI

[MMK98] Suman K. Mitra, C. A. Murthy, and Malay K. Kundu. A study on partitioned iterative function systems for image compression. Fundamenta Informaticae, 34(4):413–428, July 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Memon:1991:PTL

[MMS91] N. D. Memon, S. S. Magliveras, and K. Sayood. Prediction trees and lossless image compression. In Storer and Reif [SR91b], pages 83–92. ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1990. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=213375. IEEE

Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Matei:2012:WET

[MMZ12] Basarab Mateï, Sylvain Meignen, and Anastasia Zakharova. On a (W)ENO-type multiscale representation based on quincunx refinement: Application to image compression. Lecture Notes in Computer Science, 6920:473–487, 2012. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-27413-8_30.

Makinen:2008:SII

[MN08a] V. Makinen and G. Navarro. On self-indexing images — image compression with added value. In Storer and Marcellin [SM08], pages 422–431. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008;

QA76.9.D33 INTERNET. URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=4483320.

Makinen:2008:DEC

[MN08b] Veli Mäkinen and Gonzalo Navarro. Dynamic entropycompressed sequences and full-text indexes. ACM Transactions on Algorithms, 4(3):32:1–32:??, June 2008. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Minoo:2008:MLR

[MN08c] K. Minoo and Truong Nguyen. Maximum likelihood rate estimation: With applications in image and video compression. In Storer and Marcellin [SM08], page 535. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33 INTER-

NET. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=4483362.

Minoo:2009:ECP

[MN09] K. Minoo and Truong Nguyen. Entropy coding via parametric source model with applications in fast and efficient compression of image and video data. In Storer and Marcellin [SM09], page 461. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=4976515. IEEE

Computer Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

Marchetti:2018:SDC

[MNBC18] Yuliya Marchetti, Hai Nguyen, Amy Braverman, and Noel Cressie. Spatial data compression via adaptive dispersion clustering. Computational Statistics & Data Analysis, 117(??):138–153, January 2018. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0167947317301731.

MNG:2003:MMI

[MNG03] MNG Group. MNG (Multiple-image Network Graphics) for-

mat version 1.0, 2003. URL http://www.libpng.org/pub/
mng/spec/.

Milward:2004:DIL

[MNM04] Mark Milward, José Luis Núñez, and David Mulvaney. Design and implementation of a lossless parallel high-speed data compression system. IEEE Transactions on Parallel and Distributed Systems, 15(6):481–490, June 2004. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://csdl.computer.org/dl/trans/td/2004/06/
l0481.htm;   http://csdl.computer.org/dl/trans/td/
2004/06/l0481.pdf.

Moffat:1995:ACR

[MNW95] A. Moffat, R. Neal, and I. H. Witten. Arithmetic coding revisited. In Storer and Cohn [SC95], pages 202–211. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=515510. IEEE catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Moffat:1998:ACR

[MNW98] Alistair Moffat, Radford M. Neal, and Ian H. Witten. Arithmetic coding revisited. ACM Transactions on Information Systems, 16(3):256–294, July 1998. CODEN ATISET. ISSN

1046-8188. URL http://www.acm.org:80/pubs/citations/
journals/tois/1998-16-3/p256-moffat/.

Moreno:2011:ICB

[MO11] J. Moreno and X. Otazu. Image coder based on Hilbert scanning of embedded QuadTrees. In Storer and Marcellin [SM11b], page 470. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

5749527. IEEE Computer Society Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Moayeri:1986:FVQ

[Moa86] Nader Moayeri. Fast Vector Quantization (Communication Theory, Source Coding, Data Compression). Ph.D. thesis, University of Michigan, Ann Arbor, MI, USA, 1986. 178 pp. URL

http://search.proquest.com/docview/303494436.

Moayeri:1998:LCF

[Moa98] Nader Moayeri. A low-complexity, fixed-rate compression scheme for color images and documents. Hewlett–Packard Journal, 50(1):46–??, November 1998. CODEN HPJOAX.

ISSN 0018-1153. URL http://www.hp.com/hpj/98nov/tc-
11-98.htm.

Mock:2004:ME

[Moc04] T. Mock. Music everywhere. IEEE Spectrum, 41(9):42–47, September 2004. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Modha:2003:CPQ

[Mod03] D. S. Modha. Codelet parsing: quadratic-time, sequential, adaptive algorithms for lossy compression. In Storer and Cohn [SC03], pages 223–232. ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2003. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1194013. IEEE Computer Society Order number PR01896.

Moffat:1990:IPD

[Mof90] Alistair Moffat. Implementing the PPM data compression scheme. IEEE Transactions on Communications, COM-38 (11):1917–1921, November 1990. CODEN IECMBT. ISSN 0090-6778 (print), 1558-0857 (electronic).

Moffat:1991:TLC

[Mof91] Alistair Moffat. Two-level context based compression of binary images. In Storer and Reif [SR91b], pages 382–391. ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1990. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=213342. IEEE Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Moffat:1997:ESI

[Mof97] A. Moffat. Editorial for special issue on lossless compression. The Computer Journal, 40(2–3):65–66, ???? 1997. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www.oup.co.uk/computer_journal/Volume_
40/Issue_02/Vol40_02.body.html#AbstractEditorial;
http://www3.oup.co.uk/computer_journal/Volume_40/Issue_
02/Vol40_02.body.html#AbstractEditorial;   http://
www3.oup.co.uk/computer_journal/Volume_40/Issue_02/
Vol40_03.body.html#AbstractEditorial.

Moffat:2000:CIS

[Mof00] Alistair Moffat. Compressing integer sequences and sets. In Ming-Yang Kao, editor, Encyclopedia of Algorithms, page ?? Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000.

Moffat:2019:HC

[Mof19] Alistair Moffat. Huffman coding. ACM Computing Surveys, 52(4):85:1–85:35, September 2019. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic). URL https://dl.

acm.org/ft_gateway.cfm?id=3342555.

Matloub:2005:OQP

[MOG05] S. Matloub, D. B. O’Brien, and R. M. Gray. Optimal quantizer performance and the Wasserstein distortion. In Storer and Cohn [SC05], pages 243–250. ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=1402185.

Mohr:2002:BAS

[Moh02] A. E. Mohr. Bit allocation in sub-linear time and the multiple-choice knapsack problem. In Storer and Cohn [SC02], pages 352–361. ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2002. URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

999973. IEEE Computer Society Order Number PR01477.

Mohammed:2004:HDR

[Moh04] Falah H. M. Ahmad Mohammed. High Data Rate Pulse Compression Using Non-Linear Transmission Line (NLTL) Technology. Ph.D. thesis, Queen’s University of Belfast, Belfast, Northern Ireland, 2004. 249 pp. URL http://search.

proquest.com/docview/305088588.

Murata:2000:WCD

[MOK00] D. Murata, T. Otake, and A. Kawanaka. Wavelet coding of 3-D shape data using space-frequency quantization. In Storer and Cohn [SC00], page ?? ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2000. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=838198. This

millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Monro:1993:GFT

[Mon93] D. M. Monro. Generalized fractal transforms: complexity issues. In Storer and Cohn [SC93], pages 254–261. ISBN 0-8186-3391-3 (microfiche), 0-8186-3392-1 (casebound). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1993. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=253124. IEEE Computer Society Press order number 3392-02. IEEE catalog number 93TH0536-3.

Morrin:1976:CLC

[Mor76] T. H. Morrin, II. Chain-link compression of arbitrary blackwhite images. Computer Graphics and Image Processing, 5(2): 172–189, June 1976. CODEN CGIPBG. ISSN 0146-664X.

Moreno:2012:PPF

[Mor12] J. Moreno. P^{2}SNR: Perceptual full-reference image quality assessment for JPEG2000. In Storer and Marcellin [SM12], page 406. ISBN 1-4673-0715-7 (paperback). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

6189287. IEEE Computer Society order number P4656.

Moss:1998:RDE

[Mos98] K. N. Moss. Recovery from dropout errors in LZ77 compressed text. In Storer and Cohn [SC98], page ?? ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1998. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=672307. IEEE Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Motte:1998:OPP

[Mot98] Warren F. Motte. Oulipo, a Primer of Potential Literature. Daleky Archive Press, Normal, IL, USA, 1998. ?? pp.

Matsuo:2003:IPC

[MOT^{+}03] Yoshihiro Matsuo, Shigeyuki Okada, Kazuhiko Taketa, Tatsushi Ohyama, Yuh Matsuda, Tsugio Mori, Shin’ichiro Okada, Tsuyoshi Watanabe, Tatsuya Ichikawa Yuji Yamada, Hideki Yamauchi, and Yoshifumi Matsushita. An image processor capable of block-noise-free JPEG2000 compression with 30 frames/s for digital camera applications. In Anonymous [Ano03], page ?? ISBN ???? LCCN ????

Motil:2007:PC

[Mot07] John Motil. Private communication, 2007.

Motta:2008:DLC

[MOW08] G. Motta, E. Ordentlich, and M. J. Weinberger. Defect list compression. In Storer and Marcellin [SM08], pages 3–12.

ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33 INTER-

NET. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=4483278.

McIntyre:1985:DCU

[MP85] David R. McIntyre and Michael A. Pechura. Data compression using static Huffman code-decode tables. Communications of the ACM, 28(6):612–616, June 1985. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-0782/3815.html.

Moghaddam:1995:ASM

[MP95] B. Moghaddam and A. Pentland. An automatic system for model-based coding of faces. In Storer and Cohn [SC95], pages 362–370. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=515526. IEEE

catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Myers:2006:PBC

[MP06] Wayne L. Myers and Ganapati P. Patil. Pattern-Based Compression of Multi-Band Image Data for Landscape Analysis, volume 2 of Environmental and Ecological Statistics. Springer Science+Business Media, LLC, Boston, MA, USA, 2006. ISBN 0-387-44434-3, 0-387-44439-4. xiii + 186 pp. LCCN QH541.15 L35 M94 2006.

Mavridis:2012:PTM

[MP12] Pavlos Mavridis and Georgios Papaioannou. Parameterization and texture mapping: Texture compression using wavelet decomposition. Computer Graphics Forum, 31(7pt1):2107–2116, September 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Martinez-Prieto:2010:HOT

[MPAFF10] M. A. Martinez-Prieto, J. Adiego, P. Fuente, and J. D. Fernandez. High-order text compression on hierarchical edgeguided. In Storer and Marcellin [SM10b], page 543. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5453496.

Monteagudo-Pereira:2012:ETJ

[MPALSS^{+}12] J. L. Monteagudo-Pereira, F. Aulí-Llinás, J. Serra-Sagristà, A. Zabala, J. Maso, and X. Pons. Enhanced transmission of JPEG2000 imagery through JPIP proxy and user-navigation model. In Storer and Marcellin [SM12], pages 22–31. ISBN 1-4673-0715-7 (paperback). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=6189233. IEEE Com-

puter Society order number P4656.

Monteagudo-Pereira:2010:SJP

[MPALSSBR10] J. L. Monteagudo-Pereira, F. Aulí-Llinás, J. Serra-Sagristà, and J. Bartrina-Rapesta. Smart JPIP proxy server with prefetching strategies. In Storer and Marcellin [SM10b], pages 99–108. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=5453452.

Martinez-Prieto:2009:UWA

[MPASM^{+}09] M. A. Martinez-Prieto, J. Adiego, F. Sanchez-Martinez, P. Fuente, and R. C. Carrasco. On the use of word alignments to enhance bitext compression. In Storer and Marcellin [SM09], page 459. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=4976513. IEEE

Computer Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

Min:2003:XQC

[MPC03] Jun-Ki Min, Myung-Jae Park, and Chin-Wan Chung. XPRESS: a queriable compression for XML data. In ACM [ACM03], pages 122–133. ISBN 1-58113-634-X. LCCN ????

MPEG:1998:RWM

[MPE00] MPEG. The reference website for MPEG!, 1998, 2000. URL

http://www.mpeg.org/.

Mitchell:1997:MVC

[MPFL97] Joan L. Mitchell, W. B. Pennebaker, C. E. Fogg, and D. J. LeGall, editors. MPEG Video Compression Standard. Chapman and Hall and International Thomson Publishing, New York, NY, USA, 1997. ?? pp.

Munoz-Perez:2019:EBS

[MPGMD19] Luis F. Muñoz-Pérez, José A. Guerrero, and Jorge E. Macías-Díaz. Entropy-based selection of cluster representatives for document image compression. SIAM Journal on Imaging Sciences, 12(4):1720–1738, ???? 2019. CODEN SJISBI. ISSN 1936-4954.

Milidiu:1999:TSE

[MPL99] R. L. Milidiu, A. A. Pessoa, and E. S. Laber. Two spaceeconomical algorithms for calculating minimum redundancy prefix codes. In Storer and Cohn [SC99], pages 267–276. ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1999. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=755676. IEEE Computer Society Or-

der Number PR00096. IEEE Order Plan Catalog Number PR00096.

Marion-Poty:2002:WDS

[MPL02] V. Marion-Poty and Wilfrid Lefer. A wavelet decomposition scheme and compression method for streamline-based vector field visualizations. Computers and Graphics, 26(6):899–906, December 2002. CODEN COGRD2. ISSN 0097-8493 (print),

1873-7684 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0097849302001784.

Markas:1991:ICM

[MR91] T. Markas and J. Reif. Image compression methods with distortion controlled capabilities. In Storer and Reif [SR91b], pages 93–102. ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1990. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=213376. IEEE

Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Markas:1993:MIC

[MR93a] T. Markas and J. Reif. Multispectral image compression algorithms. In Storer and Cohn [SC93], pages 391– 400. ISBN 0-8186-3391-3 (microfiche), 0-8186-3392-1 (casebound). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 1993. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=253110. IEEE Computer

Society Press order number 3392-02. IEEE catalog number 93TH0536-3.

Meyer:1993:PWA

[MR93b] Yves Meyer and Sylvie Roques, editors. Progress in Wavelet Analysis and Applications, Proceedings of the International Conference “Wavelets and Applications,” Toulouse, France, 8– 13 June 1992. Observatoire Midi-Pyrénées de l’Université Paul Sabatier, Editions Frontières, Gif-sur-Yvette, France, 1993. ISBN 2-86332-130-7.

Miller:1993:ISS

[MR93c] D. Miller and K. Rose. An improved sequential search multistage vector quantizer. In Storer and Cohn [SC93], pages 12–21. ISBN 0-8186-3391-3 (microfiche), 0-8186-3392-1 (casebound). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 1993. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=253149. IEEE Computer

Society Press order number 3392-02. IEEE catalog number 93TH0536-3.

Marchand:1997:NIP

[MR97] J. F. P. Marchand and H. E. Rhody. Noncausal image prediction and reconstruction. In Storer and Cohn [SC97], page ?? ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1997. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=582114. IEEE

Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Miguel:2000:PRI

[MR00] A. C. Miguel and E. A. Riskin. Protection of regions of interest against data loss in a generalized multiple description framework. In Storer and Cohn [SC00], page ?? ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2000. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=838209. This millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Manzini:2004:SFD

[MR04] Giovanni Manzini and Marcella Rastero. A simple and fast DNA compressor. Software — Practice and Experience, 34 (14):1397–1411, November 25, 2004. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Murthy:2006:HRA

[MR06] C. R. Murthy and B. D. Rao. High-rate analysis of source coding for symmetric error channels. In Storer and Cohn [SC06], pages 163–172. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ????

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=
&arnumber=1607251.

Mohsenian:1999:SPC

[MRG99] N. Mohsenian, R. Rajagopalan, and C. A. Gonzales. Singlepass constant- and variable-bit-rate MPEG-2 video compression. IBM Journal of Research and Development, 43(4): 489–509, July 1999. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http://www.research.

ibm.com/journal/rd/434/mohsenian.html.

Matela:2011:EJE

[MRH11] Jiří Matela, Vít Rusňák, and Petr Holub. Efficient JPEG2000 EBCOT context modeling for massively parallel architectures. In Storer and Marcellin [SM11b], pages 423–432. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5749500. IEEE Computer Society

Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Mohr:1999:GDP

[MRL99] A. E. Mohr, E. A. Riskin, and R. E. Ladner. Graceful degradation over packet erasure channels through forward error correction. In Storer and Cohn [SC99], pages 92–101. ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1999. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=755658. IEEE Computer Society Or-

der Number PR00096. IEEE Order Plan Catalog Number PR00096.

Matias:1999:EFP

[MRS99] Y. Matias, N. Rajpoot, and S. C. Sahinalp. The effect of flexible parsing for dynamic dictionary-based data compression. In Storer and Cohn [SC99], pages 238–246. ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1999. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=755673. IEEE Computer Society Or-

der Number PR00096. IEEE Order Plan Catalog Number PR00096.

Matias:2001:EFP

[MRS01] Yossi Matias, Nasir Rajpoot, and Cenk Sahinalp. The effect of flexible parsing for dynamic dictionary-based data compression. ACM Journal of Experimental Algorithmics, 6:10:1– 10:??, ???? 2001. CODEN ???? ISSN 1084-6654.

Motta:2003:CHI

[MRS03] G. Motta, F. Rizzo, and J. A. Storer. Compression of hyperspectral imagery. In Storer and Cohn [SC03], pages 333–342. ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2003. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1194024. IEEE Computer Society Order number PR01896.

Mantaci:2005:EBW

[MRS05] S. Mantaci, A. Restivo, and M. Sciortino. An extension of the Burrows–Wheeler transform to k words. In Storer and Cohn [SC05], page ?? ISBN 0-7695-2309-

9. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=1402226.

Motta:2006:HDC

[MRS06] Giovanni Motta, Francesco Rizzo, and James A. (James Andrew) Storer, editors. Hyperspectral Data Compression. Springer Science+Business Media, Inc., Boston, MA, USA, 2006. ISBN 0-387-28579-2 , 0-387-28600-4 (e-book). xi + 417 pp. LCCN TA1632 .H97 2006. URL http://www.loc.

gov/catdir/enhancements/fy0663/2005051678-d.html;
http://www.loc.gov/catdir/enhancements/fy0824/2005051678-

b.html;   http://www.loc.gov/catdir/enhancements/fy0824/
2005051678-t.html.

Monro:1993:ODQ

[MS93] D. M. Monro and B. G. Sherlock. Optimum DCT quantization. In Storer and Cohn [SC93], pages 188–194. ISBN 0-8186-3391-3 (microfiche), 0-8186-3392-1 (casebound). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1993. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=253131. IEEE Computer Society Press order number 3392-02. IEEE catalog number 93TH0536-3.

Memon:1995:ALI

[MS95a] N. D. Memon and K. Sayood. Asymmetric lossless image compression. In Storer and Cohn [SC95], page 457. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=515567. IEEE catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Mohammed:1995:QWC

[MS95b] A. Mohammed and K. Sayood. Quantization of wavelet coefficients for image compression. In Storer and Cohn [SC95], page

483. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=515593. IEEE catalog

number 95TH8037. IEEE Computer Society Press order number PR07010.

Moffat:1996:ECI

[MS96] A. Moffat and L. Stuiver. Exploiting clustering in inverted file compression. In Storer and Cohn [SC96], pages 82–91. ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1996. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=488313. IEEE Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Moni:1997:FAS

[MS97] S. Moni and S. Sista. A framework for application specific image compression. In Storer and Cohn [SC97], page ?? ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1997. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=582119. IEEE Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Moffat:2000:BIC

[MS00] Alistair Moffat and Lang Stuiver. Binary interpolative coding for effective index compression. Information Retrieval, 3(1): 25–47, July 2000. CODEN IFRTFY. ISSN 1386-4564 (print), 1573-7659 (electronic).

Mecking:2002:MDM

[MS02] M. Mecking and T. Stockhammer. Minimizing distortion via multiuser resource allocation. In Storer and Cohn [SC02], page ?? ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2002. URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

1000007. IEEE Computer Society Order Number PR01477.

Mary:2003:SOT

[MS03] D. Mary and D. T. M. Slock. On the suboptimality of orthogonal transforms for single- or multi-stage lossless transform coding. In Storer and Cohn [SC03], pages 303–312. ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 2003. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1194021. IEEE Computer

Society Order number PR01896.

Menon:2005:ISA

[MS05] S. K. Menon and P. Shankar. An instruction set architecture based code compression scheme for embedded processors. In Storer and Cohn [SC05], page ?? ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=1402227.

Modha:2006:MCM

[MS06] D. S. Modha and N. P. Santhanam. Making the correct mistakes. In Storer and Cohn [SC06], pages 302–311. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1607265.

Malvar:2012:PLC

[MS12] H. S. Malvar and G. J. Sullivan. Progressive-to-lossless compression of color-filter-array images using macropixel spectralspatial transformation. In Storer and Marcellin [SM12], pages 3–12. ISBN 1-4673-0715-7 (paperback). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

6189231. IEEE Computer Society order number P4656.

Motta:1999:ALP

[MSC99] G. Motta, J. A. Storer, and B. Carpentieri. Adaptive linear prediction lossless image coding. In Storer and Cohn [SC99], pages 491–500. ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1999. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=755699. IEEE

Computer Society Order Number PR00096. IEEE Order Plan Catalog Number PR00096.

Motta:2000:ISC

[MSC00a] G. Motta, J. A. Storer, and B. Carpentieri. Improving scene cut quality for real-time video decoding. In Storer and Cohn [SC00], pages 470–479. ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2000. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=838187. This

millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Motta:2000:LIC

[MSC00b] Giovanni Motta, James A. Storer, and Bruno Carpentieri. Lossless image coding via adaptive linear prediction and classification. Proceedings of the IEEE, 88(11):1790–1796, November 2000. CODEN IEEPAD. ISSN 0018-9219 (print), 1558-2256 (electronic).

Ma:2017:QAV

[MSE17] Bo Ma, Susanne K. Suter, and Alireza Entezari. Quality assessment of volume compression approaches using isovalue clustering. Computers and Graphics, 63(??):18–27, April 2017. CODEN COGRD2. ISSN 0097-8493 (print), 1873-

7684 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0097849317300158.

Meron:2008:LBR

[MSFZ08] E. Meron, O. Shayevitz, M. Feder, and R. Zamir. A lower bound on the redundancy of arithmetic-type delay constrained coding. In Storer and Marcellin [SM08], pages 489–498. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33 INTER-

NET. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=4483327.

Matela:2012:LGO

[MŠH12] Jiří Matela, Martin Šrom, and Petr Holub. Low GPU occupancy approach to fast arithmetic coding in JPEG2000. Lecture Notes in Computer Science, 7119:136–145, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-25929-6_13/.

Miranda:2009:PVQ

[MSM09] E. Miranda, Guoqiang Shan, and V. Megalooikonomou. Performing vector quantization using reduced data representation. In Storer and Marcellin [SM09], page 462. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4976516. IEEE Computer Society

Order Number P3592. BMS Part Number CFP09DCC-PRT. Murashita:1996:HSS

[MSOY96] K. Murashita, N. Satoh, Y. Okada, and S. Yoshida. High-speed statistical compression using self-organized rules and predetermined code tables. In Storer and Cohn [SC96], page ?? ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1996. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=488381. IEEE Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Mehrotra:2004:SAI

[MSR04] A. Mehrotra, R. Srikanth, and A. G. Ramakrishnan. Shape adaptive integer wavelet transform based coding scheme for 2-D/3-D MR images. In Storer and Cohn [SC04], page ?? ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1281528. IEEE Computer

Society Order Number: P2082.

Merlini:1998:ACA

[MSV98] Donatella Merlini, Renzo Sprugnoli, and M. Cecilia Verri. Average-case analysis for a simple compression algorithm. Algorithmica, 22(4):585–599, December 1998. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

URL http://link.springer.de/link/service/journals/
00453/bibs/22n4p585.html;   http://www.springerlink.
com/openurl.asp?genre=article&issn=0178-4617&volume=

22&issue=4&spage=585. Average-case analysis of algorithms.

Malvar:1996:LOV

[MSW96] H. S. Malvar, G. J. Sullivan, and G. W. Wornell. Lapped orthogonal vector quantization. In Storer and Cohn [SC96], pages 320–329. ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1996. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=488337. IEEE

Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Martin:2004:LTU

[MSW04] A. Martin, G. Seroussi, and M. J. Weinberger. Linear time universal coding of tree sources via FSM closure. In Storer and Cohn [SC04], pages 372–381. ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1281482. IEEE Computer Society Order Number: P2082.

Matthews:2010:TNA

[MSW10] S. J. Matthews, Seung-Jin Sul, and T. L. Williams. TreeZip: a new algorithm for compressing large collections of evolutionary trees. In Storer and Marcellin [SM10b], page 544. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=5453499.

Moffat:1993:EEC

[MSWB93] A. Moffat, N. Sharman, I. H. Witten, and T. C. Bell. An empirical evaluation of coding methods for multi-symbol alphabets. In Storer and Cohn [SC93], pages 108–117. ISBN 0-8186-3391-3 (microfiche), 0-8186-3392-1 (casebound). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1993. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=253139. IEEE Computer Society Press order number 3392-02. IEEE catalog number 93TH0536-3.

Morihara:1998:JTC

[MSYY98] T. Morihara, N. Satoh, H. Yahagi, and S. Yoshida. Japanese text compression using word-based coding. In Storer and Cohn [SC98], page ?? ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1998. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=672306. IEEE

Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Moffat:1994:SCD

[MSZ94] A. Moffat, N. Sharman, and J. Zobel. Static compression for dynamic texts. In Storer and Cohn [SC94], pages 126–135. ISBN 0-8186-5636-0, 0-8186-5637-9 (case). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1994.

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=305920. IEEE Catalog Number 93TH0626-2. IEEE Computer Society Press Order Number 5637-02.

Manohar:1992:PVQ

[MT92] M. Manohar and J. C. Tilton. Progressive vector quantization of multispectral image data using a massively parallel SIMD machine. In Storer and Cohn [SC92], pages 181– 190. ISBN 0-8186-2717-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=227463. IEEE catalog

number 91TH0436-6.

Moffat:1996:IMR

[MT96] A. Moffat and A. Turpin. On the implementation of minimumredundancy prefix codes. In Storer and Cohn [SC96], pages 170–179. ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN QA76.9.D33 D37 1996. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=488322. IEEE

Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Meyer:1998:ETN

[MT98] B. Meyer and P. Tischer. Extending TMW for near lossless compression of greyscale images. In Storer and Cohn [SC98], pages 458–470. ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1998. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=672194. IEEE

Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Moffat:2002:CCA

[MT02] Alistair Moffat and Andrew Turpin. Compression and Coding Algorithms. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2002. ?? pp.

Mallikarachchi:2020:DCR

[MTA^{+}20] Thanuja Mallikarachchi, Dumidu Talagala, Hemantha Kodikara Arachchi, Chaminda Hewage, and Anil Fernando. A decodingcomplexity and rate-controlled video-coding algorithm for HEVC. Future Internet, 12(7):120, July 16, 2020. CODEN

???? ISSN 1999-5903. URL https://www.mdpi.com/1999-
5903/12/7/120.

Mohajer:2008:AML

[MTD08] S. Mohajer, Chao Tian, and S. N. Diggavi. Asymmetric multilevel diversity coding. In Storer and Marcellin [SM08], pages 412–421. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33

INTERNET. URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4483319.

Moffat:1995:SEC

[MTK95] A. Moffat, A. Turpin, and J. Katajainen. Space-efficient construction of optimal prefix codes. In Storer and Cohn [SC95], pages 192–201. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

515509. IEEE catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Mei:2013:NLS

[MTTH13] Tao Mei, Lin-Xie Tang, Jinhui Tang, and Xian-Sheng Hua. Near-lossless semantic video summarization and its applications to video analysis. ACM Transactions on Multimedia Computing, Communications, and Applications, 9(3):16:1– 16:??, June 2013. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Mukerjee:2004:DTL

[Muk04] K. Mukerjee. Decaying tree lossless coding for image and video. In Storer and Cohn [SC04], page ?? ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1281529. IEEE Computer Society Order Number: P2082.

Mullin:1987:ATD

[Mul87] J. K. Mullin. Accessing textual documents using compressed indexes of arrays of small Bloom filters. The Computer Journal, 30(4):343–348, August 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/30/4/343.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_04/tiff/343.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_30/Issue_04/
tiff/344.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_04/tiff/345.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_30/Issue_04/
tiff/346.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_30/Issue_04/tiff/347.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_30/Issue_04/
tiff/348.tif.

Mulcahy:1996:PSW

[Mul96] Colm Mulcahy. Plotting and scheming with wavelets. Mathematics Magazine, 69(5):323–343, December 1996. CODEN MAMGA8. ISSN 0025-570X. URL http://www.spelman.

edu/~colm/csam.ps.

Mulcahy:1997:ICU

[Mul97] Colm Mulcahy. Image compression using the Haar wavelet transform. Spelman Science and Mathematics Journal, 1(1): 22–31, April 1997. ISSN 1098-4593. URL http://www.

spelman.edu/~colm/wav.ps. It has been claimed that any

smart 15-year-old could follow this introduction to wavelets.

Mund:1991:ZLC

[Mun91] Sibylle Mund. Ziv–Lempel complexity for periodic sequences and its cryptographic application. Lecture Notes in Computer Science, 547:114–??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer-

ny.com/link/service/series/0558/bibs/0547/05470114.
htm;   http://link.springer-ny.com/link/service/series/
0558/papers/0547/05470114.pdf.

Murray:1996:SSP

[Mur96] James D. Murray. SPIFF: Still picture interchange file format. Dr. Dobb’s Journal of Software Tools, 21(7):34, 36, 38, 40–41, July 1996. CODEN DDJOEB. ISSN 1044-789X.

Murray:1994:EGF

[Mv94] James D. Murray and William vanRyper. Encyclopedia of Graphics File Formats. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, July 1994. ISBN 1-56592-058-9. xxxii + 894 pp. LCCN T385 .M87 1994. US$59.95. URL http://www.

oreilly.com/catalog/9781565920583.

Mittal:2016:SAA

[MV16] Sparsh Mittal and Jeffrey S. Vetter. A survey of architectural approaches for data compression in cache and main memory systems. IEEE Transactions on Parallel and Distributed Systems, 27(5):1524–1536, May 2016. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL http://www.computer.org/csdl/trans/td/
2016/05/07110612-abs.html.

Mahammad:2020:PAD

[MV20] Farooq Sunar Mahammad and V. Madhu Viswanatham. Performance analysis of data compression algorithms for heterogeneous architecture through parallel approach. The Journal of Supercomputing, 76(4):2275–2288, April 2020. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Mittal:2017:ARD

[MVJ17] Sparsh Mittal, Jeffrey S. Vetter, and Lei Jiang. Addressing read-disturbance issue in STT–RAM by data compression and selective duplication. IEEE Computer Architecture Letters, 16 (2):94–98, July/December 2017. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Miller:1985:VTZ

[MW85] V. S. Miller and M. N. Wegman. Variations on a theme by Ziv and Lempel. In A. Apostolico and Z. Galil, editors, Combinatorial Algorithms on Words, volume F12 of NATO ASI series, pages 131–140. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985.

McMillan:1992:FMR

[MW92] L. McMillan and L. Westover. A forward-mapping realization of the inverse discrete cosine transform. In Storer and Cohn [SC92], pages 219–228. ISBN 0-8186-2717-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ????

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=227459. IEEE catalog number 91TH0436-6.

Memon:1997:RDC

[MW97] N. Memon and X. Wu. Recent developments in contextbased predictive techniques for lossless image compression. The Computer Journal, 40(2–3):127–136, ???? 1997. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/
content/40/2_and_3/127.full.pdf+html;   http://www.
oup.co.uk/computer_journal/Volume_40/Issue_02/Vol40_
02.body.html#AbstractMemon;   http://www3.oup.co.uk/
computer_journal/Volume_40/Issue_02/Vol40_03.body.
html#AbstractMemon.

Moo:1999:RPA

[MW99] P. W. Moo and X. Wu. Resynchronization properties of arithmetic coding. In Storer and Cohn [SC99], page ?? ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1999. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=785697. IEEE Computer Society Or-

der Number PR00096. IEEE Order Plan Catalog Number PR00096.

Martinian:2006:LDC

[MW06] E. Martinian and M. Wainwright. Low density codes achieve the rate-distortion bound. In Storer and Cohn [SC06], pages 153–162. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=1607250.

Majumdar:2010:MCA

[MW10] A. Majumdar and R. K. Ward. A matrix completion approach to reduce energy consumption in wireless sensor networks. In Storer and Marcellin [SM10b], page 542. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5453497.

Marpe:2001:WBV

[MWC01] Detlev Marpe, Thomas Wiegand, and Hans L. Cycon. Wavelet-based video compression using long-term memory motion-compensated prediction and context-based adaptive arithmetic coding. Lecture Notes in Computer Science, 2251: 76–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer-ny.com/
link/service/series/0558/bibs/2251/22510076.htm;
http://link.springer-ny.com/link/service/series/0558/
papers/2251/22510076.pdf.

Ma:2007:IMI

[MWCG07] Jing Ma, Chengke Wu, Dong Chen, and Jie Guo. Interference multispectral image compression with adaptive distortion control in Fourier domain. In Storer and Cohn [SC07], page 395. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=4148796. IEEE Computer

Society Order Number P2791.

Ma:2009:DDP

[MWLW09] Jing Ma, Chengke Wu, Yunsong Li, and Keyan Wang. Dualdirection prediction vector quantization for lossless compression of LASIS data. In Storer and Marcellin [SM09], page 458.

ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=4976512. IEEE Computer

Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

Martinian:2004:SCD

[MWZ04] E. Martinian, G. W. Wornell, and R. Zamir. Source coding with distortion side information at the encoder. In Storer and Cohn [SC04], pages 172–181. ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1281462. IEEE Computer Society Order Number: P2082.

Moran:2016:SCS

[MY16] Shay Moran and Amir Yehudayoff. Sample compression schemes for VC classes. Journal of the ACM, 63(3):21:1–21:??, September 2016. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Myers:1986:DAV

[Mye86] Eugene W. Myers. An O(ND) difference algorithm and its variations. Algorithmica, 1(2):251–266, 1986. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Masala:2004:RDO

[MYRD04] E. Masala, H. Yang, K. Rose, and J. C. De Martin. Ratedistortion optimized slicing, packetization and coding for error resilient video transmission. In Storer and Cohn [SC04], pages 182–191. ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=1281463. IEEE

Computer Society Order Number: P2082.

Mansouri:2020:NLD

[MYS20] Deloula Mansouri, Xiaohui Yuan, and Abdeldjalil Saidani. A new lossless DNA compression algorithm based on a singleblock encoding scheme. Algorithms (Basel), 13(4), April 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https:

//www.mdpi.com/1999-4893/13/4/99.

Minami:1991:OAR

[MZ91] S. Minami and A. Zakhor. An optimization approach for removing blocking effects in transform coding. In Storer and Reif [SR91b], page ?? ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1990. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=213319. IEEE

Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Moffat:1992:CCF

[MZ92] A. Moffat and J. Zobel. Coding for compression in full-text retrieval systems. In Storer and Cohn [SC92], pages 72– 81. ISBN 0-8186-2717-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=227474. IEEE catalog

number 91TH0436-6.

Mehes:1995:PAI

[MZ95] A. Mehes and K. Zeger. On the performance of affine index assignments for redundancy free source-channel coding. In Storer and Cohn [SC95], page 433. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=515543. IEEE catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Moffat:1996:SII

[MZ96] Alistair Moffat and Justin Zobel. Self-indexing inverted files for fast text retrieval. ACM Transactions on Information Systems, 14(4):349–379, October 1996. CODEN ATISET. ISSN

1046-8188. URL http://www.acm.org:80/tois/abstracts/
moffat.html.

Mehes:1999:PQN

[MZ99] A. Mehes and K. Zeger. Performance of quantizers on noisy channels using structured families of codes. In Storer and Cohn [SC99], pages 473–482. ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1999. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=755697. IEEE

Computer Society Order Number PR00096. IEEE Order Plan Catalog Number PR00096.

Mamou:2006:SAD

[MZP06] Khaled Mamou, Titus Zaharia, and Françoise Prêteux. A skinning approach for dynamic 3D mesh compression. Computer Animation and Virtual Worlds, 17(3–4):337–346, ???? 2006. CODEN ???? ISSN 1546-427X.

Mathew:2012:HSC

[MZT12] R. Mathew, P. Zanuttigh, and D. Taubman. Highly scalable coding of depth maps with arc breakpoints. In Storer and Marcellin [SM12], pages 42–51. ISBN 1-4673-0715-7 (paperback). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=6189235. IEEE Computer Society order number P4656.

Nait-Ali:2008:CIS

[NACM08] Amine Nait-Ali and Christine Cavaro-Menard, editors. Compression des images et des signaux medicaux. (French) [Compression of biomedical images and signals]. ISTE, London, UK, 2008. ISBN 1-84821-028-0. ???? pp. LCCN RC78.7.D35 C63813 2008. URL http://www.loc.

gov/catdir/enhancements/fy0808/2008003130-b.html;
http://www.loc.gov/catdir/enhancements/fy0808/2008003130-
d.html;   http://www.loc.gov/catdir/enhancements/fy0808/
2008003130-t.html.

Nadim:1983:DCT

[Nad83] Galal Ezz E. Nadim. Data compression techniques as applied to the electroencephalograph signal. M.Sc. thesis, The University of Manitoba, Winnipeg, MB, Canada, 1983.

???? pp. URL http://search.proquest.com/docview/
303201518?accountid=14677. 2 microfiches (150 frames).

Na:2003:TST

[NAIP03] Joong Chae Na, Alberto Apostolico, Costas S. Iliopoulos, and Kunsoo Park. Truncated suffix trees and their application to data compression. Theoretical Computer Science, 304(1–3): 87–101, July 28, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Namphol:1991:NNB

[Nam91] Aran Namphol. Neural network based data compression. Ph.D. thesis, The Catholic University of America, Washington, DC,

USA, 1991. 276 pp. URL http://search.proquest.com/
docview/303929431.

Nasim:2013:RVL

[Nas13] Farhan Nasim. Review of Variable-length codes for data compression by David Salomon. ACM SIGACT News, 44(4):24–26, December 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Natarajan:1993:FRN

[Nat93] B. K. Natarajan. Filtering random noise via data compression. In Storer and Cohn [SC93], pages 60–69. ISBN 0-8186-3391-3 (microfiche), 0-8186-3392-1 (casebound). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1993. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=253144. IEEE Computer Society Press order number 3392-02. IEEE catalog number 93TH0536-3.

Natarajan:1994:SBO

[Nat94] B. K. Natarajan. Sharper bounds on Occam filters and application to digital video. In Storer and Cohn [SC94], pages 440–449. ISBN 0-8186-5636-0, 0-8186-5637-9 (case). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 1994. URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=305951. IEEE Cata-

log Number 93TH0626-2. IEEE Computer Society Press Order Number 5637-02.

Natarajan:1995:VST

[Nat95] K. Natarajan. Video servers take root. IEEE Spectrum, 32(4): 66–69, April 1995. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Navarro:2001:RES

[Nav01] Gonzalo Navarro. Regular expression searching over Ziv– Lempel compressed text. Lecture Notes in Computer Science, 2089:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer-

ny.com/link/service/series/0558/bibs/2089/20890001.
htm;   http://link.springer-ny.com/link/service/series/
0558/papers/2089/20890001.pdf.

Navarro:2002:ITU

[Nav02] Gonzalo Navarro. Indexing text using the Ziv–Lempel trie. Lecture Notes in Computer Science, 2476:325–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.de/link/
service/series/0558/bibs/2476/24760325.htm;   http:
//link.springer.de/link/service/series/0558/papers/
2476/24760325.pdf.

Nakamura:2007:SLT

[NBIT07] R. Nakamura, H. Bannai, S. Inenaga, and M. Takeda. Simple linear-time off-line text compression by longest-first substitution. In Storer and Cohn [SC07], pages 123–132. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4148751. IEEE Computer Society

Order Number P2791.

Nathan:2016:SRO

[NBK16] Senthil Nathan, Umesh Bellur, and Purushottam Kulkarni. On selecting the right optimizations for virtual machine migration. ACM SIGPLAN Notices, 51(7):37–49, July 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ng:1997:WBM

[NC97] K. S. Ng and L. M. Cheng. Word based multiple dictionary scheme for text compression with application to 2D bar code. In Storer and Cohn [SC97], page ?? ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1997. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=582120. IEEE Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Nagarkar:2015:CSH

[NCB15] Parth Nagarkar, K. Selçuk Candan, and Aneesha Bhat. Compressed spatial hierarchical bitmap (cSHB) indexes for efficiently processing spatial range query workloads. Proceedings of the VLDB Endowment, 8(12):1382–1393, August 2015. CO-DEN VLDBFR. ISSN 2150-8097.

Neogi:2002:CTA

[NcC02] A. Neogi and Tzi cker Chiueh. Compression techniques for active video content. In Storer and Cohn [SC02], page ?? ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2002. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=1000009. IEEE

Computer Society Order Number PR01477.

Nguyen:2004:RCV

[ND04] H. Nguyen and P. Duhamel. Reduced complexity VLC sequence decoder. In Storer and Cohn [SC04], page ?? ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1281530. IEEE Computer Society

Order Number: P2082.

Natarajan:2013:ACIb

[NDN13a] Meghanathan Natarajan, Nagamalai Dhinaharan, and Chaki Nabendu, editors. Advances in computing and information technology: Proceedings of the Second International Conference on Advances in Computing and Information Technology (ACITY), July 13–15, 2012, Chennai, India, Volume 2, volume 177 of Advances in intelligent systems and computing. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 3-642-31551-8, 3-642-31552-6 (e-book). LCCN ????

Natarajan:2013:ACIc

[NDN13b] Meghanathan Natarajan, Nagamalai Dhinaharan, and Chaki Nabendu, editors. Advances in computing and information technology: Proceedings of the Second International Conference on Advances in Computing and Information Technology (ACITY), July 13–15, 2012, Chennai, India, Volume 3, volume 178 of Advances in intelligent systems and computing. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 3-642-31599-2, 3-642-31600-X (e-book). LCCN ????

Neve:2000:ACR

[NDPL00] P. De Neve, K. Denecker, W. Philips, and I. Lemahieu. An advanced color representation for lossy compression of CMYK prepress images. Computer Graphics Forum, 19(1): ??, March 2000. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). URL aid=384&amp; http://www.

blackwellpublishers.co.uk/asp/journal.asp?ref=0167-
7055&amp;   iid=1&amp; src=ard&amp; vid=19.

Nuallain:2004:OST

[NdR04] B. O. Nuallain and S. de Rooij. Online suffix trees with counts. In Storer and Cohn [SC04], page ?? ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1281531. IEEE Computer Society

Order Number: P2082.

Nastea:1996:PIO

[NEGF96] S. G. Nastea, T. El-Ghazawi, and O. Frieder. Parallel input/ output impact on sparse matrix compression. In Storer and Cohn [SC96], page ?? ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1996. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=488382. IEEE

Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Nekritch:2000:DCH

[Nek00] Y. Nekritch. Decoding of canonical Huffman codes with lookup tables. In Storer and Cohn [SC00], page ?? ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2000. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=838213. This millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Nekrich:2007:EIA

[Nek07] Y. Nekrich. An efficient implementation of adaptive prefix coding. In Storer and Cohn [SC07], page 396. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4148797. IEEE Computer Society

Order Number P2791.

Nelson:1987:AWM

[Nel87] Russell Nelson. ARC wars: MS-DOS archiving utilities. Dr. Dobb’s Journal of Software Tools, 12(3):26–28, 30, March 1987. CODEN DDJOEB. ISSN 0888-3076.

Nelson:1989:LDC

[Nel89] Mark R. Nelson. LZW data compression. Dr. Dobb’s Journal of Software Tools, 14(10):29, 32, 34, 36, 86–87, October 1989. CODEN DDJOEB. ISSN 0888-3076.

Nelson:1991:DDC

[Nel91] Mark R. Nelson. DDJ data compression contest results. Dr. Dobb’s Journal of Software Tools, 16(11):62–??, November 1991. CODEN DDJOEB. ISSN 1044-789X.

Nelson:1992:DCB

[Nel92] Mark Nelson. The data compression book: featuring fast, efficient data compression techniques in C. M&T Books, M&T Publishing, Inc., 501 Galveston Drive, Redwood City, CA 94063, USA, 1992. ISBN 1-55851-214-4, 1-55851-216-0 (Book disk set). 527 pp. LCCN QA76.9.D33 N46 1992.

Nelson:2002:IIJ

[Nel02] Mark R. Nelson. Inside Intel’s JPEG library. Dr. Dobb’s Journal of Software Tools, 27(7):38, 40, 42, 44, July 2002. CODEN DDJOEB. ISSN 1044-789X. URL http://www.ddj.com/ftp/

2002/2002_07/slide.zip.

Neuburger:2010:BRB

[Neu10] Shoshana Neuburger. Book review: The Burrows–Wheeler Transform: Data Compression, Suffix Arrays, and Pattern Matching, by Donald Adjeroh, Timothy Bell and Amar Mukherjee Springer, 2008. ACM SIGACT News, 41(1):21– 24, March 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [ABM08].

Nelson:1996:DCB

[NG96a] Mark Nelson and Jean-Loup Gailly. The data compression book. M&T Books, M&T Publishing, Inc., 501 Galveston Drive, Redwood City, CA 94063, USA, second edition, 1996. ISBN 1-55851-434-1. xiii + 557 pp. LCCN QA76.9.D33 N46

Ng:1996:LCV

[Ng96b] Kent Wai Ng. Lossless compression of voiceband data signals. Dissertation/thesis, The University of Manitoba, Winnipeg,

MB, Canada, 1996. 79 pp. URL http://search.proquest.
com/docview/304322906.

Ng:2010:PSI

[Ng10] Jason Ng, editor. Practical signal and image processing in clinical cardiology. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 1-84882-515-3, 1-84882-514-5. xv + 400 pp. LCCN RC683 .P73 2010.

Ngai:1995:VAF

[Nga95] A. Ngai. VLSI architecture of the I-frame encoder for the MPEG-2 video compression. In IEEE [IEE95a], pages 103– 110. ISBN ???? LCCN ????

Niedermeier:2002:MBF

[NHHS02] U. Niedermeier, J. Heuer, A. Hutter, and W. Stechele. MPEG-7 binary format for XML data. In Storer and Cohn [SC02], page ?? ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2002. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1000010. IEEE Computer Society Order Number PR01477.

NHK:2006:NOE

[NHK06] NHK. NHK Online English, 2006. URL http://www.nhk.or.

jp/english/.

Nakamura:2009:LTT

[NIB^{+}09] Ryosuke Nakamura, Shunsuke Inenaga, Hideo Bannai, Takashi Funamoto, Masayuki Takeda, and Ayumi Shinohara. Lineartime text compression by longest-first substitution. Algorithms (Basel), 2(4):1429–1448, December 2009. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/2/4/1429.

Niu:2002:DBE

[Niu02] Yan Niu. DWT-based error concealment techniques. In Storer and Cohn [SC02], page ?? ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2002. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1000023. IEEE Computer Society Order Number PR01477.

Nix:1981:ESE

[Nix81] Robert P. Nix. Experience with a space efficient way to store a dictionary. Communications of the ACM, 24(5):297–298, 1981. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Novoselsky:2016:RAD

[NK16] Alexander Novoselsky and Eugene Kagan. Remark on “Algorithm 673: Dynamic Huffman Coding”. ACM Transactions on Mathematical Software, 42(1):10:1, February 2016. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [Vit89].

Newman:2020:NTP

[NK20] Elizabeth Newman and Misha E. Kilmer. Nonnegative tensor patch dictionary approaches for image compression and deblurring applications. SIAM Journal on Imaging Sciences, 13 (3):1084–1112, ???? 2020. CODEN SJISBI. ISSN 1936-4954.

Na:2016:FIA

[NKP^{+}16] Joong Chae Na, Hyunjoon Kim, Heejin Park, Thierry Lecroq, Martine Léonard, Laurent Mouchard, and Kunsoo Park. FMindex of alignment: a compressed index for similar strings. Theoretical Computer Science, 638(??):159–170, July 25, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304397515007288.

Navarro:2001:FAS

[NKT^{+}01] G. Navarro, T. Kida, M. Takeda, A. Shinohara, and S. Arikawa. Faster approximate string matching over compressed text. In Storer and Cohn [SC01c], pages 459–468. ISBN 0-7695-1031-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2001. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=917177. IEEE

CSP number 01PR1031.

Netravali:1980:PCP

[NL80] A. Netravali and J. O. Limb. Picture coding: a preview. Proceedings of the IEEE, 68:366–406, 1980. CODEN IEEPAD. ISSN 0018-9219 (print), 1558-2256 (electronic).

Nagy:2003:ESB

[NL03] D. A. Nagy and T. Linder. Experimental study of a binary block sorting compression scheme. In Storer and Cohn [SC03], page ?? ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2003. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1194058. IEEE Computer Society Order number PR01896.

Nielsen:2006:MLC

[NL06] C. Nielsen and Xiobo Li. MST for lossy compression of image sets. In Storer and Cohn [SC06], page 463. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1607306.

Neidecker-Lutz:1993:SMP

[NLU93] Burkhard K. Neidecker-Lutz and Robert Ulichney. Software motion pictures. Digital Technical Journal, 5(2):19–27, Spring 1993. CODEN DTJOEL. ISSN 0898-901X. URL ftp://

ftp.digital.com/pub/Digital/info/DTJ/v5n2/Software_
Motion_Pictures_01oct1993DTJA02P8.ps;   http://www.
digital.com:80/info/DTJA02/DTJA02SC.TXT.

Nagumo:1995:PAS

[NLW95] H. Nagumo, Mi Lu, and K. Watson. Parallel algorithms for the static dictionary compression. In Storer and Cohn [SC95], pages 162–171. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

515506. IEEE catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Nagumo:1996:LLF

[NLW96] Hideo Nagumo, Mi Lu, and Karan Watson. On-line longest fragment first parsing algorithm. Information Processing Letters, 59(2):91–96, July 22, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nagumo:1999:PPA

[NLW99] H. Nagumo, M. Lu, and K. L. Watson. Parallel parsing algorithms for static dictionary compression. IEEE Transactions on Parallel and Distributed Systems, 10(12): 241–??, December 1999. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://dlib.

computer.org/td/books/td1999/pdf/l1241.pdf;   http:
//www.computer.org/tpds/td1999/l1241abs.htm.

Nakamura:1991:DCB

[NM91a] Hirofumi Nakamura and Sadayuki Murashima. The data compression based on concatenation of frequentative code neighbor. In Proceedings of the 14th Symposium on Information Theory and its Applications (SITA ’91), December 11–14, 1991, Ibusuki, Japan, pages 701–704. ????, ????, December 11, 1991. In Japanese.

Nguyen:1991:CSP

[NM91b] H. Nguyen and J. W. Mark. Concentric-shell partition vector quantization with application to image coding. In Storer and Reif [SR91b], pages 119–128. ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1990. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=213379. IEEE Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Nakamura:1996:DCC

[NM96a] Hirofumi Nakamura and Sadayuki Murashima. Data compression by concatenations of symbol pairs. In Proceedings of the IEEE International Symposium on Information Theory and its Applications, Victoria, BC, Canada, pages 496–499. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, September 1996.

Nevill-Manning:1996:ISS

[NM96b] C. G. Nevill-Manning. Inferring Sequential Structure. Ph.D. thesis, Department of Computer Science, University of Waikato, Waikato, NZ, 1996. ?? pp.

Navarro:2007:CFT

[NM07] Gonzalo Navarro and Veli Mäkinen. Compressed full-text indexes. ACM Computing Surveys, 39(1):2:1–2:61, April 2007.

CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

Nomura:2010:OPF

[NM10] R. Nomura and T. Matsushima. On the overflow probability of fixed-to-variable length codes with side information. In Storer and Marcellin [SM10b], page 548. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=5453525.

Nagano:1999:EIM

[NMN99] H. Nagano, A. Matsuura, and A. Nagoya. An efficient implementation method of fractal image compression on dynamically reconfigurable architecture. Lecture Notes in Computer Science, 1586:670–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nevill-Manning:1997:LTI

[NMW97a] C. G. Nevill-Manning and I. H. Witten. Linear-time, incremental hierarchy inference for compression. In Storer and Cohn [SC97], pages 3–11. ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1997. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=581951. IEEE

Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Nevill-Manning:1997:CEU

[NMW97b] C. G. Nevill-Manning and Ian H. Witten. Compression and explanation using hierarchical grammars. The Computer Journal, 40(2–3):103–116, ???? 1997. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/40/2_and_3/103.
full.pdf+html;   http://www.oup.co.uk/computer_journal/
Volume_40/Issue_02/Vol40_02.body.html#AbstractNevill;
 http://www3.oup.co.uk/computer_journal/Volume_40/
Issue_02/Vol40_02.body.html#AbstractNevill;   http:
//www3.oup.co.uk/computer_journal/Volume_40/Issue_
02/Vol40_03.body.html#AbstractNevill.

Nevill-Manning:1998:PHI

[NMW98] C. G. Nevill-Manning and I. H. Witten. Phrase hierarchy inference and compression in bounded space. In Storer and Cohn [SC98], pages 179–188. ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1998. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=672146. IEEE Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Nevill-Manning:1999:PI

[NMW99] C. G. Nevill-Manning and I. H. Witten. Protein is incompressible. In Storer and Cohn [SC99], pages 257–266. ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1999. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=755675. IEEE Computer Society Or-

der Number PR00096. IEEE Order Plan Catalog Number PR00096.

Nevill-Manning:1994:CIH

[NMWM94] C. G. Nevill-Manning, I. H. Witten, and D. L. Maulsby. Compression by induction of hierarchical grammars. In Storer and Cohn [SC94], pages 244–253. ISBN 0-8186-5636-0, 0-8186-5637-9 (case). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1994. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=305932. IEEE

Catalog Number 93TH0626-2. IEEE Computer Society Press Order Number 5637-02.

Nevill-Manning:1996:CSS

[NMWO96] C. G. Nevill-Manning, I. H. Witten, and D. R. Olsen, Jr. Compressing semi-structured text using hierarchical phrase identifications. In Storer and Cohn [SC96], pages 63–72. ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1996. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=488311. IEEE Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Nelson:1997:JZF

[NN97] Mark R. Nelson and Mark Nelson. Java and the zip file format: Using Java to compress and extract files. Dr. Dobb’s Journal of Software Tools, 22(12):50, 52–54, 102, December 1997. CODEN DDJOEB. ISSN 1044-789X.

Nosratinia:1995:NRO

[NO95] A. Nosratinia and M. T. Orchard. New relationships in operator-based backward motion compensation. In Storer and Cohn [SC95], pages 391–400. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ????

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=515529. IEEE catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Nash:1996:PVQ

[NOG96] C. L. Nash, R. A. Olshen, and R. M. Gray. Predictive vector quantization with ridge regression. In Storer and Cohn [SC96], pages 310–319. ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1996. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=488336. IEEE

Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Nosratinia:1997:PWM

[Nos97] A. Nosratinia. Parametric warping for motion estimation. In Storer and Cohn [SC97], page ?? ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1997. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=582124. IEEE Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Nosratinia:1999:EPP

[Nos99] A. Nosratinia. Embedded post-processing for enhancement of compressed images. In Storer and Cohn [SC99], pages 62–71. ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1999. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=755655. IEEE Computer Society Or-

der Number PR00096. IEEE Order Plan Catalog Number PR00096.

Novak:1998:TER

[Nov98] M. Novak. Transmission error robust fractal coding using a model-residual approach. In Storer and Cohn [SC98], pages 349–358. ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1998. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=672164. IEEE

Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Novak:2000:CIR

[Nov00] M. Novak. Coding of image residuals with tailbiting convolutional codes and BCJR decoding. In Storer and Cohn [SC00], page ?? ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2000. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=838214. This

millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Na:2000:DCT

[NP00] Joong Chae Na and Kunsoo Park. Data compression with truncated suffix trees. In Storer and Cohn [SC00], page ?? ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2000. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=838212. This millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Navarro:2016:FCS

[NP16] Gonzalo Navarro and Alberto Ordóñez Pereira. Faster compressed suffix trees for repetitive collections. ACM Journal of Experimental Algorithmics, 21(1):1.8:1–1.8:??, November 2016. CODEN ???? ISSN 1084-6654.

Navarro:2015:GDR

[NPV15] Gonzalo Navarro, Simon J. Puglisi, and Daniel Valenzuela. General document retrieval in compact space. ACM Journal of Experimental Algorithmics, 19(??):2.3:1–2.3:??, February 2015. CODEN ???? ISSN 1084-6654.

Nguyen-Phi:1997:BLI

[NPW97a] K. Nguyen-Phi and H. Weinrichter. Bi-level image compression using adaptive tree model. In Storer and Cohn [SC97], page ?? ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1997. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=582122. IEEE

Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Nguyen-Phi:1997:FIT

[NPW97b] K. Nguyen-Phi and H. Weinrichter. Fast implementation of two-level compression method using QM-coder. In Storer and Cohn [SC97], page ?? ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1997. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=582123. IEEE

Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Ng:1995:VQL

[NR95] W. K. Ng and C. V. Ravishankar. Vector quantization for lossless textual data compression. In Storer and Cohn [SC95], page 474. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=515584. IEEE catalog

number 95TH8037. IEEE Computer Society Press order number PR07010.

Navarro:1999:GPA

[NR99] Gonzalo Navarro and Mathieu Raffinot. A general practical approach to pattern matching over Ziv–Lempel compressed text. Lecture Notes in Computer Science, 1645:14–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer-ny.com/link/
service/series/0558/bibs/1645/16450014.htm;   http:
//link.springer-ny.com/link/service/series/0558/papers/
1645/16450014.pdf.

Navarro:2008:RPA

[NR08] G. Navarro and L. Russo. Re-pair achieves high-order entropy. In Storer and Marcellin [SM08], page 537. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33 INTER-

NET. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=4483364.

Nanjundaswamy:2014:CLT

[NR14] T. Nanjundaswamy and K. Rose. Cascaded long term prediction for enhanced compression of polyphonic audio signals. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 22(3):697–710, March 2014. CODEN ???? ISSN 2329-9290.

Nguyen:2001:RHQ

[NS01] Ky Giang Nguyen and Dietmar Saupe. Rapid high quality compression of volume data for visualization. Computer Graphics Forum, 20(3):??, September 2001. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ndzana:2006:BWT

[NSA06] B. N. Ndzana, A. Shokrollahi, and J. Abel. Burrows–Wheeler text compression with fountain codes. In Storer and Cohn [SC06], page 462. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=
1607305.

Nieto-Santisteban:1999:DCN

[NSFO^{+}99] M. A. Nieto-Santisteban, D. J. Fixsen, J. D. Offenberg, R. J. Hanisch, and H. S. Stockman. Data compression for the next-generation space telescope. In Storer and Cohn [SC99], page ?? ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1999. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=785699. IEEE

Computer Society Order Number PR00096. IEEE Order Plan Catalog Number PR00096.

Nickels:1991:SDA

[NT91] K. Nickels and C. Thacker. Satellite data archives algorithm. In Storer and Reif [SR91b], page ?? ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1990. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=213314. IEEE Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Nilsson:2002:ESC

[NT02] S. Nilsson and M. Tikkanen. An experimental study of compression methods for dynamic tries. Algorithmica, 33(1):19–33, January 2002. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0178-4617&volume=

33&issue=1&spage=19. Experimental algorithms.

Navarro:2005:LBM

[NT05a] Gonzalo Navarro and Jorma Tarhio. LZgrep: a Boyer–Moore string matching tool for Ziv–Lempel compressed text. Software — Practice and Experience, 35(12):1107–1130, October 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X

(electronic). URL http://www.dcc.uchile.cl/~gnavarro/
pubcode/.

Nourani:2005:RHE

[NT05b] Mehrdad Nourani and Mohammad H. Tehranipour. RL-Huffman encoding for test compression and power reduction in scan applications. ACM Transactions on Design Automation of Electronic Systems (TODAES), 10(1):91–115, January 2005. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Niemi:2020:BWP

[NT20] Arto Niemi and Jukka Teuhola. Burrows–Wheeler posttransformation with effective clustering and interpolative coding. Software — Practice and Experience, 50(9):1858–1874, September 2020. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

NTFS:2006:NCH

[NTF06] NTFS. NTFS.com home page: Complete info source: NTFS & FAT file systems and data recovery, 2006. URL http://

www.ntfs.com/.

Norcen:2005:HPJ

[NU05] Roland Norcen and Andreas Uhl. High performance JPEG 2000 and MPEG-4 VTC on SMPs using OpenMP. Parallel Computing, 31(10–12):1082–1098, October/December 2005. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Nussbaum:1976:SCA

[Nus76] Howard Steven Nussbaum. Source Coding and Adaptive Data Compression for Communication Networks. Ph.D. thesis, University of California, Los Angeles, Los Angeles, CA, USA,

1976. 220 pp. URL http://search.proquest.com/docview/
302826487.

Nand:1998:MSE

[NY98] A. Nand and Tong Lai Yu. Mail servers with embedded data compression mechanisms. In Storer and Cohn [SC98], page ?? ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1998. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=672308. IEEE

Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Nunez-Yanez:2005:CSL

[NYC05] J. L. Nunez-Yanez and V. A. Chouliaras. A configurable statistical lossless compression core based on variable order Markov modeling and arithmetic coding. IEEE Transactions on Computers, 54(11):1345–1359, November 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=1514415.

Nejati:2008:WVT

[NYJ08] N. Nejati, H. Yousefi’zadeh, and H. Jafarkhani. Wireless video transmission: a distortion-optimal approach. In Storer and Marcellin [SM08], pages 202–211. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33 INTERNET.

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=
&arnumber=4483298.

Nejati:2009:WVT

[NYJ09] N. Nejati, H. Yousefi’zadeh, and H. Jafarkhani. Wireless video transmission: a single layer distortion optimal approach. In Storer and Marcellin [SM09], pages 362–371. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=4976480. IEEE Computer Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

Nakano:1994:HEU

[NYOY94] Y. Nakano, H. Yahagi, Y. Okada, and S. Yoshida. Highly efficient universal coding with classifying to subdictionaries for text compression. In Storer and Cohn [SC94], pages 234–243. ISBN 0-8186-5636-0, 0-8186-5637-9 (case). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1994.

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=305931. IEEE Catalog Number 93TH0626-2. IEEE Computer Society Press Order Number 5637-02.

Nyquist:1928:CTT

[Nyq28] Harry Nyquist. Certain topics in telegraph transmission theory. Transactions of the American Institute of Electrical Engineers, 47:617–644, 1928. CODEN TAEEA5. ISSN 0096-3860.

URL http://www.xiph.org/ogg/flac.html.

Neff:1995:MPV

[NZ95] R. Neff and A. Zakhor. Matching pursuit video coding at very low bit rates. In Storer and Cohn [SC95], pages 411– 420. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=515531. IEEE catalog

number 95TH8037. IEEE Computer Society Press order number PR07010.

Nong:2006:UBW

[NZ06] Ge Nong and Sen Zhang. Unifying the Burrows–Wheeler and the Schindler transforms. In Storer and Cohn [SC06], page 464. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=1607307.

Nong:2007:EAI

[NZ07] Ge Nong and Sen Zhang. An efficient algorithm for the inverse ST problem. In Storer and Cohn [SC07], page 397. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=4148798. IEEE Computer

Society Order Number P2791.

Nong:2009:LSA

[NZC09] Ge Nong, Sen Zhang, and Wai Hong Chan. Linear suffix array construction by almost pure induced-sorting. In Storer and Marcellin [SM09], pages 193–202. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=4976463. IEEE Computer Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

Olano:2011:ETV

[OBGB11] M. Olano, D. Baker, W. Griffin, and J. Barczak. Efficient textures: Variable bit rate GPU texture decompression. Computer Graphics Forum, 30(4):1299–1308, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Oehler:1993:CIC

[OG93] K. L. Oehler and R. M. Gray. Combining image classification and image compression using vector quantization. In Storer and Cohn [SC93], pages 2–11. ISBN 0-8186-3391-3 (microfiche), 0-8186-3392-1 (casebound). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1993. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=253150. IEEE Computer Society Press order number 3392-02. IEEE catalog number 93TH0536-3.

Ozonat:2004:ICU

[OG04] K. M. Ozonat and R. M. Gray. Image classification using adaptive-boosting and tree-structured discriminant vector quantization. In Storer and Cohn [SC04], page ?? ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1281532. IEEE Computer Society

Order Number: P2082.

Ozonat:2005:VQC

[OG05] K. M. Ozonat and R. M. Gray. Vector quantization for classification in a simple network. In Storer and Cohn [SC05], page ?? ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1402229.

Ozonat:2006:GMM

[OG06] K. Ozonat and R. M. Gray. Gauss mixture model-based classification for sensor networks. In Storer and Cohn [SC06], pages 322–331. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=1607267.

Osorio:2012:FCN

[OG12] R. R. Osorio and P. Gonzalez. Fast construction of nearlyoptimal prefix codes without probability sorting. In Storer and Marcellin [SM12], page 407. ISBN 1-4673-0715-7 (paperback). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=6189288. IEEE Computer Society order number P4656.

Osorio:2016:FAC

[OG16] Roberto R. Osorio and Patricia González. A fast algorithm for constructing nearly optimal prefix codes. Software — Practice and Experience, 46(10):1299–1316, October 2016. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Oh:2010:JJI

[Oh10] Tick H. Oh. JPEG2000 and JPEG: Image quality comparison of compressed medical modalities. International Journal of Computer Applications, 32(4):393–398, 2010. CO-DEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.tandfonline.com/doi/full/10.
2316/Journal.202.2010.4.202-1503.

Ohio-State:2001:U

[Ohi01] Ohio-State. [unknown], 2001. URL http://www.cis.ohio-

state.edu/htbin/rfc/rfc804.html.

Oswald:2021:END

[OHRS21] C. Oswald, E. Haritha, A. Akash Raja, and B. Sivaselvan. An efficient and novel data clustering and run length encoding approach to image compression. Concurrency and Computation: Practice and Experience, 33(10):e6185:1–e6185:??, May 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

ONeill:2012:ACT

[OHSS12] A. O’Neill, M. Hutter, Wen Shao, and P. Sunehag. Adaptive context tree weighting. In Storer and Marcellin [SM12], pages 317–326. ISBN 1-4673-0715-7 (paperback). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

6189263. IEEE Computer Society order number P4656.

Odagiri:2010:DCT

[OINC10] J. Odagiri, N. Itani, Y. Nakano, and D. E. Culler. Data compression technology dedicated to distribution and embedded systems. In Storer and Marcellin [SM10b], page 549. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5453524.

Ostergaard:2005:CSM

[ØJH05] J. Østergaard, J. Jensen, and R. Heusdens. n-channel symmetric multiple-description lattice vector quantization. In Storer and Cohn [SC05], pages 378–387. ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=1402199.

Okanohara:2005:PDC

[Oka05] D. Okanohara. Partially decodable compression with static PPM. In Storer and Cohn [SC05], page ?? ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=1402228.

Okumura:1998:HDC

[Oku98]  Haruhiko Okumura.  History of data compression in

Japan, 1998. URL http://oku.edu.mie-u.ac.jp/~okumura/

compression.html;   http://oku.edu.mie-u.ac.jp/~okumura/
compression/history.html.

Ou:2008:LIT

[OLHL08] Chien-Min Ou, Hui-Ya Li, Wen-Jyi Hwang, and Mei-Hwa Liu. Layered image transmission with quality pre-specifiable JPEG2000. International Journal of Image and Graphics (IJIG), 8(4):629–641, October 2008. CODEN ???? ISSN 0219-4678.

Oliver:2003:FES

[OM03] J. Oliver and M. P. Malumbres. Fast and efficient spatial scalable image compression using wavelet lower trees. In Storer and Cohn [SC03], pages 133–142. ISBN 0-7695-1896-

6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 2003. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1194004. IEEE Computer

Society Order number PR01896.

Ota:2007:CAB

[OM07] Takahiro Ota and Hiroyoshi Morita. On the construction of an antidictionary of a binary string with linear complexity. IEICE Transactions on Fundamentals of Electronics Communications and Computer Sciences, E90-A(11):2533–2539, November 2007. CODEN IFESEX. ISSN 0916-8508 (print), 1745-1337 (electronic).

Omodeo:1991:IJA

[OPS91] A. Omodeo, M. Pugassi, and N. Scarabottolo. Implementing JPEG algorithm on INMOS transputer equipped machines. In Storer and Reif [SR91b], page ?? ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1990. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=213326. IEEE Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Orchard:1994:IWB

[OR94] M. T. Orchard and K. Ramchandran. An investigation of wavelet-based image coding using an entropy-constrained quantization framework. In Storer and Cohn [SC94], pages 341–350. ISBN 0-8186-5636-0, 0-8186-5637-9 (case). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33

D37 1994. URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=305942. IEEE Catalog Number

93TH0626-2. IEEE Computer Society Press Order Number 5637-02.

OConnor:1999:RF

[OR99] John J. O’Connor and Edmund F. Robertson. References for

Fibonacci, 1999. URL http://www-groups.dcs.st-and.ac.
uk/~history/References/Fibonacci.html.

Ortega:1996:OBA

[Ort96] A. Ortega. Optimal bit allocation under multiple rate constraints. In Storer and Cohn [SC96], pages 349–358. ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1996. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=488340. IEEE Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Ornstein:1990:UAS

[OS90] Donald S. Ornstein and Paul C. Shields. Universal almost sure data compression. Annals of Probability, 18(2):441–452, April 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X

(electronic). URL http://projecteuclid.org/euclid.aop/
1176990840.

Ohnesorge:1996:PMM

[OS96] K. W. Ohnesorge and R. Sennhauser. Parameterized Markov models for efficient compression of grayscale images. In Storer and Cohn [SC96], page ?? ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1996. URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

488383. IEEE Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Ogunbona:1997:CBR

[OS97] P. O. Ogunbona and P. Sangassapaviriya. Content-based retrieval from compressed-image databases. In Storer and Cohn [SC97], page ?? ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1997. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=582126. IEEE

Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Oggier:2001:SPD

[OS01] F. E. Oggier and S. D. Servetto. Semidefinite programs for the design of codes for delay-constrained communication in networks. In Storer and Cohn [SC01c], pages 183–192. ISBN 0-7695-1031-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2001. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=917149. IEEE

CSP number 01PR1031.

Orlitsky:2003:PUC

[OS03] A. Orlitsky and N. P. Santhanam. Performance of universal codes over infinite alphabets. In Storer and Cohn [SC03], pages 402–410. ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2003. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1194031. IEEE Computer Society Order number PR01896.

Ozsoy:2014:OLC

[OSC14] Adnan Ozsoy, Martin Swany, and Arun Chauhan. Optimizing LZSS compression on GPGPUs. Future Generation Computer Systems, 30(??):170–178, January 2014. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167739X13001362.

Okazaki:2002:DCM

[OSI02] Takumi Okazaki, Kunihiko Sadakane, and Hiroshi Imai. Data compression method combining properties of PPM and CTW. Lecture Notes in Computer Science, 2281:268–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer-ny.com/link/
service/series/0558/bibs/2281/22810268.htm;   http:
//link.springer-ny.com/link/service/series/0558/papers/
2281/22810268.pdf.

Okada:1995:SOD

[OSMY95] Y. Okada, N. Satoh, K. Murashita, and S. Yoshida. Selforganized dynamic Huffman coding without frequency counts. In Storer and Cohn [SC95], page 473. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=515583. IEEE catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Osterberg:1935:TLR

[Ost35] G. Osterberg. Topography of the layer of rods and cones in the human retina. Acta Ophthalmologica, (suppl. 6)(??):1–103, ???? 1935.

Okada:2003:NDC

[OT03] Y. Okada and K. Terazono. A new delta compression algorithm suitable for program updating in embedded systems. In Storer and Cohn [SC03], page ?? ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 2003. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1194059. IEEE Computer

Society Order number PR01896.

Ordentlich:1999:MES

[OTW^{+}99] E. Ordentlich, D. Taubman, M. Weinberger, G. Seroussi, and M. W. Marcellin. Memory-efficient scalable line-based image coding. In Storer and Cohn [SC99], pages 218–227. ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1999. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=755671. IEEE Computer Society Or-

der Number PR00096. IEEE Order Plan Catalog Number PR00096.

Okada:1997:EMC

[OTY^{+}97] Y. Okada, M. Tokuyo, S. Yoshida, N. Okayasu, and H. Shimoi. Effective management of compressed data with packed file systems. In Storer and Cohn [SC97], page ?? ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1997. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=582127. IEEE Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Ochoa:2009:SPH

[OVCS^{+}09] H. Ochoa, O. Vergara, V. Cruz-Sanchez, G. Rosiles, and J. Vega-Pineda. Set Partitioning in Hierarchical Frequency Bands (SPHFB). In Storer and Marcellin [SM09], page 463. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=4976517. IEEE Computer

Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

OConnell:2003:PJD

[OW03a] S. J. O’Connell and N. Winterbottom. Performing joins without decompression in a compressed database system. ACM SIGMOD Record, 32(1):6–11, March 2003. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Orpaz:2003:PMM

[OW03b] A. Orpaz and S. Weiss. Pattern matching by means of multi-resolution compression. In Storer and Cohn [SC03], page ?? ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2003. URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

1194060. IEEE Computer Society Order number PR01896.

Ordentlich:1998:LCM

[OWS98] E. Ordentlich, M. Weinberger, and G. Seroussi. A lowcomplexity modeling approach for embedded coding of wavelet coefficients. In Storer and Cohn [SC98], pages 408–417. ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1998. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=672180. IEEE Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Ordentlich:2004:DUF

[OWW^{+}04] E. Ordentlich, T. Weissman, M. J. Weinberger, A. Somekh-Baruch, and N. Merhav. Discrete universal filtering through incremental parsing. In Storer and Cohn [SC04], pages 352– 361. ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=1281480. IEEE Com-

puter Society Order Number: P2082.

Ong:2006:BPA

[OY06] Ghim Hwee Ong and Kai Yang. A binary partitioning approach to image compression using weighted finite automata for large images. Computers and Mathematics with Applications, 51(11):1705–1714, June 2006. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0898122106001374.

Ostergaard:2007:MDC

[ØZ07] J. Østergaard and R. Zamir. Multiple-description coding by dithered delta-sigma quantization. In Storer and Cohn [SC07], pages 63–72. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

4148745. IEEE Computer Society Order Number P2791.

Pan:2013:RAI

[P^{+}13] Jeng-Shyang Pan et al., editors. Recent advances in information hiding and applications, volume 40 of Intelligent systems reference library. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 3-642-28579-1, 3-642-28580-5 (e-book). LCCN ????

Puga:1997:CGG

[PA97] A. T. Puga and A. P. Alves. Compression of generalised Gaussian sources. In Storer and Cohn [SC97], page ?? ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1997. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=582131. IEEE Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Payan:2007:TWB

[PA07] Frédéric Payan and Marc Antonini. Temporal waveletbased compression for 3D animated models. Computers and Graphics, 31(1):77–88, January 2007. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0097849306001774.

Pereira:2003:MDC

[PAB03] M. Pereira, M. Antonini, and M. Barlaud. Multiple description coding for noisy-varying channels. In Storer and Cohn [SC03], page ?? ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2003. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1194062. IEEE Computer Society Order number PR01896.

Paeth:1991:IFC

[Pae91] Alan W. Paeth. Image file compression made easy. In James Arvo, editor, Graphics Gems II, page ?? Academic Press, New York, USA, 1991.

Pan:1993:DAC

[Pan93] Davis Yen Pan. Digital audio compression. Digital Technical Journal, 5(2):28–33 (or 28–40??), Spring 1993.

CODEN DTJOEL.  ISSN 0898-901X.  URL ftp://

ftp.digital.com/pub/Digital/info/DTJ/v5n2/Digital_
Audio_Compression_01oct1993DTJA03P8.ps;   http://www.
digital.com:80/info/DTJA03/DTJA03SC.TXT.

Pan:1995:TMA

[Pan95] Davis Yen Pan. A tutorial on MPEG/audio compression. IEEE MultiMedia, 2(2):60–74, Summer 1995. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http://

dlib.computer.org/mu/books/mu1995/pdf/u20060.pdf;
http://www.computer.org/multimedia/mu1995/u2060abs.
htm.

Papaefstathiou:1999:CAS

[Pap99] I. Papaefstathiou. Compressing ATM streams on-line. In Storer and Cohn [SC99], page ?? ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1999. URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

785700. IEEE Computer Society Order Number PR00096. IEEE Order Plan Catalog Number PR00096.

Parsons:1987:VSP

[Par87] Thomas W. Parsons. Voice and Speech Processing. McGraw-Hill, New York, NY, USA, 1987. ?? pp.

Pasco:1976:SCA

[Pas76] Richard Clark Pasco. Source Coding Algorithms for Fast Data Compression. Ph.D. dissertation, Department of Electrical Engineering, Stanford University, Stanford, CA, USA,

1976. 113 pp. URL http://search.proquest.com/docview/
302770683.

Patankar:1994:SHA

[Pat94] Rashmi Arun Patankar. Software and hardware approaches to data compression of IEEE 64-bit floating point data. Thesis (M.S.), Iowa State University, Ames, IA, USA, 1994. 55 pp.

Pavlev:2006:LSS

[Pav06] Igor Pavlev. LZMA SDK (software development kit) home

page, 2006. URL http://www.7-zip.org/sdk.html.

Promhouse:1991:SDC

[PB91] G. Promhouse and M. Bennett. Semantic data compression. In Storer and Reif [SR91b], pages 323–331. ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1990. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=213348. IEEE Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Pigeon:1998:MEA

[PB98a] S. Pigeon and Y. Bengio. A memory-efficient adaptive Huffman coding algorithm for very large sets of symbols. In Storer and Cohn [SC98], page ?? ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1998. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=672310. IEEE

Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Pigeon:1998:AAM

[PB98b] Steven Pigeon and Yoshua Bengio. Algorithm alley: Memoryefficient adaptive Huffman coding. Dr. Dobb’s Journal of Software Tools, 23(10):131–132, 134–135, October 1998. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/ddj/1998/1998_10/Shas/Shas.htm;   http:
//www.ddj.com/ftp/1998/1998_10/aa108.txt;   http://
www.ddj.com/ftp/1998/1998_10/aa108.zip.

Pigeon:1999:BPA

[PB99] S. Pigeon and Y. Bengio. Binary pseudowavelets and applications to bilevel image processing. In Storer and Cohn [SC99], pages 364–373. ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1999. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=755686. IEEE

Computer Society Order Number PR00096. IEEE Order Plan Catalog Number PR00096.

Pereira:2011:CBA

[PB11] R. Pereira and K. Breitman. A cloud based architecture for improving video compression time efficiency: The split & merge approach. In Storer and Marcellin [SM11b], page 471. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5749528. IEEE Computer Society

Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Pan:2005:FTB

[PBC05] Xiang Pan, A. H. Banihashemi, and A. Cuhadar. A fast trellis-based rate-allocation algorithm for robust transmission of progressively coded images over noisy channels. In Storer and Cohn [SC05], page ?? ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=1402230.

Pawlikowski:1995:CDT

[PBEA95] K. Pawlikowski, T. Bell, H. Emberson, and P. Ashton. Compression of data traffic in packet-based LANs. In Storer and Cohn [SC95], page 430. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=515540. IEEE catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Pavlovic:2019:DCP

[PBGA19] Mirjana Pavlovic, Kai-Niklas Bastian, Hinnerk Gildhoff, and Anastasia Ailamaki. Dictionary compression in point cloud data management. ACM Transactions on Spatial Algorithms and Systems (TSAS), 5(1):3:1–3:??, June 2019. CODEN ????

ISSN 2374-0353. URL https://dl.acm.org/citation.cfm?
id=3299770.

Park:2017:HHC

[PBL^{+}17] Jaehyun Park, Seungcheol Baek, Hyung Gyu Lee, Chrysostomos Nicopoulos, Vinson Young, Junghee Lee, and Jongman Kim. HoPE: Hot-cacheline prediction for dynamic early decompression in compressed LLCs. ACM Transactions on Design Automation of Electronic Systems (TODAES), 22(3): 40:1–40:??, May 2017. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Pekhimenko:2015:TAC

[PBO^{+}15] Gennady Pekhimenko, Evgeny Bolotin, Mike O’Connor, Onur Mutlu, Todd C. Mowry, and Stephen W. Keckler. Toggleaware compression for GPUs. IEEE Computer Architecture Letters, 14(2):164–168, July/December 2015. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Prasanna:1995:FIP

[PBPT95] Viktor K. Prasanna, V. P. Bhatkar, L. M. Patnaik, and S. K. Tripathi, editors. First IWPP parallel processing: proceedings of the First International Workshop on Parallel Processing (IWPP-94): December 26–31, 1994, Bangalore, India. Taka McGraw-Hill Pub. Co, New Delhi; New York, 1995. ISBN 0-07-462332-X. LCCN QA 76.58 I587 1994.

Pigeon:2008:VLC

[PC08] S. Pigeon and S. Coulombe. Very low cost algorithms for predicting the file size of JPEG images subject to changes of quality factor and scaling. In Storer and Marcellin [SM08], page 538. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33

INTERNET. URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4483365.

Pizzolante:2017:AVQ

[PCD17] Raffaele Pizzolante, Bruno Carpentieri, and Sergio De Agostino. Adaptive vector quantization for lossy compression of image sequences. Algorithms (Basel), 10(2), June 2017. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https:

//www.mdpi.com/1999-4893/10/2/51.

Perez-Cordoba:2000:HDC

[PCRLSB00] J. L. Perez-Cordoba, A. J. Rubio, J. M. Lopez-Soler, and M. C. Benitez. Hard-decision in COVQ over waveform channels. In Storer and Cohn [SC00], pages 53–62. ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2000. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=838145. This millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Peterson:2007:CNS

[PD07] Larry L. Peterson and Bruce S. Davie. Computer networks: a systems approach. Morgan Kaufmann, Amsterdam, The Netherlands, fourth edition, 2007. ISBN 0-12-370548-7 (hardcover), 0-12-374013-4 (paperback). xxvii + 806 pp. LCCN

TK5105.5 .P479 2007. URL http://www.loc.gov/catdir/
enhancements/fy0704/2006102454-d.html.

Ponchio:2016:MFD

[PD16] Federico Ponchio and Matteo Dellepiane. Multiresolution and fast decompression for optimal web-based rendering. Graphical Models, 88(??):1–11, November 2016. CODEN GR-MOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S1524070316300285.

Pereira:2002:MB

[PE02] Fernando Pereira and Touradj Ebrahimi. The MPEG-4 Book. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2002. ?? pp.

Peano:1890:CQR

[Pea90] G. Peano. Sur une courbe qui remplit toute une aire plane. (French) [On a curve that completely fills a planar area]. Mathematische Annalen, 36:157–160, 1890. CODEN MAANA3. ISSN 0025-5831 (print), 1432-1807 (electronic).

Pechura:1982:FAT

[Pec82] Michael A. Pechura. File archival techniques using data compression. Communications of the ACM, 25(9):605–609, 1982.

CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ponomarenko:2002:CIB

[PELA02] N. N. Ponomarenko, K. Egiazarian, V. V. Lukin, and J. T. Astola. Compression of image block means for nonequal block size partition schemes using Delaunay triangulation and prediction. In Storer and Cohn [SC02], page ?? ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2002. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=1000011. IEEE

Computer Society Order Number PR01477.

Perkins:1988:DCS

[Per88] Michael G. Perkins. Data compression of stereopairs. Ph.D. thesis, Stanford University, Stanford, CA, USA, 1988. 187 pp.

URL http://search.proquest.com/docview/303695164.

Perkins:1989:RPP

[Per89] D. Perkins. RFC 1134: Point-to-point protocol: a proposal for multi-protocol transmission of datagrams over point-to-point links, November 1, 1989. URL ftp://ftp.internic.

net/rfc/rfc1134.txt;   ftp://ftp.internic.net/rfc/
rfc1171.txt;   https://www.math.utah.edu/pub/rfc/rfc1134.
txt;   https://www.math.utah.edu/pub/rfc/rfc1171.txt.

Obsoleted by RFC1171 [Per90]. Status: PROPOSED STAN-DARD.

Perkins:1990:RPPa

[Per90] D. Perkins. RFC 1171: Point-to-point protocol for the transmission of multi-protocol datagrams over point-to-point links, July 1, 1990. URL ftp://ftp.internic.

net/rfc/rfc1134.txt;   ftp://ftp.internic.net/rfc/
rfc1171.txt;   ftp://ftp.internic.net/rfc/rfc1331.
txt;   https://www.math.utah.edu/pub/rfc/rfc1134.txt;
 https://www.math.utah.edu/pub/rfc/rfc1171.txt;   https:
//www.math.utah.edu/pub/rfc/rfc1331.txt. Obsoleted

by RFC1331 [Sim92]. Obsoletes RFC1134 [Per89]. Status: DRAFT STANDARD.

Percival:2003:ABS

[Per03a] Colin Percival. An automated binary security update system for FreeBSD. In Proceedings of BSDCon ’03, pages 29–34. USENIX, Berkeley, CA, USA, 2003.

Percival:2003:NDE

[Per03b] Colin Percival. Naïve differences of executable code. Technical report, Computing Lab, Oxford University, Oxford, UK, 2003.

3 pp. URL http://www.daemonology.net/papers/bsdiff.
pdf.

Percival:2005:BDP

[Per05] Colin Percival. Binary diff/patch utility, 2005. URL http://

www.daemonology.net/bsdiff/bsdiff-4.3.tar.gz.

Percival:2006:MMA

[Per06] Colin Percival. Matching with Mismatches and Assorted Applications. Ph.D. thesis, ????, ????, 2006. ?? pp. URL http:

//www.daemonology.net/papers/thesis.pdf.

Perry:2021:CCW

[Per21] Tekla S. Perry. Conjurer of compression: From WinZips to cat GIFs, Jacob Ziv’s algorithms have been making data disappear and reappear for decades. IEEE Spectrum, 58(5):48–53, May 2021. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Peshkin:2007:SIL

[Pes07] L. Peshkin. Structure induction by lossless graph compression. In Storer and Cohn [SC07], pages 53–62. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4148744. IEEE Computer Society

Order Number P2791.

Peterson:1980:CPD

[Pet80] James Lyle Peterson. Computer programs for detecting and correcting spelling errors. Communications of the ACM, 23 (12):676–687, December 1980. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remarks [Dun81, Mil81, Nix81, Pet81].

Peterson:1981:LES

[Pet81] James L. Peterson. Letter to the Editor: On spelling correction and beyond. Communications of the ACM, 24(9):609, September 1981. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Pet80, Dun81, Mil81, Nix81].

Press:1988:NRC

[PF^{+}88] W. H. Press, B. P. Flannery, et al. Numerical Recipes in C: The Art of Scientific Computing. Cambridge University Press, Cambridge, UK, 1988. ISBN 0-521-35465-X, 0-521-35466-8 (diskette). xxii + 735 pp. LCCN QA76.73.C15 N865 1988.

URL http://www.nr.com/.

Patchoo:2010:AAQ

[PFAK10] W. Patchoo, T. R. Fischer, Changho Ahn, and Sangwon Kang. Analysis of amplitude quantization in ACELP excitation coding. In Storer and Marcellin [SM10b], page 550. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5453527.

Pinho:1999:FMM

[PFP99] M. S. Pinho, W. A. Finamore, and W. A. Pearlman. Fast multi-match Lempel–Ziv. In Storer and Cohn [SC99], page ?? ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1999. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=785702. IEEE Computer Society Or-

der Number PR00096. IEEE Order Plan Catalog Number PR00096.

Paez:1972:MMS

[PG72] M. D. Paez and T. H. Glisson. Minimum mean squared error quantization in speech PCM and DPCM systems. IEEE Transactions on Communications, COM-20(2):225–230, 1972. CODEN IECMBT. ISSN 0090-6778 (print), 1558-0857 (electronic).

Perlmutter:1994:BRW

[PGOO94] K. O. Perlmutter, R. M. Gray, K. L. Oehler, and R. A. Olshen. Bayes risk weighted tree-structured vector quantization with posterior estimation. In Storer and Cohn [SC94], pages 274–283. ISBN 0-8186-5636-0, 0-8186-5637-9 (case). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1994. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=305935. IEEE

Catalog Number 93TH0626-2. IEEE Computer Society Press Order Number 5637-02.

Parodi:2010:PLA

[PGS10] Mauro Parodi, Mauro Gaggero, and Marco Storace. Piecewise linear approximations of multivariate functions: a multiresolution-based compression algorithm suitable for circuit implementation. Applied Numerical Mathematics: Transactions of IMACS, 60(9):924–933, September 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Peng:2017:PMT

[PGW^{+}17] Min Peng, Wang Gao, Hua Wang, Yanchun Zhang, Jiajia Huang, Qianqian Xie, Gang Hu, and Gang Tian. Parallelization of massive textstream compression based on compressed sensing. ACM Transactions on Information Systems, 36(2): 17:1–17:??, September 2017. CODEN ATISET. ISSN 1046-8188.

Perkins:1990:RPPb

[PH90] D. Perkins and R. Hobby. RFC 1172: Point-to-point protocol (PPP) initial configuration options, July 1, 1990.

URL ftp://ftp.internic.net/rfc/rfc1172.txt;   ftp:
//ftp.internic.net/rfc/rfc1331.txt;   https://www.
math.utah.edu/pub/rfc/rfc1172.txt;   https://www.math.
utah.edu/pub/rfc/rfc1331.txt. Obsoleted by RFC1331

[Sim92]. Status: PROPOSED STANDARD.

Pentland:1991:PAF

[PH91] A. Pentland and B. Horowitz. A practical approach to fractal-based image compression. In Storer and Reif [SR91b], pages 176–185. ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1990. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=213385. IEEE

Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Prochazka:2011:BOD

[PH11] Petr Procházka and Jan Holub. Block-oriented dense compressor. In Storer and Marcellin [SM11b], page 472. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5749529. IEEE Computer Society

Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Posner:1986:CES

[PHG86] E. C. Posner, R. L. Horttor, and T. L. Grant. Communicating from the edge of the solar system: Centimeter-band transmissions, three-dimensional data compression, and laser-based systems may launch unmanned missions to deep space. IEEE Spectrum, 23(6):60–65, June 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Phillips:1991:IPPa

[Phi91] Dwayne Phillips. Image processing — part 1: Reading the Tag Image File Format. C Users Journal, 9(3):92–??, March 1991. ISSN 0898-9788.

Phillips:1992:DCU

[Phi92a] Dwayne Phillips. Data compression using Huffman coding. C Users Journal, 10(2):55–??, February 1992. ISSN 0898-9788.

Phillips:1992:LDC

[Phi92b] Dwayne Phillips. LZW data compression. Circuit Cellar INK — The Computer Applications Journal, 27:36–48, June/July 1992. CODEN CCIEBN. ISSN 0896-8985.

Pajak:2014:IIP

[PHM^{+}14] Dawid Pajak, Robert Herzog, Radosław Mantiuk, Piotr Didyk, Elmar Eisemann, Karol Myszkowski, and Kari Pulli. Images II: Perceptual depth compression for stereo applications. Computer Graphics Forum, 33(2):195–204, May 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pierzchala:2003:CWO

[Pie03] Stephen Pierzchala. Compressing Web output using mod gzip and Apache. ;login: the USENIX Association newsletter, 28(2):??, April 2003. CODEN LOGNEM. ISSN 1044-

6397. URL http://www.usenix.org/publications/login/
2003-04/pdfs/pierzchala.pdf.

Pigeon:1999:AAI

[Pig99] Steven Pigeon. Algorithm alley: Image compression with wavelets. Dr. Dobb’s Journal of Software Tools, 24(8): 111–115, August 1999. CODEN DDJOEB. ISSN 1044-

789X. URL 9908n/9908n.htm;   http://www.ddj.com/
ftp/1999/1999_08/aa899.txt;   http://www.ddj.com/ftp/
1999/1999_08/aa899.zip.

Pigeon:2001:CCD

[Pig01a] Steven Pigeon. Contributions à la compression de données. (French) [Contributions to Data Compression]. Ph.D. thesis, Département d’informatique et de recherche opérationnelle, Faculté des Arts & Sciences, Université de Montréal, Montréal, QC, Canada, December 2001. xvi + 296 pp. URL http://

www.stevenpigeon.org/Publications/publications/phd.

pdf. The part on taboo codes is “Taboo Codes, New Classes of Universal Codes”, and is available as a separate report [Pig02].

Pigeon:2001:SSC

[Pig01b] Steven Pigeon. Start/stop codes. In Proceedings of the Data Compression Conference (DCC ’01), page 511. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2001. URL http://

www.stevenpigeon.org/Publications/publications/ssc_
full.pdf.

Pigeon:2002:TCN

[Pig02] Steven Pigeon. Taboo codes: New classes of universal codes. Report, Université de Montréal, Montréal, QC, Canada, January 17, 2002. http://www.iro.umontreal.ca/ brassard/SEMINAIRES/taboo.ps.

Piner:1999:CSN

[Pin99] Mary-Louise G. Piner. Computer Society notes: David Huffman (1925–1999). Computer, 32(12):86, December 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814

(electronic). URL http://dlib.computer.org/co/books/
co1999/pdf/rz086.pdf.

Puschel:2005:RTF

[PK05] M. Puschel and J. Kovacevic. Real, tight frames with maximal robustness to erasures. In Storer and Cohn [SC05], pages 63– 72. ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=1402167.

Park:2010:DCO

[PK10] Sang-Young Park and Seong-Dae Kim. Depth compression of 3D object represented by layered depth image. In Storer and Marcellin [SM10b], pages 504–513. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=5453470.

Pattem:2008:ISC

[PKG08] Sundeep Pattem, Bhaskar Krishnamachari, and Ramesh Govindan. The impact of spatial correlation on routing with compression in wireless sensor networks. ACM Transactions on Sensor Networks, 4(4):24:1–24:??, August 2008. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

PKWare:2003:PHP

[PKW03] PKWare. PKWare home page, 2003. URL http://www.

pkware.com/.

Prusinkiewicz:1990:ABP

[PL90] P. Prusinkiewicz and A. Lindenmayer. The Algorithmic Beauty of Plants. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. ???? pp.

Prusinkiewicz:1991:SSF

[PLF91] P. Prusinkiewicz, A. Lindenmayer, and F. D. Fracchia. Synthesis of space-filling curves on the square grid. In H.-O. Peitgen et al., editors, Fractals in the Fundamental and Applied Sciences, pages 341–366. Elsevier Science, Inc., Amsterdam, The Netherlands, 1991.

Pennebaker:1988:PEC

[PM88] William B. Pennebaker and Joan L. Mitchell. Probability estimation for the Q-coder. IBM Journal of Research and Development, 32(6):737–752, November 1988. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Perl:1991:CLA

[PM91] Yehoshua Perl and Ashish Mehta. Cascading LZW algorithm with Huffman coding: a variable to variable length compression algorithm. Lecture Notes in Computer Science, 507:170– ??, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pennebaker:1992:JSI

[PM92] William B. Pennebaker and Joan L. Mitchell. JPEG Still Image Data Compression Standard. Van Nostrand Reinhold, New York, NY, USA, 1992. ISBN 0-442-01272-1. xviii + 638 pp. LCCN TA1632 .P45 1993.

Prashant:1995:MPA

[PM95] K. S. Prashant and V. J. Mathews. A massively parallel algorithm for vector quantization. In Storer and Cohn [SC95], page ?? ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=515604. IEEE

catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Palau:1998:FSC

[PM98] A. Palau and G. Mirchandani. A fast symbol coding scheme with specific application in bulk compression. In Storer and Cohn [SC98], page ?? ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1998. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=672309. IEEE

Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Park:1999:IJS

[PM99] M. S. Park and D. J. Miller. Improved joint source-channel decoding for variable-length encoded data using soft decisions and MMSE estimation. In Storer and Cohn [SC99], page ?? ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1999. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=785701. IEEE Computer Society Or-

der Number PR00096. IEEE Order Plan Catalog Number PR00096.

Pennec:2005:BIA

[PM05] E. Le Pennec and S. Mallat. Bandelet image approximation and compression. Multiscale Modeling & Simulation, 4(3):992– 1039, 2005. CODEN MMSUBT. ISSN 1540-3459 (print), 1540-

3467 (electronic). URL http://epubs.siam.org/sam-bin/
dbq/article/61945.

Palangappa:2017:CCE

[PM17] Poovaiah M. Palangappa and Kartik Mohanram. CompEx++: Compression-expansion coding for energy, latency, and lifetime improvements in MLC/TLC NVMs. ACM Transactions on Architecture and Code Optimization, 14(1):10:1–10:??, April 2017. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Petri:2018:CII

[PM18] Matthias Petri and Alistair Moffat. Compact inverted index storage using general-purpose compression libraries. Software — Practice and Experience, 48(4):974–982, April 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/abs/
10.1002/spe.2556.

Prashant:1995:NMP

[PMH95] K. S. Prashant, V. J. Mathews, and P. J. Hahn. A new model of perceptual threshold functions for application in image compression systems. In Storer and Cohn [SC95], pages 371– 380. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=515527. IEEE catalog

number 95TH8037. IEEE Computer Society Press order number PR07010.

Perl:1991:CLC

[PMK91] Y. Perl, V. Maram, and N. Kadakuntla. The cascading of the LZW compression algorithm with arithmetic coding. In Storer and Reif [SR91b], pages 277–286. ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1990. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=213353. IEEE Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Pennebaker:1988:OBP

[PMLA88] William B. Pennebaker, Joan L. Mitchell, Glen G. Langdon, Jr., and Ronald B. Arps. An overview of the basic principles of the Q-coder adaptive binary arithmetic coder. IBM Journal of Research and Development, 32(6):717–726 (or 206–211??), November 1988. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Pande:2013:SMC

[PMZ13] Amit Pande, Prasant Mohapatra, and Joseph Zambreno. Securing multimedia content using joint compression and encryption. IEEE MultiMedia, 20(4):50–61, October/December 2013. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Polit:1991:NTD

[PN91] P. P. Polit and N. M. Nasrabadi. A new transform domain vector quantization technique for image data compression in an asynchronous transfer mode network. In Storer and Reif [SR91b], pages 149–158. ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1990. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=213382. IEEE Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Pan:2000:CST

[PO00] W. Pan and A. Ortega. Complexity-scalable transform coding using variable complexity algorithms. In Storer and Cohn [SC00], pages 263–272. ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2000. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=838166. This

millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Poehler:1983:DCI

[Poe83] Paul Leonard Poehler. Data Compression of Imagery Using Linear Predictive Coding Techniques. Ph.D. thesis, Florida Institute of Technology, Melbourne, FL, USA, 1983. 268 pp.

URL http://search.proquest.com/docview/303271798.

Pohlmann:1985:PDA

[Poh85] Ken Pohlmann. Principles of Digital Audio. SAMS Publishing, Indianapolis, IN, USA, 1985. ?? pp.

Pomeranz:2016:PSS

[Pom16] Irith Pomeranz. Periodic scan-in states to reduce the input test data volume for partially functional broadside tests. ACM Transactions on Design Automation of Electronic Systems (TODAES), 22(1):7:1–7:??, December 2016. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Popa:2005:KTD

[Pop05] Ionuţ Popa. KDS-transformation for data compression. Fundamenta Informaticae, 67(4):371–375, March 2005. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Popa:2011:SCG

[Pop11] Ionut. Popa. SE-compression: a generalization of dictionarybased compression. The Computer Journal, 54(11):1876–1881, November 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Pore:1973:FAD

[Por73] Michael David Pore. Factor Analysis as a Data Compression Technique. Ph.D. thesis, Texas Tech University, Lubbock, TX,

USA, 1973. 381 pp. URL http://search.proquest.com/
docview/302771096.

Pournelle:1993:QOC

[Pou93] Jerry Pournelle. The DOS 6 question: Our columnist finds DOS 6 is the least expensive route to disk compression and memory optimization. Byte Magazine, 18(8):209–??, July 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Powell:1993:DCE

[Pow93] E. T. Powell. Data compression enhancements to the distributed interactive simulation protocol for use by entitylevel simulations. Report UCRL-ID-114812, Lawrence Livermore National Laboratory, United States. Department of Energy, Livermore, CA, USA, 1993. 18 pp. URL http:

//www.columbia.edu/cgi-bin/cul/resolve?clio9803512;
 http://www.osti.gov/servlets/purl/10188557-PEbS9W/

native/. DE94000957. W-7405-ENG-4.

Poynton:2008:CSN

[Poy08] Charles Poynton. Chroma subsampling notation, 2008. URL

http://www.poynton.com/PDFs/Chroma_subsampling_notation.
pdf.

Poss:2003:DCO

[PP03] Meikel Pöss and Dmitry Potapov. Data compression in Oracle. In Freytag et al. [FLA^{+}03], pages 937–947. ISBN 0-12-

722442-4. LCCN ???? URL http://www.vldb.org/dblp/db/
indices/a-tree/p/P=ouml=ss:Meikel.html.

Papadopoulos:2010:TRM

[PP10] Konstantinos Papadopoulos and Ioannis Papaefstathiou. Titan-R: a multigigabit reconfigurable combined compression/ decompression unit. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 3(2):7:1–7:??, May 2010. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Policriti:2018:LCB

[PP18] Alberto Policriti and Nicola Prezza. LZ77 computation based on the run-length encoded BWT. Algorithmica, 80(7):1986– 2011, July 2018. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

Puri:2002:CMD

[PPR02] R. Puri, S. S. Pradhan, and K. Ramchandran. n-channel multiple descriptions: theory and constructions. In Storer and Cohn [SC02], pages 262–271. ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2002. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=999964. IEEE Computer Society Order Number PR01477.

Posner:1971:EED

[PR71] Edward C. Posner and Eugene R. Rodemich. Epsilon entropy and data compression. Annals of Mathematical Statistics, 42 (6):2079–2125, December 1971. CODEN AASTAD. ISSN

0003-4851. URL http://projecteuclid.org/euclid.aoms/
1177693077.

Peitgen:1986:BF

[PR86] Heinz-Otto Peitgen and Peter H. Richter. The Beauty of Fractals. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. ISBN 0-387-15851-0. xii + 199 pp. LCCN QA447 .P45 1986.

Plandowski:1998:ALZ

[PR98] Wojciech Plandowski and Wojciech Rytter. Application of Lempel–Ziv encodings to the solution of word equations. Lecture Notes in Computer Science, 1443:731–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer-ny.com/link/
service/series/0558/bibs/1443/14430731.htm;   http:
//link.springer-ny.com/link/service/series/0558/papers/
1443/14430731.pdf.

Pradhan:1999:DSC

[PR99] S. S. Pradhan and K. Ramchandran. Distributed source coding using syndromes (DISCUS): design and construction. In Storer and Cohn [SC99], pages 158–167. ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1999. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=755665. IEEE Computer Society Or-

der Number PR00096. IEEE Order Plan Catalog Number PR00096.

Pradhan:2000:DSC

[PR00] S. S. Pradhan and K. Ramchandran. Distributed source coding: symmetric rates and applications to sensor networks. In Storer and Cohn [SC00], pages 363–372. ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2000. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=838176. This millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Pradhan:2001:EAI

[PR01] S. S. Pradhan and K. Ramchandran. Enhancing analog image transmission systems using digital side information: a new wavelet-based image coding paradigm. In Storer and Cohn [SC01c], pages 63–72. ISBN 0-7695-1031-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2001. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=917137. IEEE CSP number 01PR1031.

Pugh:2011:DQO

[PR11] M. Pugh and B. D. Rao. Distributed quantization of order statistics with applications to CSI feedback. In Storer and Marcellin [SM11b], pages 323–332. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=5749490. IEEE Computer Society Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Pavlovic:1997:UTI

[PRM97] V. Pavlovic, K. Ramchandran, and P. Moulin. ‘universal’ transform image coding based on joint adaptation of filter banks, tree structures and quantizers. In Storer and Cohn [SC97], page ?? ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1997. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=582130. IEEE

Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Prountzos:1986:DIM

[Pro86] Demetrios A. Prountzos. The design and implementation of a microprocessor based data compression system. M.E.E. thesis, The Cooper Union for the Advancement of Science and Art, New York, NY, USA, 1986. 134 pp. URL http://search.

proquest.com/docview/303488933.

Promhouse:1993:TCT

[Pro93] G. Promhouse. Tree compacting transformations. In Storer and Cohn [SC93], pages 320–329. ISBN 0-8186-3391-3 (microfiche), 0-8186-3392-1 (casebound). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1993. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=253117. IEEE Computer Society Press order number 3392-02. IEEE catalog number 93TH0536-3.

Przelaskowski:1998:DPW

[Prz98] Artur Przelaskowski. Detail preserving wavelet-based compression with adaptive context-based quantisation. Fundamenta Informaticae, 34(4):369–388, July 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Przelaskowski:2000:MUQ

[Prz00] A. Przelaskowski. Modifications of uniform quantization applied in wavelet coder. In Storer and Cohn [SC00], pages 293–302. ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 2000. URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=838169. This millen-

nial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Printz:1993:MAC

[PS93] H. Printz and P. Stubley. Multialphabet arithmetic coding at 16 MBytes/sec. In Storer and Cohn [SC93], pages 128– 137. ISBN 0-8186-3391-3 (microfiche), 0-8186-3392-1 (casebound). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 1993. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=253137. IEEE Computer

Society Press order number 3392-02. IEEE catalog number 93TH0536-3.

Perkins:2005:RDC

[PS05] S. Perkins and D. H. Smith. Robust data compression: Variable length codes and burst errors. The Computer Journal, 48(3):315–322, May 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/cgi/content/abstract/48/3/315;
http://comjnl.oxfordjournals.org/cgi/reprint/48/3/
315.

Prasad:2006:PID

[PS06] M. V. N. K. Prasad and K. K. Shukla. Parallel implementation of domain decomposition based image compression algorithm. International Journal of Computer Applications, 28 (4):334–341, 2006. ISSN 1206-212X (print), 1925-7074 (elec-

tronic). URL https://www.tandfonline.com/doi/full/10.
1080/1206212X.2006.11441819.

Porat:2012:CHV

[PS12] E. Porat and B. Shalem. A cuckoo hashing variant with improved memory utilization and insertion time. In Storer and Marcellin [SM12], pages 347–356. ISBN 1-4673-0715-7 (paperback). ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=6189266. IEEE Computer Society

order number P4656.

Pettijohn:2000:JSC

[PSH00] B. D. Pettijohn, K. Sayood, and M. W. Hoffman. Joint source/channel coding using arithmetic codes. In Storer and Cohn [SC00], pages 73–82. ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2000. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=838147. This

millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Prasad:2001:IBI

[PSM01] M. V. N. K. Prasad, K. K. Shukla, and R. N. Mukherjee. Implementation of BTTC image compression algorithm using fuzzy technique. Lecture Notes in Computer Science, 2275:375–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer-

ny.com/link/service/series/0558/bibs/2275/22750375.
htm;   http://link.springer-ny.com/link/service/series/
0558/papers/2275/22750375.pdf.

Perkins:2004:RDC

[PSR04] S. Perkins, D. H. Smith, and A. Ryley. Robust data compression: Consistency checking in the synchronization of variable length codes. The Computer Journal, 47(3):309–319, May 2004. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Prakash:2003:SFS

[PSSS03] J. Prakash, C. Sandeep, P. Shankar, and Y. N. Srikant. A simple and fast scheme for code compression for VLIW processors. In Storer and Cohn [SC03], page ?? ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 2003. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1194063. IEEE Computer

Society Order number PR01896.

Panagopoulou:1997:ESC

[PST97] G. Panagopoulou, S. Sirmakessis, and A. Tsakalidis. Efficient storage compression for 3D regions. In Storer and Cohn [SC97], page ?? ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1997. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=582128. IEEE

Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Puglisi:2005:PLT

[PST05] S. J. Puglisi, W. F. Smyth, and A. Turpin. The performance of linear time suffix sorting algorithms. In Storer and Cohn [SC05], pages 358–367. ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ????

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=
&arnumber=1402197.

Pulina:2010:ESQ

[PT10] Luca Pulina and Armando Tacchella. An empirical study of QBF encodings: from treewidth estimation to useful preprocessing. Fundamenta Informaticae, 102(3–4):391–427, August 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Pu:2006:FDC

[Pu06] Ida Mengyi Pu. Fundamental data compression. Butterworth-Heinemann, Oxford, UK / Boston, MA, USA, 2006. ISBN 0-7506-6310-3 (paperback). xxi + 246 pp. LCCN QA76.9 .D33

P8 2006. URL http://site.ebrary.com/lib/ucalgary/
Doc?id=10190890.

Price:2006:CCT

[PV06] G. Price and M. Vachharajani. A case for compressing traces with BDDs. IEEE Computer Architecture Letters, 5(2):18, February 2006. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Pibiri:2021:TII

[PV21] Giulio Ermanno Pibiri and Rossano Venturini. Techniques for inverted index compression. ACM Computing Surveys, 53(6): 125:1–125:36, February 2021. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic). URL https://dl.acm.

org/doi/10.1145/3415148.

Prenter:1986:TAD

[PW86] P. M. Prenter and E. R. Westwater. Three adaptive discrete least squares cubic spline procedures for the compression of data. Computer Vision, Graphics, and Image Processing, 33 (3):327–345, March 1986. CODEN CVGPDB. ISSN 0734-189X.

Patra:2003:VMI

[PW03a] A. Patra and M. D. Wang. Volumetric medical image compression and reconstruction for interactive visualization in surgical planning. In Storer and Cohn [SC03], page ?? ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 2003. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1194061. IEEE Computer

Society Order number PR01896.

Pigeon:2003:SCR

[PW03b] S. Pigeon and Xiaolin Wu. Searchable compressed representation of very sparse bitmaps. In Storer and Cohn [SC03], pages 353–361. ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2003. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1194026. IEEE Computer Society Order number PR01896.

Pavlichin:2018:QSG

[PWM18] D. Pavlichin, T. Weissman, and G. Mably. The quest to save genomics: Unless researchers solve the looming data compression problem, biomedical science could stagnate. IEEE Spectrum, 55(9):27–31, September 2018. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Pence:2010:OCF

[PWS10] W. D. Pence, R. L. White, and R. Seaman. Optimal compression of floating-point astronomical images without significant loss of information. arXiv.org, ??(??):??, July 7, 2010. CODEN

???? ISSN ???? URL http://arxiv.org/abs/1007.1179.

Published in PASP 122, 1096 (2010).

Pan:2012:SST

[PX12] Zhiming Pan and Hongkai Xiong. Sparse spatio-temporal representation with adaptive regularized dictionaries for superresolution based video coding. In Storer and Marcellin [SM12], pages 139–148. ISBN 1-4673-0715-7 (paperback). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ????

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=6189245. IEEE Computer Society order number P4656.

Pylak:1990:EML

[Pyl90] Paweł Pylak. Efficient modification of LZSS compression algorithm. Annales UMCS Informatica, section AI, pages 61–72,

1990. URL http://www.annales.umcs.lublin.pl/AI/2003/
07.pdf.

Pande:2012:PDP

[PZ12] Amit Pande and Joseph Zambreno. Poly-DWT: Polymorphic wavelet hardware support for dynamic image compression. ACM Transactions on Embedded Computing Systems, 11(1):6:1–6:??, March 2012. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Pan:2022:EDA

[PZZ^{+}22] Zaifeng Pan, Feng Zhang, Yanliang Zhou, Jidong Zhai, Xipeng Shen, Onur Mutlu, and Xiaoyong Du. Exploring data analytics without decompression on embedded GPU systems. IEEE Transactions on Parallel and Distributed Systems, 33(7):1553– 1568, July 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Qiu:2009:PAT

[QB09] Jinshi Qiu and A. H. Banihashemi. Policy allocation for transmission of embedded bit streams over noisy channels with feedback. In Storer and Marcellin [SM09], page 464. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4976518. IEEE Computer Society

Order Number P3592. BMS Part Number CFP09DCC-PRT.

Qian:1999:GJS

[QJRA99] L. Qian, D. L. Jones, K. Ramchandran, and S. Appadwedula. A general joint source-channel matching method for wireless video transmission. In Storer and Cohn [SC99], pages 414–423. ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1999. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=755691. IEEE Computer Society Or-

der Number PR00096. IEEE Order Plan Catalog Number PR00096.

Qu:2012:ISD

[QLH12] Zhenhua Qu, Weiqi Luo, and Jiwu Huang. Identifying shifted double JPEG compression artifacts for non-intrusive digital image forensics. Lecture Notes in Computer Science, 7633:1– 8, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34263-9_1/.

Qi:2011:UHC

[QLQ11] Zhao-Hui Qi, Ling Li, and Xiao-Qin Qi. Using Huffman coding method to visualize and analyze DNA sequences. Journal of Computational Chemistry, 32(15):3233–3240, November 30, 2011. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Qiu:2019:DCR

[QMCX19] Han Qiu, Gerard Memmi, Xuan Chen, and Jian Xiong. DC coefficient recovery for JPEG images in ubiquitous communication systems. Future Generation Computer Systems, 96(??): 23–31, July 2019. CODEN FGSEVI. ISSN 0167-739X (print),

1872-7115 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0167739X1832942X.

Qi:2003:DBM

[QT03] Xiaojun Qi and J. M. Tyler. Differentiation-based multiresolution approach for lossless image compression. In Storer and Cohn [SC03], page ?? ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2003. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1194064. IEEE Computer Society Order number PR01896.

Quan:2012:NWB

[QWH12] Jin Quan, W. G. Wee, and C. Y. Han. A new wavelet based image denoising method. In Storer and Marcellin [SM12], page 408. ISBN 1-4673-0715-7 (paperback). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

6189289. IEEE Computer Society order number P4656.

Qian:2012:LDH

[QZ12] Zhenxing Qian and Xinpeng Zhang. Lossless data hiding in JPEG bitstream. The Journal of Systems and Software, 85(2):309–313, February 2012. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0164121211002123.

Qian:2014:IAF

[QZ14]  Zhenxing Qian and Xinpeng Zhang.  Improved anti-

forensics of JPEG compression. The Journal of Systems and Software, 91(??):100–108, May 2014. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0164121214000168.

Qadir:2010:EJC

[QZH10] G. Qadir, X. Zhao, and A. T. S. Ho. Estimating JPEG2000 compression for image forensics using Benford’s Law. In Society of Photo-Optical Instrumentation Engineers (SPIE) Conference Series, volume 7723 of Society of Photo-Optical Instrumentation Engineers (SPIE) Conference Series. SPIE Optical Engineering Press, Bellingham, WA, USA, April 2010. URL

http://adsabs.harvard.edu/abs/2010SPIE.7723E..14Q.

Ritter:2006:MDI

[R^{+}06] G. X. Ritter et al., editors. Mathematics of data/image pattern recognition, compression, and encryption with applications IX: 15–16 August 2006, San Diego, California, USA, volume 6315 of Proceedings of SPIE. SPIE Optical Engineering Press, Bellingham, WA, USA, 2006. ISBN 0-8194-6394-9. LCCN TA1637 .M389 2006; TS510 .S62 V.6315.

Ritter:2007:MDI

[R^{+}07] G. X. Ritter et al., editors. Mathematics of data/image pattern recognition, compression, coding, and encryption X, with applications: 26–27 August, 2007, San Diego, California, USA, volume 6700 of Proceedings of SPIE. SPIE Optical Engineering Press, Bellingham, WA, USA, 2007. ISBN 0-8194-6848-7. LCCN TS510 .S62 V.6700.

Reznik:2003:ARR

[RA03] Y. A. Reznik and A. V. Anisimov. On the average redundancy rate of adaptive block codes under mixed sources. In Storer and Cohn [SC03], page ?? ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 2003. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1194066. IEEE Computer

Society Order number PR01896.

Ryabko:2004:ACP

[RA04a] B. Ryabko and J. Astola. Adaptive coding and prediction of sources with large and infinite alphabets. In Storer and Cohn [SC04], page ?? ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1281536. IEEE Computer Society Order Number: P2082.

Ryabko:2004:UCF

[RA04b] B. Ryabko and J. Astola. Universal coding of function spaces as a model signal compression. In Storer and Cohn [SC04], pages 382–388. ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ????

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=1281483. IEEE Computer Society Order Number: P2082.

Ryabkot:2004:FCL

[RA04c] B. Ryabkot and J. Astola. Fast codes for large alphabet sources and its application to block encoding. In Storer and Cohn [SC04], page ?? ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1281537. IEEE Computer Society Order Number: P2082.

Richter:2016:JXN

[RAE16] Thomas Richter, Alessandro Artusi, and Touradj Ebrahimi. JPEG XT: A new family of JPEG backward-compatible standards. IEEE MultiMedia, 23(3):80–88, July/September 2016. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (elec-

tronic). URL https://www.computer.org/csdl/mags/mu/
2016/03/mmu2016030080-abs.html.

Reed:1992:PBC

[RAFH92] T. R. Reed, V. R. Algazi, G. E. Ford, and I. Hussain. Perceptually based coding of monochrome and color still images. In Storer and Cohn [SC92], pages 142–151. ISBN 0-8186-2717-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=227467. IEEE catalog number 91TH0436-6.

Raita:1987:ASF

[Rai87] T. Raita. An automatic system for file compression. The Computer Journal, 30(1):80–86, February 1987. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/30/1/80.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_30/Issue_01/tiff/80.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_01/tiff/
81.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_30/Issue_01/tiff/82.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_01/tiff/
83.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_30/Issue_01/tiff/84.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_30/Issue_01/tiff/
85.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_30/Issue_01/tiff/86.tif.

Rajpoot:2004:MBO

[Raj04] N. M. Rajpoot. Model based optimal bit allocation. In Storer and Cohn [SC04], page ?? ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1281534. IEEE Computer Society Order Number: P2082.

Raja:2020:WBI

[Raj20] S. P. Raja. Wavelet-based image compression encoding techniques — a complete performance analysis. International Journal of Image and Graphics (IJIG), 20(02):??, April 2020. ISSN

0219-4678. URL https://www.worldscientific.com/doi/
10.1142/S0219467820500084.

Ramamurthi:1985:VQI

[Ram85] Bhaskar Ramamurthi. Vector Quantization of Images Based on a Composite Source Model (Coding, Data Compression, Picture). Ph.D. thesis, University of California, Santa Barbara, Santa Barbara, USA, 1985. 377 pp. URL http://search.

proquest.com/docview/303336232.

Ramabadran:1987:ASM

[Ram87] Tenkasi V. Ramabadran. Adaptive Source Models for Data Compression. Ph.D. thesis, University of Notre Dame, Notre Dame, IN, USA, 1987. 162 pp. URL http://search.

proquest.com/docview/303617541.

Ramabadran:1991:SRA

[Ram91] T. V. Ramabadran. Some results on adaptive statistics estimation for the reversible compression of sequences. In Storer and Reif [SR91b], page ?? ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1990. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=213312. IEEE

Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Ramlaoui:1992:DCR

[Ram92] Hussam Ibrahim Ramlaoui. Data compression review and implementation. M.S. thesis, California State University, Long Beach, Long Beach, CA, USA, 1992. 72 pp. URL http://

search.proquest.com/docview/304033789.

Ranganathan:1988:HAD

[Ran88] N. Ranganathan. Hardware algorithms for data compression. Ph.D. thesis, University of Central Florida, Orlando, FL, USA,

1988. 181 pp. URL http://search.proquest.com/docview/
303705163.

Rand:1996:RPC

[Ran96] D. Rand. RFC 1962: The PPP compression control proto-

col (CCP), June 1996. URL ftp://ftp.internic.net/rfc/
rfc1962.txt;   ftp://ftp.internic.net/rfc/rfc2153.
txt;   https://www.math.utah.edu/pub/rfc/rfc1962.txt;
 https://www.math.utah.edu/pub/rfc/rfc2153.txt. Up-

dated by RFC2153 [Sim97]. Status: PROPOSED STAN-DARD.

Ratcliff:1992:AC

[Rat92] John W. Ratcliff. Audio compression. Dr. Dobb’s Journal of Software Tools, 17(7):32, 37–39, 96, 98–100, July 1992. CO-DEN DDJOEB. ISSN 1044-789X.

Ratushnyak:2008:U

[Rat08] Ratushnyak. [unknown], 2008. URL http://www.geocities.

com/SiliconValley/Bay/1995/texts22.html.

Rao:1998:WTI

[RB98] Raghuveer M. Rao and Ajit S. Bopardikar. Wavelet Transforms: Introduction to Theory and Applications. Addison-Wes-ley, Reading, MA, USA, 1998. ISBN 0-201-63463-5. xiii + 310 pp. LCCN QA403.3 .R36 1998.

Reibman:2001:MDD

[RB01] A. R. Reibman and L. Bottou. Managing drift in DCTbased scalable video coding. In Storer and Cohn [SC01c], pages 351–360. ISBN 0-7695-1031-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2001. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=917166. IEEE CSP number 01PR1031.

Razzaque:2013:CWS

[RBD13] M. A. Razzaque, Chris Bleakley, and Simon Dobson. Compression in wireless sensor networks: a survey and comparative evaluation. ACM Transactions on Sensor Networks, 10 (1):5:1–5:??, November 2013. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Rubinstein:2009:SOI

[RBR09] Benjamin I. P. Rubinstein, Peter L. Bartlett, and J. Hyam Rubinstein. Shifting: One-inclusion mistake bounds and sample compression. Journal of Computer and System Sciences, 75(1):37–59, January 2009. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022000008000676.

See corrigendum [RBR10].

Rubinstein:2010:CSO

[RBR10] Benjamin I. P. Rubinstein, Peter L. Bartlett, and J. Hyam Rubinstein. Corrigendum to “Shifting: One-inclusion mistake bounds and sample compression” [J. Comput. System Sci. 75 (1) (2009) 37–59]. Journal of Computer and System Sciences, 76(3–4):278–280, May/June 2010. CO-DEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0022000010000280. See [RBR09].

Rollins:1992:PHW

[RC92] M. Rollins and F. Carden. Possible harmonic-wavelet hybrids in image compression. In Storer and Cohn [SC92], pages 191–

199. ISBN 0-8186-2717-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=227462. IEEE catalog

number 91TH0436-6.

Roger:1996:LCA

[RC96] R. E. Roger and M. C. Cavenor. Lossless compression of AVIRIS images. IEEE Transactions on Image Processing, 5 (5):713–719, May 1996. CODEN IIPRE4. ISSN 1057-7149 (print), 1941-0042 (electronic).

Rogers:1998:RWZ

[RC98] J. K. Rogers and P. C. Cosman. Robust wavelet zerotree image compression with fixed-length packetization. In Storer and Cohn [SC98], pages 418–427. ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1998. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=672184. IEEE Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Reznik:2009:FTI

[RC09] Y. A. Reznik and R. K. Chivukula. Fast 15 × 15 transform for image and video coding applications. In Storer and Marcellin [SM09], page 465. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ????

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=4976519. IEEE Computer Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

Roder:2004:CRD

[RCH04] M. Roder, J. Cardinal, and R. Hamzaoui. On the complexity of rate-distortion optimal streaming of packetized media. In Storer and Cohn [SC04], pages 192–201. ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1281464. IEEE Computer Society Order Number: P2082.

Rizzo:2004:HPC

[RCMS04] F. Rizzo, B. Carpentieri, G. Motta, and J. A. Storer. High performance compression of hyperspectral imagery with reduced search complexity in the compressed domain. In Storer and Cohn [SC04], pages 479–488. ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1281493. IEEE Computer Society Order Number: P2082.

Ronson:1982:ABT

[RD82] J. Ronson and J. Dewitte. Adaptive block truncation coding scheme using an edge following algorithm. In Proceedings of the IEEE International Conference on Acoustics, Speech, and Signal Processing, pages 1235–1238. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1982.

Rivals:1996:GCS

[RDDD96] E. Rivals, J.-P. Delahaye, M. Dauchet, and O. Delgrange. A guaranteed compression scheme for repetitive DNA sequences. In Storer and Cohn [SC96], page ?? ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1996. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=488385. IEEE Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Ready:1972:MDC

[Rea72] Patrick John Ready. Multispectral Data Compression Through Transform Coding and Block Quantization. Ph.D. thesis, Purdue University, West Lafayette, IN, USA, 1972. 164 pp. URL

http://search.proquest.com/docview/302618152.

Ramstad:1995:SCI

[Rea95] T. A. Ramstad and et al. Subband Compression of Images: Principles and Examples. Elsevier Science Publishers B.V., Amsterdam, The Netherlands, 1995. ?? pp.

Rebaza:2012:FCA

[Reb12] Jorge Rebaza. A first course in applied mathematics. John Wiley, New York, NY, USA, 2012. ISBN 1-118-22962-2. xvi + 439 pp. LCCN TA342 .R43 2012. URL http://www.loc.

gov/catdir/enhancements/fy1201/2011043340-d.html;
http://www.loc.gov/catdir/enhancements/fy1201/2011043340-
t.html;   http://www.loc.gov/catdir/enhancements/fy1210/
2011043340-b.html.

Reghbati:1981:ODC

[Reg81] H. K. Reghbati. An overview of data compression techniques. Computer, 14(4):71–76, 1981. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Regan:1990:LRS

[Reg90] Shawn M. Regan. LZW revisited (speeding up data compression). Dr. Dobb’s Journal of Software Tools, 15(6):126–127, 167, June 1990. CODEN DDJOEB. ISSN 1044-789X.

Reimer:1994:DAC

[Rei94] Jay B. Reimer. DSP and audio compression. Dr. Dobb’s Journal of Software Tools, 19(??):63–??, ???? 1994. CODEN DDJOEB. ISSN 1044-789X.

Reznik:1998:LH

[Rez98] Y. A. Reznik. LZRW1 without hashing. In Storer and Cohn [SC98], page ?? ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1998. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=672311. IEEE

Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Reznik:2004:CPR

[Rez04] Yuriy Reznik. Coding of prediction residual in MPEG-4 standard for lossless audio coding (MPEG-4 ALS). In IEEE [IEE04], pages 1024–1027. ISBN 0-7803-8484-9. LCCN

TK7882.S65 I61 2004. URL http://ieeexplore.ieee.
org/servlet/opac?punumber=9248. IEEE Catalog Number:

04CH37568.

Reznik:2005:APS

[Rez05] Y. A. Reznik. Asymptotic properties of sample-based entropy, information divergence, and related metrics. In Storer and Cohn [SC05], page ?? ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=1402232.

Reznik:2007:MED

[Rez07a] Y. A. Reznik. Memory-efficient decoding of variable length codes for monotonic sources. In Storer and Cohn [SC07], page 398. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=4148799. IEEE Com-

puter Society Order Number P2791.

Reznik:2007:PBA

[Rez07b] Y. A. Reznik. Practical binary adaptive block coder. In Storer and Cohn [SC07], page 399. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=4148800. IEEE Computer Society Order Number P2791.

Reznik:2008:PRR

[Rez08] Y. A. Reznik. On precision-redundancy relation in the design of source coding algorithms. In Storer and Marcellin [SM08], page 539. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008;

QA76.9.D33 INTERNET. URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=4483366.

Reznik:2011:AQD

[Rez11a] Y. A. Reznik. An algorithm for quantization of discrete probability distributions. In Storer and Marcellin [SM11b], pages 333–342. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=5749491. IEEE

Computer Society Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Reznik:2011:CSW

[Rez11b] Y. A. Reznik. Coding of sets of words. In Storer and Marcellin [SM11b], pages 43–52. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=5749462. IEEE Computer Society Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Rajwan:2000:UFM

[RF00] D. Rajwan and M. Feder. Universal finite memory machines for coding binary sequences. In Storer and Cohn [SC00], pages 113–122. ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2000. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=838151. This

millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Radulovic:2005:FIA

[RF05] I. Radulovic and P. Frossard. Fast index assignment for balanced N-description scalar quantization. In Storer and Cohn [SC05], page ?? ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=
1402231.

Ramabadran:1988:TCC

[RG88] Tenkasi V. Ramabadran and Sunil S. Gaitonde. A tutorial on CRC computations. IEEE Micro, 8(4):62–75, July/August 1988. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Rath:2002:SDP

[RG02] G. Rath and C. Guillemot. Syndrome decoding and performance analysis of DFT codes with bursty erasures. In Storer and Cohn [SC02], pages 282–291. ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2002. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=999966. IEEE Computer Society Order Number PR01477.

Rein:2006:LCC

[RGF06] S. Rein, C. Guhmann, and F. H. P. Fitzek. Low-complexity compression of short messages. In Storer and Cohn [SC06], pages 123–132. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=
1607247.

Ranganathan:1991:SAL

[RH91] N. Ranganathan and S. Henriques. A systolic architecture for LZ based decompression. In Storer and Reif [SR91b], page ?? ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1990. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=213311. IEEE

Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Rao:1996:TSI

[RH96] K. R. Rao and J. J. Hwang. Techniques and Standards for Image, Video, and Audio Coding. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1996. ?? pp.

Ramos:1997:RIC

[RH97a] M. G. Ramos and S. S. Hemami. Robust image coding with perceptual-based scalability. In Storer and Cohn [SC97], page ?? ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1997. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=582133. IEEE

Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Ruhl:1997:OFC

[RH97b] M. Ruhl and H. Hartenstein. Optimal fractal coding is NPhard. In Storer and Cohn [SC97], pages 261–270. ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1997. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=582049. IEEE Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Romines:2007:HIC

[RH07] K. Romines and E. S. Hong. Hyperspectral image compression with optimization for spectral analysis. In Storer and Cohn [SC07], page 400. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

4148801. IEEE Computer Society Order Number P2791.

Ryabko:1987:CLA

[RHC87] B. Ya. Ryabko, R. Nigel Horspool, and Gordon V. Cormack. Comments to: “A locally adaptive data compression scheme” [Comm. ACM 29, 1986, 4, 320–330, MR 87e:68020] by J. L. Bentley, D. D. Sleator, R. E. Tarjan and V. K. Wei. Communications of the ACM, 30(9):792–794, September 1987. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rice:1979:SPU

[Ric79] Robert F. Rice. Some practical universal noiseless coding techniques, March 1979.

Richards:1986:DCG

[Ric86] Dana Richards. Data compression and Gray-code sorting. Information Processing Letters, 22(4):201–205, April 17, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rice:1991:SPU

[Ric91] Robert F. Rice. Some practical universal noiseless coding techniques — Part III. Module PSI14.K. JPL Publication 91-3, Jet Propulsion Laboratory, Pasadena, CA, USA, November 1991. ?? pp.

Richardson:2003:HMV

[Ric03] Iain G. Richardson. H.264 and MPEG-4 Video Compression Video Coding for Next-generation Multimedia. John Wiley, New York, NY, USA, 2003. ?? pp.

Richardson:2006:HAV

[Ric06] Iain Richardson. H.264 Advanced Video Coding, 2006. URL

http://www.vcodex.com/h264.html.

Richter:2008:EVM

[Ric08] T. Richter. Effective visual masking techniques in JPEG2000. In Storer and Marcellin [SM08], page 540. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33 INTERNET.

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=
&arnumber=4483367.

Richter:2010:SCQ

[Ric10] T. Richter. Spatial constant quantization in JPEG XR is nearly optimal. In Storer and Marcellin [SM10b], pages 79– 88. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=5453436.

Richter:2011:DBR

[Ric11] T. Richter. Deadzone based rate allocation for JPEG XR. In Storer and Marcellin [SM11b], page 474. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=5749531. IEEE Computer Society Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Richter:2012:CJJ

[Ric12] T. Richter. Compressing JPEG 2000 JPIP cache state information. In Storer and Marcellin [SM12], pages 13–21. ISBN 1-4673-0715-7 (paperback). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=6189232. IEEE Com-

puter Society order number P4656.

Rissanen:1976:GKI

[Ris76] J. J. Rissanen. Generalized Kraft inequality and arithmetic coding. IBM Journal of Research and Development, 20(3):198– 203, May 1976. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Rabbani:1991:DIC

[RJ91] Majid Rabbani and Paul W. Jones. Digital Image Compression Techniques. SPIE Optical Engineering Press, Bellingham, WA, USA, 1991. ?? pp.

Ramirez-Javega:2011:VLS

[RJL11] F. Ramirez-Javega and M. Lamarca. Variable-length source compression using successive refinement and non-linear graphbased codes. In Storer and Marcellin [SM11b], page 473. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5749530. IEEE Computer Society

Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Raittinen:1993:FBI

[RK93a] H. Raittinen and K. Kaski. Fractal based image compression with affine transformations. In Storer and Cohn [SC93], pages 244–253. ISBN 0-8186-3391-3 (microfiche), 0-8186-3392-1 (casebound). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1993. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=253125. IEEE

Computer Society Press order number 3392-02. IEEE catalog number 93TH0536-3.

Rios:1993:HPA

[RK93b] A. Rios and M. R. Kabuka. A high performance adaptive image compression system using a generative neural network: DynAmic Neural Network II (DANN II). In Storer and Cohn [SC93], pages 204–213. ISBN 0-8186-3391-3 (microfiche), 0-8186-3392-1 (casebound). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1993. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=253129. IEEE Computer Society Press order number 3392-02. IEEE catalog number 93TH0536-3.

Raittinen:1995:FSC

[RK95] H. Raittinen and K. Kaski. Fast subband coder for telephone quality audio. In Storer and Cohn [SC95], page 471. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=515581. IEEE catalog

number 95TH8037. IEEE Computer Society Press order number PR07010.

Richter:2009:MSO

[RK09a] T. Richter and Kil Joong Kim. A MS-SSIM optimal JPEG 2000 encoder. In Storer and Marcellin [SM09], pages 401–410. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=4976484. IEEE Computer

Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

Ruiters:2009:AMR

[RK09b] Roland Ruiters and Reinhard Klein. Appearance measurement and rendering: BTF compression via sparse tensor decomposition. Computer Graphics Forum, 28(4):1181–1188, June 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ristov:2015:USS

[RK15] Strahil Ristov and Damir Korencić. Using static suffix array in dynamic application: Case of text compression by longest first substitution. Information Processing Letters, 115(2):175–181, February 2015. CODEN IFPLAT. ISSN 0020-0190 (print),

1872-6119 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0020019014001835.

Ratanaworabhan:2006:FLC

[RKB06] P. Ratanaworabhan, Jian Ke, and M. Burtscher. Fast lossless compression of scientific floating-point data. In Storer and Cohn [SC06], pages 133–142. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=1607248.

Ringh:2016:MRC

[RKL16] Axel Ringh, Johan Karlsson, and Anders Lindquist. Multidimensional rational covariance extension with applications to spectral estimation and image compression. SIAM Journal on Control and Optimization, 54(4):1950–1982, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Rebmann:2003:ACC

[RKSM03] R. Rebmann, G. Krell, U. Seiffert, and B. Michaelis. Associative correction of compression artefacts with a selforganizing map classifying the image content. In Storer and Cohn [SC03], page ?? ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2003. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1194065. IEEE Computer Society Order number PR01896.

Ratnakar:1995:ROE

[RL95] Viresh Ratnakar and M. Livny. RD-OPT: an efficient algorithm for optimizing DCT quantization tables. In Storer and Cohn [SC95], pages 332–341. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ????

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=515523. IEEE catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Ratnakar:1996:ERO

[RL96] V. Ratnakar and M. Livny. Extending RD-OPT with global thresholding for JPEG optimization. In Storer and Cohn [SC96], pages 379–386. ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1996. URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

488343. IEEE Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Ristov:1999:ZLC

[RL99] Strahil Ristov and Eric Laporte. Ziv–Lempel compression of huge natural language data tries using suffix arrays. Lecture Notes in Computer Science, 1645:196–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer-ny.com/link/
service/series/0558/bibs/1645/16450196.htm;   http:
//link.springer-ny.com/link/service/series/0558/papers/
1645/16450196.pdf.

Ristov:2002:MCL

[RL02] S. Ristov and E. Laporte. A method for compressing lexicons. In Storer and Cohn [SC02], page ?? ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2002. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1000013. IEEE Computer Society Order Number PR01477.

Racz:2004:HDC

[RL04] B. Racz and A. Lukacs. High density compression of log files. In Storer and Cohn [SC04], page ?? ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1281533. IEEE Computer Society Order Number: P2082.

Richter:2008:SOA

[RL08] T. Richter and C. Larabi. Subjective and objective assessment of visual image quality metrics and still image codecs. In Storer and Marcellin [SM08], page 541. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33 INTER-

NET. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=4483368.

Rossling:2009:VCA

[RL09] Guido Rößling and Florian Lindner. Visualizing compression algorithms on-the-fly. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):376, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Rein:2009:WIT

[RLG09] S. Rein, S. Lehmann, and C. Guhmann. Wavelet image twoline coder for wireless sensor node with extremely little RAM. In Storer and Marcellin [SM09], pages 252–261. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4976469. IEEE Computer Society

Order Number P3592. BMS Part Number CFP09DCC-PRT.

Ringenburg:2004:GMI

[RLR04] M. F. Ringenburg, R. E. Ladner, and E. A. Riskin. Global MINMAX interframe bit allocation for embedded video coding. In Storer and Cohn [SC04], pages 222–231. ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1281467. IEEE Computer Society

Order Number: P2082.

Rema:2021:EHC

[RM21] N. R. Rema and P. Mythili. Extremely high compression and identification of fingerprint images using SA4 multiwavelet transform. International Journal of Image and Graphics (IJIG), 21(03):??, July 2021. ISSN 0219-

4678. URL https://www.worldscientific.com/doi/10.
1142/S0219467821500376.

Ranganathan:1991:VAD

[RMB91] N. Ranganathan, Amar Mukherjee, and M. Bassiouni. VLSI algorithms for data compression. Computer Systems Science and Engineering, 6(4):238–253, October 1991. CODEN CSSEEI. ISSN 0267-6192.

Raguram:2007:IRS

[RMB07] R. Raguram, M. W. Marcellin, and A. Bilgin. Improved resolution scalability for bi-level image data in JPEG2000. In Storer and Cohn [SC07], pages 203–212. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=4148759. IEEE Computer Society Order Number P2791.

Rose:1994:ECT

[RMG94] K. Rose, D. Miller, and A. Gersho. Entropy-constrained treestructured vector quantizer design by the minimum cross entropy principle. In Storer and Cohn [SC94], pages 12–21. ISBN 0-8186-5636-0, 0-8186-5637-9 (case). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1994.

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=305908. IEEE Catalog Number 93TH0626-2. IEEE Computer Society Press Order Number 5637-02.

Rebollo-Monedero:2005:GRD

[RMG05] D. Rebollo-Monedero and B. Girod. Generalization of the ratedistortion function for Wyner–Ziv coding of noisy sources in the quadratic-Gaussian case. In Storer and Cohn [SC05], pages 23–32. ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=1402163.

Rebollo-Monedero:2003:DOQ

[RMZG03] D. Rebollo-Monedero, Rui Zhang, and B. Girod. Design of optimal quantizers for distributed source coding. In Storer and Cohn [SC03], pages 13–22. ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2003. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1193992. IEEE Computer Society Order number PR01896.

Reyes:2010:LRC

[RN10] M. G. Reyes and D. L. Neuhoff. Lossless reduced cutset coding of Markov random fields. In Storer and Marcellin [SM10b], pages 386–395. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=
5453482.

Russo:2009:ASM

[RNOM09] Luís M. S. Russo, Gonzalo Navarro, Arlindo L. Oliveira, and Pedro Morales. Approximate string matching with compressed indexes. Algorithms (Basel), 2(3):1105–1136, September 2009. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https:

//www.mdpi.com/1999-4893/2/3/1105.

Robinson:1991:CNI

[Rob91a] J. A. Robinson. Compression of natural images using threadlike visual primitives. In Storer and Reif [SR91b], pages 307–312. ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1990. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=213350. IEEE

Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Robinson:1991:APF

[Rob91b] J. P. Robinson. Aliasing probabilities for feedback signature compression of test data. IEEE Transactions on Computers, 40(7):867–873, July 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=83625.

Robinson:1994:SSL

[Rob94a] Tony Robinson. SHORTEN: Simple lossless and nearlossless waveform compression. Technical report CUED/F-INFENG/TR.15, Cambridge University Engineering Department, Trumpington Street, Cambridge, CB2 1PZ, UK, December 1994.

Robinson:1994:SLN

[Rob94b] Tony Robinson. Simple lossless and near-lossless waveform compression. Technical Report CUED/F-INFENG/TR.156, Cambridge University, Cambridge, UK, December 1994. ??

pp. URL http://citeseer.nj.nec.com/robinson94shorten.
html;   http://citeseerx.ist.psu.edu/viewdoc/.

Robinson:1997:EGP

[Rob97] John A. Robinson. Efficient general-purpose image compression with binary tree predictive coding. IEEE Transactions on Image Processing, 6(4):601–607, April 1997. CODEN IIPRE4. ISSN 1057-7149 (print), 1941-0042 (electronic).

Rodriguez:1995:GFF

[Rod95] Karen Rodriguez. Graphics file format patent Unisys seeks royalties from GIF developers. InfoWorld, 17(2):3, January 9, 1995. CODEN INWODU. ISSN 0199-6649.

Roetling:1976:HME

[Roe76] P. G. Roetling. Halftone method with edge enhancement and moiré suppression. Journal of the Optical Society of America, 66:985–989, 1976. CODEN JOSAAH. ISSN 0030-3941.

Roetling:1977:BAC

[Roe77] P. G. Roetling. Binary approximation of continuous tone images. Photographic Science and Engineering, 21:60–65, 1977. CODEN PSENAC. ISSN 0031-8760.

Roelofs:1999:PDG

[Roe99] Greg Roelofs. PNG: The Definitive Guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-542-4. xix + 327 + 4 pp. LCCN T385 .R588 1999. US$32.95. URL

http://www.libpng.org/pub/png/pngbook.html;   http:
//www.oreilly.com/catalog/9781565925427;   http://
www.oreilly.com/catalog/pngdefg.

Roelofs:2003:PPN

[Roe03] Greg Roelofs. PNG: Portable Network Graphics: an open, extensible image format with lossless compression, 2003. URL

http://www.libpng.org/pub/png/.

Rokicki:1985:PPF

[Rok85] Tomas Rokicki. Packed (PK) font file format. TUGboat, 6 (3):115–120, November 1985. CODEN ???? ISSN 0896-

3207. URL http://www.tug.org/TUGboat/Articles/tb06-
3/tb13pk.pdf.

Rokicki:1990:GV

[Rok90] Tomas Rokicki. GFtoPK, v. 2.3, 1990. URL http://tug.org/

texlive/devsrc/Build/source/texk/web2c/gftopk.web.

Rosenfeld:1981:IM

[Ros81] Azriel Rosenfeld, editor. Image Modeling. Academic Press, New York, USA, 1981. ISBN 0-12-597320-9. LCCN T385 .I45.

Ross:1992:SDC

[Ros92] Ed Ross. A simple data-compression technique. C Users Journal, 10(10):113–??, October 1992. ISSN 0898-9788.

Ross:1994:ROD

[Ros94] John W. Ross. Record-oriented data compression. C Users Journal, 12(4):83–??, April 1994. ISSN 0898-9788.

Rossignac:1998:ECC

[Ros98] J. Rossignac. Edgebreaker: Connectivity compression for triangle meshes. Technical Report GIT-GVU-98-35, Georgia Institute of Technology, Atlanta, GA, USA, 1998. ?? pp.

Roshal:2006:RHP

[Ros06] Alexander Roshal. RARLAB home page:welcome to RAR-LAB, home of WinRAR and RAR archivers, 2006. URL http:

//www.rarlab.com/.

Roy:1987:CES

[Roy87] Bimal Chandra Roy. Compression of Equally Spaced Digital Elevation Model (Dem) Data. Ph.D. thesis, The Ohio State University, Columbus, OH, USA, 1987. 268 pp. URL http:/

/search.proquest.com/docview/303501021.

Reddy:1988:RMT

[RP88] B. R. K. Reddy and A. L. Pai. Reed–Muller transform image coding. Computer Vision, Graphics, and Image Processing, 42 (1):48–61, April 1988. CODEN CVGPDB. ISSN 0734-189X.

Rodeh:1981:LAD

[RPE81] Michael Rodeh, Vaughan R. Pratt, and Shimon Even. Linear algorithm for data compression via string matching. Journal of the ACM, 28(1):16–24, January 1981. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ramaswamy:2008:SDF

[RR08] S. Ramaswamy and K. Rose. Shared descriptions fusion coding for storage and selective retrieval of correlated sources. In Storer and Marcellin [SM08], pages 262–271. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33 INTERNET.

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=
&arnumber=4483304.

Ramakrishnan:1995:CSV

[RRG95a] Sangeeta Ramakrishnan, K. Rose, and A. Gersho. Constrainedstorage vector quantization with a universal codebook. In Storer and Cohn [SC95], pages 42–51. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ????

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=515494. IEEE catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Ruffer:1995:EIL

[RRG95b] P. Ruffer, F. Rabe, and F. Gliem. Enhancement of IMP lossy image data compression using LCT. In Storer and Cohn [SC95], page ?? ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

515602. IEEE catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Regunathan:1997:MIC

[RRG97] S. L. Regunathan, K. Rose, and S. Gadkari. Multimode image coding for noisy channels. In Storer and Cohn [SC97], pages 82–90. ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1997. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=581974. IEEE

Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Rane:2007:HRA

[RRMG07] S. Rane, D. Rebollo-Monedero, and B. Girod. High-rate analysis of systematic lossy error protection of a predictively encoded source. In Storer and Cohn [SC07], pages 263–272. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=4148765. IEEE Computer

Society Order Number P2791.

Rabiner:1978:DPS

[RS78] Lawrence R. Rabiner and Ronald W. Schafer. Digital Processing of Speech Signals. Prentice-Hall Series in Signal Processing. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1978. ?? pp.

Robinson:1981:ASC

[RS81] P. J. Robinson and Dave Singer. Another spelling correction program. Communications of the ACM, 24(5):296–297, 1981. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Reif:1998:OLC

[RS98] J. H. Reif and J. A. Storer. Optimal lossless compression of a class of dynamic sources. In Storer and Cohn [SC98], pages 501–510. ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1998. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=672221. IEEE

Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Rosiles:1999:CSU

[RS99] J.-G. Rosiles and M. J. T. Smith. Compression of SAR and ultrasound imagery using texture models. In Storer and Cohn [SC99], page ?? ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1999. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=785704. IEEE

Computer Society Order Number PR00096. IEEE Order Plan Catalog Number PR00096.

Reznik:2000:ARR

[RS00] Y. A. Reznik and W. Szpankowski. On the average redundancy rate of the Lempel–Ziv code with k-error protocol. In Storer and Cohn [SC00], pages 373–382. ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2000. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=838177. This millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Rizzo:2001:OAV

[RS01] F. Rizzo and J. A. Storer. Overlap in adaptive vector quantization. In Storer and Cohn [SC01c], pages 401–410. ISBN 0-7695-1031-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 2001. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=917171. IEEE CSP num-

ber 01PR1031.

Reznik:2002:IBT

[RS02] Y. A. Reznik and W. Szpankowski. Improved behaviour of tries by the “symmetrization” of the source. In Storer and Cohn [SC02], pages 372–381. ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2002. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=999975. IEEE Computer Society Order Number PR01477.

Ros:2004:CCB

[RS04] M. Ros and P. Sutton. Code compression based on operandfactorization for VLIW processors. In Storer and Cohn [SC04], page ?? ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=1281535. IEEE

Computer Society Order Number: P2082.

Raja:2016:EMT

[RS16] S. P. Raja and A. Suruliandi. Evaluating multiscale transform based image compression using encoding techniques. International Journal of Image and Graphics (IJIG), 16(4):1650018, October 2016. CODEN ???? ISSN 0219-4678.

RodriguesCarvalho:2021:UCC

[RS21] Daniel Rodrigues Carvalho and André Seznec. Understanding cache compression. ACM Transactions on Architecture and Code Optimization, 18(3):36:1–36:27, June 2021. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:

//dl.acm.org/doi/10.1145/3457207.

Rizzo:1999:ESP

[RSC99] F. Rizzo, J. A. Storer, and B. Carpentieri. Experiments with single-pass adaptive vector quantization. In Storer and Cohn [SC99], page ?? ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1999. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=785703. IEEE

Computer Society Order Number PR00096. IEEE Order Plan Catalog Number PR00096.

Reddy:1988:DCT

[RSK88] S. M. Reddy, K. K. Saluja, and M. G. Karpovsky. A data compression technique for built-in self-test. IEEE Transactions on Computers, 37(9):1151–1156, September 1988. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=2271.

Resende:2019:BMI

[RSMA19] João S. Resende, Patrícia R. Sousa, Rolando Martins, and Luís Antunes. Breaking MPC implementations through compression. International Journal of Information Security, 18 (4):505–518, August 2019. CODEN ???? ISSN 1615-5262 (print), 1615-5270 (electronic). URL http://link.springer.

com/article/10.1007/s10207-018-0424-2.

Rishe:2000:HPL

[RSV^{+}00] N. Rishe, A. Shaposhnikov, A. Vaschillo, D. Vasilevsky, and Shu-Ching Chen. High performance Lempel–Ziv compression using optimized longest string parsing and adaptive Huffman window size. In Storer and Cohn [SC00], page ?? ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2000. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=838215. This millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Randall:2001:LDF

[RSWW01] Keith Randall, Raymie Stata, Rajiv Wickremesinghe, and Janet L. Wiener. The LINK database: Fast access to graphs of the Web. Research Report 175, Compaq Systems Research Center, Palo Alto, CA, USA, 2001. ?? pp.

Randall:2002:LDF

[RSWW02] K. H. Randall, R. Stata, R. G. Wickremesinghe, and J. L. Wiener. The link database: fast access to graphs of the Web. In Storer and Cohn [SC02], pages 122–131. ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2002. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=999950. IEEE Computer Society Order Number PR01477.

Raite:1987:PTC

[RT87] T. Raita and J. Teuhola. Predictive test compression by hashing. In Yu and Rijsbergen [YR87], pages 223–233. ISBN 0-89791-232-2. LCCN Z 699 A1 I58 1987. US$19.

Ristad:1995:CMM

[RT95] E. S. Ristad and R. G. Thomas III. Context models in the MDL framework. In Storer and Cohn [SC95], pages 62– 71. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=515496. IEEE catalog

number 95TH8037. IEEE Computer Society Press order number PR07010.

Retvari:2016:CIF

[RTK^{+}16] Gábor Rétvári, János Tapolcai, Attila Kőrösi, András Majdán, and Zalán Heszberger. Compressing IP forwarding tables: towards entropy bounds and beyond. IEEE/ACM Transactions on Networking, 24(1):149–162, February 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Rautio:2002:SMS

[RTT02] J. Rautio, J. Tanninen, and J. Tarhio. String matching with stopper compression. In Storer and Cohn [SC02], page ?? ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2002. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=1000012. IEEE

Computer Society Order Number PR01477.

Rubin:1976:ETF

[Rub76] Frank Rubin. Experiments in text file compression. Communications of the ACM, 19(11):617–623, November 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rubin:1979:ASC

[Rub79a] F. Rubin. Arithmetic stream coding using fixed precision registers. IEEE Transactions on Information Theory, 25(6):672– 675, November 1979. CODEN IETTAW. ISSN 0018-9448 (print), 1557-9654 (electronic).

Rubin:1979:CAD

[Rub79b] Frank Rubin. Cryptographic aspects of data compression codes. Cryptologia, 3(4):202–205, October 1979. CODEN CRYPE6. ISSN 0161-1194 (print), 1558-1586

(electronic). URL http://www.informaworld.com/smpp/
content~content=a741902806~db=all~order=page. See also

Comm. ACM, 19 (November 1976): 616–623.

Rubinstein:1988:SCC

[Rub88] Roy Steven Rubinstein. Structural complexity classes of sparse sets: Intractability, data compression and printability. Ph.D. thesis, Northeastern University, Boston, MA, USA,

1988. 79 pp. URL http://search.proquest.com/docview/
303720650.

Rudner:1971:CMR

[Rud71] B. Rudner. Construction of minimum-redundancy codes with an optimum synchronization property. IEEE Transactions on Information Theory, 17(4):478–487, July 1971. CODEN IET-TAW. ISSN 0018-9448 (print), 1557-9654 (electronic).

Ruf:1994:HPF

[Ruf94] M. J. Ruf. A high performance fixed rate compression scheme for still image transmission. In Storer and Cohn [SC94], pages 294–303. ISBN 0-8186-5636-0, 0-8186-5637-9 (case). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1994. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=305937. IEEE

Catalog Number 93TH0626-2. IEEE Computer Society Press Order Number 5637-02.

Ruf:1995:TAS

[Ruf95] M. J. Ruf. Trade-off and applications of source-controlled channel decoding to still images. In Storer and Cohn [SC95], page 443. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=515553. IEEE

catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Roth:1993:DC

[RV93] Mark A. Roth and Scott J. Van Horn. Database compression. ACM SIGMOD Record, 22(3):31–39, September 1993. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Ramchandran:1994:SCE

[RV94] K. Ramchandran and M. Vetterli. Syntax-constrained encoder optimization using adaptive quantization thresholding for JPEG/MPEG coders. In Storer and Cohn [SC94], pages 146–155. ISBN 0-8186-5636-0, 0-8186-5637-9 (case). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 1994. URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=305922. IEEE Cata-

log Number 93TH0626-2. IEEE Computer Society Press Order Number 5637-02.

Ruch:2009:DDT

[RV09] David K. Ruch and Patrick J. Van Fleet. The Discrete Daubechies Transformation and Applications, chapter 7, pages 277–323. John Wiley, New York, NY, USA, 2009. ISBN 1-118-16565-9.

Rhim:2011:CDHa

[RVG11a] Joong Bum Rhim, L. R. Varshney, and V. K. Goyal. Collaboration in distributed hypothesis testing with quantized prior probabilities. In Storer and Marcellin [SM11b], pages 303–312. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=5749488. IEEE

Computer Society Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Rhim:2011:CDHb

[RVG11b] Joong Bum Rhim, L. R. Varshney, and V. K. Goyal. Conflict in distributed hypothesis testing with quantized prior probabilities. In Storer and Marcellin [SM11b], pages 313–322. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5749489. IEEE Computer Society

Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Rao:1990:DCT

[RY90] K. R. Rao and P. Yip. Discrete Cosine Transform— Algorithms, Advantages, Applications. Academic Press, New York, USA, 1990. ?? pp.

Reif:1992:OTI

[RY92] J. H. Reif and A. Yoshida. Optical techniques for image compression. In Storer and Cohn [SC92], pages 32–40. ISBN 0-8186-2717-4. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=227478. IEEE catalog number

91TH0436-6.

Rao:2001:TDC

[RY01] K. Ramamohan (Kamisetty Ramamohan) Rao and P. C. (Pat C.) Yip, editors. The Transform and Data Compression Handbook. Electrical engineering and signal processing series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2001. ISBN 0-8493-3692-9, 1-315-22052-0, 1-351-82772-3, 1-4200-3738-2 (e-book). xix + 388 pp. LCCN TK5105 .T72 2001. URL http://www.loc.gov/

catdir/enhancements/fy0646/00057149-d.html.

Ryabko:2019:TUD

[Rya19] Boris Ryabko. Time-universal data compression. Algorithms (Basel), 12(6), June 2019. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/12/6/
116.

Rao:2000:TDC

[RYe00] Kamisetty Ramamohan Rao, Patrick C. Yip, and editors. The Transform and Data Compression Handbook. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2000. ?? pp.

Rytter:2002:ALZ

[Ryt02] Wojciech Rytter. Application of Lempel–Ziv factorization to the approximation of grammar-based compression. Lecture Notes in Computer Science, 2373:20–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer-ny.com/link/
service/series/0558/bibs/2373/23730020.htm;   http:
//link.springer-ny.com/link/service/series/0558/papers/
2373/23730020.pdf.

Rytter:2003:ALZ

[Ryt03] Wojciech Rytter. Application of Lempel–Ziv factorization to the approximation of grammar-based compression. Theoretical Computer Science, 302(1–3):211–222, June 13, 2003. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sacco:1988:ITS

[S^{+}88] William Sacco et al. Information Theory, Saving Bits. Janson Publications, Providence, RI, USA, 1988. ?? pp.

Stuart:1999:MLC

[S^{+}99] J. R. Stuart et al. MLP lossless compression. In AES 9th Regional Convention, Tokyo, page ?? ????, ????, 1999.

URL http://www.meridian-audio.com/w_paper/mlp_jap_
new.PDF.

Self:2009:AE

[S^{+}09] Douglas Self et al. Audio engineering. The Newnes know it all series. Newnes/Elsevier, Amsterdam, The Netherlands, 2009. ISBN 1-85617-526-X. xvii + 907 pp. LCCN TK7881.4 .A9235 2009.

Simoncelli:1990:ST

[SA90] Eero P. Simoncelli and Edward. H. Adelson. Subband transforms. In John Woods, editor, Subband Coding, pages 143–192. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1990.

Sayood:1992:DLI

[SA92] Khalid Sayood and K. Anderson. A differential lossless image compression scheme. IEEE Transactions on Signal Processing, 40(1):236–241, January 1992. CODEN ITPRED. ISSN 1053-587X (print), 1941-0476 (electronic).

Sampath:1993:CPS

[SA93] Ashwin Sampath and Ahmad C. Ansari. Combined Peano scan and VQ approach to image compression. Image and Video Processing, 1903:175–186, 1993. ISSN 1088-0356.

Sehgal:2003:RPC

[SA03] A. Sehgal and N. Ahuja. Robust predictive coding and the Wyner–Ziv problem. In Storer and Cohn [SC03], pages 103–112. ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2003. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1194001. IEEE Computer Society Order number PR01896.

Sabourin:1994:CCP

[Sab94] Conrad F. Sabourin. Computational character processing: character coding, input, output, synthesis, ordering, conversion, text compression, encryption, display hashing, literate programming: bibliography. Infolingua, Montréal, PQ, Canada, 1994. ISBN 2-921173-18-2. vii + 579 pp. LCCN ????

Sadeh:1993:ASM

[Sad93] I. Sadeh. On approximate string matching. In Storer and Cohn [SC93], pages 148–157. ISBN 0-8186-3391-3 (microfiche), 0-8186-3392-1 (casebound). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1993. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=253135. IEEE Computer Society Press order number 3392-02. IEEE catalog number 93TH0536-3.

Sadeh:1996:ODC

[Sad96a] I. Sadeh. Optimal data compression algorithm. Computers and Mathematics with Applications, 32(5):57–72, September 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-

7668 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0898122196001356.

Sadeh:1996:UDC

[Sad96b] Ilan Sadeh. Universal data compression algorithm based on approximate string matching. Probability in the Engineering and Informational Sciences, 10(4):465–486, October 1996. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
F03252060CD00649B02534E0CF74CD68.

Sadakane:1998:FAM

[Sad98a] K. Sadakane. A fast algorithm for making suffix arrays and for Burrows–Wheeler transformation. In Storer and Cohn [SC98], pages 129–138. ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1998. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=672139. IEEE

Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Sadakane:1998:OVB

[Sad98b] K. Sadakane. On optimality of variants of the block sorting compression. In Storer and Cohn [SC98], page ?? ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1998. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=672312. IEEE Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Sadeh:1998:BDC

[Sad98c] Ilan Sadeh. Bounds on data compression ratio with a given tolerable error probability. Probability in the Engineering and Informational Sciences, 12(2):189–210, April 1998. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
A2EDAF726A22B80AE89F80B7B15BC62B.

Sadakane:1999:MBW

[Sad99] K. Sadakane. A modified Burrows–Wheeler transformation for case-insensitive search with application to suffix array compression. In Storer and Cohn [SC99], page ?? ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1999. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=785705. IEEE Computer Society Or-

der Number PR00096. IEEE Order Plan Catalog Number PR00096.

Sagan:1994:SFC

[Sag94] Hans Sagan. Space-Filling Curves. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ?? pp.

Said:2004:CAA

[Sai04] A. Said. Comparative analysis of arithmetic coding computational complexity. In Storer and Cohn [SC04], page ?? ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1281538. IEEE Computer

Society Order Number: P2082.

Said:2005:EAP

[Sai05a] A. Said. Efficient alphabet partitioning algorithms for lowcomplexity entropy coding. In Storer and Cohn [SC05], pages 183–192. ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=1402179.

Said:2005:IGR

[Sai05b] A. Said. On the inadequacy of Golomb–Rice codes for adaptive coding. In Storer and Cohn [SC05], page ?? ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1402233.

Salomonson:1991:MRI

[Sal91] V. V. Salomonson. The Moderate Resolution Imaging Spectrometer. an EOS facility instrument candidate for application of data compression methods. In Storer and Reif [SR91b], pages 468–471. ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1990. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=213296. IEEE

Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Salomon:1998:DCC

[Sal98] David Salomon. Data Compression: The Complete Reference. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. ISBN 0-387-98280-9. xx + 427 pp. LCCN QA76.9.D33S25 1997. US$39.95.

URL http://www.booksbydavidsalomon.com/;   http://
www.ecs.csun.edu/~dxs/DCadvertis/DcompAd.html.

Salomon:1999:CGG

[Sal99] David Salomon. Computer Graphics and Geometric Modeling. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. ISBN 0-387-98682-0. xviii + 851 pp. LCCN T385 S243 1999. URL

http://www.booksbydavidsalomon.com/;   http://www.
ecs.csun.edu/~dxs/CGGMadvertis/CGad.html.

Salomon:2000:DCC

[Sal00a] David Salomon. Data Compression: The Complete Reference. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2000. ISBN 0-387-95045-1. xvi + 823 pp. LCCN QA76.9.D33 S25 2000. US$39.95.

URL http://www.booksbydavidsalomon.com/;   http://
www.ecs.csun.edu/~dxs/DCadvertis/Dcomp2Ad.html.

Salomon:2000:PCS

[Sal00b] David Salomon. Prefix compression of sparse binary strings. ACM Crossroads, 6(3):??, February 2000. ISSN 1528-4972 (print), 1528-4980 (electronic).

Salomon:2002:CSD

[Sal02a] David Salomon. Correlation in statistics and in data

compression, 2002. URL http://www.davidsalomon.name/
DC2advertis/Corr.pdf.

Salomon:2002:GDC

[Sal02b] David Salomon. A Guide to Data Compression Methods. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. ISBN 0-387-95260-8 (paperback). xii + 295 pp. LCCN QA76.9.D33 S28 2001. URL http://

www.booksbydavidsalomon.com/.

Salomon:2003:DPS

[Sal03] David Salomon. Data Privacy and Security. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. ISBN 0-387-00311-8. xiv + 465 pp. LCCN QA76.9.A25 S265 2003. US$59.95. URL http://www.

booksbydavidsalomon.com/.

Salomon:2004:DCC

[Sal04] David Salomon. Data Compression: The Complete Reference. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 2004. ISBN 0-387-40697-2. xx + 898 pp. LCCN QA76.9.D33 S25 2004. URL

http://www.booksbydavidsalomon.com/;   http://www.
ecs.csun.edu/~dsalomon/DC3advertis/DComp3Ad.html.

Salomon:2005:CDC

[Sal05] David Salomon. Coding for data and computer communications. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. ISBN 0-387-21245-0. xv + 548 pp. LCCN TK5102.94 .S35 2005. URL http:

//www.DavidSalomon.name/Codes/Codes.html;   http://
www.ecs.csun.edu/~dsalomon/.

Salomon:2006:CSC

[Sal06] David Salomon. Curves and Surfaces for Computer Graphics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. ISBN 0-387-24196-5, 0-387-28452-4 (e-book). xvi + 460 pp. LCCN T385 .S2434 2005.

Salomon:2007:DCC

[Sal07a] David Salomon. Data Compression: The Complete Reference. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., fourth edition, 2007. ISBN 1-84628-602-6, 1-84628-603-4 (e-book). xxv + 1092 pp. LCCN QA76.9.D33 S25 2007. With contributions by Giovanni Motta and David Bryant.

Salomon:2007:VLC

[Sal07b] David Salomon. Variable-length Codes for Data Compression. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 1-84628-958-0 (paperback), 1-84628-959-9 (ebook). xii + 191 pp. LCCN QA76.9.D33 S25 2007.

Salomon:2008:CID

[Sal08] David Salomon. A Concise Introduction to Data Compression. Undergraduate topics in computer science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. ISBN 1-84800-071-5 (paperback). xiii + 310 pp. LCCN QA76.9.D33 S34 2008. EUR 32.

Salman:2013:HDD

[Sal13] Raied Salman. Hamming distance and data compression of 1-D CA. In Computer Science & Information Technology (CS & IT). Academy & Industry Research Collaboration Center (AIRCC), ????, May 2013.

Salikhov:2018:ICD

[Sal18] Kamil Salikhov. Improved compression of DNA sequencing data with cascading Bloom filters. International Journal of Foundations of Computer Science (IJFCS), 29(8):??, December 2018. ISSN 0129-0541.

Samet:1985:DSQ

[Sam85] Hanan Samet. Data structures for quadtree approximation and compression. Communications of the ACM, 28(9):973– 993, September 1985. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.org/

pubs/toc/Abstracts/0001-0782/4290.html.

Samples:1989:MNL

[Sam89] A. D. Samples. Mache: no-loss trace compaction. ACM SIG-METRICS Performance Evaluation Review, 17(1):89–97, May 1989. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Samet:1990:ASD

[Sam90a] Hanan Samet. Applications of Spatial Data Structures: Computer Graphics, Image Processing, and GIS. Addison-Wesley, Reading, MA, USA, 1990. ?? pp.

Samet:1990:DAS

[Sam90b] Hanan Samet. The Design and Analysis of Spatial Data Structures. Addison-Wesley, Reading, MA, USA, 1990. ?? pp.

Sandman:1989:EPA

[San89] Aubrey Max Sandman. Errors — a Positive Approach (Polynomial Data Compression). Ph.D. thesis, The City University, London, UK, 1989. 187 pp. URL http://search.proquest.

com/docview/303820371.

Sachs:2000:WIT

[SAR00] D. G. Sachs, R. Anand, and K. Ramchandran. Wireless image transmission using multiple-description based concatenated codes. In Storer and Cohn [SC00], page ?? ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2000. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=838216. This millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Sartipi:2010:LCI

[Sar10] M. Sartipi. LDPC codes for information embedding and lossy distributed source coding. In Storer and Marcellin [SM10b], page 551. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=5453526.

Sartipi:2012:LCD

[Sar12] M. Sartipi. Low-complexity distributed compression in wireless sensor networks. In Storer and Marcellin [SM12], pages 227–236. ISBN 1-4673-0715-7 (paperback). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

6189254. IEEE Computer Society order number P4656.

Shermer:2010:NMP

[SAS10] E. Shermer, M. Avigal, and D. Shapira. Neural Markovian predictive compression: An algorithm for online lossless data compression. In Storer and Marcellin [SM10b], pages 209– 218. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=5453465.

Saupe:1994:BTC

[Sau94] Dietmar Saupe. Breaking the time complexity of fractal image compression. Technical Report 53, Institut für Informatik, Universität Freiburg, Freiburg, Germany,

1994.  URL ftp://ftp.informatik.uni-freiburg.de/
papers/fractal/Saup94a.ps.Z.

Saupe:1995:AFI

[Sau95] D. Saupe. Accelerating fractal image compression by multidimensional nearest neighbor search. In Storer and Cohn [SC95], pages 222–231. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

515512. IEEE catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Saupe:1996:LDP

[Sau96] Dietmar Saupe. Lean domain pools for fractal image compression. Proceedings of the SPIE — The International Society for Optical Engineering, 2669(??):150–157, January 1996. CODEN PSISDG. ISSN ????

Savari:1997:RLZ

[Sav97] S. A. Savari. Redundancy of the Lempel–Ziv–Welch code. In Storer and Cohn [SC97], pages 191–200. ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1997. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=582011. IEEE Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Savari:1998:VFL

[Sav98] S. A. Savari. Variable-to-fixed length codes for predictable sources. In Storer and Cohn [SC98], pages 481–490. ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1998. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=672213. IEEE Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Savari:1999:VFL

[Sav99] S. A. Savari. Variable-to-fixed length codes and plurally parsable dictionaries. In Storer and Cohn [SC99], pages 453–462. ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 1999. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=755695. IEEE Computer

Society Order Number PR00096. IEEE Order Plan Catalog Number PR00096.

Savari:2003:CIC

[Sav03] S. A. Savari. On compressing interchange classes of events in a concurrent system. In Storer and Cohn [SC03], pages 153–162. ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2003. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1194006. IEEE Computer Society Order number PR01896.

Savari:2009:MRF

[Sav09] S. A. Savari. On minimum-redundancy fix-free codes. In Storer and Marcellin [SM09], pages 3–12. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=4976444. IEEE Computer Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

Sayano:1992:ACC

[Say92] Masahiro Sayano. Analyses of coding and compression strategies for data storage and transmission. Ph.D. thesis, California Institute of Technology, Pasadena, CA, USA, 1992. 121 pp.

URL http://search.proquest.com/docview/303987736.

Sayood:1996:IDC

[Say96] Khalid Sayood. Introduction to Data Compression. Morgan Kaufmann Publishers, San Francisco, CA, USA, 1996. ISBN 1-55860-346-8. xx + 636 pp. LCCN TK5102.92 .S39 1996.

Sayir:1998:CSC

[Say98] J. Sayir. Conditional source coding with competitive lists. In Storer and Cohn [SC98], page ?? ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1998. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=672313. IEEE Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Sayood:2000:IDC

[Say00] Khalid Sayood. Introduction to Data Compression. Morgan Kaufmann Publishers, 2929 Campus Drive, Suite 260, San Mateo, CA 94403, USA, second edition, 2000. ISBN 1-55860-558-4. xx + 636 pp. LCCN TK5102.92 .S39

2000. US$69.95. URL http://www.mkp.com/books_catalog/
catalog.asp?ISBN=1-55860-558-4.

Sayood:2003:LCH

[Say03] Khalid Sayood, editor. Lossless Compression Handbook. Academic Press series in communications, networking and multimedia. Academic Press, New York, USA, 2003. ISBN 0-12-620861-1. xx + 455 pp. LCCN TK5102.5 .L69 2003.

URL http://proquest.safaribooksonline.com/?uiCode=
stanford\%26xmlId=9780126208610;   http://www.loc.
gov/catdir/description/els031/2002104270.html;   http:
//www.loc.gov/catdir/toc/els031/2002104270.html;
http://www.myilibrary.com?id=104988;   http://www.
sciencedirect.com/science/book/9780126208610.

Sayood:2005:IDC

[Say05] Khalid Sayood. Introduction to data compression. Elsevier, Amsterdam, The Netherlands, third edition, 2005. ISBN 0-12-620862-X. xxii + 680 pp. LCCN TK5102.92 .S39 2005.

Sayood:2006:IDC

[Say06] Khalid Sayood. Introduction to Data Compression. Morgan Kaufmann series in multimedia information and systems. Elsevier, Amsterdam, The Netherlands, third edition, 2006. ISBN 0-12-620862-X. xxii + 680 pp. LCCN TK5102.92 .S39 2006.

URL http://www.loc.gov/catdir/enhancements/fy0625/
2005052759-d.html;   http://www.loc.gov/catdir/enhancements/
fy0625/2005052759-t.html.

Sayood:2012:IDC

[Say12] Khalid Sayood. Introduction to data compression. Morgan Kaufmann series in multimedia information and systems. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, fourth edition, 2012. ISBN 0-12-415796-3, 0-12-415796-3 (e-book). xxiii + 740 pp. LCCN QA76.9.D33 S29 2012; TK5102.92 .S39

2012. URL http://www.sciencedirect.com/science/book/
9780124157965.

Savoji:1985:DMB

[SB85] M. H. Savoji and R. E. Burge. On different methods based on the Karhunen-Loeve expansion and used in image analysis. Computer Vision, Graphics, and Image Processing, 29(2):259– 269, February 1985. CODEN CVGPDB. ISSN 0734-189X.

Sedgewick:1997:FAS

[SB97a] R. Sedgewick and J. Bentley. Fast algorithms for sorting and searching strings. In ACM [ACM97], pages 360–369. CODEN PAAAF2. ISBN 0-89871-390-0. LCCN ???? URL http://

www.cs.princeton.edu/~rs/talks/strings.ps. This is the

fourth of four key papers behind the bzip2 compression tools. The others are [HL90, BW94b, Whe97].

Subrahmanya:1997:MDE

[SB97b] P. Subrahmanya and T. Berger. Multiple descriptions encoding of images. In Storer and Cohn [SC97], page ?? ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1997. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=582138. IEEE Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Sculley:2006:CML

[SB06] D. Sculley and C. E. Brodley. Compression and machine learning: a new perspective on feature space vectors. In Storer and Cohn [SC06], pages 332–341. ISBN 0-7695-2545-

8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=1607268.

Scandolo:2016:ERM

[SBE16] Leonardo Scandolo, Pablo Bauszat, and Elmar Eisemann. Efficient rendering: Merged multiresolution hierarchies for shadow map compression. Computer Graphics Forum, 35(7):383–390, October 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Storer:1992:DCC

[SC92] James A. (James Andrew) Storer and Martin Cohn, editors. Data Compression Conference: DDC ’92, March 24–27, 1992 at Cliff Lodge, Snowbird, Utah. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1992. ISBN 0-8186-2717-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? IEEE catalog number 91TH0436-6.

Storer:1993:DDC

[SC93] James A. (James Andrew) Storer and Martin Cohn, editors. DCC ’93: Data Compression Conference: [March 30– April 2, 1993, Snowbird, Utah: proceedings]. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN 0-8186-3391-3 (microfiche), 0-8186-3392-1 (casebound). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1993. URL http://

ieeexplore.ieee.org/servlet/opac?punumber=452. IEEE

Computer Society Press order number 3392-02. IEEE catalog number 93TH0536-3.

Storer:1994:DDC

[SC94] James A. (James Andrew) Storer and Martin Cohn, editors. DCC ’94: Data Compression Conference, proceedings: March 29–31, 1994, Snowbird, Utah. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-5636-0, 0-8186-5637-9 (case). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1994. URL http://ieeexplore.ieee.org/servlet/opac?

punumber=961. IEEE Catalog Number 93TH0626-2. IEEE Computer Society Press Order Number 5637-02.

Storer:1995:DDC

[SC95] James A. (James Andrew) Storer and Martin Cohn, editors. DCC ’95, Data Compression Conference: March 28–30, 1995, Snowbird, Utah. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/servlet/

opac?punumber=3874. IEEE catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Storer:1996:DDC

[SC96] James A. (James Andrew) Storer and Martin Cohn, editors. DCC ’96: Data Compression Conference, March 31–April 3, 1996, Snowbird, Utah. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33

D37 1996. URL http://ieeexplore.ieee.org/servlet/

opac?punumber=3509. IEEE Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Storer:1997:DDC

[SC97] James A. (James Andrew) Storer and Martin Cohn, editors. DCC ’97: Data Compression Conference, March 25–27, 1997, Snowbird, Utah. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33

D232 1997.  URL http://ieeexplore.ieee.org/xpl/
mostRecentIssue.jsp?punumber=4456. IEEE Computer So-

ciety Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Storer:1998:PDD

[SC98] James A. (James Andrew) Storer and Martin Cohn, editors. Proceedings: DCC ’98: Data Compression Conference: March 30–April 1, 1998, Snowbird, Utah. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1998. ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1998. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=672118. IEEE

Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Storer:1999:DPD

[SC99] James A. Storer and Martin Cohn, editors. DCC ’99: proceedings: Data Compression Conference: March 29–31, 1999, Snowbird, Utah. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1999. ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1999. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=755647. IEEE Computer Society Or-

der Number PR00096. IEEE Order Plan Catalog Number PR00096.

Storer:2000:DPD

[SC00] James A. (James Andrew) Storer and Martin Cohn, editors. DCC 2000: Proceedings, Data Compression Conference, March 28–30, 2000, Snowbird, Utah. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2000. ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2000. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=838139. This

millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Schilling:2001:FPI

[SC01a] D. Schilling and P. Cosman. Feature-preserving image coding for very low bit rates. In Storer and Cohn [SC01c], pages 103–112. ISBN 0-7695-1031-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2001. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=917141. IEEE CSP number 01PR1031.

Shamir:2001:ULC

[SC01b] G. I. Shamir and D. J. Costello, Jr. Universal lossless compression of piecewise stationary slowly varying sources. In Storer and Cohn [SC01c], pages 371–380. ISBN 0-7695-1031-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 2001. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=917168. IEEE CSP num-

ber 01PR1031.

Storer:2001:DPD

[SC01c] James A. (James Andrew) Storer and Martin Cohn, editors. DCC 2001: proceedings, Data Compression Conference, March 27–29, 2001, Snowbird, Utah. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2001. ISBN 0-7695-1031-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2001. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=917130. IEEE CSP number 01PR1031.

Storer:2002:DPD

[SC02] James A. (James Andrew) Storer and Martin Cohn, editors. DCC 2002: proceedings: Data Compression Conference: April 2–4, 2002, Snowbird, Utah. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2002. ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2002. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=999937. IEEE

Computer Society Order Number PR01477.

Storer:2003:DPD

[SC03] James A. (James Andrew) Storer and Martin Cohn, editors. DCC 2003: proceedings: Data Compression Conference: March 25–27, 2003, Snowbird, Utah. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2003. ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2003. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1193990. IEEE Computer Society Order number PR01896.

Storer:2004:DCC

[SC04] James A. (James Andrew) Storer and Martin Cohn, editors. Data Compression Conference, 2004. Proceedings. DCC 2004: Snowbird, Utah, USA: 23–25 March 2004. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2004. ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1281440. IEEE Computer Society Order Number: P2082.

Storer:2005:DCC

[SC05] James A. (James Andrew) Storer and Martin Cohn, editors. Data Compression Conference, 2005. Proceedings. DCC 2005. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2005. ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/servlet/opac?
punumber=9633;   http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1402158.

Storer:2006:DPD

[SC06] James A. (James Andrew) Storer and Martin Cohn, editors. DCC 2006: Proceedings Data Compression Conference: 28–30 March 2006, Snowbird, Utah. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2006. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=1607232.

Storer:2007:DPD

[SC07] James A. (James Andrew) Storer and Martin Cohn, editors. DCC 2007: Proceedings: Data Compression Conference: 27–29 March 2007, Snowbird, Utah. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2007. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/servlet/opac?

punumber=4148731. IEEE Computer Society Order Number P2791.

Soatto:2011:CRB

[SC11] S. Soatto and A. Chiuso. Controlled recognition bounds for scaling and occlusion channels. In Storer and Marcellin [SM11b], page 477. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

5749534. IEEE Computer Society Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Storer:20xx:DXD

[SCxx] James A. Storer and Martin Cohn, editors. DCC ’XX: Data Compression Conference. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 20xx.

ISSN 1068-0314 (print), 2375-0359 (electronic). Annual conference proceedings series, mostly held at Snowbird, UT, USA.

Shih:2012:APB

[SCC12] Y.-T. Shih, C.-S. Chien, and C.-Y. Chuang. An adaptive parameterized block-based singular value decomposition for image de-noising and compression. Applied Mathematics and Computation, 218(21):10370–10385, July 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0096300312003281.

Schneider:1970:RDC

[Sch70] Kenneth Stuart Schneider. Reliable Data Compression of Constant Rate Markov Sources for Fixed Rate Channels. Ph.D. thesis, Cornell University, Ithaca, NY, USA, 1970. 343 pp.

URL http://search.proquest.com/docview/302518830.

Schalkwijk:1972:ASC

[Sch72] J. Pieter M. Schalkwijk. An algorithm for source coding. IEEE Transactions on Information Theory, 18(3):395–399, May 1972. CODEN IETTAW. ISSN 0018-9448 (print), 1557-9654 (electronic).

Schmidt:1986:SRC

[Sch86] Alfred Rudolf Schmidt. Secondary range compression for improved range/Doppler processing of SAR data with high squint. M.A.Sc, The University of British Columbia, Vancouver, BC, Canada, 1986. xviii + 213 pp. URL http://search.

proquest.com/docview/303539025.

Schindler:1997:FBS

[Sch97] M. Schindler. A fast block-sorting algorithm for lossless data compression. In Storer and Cohn [SC97], page ?? ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1997. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=582137. IEEE Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Schindler:1998:FRA

[Sch98] Michael Schindler. A fast renormalisation for arithmetic coding. In Storer and Cohn [SC98], page ?? ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1998.

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=
&arnumber=672314;   http://www.compressconsult.com/

rangecoder/. A poster in the Data Compression Conference, 1998.

Schmalz:2003:MDI

[Sch03] Mark S. Schmalz, editor. Mathematics of data/image coding, compression, and encryption VI, with applications, 5–7 August 2003, San Diego, California, USA, volume 5208 of Proceedings of SPIE. SPIE Optical Engineering Press, Bellingham, WA, USA, 2003. ISBN 0-8194-5081-2. LCCN ????

Schmalz:2004:MDI

[Sch04] Mark S. Schmalz, editor. Mathematics of data/image coding, compression, and encryption VII, with applications: 4–5 August 2004, Denver, Colorado, USA, volume 5561 of Proceedings of SPIE. SPIE Optical Engineering Press, Bellingham, WA, USA, 2004. ISBN 0-8194-5499-0. LCCN ????

Schmalz:2005:RAO

[Sch05] M. S. Schmalz. Recent advances in object-based image compression. In Storer and Cohn [SC05], page ?? ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1402235.

Schmalz:2006:MMM

[Sch06] M. S. Schmalz. Multi-modal, multi-fractal boundary encoding in object-based image compression. In Storer and Cohn [SC06], page 465. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=1607308.

Schlick:2010:MMS

[Sch10] Tamar Schlick. Molecular modeling and simulation: an interdisciplinary guide, volume 21 of Interdisciplinary applied mathematics. Springer Science+Business Media, LLC, New York, NY, USA, second edition, 2010. ISBN 1-4419-6351-0, 1-4419-6350-2. xlv + 723 pp. LCCN QD480 .S37 2010.

Schneider:2011:SRP

[Sch11] J. Schneider. Set reordering for paletted data. In Storer and Marcellin [SM11b], page 475. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ????

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=5749532. IEEE Computer Society Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Schneider:2013:NCC

[Sch13] D. Schneider. New camera chip captures only what it needs. IEEE Spectrum, 50(3):13–14, March 2013. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Son:2014:DCE

[SCH^{+}14] Seung Woo Son, Zhengzhang Chen, William Hendrix, Ankit Agrawal, Wei keng Liao, and Alok Choudhary. Data compression for the exascale computing era — survey. Supercomputing Frontiers and Innovations, 1(2):76–88, ???? 2014. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (electronic). URL

http://superfri.org/superfri/article/view/13.

Sermadevi:2005:WBA

[SCHB05] Y. Sermadevi, Jun Chen, S. S. Hemami, and T. Berger. When is bit allocation for predictive video coding easy? In Storer and Cohn [SC05], pages 289–298. ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=1402190.

Schiller:1997:SRS

[SCKS97] I. Schiller, C.-K. Chuang, S. M. King, and J. A. Storer. Selective resolution for surveillance video compression. In Storer and Cohn [SC97], page ?? ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1997.

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=582136. IEEE Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Shen:2005:VCT

[SCM05] Yushi Shen, P. C. Cosman, and L. B. Milstein. Video coding for a time varying tandem channel with feedback. In Storer and Cohn [SC05], page ?? ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=1402237.

Scott:2002:AAO

[Sco02] David A. Scott. Algorithm alley: Optimal EOF handling with arithmetic compression. Dr. Dobb’s Journal of Software Tools, 27(1):99, 102, January 2002. CODEN DDJOEB. ISSN

1044-789X. URL http://www.ddj.com/ftp/2002/2002_01/
aa0102.zip.

Sezer:2011:RLD

[SCV11] O. G. Sezer, R. Cohen, and A. Vetro. Robust learning of 2-D separable transforms for next-generation video coding. In Storer and Marcellin [SM11b], pages 63–72. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=5749464. IEEE Computer Society Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Shen:2011:EDV

[SCW11] Yun-Chung Shen, Han-Ping Cheng, and Ja-Ling Wu. An efficient distributed video coding with parallelized design for concurrent computing. In Storer and Marcellin [SM11b], page 476. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=5749533. IEEE

Computer Society Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Scherer:2002:BMC

[SD02] Markus W. Scherer and Mark Davis. BOCU-1: MIMEcompatible Unicode compression. Unicode Technical Note 6, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-mail: unicode-inc@unicode.org, August 9, 2002.

URL http://www.unicode.org/notes/tn6/. Last revision:

2006-02-04.

Scherer:2006:PDU

[SD06] Markus W. Scherer and Mark Davis. Proposed draft Unicode technical standard #40: BOCU-1: MIME-compatible Unicode compression. Technical report, The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-mail: unicode-inc@

unicode.org, February 3, 2006. URL http://www.unicode.
org/reports/tr40/tr40-1.html.

Sebe:2000:SDI

[SDFH00] Francesc Sebé, Josep Domingo-Ferrer, and Jordi Herrera. Spatial-domain image watermarking robust against compression, filtering, cropping and scaling. Lecture Notes in Computer Science, 1975:44–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/0558/bibs/1975/
19750044.htm;   http://link.springer-ny.com/link/service/
series/0558/papers/1975/19750044.pdf.

Senecal:2004:LLV

[SDJ04a] J. Senecal, M. Duchaineau, and K. I. Joy. Length-limited variable-to-variable length codes for high-performance entropy coding. In Storer and Cohn [SC04], pages 389–398. ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1281484. IEEE Computer Society

Order Number: P2082.

Senecal:2004:RBB

[SDJ04b] J. Senecal, M. Duchaineau, and K. I. Joy. Reversible n-bit to n-bit integer Haar-like transforms. In Storer and Cohn [SC04], page ?? ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=1281540. IEEE

Computer Society Order Number: P2082.

Staller:2002:IJS

[SDM02] Alexander Staller, Peter Dillinger, and Reinhard Männer. Implementation of the JPEG 2000 standard on a virtex 1000 FPGA. Lecture Notes in Computer Science, 2438:503–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer-ny.com/link/

service/series/0558/bibs/2438/24380503.htm;   http:
//link.springer-ny.com/link/service/series/0558/papers/
2438/24380503.pdf.

Stollnitz:1995:WCGa

[SDS95] Eric J. Stollnitz, Tony D. DeRose, and David H. Salesin. Wavelets for computer graphics: a primer, part 1. IEEE Computer Graphics and Applications, 15(3):76–84, May 1995. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Stollnitz:1996:WCG

[SDS96] E. J. Stollnitz, T. D. DeRose, and D. H. Salesin. Wavelets for Computer Graphics. Morgan Kaufmann Publishers, San Francisco, CA, USA, 1996. ?? pp.

Syahrul:2011:CTA

[SDV11] E. Syahrul, J. Dubois, and V. Vajnovszki. Combinatorial transforms: Applications in lossless image compression. Journal of Discrete Mathematical Sciences and Cryptography, 14(2):129–147, April 2011. CODEN ???? ISSN 0972-

0529. URL http://www.connectjournals.com/achivestoc.
php?bookmark=CJ-003072&volume=14&issue_id=02.

Searfoss:1990:BBD

[Sea90] Glenn Searfoss. Bounding box data compression. Dr. Dobb’s Journal of Software Tools, 15(4):56–64, 108, April 1990. CO-DEN DDJOEB. ISSN 1044-789X.

SantaCruz:2001:JIC

[SEC01] Diego Santa Cruz, Touradj Ebrahimi, and Charilaos Christopoulos. The JPEG 2000 image coding standard. Dr. Dobb’s Journal of Software Tools, 26(4):46, 48, 50–52, 54, April 2001. CODEN DDJOEB. ISSN 1044-789X.

URL http://jj2000.epfl.ch/;   http://www.ddj.com/
; http://www.ece.ubc.ca/~mdadams/jasper/;   http://
www.imagepower.com/products/ubcform.htm; http://www.
jpeg.org/;   http://www.jpeg.org/public/wavedemo.zip.

Senft:2006:CST

[Sen06] M. Senft. Compressed by the suffix tree. In Storer and Cohn [SC06], pages 183–192. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ????

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=
&arnumber=1607253.

Servetto:2000:LQS

[Ser00] S. D. Servetto. Lattice quantization with side information. In Storer and Cohn [SC00], pages 510–519. ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2000. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=838191. This millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Seward:2000:PBS

[Sew00] J. Seward. On the performance of BWT sorting algorithms. In Storer and Cohn [SC00], pages 173–182. ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2000. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=838157. This millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Seward:2001:STT

[Sew01] J. Seward. Space-time tradeoffs in the inverse B-W transform. In Storer and Cohn [SC01c], pages 439–448. ISBN 0-7695-1031-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 2001. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=917175. IEEE CSP num-

ber 01PR1031.

Sartipi:2005:DSC

[SF05] M. Sartipi and F. Fekri. Distributed source coding in wireless sensor networks using LDPC codes: a non-uniform framework. In Storer and Cohn [SC05], page ?? ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=1402234.

Shayevitz:2009:PMF

[SF09] O. Shayevitz and M. Feder. The posterior matching feedback scheme for joint source-channel coding with bandwidth expansion. In Storer and Marcellin [SM09], pages 83–92. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4976452. IEEE Computer Society

Order Number P3592. BMS Part Number CFP09DCC-PRT.

Sartipi:2011:EED

[SF11] M. Sartipi and R. Fletcher. Energy-efficient data acquisition in wireless sensor networks using compressed sensing. In Storer and Marcellin [SM11b], pages 223–232. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5749480. IEEE Computer Society

Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Saponara:2003:LPV

[SFT03] Sergio Saponara, Luca Fanucci, and Pierangelo Terren. Lowpower VLSI architectures for 3D discrete cosine transform (DCT). In Midwest Symposium on Circuits and Systems (MWSCAS), page ?? ????, ????, 2003.

Savari:1992:ACM

[SG92] S. A. Savari and R. G. Gallager. Arithmetic coding for memoryless cost channels. In Storer and Cohn [SC92], pages 92– 101. ISBN 0-8186-2717-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=227472. IEEE catalog

number 91TH0436-6.

Skibinski:2004:VLC

[SG04] P. Skibiński and S. Grabowski. Variable-length contexts for PPM. In Storer and Cohn [SC04], pages 409–418. ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1281486. IEEE Computer Society

Order Number: P2082.

Seraco:2009:AMC

[SG09] E. P. Seraco and J. G. R. C. Gomes. Affine modeling for the complexity of vector quantizers. In Storer and Marcellin [SM09], page 466. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ????

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=4976520. IEEE Computer Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

Sun:2011:SQR

[SG11] J. Z. Sun and V. K. Goyal. Scalar quantization for relative error. In Storer and Marcellin [SM11b], pages 293–302. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5749487. IEEE Computer Society

Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Seraco:2012:PVQ

[SG12] E. P. Seraco and J. G. R. C. Gomes. On the performance of vector quantization under constraint of complexity functionals. In Storer and Marcellin [SM12], page 409. ISBN 1-4673-0715-7 (paperback). ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=6189290. IEEE Computer

Society order number P4656.

Shkurti:2016:PPB

[SGA^{+}16] Ardita Shkurti, Ramon Goni, Pau Andrio, Elena Breitmoser, Iain Bethune, Modesto Orozco, and Charles A. Laughton. pyPcazip: a PCA-based toolkit for compression and analysis of molecular simulation data. SoftwareX, 5(??):37–43, ???? 2016. CODEN ???? ISSN

2352-7110. URL http://www.sciencedirect.com/science/
article/pii/S2352711016300036.

Skibinski:2005:RDB

[SGD05] Przemysław Skibiński, Szymon Grabowski, and Sebastian Deorowicz. Revisiting dictionary-based compression. Software — Practice and Experience, 35(15):1455–1476, December 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Soulodre:1998:SES

[SGLT98] G. A. Soulodre, T. Grusec, M. Lavoie, and L. Thibault. Subjective evaluation of state-of-the-art 2-channel audio codecs. Journal of the Audio Engineering Society, 46(3):164–176, March 1998. CODEN ADIOA3. ISSN 1549-4950.

Stauffer:1992:TCS

[SH92] L. M. Stauffer and D. S. Hirschberg. Transpose coding on the systolic array. In Storer and Cohn [SC92], pages 162– 171. ISBN 0-8186-2717-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=227465. IEEE catalog

number 91TH0436-6.

Saupe:1994:CRM

[SH94a] Dietmar Saupe and Raouf Hamzaoui. Complexity reduction methods for fractal image compression. In J. M. Blackledge, editor, Proceedings of the IMA Conference on Image Processing; Mathematical Methods and Applications. ????,

????, September 1994. URL ftp://ftp.informatik.uni-
freiburg.de/papers/fractal/SaHa95a.ps.gz.

Shamoon:1994:RAL

[SH94b] T. Shamoon and C. Heegard. A rapidly adaptive lossless compression algorithm for high fidelity audio coding. In Storer and Cohn [SC94], pages 430–439. ISBN 0-8186-5636-0, 0-8186-5637-9 (case). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1994. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=305950. IEEE

Catalog Number 93TH0626-2. IEEE Computer Society Press Order Number 5637-02.

Storer:1997:LIC

[SH97] James A. Storer and Harald Helfgott. Lossless image compression by block matching. The Computer Journal, 40(2– 3):137–145, ???? 1997. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/40/2_and_3/137.full.pdf+
html;   http://www.oup.co.uk/computer_journal/Volume_
40/Issue_02/Vol40_02.body.html#AbstractStorer;   http:
//www3.oup.co.uk/computer_journal/Volume_40/Issue_
02/Vol40_03.body.html#AbstractStorer.

Sermadevi:2003:LPO

[SH03] Y. Sermadevi and S. S. Hemami. Linear programming optimization for video coding under multiple constraints. In Storer and Cohn [SC03], pages 53–62. ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 2003. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1193996. IEEE Computer

Society Order number PR01896.

Sermadevi:2004:EBA

[SH04] Y. Sermadevi and S. S. Hemami. Efficient bit allocation for dependent video coding. In Storer and Cohn [SC04], pages 232–241. ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=1281468. IEEE

Computer Society Order Number: P2082.

Shannon:1948:MTCa

[Sha48a] Claude E. Shannon. A mathematical theory of communication. The Bell System Technical Journal, 27(3):379–423, July 1948. CODEN BSTJAN. ISSN 0005-8580. URL http://bstj.

bell-labs.com/BSTJ/images/Vol27/bstj27-3-379.pdf;
http://www.alcatel-lucent.com/bstj/vol27-1948/articles/

bstj27-3-379.pdf. Reprinted in [Sle74]. From the first page: “If the base 2 is used the resulting units may be called binary digits, or more briefly, bits, a word suggested by J. W. Tukey.”. This is the first known printed instance of the word ‘bit’ with the meaning of binary digit.

Shannon:1948:MTCb

[Sha48b] Claude E. Shannon. A mathematical theory of communication (continued). The Bell System Technical Journal, 27(4):623– 656, October 1948. CODEN BSTJAN. ISSN 0005-8580. URL

http://bstj.bell-labs.com/BSTJ/images/Vol27/bstj27-
4-623.pdf;   http://www.alcatel-lucent.com/bstj/vol27-
1948/articles/bstj27-4-623.pdf. Reprinted in [Sle74].

Shannon:1951:PEP

[Sha51] Claude E. Shannon. Prediction and entropy of printed English. The Bell System Technical Journal, 30(1):50– 64, January 1951. CODEN BSTJAN. ISSN 0005-

8580. URL http://bstj.bell-labs.com/BSTJ/images/
Vol30/bstj30-1-50.pdf;   http://www.alcatel-lucent.
com/bstj/vol30-1951/articles/bstj30-1-50.pdf. Reprinted

in [Sle74].

Shanmugam:1975:CDC

[Sha75] K. S. Shanmugam. Comments on “Discrete Cosine Transform”. IEEE Transactions on Computers, C-24(7):759, July 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956

(electronic). URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1672894. See [ANR74].

Shalaby:1991:DHT

[Sha91] Hossam M. H. Shalaby. Distributed hypothesis testing with data compression. Ph.D. thesis, University of Maryland, College Park, MD, USA, 1991. 136 pp. URL http://search.

proquest.com/docview/304016569.

Shapiro:1993:EIC

[Sha93a] J. Shapiro. Embedded image coding using zerotrees of wavelet coefficients. IEEE Transactions on Signal Processing, 41(12): 3445–3462, October 1993. CODEN ITPRED. ISSN 1053-587X (print), 1941-0476 (electronic).

Shapiro:1993:EHI

[Sha93b] J. M. Shapiro. An embedded hierarchical image coder using zerotrees of wavelet coefficients. In Storer and Cohn [SC93], pages 214–223. ISBN 0-8186-3391-3 (microfiche), 0-8186-3392-1 (casebound). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1993. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=253128. IEEE

Computer Society Press order number 3392-02. IEEE catalog number 93TH0536-3.

Shah:1994:EDC

[Sha94] Jaimin B. Shah. An evaluation of data compression algorithms using the communication network simulation. M.S.E. thesis, The University of Alabama in Huntsville, Huntsville, AL, USA,

1994. 111 pp. URL http://search.proquest.com/docview/
304135847.

Shamir:2004:SUL

[Sha04] G. I. Shamir. Sequential universal lossless techniques for compression of patterns and their description length. In Storer and Cohn [SC04], pages 419–428. ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1281487. IEEE Computer Society Order Number: P2082.

Shalinie:2006:MPF

[Sha06] S. Mercy Shalinie. Modeling parallel feed-forward based compression network. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 21(4):227–237, ???? 2006. CODEN ???? ISSN 1744-5760 (print), 1744-5779

(electronic). URL http://www.informaworld.com/smpp/
content~content=a747880320.

Shapira:2009:CTD

[Sha09] D. Shapira. Compressed transitive delta encoding. In Storer and Marcellin [SM09], pages 203–212. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=4976464. IEEE Computer Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

Shayevitz:2010:SDS

[Sha10] O. Shayevitz. A symbolic dynamical system approach to lossy source coding with feedforward. In Storer and Marcellin [SM10b], pages 317–326. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ????

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=
&arnumber=5453450.

Shayevitz:2011:GEC

[Sha11] O. Shayevitz. Graph entropy characterization of relay-assisted zero-error source coding with side information. In Storer and Marcellin [SM11b], pages 243–252. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=5749482. IEEE Computer Society Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Sodsong:2016:DPB

[SHC^{+}16] Wasuwee Sodsong, Jingun Hong, Seongwook Chung, Yeongkyu Lim, Shin-Dug Kim, and Bernd Burgstaller. Dynamic partitioning-based JPEG decompression on heterogeneous multicore architectures. Concurrency and Computation: Practice and Experience, 28(2):517–536, February 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Shekhar:2002:DDA

[SHD02] S. Shekhar, Yan Huang, and J. Djugash. Dictionary design algorithms for vector map compression. In Storer and Cohn [SC02], page ?? ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2002. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1000014. IEEE Computer Society Order Number PR01477.

Shepard:1990:DCM

[She90] Ron Shepard. A data compression method applicable to firstorder convergent iterative procedures. Journal of Computational Chemistry, 11(1):45–57, January 1990. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Sheinwald:1992:BAC

[She92] D. Sheinwald. On binary alphabetical codes. In Storer and Cohn [SC92], pages 112–121. ISBN 0-8186-2717-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ????

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=227470. IEEE catalog number 91TH0436-6.

Shenoi:1995:DSP

[She95] Kishan Shenoi. Digital Signal Processing in Telecommunications. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1995. ?? pp.

Shohdohji:1999:OQT

[SHK99] T. Shohdohji, Y. Hoshino, and N. Kutsuwada. Optimization of quantization table based on visual characteristics in DCT image coding. Computers and Mathematics with Applications, 37(11–12):225–232, June 1999. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0898122199001601.

Shkarin:2002:POS

[Shk02] D. Shkarin. PPM: one step to practicality. In Storer and Cohn [SC02], pages 202–211. ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2002. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=999958. IEEE Computer Society Order Number PR01477.

Shlien:1994:GMA

[Shl94] Seymour Shlien. Guide to MPEG-1 audio standard. IEEE Transactions on Broadcasting, 40(4):206–218, December 1994. CODEN IETBAC. ISSN 0018-9316 (print), 1557-9611 (electronic).

Shoa:2008:VLC

[Sho08] A. Shoa. Variable length coding for fixed rate, low latency, low complexity compression applications. In Storer and Marcellin [SM08], page 544. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008;

QA76.9.D33 INTERNET. URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=4483371.

Simpson:2001:DRT

[SHW^{+}01] Albert Simpson, Jill Hunter, Moira Wylie, Yi Hu, and David Mann. Demonstrating real-time JPEG image compressiondecompression using standard component IP cores on a programmable logic based platform for DSP and image processing. Lecture Notes in Computer Science, 2147:441–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer-ny.com/link/
service/series/0558/bibs/2147/21470441.htm;   http:
//link.springer-ny.com/link/service/series/0558/papers/
2147/21470441.pdf.

Shilane:2012:WOR

[SHWH12] Phlip Shilane, Mark Huang, Grant Wallace, and Windsor Hsu. WAN-optimized replication of backup datasets using streaminformed delta compression. ACM Transactions on Storage, 8 (4):13:1–13:??, November 2012. CODEN ???? ISSN 1553-3077 (print), 1553-3093 (electronic).

Sadakane:1999:CDT

[SI99] K. Sadakane and H. Imai. A cooperative distributed text database management method unifying search and compression based on the Burrows–Wheeler transformation. Lecture Notes in Computer Science, 1552:434–445, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sierpinski:1912:CQR

[Sie12] W. Sierpiński. Sur une courbe qui remplit toute une aire plane. (French) [On a curve that completely fills a planar area]. Bull. Acad. Sci. Cracovie, Serie A, pages 462–478, 1912. Reprinted in [Sie90].

Sieminski:1988:FDH

[Sie88] Andrzej Sieminski. Fast decoding of the Huffman codes. Information Processing Letters, 26(5):237–241, January 11, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sierpinski:1990:NCQ

[Sie90] W. Sierpiński. Sur une nouvelle courbe qui remplit toute une aire plaine,” Bull. Acad. Sci. Cracovie, Serie A:462–478. Simoncelli, Eero P., and Edward. H. Adelson (1990) “Subband Transforms”. In John Woods, editor, Subband Coding, pages 143–192. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1990. Reprint of [Sie12].

Seo:2011:CDM

[SILN11] Jaewoo Seo, Geoffrey Irving, J. P. Lewis, and Junyong Noh. Compression and direct manipulation of complex blendshape models. ACM Transactions on Graphics, 30(6):164:1–164:??, December 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Simpson:1992:RPP

[Sim92] W. Simpson. RFC 1331: The point-to-point protocol (PPP) for the transmission of multi-protocol datagrams over point-to-point links, May 1992. URL ftp://ftp.

internic.net/rfc/rfc1171.txt;   ftp://ftp.internic.
net/rfc/rfc1172.txt;   ftp://ftp.internic.net/rfc/
rfc1331.txt;   ftp://ftp.internic.net/rfc/rfc1548.
txt;   https://www.math.utah.edu/pub/rfc/rfc1171.txt;
 https://www.math.utah.edu/pub/rfc/rfc1172.txt;   https:
//www.math.utah.edu/pub/rfc/rfc1331.txt;   https://
www.math.utah.edu/pub/rfc/rfc1548.txt. Obsoleted by

RFC1548 [Sim93a]. Obsoletes RFC1171, RFC1172 [Per90, PH90]. Status: PROPOSED STANDARD.

Simpson:1993:RPP

[Sim93a] W. Simpson. RFC 1548: The point-to-point protocol (PPP), December 1993. URL ftp://ftp.internic.

net/rfc/rfc1331.txt;   ftp://ftp.internic.net/rfc/
rfc1548.txt;   ftp://ftp.internic.net/rfc/rfc1570.
txt;   ftp://ftp.internic.net/rfc/rfc1661.txt;   https:
//www.math.utah.edu/pub/rfc/rfc1331.txt;   https:/
/www.math.utah.edu/pub/rfc/rfc1548.txt;   https://
www.math.utah.edu/pub/rfc/rfc1570.txt;   https://www.
math.utah.edu/pub/rfc/rfc1661.txt. Obsoleted by RFC1661

[Sim94b]. Obsoletes RFC1331 [Sim92]. Updated by RFC1570 [Sim94a]. Status: DRAFT STANDARD.

Simpson:1993:RPH

[Sim93b] W. Simpson. RFC 1549: PPP in HDLC framing, Decem-

ber 1993. URL ftp://ftp.internic.net/rfc/rfc1549.txt;
 ftp://ftp.internic.net/rfc/rfc1662.txt; ftp://ftp.
internic.net/rfc/std51.txt;   https://www.math.utah.
edu/pub/rfc/rfc1549.txt;   https://www.math.utah.edu/
pub/rfc/rfc1662.txt;   https://www.math.utah.edu/pub/

rfc/std51.txt. Obsoleted by RFC1662, STD0051 [Sim94c, Sim94d]. Status: DRAFT STANDARD.

Simpson:1994:RPL

[Sim94a] W. Simpson. RFC 1570: PPP LCP extensions, January

1994. URL ftp://ftp.internic.net/rfc/rfc1548.txt;
 ftp://ftp.internic.net/rfc/rfc1570.txt;   https:/
/www.math.utah.edu/pub/rfc/rfc1548.txt;   https://
www.math.utah.edu/pub/rfc/rfc1570.txt. Updates RFC1548

[Sim93a]. Status: PROPOSED STANDARD.

Simpson:1994:RPP

[Sim94b] W. Simpson. RFC 1661: The point-to-point protocol

(PPP), July 1994. URL ftp://ftp.internic.net/rfc/
rfc1548.txt;   ftp://ftp.internic.net/rfc/rfc1661.
txt;   ftp://ftp.internic.net/rfc/rfc2153.txt; ftp:
//ftp.internic.net/rfc/std51.txt;   https://www.math.
utah.edu/pub/rfc/rfc1548.txt;   https://www.math.utah.
edu/pub/rfc/rfc1661.txt;   https://www.math.utah.edu/
pub/rfc/rfc2153.txt;   https://www.math.utah.edu/pub/

rfc/std51.txt. See also STD0051 [Sim94d]. Obsoletes RFC1548 [Sim93a]. Updated by RFC2153 [Sim97]. Status: STANDARD.

Simpson:1994:RPH

[Sim94c] W. Simpson. RFC 1662: PPP in HDLC-like fram-

ing, July 1994.  URL ftp://ftp.internic.net/rfc/
rfc1549.txt;   ftp://ftp.internic.net/rfc/rfc1662.
txt; ftp://ftp.internic.net/rfc/std51.txt;   https:
//www.math.utah.edu/pub/rfc/rfc1549.txt;   https:/
/www.math.utah.edu/pub/rfc/rfc1662.txt;   https://
www.math.utah.edu/pub/rfc/std51.txt. See also STD0051

[Sim94d]. Obsoletes RFC1549 [Sim93b]. Status: STANDARD.

Simpson:1994:SPP

[Sim94d] W. Simpson. STD 51: The Point-to-Point Protocol

(PPP), July 1994. URL ftp://ftp.isi.edu/in-notes/
rfc1549.txt;   ftp://ftp.isi.edu/in-notes/rfc1661.
txt;   ftp://ftp.isi.edu/in-notes/rfc1662.txt;   ftp:
//ftp.isi.edu/in-notes/std/std51.txt;   https://www.
math.utah.edu/pub/rfc/rfc1549.txt;   https://www.math.
utah.edu/pub/rfc/rfc1661.txt;   https://www.math.utah.
edu/pub/rfc/rfc1662.txt;   https://www.math.utah.edu/

pub/rfc/std/std51.txt. See also RFC1661, RFC1662 [Sim94b, Sim94c]. Obsoletes RFC1549 [Sim93b].

Simpson:1997:RPV

[Sim97] W. Simpson. RFC 2153: PPP vendor extensions, May 1997.

URL ftp://ftp.internic.net/rfc/rfc1661.txt;   ftp:
//ftp.internic.net/rfc/rfc1962.txt;   ftp://ftp.internic.
net/rfc/rfc2153.txt;   https://www.math.utah.edu/pub/
rfc/rfc1661.txt;   https://www.math.utah.edu/pub/rfc/
rfc1962.txt;   https://www.math.utah.edu/pub/rfc/rfc2153.

txt. Updates RFC1661, RFC1962 [Sim94b, Ran96]. Status: INFORMATIONAL.

Simon:1998:SMC

[Sim98] S. F. Simon. On suboptimal multidimensional companding. In Storer and Cohn [SC98], pages 438–447. ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1998. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=672191. IEEE Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Schieber:1994:RRT

[SJ94] Colleen D. Schieber and Eric E. Johnson. RATCHET: realtime address trace compression hardware for extended traces. ACM SIGMETRICS Performance Evaluation Review, 21(3– 4):22–32, April 1994. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Sodsong:2017:JPE

[SJPB17] Wasuwee Sodsong, Minyoung Jung, Jinwoo Park, and Bernd Burgstaller. JParEnt: Parallel entropy decoding for JPEG decompression on heterogeneous multicore architectures. Concurrency and Computation: Practice and Experience, 29(15), August 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sun:2008:MLR

[SjWL08] Chang Sun, Hong jun Wang, and Hua Li. Macroblock-level rate-distortion optimization with perceptual adjustment for video coding. In Storer and Marcellin [SM08], page 546. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33 INTER-

NET. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=4483373.

Srinivasan:1999:EWH

[SK99] S. Srinivasan and L. N. Kanal. Eigen wavelet: hyperspectral image compression algorithm. In Storer and Cohn [SC99], page ?? ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1999. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=785707. IEEE

Computer Society Order Number PR00096. IEEE Order Plan Catalog Number PR00096.

Schonberg:2004:ICC

[SK04] D. Schonberg and D. Kirovski. Iris compression for cryptographically secure person identification. In Storer and Cohn [SC04], pages 459–468. ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

1281491. IEEE Computer Society Order Number: P2082.

Singh:2009:IIC

[SK09] Satish Kumar Singh and Shishir Kumar. Improved image compression based on feed-forward adaptive downsampling algorithm. International Journal of Image and Graphics (IJIG), 9 (4):575–589, October 2009. CODEN ???? ISSN 0219-4678.

Savari:2010:WHM

[SK10a] S. A. Savari and J. Kliewer. When Huffman meets Hamming: a class of optimal variable-length error correcting codes. In Storer and Marcellin [SM10b], pages 327–336. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5453476.

Shapira:2010:BDF

[SK10b] D. Shapira and M. Kats. Bidirectional delta files. In Storer and Marcellin [SM10b], pages 249–258. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=5453445.

Skarbek:1998:PSI

[Ska98] Władysław Skarbek. Preface: Special issue on image compression. Fundamenta Informaticae, 34(4):??, July 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Sri-Krishna:1995:FBF

[SKCB95a] A. Sri-Krishna, Cheehung Chu, and M. Bayoumi. FFT based fast architecture & algorithm for discrete wavelet transforms. In Storer and Cohn [SC95], page 441. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=515550. IEEE catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Sri-Krishna:1995:PBF

[SKCB95b] A. Sri-Krishna, Cheehung Chu, and M. Bayoumi. FFT based fast architecture & algorithm for discrete wavelet transforms. In Storer and Cohn [SC95], page 441. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=515551. IEEE catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Shibata:2000:SPM

[SKF^{+}00] Yusuke Shibata, Takuya Kida, Shuichi Fukamachi, Masayuki Takeda, Ayumi Shinohara, Takeshi Shinohara, and Setsuo Arikawa. Speeding up pattern matching by text compression. Lecture Notes in Computer Science, 1767:306–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer-ny.com/link/
service/series/0558/bibs/1767/17670306.htm;   http:
//link.springer-ny.com/link/service/series/0558/papers/
1767/17670306.pdf.

Skibinski:2005:TLD

[Ski05] P. Skibiński. Two-level directory based compression. In Storer and Cohn [SC05], page ?? ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=1402238.

Skibinski:2008:IHC

[Ski08] P. Skibiński. Improving HTML compression. In Storer and Marcellin [SM08], page 545. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33 INTERNET. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=4483372.

Kuo:2001:EIC

[sKJ01] Shyh shiaw Kuo and J. D. Johnston. Enhancing image coders by using spatial noise shaping (SNS). In Storer and Cohn [SC01c], pages 321–330. ISBN 0-7695-1031-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2001. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=917163. IEEE CSP number 01PR1031.

Srisooksai:2012:PDC

[SKLA12] Tossaporn Srisooksai, Kamol Keamarungsi, Poonlap Lamsrichan, and Kiyomichi Araki. Practical data compression in wireless sensor networks: a survey. Journal of Network and Computer Applications, 35(1):37–59, January 2012. CO-DEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804511000555.

Sano:2010:SPP

[SKY10] K. Sano, K. Katahira, and S. Yamamoto. Segment-parallel predictor for FPGA-based hardware compressor and decompressor of floating-point data streams to enhance memory I/O bandwidth. In Storer and Marcellin [SM10b], pages 416–425. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=5453487.

Samarawickrama:2008:MCM

[SL08] U. Samarawickrama and Jie Liang. M-channel multiple description coding with two-rate predictive coding and staggered quantization. In Storer and Marcellin [SM08], page 542. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33

INTERNET. URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4483369.

Shamaie:2010:SQD

[SL10] S. Shamaie and F. Lahouti. Scalar quantizer design for noisy channel with decoder side information. In Storer and Marcellin [SM10b], pages 307–316. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=5453451.

Senecal:2005:ILI

[SLDJ05] J. G. Senecal, P. Lindstrom, M. A. Duchaineau, and K. I. Joy. Investigating lossy image coding using the PLHaar transform. In Storer and Cohn [SC05], page ?? ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=1402236.

Slepian:1974:KPD

[Sle74] David Slepian, editor. Key papers in the development of information theory. IEEE Press selected reprint series. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1974. ISBN 0-87942-027-8, 0-87942-028-6 (paperback). vi + 463 pp. LCCN Q360 .S54.

Sofia:2015:IHP

[SLJ^{+}15] A. T. Sofia, C. C. Lewis, C. Jacobi, D. A. Jamsek, D. F. Riedy, J. Vogt, P. Sutton, and R. W. St. John. Integrated highperformance data compression in the IBM z13. IBM Journal of Research and Development, 59(4–5):7:1–7:10, July/September 2015. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Selgrad:2017:CRR

[SLM^{+}17] Kai Selgrad, Alexander Lier, Magdalena Martinek, Christoph Buchenau, Michael Guthe, Franziska Kranz, Henry Schäfer, and Marc Stamminger. A compressed representation for ray tracing parametric surfaces. ACM Transactions on Graphics, 36(1):5:1–5:??, February 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sloane:2006:LEI

[Slo06] Neil J. A. Sloane. The on-line encyclopedia of integer sequences, 2006. URL http://oeis.org/.

Shavit:2005:MBA

[SLR05] G. Shavit, R. E. Ladner, and E. A. Riskin. MINMAX bit allocation for quantization-based video coders. In Storer and Cohn [SC05], pages 299–308. ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ????

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=
&arnumber=1402191.

Stankovic:2004:DSW

[SLXG04] V. Stankovic, A. D. Liveris, Zixiang Xiong, and C. N. Georghiades. Design of Slepian–Wolf codes by channel code partitioning. In Storer and Cohn [SC04], pages 302–311. ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1281475. IEEE Computer Society

Order Number: P2082.

Sheinwald:1991:CTW

[SLZ91] D. Sheinwald, A. Lempel, and J. Ziv. On compression with two-way head machines. In Storer and Reif [SR91b], pages 218–227. ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN QA76.9.D33 D37 1990. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=213359. IEEE

Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Sestak:2008:SAL

[SLZ08] R. Sestak, J. Lansky, and M. Zemlicka. Suffix array for large alphabet. In Storer and Marcellin [SM08], page 543. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33 INTER-

NET. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=4483370.

Sun:2019:CFP

[SLZ^{+}19] Xiuwen Sun, Hao Li, Dan Zhao, Xingxing Lu, Kaiyu Hou, and Chengchen Hu. COIN: a fast packet inspection method over compressed traffic. Journal of Network and Computer Applications, 127(??):122–134, February 1, 2019. CO-DEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804518303904.

Sherlock:1995:AFD

[SM95a] Barry G. Sherlock and Donald M. Monro. Algorithm 749: Fast discrete cosine transform. ACM Transactions on Mathematical Software, 21(4):372–378, December 1995. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295

(electronic). URL http://www.acm.org/pubs/citations/
journals/toms/1995-21-4/p372-sherlock/.

Sum:1995:GRB

[SM95b] Kwok Wing Sum and R. D. Murch. Generalized region based transform coding for video compression. In Storer and Cohn [SC95], page 478. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=515588. IEEE catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Sarma:1996:HRT

[SM96] D. Das Sarma and D. W. Matula. Hardware reciprocal table compression/decompression techniques. In Alefeld et al.

[AFL96], pages 11–17. ISBN 3-05-501737-4. ISSN 0138-3019. LCCN QA76.95 .I575 1995.

mme:1997:CFR

[SM97] Øyvind Strømme and Douglas R. McGregor. Comparison of fidelity of reproduction of images after lossy compression using standard and nonstandard wavelet decompositions. In Proceedings of The First European Conference on Signal Analysis and Prediction (ECSAP 97), Prague, Czechoslovakia, page ?? ????, ????, June 1997.

Stuiver:1998:PIM

[SM98] L. Stuiver and A. Moffat. Piecewise integer mapping for arithmetic coding. In Storer and Cohn [SC98], pages 3–12. ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1998. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=672121. IEEE Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Sellis:2001:PAS

[SM01a] Timos Sellis and Sharad Mehrotra, editors. Proceedings of the 2001 ACM SIGMOD International Conference on Management of Data 2001, Santa Barbara, California, United States, May 21–24, 2001. ACM Press, New York, NY 10036, USA, 2001. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN ???? ACM order number 472010.

Simard:2001:WCM

[SM01b] P. Y. Simard and H. S. Malvar. A wavelet coder for masked images. In Storer and Cohn [SC01c], pages 93–102. ISBN 0-7695-1031-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 2001. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=917140. IEEE CSP num-

ber 01PR1031.

Storer:2008:DDC

[SM08] James A. (James Andrew) Storer and Michael W. Marcellin, editors. DCC 2008: 2008 Data Compression Conference: March 25–27, 2008, Snowbird, Utah: proceedings. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2008. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33 INTER-

NET. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=4483271.

Storer:2009:DPD

[SM09] James A. (James Andrew) Storer and Michael W. Marcellin, editors. DCC 2009: Proceedings Data Compression Conference: Snowbird, Utah, USA, 16–28 March 2011. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2009. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=4976436. IEEE Computer Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

Salomon:2010:HDC

[SM10a] David Salomon and Giovanni Motta. Handbook of Data Compression. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., fifth edition, 2010. ISBN 1-84882-902-7 (hardcover), 1-84882-903-5. 300 pp. LCCN QA76.9.D33 S282 2010. With contributions by David Bryant.

Storer:2010:DPD

[SM10b] James A. (James Andrew) Storer and Michael W. Marcellin, editors. DDC 2010: proceedings: Data Compression Conference: 24–26 March 2010, Snowbird, Utah. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2010. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=5453521.

Sharafeddine:2011:LAC

[SM11a] Sanaa Sharafeddine and Rakan Maddah. A lightweight adaptive compression scheme for energy-efficient mobile-to-mobile file sharing applications. Journal of Network and Computer Applications, 34(1):52–61, January 2011. CO-DEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1084804510001682.

Storer:2011:DDC

[SM11b] James A. (James Andrew) Storer and Michael W. Marcellin, editors. DCC 2011: Data Compression Conference: Snowbird, Utah, USA, 29–31 March 2011. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2011. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/servlet/opac?punumber=5749456. IEEE Computer So-

ciety Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Storer:2012:DDC

[SM12] James A. Storer and Michael W. Marcellin, editors. DCC 2012: 2012 Data Compression Conference: proceedings, 10–12 April 2012, Snowbird, Utah. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2012. ISBN 1-4673-0715-7 (paperback). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL

http://ieeexplore.ieee.org/xpl/mostRecentIssue.jsp?

punumber=6189093. IEEE Computer Society order number P4656.

Senapati:2014:ILE

[SM14] Ranjan Kumar Senapati and Prasanth Mankar. Improved listless embedded block partitioning algorithms for image compression. International Journal of Image and Graphics (IJIG), 14(4):1450020, October 2014. CODEN ???? ISSN 0219-4678.

Starck:1998:IPD

[SMB98] J. L. Starck, F. Murtagh, and A. Bijaoui. Image Processing and Data Analysis: The Multiscale Approach. Cambridge University Press, Cambridge, UK, 1998. ?? pp.

Schelkens:1999:WBC

[SMC99] Peter Schelkens, Adrian Munteanu, and Jan Cornelis. Waveletbased compression of medical images: Protocols to improve resolution and quality scalability and region-of-interest coding. Future Generation Computer Systems, 15(2):171–184, March 11, 1999. CODEN FGSEVI. ISSN 0167-739X (print), 1872-

7115 (electronic). URL http://www.elsevier.com/gej-ng/
10/19/19/30/17/20/abstract.html.

Shayevitz:2007:BRC

[SMFZ07] O. Shayevitz, E. Meron, M. Feder, and R. Zamir. Bounds on redundancy in constrained delay arithmetic coding. In Storer and Cohn [SC07], pages 133–142. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=4148752. IEEE Computer Society Order Number P2791.

Smith:1979:SPA

[Smi79] George Heavener Smith. Spectral Peak Amplitude Coding for Voice Data Compression. Ph.D. thesis, Stanford University, Stanford, CA, USA, 1979. 101 pp. URL http://search.

proquest.com/docview/302972664.

Smith:1984:PFF

[Smi84] Alvy Ray Smith III. Plants, fractals, and formal languages. Computers and Graphics, 18(3):1–10, July 1984. CODEN CGRADI, CPGPBZ. ISSN 0097-8930 (print), 1558-4569 (electronic). Applications of formal grammars for generating graphic fractals — “graftals” — in the modeling of plants.

Smith:2009:PLS

[Smi09] Alvy Ray Smith. A pixel is not a little square, 2009. URL

ftp://ftp.alvyray.com/Acrobat/6_Pixel.pdf;   http:/
/www.cs.cmu.edu/afs/cs/academic/class/15869-f11/www/
readings/smith95_pixelsquare.pdf.

Seo:2015:DDF

[SMLS15] Bon-Keun Seo, Seungryoul Maeng, Joonwon Lee, and Euiseong Seo. DRACO: A deduplicating FTL for tangible extra capacity. IEEE Computer Architecture Letters, 14(2):123–126, July/ December 2015. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Shinya:1998:PMS

[SMO98] Mikio Shinya, Takeaki Mori, and Noriyoshi Osumi. Periodic motion synthesis and Fourier compression. Journal of Visualization and Computer Animation, 9(2):95–107, April–June 1998. CODEN JVCAEO. ISSN 1049-8907. URL http:

//www3.interscience.wiley.com/cgi-bin/abstract?ID=
5508;   http://www3.interscience.wiley.com/cgi-bin/
fulltext?ID=5508&PLACEBO=IE.pdf.

Satoh:1997:SJT

[SMOY97] N. Satoh, T. Morihara, Y. Okada, and S. Yoshida. Study of Japanese text compression. In Storer and Cohn [SC97], page ?? ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1997. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=582134. IEEE

Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Simard:2004:FBS

[SMRR04] P. Y. Simard, H. S. Malvar, J. Rinker, and E. Renshaw. A foreground-background separation algorithm for image compression. In Storer and Cohn [SC04], pages 498–507. ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1281495. IEEE Computer Society

Order Number: P2082.

Sriram:1991:VPC

[SMS91] Kotikalapudi Sriram, R. Scott McKinney, and Mostafa Hashem Sherif. Voice packetization and compression in broadband ATM networks. IEEE Journal on Selected Areas in Communications, 9(3):294–304, April 1991. CODEN ISACEM. ISSN 0733-8716 (print), 1558-0008 (electronic). IEEE Log Number 9042034.

Sairam:2008:NPP

[SMS08] Y. N. Sairam, Nan Ma, and N. Sinha. A novel partial prediction algorithm for fast 4 × 4 intra prediction mode decision in H.264/AVC. In Storer and Marcellin [SM08], pages 232–241. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33

INTERNET. URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4483301.

Schwarz:2007:OSV

[SMW07] Heiko Schwarz, Detlev Marpe, and Thomas Wiegand. Overview of the scalable video coding extension of the H.264/ AVC standard. IEEE Transactions on Circuits and Systems for Video Technology, 17(9):1103–1120, September 2007. CO-DEN ITCTEM. ISSN 1051-8215 (print), 1558-2205 (electronic).

Slyz:1994:NVB

[SN94] M. J. Slyz and D. L. Neuhoff. A nonlinear VQ-based predictive lossless image coder. In Storer and Cohn [SC94], pages 304–310. ISBN 0-8186-5636-0, 0-8186-5637-9 (case). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 1994. URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=305938. IEEE Cata-

log Number 93TH0626-2. IEEE Computer Society Press Order Number 5637-02.

Slyz:1996:PLT

[SN96a] M. J. Slyz and D. L. Neuhoff. Piecewise linear tree-structured models for lossless image compression. In Storer and Cohn [SC96], pages 260–269. ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1996. URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

488331. IEEE Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Strang:1996:WFB

[SN96b] Gilbert Strang and Truong Nguyen. Wavelets and Filter Banks. Wellesley-Cambridge Press, Wellesley, MA, USA, 1996. ISBN 0-9614088-7-1. xxi + 490 pp. LCCN TK7872.F5 S79 1996.

Snow:1976:ICT

[Sno76] Roderick Charles Snow. An investigation of clustering techniques, query methods and data base compression with respect to the enlist system. M.Sc.C.S., The University of New Brunswick, Fredericton, NB, Canada, 1976. ???? pp. URL

http://search.proquest.com/docview/302820785.

Saxena:2006:EQD

[SNR06] A. Saxena, J. Nayak, and K. Rose. On efficient quantizer design for robust distributed source coding. In Storer and Cohn [SC06], pages 63–72. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=
1607241.

SilvadeMoura:2000:FFW

[SNZBY00] Edleno Silva de Moura, Gonzalo Navarro, Nivio Ziviani, and Ricardo Baeza-Yates. Fast and flexible word searching on compressed text. ACM Transactions on Information Systems, 18(2):113–139, April 2000. CODEN ATISET. ISSN

1046-8188. URL http://www.acm.org/pubs/citations/
journals/tois/2000-18-2/p113-silva_de_moura/.

Sennhauser:1994:DIC

[SO94] Rene Sennhauser and Krystyna W. Ohnesorge. Document image compression using document analysis and block-class-specific data compression methods. Proceedings of SPIE — The International Society for Optical Engineering, 2186:146– 155, 1994. CODEN PSISDG. ISBN 0-8194-1481-6. ISSN 0277-786X (print), 1996-756X (electronic).

Srinivasamurthy:2002:RCQ

[SO02] N. Srinivasamurthy and A. Ortega. Reduced complexity quantization under classification constraints. In Storer and Cohn [SC02], pages 402–411. ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2002. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=999980. IEEE Computer Society Order Number PR01477.

SMPTE:2008:SHP

[Soc08a] Society of Motion Picture and Television Engineers (SMPTE). SMPTE home page, 2008. URL http://www.smpte.org/.

SMPTE-421M:2008:SFI

[Soc08b] Society of Motion Picture and Television Engineers (SMPTE). Standards facilitate interoperability, 2008. URL http://www.

smpte.org/standards.

Sofokleous:2005:RBM

[Sof05] Anastasis Sofokleous. Review: H.264 and MPEG-4 Video Compression: Video Coding for Next-generation Multimedia. The Computer Journal, 48(5):563, September 2005. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/cgi/
reprint/48/5/563.

Softsound:2003:U

[Sof07] Softsound. [unknown], 2003, 2007. URL http://mi.

eng.cam.ac.uk/reports/ajr/TR156/tr156.html;   http:
//www.softsound.com/Shorten.html. See also [Rob94a].

Sadakane:2000:ICT

[SOI00] K. Sadakane, T. Okazaki, and H. Imai. Implementing the context tree weighting method for text compression. In Storer and Cohn [SC00], pages 123–132. ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2000. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=838152. This millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Southard:1995:VQN

[Sou95] David Alan Southard. Vector quantization and nearest neighbor clustering with applications to image compression and data visualization. Sc.D. thesis, University of Massachusetts Lowell, Lowell, MA, USA, 1995. 169 pp. URL http://search.

proquest.com/docview/304200450.

Said:1996:NFE

[SP96] A. Said and W. A. Pearlman. A new fast and efficient image codec based on set partitioning in hierarchical trees. IEEE Transactions on Circuits and Systems for Video Technology, 6 (6):243–250, June 1996. CODEN ITCTEM. ISSN 1051-8215 (print), 1558-2205 (electronic).

Smith:1998:FCG

[SP98] T. C. Smith and R. Peeters. Fast convergence with a greedy tag-phrase dictionary. In Storer and Cohn [SC98], pages 33–42. ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1998. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=672128. IEEE

Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Schwartz:2011:IFI

[SP11] William Robson Schwartz and Helio Pedrini. Improved fractal image compression based on robust feature descriptors. International Journal of Image and Graphics (IJIG), 11(4):571– 587, October 2011. CODEN ???? ISSN 0219-4678.

Spaan:2000:ERC

[Spa00] Franciscus Henricus Petrus Spaan. Error resilient compression of digital video data. Doctoral, Technische Universiteit Delft, Delft, The Netherlands, 2000. 157 pp. URL http://search.

proquest.com/docview/304672016.

Spinellis:1993:IMAd

[Spi93] Diomidis Spinellis. Implementing multimedia applications: File storage and compression methods. Win magazine, pages

80–83, July 1993. URL http://kerkis.math.aegean.gr/~
dspin/pubs/trade/1993-Winmag/fastwin4/html/win4.html.

In Greek.

Seo:2020:EGS

[SPL20] Hojin Seo, Kisung Park, and Young-Koo Lee. An effective graph summarization and compression technique for a large-scaled graph. The Journal of Supercomputing, 76(10): 7906–7920, October 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https://link.

springer.com/article/10.1007/s11227-018-2245-5.

Skarbek:1998:MOR

[SPS98] Władysław Skarbek, Adam Pietrowcew, and Radosław Sikora. The modified oja-RLS algorithm, stochastic convergence analysis and application for image compression. Fundamenta Informaticae, 36(4):345–365, December 1998. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Skarbek:1999:MOR

[SPS99] W. Skarbek, A. Pietrowcew, and R. Sikora. Modified Oja-RLS algorithm — stochastic convergence analysis and application for image compression. Lecture Notes in Computer Science, 1609:402–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Saleh:2016:PAN

[SQ16] Bassel Saleh and Dongyu Qiu. Performance analysis of network-coding-based P2P live streaming systems. IEEE/ ACM Transactions on Networking, 24(4):2140–2153, August 2016. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Storer:1991:PAH

[SR91a] James A. Storer and John H. Reif. A parallel architecture for high-speed data compression. Journal of Parallel and Distributed Computing, 13(2):222–227, October 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Storer:1991:DDC

[SR91b] James A. (James Andrew) Storer and J. H. (John H.) Reif, editors. DCC ’91: Data Compression Conference: April 8– 11, 1991, Cliff Lodge, Snowbird, Utah. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1991. ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1990. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=213364. IEEE

Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Smith:1993:AMC

[SR93] Brian C. Smith and Lawrence A. Rowe. Algorithms for manipulating compressed images. IEEE Computer Graphics and Applications, 13(5):34–42, September 1993. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Storer:1997:LCP

[SR97a] J. A. Storer and J. Reif. Low-cost prevention of errorpropagation for data compression with dynamic dictionaries. In Storer and Cohn [SC97], pages 171–180. ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1997. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=582007. IEEE Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Storer:1997:ERO

[SR97b] James A. Storer and John H. Reif. Error-resilient optimal data compression. SIAM Journal on Computing, 26(4):934– 949, August 1997. CODEN SMJCAT. ISSN 0097-5397 (print),

1095-7111 (electronic). URL http://epubs.siam.org/sam-
bin/dbq/article/24078.

Srinivasan:2005:CET

[SR05] Sridhar Srinivasan and Shankar Regunathan. Computationally efficient transforms for video coding. In IEEE [IEE05], pages 325–328. ISBN 0-7803-9134-9. LCCN ???? Five volumes. IEEE catalog number 05CH37687.

Saxena:2008:DMS

[SR08] A. Saxena and K. Rose. Distributed multi-stage coding of correlated sources. In Storer and Marcellin [SM08], pages 312–321. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33

INTERNET. URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4483309.

Steger:2009:URT

[SR09] S. Steger and T. Richter. Universal refinable Trellis coded quantization. In Storer and Marcellin [SM09], pages 312–321. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=4976475. IEEE Computer

Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

Sowmyayani:2016:ETR

[SR16] S. Sowmyayani and P. Arockia Jansi Rani. An efficient temporal redundancy transformation for wavelet based video compression. International Journal of Image and Graphics (IJIG), 16(3):1650015, July 2016. CODEN ???? ISSN 0219-4678.

Stein:2021:LAA

[SRG^{+}21] Charles M. Stein, Dinei A. Rockenbach, Dalvan Griebler, Massimo Torquati, Gabriele Mencagli, Marco Danelutto, and Luiz G. Fernandes. Latency-aware adaptive micro-batching techniques for streamed data compression on graphics processing units. Concurrency and Computation: Practice and Experience, 33(11):e5786:1–e5786:??, June 10, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Srinivasan:1999:IDM

[Sri99] M. Srinivasan. Iterative decoding of multiple descriptions. In Storer and Cohn [SC99], pages 463–472. ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1999. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=755696. IEEE Computer Society Or-

der Number PR00096. IEEE Order Plan Catalog Number PR00096.

Scales:1995:LCU

[SRKS95] A. Scales, W. Roark, F. Kossentini, and M. J. T. Smith. Lossless compression using conditional entropy-constrained subband quantization. In Storer and Cohn [SC95], page ?? ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ????  URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=515603. IEEE catalog

number 95TH8037. IEEE Computer Society Press order number PR07010.

Schonberg:2004:DCC

[SRP04] D. Schonberg, K. Ramchandran, and S. S. Pradhan. Distributed code constructions for the entire Slepian–Wolf rate region for arbitrarily correlated sources. In Storer and Cohn [SC04], pages 292–301. ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

1281474. IEEE Computer Society Order Number: P2082.

Shavit:2004:GTV

[SRW^{+}04] G. Shavit, M. F. Ringenburg, J. West, R. E. Ladner, and E. A. Riskin. Group testing for video compression. In Storer and Cohn [SC04], pages 212–221. ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1281466. IEEE Computer Society Order Number: P2082.

Storer:1978:MMD

[SS78] James A. Storer and Thomas G. Szymanski. The macro model for data compression (extended abstract). In ACM [ACM78], pages 30–39. LCCN QA76.6.A13 1978.

Storer:1982:DCT

[SS82] James A. Storer and Thomas G. Szymanski. Data compression via textual substitution. Journal of the ACM, 29(4):928–951, October 1982. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Sproull:1992:CCG

[SS92] R. F. Sproull and I. E. Sutherland. A comparison of codebook generation techniques for vector quantization. In Storer and Cohn [SC92], pages 122–131. ISBN 0-8186-2717-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=227469. IEEE catalog number 91TH0436-6.

Sweldens:1996:BYO

[SS96] Wim Sweldens and Peter Schröder. Building your own wavelets at home. SIGGRAPH 96 Course Notes. Available on the WWW., 1996.

Salih:1999:ETR

[SS99] I. Salih and S. H. Smith. Encoding time reduction in fractal image compression. In Storer and Cohn [SC99], page ?? ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1999. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=785706. IEEE Computer Society Or-

der Number PR00096. IEEE Order Plan Catalog Number PR00096.

Shlepakov:2000:ZMA

[SS00] D. V. Shlepakov and L. N. Shlepakov. On zonal morphological approach to natural language texts processing. In Storer and Cohn [SC00], page ?? ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2000. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=838218. This

millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Shapira:2003:PDF

[SS03] D. Shapira and J. A. Storer. In-place differential file compression. In Storer and Cohn [SC03], pages 263–272. ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 2003. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1194017. IEEE Computer

Society Order number PR01896.

Shapira:2004:PDF

[SS04] D. Shapira and J. A. Storer. In-place differential file compression of nonaligned files with applications to file distribution, backups, and string similarity. In Storer and Cohn [SC04], pages 82–91. ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

1281453. IEEE Computer Society Order Number: P2082.

Seidel:2005:TAP

[SS05a] Raimund Seidel and Micha Sharir. Top-down analysis of path compression. SIAM Journal on Computing, 34(3):515–525, June 2005. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-

7111 (electronic). URL http://epubs.siam.org/sam-bin/
dbq/article/43908.

Shapira:2005:PDF

[SS05b] Dana Shapira and James A. Storer. In place differential file compression. The Computer Journal, 48(6):677– 691, November 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/cgi/content/abstract/48/6/677;
http://comjnl.oxfordjournals.org/cgi/content/full/
48/6/677;   http://comjnl.oxfordjournals.org/cgi/reprint/
48/6/677.

Shoa:2006:DMP

[SS06] A. Shoa and S. Shirani. Distortion of matching pursuit: modeling and optimization. In Storer and Cohn [SC06], page 466. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1607309.

Segall:2007:SSW

[SS07] C. Andrew Segall and Gary J. Sullivan. Spatial scalability within the H.264/AVC scalable video coding extension. IEEE Transactions on Circuits and Systems for Video Technology, 17(9):1121–1135, September 2007. CODEN ITCTEM. ISSN 1051-8215 (print), 1558-2205 (electronic).

Simard:2002:LAI

[SSM02] P. Simard, D. Steinkraus, and H. Malvar. On-line adaptation in image coding with a 2-D tarp filter. In Storer and Cohn [SC02], pages 23–32. ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2002. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=999940. IEEE Computer Society Order Number PR01477.

Schmiedekamp:2006:CCS

[SSP06] Mendel Schmiedekamp, Aparna Subbu, and Shashi Phoha. The clustered causal state algorithm: Efficient pattern discovery for lossy data-compression applications. Computing in Science and Engineering, 8(5):59–67, September/October 2006. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Sharma:2021:RIB

[SSP21] Urvashi Sharma, Meenakshi Sood, and Emjee Puthooran. Region of interest-based coding technique of medical images using varying grading of compression. International Journal of Image and Graphics (IJIG), 21(02):??, April 2021. ISSN

0219-4678. URL https://www.worldscientific.com/doi/
10.1142/S0219467821500212.

Singh:2009:OCS

[SSRM09] J. Singh, A. Saxena, K. Rose, and U. Madhow. Optimization of correlated source coding for event-based monitoring in sensor networks. In Storer and Marcellin [SM09], pages 73–82. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=4976451. IEEE Computer

Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

Salami:1998:LCH

[SST^{+}98] M. Salami, H. Sakanashi, M. Tanaka, M. Iwata, T. Kurita, and T. Higuchi. On-line compression of high precision printer images by evolvable hardware. In Storer and Cohn [SC98], pages 219–228. ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1998. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=672150. IEEE

Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Singla:2012:APS

[SST^{+}12] A. Singla, J. Shukla, A. K. Tiwari, S. P. Jaiswal, and V. Jakhetiya. Adaptive predictor structures for lossless compression of videos. In Storer and Marcellin [SM12], page 410. ISBN 1-4673-0715-7 (paperback). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

6189291. IEEE Computer Society order number P4656.

Sardashti:2016:YAC

[SSW16] Somayeh Sardashti, Andre Seznec, and David A. Wood. Yet another compressed cache: a low-cost yet effective compressed cache. ACM Transactions on Architecture and Code Optimization, 13(3):27:1–27:??, September 2016. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Song:2014:PNF

[SSZZ14] Renchu Song, Weiwei Sun, Baihua Zheng, and Yu Zheng. PRESS: a novel framework of trajectory compression in road networks. Proceedings of the VLDB Endowment, 7(9):661–672, May 2014. CODEN ???? ISSN 2150-8097.

Seemann:1997:GLA

[ST97] T. Seemann and P. Tischer. Generalised locally adaptive DPCM. In Storer and Cohn [SC97], page ?? ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1997. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=582142. IEEE Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Stanley:1967:CMD

[Sta67] Thomas Patrick Stanley. Compression of Markov Data. Ph.D. thesis, Princeton University, Princeton, NJ, USA,

1967. 106 pp. URL http://search.proquest.com/docview/
302264732.

Stauffer:1994:PHS

[Sta94] Lynn Marie Stauffer. Parallel and high-speed data compression. Ph.D. thesis, University of California, Irvine, Irvine, CA, USA,

1994. 113 pp. URL http://search.proquest.com/docview/
304084505.

Starosolski:2007:SFA

[Sta07] Roman Starosolski. Simple fast and adaptive lossless image compression algorithm. Software — Practice and Experience, 37(1):65–91, January 2007. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Stamplecoskie:2009:DCB

[Sta09] Jennifer Krista Stamplecoskie. Data compression of Blainville’s beaked whale vocalizations = La compression de vocalizations de baleine à bec de Blainville. Master of Applied Science in Signal Processing, Division of Graduate Studies, Royal Military College of Canada, St. Jean, QC, Canada, April 2009. ix

+ 79 pp. URL http://www.collectionscanada.gc.ca/obj/
thesescanada/vol2/002/MR52296.PDF.

Stewart:1981:TDC

[Ste81] Lawrence Colm Stewart. Trellis Data Compression. Ph.D. thesis, Stanford University, Stanford, CA, USA, 1981. 137 pp.

URL http://search.proquest.com/docview/303177936.

Stearns:1991:WDC

[Ste91] S. D. Stearns. Waveform data compression with exact recovery. In Storer and Reif [SR91b], page ?? ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1990. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=213297. IEEE Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Stevenson:2006:PPP

[Ste06] Daniel E. Stevenson. PNG palette permuter. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):143–147, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sterne:2012:MMF

[Ste12] Jonathan Sterne. MP3: the meaning of a format. Duke University Press, Durham, NC, USA, 2012. ISBN 0-8223-5287-7, 0-8223-5283-4. 360 pp. LCCN ????

Stout:1980:IPE

[Sto80] Quentin F. Stout. Improved prefix encodings of the natural numbers. IEEE Transactions on Information Theory, 26(5): 607–609, September 1980. CODEN IETTAW. ISSN 0018-9448 (print), 1557-9654 (electronic).

Storer:1983:TAT

[Sto83] J. A. Storer. Toward an abstract theory of data compression. Theoretical Computer Science, 24(3):221–237, August 1983. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Storer:1988:DCM

[Sto88] James A. (James Andrew) Storer. Data Compression: Methods and Theory. Principles of computer science series. Computer Science Press, Inc., 11 Taft Court, Rockville, MD 20850, USA, 1988. ISBN 0-88175-161-8. x + 413 pp. LCCN QA76.9.D33 S76 1988.

Storer:1996:LIC

[Sto96] J. A. Storer. Lossless image compression using generalized LZ1-type methods. In Storer and Cohn [SC96], pages 290–299. ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1996. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=488334. IEEE

Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Storer:1998:PEP

[Sto98] J. A. Storer. The prevention of error propagation in dictionary compression with update and deletion. In Storer and Cohn [SC98], pages 199–208. ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1998.

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=672148. IEEE Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Strang:1999:DCT

[Str99a] Gilbert Strang. The discrete cosine transform. SIAM Review, 41(1):135–147, March 1999. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://epubs.

siam.org/sam-bin/dbq/article/33674.

Strohbeck:1999:NAI

[Str99b] Uwe Strohbeck. A new approach in image data compression by multiple-resolution frame processing. Ph.D. thesis, University of Northumbria at Newcastle, Newcastle, UK, 1999. 140 pp.

URL http://search.proquest.com/docview/304587067.

mme:1999:AWT

[Str99c] Øyvind Strømme. On The Applicability of Wavelet Transforms to Image and Video Compression. Ph.D. thesis, University of Strathclyde, ????, February 1999. ?? pp.

Stubley:1994:AVV

[Stu94a] P. R. Stubley. Adaptive variable-to-variable length codes. In Storer and Cohn [SC94], pages 98–105. ISBN 0-8186-5636-0, 0-8186-5637-9 (case). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1994. URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

305917. IEEE Catalog Number 93TH0626-2. IEEE Computer Society Press Order Number 5637-02.

Stubley:1994:ROF

[Stu94b] P. R. Stubley. On the redundancy of optimum fixed-to-variable length codes. In Storer and Cohn [SC94], pages 90– 97. ISBN 0-8186-5636-0, 0-8186-5637-9 (case). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1994.

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=305916. IEEE Catalog Number 93TH0626-2. IEEE Computer Society Press Order Number 5637-02.

Subrahmanya:1999:AOD

[Sub99] P. Subrahmanya. An asymptotically optimal data compression algorithm based on an inverted index. In Storer and Cohn [SC99], page ?? ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1999. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=785708. IEEE

Computer Society Order Number PR00096. IEEE Order Plan Catalog Number PR00096.

Sun:2016:NEB

[Sun16] Shuliang Sun. A novel edge based image steganography with 2^{k} correction and Huffman encoding. Information Processing Letters, 116(2):93–99, February 2016. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0020019015001738.

Suzuki:1996:CSN

[Suz96] J. Suzuki. A CTW scheme for non-tree sources. In Storer and Cohn [SC96], page ?? ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1996. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=488386. IEEE

Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

suzannevega:2006:U

[suz06] suzannevega. [unknown], 2006. URL http://www.suzannevega.

com/suzanne/funfacts/music.aspx.

Suzuki:2011:UMG

[Suz11] J. Suzuki. The universal measure for general sources and its application to MDL/Bayesian criteria. In Storer and Marcellin [SM11b], page 478. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ????

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=5749535. IEEE Computer Society Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Suzuki:2012:BNS

[Suz12] J. Suzuki. Bayesian network structure estimation based on the Bayesian/MDL criteria when both discrete and continuous variables are present. In Storer and Marcellin [SM12], pages 307–316. ISBN 1-4673-0715-7 (paperback). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ????

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=6189262. IEEE Computer Society order number P4656.

Schneider:1995:DSC

[SV95] K. W. Schneider and W. S. Venters. The development of a standard for compression of synchronous data in DSU/CSU’s. In Storer and Cohn [SC95], page ?? ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=515598. IEEE catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Subbalakshmi:1998:ODE

[SV98] K. P. Subbalakshmi and J. Vaisey. Optimal decoding of entropy coded memoryless sources over binary symmetric channels. In Storer and Cohn [SC98], page ?? ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1998. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=672315. IEEE Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Saha:2000:EIA

[SV00] S. Saha and R. Vemuri. Effect of image activity on lossy and lossless coding performance. In Storer and Cohn [SC00], page ?? ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 2000. URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=838217. This millen-

nial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Soundararajan:2009:CDC

[SV09] R. Soundararajan and S. Vishwanath. Communicating the difference of correlated Gaussian sources over a MAC. In Storer and Marcellin [SM09], pages 282–291. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=4976472. IEEE Computer Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

Servetto:1999:MDL

[SVS99] S. D. Servetto, V. A. Vaishampayan, and N. J. A. Sloane. Multiple description lattice vector quantization. In Storer and Cohn [SC99], pages 13–22. ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1999. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=755649. IEEE

Computer Society Order Number PR00096. IEEE Order Plan Catalog Number PR00096.

Salari:1991:CSI

[SW91] E. Salari and W. A. Whyte, Jr. Compression of stereoscopic image data. In Storer and Reif [SR91b], page ?? ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1990. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=213336. IEEE Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Seroussi:1997:ASE

[SW97] G. Seroussi and M. J. Weinberger. On adaptive strategies for an extended family of Golomb-type codes. In Storer and Cohn [SC97], pages 131–140. ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1997. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=581993. IEEE Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Sarshar:2004:MMS

[SW04a] N. Sarshar and X. Wu. Minimax multiresolution scalar quantization. In Storer and Cohn [SC04], pages 52–61. ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1281450. IEEE Computer Society

Order Number: P2082.

Sarshar:2004:WCS

[SW04b] N. Sarshar and X. Wu. On wavelet compression of self-similar processes. In Storer and Cohn [SC04], page ?? ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1281539. IEEE Computer Society

Order Number: P2082.

Sarshar:2006:PAJ

[SW06] N. Sarshar and Xiaolin Wu. A practical approach to joint network-source coding. In Storer and Cohn [SC06], pages 93– 102. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=1607244.

Sasaki:2012:DBT

[SW12a] Yu Sasaki and Lei Wang. Distinguishers beyond three rounds of the RIPEMD-128/-160 compression functions. Lecture Notes in Computer Science, 7341:275–292, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/chapter/10.1007/978-
3-642-31284-7_17/.

Shcherbakov:2012:PAR

[SW12b] I. Shcherbakov and N. Wehn. A parallel adaptive range coding compressor: Algorithm, FPGA prototype, evaluation. In Storer and Marcellin [SM12], pages 119–128. ISBN 1-4673-0715-7 (paperback). ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=6189243. IEEE Computer

Society order number P4656.

Sardashti:2017:CCG

[SW17] Somayeh Sardashti and David A. Wood. Could compression be of general use? Evaluating memory compression across domains. ACM Transactions on Architecture and Code Optimization, 14(4):44:1–44:??, December 2017. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Swan:1993:IWF

[Swa93] Tom Swan. Inside Windows File Formats. SAMS Publishing, Indianapolis, IN, USA, 1993. ?? pp.

Solomon:1994:VDB

[SWA94] J. A. Solomon, A. B. Watson, and A. Ahumada. Visibility of DCT basis functions: effects of contrast masking. In Storer and Cohn [SC94], pages 361–370. ISBN 0-8186-5636-0, 0-8186-5637-9 (case). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1994. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=305944. IEEE

Catalog Number 93TH0626-2. IEEE Computer Society Press Order Number 5637-02.

Swacha:2004:PSC

[Swa04a] J. Swacha. The predictive-substitutional compression scheme and its effective implementation. In Storer and Cohn [SC04], page ?? ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=1281542. IEEE

Computer Society Order Number: P2082.

Swacha:2004:PSM

[Swa04b] Jakub Swacha. Predykcyjno-substytucyjna metoda bezstratnej kompresji danych. (Polish) [A predictive-substitutional lossless data compression method]. Ph.D. thesis, Politechnika Wrocławska, Wrocław, Poland, 2004. 146 pp. URL http:/

/search.proquest.com/docview/305039563.

Swacha:2008:CST

[Swa08] J. Swacha. CoTe: a software tool for compression benchmarking. In Storer and Marcellin [SM08], page 547. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33 INTER-

NET. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=4483374.

Song:2009:BAP

[SWGZ09] Chuan-Ming Song, Xiang-Hai Wang, Yanwen Guo, and Fu-Yan Zhang. Binary alpha-plane assisted fast motion estimation of video objects in wavelet domain. In Storer and Marcellin [SM09], page 467. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ????

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=4976521. IEEE Computer Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

Shen:2012:PSI

[SWW12] Yun-Chung Shen, Pin-Shiang Wang, and Ja-Ling Wu. Progressive side information refinement with non-local means based denoising process for Wyner–Ziv video coding. In Storer and Marcellin [SM12], pages 219–226. ISBN 1-4673-0715-7 (paperback). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=6189253. IEEE Computer Society order number P4656.

Song:2015:VRF

[SWWW15] Ying Song, Jiaping Wang, Li-Yi Wei, and Wencheng Wang. Vector regression functions for texture compression. ACM Transactions on Graphics, 35(1):5:1–5:??, December 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Stankovic:2004:PEP

[SX04] V. Stankovic and Zixiang Xiong. Packet erasure protection for multicasting. In Storer and Cohn [SC04], pages 142–151. ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1281459. IEEE Computer

Society Order Number: P2082.

Suel:2001:CGS

[SY01] T. Suel and Jun Yuan. Compressing the graph structure of the Web. In Storer and Cohn [SC01c], pages 213–222. ISBN 0-7695-1031-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2001. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=917152. IEEE

CSP number 01PR1031.

Sun:2004:OBC

[SY04] Fangting Sun and Wei Yu. Optimal buffer-constrained bit allocation using marginal analysis. In Storer and Cohn [SC04], page ?? ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=1281541. IEEE

Computer Society Order Number: P2082.

Sui:2018:PPC

[SY18] Peipei Sui and Xiaoyu Yang. A privacy-preserving compression storage method for large trajectory data in road network. Journal of Grid Computing, 16(2):229–245, June 2018. CODEN ???? ISSN 1570-7873 (print), 1572-9184

(electronic). URL https://link.springer.com/article/
10.1007/s10723-018-9435-5.

Schonberg:2007:CEV

[SYDR07] D. Schonberg, Chuohao Yeo, S. C. Draper, and K. Ramchandran. On compression of encrypted video. In Storer and Cohn [SC07], pages 173–182. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ????

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=4148756. IEEE Computer Society Order Number P2791.

Symes:2003:MD

[Sym03] Peter D. Symes. MPEG-4 Demystified. McGraw-Hill Professional, New York, NY, USA, 2003. ?? pp.

Scheuermann:1994:CBI

[SYO94] P. Scheuermann, A. Yaagoub, and M. Aris Ouksel. Compression of binary images on a hypercube machine. Journal of Parallel and Distributed Computing, 23(1):49–59, October 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.idealibrary.

com/links/doi/10.1006/jpdc.1994.1118/production;
http://www.idealibrary.com/links/doi/10.1006/jpdc.
1994.1118/production/pdf.

Schaeffer:2015:STR

[SYO15] Hayden Schaeffer, Yi Yang, and Stanley Osher. Space-time regularization for video decompression. SIAM Journal on Imaging Sciences, 8(1):373–402, ???? 2015. CODEN SJISBI. ISSN 1936-4954.

Said:2004:CFA

[SYP04] A. Said, S. Yea, and W. A. Pearlman. Coding for fast access to image regions defined by pixel range. In Storer and Cohn [SC04], pages 489–497. ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ????

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=1281494. IEEE Computer Society Order Number: P2082.

Schaeffer:2015:RTA

[SYZO15] Hayden Schaeffer, Yi Yang, Hongkai Zhao, and Stanley Osher. Real-time adaptive video compression. SIAM Journal on Scientific Computing, 37(6):B980–B1001, ???? 2015. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

Sherwood:1997:PIC

[SZ97] P. G. Sherwood and K. Zeger. Progressive image coding on noisy channels. In Storer and Cohn [SC97], pages 72–81. ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1997. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=581971. IEEE Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Slyz:2005:BBI

[SZ05] M. Slyz and Lei Zhang. A block-based inter-band lossless hyperspectral image compressor. In Storer and Cohn [SC05], pages 427–436. ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=
1402204.

Shen:2013:MSA

[SZ13] Xiaoping Shen and Ahmed I. Zayed, editors. Multiscale Signal Analysis and Modeling. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 1-4614-4144-7, 1-4614-4145-5 (e-book). xvii + 378 pp. LCCN TA342 .M855 2013.

Sun:2003:DBM

[SZM03] Welfeng Sun, Nan Zhang, and A. Mukherjee. A dictionarybased multi-corpora text compression system. In Storer and Cohn [SC03], page ?? ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2003. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1194067. IEEE Computer Society Order number PR01896.

Szpankowski:1991:TBS

[Szp91] W. Szpankowski. A typical behavior of some data compression schemes. In Storer and Reif [SR91b], pages 247–256. ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1990. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=213356. IEEE Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Szymczak:2002:OEE

[Szy02] A. Szymczak. Optimized edgebreaker encoding for large and regular triangle meshes. In Storer and Cohn [SC02], page ?? ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2002. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=1000015. IEEE

Computer Society Order Number PR01477.

Tadayon:1998:ALD

[Tad98] Nasser Tadayon. Adaptive lossless data compression algorithms with new approaches and modeling techniques. Ph.D. thesis, University of Southwestern Louisiana, Lafayette, LA, USA,

1998. 122 pp. URL http://search.proquest.com/docview/
304454825.

Taddei:2020:RMM

[Tad20] Tommaso Taddei. A registration method for model order reduction: Data compression and geometry reduction. SIAM Journal on Scientific Computing, 42(2):A997–A1027, ???? 2020. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

Tai:1995:EMS

[Tai95] Shen-Chuan Tai. An extensive Markov system for ECG exact coding. In Storer and Cohn [SC95], page 467. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=515577. IEEE catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Tan:1993:DCA

[Tan93] Roland K. D. Tan. Data Compression Algorithm Using Multi-Pulse Adaptive Sub-Band Coding (MASC). Ph.D. thesis, University of Essex, Southend-on-Sea, Essex, UK, 1993. ???? pp.

URL http://search.proquest.com/docview/304081512.

Tanenbaum:2002:CN

[Tan02] Andrew S. Tanenbaum. Computer Networks. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2002. ?? pp.

Tani:2008:PHP

[Tan08] Giorgio Tani. PeaZip home page, 2008. URL http://

sourceforge.net/projects/peazip/.

Tao:1982:RTS

[Tao82] Bertram P. M. Tao. Real-time speech data compression system based upon vector quantization. M.S. thesis, California State University, Long Beach, Long Beach, CA, USA, 1982. 189 pp.

URL http://search.proquest.com/docview/303109401.

Tarjan:1979:APC

[Tar79] Robert Endre Tarjan. Applications of path compression on balanced trees. Journal of the ACM, 26(4):690–715, October 1979. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Tarhio:1995:CCP

[Tar95] J. Tarhio. Context coding of parse trees. In Storer and Cohn [SC95], page 442. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

515552. IEEE catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Tate:1994:BOL

[Tat94] S. R. Tate. Band ordering in lossless compression of multispectral images. In Storer and Cohn [SC94], pages 311–320. ISBN 0-8186-5636-0, 0-8186-5637-9 (case). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1994.

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=305939. IEEE Catalog Number 93TH0626-2. IEEE Computer Society Press Order Number 5637-02.

Taubman:2000:HPS

[Tau00] David Taubman. High performance scalable image compression with EBCOT. IEEE Transactions on Image Processing, 9(7):1158–1170, July 2000. CODEN IIPRE4. ISSN 1057-7149 (print), 1941-0042 (electronic).

Tawel:1993:RTF

[Taw93] R. Tawel. Real-time focal-plane image compression. In Storer and Cohn [SC93], pages 401–409. ISBN 0-8186-3391-3 (microfiche), 0-8186-3392-1 (casebound). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1993. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=253109. IEEE Computer Society Press order number 3392-02. IEEE catalog number 93TH0536-3.

Tokdemir:2011:PPD

[TB11] S. Tokdemir and S. Belkasim. Parallel processing of DCT on GPU. In Storer and Marcellin [SM11b], page 479. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5749536. IEEE Computer Society

Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Tabesh:2005:JMJ

[TBKM05] A. Tabesh, A. Bilgin, K. Krishnan, and M. W. Marcellin. JPEG2000 and Motion JPEG2000 content analysis using codestream length information. In Storer and Cohn [SC05], pages 329–337. ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=1402194.

Tam:1998:BSR

[TBMM98] D. Tam, W. Barrett, B. Morse, and E. Mortensen. Breakpoint skeletal representation and compression of document images. In Storer and Cohn [SC98], page ?? ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1998. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=672317. IEEE Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Teahan:1996:EEU

[TC96] W. J. Teahan and J. G. Cleary. The entropy of English using PPM-based models. In Storer and Cohn [SC96], pages 53–62. ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1996. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=488310. IEEE

Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Teahan:1997:MET

[TC97] W. J. Teahan and J. G. Cleary. Models of English text. In Storer and Cohn [SC97], pages 12–21. ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1997. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=581953. IEEE Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Teahan:1998:TBM

[TC98] W. J. Teahan and J. G. Cleary. Tag based models of English text. In Storer and Cohn [SC98], pages 43–52. ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1998. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=672130. IEEE Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Tiwari:2006:DFV

[TC06a] M. Tiwari and P. Cosman. Dual frame video coding with pulsed quality and a lookahead window. In Storer and Cohn [SC06], pages 372–381. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ????

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=
&arnumber=1607272.

Tseng:2006:ERL

[TC06b] Hsien-Wen Tseng and Chin-Chen Chang. Error resilient locally adaptive data compression. The Journal of Systems and Software, 79(8):1156–1160, August 2006. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Tsai:2002:COM

[TCCT02] Chang-Ming Tsai, Wen-Yan Chang, Chu-Song Chen, and G. Y. Tang. Compression of 3D objects with multistage colordepth panoramic maps. In Storer and Cohn [SC02], page ?? ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2002. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=1000018. IEEE

Computer Society Order Number PR01477.

Taylor:1993:MEI

[TCJ93] H. H. Taylor, D. Chin, and A. W. Jessup. A MPEG encoder implementation on the Princeton Engine video supercomputer. In Storer and Cohn [SC93], pages 420–429. ISBN 0-8186-3391-3 (microfiche), 0-8186-3392-1 (casebound). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1993. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=253107. IEEE Computer Society Press order number 3392-02. IEEE catalog number 93TH0536-3.

Testoni:2010:ACS

[TCKM10] V. Testoni, M. H. M. Costa, D. Kirovski, and H. S. Malvar. On the adaptive coefficient scanning of JPEG XR/HD photo. In Storer and Marcellin [SM10b], pages 69–78. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5453437.

Tan:2017:JDC

[TCN^{+}17] Rui Tan, Sheng-Yuan Chiu, Hoang Hai Nguyen, David K. Y. Yau, and Deokwoo Jung. A joint data compression and encryption approach for wireless energy auditing networks. ACM Transactions on Sensor Networks, 13(2):9:1–9:??, June 2017. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Tang:2005:EIB

[TCOR05] Caimu Tang, Ngai-Man Cheung, A. Ortega, and C. S. Raghavendra. Efficient inter-band prediction and wavelet based compression for hyperspectral imagery: a distributed source coding approach. In Storer and Cohn [SC05], pages 437–446. ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=1402205.

Tang:2003:CSP

[TCP03] Xiaoli Tang, Sungdae Cho, and W. A. Pearlman. Comparison of 3D set partitioning methods in hyperspectral image compression featuring an improved 3D-SPIHT. In Storer and Cohn [SC03], page ?? ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2003. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1194068. IEEE Computer Society Order number PR01896.

Tai:1991:DCE

[TD91] W. Tai and J. Dozier. Data compression in end-to-end information systems for NASA space science missions. In Storer and Reif [SR91b], page ?? ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1990. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=213294. IEEE

Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Tao:2019:ZCF

[TDG^{+}19] Dingwen Tao, Sheng Di, Hanqi Guo, Zizhong Chen, and Franck Cappello. Z-checker: a framework for assessing lossy compression of scientific data. The International Journal of High Performance Computing Applications, 33(2):285–303, March 1, 2019. CODEN IHPCFL. ISSN 1094-3420 (print),

1741-2846 (electronic). URL https://journals.sagepub.
com/doi/full/10.1177/1094342017737147.

Tao:2019:OLC

[TDL^{+}19] D. Tao, S. Di, X. Liang, Z. Chen, and F. Cappello. Optimizing lossy compression rate-distortion from automatic online selection between SZ and ZFP. IEEE Transactions on Parallel and Distributed Systems, 30(8):1857–1871, August 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tomei:2021:BSC

[TDS^{+}21] Matthew Tomei, Shomit Das, Mohammad Seyedzadeh, Philip Bedoukian, Bradford Beckmann, Rakesh Kumar, and David Wood. Byte-select compression. ACM Transactions on Architecture and Code Optimization, 18(4):49:1–49:27, December 2021. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3462209.

Tischler:2008:CCI

[TEA08] G. Tischler, M. Ebner, and J. Albert. Color constancy from image transformations in JPEG and JPEG2000. In Storer and Marcellin [SM08], page 549. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33 INTERNET. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=4483376.

Tsaig:2003:OFL

[TEGM03] Yaakov Tsaig, Michael Elad, Gene H. Golub, and Peyman Milanfar. Optimal framework for low bit-rate block coders. In 2003 International Conference on Image Processing, 2003. ICIP 2003. Proceedings, volume 2, pages II– 219–II–222 (vol. 3). IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA,

2003. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=1246656.

Teixeira:1998:AJV

[Tei98] L. Teixeira. Analysis of a joint video coding system. In Storer and Cohn [SC98], page ?? ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1998.

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=672318. IEEE Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Teng:2000:CSD

[Ten00] Chia-Yuan Teng. Compression of SMIL documents. In Storer and Cohn [SC00], page ?? ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2000. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=838219. This

millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Teng:2004:IBP

[Ten04] C.-Y. Teng. An improved block prediction mode for H.264/ AVC intra-frame prediction. In Storer and Cohn [SC04], page ?? ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=1281545. IEEE

Computer Society Order Number: P2082.

Tetrachromat:2007:T

[Tet07] Tetrachromat. Telegraphy, 2007. URL http://en.wikipedia.

org/wiki/Tetrachromat.

Teuhola:1978:CMC

[Teu78] Jukka Teuhola. A compression method for clustered bitvectors. Information Processing Letters, 7(6):308–311, October ??, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Teuhola:2001:GAC

[Teu01] Jukka Teuhola. A general approach to compression of hierarchical indexes. Lecture Notes in Computer Science, 2113:775– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer-ny.com/
link/service/series/0558/bibs/2113/21130775.htm;
http://link.springer-ny.com/link/service/series/0558/
papers/2113/21130775.pdf.

Tramel:2011:VCS

[TF11] E. W. Tramel and J. E. Fowler. Video compressed sensing with multihypothesis. In Storer and Marcellin [SM11b], pages 193–202. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=5749477. IEEE

Computer Society Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Tuttle:1993:LPA

[TFA93] G. T. Tuttle, S. Fallahi, and A. A. Abidi. A low-power analog CMOS vector quantizer. In Storer and Cohn [SC93], pages 410–419. ISBN 0-8186-3391-3 (microfiche), 0-8186-3392-1 (casebound). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1993. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=253108. IEEE

Computer Society Press order number 3392-02. IEEE catalog number 93TH0536-3.

Titchener:1999:TCC

[TFC99] M. R. Titchener, P. M. Fenwick, and M. C. Chen. Towards a calibrated corpus for compression testing. In Storer and Cohn [SC99], page ?? ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1999. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=785711. IEEE

Computer Society Order Number PR00096. IEEE Order Plan Catalog Number PR00096.

TFD:2006:FDR

[TFD06] TFD. The Free Dictionary: RAR, 2006. URL http:/

/acronyms.thefreedictionary.com/Refund-Anticipated+
Return.

Tadayon:1998:GAL

[TFRH98] N. Tadayon, Gui-Liang Feng, T. R. N. Rao, and E. Hinds. Grouping algorithm for lossless data compression. In Storer and Cohn [SC98], page ?? ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1998. URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

672316. IEEE Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Tosic:2005:PLB

[TFV05] I. Tosic, P. Frossard, and P. Vandergheynst. Progressive low bit rate coding of simple 3D objects with matching pursuit. In Storer and Cohn [SC05], page ?? ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=1402240.

Tewari:2017:CNU

[TG17] Aakanksha Tewari and B. B. Gupta. Cryptanalysis of a novel ultra-lightweight mutual authentication protocol for IoT devices using RFID tags. The Journal of Supercomputing, 73(3): 1085–1102, March 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Tian:2003:CCS

[TGFZ03] Tao Tian, J. Garcia-Frias, and Wei Zhong. Compression of correlated sources using LDPC codes. In Storer and Cohn [SC03], page ?? ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2003. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1194069. IEEE Computer Society Order number PR01896.

Taubin:1998:PFS

[TGHL98] Gabriel Taubin, André Gueziec, William Horn, and Francis Lazarus. Progressive forest split compression. Computer Graphics, 32(Annual Conference Series):123–132, August 1998. CODEN CGRADI, CPGPBZ. ISSN 0097-8930 (print), 1558-4569 (electronic). URL http://w3imagis.imag.

fr/Membres/Frederic.Durand/Book/sig98.html;   http:
//www.acm.org:80/pubs/citations/proceedings/graph/
280814/p123-taubin/.

Turner:1997:FIF

[TH97] M. J. Turner and K. C. Halton. Facsimile-images of the future. In Storer and Cohn [SC97], page ?? ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1997. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=582143. IEEE Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Teahan:2001:CPM

[TH01] W. J. Teahan and D. J. Harper. Combining PPM models using a text mining approach. In Storer and Cohn [SC01c], pages 153–162. ISBN 0-7695-1031-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2001. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=917146. IEEE CSP number 01PR1031.

Tian:2003:UMD

[TH03] Chao Tian and S. S. Hemami. Universal multiple description scalar quantization: analysis and design. In Storer and Cohn [SC03], pages 183–192. ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2003. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1194009. IEEE Computer Society Order number PR01896.

Tian:2004:SDM

[TH04] C. Tian and S. S. Hemami. Sequential design of multiple description scalar quantizers. In Storer and Cohn [SC04], pages 32–41. ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=1281448. IEEE

Computer Society Order Number: P2082.

Tian:2005:SLM

[TH05] Chao Tian and S. S. Hemami. Staggered lattices in multiple description quantization. In Storer and Cohn [SC05], pages 398–407. ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=1402201.

Tu:2010:BPH

[TH10] Tang-Hsun Tu and Chih-Wen Hsueh. Batch-pipelining for H.264 decoding on multicore systems. In Storer and Marcellin [SM10b], page 553. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=
5453528.

Than:1994:CID

[Tha94] Soe Than. Compression of input data in machine learning from examples. Ph.D. thesis, University of Kansas, Lawrence,

KS, USA, 1994. 101 pp. URL http://search.proquest.com/
docview/304140404.

Thede:1989:HID

[The89] Leslie D. Thede. A Hybrid Image Data Compression Scheme Using Quaternary Decomposition and Vector Quantization of DCT Coefficients. Ph.D. thesis, The University of Toledo, Toledo, OH, USA, 1989. 244 pp. URL http://search.

proquest.com/docview/303826006.

Tang:2003:RAS

[THG03] Wenye Tang, Tao He, and J. D. Gibson. Realization adaptive strategies and a comparison of Fourier bandwidth, Shannon bandwidth, and Campbell bandwidth. In Storer and Cohn [SC03], pages 243–252. ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2003. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1194015. IEEE Computer Society Order number PR01896.

Tawbi:1993:VCS

[THHS93] W. Tawbi, F. Horn, E. Horlait, and J.-B. Stéfani. Video compression standards and quality of service. The Computer Journal, 36(1):43–54, February 1993. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/36/1/
43.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/Volume_36/Issue_01/Vol36_01.body.html#AbstractTawbi.

Tilton:1991:CEA

[THM91] J. C. Tilton, D. Han, and M. Manohar. Compression experiments with AVHRR data. In Storer and Reif [SR91b], pages 411–420. ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1990. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=213339. IEEE

Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Thomas:1991:EKD

[Tho91a] Kas Thomas. Entropy: the key to data compression. Dr. Dobb’s Journal of Software Tools, 16(2):32, 34, 110, February 1991. CODEN DDJOEB. ISSN 1044-789X.

Thomborson:1991:VOZ

[Tho91b] C. Thomborson. V.42bis and other Ziv–Lempel variants. In Storer and Reif [SR91b], page ?? ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1990. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=213301. IEEE Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Thomborson:1992:VSD

[Tho92]  Clark Thomborson.  The V.42bis standard for data-

compressing modems. IEEE Micro, 12(5):41–53, September/ October 1992. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Thyagarajan:2010:SIV

[Thy10] K. S. Thyagarajan. Still image and video compression with MATLAB. John Wiley, New York, NY, USA, 2010. ISBN 0-470-48416-0. ???? pp. LCCN TA1638 .T48 2010.

Teahan:1998:CET

[TICH98] W. J. Teahan, S. Inglis, J. G. Cleary, and G. Holmes. Correcting English text using PPM models. In Storer and Cohn [SC98], pages 289–298. ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1998.

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=672157. IEEE Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Ting:1977:CHT

[Tin77] T. C. Ting. Compacting homogeneous text for minimizing storage space. International Journal of Computer and Information Sciences, 6(3):211–221, September 1977. CODEN IJ-CIAH. ISSN 0091-7036.

Tinker:1989:DPI

[Tin89] Michael Tinker. DVI parallel image compression. Communications of the ACM, 32(7):844–851, July 1989. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org/pubs/toc/Abstracts/
0001-0782/65451.html.

Tischer:1987:MLD

[Tis87] Peter Tischer. A modified LZW data compression scheme. Australian Computer Science Communications, 9(1):262–272, 1987. CODEN ACSCDD. ISSN 0157-3055.

Titchener:2000:MI

[Tit00] M. R. Titchener. A measure of information. In Storer and Cohn [SC00], pages 353–362. ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2000. URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

838175. This millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Titchener:2002:CPA

[Tit02] M. R. Titchener. Compressor performance, absolutely! In Storer and Cohn [SC02], page ?? ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2002. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1000017. IEEE Computer Society Order Number PR01477.

Tilgner:1997:RBS

[TIY97] M. Tilgner, M. Ishida, and T. Yamaguchi. Recursive block structured data compression. In Storer and Cohn [SC97], page ?? ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1997. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=582139. IEEE

Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Tjalkens:2005:ICH

[Tja05] T. Tjalkens. Implementation cost of the Huffman–Shannon– Fano code. In Storer and Cohn [SC05], pages 123–132. ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1402173.

Tompkins:1999:LJC

[TK99] D. A. D. Tompkins and F. Kossentini. Lossless JBIG2 coding performance. In Storer and Cohn [SC99], page ?? ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1999. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=785710. IEEE Computer Society Or-

der Number PR00096. IEEE Order Plan Catalog Number PR00096.

Tanaka:2007:LER

[TK07] Kiyofumi Tanaka and Takahiro Kawahara. Leakage energy reduction in cache memory by data compression. ACM SIGARCH Computer Architecture News, 35(5):17–24, December 2007. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Tabus:2000:TCB

[TKR00] I. Tabus, G. Korodi, and J. Rissanen. Text compression based on variable-to-fixed codes for Markov sources. In Storer and Cohn [SC00], pages 133–142. ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2000. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=838153. This millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Tabus:2003:DSC

[TKR03] I. Tabus, G. Korodi, and J. Rissanen. DNA sequence compression using the normalized maximum likelihood model for discrete regression. In Storer and Cohn [SC03], pages 253– 262. ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2003. URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

1194016. IEEE Computer Society Order number PR01896.

Tuncel:2002:ZES

[TKRR02] E. Tuncel, P. Koulgi, S. Regunathan, and K. Rose. Zeroerror source coding with maximum distortion criterion. In Storer and Cohn [SC02], pages 92–101. ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2002. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=999947. IEEE Computer Society Order Number PR01477.

Takeda:2016:CBE

[TKYS16] Kotaro Takeda, Naoki Kobayashi, Kazuya Yaguchi, and Ayumi Shinohara. Compact bit encoding schemes for simplytyped lambda-terms. ACM SIGPLAN Notices, 51(9):146–157, September 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Thiebaut:2006:BCA

[TLLB06] C. Thiebaut, D. Lebedeff, C. Latry, and Y. Bobichon. Onboard compression algorithm for satellite multispectral images. In Storer and Cohn [SC06], page 467. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=1607310.

Todd:1985:PRC

[TLR85] Stephen Todd, Glen G. Langdon, Jr., and Jorma Rissanen. Parameter reduction and context selection for compression of gray-scale images. IBM Journal of Research and Development, 29(2):188–193, March 1985. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Tong:1991:TCM

[TM91] K.-L. Tong and M. W. Marcellin. Transform coding of monochrome and color images using trellis coded quantization. In Storer and Reif [SR91b], page ?? ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1990. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=213307. IEEE Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Turpin:1997:EAA

[TM97] A. Turpin and A. Moffat. Efficient approximate adaptive coding. In Storer and Cohn [SC97], pages 357–366. ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1997. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=582059. IEEE Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Turpin:1998:CEH

[TM98] Andrew Turpin and Alistair Moffat. Comment on “Efficient Huffman decoding” and “An efficient finite-state machine implementation of Huffman decoders”. Information Processing Letters, 68(1):1–2, October 15, 1998. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Chu97, IC98].

Taubman:2002:JIC

[TM02] David S. Taubman and Michael W. Marcellin. JPEG 2000, Image Compression Fundamentals, Standards and Practice. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2002. ?? pp.

Tao:2004:LBC

[TM04] T. Tao and A. Mukherjee. LZW based compressed pattern matching. In Storer and Cohn [SC04], page ?? ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1281544. IEEE Computer Society

Order Number: P2082.

Tao:2005:PML

[TM05a] T. Tao and Amar Mukherjee. Pattern matching in LZW compressed files. IEEE Transactions on Computers, 54(8):929– 938, August 2005. CODEN ITCOB4. ISSN 0018-9340 (print),

1557-9956 (electronic). URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1453495.

Tao:2005:MPM

[TM05b] Tao Tao and A. Mukherjee. Multiple-pattern matching in LZW compressed files using Aho–Corasick algorithm. In Storer and Cohn [SC05], page ?? ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=1402239.

Tian:2008:SGM

[TMD08] Chao Tian, S. Mohajer, and S. Diggavi. On the symmetric Gaussian multiple description rate-distortion function. In Storer and Marcellin [SM08], pages 402–411. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33 INTER-

NET. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=4483318.

Trott:1996:AWLb

[TMM96] A. Trott, R. Moorhead, and J. Mcginley. The application of wavelets to lossless compression and progressive transmission of floating point data in 3-D curvilinear grids. In Storer and Cohn [SC96], page ?? ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1996. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=488390. IEEE

Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Trendafilov:2002:ZED

[TMS02] Dimitre Trendafilov, Nasir Memon, and Torsten Suel. Zdelta: An efficient delta compression tool. Technical Report TR-CIS-2002-02, Polytechnic Institute of New York University, New York, NY, USA, 2002. ?? pp.

Teng:1995:IHL

[TN95] Chia-Yuan Teng and D. L. Neuhoff. An improved hierarchical lossless text compression algorithm. In Storer and Cohn [SC95], pages 292–301. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ????

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=515519. IEEE catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Teng:1996:QGW

[TN96] Chia-Yuan Teng and D. L. Neuhoff. Quadtree-guided wavelet image coding. In Storer and Cohn [SC96], pages 406–415. ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1996. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=488346. IEEE Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Tanskanen:1999:PMV

[TN99] J. Tanskanen and J. Niittylahti. Parallel memories in video encoding. In Storer and Cohn [SC99], page ?? ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1999. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=785709. IEEE Computer Society Or-

der Number PR00096. IEEE Order Plan Catalog Number PR00096.

Terazono:2004:EDC

[TO04] K. Terazono and Y. Okada. An extended delta compression algorithm and the recovery of failed updating in embedded systems. In Storer and Cohn [SC04], page ?? ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1281546. IEEE Computer Society

Order Number: P2082.

Todd:1989:IDC

[Tod89] Martin Peter Todd. Image Data Compression Based on a Multiresolution Signal Model (Data Compression). Ph.D. thesis, University of Warwick, Warwick, UK, 1989. 204 pp. URL

http://search.proquest.com/docview/303789294.

Timo:2011:TWR

[TOL11] R. Timo, L. Ong, and G. Lechner. The two-way relay network with arbitrarily correlated sources and an orthogonal MAC. In Storer and Marcellin [SM11b], pages 253–262. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5749483. IEEE Computer Society

Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Tomic:2006:QIB

[Tom06] R. V. Tomic. Quantized indexing: beyond arithmetic coding. In Storer and Cohn [SC06], page 468. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=1607311.

Tong:1990:DCS

[Ton90] Tak Yen Tong. Diversity and coding in spread spectrum systems. Master’s thesis, University of Waterloo, Waterloo, ON,

Canada, 1990. 1 pp. URL http://search.proquest.com/
docview/303906263.

Tong:1993:DCA

[Ton93] Tak Yen Tong. Data compression with arithmetic coding and source modeling. Ph.D. thesis, University of Waterloo, Waterloo, ON, Canada, 1993. 161 pp. URL http://search.

proquest.com/docview/304101874. 2 microfiches.

Topiwala:1996:HPW

[Top96] P. Topiwala. High performance wavelet image compression optimized for MSE and HVS metrics. In Storer and Cohn [SC96], page ?? ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1996. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=488389. IEEE

Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Toussaint:2006:CRS

[Tou06] Marc Toussaint. Compact representations as a search strategy: Compression EDAs. Theoretical Computer Science, 361(1): 57–71, August 28, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tamir:1996:ECC

[TPAK96] D. E. Tamir, K. Phillips, and A. Abdul-Karim. Efficient chain-code encoding for segmentation-based image compression. In Storer and Cohn [SC96], page ?? ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1996. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=488387. IEEE Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Tschumperle:2020:RSC

[TPM20] David Tschumperlé, Christine Porquet, and Amal Mahboubi. Reconstruction of smooth 3D color functions from keypoints: Application to lossy compression and exemplar-based generation of color LUTs. SIAM Journal on Imaging Sciences, 13 (3):1511–1535, ???? 2020. CODEN SJISBI. ISSN 1936-4954.

Tamir:1995:VQC

[TPY95] D. E. Tamir, Chi-Yeon Park, and Wook-Sung Yoo. Vector quantization and clustering: a pyramid approach. In Storer and Cohn [SC95], page 482. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=515592. IEEE catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Tung:1988:FAO

[TR88] C.-H. Tung and J. P. Robinson. A fast algorithm for optimum syndrome space compression. IEEE Transactions on Computers, 37(2):228–232, February 1988. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=2153.

Teuhola:1991:PAC

[TR91] J. Teuhola and T. Raita. Piecewise arithmetic coding. In Storer and Reif [SR91b], pages 33–42. ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1990. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=213371. IEEE Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Teuhola:1993:AFS

[TR93] Jukka Teuhola and Timo Raita. Application of a finitestate model to text compression. The Computer Journal, 36(7):607–614, ???? 1993. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/36/7/607.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/Volume_36/
Issue_07/Vol36_07.body.html#AbstractTeuhola.

Taubin:1998:GCT

[TR98] Gabriel Taubin and Jarek Rossignac. Geometric compression through topological surgery. ACM Transactions on Graphics, 17(2):84–115, April 1998. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL http://

www.acm.org:80/pubs/citations/journals/tog/1998-17-
2/p84-taubin/.

Tran:1987:DCB

[Tra87] Thao Duy Tran. Data Compression of Bauded Signals Using the Baseband Residual Vector Quantization Algorithm. Ph.D. thesis, The University of Utah, Salt Lake City, UT 84112, USA,

1987. 167 pp. URL http://search.proquest.com/docview/
303609161.

Trott:1996:AWLa

[Tro96] Aaron Gregory Trott. The application of wavelets to lossless compression and progressive transmission of floating point data in three-dimensional curvilinear grids. M.S.E.E. thesis, Mississippi State University, Mississippi State, MS 39762,

USA, 1996. 103 pp. URL http://search.proquest.com/
docview/304273691.

Theisel:2003:CCV

[TRS03] H. Theisel, Ch. Rössl, and H.-P. Seidel. Compression: Compression of 2D vector fields under guaranteed topology preservation. Computer Graphics Forum, 22(3):333–342, September 2003. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Thayammal:2017:EPM

[TS17] S. Thayammal and D. Selvathi. Edge preserved multispectral image compression using extended shearlet transform.

The Computer Journal, 60(7):986–994, July 1, 2017. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL https://academic.oup.com/comjnl/article/
60/7/986/2608041.

Tsai:1991:SPC

[Tsa91] Y. T. Tsai. Signal processing and compression for image capturing systems using a single-chip sensor. In Storer and Reif [SR91b], page ?? ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1990. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=213313. IEEE

Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Tsai:1998:CIC

[Tsa98] Min-Jen Tsai. Color image compression by stack-run-end coding. In Storer and Cohn [SC98], page ?? ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1998. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=672319. IEEE Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Tsekouras:2005:FVQ

[Tse05] George E. Tsekouras. A fuzzy vector quantization approach to image compression. Applied Mathematics and Computation, 167(1):539–560, August 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0096300304005120.

Tsai:2021:LST

[TSFS21] Po-An Tsai, Andres Sanchez, Christopher W. Fletcher, and Daniel Sanchez. Leaking secrets through compressed caches. IEEE Micro, 41(3):27–33, May/June 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Tanaka:1998:DCD

[TSS^{+}98] Masaharu Tanaka, Hidenori Sakanashi, Mehrdad Salami, Masaya Iwata, Takio Kurita, and Tetsuya Higuchi. Data compression for digital color electrophotographic printer with evolvable hardware. Lecture Notes in Computer Science, 1478:106–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer-

ny.com/link/service/series/0558/bibs/1478/14780106.
htm;   http://link.springer-ny.com/link/service/series/
0558/papers/1478/14780106.pdf.

Tan:1999:PIB

[TSS99] Min Tan, Janet M. Siegel, and Howard Jay Siegel. Parallel implementations of block-based motion vector estimation for video compression on four parallel processing systems. International Journal of Parallel Programming, 27(3):195–225, June 1999. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=0885-7458&volume=27&issue=3&
spage=195.

Tsui:1972:DCP

[Tsu72] Stephen S. T. Tsui. Data compression with particular applications to video signals. M.Eng., Department of Electrical Engineering, McGill University, Montréal, QC, Canada, December 1972. iii + 151 pp. URL http:

//digitool.library.mcgill.ca/thesisfile51654.pdf;
http://www.collectionscanada.ca/obj/thesescanada/vol1/
QMM/TC-QMM-51654.pdf.

Takamura:1994:LIC

[TT94] S. Takamura and M. Takagi. Lossless image compression with lossy image using adaptive prediction and arithmetic coding. In Storer and Cohn [SC94], pages 166–174. ISBN 0-8186-5636-0, 0-8186-5637-9 (case). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1994.

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=305924. IEEE Catalog Number 93TH0626-2. IEEE Computer Society Press Order Number 5637-02.

Tay:2007:BDB

[TTW07] Wee-Peng Tay, J. N. Tsitsiklis, and M. Z. Win. Bayesian detection in bounded height tree networks. In Storer and Cohn [SC07], pages 243–252. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ????

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=4148763. IEEE Computer Society Order Number P2791.

Tunstall:1967:SNC

[Tun67] B. P. Tunstall. Synthesis of Noiseless Compression Codes. Ph.D. dissertation, Georgia Institute of Technology, Atlanta, GA, USA, 1967. ?? pp.

Tuncel:2004:OMS

[Tun04] E. Tuncel. On optimal multiresolution source-channel coding across degraded broadcast channels. In Storer and Cohn [SC04], pages 122–131. ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

1281457. IEEE Computer Society Order Number: P2082.

Turner:1975:GCS

[Tur75] L. F. Turner. The ‘on-ground’ compression of satellite data. The Computer Journal, 18(3):243–247, August 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/
content/18/3/243.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_18/Issue_03/tiff/
243.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_03/tiff/244.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_18/Issue_03/
tiff/245.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_03/tiff/246.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_18/Issue_03/
tiff/247.tif.

Turner:1995:LRC

[Tur95] M. J. Turner. Lossless region coding schemes. In Storer and Cohn [SC95], page 481. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

515591. IEEE catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Turner:1996:IOU

[Tur96] M. J. Turner. Image operations using a semi-compressed contour tree image definition. In Storer and Cohn [SC96], page ?? ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN QA76.9.D33 D37 1996. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=488391. IEEE

Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Tabus:2009:LBR

[TV09] I. Tabus and A. Vasilache. Low bit rate vector quantization of outlier contaminated data based on shells of Golay codes. In Storer and Marcellin [SM09], pages 302–311. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4976474. IEEE Computer Society

Order Number P3592. BMS Part Number CFP09DCC-PRT.

Tahoces:2008:ICM

[TVLS08] Pablo G. Tahoces, J. Ramón Varela, María J. Lado, and Miguel Souto. Image compression: Maxshift ROI encoding options in JPEG2000. Computer Vision and Image Understanding: CVIU, 109(2):139–145, February 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Triantafyllidis:2003:EAE

[TVS^{+}03] G. A. Triantafyllidis, M. Varnuska, D. Sampson, D. Tzovaras, and M. G. Strintzis. An efficient algorithm for the enhancement of JPEG-coded images. Computers and Graphics, 27(4): 529–534, August 2003. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Tjalkens:1997:CTW

[TVW97] T. J. Tjalkens, P. A. J. Volf, and F. M. J. Willems. A context-tree weighting method for text generating sources. In Storer and Cohn [SC97], page ?? ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1997. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=582140. IEEE Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Tjalkens:1992:UVF

[TW92] T. Tjalkens and Frans M. Willems. A universal variable-to-fixed length source code based on Lawrence’s algorithm. IEEE Transactions on Information Theory, 38(2):247–253, March 1992. CODEN IETTAW. ISSN 0018-9448 (print), 1557-9654 (electronic).

Tian:1996:LIC

[TW96a] Jun Tian and R. O. Wells, Jr. A lossy image codec based on index coding. In Storer and Cohn [SC96], page ?? ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1996. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=488388. IEEE Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Turner:1996:ELI

[TW96b] Martin J. Turner and Neil E. Wiseman. Efficient lossless image contour coding. Computer Graphics Forum, 15(2):107–117, June 1996. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tjalkens:2000:VFL

[TW00] T. Tjalkens and F. Willems. Variable-to-fixed length codes: a geometrical approach to low-complexity source codes. In Storer and Cohn [SC00], page ?? ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2000. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=838220. This

millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Tsai:2001:CHC

[TW01a] C. W. Tsai and J. L. Wu. On constructing the Huffman-code based reversible variable length codes. IEEE Transactions on Communications, 49(9):1506–1509, September 2001. CODEN IECMBT. ISSN 0090-6778 (print), 1558-0857 (electronic).

Tsai:2001:MSR

[TW01b] Chien-Wu Tsai and Ja-Ling Wu. Modified symmetrical reversible variable-length code and its theoretical bounds. IEEE Transactions on Information Theory, 47(6):2543–2548, September 2001. CODEN IETTAW. ISSN 0018-9448 (print), 1557-9654 (electronic).

Tournier:2009:MCM

[TWC^{+}09] M. Tournier, X. Wu, N. Courty, E. Arnaud, and L. Revéret. Motion compression: Motion compression using principal geodesics analysis. Computer Graphics Forum, 28(2):355–364, April 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tham:2002:DLC

[TWCA02] Wei-Ching Tham, S. I. Woolley, S. Cribbs, and D. Anderson. Diagnostically lossless compression of pipeline inspection data. In Storer and Cohn [SC02], page ?? ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2002. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1000016. IEEE Computer Society Order Number PR01477.

Takishima:1995:RVL

[TWM95] Y. Takishima, M. Wada, and H. Murakami. Reversible variable-length codes. IEEE Transactions on Communications, 43(2,3,4):158–162, February/March/April 1995. CO-DEN IECMBT. ISSN 0090-6778 (print), 1558-0857 (electronic).

Tischer:1993:CBL

[TWMG93] P. E. Tischer, R. T. Worley, A. J. Maeder, and M. Goodwin. Context-based lossless image compression. The Computer Journal, 36(1):68–77, February 1993. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/36/1/
68.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/Volume_36/Issue_01/Vol36_01.body.html#AbstractTischer.

Tao:2010:HSP

[TWW^{+}10] Pin Tao, Wenting Wu, Chao Wang, Mou Xiao, and Jiangtao Wen. Horizontal spatial prediction for high dimension intra coding. In Storer and Marcellin [SM10b], page 552. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=5453529.

Takamura:2004:MVC

[TY04] S. Takamura and Y. Yashima. Multiband video coding using H.264/AVC, MPEG-4 studio profile and JPEG 2000. In Storer and Cohn [SC04], page ?? ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1281543. IEEE Computer Society Order Number: P2082.

Takamura:2007:GGC

[TY07] S. Takamura and Y. Yashima. Gaussian Golomb codes. In Storer and Cohn [SC07], page 401. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=4148802. IEEE Computer Society Order Number P2791.

Takamura:2008:PEP

[TY08] S. Takamura and Y. Yashima. On performance evaluation of predictive coding using a residue-free approach. In Storer and Marcellin [SM08], page 548. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33 INTERNET. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=4483375.

Tu:2020:RBS

[TYD^{+}20] Weiping Tu, Yuhong Yang, Bo Du, Wanzhao Yang, Xiong Zhang, and Jiaxi Zheng. RNN-based signal classification for hybrid audio data compression. Computing: Archiv für informatik und numerik, 102(3):813–827, March 2020. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

UCLC:2008:U

[UCL08] UCLC. [unknown], 2008. URL http://uclc.info.

UDA:2008:U

[UDA08] UDA. [unknown], 2008. URL http://wex.cn/dwing/mycomp.

htm.

Udell:1993:FIE

[Ude93] Jon Udell. First impressions: Easy does it with MS-DOS 6.0: Microsoft adds compression and memory management to the venerable operating system. Byte Magazine, 18(4):44–??, April 1993. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Uhl:1996:PIC

[UH96] A. Uhl and J. Hammerle. Parallel image compression on a workstation cluster using PVM. In Bode et al. [BDLS96], pages 301–?? ISBN 3-540-61779-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.E975 1996.

Unat:2009:ASS

[UHB09] D. Unat, T. Hromadka, and S. B. Baden. An adaptive subsampling method for in-memory compression of scientific data. In Storer and Marcellin [SM09], pages 262–271. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4976470. IEEE Computer Society

Order Number P3592. BMS Part Number CFP09DCC-PRT.

Uhl:1995:PCC

[Uhl95] A. Uhl. Parallel compact coding of satellite images with wavelet packets using PVM. In Prasanna et al. [PBPT95], pages 382–387. ISBN 0-07-462332-X. LCCN QA 76.58 I587 1994.

Uhl:1996:PAU

[Uhl96] A. Uhl. Parallel algorithms for using non-stationary MRA in image compression. Lecture Notes in Computer Science, 1124: 151–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

uklinux:2007:U

[ukl07] uklinux. [unknown], 2007. URL http://www.bckelk.

uklinux.net/menu.html.

Unisys:2003:UCH

[Uni03] Unisys. UNISYS Corporation home page, 2003. URL http:/

/www.unisys.com.

Unicode:2003:UCH

[Uni07a] Unicode. The Unicode Consortium home page, 2003, 2007.

URL http://unicode.org/.

Consortium:2007:GC

[Uni07b] Unicode Consortium. Greek and Coptic. Technical report, The Unicode Consortium, Mountain View, CA, USA, 2007. 5

pp. URL http://www.unicode.org/charts/PDF/U0370.pdf.

From The Unicode Standard, Version 6.2.

Udupa:1999:PC

[UPR99] Raghavendra U. Udupa, Vinayaka D. Pandit, and Ashok Rao. Private communication, 1999.

UPX:2003:UUP

[UPX03] UPX. UPX: Ultimate Packer for Executables, 2003. URL

http://upx.sourceforge.net/.

Rahman:1996:WCR

[uR96] Z. u. Rahman. Wavelet coding for remote sensed data. In Storer and Cohn [SC96], page ?? ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1996. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=488384. IEEE Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

USENIX:1990:CSSa

[USE90] USENIX, editor. Computing Systems, Spring, 1990. USENIX, Berkeley, CA, USA, Spring 1990.

USENIX:1993:PWU

[USE93] USENIX, editor. Proceedings of the Winter 1993 USENIX Conference: January 25–29, 1993, San Diego, California, USA. USENIX, Berkeley, CA, USA, 1993. ISBN 1-880446-48-0. LCCN QA 76.76 O63 U84 1993. URL http://www.

usenix.org/publications/library/proceedings/sd93/.

Usevitch:1996:OBA

[Use96] B. Usevitch. Optimal bit allocation for biorthogonal wavelet coding. In Storer and Cohn [SC96], pages 387–395. ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1996. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=488344. IEEE Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Usevitch:2003:JEB

[Use03] B. E. Usevitch. JPEG2000 extensions for bit plane coding of floating point data. In Storer and Cohn [SC03], page ?? ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2003. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=1194070. IEEE

Computer Society Order number PR01896.

Usevitch:2005:JCL

[Use05] B. Usevitch. JPEG2000 compliant lossless coding of floating point data. In Storer and Cohn [SC05], page ?? ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1402241.

Ueno:2017:BCF

[USY17] Tomohiro Ueno, Kentaro Sano, and Satoru Yamamoto. Bandwidth compression of floating-point numerical data streams for FPGA-based high-performance computing. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 10(3):18:1–18:??, July 2017. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Uno:2012:BNR

[UUiN12] Takeaki Uno, Ryuhei Uehara, and Shin ichi Nakano. Bounding the number of reduced trees, cographs, and series-parallel graphs by compression. Lecture Notes in Computer Science, 7157:5–16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/
chapter/10.1007/978-3-642-28076-4_4/.

vanderVleuten:1999:NMM

[van99] R. J. van der Vleuten. New methods for multiplicationfree arithmetic coding. In Storer and Cohn [SC99], page ?? ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1999. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=785712. IEEE Computer Society Or-

der Number PR00096. IEEE Order Plan Catalog Number PR00096.

vanderVleuten:2002:IES

[van02a] R. J. van der Vleuten. Improved elastic storage of digital still images. In Storer and Cohn [SC02], page ?? ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2002. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=1000019. IEEE

Computer Society Order Number PR01477.

vanderVleuten:2002:LCL

[van02b] R. J. van der Vleuten. Low-complexity lossless and finegranularity scalable near-lossless compression of color images. In Storer and Cohn [SC02], page ?? ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2002. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1000020. IEEE Computer Society Order Number PR01477.

Vasudevan:2007:BRD

[Vas07] D. Vasudevan. Bounds to the rate distortion tradeoff of the binary Markov source. In Storer and Cohn [SC07], pages 343– 352. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=4148773. IEEE Com-

puter Society Order Number P2791.

vonBerg:1997:EBA

[vB97] D. C. Linne von Berg. Effects of boundary artifacts in parallel implementations of wavelet compression for large-format digital framing reconnaissance camera systems. In Storer and Cohn [SC97], page ?? ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1997. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=582113. IEEE

Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

vanderVleuten:1998:LCB

[vB98] R. J. van der Vleuten and F. Bruekers. Lossless compression of binary audio signals. In Storer and Cohn [SC98], page ?? ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN QA76.9.D33 D232 1998. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=672321. IEEE

Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Vasa:2014:AIM

[VB14] L. Váša and G. Brunnett. Animation II (motion data processing): Rate-distortion optimized compression of motion capture data. Computer Graphics Forum, 33(2):283–292, May 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

VanMaren:1989:MTC

[VBK89] David J. Van Maren, Mark J. Bianchi, and Jeffery J. Kato. Maximizing tape capacity by super-blocking. Hewlett–Packard Journal, 40(3):32–34, June 1989. CODEN HPJOAX. ISSN 0018-1153.

Villasenor:1994:FES

[VBL94] J. D. Villasenor, B. Belzer, and J. Liao. Filter evaluation and selection in wavelet image compression. In Storer and Cohn [SC94], pages 351–360. ISBN 0-8186-5636-0, 0-8186-5637-9 (case). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1994. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=305943. IEEE

Catalog Number 93TH0626-2. IEEE Computer Society Press Order Number 5637-02.

Valette:2009:HCP

[VCP09] Sébastien Valette, Raphaëlle Chaine, and Rémy Prost. Hierarchies and compression: Progressive lossless mesh compression via incremental parametric refinement. Computer Graphics Forum, 28(5):1301–1310, July 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Vucetic:2009:CKP

[VCW09] S. Vucetic, V. Coric, and Zhuang Wang. Compressed kernel perceptrons. In Storer and Marcellin [SM09], pages 153–162. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=4976459. IEEE Computer

Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

Vasa:2018:EPC

[VD18] L. Váša and J. Dvořák. Error propagation control in Laplacian mesh compression. Computer Graphics Forum, 37(5):61–70, August 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

VanAssche:2000:EIR

[VDPL00] S. Van Assche, D. De Rycke, W. Philips, and I. Lemahieu. Exploiting interframe redundancies in the lossless compression of 3D medical images. In Storer and Cohn [SC00], page ?? ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2000. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=838222. This millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Villasenor:1996:SDC

[VED96] J. D. Villasenor, R. A. Ergas, and P. L. Donoho. Seismic data compression using high-dimensional wavelet transforms. In Storer and Cohn [SC96], pages 396–405. ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1996. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=488345. IEEE Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Venkatesan:2007:GSF

[Ven07] R. C. Venkatesan. Generalized statistics framework for rate distortion theory. In Storer and Cohn [SC07], page 402. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4148803. IEEE Computer Society

Order Number P2791.

Ven:2017:LDC

[Ven17] Arjan Van De Ven. Linux OS data compression options: Comparing behavior. Web site, January 3, 2017. URL https://

clearlinux.org/blogs/linux-os-data-compression-options-
comparing-behavior.

Verdu:1998:FYS

[Ver98] S. Verdu. Fifty years of Shannon theory. IEEE Transactions on Information Theory, 44(6):2057–2078, October 1998. CODEN IETTAW. ISSN 0018-9448 (print), 1557-9654 (electronic).

Varnica:2000:MRA

[VFE00] N. Varnica, M. Fleming, and M. Effros. Multi-resolution adaptation of the SPIHT algorithm for multiple description. In Storer and Cohn [SC00], pages 303–312. ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2000. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=838170. This millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Vlachos:2015:CMF

[VFK15] Michail Vlachos, Nikolaos M. Freris, and Anastasios Kyrillidis. Compressive mining: fast and optimal data mining in the compressed domain. VLDB Journal: Very Large Data Bases, 24 (1):1–24, February 2015. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Kam:1995:LCC

[VG95] R. A. Vander Kam and R. M. Gray. Lossy compression of clustered-dot halftones using sub-cell prediction. In Storer and Cohn [SC95], pages 112–121. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=515501. IEEE catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Varshney:2006:TSC

[VG06] L. R. Varshney and V. K. Goyal. Toward a source coding theory for sets. In Storer and Cohn [SC06], pages 13–22. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1607236.

Villasenor:1993:FFC

[Vil93] J. D. Villasenor. Full-frame compression of tomographic images using the discrete Fourier transform. In Storer and Cohn [SC93], pages 195–203. ISBN 0-8186-3391-3 (microfiche), 0-8186-3392-1 (casebound). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1993. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=253130. IEEE Computer Society Press order number 3392-02. IEEE catalog number 93TH0536-3.

Vitter:1987:DAD

[Vit87] Jeffrey Scott Vitter. Design and analysis of dynamic Huffman codes. Journal of the ACM, 34(4):825–845, October 1987. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (elec-

tronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/42227.html.

Vitter:1989:ADH

[Vit89] Jeffrey Scott Vitter. Algorithm 673: Dynamic Huffman coding. ACM Transactions on Mathematical Software, 15(2):158– 167, June 1989. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL http://www.acm.org/pubs/

citations/journals/toms/1989-15-2/p158-vitter/. See

remark [NK16].

Villasenor:1995:WVC

[VJB^{+}95] J. Villasenor, R. Jain, B. Belzer, W. Boring, C. Chien, C. Jones, J. Liao, S. Molloy, S. Nazareth, B. Schoner, and J. Short. Wireless video coding system demonstration. In Storer and Cohn [SC95], page 448. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=515558. IEEE catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Vitter:1991:OPD

[VK91] J. S. Vitter and P. Krishnan. Optimal prefetching via data compression. In IEEE [IEE91], pages 121–130. CODEN ASFPDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Vetterli:1995:WSC

[VK95] M. Vetterli and J. Kovacevic. Wavelets and Subband Coding. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1995. ?? pp.

Vitter:1996:OPD

[VK96] Jeffrey Scott Vitter and P. Krishnan. Optimal prefetching via data compression. Journal of the ACM, 43(5):771–793, September 1996. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.org/pubs/

toc/Abstracts/jacm/234753.html.

vanderVleuten:2000:LCS

[vK00] R. J. van der Vleuten and R. P. Kleihorst. Low-complexity scalable image compression. In Storer and Cohn [SC00], pages 23–32. ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 2000. URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=838142. This millen-

nial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Vasylevska:2017:CVF

[VK17] Khrystyna Vasylevska and Hannes Kaufmann. Compressing VR: Fitting large virtual environments within limited physical space. IEEE Computer Graphics and Applications, 37(5):85–91, September/October 2017. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

URL https://www.computer.org/csdl/mags/cg/2017/05/
mcg2017050085-abs.html.

Venkataramani:2001:SRT

[VKG01] R. Venkataramani, G. Kramer, and V. K. Goyal. Successive refinement on trees: a special case of a new MD coding region. In Storer and Cohn [SC01c], pages 293–301. ISBN 0-7695-1031-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 2001. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=917160. IEEE CSP num-

ber 01PR1031.

Venbrux:1991:VHS

[VL91] J. Venbrux and N. Liu. A very high speed lossless compression/decompression chip set. In Storer and Reif [SR91b], page ?? ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1990. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=213300. IEEE

Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

vanLeeuwen:1994:GTC

[vL94] Jan van Leeuwen, editor. Graph-theoretic concepts in computer science: 19th International Workshop, WG ’93, Utrecht, The Netherlands, June 16–18, 1993: proceedings, volume 790 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. CODEN LNCSD9. ISBN 0-86636-292-4 (New York), 3-540-57899-4 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

QA75.5 .I647 1993. URL http://link.springer-ny.com/
link/service/series/0558/tocs/t0790.htm;   http://
www.springerlink.com/content/978-0-86636-292-4;   http:
//www.springerlink.com/openurl.asp?genre=issue&issn=
0302-9743&volume=790.

Vasconcelos:1997:LBC

[VL97] N. Vasconcelos and A. Lippman. Library-based coding: a representation for efficient video compression and retrieval. In Storer and Cohn [SC97], pages 121–130. ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1997. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=581989. IEEE Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Vasconcelos:1998:BFC

[VL98] N. Vasconcelos and A. Lippman. A Bayesian framework for content-based indexing and retrieval. In Storer and Cohn [SC98], page ?? ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1998. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=672322. IEEE

Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Verma:1998:SEC

[VM98] T. Verma and T. Meng. A scalable entropy code. In Storer and Cohn [SC98], page ?? ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN QA76.9.D33 D232 1998. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=672323. IEEE

Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Varodayan:2007:DGS

[VMFG07] D. Varodayan, A. Mavlankar, M. Flierl, and B. Girod. Distributed grayscale stereo image coding with unsupervised learning of disparity. In Storer and Cohn [SC07], pages 143– 152. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=4148753. IEEE Com-

puter Society Order Number P2791.

Valero:2016:ELD

[VMP^{+}16] Alejandro Valero, Negar Miralaei, Salvador Petit, Julio Sahuquillo, and Timothy M. Jones. Enhancing the L1 data cache design to mitigate HCI. IEEE Computer Architecture Letters, 15(2):93–96, July/December 2016. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Vaughan-Nichols:1990:GYBb

[VN90a] S. J. Vaughan-Nichols. Getting your byte’s worth: Hardwarebased data compression gives you more bang for your QIC, DAT, and hard disk buck. Byte Magazine, 15(12):331–332, 334–336, November 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Vaughan-Nichols:1990:SSW

[VN90b] S. J. Vaughan-Nichols. Saving space: Whatever size hard disk you have, it’s probably nearly full. data compression can help. Byte Magazine, 15(3):237–238, 240, 242–243, March 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Verlani:2008:PPB

[VN08] P. Verlani and P. J. Narayanan. A parametric proxy-based compression of depth movies. In Storer and Marcellin [SM08], page 550. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33

INTERNET. URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4483377.

Veness:2012:CTS

[VNHB12] J. Veness, Kee Siong Ng, M. Hutter, and M. Bowling. Context tree switching. In Storer and Marcellin [SM12], pages 327–336. ISBN 1-4673-0715-7 (paperback). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

6189264. IEEE Computer Society order number P4656.

VanHouten:1991:ATS

[VO91] K. Van Houten and P. W. Oman. An algorithm for tree structure compression. In Storer and Reif [SR91b], page ?? ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1990. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=213337. IEEE Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Vo:2007:CDT

[Vo07] Kiem-Phong Vo. Compression as data transformation. In Storer and Cohn [SC07], page 403. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=4148804. IEEE Computer Society Order Number P2791.

Varpaaniemi:2008:MSQ

[VO08] Kimmo Varpaaniemi and Leo Ojala. Modelling and simulation of quantum teleportation and dense coding using predicate/ transition-nets. Fundamenta Informaticae, 85(1–4):465–479, September 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Volf:1997:CTW

[Vol97] Paul A. J. Volf. A context-tree weighting algorithm for text generating sources [poster]. In James A. Storer, editor, DCC ’97: Data Compression Conference, pages 132–139. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997.

Volf:2002:WTD

[Vol02] Paulus Adrianus Jozef Volf. Weighting techniques in data compression: Theory and algorithms. Doctoral thesis, Technische Universiteit Eindhoven, Eindhoven, The Netherlands,

2002. 175 (or 185??) pp. URL http://search.proquest.
com/docview/305446928.

Volkerink:2004:DTT

[Vol04] Erik H. Volkerink. Design-for-testability for test data compression. Ph.D. thesis, Stanford University, Stanford, CA, USA,

2004. 95 pp. URL http://search.proquest.com/docview/
305130276.

Vontobel:2004:FGA

[Von04] P. O. Vontobel. A factor-graph approach to the contexttree weighting method. In Storer and Cohn [SC04], page ?? ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1281547. IEEE Computer

Society Order Number: P2082.

Vorobev:1961:FN

[Vor61] Nikolai N. Vorobev. Fibonacci numbers. In Ian Naismith Sneddon, editor, Fibonacci Numbers, volume 2 of Popular lectures in mathematics series, pages viii + 66. Pergamon Press, Oxford, UK, 1961. ISBN 0-932750-03-6. LCCN QA241 .V613 1961. Translation by Halina Moss. Reprinted in [Vor83].

Vorobev:1983:FN

[Vor83] Nikolai N. Vorobev. Fibonacci numbers. In Ian Naismith Sneddon, editor, Fibonacci Numbers, page ?? New Classics Library, ????, 1983. Translation by Halina Moss. Reprint of [Vor61].

Vosoughi:2006:TEC

[Vos06] Azadeh Vosoughi. Tracking and equalizing channels in communication systems and in data compression. Ph.D. thesis, Cornell University, Ithaca, NY, USA, 2006. 172 pp. URL

http://search.proquest.com/docview/305322063.

Vovk:2006:WCP

[Vov06] Vladimir Vovk. Well-calibrated predictions from on-line compression models. Theoretical Computer Science, 364(1):10–26, November 2, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Villafranca:2010:DOL

[VPGB10] A. G. Villafranca, J. Portell, and E. Garcia-Berro. Development of optimum lossless compression systems for space missions. In Storer and Marcellin [SM10b], page 554. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5453531.

VanAssche:1998:LCP

[VPL98] S. Van Assche, W. Philips, and I. Lemahiue. Lossless compression of pre-press images using linear color decorrelation. In Storer and Cohn [SC98], page ?? ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1998. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=672320. IEEE Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Vargas-Perez:2017:HMO

[VPS17] Sandino Vargas-Perez and Fahad Saeed. A hybrid MPI-OpenMP strategy to speedup the compression of big nextgeneration sequencing datasets. IEEE Transactions on Parallel and Distributed Systems, 28(10):2760–2769, October 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL https://www.computer.org/csdl/trans/td/
2017/10/07895161-abs.html.

Vanam:2007:DCO

[VRHL07] R. Vanam, E. A. Riskin, S. S. Hemami, and R. E. Ladner. Distortion-complexity optimization of the H.264/MPEG-4 AVC encoder using the GBFOS algorithm. In Storer and Cohn [SC07], pages 303–312. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ????

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=4148769. IEEE Computer Society Order Number P2791.

Vanam:2009:HMA

[VRL09] R. Vanam, E. A. Riskin, and R. E. Ladner. H.264/MPEG-4 AVC encoder parameter selection algorithms for complexity distortion tradeoff. In Storer and Marcellin [SM09], pages 372–381. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=4976481. IEEE

Computer Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

Vidhyaa:2016:HED

[VRR^{+}16] V. G. Vidhyaa, S. A. Rajalakshmi, R. Raghavan, G. S. V. Venu Gopal, and R. Gandhiraj. Huffman encoding and decoding algorithm using IJulia. In 2016 International Conference on Communication and Signal Processing (ICCSP), pages 0587– 0591. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2016.

Vasa:2010:GDL

[VS10] Libor Váša and Václav Skala. Geometry-driven local neighbourhood based predictors for dynamic mesh compression. Computer Graphics Forum, 29(6):1921–1933, September 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Veaux:2000:CDU

[VSG00] C. Veaux, P. Scalart, and A. Gilloire. Channel decoding using inter- and intra-correlation of source encoded frames. In Storer and Cohn [SC00], pages 103–112. ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2000. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=838150. This millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Vaishnav:2011:NCE

[VST11] M. Vaishnav, A. Sharma, and A. K. Tiwari. A novel computationally efficient motion compensation method based on pixel by pixel prediction. In Storer and Marcellin [SM11b], page 480. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=5749537. IEEE

Computer Society Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Vaisey:1995:BMS

[VT95] J. Vaisey and M. Trumbo. Bitgroup modeling of signal data for image compression. In Storer and Cohn [SC95], page 466. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=515576. IEEE catalog

number 95TH8037. IEEE Computer Society Press order number PR07010.

Vucetic:2005:AOQ

[Vuc05] S. Vucetic. Accuracy-optimized quantization for highdimensional data fusion. In Storer and Cohn [SC05], page ?? ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1402242.

Vucetic:2006:FAL

[Vuc06] S. Vucetic. A fast algorithm for lossless compression of data tables by reordering. In Storer and Cohn [SC06], page 469. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1607312.

Vo:2004:UCD

[VV04] B. D. Vo and K.-P. Vo. Using column dependency to compress tables. In Storer and Cohn [SC04], pages 92–101. ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1281454. IEEE Computer Society

Order Number: P2082.

VanderKam:1994:CJC

[VW94] R. A. Vander Kam and P. W. Wong. Customized JPEG compression for grayscale printing. In Storer and Cohn [SC94], pages 156–165. ISBN 0-8186-5636-0, 0-8186-5637-9 (case). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1994. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=305923. IEEE

Catalog Number 93TH0626-2. IEEE Computer Society Press Order Number 5637-02.

Volf:1998:SBT

[VW98] P. A. J. Volf and F. M. J. Willems. Switching between two universal source coding algorithms. In Storer and Cohn [SC98], pages 491–500. ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1998. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=672217. IEEE

Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Wolf:2000:SCS

[W^{+}00] Misha Wolf et al. A standard compression scheme for Unicode. Unicode Technical Report 6, The Unicode Consortium, Mountain View, CA, USA, 2000. ?? pp. URL http://unicode.

org/unicode/reports/tr6/index.html.

Wakatani:2005:PVC

[Wak05] A. Wakatani. Parallelization of VQ codebook generation by two algorithms: parallel LBG and aggressive PNN [image compression applications]. In Storer and Cohn [SC05], page ?? ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1402243.

Wakatani:2006:VCA

[Wak06] A. Wakatani. VQ compression algorithms on a shared-memory multiprocessor system. In Storer and Cohn [SC06], page 470. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1607313.

Wakatani:2007:AOV

[Wak07] A. Wakatani. Advanced optimizations for VQ compression on parallel systems. In Storer and Cohn [SC07], page 404. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=4148805. IEEE Computer

Society Order Number P2791.

Wallace:1991:JSI

[Wal91a] Gregory K. Wallace. The JPEG still image compression standard. Communications of the ACM, 34(4):30–44, April 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wallace:1991:JSP

[Wal91b] Gregory K. Wallace. The JPEG still picture compression standard. Communications of the ACM, 34(4):30–44, April 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://www.acm.org/pubs/toc/Abstracts/
0001-0782/103089.html.

Wang:2005:DVB

[WAL05] Yadong Wang, F. Alajaji, and T. Linder. Design of VQbased hybrid digital-analog joint source-channel codes for image communication. In Storer and Cohn [SC05], pages 193– 202. ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=1402180.

Wang:1973:SST

[Wan73] Robert Tien Pei Wang. A Study of the Structure of Television Pictures as Applied to Practical Data Compression Techniques. Ph.D. thesis, Stanford University, Stanford, CA, USA,

1973. 106 pp. URL http://search.proquest.com/docview/
302674792.

Wang:1988:AAO

[Wan88] Muzhong Wang. Almost asymptotically optimal flag encoding of the integers. IEEE Transactions on Information Theory, 34 (2):324–326, March 1988. CODEN IETTAW. ISSN 0018-9448 (print), 1557-9654 (electronic).

Wang:1995:WWC

[Wan95] J. Zhicheng Wang. Winning-weighted competitive learning for image compression. In Storer and Cohn [SC95], page 438. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=515548. IEEE catalog

number 95TH8037. IEEE Computer Society Press order number PR07010.

Wang:2006:RIC

[Wan06] Qing Wang. Research of image compression using vector quantization applied in the domain of data security. M.S., Tsinghua University, Tsinghua, Peoples Republic of China, 2006. URL

http://search.proquest.com/docview/1026939692.

Wang:2008:TSN

[Wan08] Bin Wang. Taxation system network construction and data compression technology research. M.Eng., Tianjin University, Tianjin, Peoples Republic of China, 2008. 142 pp. URL http:

//search.proquest.com/docview/1027902661.

Warmuth:1997:SCL

[War97] M. Warmuth. Sample compression, learnability, and the Vapnik–Chervonenkis dimension. Lecture Notes in Computer Science, 1208:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Watson:1993:VOD

[Wat93] A. B. Watson. Visually optimal DCT quantization matrices for individual images. In Storer and Cohn [SC93], pages 178– 187. ISBN 0-8186-3391-3 (microfiche), 0-8186-3392-1 (casebound). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 1993. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=253132. IEEE Computer

Society Press order number 3392-02. IEEE catalog number 93TH0536-3.

Watson:1994:ICU

[Wat94] Andrew B. Watson. Image compression using the discrete cosine transform. Mathematica Journal, 4(1):81–88, Winter 1994. CODEN ???? ISSN 1047-5974 (print), 1097-

1610 (electronic). URL http://www.mathematica-journal.
com/issue/v4i1/article/81-88_Watson.mj.pdf;   http:
//www.mathematica-journal.com/issue/v4i1/article/index.
html.

WavPack:2006:WHP

[Wav06] WavPack. WavPack home page: Hybrid lossless audio com-

pression, 2006. URL http://www.wavpack.com/.

Wayner:2009:DCI

[Way09] Peter Wayner. Disappearing cryptography: information hiding: steganography and watermarking. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, third edition, 2009. ISBN 0-12-374479-2, 0-08-092270-8 (e-book). xv + 439 pp. LCCN

TK5105.59 .W39 2009. URL http://www.sciencedirect.
com/science/book/9780123744791.

Wu:1982:ABA

[WB82] J. K. Wu and R. E. Burge. Adaptive bit allocation for image compression. Computer Graphics and Image Processing, 19(4): 392–400, August 1982. CODEN CGIPBG. ISSN 0146-664X.

Witten:1991:ZFP

[WB91] Ian H. Witten and Timothy C. Bell. The zero-frequency problem: Estimating the probabilities of novel events in adaptive text compression. IEEE Transactions on Information Theory, IT-37(4):1085–1094, 1991. CODEN IETTAW. ISSN 0018-9448 (print), 1557-9654 (electronic).

Wilson:2000:CDT

[WB00] B. Wilson and M. A. Bayoumi. Compressed domain texture classification from a modified EZW symbol stream. In Storer and Cohn [SC00], page ?? ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2000. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=838226. This

millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Wu:2003:CHD

[WB03] Xintao Wu and D. Barbara. Compressing high dimensional datasets by fractals. In Storer and Cohn [SC03], page ?? ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2003. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=1194071. IEEE

Computer Society Order number PR01896.

Wang:2017:SNP

[WB17] Changda Wang and Elisa Bertino. Sensor network provenance compression using dynamic Bayesian networks. ACM Transactions on Sensor Networks, 13(1):5:1–5:??, February 2017. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Witten:1994:TIC

[WBE^{+}94] Ian H. Witten, T. C. Bell, H. Emberson, S. Inglis, and A. Moffat. Textual image compression: two-stage lossy/lossless encoding of textual images. Proceedings of the IEEE, 82(6):878– 888, June 1994. CODEN IEEPAD. ISSN 0018-9219 (print), 1558-2256 (electronic).

Witten:1992:TIC

[WBH^{+}92] Ian H. Witten, T. C. Bell, M.-E. E. Harrison, M. L. James, and A. Moffat. Textual image compression. In Storer and Cohn [SC92], pages 42–51. ISBN 0-8186-2717-

4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=227477. IEEE catalog number 91TH0436-6.

Witten:1994:SGM

[WBM^{+}94] Ian H. Witten, Timothy C. Bell, Alistair Moffat, Craig G. Nevill-Manning, Tony C. Smith, and Harold Thimbleby. Semantic and generative models for lossy text compression. The Computer Journal, 37(2):83–87, ???? 1994. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/37/2/
83.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/Volume_37/Issue_02/Vol37_02.body.html#AbstractWitten.

Wu:2003:EJS

[WBM03] Zhenyu Wu, A. Bilgin, and M. W. Marcellin. An efficient joint source-channel rate allocation scheme for JPEG2000 codestreams. In Storer and Cohn [SC03], pages 113–122. ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2003. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=1194002. IEEE

Computer Society Order number PR01896.

Witten:1999:TMN

[WBMT99] I. H. Witten, Z. Bray, M. Mahoui, and B. Teahan. Text mining: a new frontier for lossless compression. In Storer and Cohn [SC99], pages 198–207. ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1999. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=755669. IEEE

Computer Society Order Number PR00096. IEEE Order Plan Catalog Number PR00096.

Witten:1991:MCF

[WBN91] I. H. Witten, T. C. Bell, and C. G. Nevill. Models for compression in full-text retrieval systems. In Storer and Reif [SR91b], pages 23–32. ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1990. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=213370. IEEE

Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Wang:2005:LHI

[WBS05] Hongqiang Wang, S. D. Babacan, and K. Sayood. Lossless hyperspectral image compression using context-based conditional averages. In Storer and Cohn [SC05], pages 418–426. ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1402203.

Witten:1988:PAA

[WC88] Ian H. Witten and John G. Cleary. On the privacy afforded by adaptive text compression. Computers & Security, 7(4):397– 408, August 1988. CODEN CPSEDU. ISSN 0167-4048 (print),

1872-6208 (electronic). URL https://www.sciencedirect.
com/science/article/pii/0167404888905809.

Wei:2002:RSV

[WC02a] Ming Wei and Hongyang Chao. Rate scalable video compression based on flexible block wavelet coding technique. In Storer and Cohn [SC02], page ?? ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2002. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1000021. IEEE Computer Society Order Number PR01477.

Wu:2002:DRT

[WC02b] Hsien-Chu Wu and Chin-Chen Chang. Detection and restoration of tampered JPEG compressed images. The Journal of Systems and Software, 64(2):151–161, November 15, 2002. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wu:1997:CHF

[WCB97] Xiaolin Wu, Wai Kin Choi, and P. Bao. L_{∞}-constrained high-fidelity image compression via adaptive context modeling. In Storer and Cohn [SC97], pages 91–100. ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1997. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=581978. IEEE Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Wang:2003:DRT

[WCH03] Ching-Te Wang, Tung-Shou Chen, and Shao-Hau He. Detecting and restoring the tampered images based on iteration-free fractal compression. The Journal of Systems and Software, 67 (2):131–140, August 15, 2003. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wu:1998:LII

[WCM98] X. Wu, Wai-Kin Choi, and N. Memon. Lossless interframe image compression via context modeling. In Storer and Cohn [SC98], pages 378–387. ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1998.

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=672169. IEEE Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Wen:2010:RSB

[WCY^{+}10] Jiangtao Wen, Zhuoyuan Chen, Shiqiang Yang, Yuxing Han, and J. D. Villasenor. Reconstruction of sparse binary signals using compressive sensing. In Storer and Marcellin [SM10b], page 556. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=5453533.

Wu:2002:OMR

[WD02] Xiaolin Wu and S. Dumitrescu. On optimal multi-resolution scalar quantization. In Storer and Cohn [SC02], pages 322– 331. ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2002. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=999970. IEEE

Computer Society Order Number PR01477.

Wu:2011:SAS

[WDR11] Hongzhi Wu, Julie Dorsey, and Holly Rushmeier. Surface appearance: a sparse parametric mixture model for BTF compression, editing and rendering. Computer Graphics Forum, 30 (2):465–473, April 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wegener:1992:RDC

[Weg92] Al Wegener. Review of Data Compression, 3rd edn. (Held, G; 1991). Computer, 25(8):110–111, August 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wegener:2007:SRT

[Weg07] A. Wegener. Samplify: Real-time compression for electronic measurements. In Storer and Cohn [SC07], page 407. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4148808. IEEE Computer Society

Order Number P2791.

Weiss:1990:SRI

[Wei90] Peter Weiss. Size reduction of inverted files using data compression and data structure reorganization. D.Sc. thesis, The George Washington University, Washington, DC, USA,

1990. 174 pp. URL http://search.proquest.com/docview/
303869016.

Weisstein:2006:US

[Wei06] Eric W. Weisstein. Ulam sequence, 2006. URL http://

mathworld.wolfram.com/UlamSequence.html. From Math-

World — a Wolfram Web resource.

Weisstein:2007:RNP

[Wei07] Eric W. Weisstein. Real number picking, 2007. URL http:/

/mathworld.wolfram.com/RealNumberPicking.html. From

MathWorld — a Wolfram Web resource.

Weinstein:2015:ICQ

[Wei15] O. Weinstein. Information complexity and the quest for interactive compression. ACM SIGACT News, 46(2):41–64, June 2015. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Wells:1972:FCU

[Wel72] M. Wells. File compression using variable length encodings. The Computer Journal, 15(4):308–313, November 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/15/4/308.full.pdf+html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_04/

150308.sgm.abs.html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_04/tiff/308.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_
04/tiff/309.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_04/tiff/310.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_
04/tiff/311.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_04/tiff/312.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_

04/tiff/313.tif. See correspondence [GW73].

Welch:1984:HSD

[Wel84a] Terry A. Welch. High speed data compression and decompression apparatus and method. US Patent 4,558,302., August 7,

1984. URL http://www.google.com/patents/US455830. US

Patent Application US19830505638 filed 20 June 1983. See also the related patent [ELZC84] issued on the same day.

Welch:1984:THP

[Wel84b] Terry A. Welch. A technique for high-performance data compression. Computer, 17(6):8–19, June 1984. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Welstead:1999:FWI

[Wel99] Stephen Welstead. Fractal and Wavelet Image Compression Techniques. SPIE Optical Engineering Press, Bellingham, WA, USA, 1999. ISBN 0-8194-3503-1 (print), 0-8194-7859-8 (e-PDF), 1-61583-713-2 (e-book). xv + 232 pp. LCCN TA1637 .W45 1999.

Wu:1992:LIC

[WF92] X. Wu and Yonggang Fang. Lossless interframe compression of medical images. In Storer and Cohn [SC92], pages 249– 258. ISBN 0-8186-2717-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=227456. IEEE catalog

number 91TH0436-6.

Wu:1993:SBP

[WF93] X. Wu and Y. Fang. Segmentation-based progressive image coding. In Storer and Cohn [SC93], pages 262–271. ISBN 0-8186-3391-3 (microfiche), 0-8186-3392-1 (casebound). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 1993. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=253123. IEEE Computer

Society Press order number 3392-02. IEEE catalog number 93TH0536-3.

Wu:1994:FBS

[WF94] X. Wu and Y. G. Fang. Fast bintree-structured image coder for high subjective quality. In Storer and Cohn [SC94], pages 284–293. ISBN 0-8186-5636-0, 0-8186-5637-9 (case). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1994. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=305936. IEEE

Catalog Number 93TH0626-2. IEEE Computer Society Press Order Number 5637-02.

Wu:1995:NAO

[WF95] Xiaolin Wu and Yonggang Fang. New algorithms for optimal binary vector quantizer design. In Storer and Cohn [SC95], pages 132–141. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

515503. IEEE catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Wiseman:2001:CEC

[WF01] Yair Wiseman and Erick Fredj. Contour extraction of compressed JPEG images. Journal of Graphics Tools: JGT, 6(3): 37–43, 2001. CODEN JGTOFD. ISSN 1086-7651. URL http:

//www.acm.org/jgt/papers/WisemanFredj01/.

Wu:2003:EVA

[WF03] Yung-Gi Wu and Kuo-Lun Fan. An efficient VQ algorithm using mean value predictive. In Storer and Cohn [SC03], page ?? ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2003. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1194073. IEEE Computer Society Order number PR01896.

Wickerhauser:1994:ECW

[WFG^{+}94] Mladen Victor Wickerhauser, Marie Farge, Eric Goirand, Eva Wesfreid, and Echeyde Cubillo. Efficiency comparison of wavelet packet and adapted local cosine bases for compression of a two-dimensional turbulent flow. In Chui et al. [CMP94].

ISBN 0-12-174575-9. URL ftp://wuarchive.wustl.edu/
doc/techreports/wustl.edu/math/papers/taormina2.ps.
Z.

Wu:2009:ARD

[WFLZ09] Feng Wu, Jingjing Fu, Zhouchen Lin, and Bing Zeng. Analysis on rate-distortion performance of compressive sensing for binary sparse source. In Storer and Marcellin [SM09], pages 113–122. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=4976455. IEEE

Computer Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

Wesel:1994:BRW

[WG94] R. D. Wesel and R. M. Gray. Bayes risk weighted VQ and learning VQ. In Storer and Cohn [SC94], pages 400–409. ISBN 0-8186-5636-0, 0-8186-5637-9 (case). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1994.

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=305948. IEEE Catalog Number 93TH0626-2. IEEE Computer Society Press Order Number 5637-02.

Ware:2000:TLT

[WG00] F. W. Ware and J. D. Gibson. A three-layer, two description image coder. In Storer and Cohn [SC00], page ?? ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2000. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=838224. This millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Wegman:1988:CSS

[WGM88] Edward J. Wegman, Donald T. Gantz, and John J. Miller, editors. Computing Science and Statistics Proceedings of the 20th Symposium on the Interface Fairfax, Virginia, April 1988. American Statistical Association, Alexandria, VA, USA,

1988. URL http://www.dtic.mil/dtic/tr/fulltext/u2/
a208838.pdf.

Wang:2018:OQA

[WGZ^{+}18] Shiqi Wang, Ke Gu, Kai Zeng, Zhou Wang, and Weisi Lin. Objective quality assessment and perceptual compression of screen content images. IEEE Computer Graphics and Applications, 38(1):47–58, January/February 2018. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (elec-

tronic). URL https://www.computer.org/csdl/mags/cg/
2018/01/mcg2018010047-abs.html.

Wei:2014:NEA

[WGZW14] Qingting Wei, Jihong Guan, Shuigeng Zhou, and Xin Wang. A new and effective approach to GML documents compression. The Computer Journal, 57(11):1723–1740, November 2014. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/57/11/1723.

Wang:2016:DBS

[WHB16] Changda Wang, Syed Rafiul Hussain, and Elisa Bertino. Dictionary based secure provenance compression for wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 27(2):405–418, February 2016. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL http://www.computer.org/csdl/trans/td/
2016/02/07038199-abs.html.

Wheeler:1997:UBM

[Whe97] David Wheeler. Upgrading bred with multiple tables. Report, The Computer Laboratory, Cambridge University, Cam-

bridge, UK, April 1997. URL http://ftp.cl.cam.ac.uk/

users/djw3/bred3.ps. This is the third of four key papers behind the bzip2 compression tools. The others are [HL90, BW94b, SB97a].

Wagner:1999:VCB

[WHH^{+}99] M. Wagner, R. Herz, H. Hartenstein, R. Hamzaoui, and D. Saupe. A video codec based on R/D-optimized adaptive vector quantization. In Storer and Cohn [SC99], page ?? ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1999. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=785713. IEEE Computer Society Or-

der Number PR00096. IEEE Order Plan Catalog Number PR00096.

White:1987:CID

[Whi87] Ronald G. White. Compressing image data with quadtrees. Dr. Dobb’s Journal of Software Tools, 12(3):16–??, March 1987. CODEN DDJOEB. ISSN 0888-3076.

Whittle:2006:RCA

[Whi06] Robin Whittle. Rice coding, AKA Rice packing, Elias Gamma codes and other approaches, 2006. URL http://www.

firstpr.com.au/audiocomp/lossless/#rice.

Wang:2010:SSV

[WHW10] Qijun Wang, Ruimin Hu, and Zhongyuan Wang. Spatially scalable video coding based on hybrid epitomic resizing. In Storer and Marcellin [SM10b], pages 139–148. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=5453456.

Wang:2012:NFS

[WHZ12] Hui Wang, Anthony T. S. Ho, and Xi Zhao. A novel fast self-restoration semi-fragile watermarking algorithm for image content authentication resistant to JPEG compression. Lecture Notes in Computer Science, 7128:72–85, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32205-1_8/.

Wickerhauser:1990:PCB

[Wic90] Mladen Victor Wickerhauser. Picture compression by bestbasis sub-band coding. Preprint, Yale University, 1990. URL

ftp://math.yale.edu/pub/wavelets/pic.tar.

Wickerhauser:1992:HRS

[Wic92] Mladen Victor Wickerhauser. High-resolution still picture compression. Digital Signal Processing: a Review Journal, 2 (4):204–226, October 1992. URL ftp://wuarchive.wustl.

edu/doc/techreports/wustl.edu/math/papers/dsp.ps.Z.

Wickerhauser:1994:CPC

[Wic94] Mladen Victor Wickerhauser. Comparison of picture compression methods: Wavelet, wavelet packet, and local cosine transform coding. In Chui et al. [CMP94]. ISBN 0-12-174575-9.

Wick:2003:OOR

[Wic03] Michael R. Wick. An object-oriented refactoring of Huffman encoding using the Java collections framework. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):283–287, January 2003. CODEN SIGSD3. ISBN 1-58113-648-X. ISSN 0097-8418 (print), 2331-

3927 (electronic). URL https://www.math.utah.edu/pub/
mirrors/ftp.ira.uka.de/bibliography/Compiler/java.
bib;   https://www.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2003.bib.

Wikipedia:2003:NSS

[Wik03] Wikipedia. Nyquist–Shannon sampling theorem, 2003.

URL http://www.wikipedia.org/wiki/Nyquist-Shannon_
sampling_theorem.

Wilkins:1970:SDC

[Wil70] Larry Clarke Wilkins. Studies on Data Compression. Part I: Picture Coding By Contours. Part II: Error Analysis of Run-Length Codes. Ph.D. thesis, Purdue University, West Lafayette, IN, USA, 1970. 318 pp. URL http://search.

proquest.com/docview/302522243.

Willems:1989:UDC

[Wil89] F. M. J. Willems. Universal data compression and repetition times. IEEE Transactions on Information Theory, IT-35(1): 54–58, January 1989. CODEN IETTAW. ISSN 0018-9448 (print), 1557-9654 (electronic).

Williams:1991:ADC

[Wil91a] Ross N. Williams. Adaptive Data Compression. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1991. ?? pp.

Williams:1991:EFZ

[Wil91b] Ross N. Williams. An extremely fast Ziv–Lempel data compression algorithm. In Storer and Reif [SR91b], pages 362–371. ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1990. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=213344. IEEE

Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Williams:1993:PGC

[Wil93] Ross N. Williams. A painless guide to CRC error detection al-

gorithms, 1993. URL http://ross.net/crc/download/crc_
v3.txt.

Wilton:1996:AFI

[Wil96] Andrew Philip Wilton. Applications of fractals to image data compression. Ph.D. thesis, Aston University, Birmingham,

UK, 1996. ???? pp. URL http://search.proquest.com/
pqdtft/docview/304336659. Report C519664 (??).

Wilson:2000:WBL

[Wil00] J. P. Wilson. Wavelet-based lossy compression of turbulence data. In Storer and Cohn [SC00], page ?? ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2000. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=838225. This millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Wilson:2002:WBL

[Wil02] J. P. Wilson. Wavelet-based lossy compression of barotropic turbulence simulation data. In Storer and Cohn [SC02], page ?? ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2002. URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

1000022. IEEE Computer Society Order Number PR01477.

Williams:2006:CP

[Wil06] Ross N. Williams. Compression patents, 2006. URL http://

www.ross.net/compression/patents.html.

WinAce:2003:WHP

[Win03] WinAce. WinAce home page, 2003. URL http://www.

winace.com/.

Wintner:2006:BBS

[Win06] R. Wintner. Bits on the big screen. IEEE Spectrum, 43(12): 42–48, December 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wirth:1976:ADS

[Wir76] Niklaus Wirth. Algorithms + Data Structures = Programs. Prentice-Hall Series in Automatic Computation. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 1976. ISBN 0-13-022418-9. xvii + 366 pp. LCCN QA76.6 .W561.

Withers:1997:ECR

[Wit97] D. Withers. The ELS-coder: a rapid entropy coder. In Storer and Cohn [SC97], page ?? ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1997.

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=582144. IEEE Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Witzgall:2008:PMA

[Wit08] H. E. Witzgall. A parametric modeling approach to image compression. In Storer and Marcellin [SM08], page 552. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33 INTER-

NET. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=4483379.

Wu:2005:JSC

[WJBM05] Zhenyu Wu, R. Jandhyala, A. Bilgin, and M. W. Marcellin. Joint source/channel coding for multiple video sequences with JPEG2000. In Storer and Cohn [SC05], page ?? ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1402246.

Wang:2010:RBR

[WJDZ10] Qifei Wang, Xiangyang Ji, Qionghai Dai, and Naiyao Zhang. Region based rate-distortion analysis for 3D video coding. In Storer and Marcellin [SM10b], page 555. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=5453530.

Wong:1992:CCT

[WK92] P. W. Wong and J. Koplowitz. Chain codes and their linear reconstruction filters. IEEE Transactions on Information Theory, IT-38(2):268–280, May 1992. CODEN IETTAW. ISSN 0018-9448 (print), 1557-9654 (electronic).

Wong:1993:FWT

[WK93] Kwo-Jyr Wong and C. C. Jay Kuo. A full wavelet transform (FWT) approach to image compression. Image and Video Processing, 1903:153–164, 1993. ISSN 1088-0356.

Wallach:1994:AMC

[WKC94] Dan S. Wallach, Sharma Kunapalli, and Michael F. Cohen. Accelerated MPEG compression of dynamic polygonal scenes. Computer Graphics, 28(Annual Conference Series): 193–196, July 1994. CODEN CGRADI, CPGPBZ. ISSN 0097-8930 (print), 1558-4569 (electronic). URL http://www.

acm.org:80/pubs/citations/proceedings/graph/192161/
p193-wallach/.

Wright:1998:CSE

[WKM^{+}98] R. G. Wright, E. Keenan, M. Mannucci, M. Rajan, T. Hanaratty, and J. Dummer. Control structure efficiency enhancement for predictive video coding. In Storer and Cohn [SC98], page ?? ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1998. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=672324. IEEE

Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Wang:1999:NDP

[WL99] N.-T. Wang and N. Ling. A novel dual-path architecture for HDTV video decoding. In Storer and Cohn [SC99], page ?? ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1999. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=785714. IEEE Computer Society Or-

der Number PR00096. IEEE Order Plan Catalog Number PR00096.

Wang:2008:FBP

[WL08] Jing Wang and Jie Liang. Filter banks for predictioncompensated multiple description coding. In Storer and Marcellin [SM08], pages 392–401. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33 INTERNET. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=4483317.

Wandelt:2013:FRC

[WL13] Sebastian Wandelt and Ulf Leser. FRESCO: Referential compression of highly similar sequences. IEEE/ACM Transactions on Computational Biology and Bioinformatics, 10(5): 1275–1288, September 2013. CODEN ITCBCY. ISSN 1545-5963 (print), 1557-9964 (electronic).

Wandelt:2015:MCS

[WL15] Sebastian Wandelt and Ulf Leser. MRCSI: compressing and searching string collections with multiple references. Proceedings of the VLDB Endowment, 8(5):461–472, January 2015. CODEN ???? ISSN 2150-8097.

Wang:2015:CSP

[WLCZ15] Chao Wang, Xi Li, Peng Chen, and Xuehai Zhou. A case study of parallel JPEG encoding on an FPGA. Journal of Parallel and Distributed Computing, 78(??):1–5, April 2015. CODEN JPDCER. ISSN 0743-7315 (print), 1096-

0848 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0743731514001737.

Wang:2017:MIL

[WLH^{+}17] Jianguo Wang, Chunbin Lin, Ruining He, Moojin Chae, Yannis Papakonstantinou, and Steven Swanson. MILC: inverted list compression in memory. Proceedings of the VLDB Endowment, 10(8):853–864, April 2017. CODEN ???? ISSN 2150-8097.

Wang:2006:APH

[WLL06] Meiqing Wang, Rong Liu, and Choi-Hong Lai. Adaptive partition and hybrid method in fractal video compression. Computers and Mathematics with Applications, 51(11):1715–1726, June 2006. CODEN CMAPDK. ISSN 0898-1221 (print),

1873-7668 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0898122106001386.

Wang:2020:CRM

[WLL^{+}20] J. Wang, T. Liu, Q. Liu, X. He, H. Luo, and W. He. Compression ratio modeling and estimation across error bounds for lossy compression. IEEE Transactions on Parallel and Distributed Systems, 31(7):1621–1635, July 2020. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wu:2006:ERL

[WLS06] Yonghui Wu, S. Lonardi, and W. Szpankowski. Error-resilient LZW data compression. In Storer and Cohn [SC06], pages 193– 202. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=1607254.

Weinberger:1992:CDG

[WLZ92] M. J. Weinberger, A. Lempel, and J. Ziv. On the coding delay of a general coder. In Storer and Cohn [SC92], pages 102– 111. ISBN 0-8186-2717-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=227471. IEEE catalog

number 91TH0436-6.

Wang:2004:DCS

[WLZW04] Ze Wang, Chi-Sing Leung, Yi-Sheng Zhu, and Tien-Tsin Wong. Data compression with spherical wavelets and wavelets for the image-based relighting. Computer Vision and Image Understanding: CVIU, 96(3):327–344, December 2004. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Woolley:1994:RDP

[WM94] Stuart J. Woolley and Donald M. Monro. Rate-distortion performance of fractal transforms for image compression. Fractals, 2(3 (or 6??)):395–398, ???? 1994. CODEN FRACEG. ISSN

0218-348X. URL http://dmsun2.bath.ac.uk:8080/people/
papers/montreal.ps.gz.

Wirth:2001:CWD

[WM01] A. I. Wirth and A. Moffat. Can we do without ranks in Burrows–Wheeler transform compression? In Storer and Cohn [SC01c], pages 419–428. ISBN 0-7695-1031-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2001. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=917173. IEEE CSP number 01PR1031.

Wang:2005:VLF

[WM05] Jue Wang and Cohen MF. Very low frame-rate video streaming for face-to-face teleconference. In Storer and Cohn [SC05], pages 309–318. ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=
1402192.

Witten:1994:MGC

[WMB94] Ian H. Witten, Alistair Moffat, and Timothy C. Bell. Managing Gigabytes: Compressing and Indexing Documents and Images. Van Nostrand Reinhold, New York, NY, USA, 1994. ISBN 0-442-01863-0. xiv + 429 pp. LCCN TA1637 .W58 1994. US$54.95. The software for full-text indexing described in this book, and errata for the book, are available for anonymous ftp

from munnari.oz.au in the directory /pub/mg.

Witten:1999:MGC

[WMB99] Ian H. Witten, Alistair Moffat, and Timothy C. Bell. Managing Gigabytes: Compressing and Indexing Documents and Images. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, second edition, 1999. ISBN 1-55860-570-3. xxxi + 519 pp. LCCN TA1637 .W58 1994. US$54.95. URL ftp:

//ftp.math.utah.edu/pub/mg/;   ftp://ftp.math.utah.
edu/pub/mg/mg-1.3x/bibsearch-1.02.tar.gz;   ftp://
munnari.oz.au:/pub/mg; http://www.cs.mu.oz.au/mg/;
 http://www.cs.mu.oz.au/~alistair/arith_coder/;   https:
//www.math.utah.edu/pub/mg/;   https://www.math.utah.
edu/pub/mg/mg-1.3x/bibsearch-1.02.tar.gz;   http://
www.mkp.com/books_catalog/1-55860-570-3.asp.

Wong:1996:OPP

[WMH96] Ping Wah Wong, N. Moayeri, and C. Herley. Optimum pre- and post-filters for robust scalar quantization. In Storer and Cohn [SC96], pages 240–249. ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1996. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=488329. IEEE Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Wendt:2012:ECR

[WMNP12] J. B. Wendt, S. Meguerdichian, H. Noshadi, and M. Potkonjak. Energy and cost reduction in localized multisensory systems through application-driven compression. In Storer and Marcellin [SM12], page 411. ISBN 1-4673-0715-7 (paperback). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=6189292. IEEE Computer Society order number P4656.

Wu:2010:CPB

[WMWW10] Xiaolin Wu, H. D. Mittelmann, Xiaohan Wang, and J. Wang. On computation of performance bounds of optimal index assignment. In Storer and Marcellin [SM10b], pages 189–198. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=5453463.

Wu:2010:PDB

[WN10] Hao Wu and Frank Noé. Probability distance based compression of hidden Markov models. Multiscale Modeling & Simulation, 8(5):1838–1861, ???? 2010. CODEN MMSUBT. ISSN 1540-3459 (print), 1540-3467 (electronic). URL http:

//epubs.siam.org/mms/resource/1/mmsubt/v8/i5/p1838_
s1.

Witten:1987:ACD

[WNC87] Ian H. Witten, Radford M. Neal, and John G. Cleary. Arithmetic coding for data compression. Communications of the ACM, 30(6):520–540, June 1987. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/0001-0782/214771.html.

Weinberger:2000:LDM

[WO00] M. J. Weinberger and E. Ordentlich. On-line decision making for a class of loss functions via Lempel–Ziv parsing. In Storer and Cohn [SC00], pages 163–172. ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2000. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=838156. This millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Wang:2001:DTC

[WO01] Xin Wang and M. T. Orchard. Design of trellis codes for source coding with side information at the decoder. In Storer and Cohn [SC01c], pages 361–370. ISBN 0-7695-1031-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 2001. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=917167. IEEE CSP num-

ber 01PR1031.

Wolff:1999:CIC

[Wol99a] G. Wolff. Computing as information compression by multiple alignment, unification and search. J.UCS: The Journal of Universal Computer Science, 5(11):777–815, November 28, 1999. CODEN ???? ISSN 0948-695X (print), 0948-6968 (elec-

tronic). URL http://www.jucs.org/jucs_5_11/computing_
as_information_compression.

Wolff:1999:U

[Wol99b] Gerry Wolff. [unknown], 1999. URL http://www.cognitionresearch.

org.uk/sp.htm.

Wolff:1999:PRI

[Wol99c] J. G. Wolff. Probabilistic reasoning as information compression by multiple alignment, unification and search: An introduction and overview. J.UCS: The Journal of Universal Computer Science, 5(7):418–462, July 28, 1999. CODEN ???? ISSN 0948-695X (print), 0948-6968 (electronic). URL http:

//www.jucs.org/jucs_5_7/probabilistic_reasoning_as_
information.

Wong:1985:FSCb

[Won85] Willard Lee Wong. Font storage, compression, and scaling for networked professional workstations. Report UIUCDCS-R-85-1201, Dept. of Computer Science, University of Illinois at Urbana-Champaign, Urbana, IL, USA, April 1985. v + 80 pp.

Wong:1991:MCV

[Won91] C. W. Wong. Motion-compensated video image compression using luminance and chrominance components for motion estimation. In Storer and Reif [SR91b], page ?? ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1990. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=213321. IEEE Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Wood:1994:GDC

[Woo94] Christopher M. Wood. General data compression algorithm for space images using fractal techniques. Proceedings of the SPIE — The International Society for Optical Engineering, 2198(??):1336–1341, ???? 1994. CODEN PSISDG. ISSN ????

Woods:1996:RPD

[Woo96] J. Woods. RFC 1979: PPP Deflate Protocol, August

1996. URL ftp://ftp.internic.net/rfc/rfc1979.txt;
 http://www.faqs.org/rfcs/rfc1979.html;   https://
www.math.utah.edu/pub/rfc/rfc1979.txt. Status: INFOR-

MATIONAL.

Woods:2000:CPT

[Woo00] John Charles Woods. Compression and packetised transport of model-based video data. Ph.D. thesis, University of Essex, Southend-on-Sea, Essex, UK, 2000. ???? pp. URL http://

search.proquest.com/docview/304662752.

Wu:2006:OBI

[WOS06] Kesheng Wu, Ekow J. Otoo, and Arie Shoshani. Optimizing bitmap indices with efficient compression. ACM Transactions on Database Systems, 31(1):1–38, March 2006. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Wharton:2008:SEC

[WPA08] E. Wharton, K. Panetta, and S. Agaian. Simultaneous encryption/compression of images using alpha rooting. In Storer and Marcellin [SM08], page 551. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33 INTERNET. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=4483378.

Wu:2009:HCI

[WPXL09] Feng Wu, Xiulian Peng, Jizheng Xu, and Shipeng Li. How can intra correlation be exploited better? In Storer and Marcellin [SM09], page 468. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ????

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=4976522. IEEE Computer Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

Wakin:2002:ICU

[WRCB02] M. Wakin, J. Romberg, Hyeokho Choi, and R. Baraniuk. Image compression using an efficient edge cartoon + texture model. In Storer and Cohn [SC02], pages 43–52. ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2002. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=999942. IEEE

Computer Society Order Number PR01477.

Wright:1939:GSW

[Wri39] Ernest Vincent Wright. Gadsby; a story of over 50,000 words without using the letter “E”. Wetzel Publishing Co., Inc., Los Angeles, CA, USA, 1939. 267 pp. URL http://gadsby. hypermart.net/. Reprinted by University Microfilms, Ann Arbor, MI, 1991.

Whyte:1991:DCF

[WS91] Wayne A. Whyte and Khalid Sayood. Data compression for full motion video transmission. NASA technical memorandum 105239, NASA, Washington, DC, USA, 1991. ???? pp. Shipping list number 92-0647-M.

Weiss:1993:PDD

[WS93] J. Weiss and D. Schremp. Putting data on a diet. IEEE Spectrum, 30(8):36–39, August 1993. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wagner:2000:ROH

[WS00] M. Wagner and D. Saupe. RD-optimization of hierarchical structured adaptive vector quantization for video coding. In Storer and Cohn [SC00], page ?? ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2000. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=838223. This millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Wang:2005:BBU

[WS05] Li Wang and G. I. Shamir. BWT based universal lossless source controlled channel decoding with low density parity check codes. In Storer and Cohn [SC05], page ?? ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1402244.

Wernersson:2006:MDC

[WS06] N. Wernersson and M. Skoglund. Multiple description coding using rotated permutation codes. In Storer and Cohn [SC06], page 471. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=1607314.

Wu:2008:ICV

[WS08] Feng Wu and Xiaoyan Sun. Image compression by Visual Pattern Vector Quantization (VPVQ). In Storer and Marcellin [SM08], pages 123–131. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008;

QA76.9.D33 INTERNET. URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=4483290.

Wennersten:2009:TVT

[WS09] Per Wennersten and Jacob Ström. Textures and video: Tablebased Alpha compression. Computer Graphics Forum, 28(2): 687–695, April 2009. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2012:ACA

[WS12] Xiaoyun Wang and Kazue Sako, editors. Advances in Cryptology — ASIACRYPT 2012: 18th International Conference on the Theory and Application of Cryptology and Information Security, Beijing, China, December 2–6, 2012. Proceedings, volume 7658 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. CODEN LNCSD9. ISBN 3-642-34960-9 (print), 3-642-34961-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). 330 (est.) pp. LCCN ???? URL http://

www.springerlink.com/content/978-3-642-34961-4.

Watson:1994:VDB

[WSA94] A. B. Watson, J. A. Solomon, and A. J. Ahumada, Jr. Visibility of DCT basis functions: effects of display resolution. In Storer and Cohn [SC94], pages 371–379. ISBN 0-8186-5636-0, 0-8186-5637-9 (case). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1994. URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

305945. IEEE Catalog Number 93TH0626-2. IEEE Computer Society Press Order Number 5637-02.

Wang:2011:LGT

[WSDTO11] Ching Ming Wang, J. Shol-Dickstein, I. Tošić, and B. A. Olshausen. Lie group transformation models for predictive video coding. In Storer and Marcellin [SM11b], pages 83–92. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5749466. IEEE Computer Society

Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Wang:2011:SRB

[WSK^{+}11] Jin Wang, Yunhui Shi, Dehui Kong, Wenpeng Ding, Chunjing Li, and Baocai Yin. Sparse representation based downsampling image compression. Journal of Computational and Applied Mathematics, 236(5):675–683, October 1, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S037704271100361X.

Wang:1994:OCV

[WSS94] X. Wang, S. M. Shende, and K. Sayood. Online compression of video sequences using adaptive VQ codebooks. In Storer and Cohn [SC94], pages 185–194. ISBN 0-8186-5636-0, 0-8186-5637-9 (case). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1994. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=305926. IEEE

Catalog Number 93TH0626-2. IEEE Computer Society Press Order Number 5637-02.

Weinberger:1996:LLC

[WSS96] M. J. Weinberger, G. Seroussi, and G. Sapiro. LOCO-I: a low complexity, context-based, lossless image compression algorithm. In Storer and Cohn [SC96], pages 140–149. ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1996. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=488319. IEEE Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Weinberger:2000:LLI

[WSS00] M. J. Weinberger, G. Seroussi, and G. Sapiro. The LOCO-I lossless image compression algorithm: Principles and standardization into JPEG-LS. IEEE Transactions on Image Processing, 9(8):1309–1324, August 2000. CODEN IIPRE4. ISSN 1057-7149 (print), 1941-0042 (electronic).

Willems:1995:CTW

[WST95] F. M. J. Willems, Y. M. Shtarkov, and Tj. J. Tjalkens. The context-tree weighting method: Basic properties. IEEE Transactions on Information Theory, IT-41:653–664, May 1995. CODEN IETTAW. ISSN 0018-9448 (print), 1557-9654 (electronic).

Wu:2005:MCT

[WT05] Peiliang Wu and W. J. Teahan. Modelling Chinese for text compression. In Storer and Cohn [SC05], page ?? ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1402245.

Wu:1992:VQD

[Wu92] Xiaolin Wu. Vector quantizer design by constrained global optimization. In Storer and Cohn [SC92], pages 132–141. ISBN 0-8186-2717-4. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=227468. IEEE catalog number

91TH0436-6.

Wu:1993:GOB

[Wu93] X. Wu. Globally optimal bit allocation. In Storer and Cohn [SC93], pages 22–31. ISBN 0-8186-3391-3 (microfiche), 0-8186-3392-1 (casebound). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1993. URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

253148. IEEE Computer Society Press order number 3392-02. IEEE catalog number 93TH0536-3.

Wu:1995:CSQ

[Wu95] Xiaolin Wu. Context selection and quantization for lossless image coding. In Storer and Cohn [SC95], page 453. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=515563. IEEE catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Wu:1996:ASL

[Wu96] Xiaolin Wu. An algorithmic study on lossless image compression. In Storer and Cohn [SC96], pages 150–159. ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1996. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=488320. IEEE Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Wu:1999:CQF

[Wu99a] X. Wu. Context quantization with Fisher discriminant for adaptive embedded wavelet image coding. In Storer and Cohn [SC99], pages 102–111. ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1999. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=755659. IEEE

Computer Society Order Number PR00096. IEEE Order Plan Catalog Number PR00096.

Wu:1999:LCH

[Wu99b] X. Wu. Low complexity high-order context modeling of embedded wavelet bit streams. In Storer and Cohn [SC99], pages 112–120. ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN QA76.9.D33 D37 1999. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=755660. IEEE

Computer Society Order Number PR00096. IEEE Order Plan Catalog Number PR00096.

Wen:1998:RVL

[WV98] Jiangtao Wen and John D. Villasenor. Reversible variable length codes for efficient and robust image and video coding. In Storer and Cohn [SC98], pages 471–480. ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1998. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=672209. IEEE Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Weidmann:1999:RDA

[WV99a] C. Weidmann and M. Vetterli. Rate-distortion analysis of spike processes. In Storer and Cohn [SC99], pages 82–91. ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1999. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=755657. IEEE Computer Society Or-

der Number PR00096. IEEE Order Plan Catalog Number PR00096.

Wen:1999:USI

[WV99b] J. Wen and J. D. Villasenor. Utilizing soft information in decoding of variable length codes. In Storer and Cohn [SC99], pages 131–139. ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1999. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=755662. IEEE

Computer Society Order Number PR00096. IEEE Order Plan Catalog Number PR00096.

Weidmann:2000:RDB

[WV00] C. Weidmann and M. Vetterli. Rate distortion behavior of threshold-based nonlinear approximations. In Storer and Cohn [SC00], pages 333–342. ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2000. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=838173. This

millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Wang:1992:VRL

[WW92] Y. Wang and J.-M. Wu. Vector run-length coding of bi-level images. In Storer and Cohn [SC92], pages 289–298. ISBN 0-8186-2717-4. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=227452. IEEE catalog number

91TH0436-6.

Wang:2010:IFV

[WW10] Jia Wang and Xiaolin Wu. Information flows in video coding. In Storer and Marcellin [SM10b], pages 149–158. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5453459.

Wang:2012:GSD

[WW12] Cong Wang and Hongzhi Wang. Graph-structured data compression based on frequent subgraph contraction. Lecture Notes in Computer Science, 7419:11–18, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/chapter/10.1007/978-
3-642-33050-6_2/.

Willis:1995:IDC

[WWA^{+}95] I. C. Willis, L. J. C. Woolliscroft, T. Averkamp, D. A. Gurnett, R. A. Johnson, D. Kirchner, W. S. Kurth, and W. Robison. The implementation of data compression in the Cassini RPWS Dedicated Compression Processor. In Storer and Cohn [SC95], page ?? ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=515601. IEEE

catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Wang:2011:PQC

[WWS11] Liangjun Wang, Xiaolin Wu, and Guangming Shi. Progressive quantization of compressive sensing measurements. In Storer and Marcellin [SM11b], pages 233–242. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=5749481. IEEE Computer Society Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Wang:2007:MSS

[WWSY07] Jia Wang, Xiaolin Wu, Jun Sun, and Songyu Yu. On multi-stage sequential coding of correlated sources. In Storer and Cohn [SC07], pages 253–262. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=4148764. IEEE Computer Society Order Number P2791.

Wu:1997:CEC

[WWW97] Xiaolin Wu, Jiang Wen, and Wing Hung Wong. Conditional entropy coding of VQ indexes for image compression. In Storer and Cohn [SC97], pages 347–356. ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1997. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=582058. IEEE Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Wu:2006:JSC

[WWW06] Xiaolin Wu, Xiaohan Wang, and Jia Wang. Joint sourcechannel decoding of multiple description quantized Markov sequences. In Storer and Cohn [SC06], pages 103–112. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1607245.

Wang:2007:WZR

[WWY^{+}07] Peng Wang, Jia Wang, Songyu Yu, Erkang Chen, and Xiaokang Yang. The Wyner–Ziv rate-distortion function of multivariate Gaussian sources and its application in distributed video coding. In Storer and Cohn [SC07], page 405. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4148806. IEEE Computer Society

Order Number P2791.

Wu:2000:OPE

[WX00] Xiaolin Wu and Zixiang Xiong. Optimal packetization of embedded bitstreams. In Storer and Cohn [SC00], page ?? ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2000. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=838227. This millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Wen:2010:FRD

[WXC^{+}10] Jiangtao Wen, Mou Xiao, Jianwen Chen, Pin Tao, and Chao Wang. Fast rate distortion optimized quantization for H.264/ AVC. In Storer and Marcellin [SM10b], page 557. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5453532.

Wang:1999:TDW

[WXCM99] A. Wang, Zixiang Xiong, P. A. Chou, and S. Mehrotra. Three-dimensional wavelet coding of video with global motion compensation. In Storer and Cohn [SC99], pages 404–413. ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1999. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=755690. IEEE Computer Society Or-

der Number PR00096. IEEE Order Plan Catalog Number PR00096.

Wu:1991:ICA

[WY91] X. Wu and C. Yao. Image coding by adaptive tree-structured segmentation. In Storer and Reif [SR91b], pages 73–82. ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1990. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=213374. IEEE Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Wu:2010:TOL

[WY10] Feng Wu and Peiwen Yu. Theoretically optimal low-density parity-check code ensemble for Gallager’s Decoding Algorithm A. In Storer and Marcellin [SM10b], page 558. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5453535.

Wang:2010:ADC

[WYM10] Chaoli Wang, Hongfeng Yu, and Kwan-Liu Ma. Applicationdriven compression for visualizing large-scale time-varying data. IEEE Computer Graphics and Applications, 30(1):59–69, January/February 2010. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Wu:1991:BTS

[WZ91a] X. Wu and K. Zhang. A better tree-structured vector quantizer. In Storer and Reif [SR91b], pages 392–401. ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1990. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=213341. IEEE Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Wyner:1991:FDB

[WZ91b] A. D. Wyner and J. Ziv. Fixed data base version of the Lempel–Ziv data compression algorithm. In Storer and Reif [SR91b], pages 202–207. ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1990. URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

213361. IEEE Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Wu:1994:SDM

[WZ94] Xiaolin Wu and Kaizhong Zhang. A subjective distortion measure for vector quantization. In Storer and Cohn [SC94], pages 22–31. ISBN 0-8186-5636-0, 0-8186-5637-9 (case). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 1994. URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=305909. IEEE Cata-

log Number 93TH0626-2. IEEE Computer Society Press Order Number 5637-02.

Weinberger:1991:OAP

[WZL91] M. J. Weinberger, J. Ziv, and A. Lempel. On the optimal asymptotic performance of universal ordering and discrimination of individual sequences. In Storer and Reif [SR91b], pages 239–246. ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1990. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=213357. IEEE

Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Wang:2017:SPD

[WZL^{+}17] Zhoukai Wang, Yinliang Zhao, Yang Liu, Zhong Chen, Cuocuo Lv, and Yuxiang Li. A speculative parallel decompression algorithm on Apache Spark. The Journal of Supercomputing, 73(9):4082–4111, September 2017. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Wang:2007:ERE

[WZMG07] Qiang Wang, Debin Zhao, Siwei Ma, and Wen Gao. An enhanced robust entropy coder for video codecs based on contextadaptive reversible VLC. In Storer and Cohn [SC07], page 406. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=4148807. IEEE Computer

Society Order Number P2791.

Wu:2009:MGA

[WZW09] Xiaolin Wu, Xiangjun Zhang, and Jia Wang. Modelguided adaptive recovery of compressive sensing. In Storer and Marcellin [SM09], pages 123–132. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=4976456. IEEE Computer Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

Wu:2011:HFI

[WZW11] Xiaolin Wu, Jiantao Zhou, and Heng Wang. High-fidelity image compression for high-throughput and energy-efficient cameras. In Storer and Marcellin [SM11b], pages 433–442. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5749501. IEEE Computer Society

Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Wei:2011:TSB

[WZY^{+}11] Hai Wei, S. Zabuawala, J. Yadegar, J. de la Cruz, and H. J. Gonzalez. Towards the synergy between compression and content-based analysis: a pattern-driven approach. In Storer and Marcellin [SM11b], page 481. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=5749538. IEEE Computer Society Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Xie:2018:QLC

[XCK18] Ting Xie, Varun Chandola, and Oliver Kennedy. Query log compression for workload analytics. Proceedings of the VLDB Endowment, 12(3):183–196, November 2018. CODEN ???? ISSN 2150-8097.

Xu:2011:FRD

[XD11] Jiayi Xu and S. Dumitrescu. Fast R-D optimal packetization of embedded bitstreams into independent source packets. In Storer and Marcellin [SM11b], pages 283–292. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5749486. IEEE Computer Society

Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Xie:2011:ULD

[XDZ11] Ningde Xie, Guiqiang Dong, and Tong Zhang. Using lossless data compression in data storage systems: Not for saving space. IEEE Transactions on Computers, 60(3):335–345, March 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xue:1997:ACI

[XG97a] Xiaohui Xue and Wen Gao. Arithmetic coding with improved solution for the carry-over problem. In Storer and Cohn [SC97], page ?? ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1997. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=582146. IEEE

Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Xue:1997:HPA

[XG97b] Xiaohui Xue and Wen Gao. High performance arithmetic coding for small alphabets. In Storer and Cohn [SC97], page ?? ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1997. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=582149. IEEE

Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Xiong:2010:TGB

[XG10] Ruiqin Xiong and Wen Gao. Tanner graph based image interpolation. In Storer and Marcellin [SM10b], pages 376–385. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=5453483.

Xu:2005:HDI

[XHS05] Lifeng Xu, M. W. Hoffman, and K. Sayood. Hard decision and iterative joint source channel coding using arithmetic codes. In Storer and Cohn [SC05], pages 203–212. ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1402181.

Xiong:2011:AID

[Xio11] Ruiqin Xiong. Adaptive image deblurring via Tanner graph representation and belief propagation. In Storer and Marcellin [SM11b], page 482. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

5749539. IEEE Computer Society Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Xie:2014:TEC

[XJ14] Ruitao Xie and Xiaohua Jia. Transmission-efficient clustering method for wireless sensor networks using compressive sensing. IEEE Transactions on Parallel and Distributed Systems, 25 (3):806–815, March 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

XMill:2003:XUC

[XMi03] XMill. XMill user-configurable XML compressor, 2003. URL

http://sourceforge.net/projects/xmill/.

XML:2003:XHP

[XML03] XML. XML home page, 2003. URL http://www.xml.com/.

Xie:2013:EHA

[XMRF^{+}13] Yulai Xie, Kiran-Kumar Muniswamy-Reddy, Dan Feng, Yan Li, and Darrell D. E. Long. Evaluation of a hybrid approach for efficient provenance storage. ACM Transactions on Storage, 9 (4):14:1–14:??, November 2013. CODEN ???? ISSN 1553-3077 (print), 1553-3093 (electronic).

Xie:2005:DNS

[XS05] Kai Xie and G. I. Shamir. Decoding of non-systematic turbo codes for stationary memoryless and piecewise stationary memoryless sequences. In Storer and Cohn [SC05], page ?? ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1402247.

Xu:2007:CBR

[XSW07] Jin Xu, Limin Sun, and Zhimei Wu. A content-based robust mode decision scheme for Internet videophone applications. In Storer and Cohn [SC07], page 408. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=4148809. IEEE Computer Society Order Number P2791.

Xu:2005:DJS

[XSX05] Qian Xu, V. Stankovic, and Zixiang Xiong. Distributed joint source-channel coding of video using Raptor codes. In Storer and Cohn [SC05], page ?? ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=1402248.

Xiao:2009:PCW

[XTH09] Jinyou Xiao, Johannes Tausch, and Yucai Hu. A-posteriori compression of wavelet-BEM matrices. Computational mechanics, 44(5):705–715, 2009. CODEN CMMEEE. ISSN 0178-7675.

Xu:2006:RDC

[Xu06] Feng Rong Xu. The research of data compression on Abis interface of GSM network. M.S., Huazhong (Central China) University of Science and Technology, Huazhong, Peoples Re-

public of China, 2006. URL http://search.proquest.com/
docview/1026570127.

Xue:1999:ICBb

[Xue99a] X. Xue. Image compression based on low-pass wavelet transform and multi-scale edge compensation. Part II: evidence and experiments. In Storer and Cohn [SC99], page ?? ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1999. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=785716. IEEE Computer Society Or-

der Number PR00096. IEEE Order Plan Catalog Number PR00096.

Xue:1999:ICBa

[Xue99b] X. Xue. Image compression based on low-pass wavelet transform and multi-scale edge compensation. Part II: MSEC model. In Storer and Cohn [SC99], page ?? ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1999. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=785715. IEEE Computer Society Or-

der Number PR00096. IEEE Order Plan Catalog Number PR00096.

Xu:2010:SAD

[XW10] Jizheng Xu and Feng Wu. Subsampling-adaptive directional wavelet transform for image coding. In Storer and Marcellin [SM10b], pages 89–98. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=
5453453.

Xiao:2017:SOO

[XWC^{+}17] Jing Xiao, Zhongyuan Wang, Yu Chen, Liang Liao, Jun Xiao, Gen Zhan, and Ruimin Hu. A sensitive object-oriented approach to big surveillance data compression for social security applications in smart cities. Software — Practice and Experience, 47(8):1061–1080, August 2017. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Xiao:2006:ERT

[XWDY06] Song Xiao, Chengke Wu, Jianchao Du, and Yadong Yang. Error resilient transmission of H.264 video over wireless network. In Storer and Cohn [SC06], page 472. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=1607315.

Xie:2003:CCU

[XWL03] Yuan Xie, W. Wolf, and H. Lekatsas. Code compression using variable-to-fixed coding based on arithmetic coding. In Storer and Cohn [SC03], pages 382–391. ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 2003. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1194029. IEEE Computer

Society Order number PR01896.

Xu:2008:DLT

[XWLZ08] Jizheng Xu, Feng Wu, Jie Liang, and Wenjun Zhang. Directional lapped transforms for image coding. In Storer and Marcellin [SM08], pages 142–151. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33 INTERNET. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=4483292.

Xu:2011:IBP

[XXZ11] Yang Xu, Hongkai Xiong, and Y. F. Zheng. Inferring BP priority order using 5D tensor voting for inpainting-based macroblock prediction. In Storer and Marcellin [SM11b], page 483. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=5749540. IEEE

Computer Society Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Xie:2015:CTC

[XZG15] Xiaojing Xie, Shuigeng Zhou, and Jihong Guan. CoGI: Towards compressing genomes as an image. IEEE/ACM Transactions on Computational Biology and Bioinformatics, 12(6):

1275–1285, November 2015. CODEN ITCBCY. ISSN 1545-5963 (print), 1557-9964 (electronic).

Y:2013:LCE

[YA13] Asnath Victy Phamila Y. and R. Amutha. Low complexity energy efficient very low bit-rate image compression scheme for wireless sensor network. Information Processing Letters, 113(18):672–676, September 15, 2013. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0020019013001750.

Yacobi:1998:FEU

[Yac98] Yacov Yacobi. Fast exponentiation using data compression. SIAM Journal on Computing, 28(2):700–703, April 1998. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-

7111 (electronic). URL http://epubs.siam.org/sam-bin/
dbq/article/23497.

Yalabik:1978:CDC

[Yal78] Nese Yalabik. Contour Data Compression Exploiting Curve-To-Curve Dependence. Ph.D. thesis, Brown University, Providence, RI, USA, 1978. 100 pp. URL http://search.

proquest.com/docview/302882442.

Yamamoto:2000:NRU

[Yam00] Hirosuke Yamamoto. A new recursive universal code of the positive integers. IEEE Transactions on Information Theory, 46(2):717–723, March 2000. CODEN IETTAW. ISSN 0018-9448 (print), 1557-9654 (electronic).

Yan:2003:RPQ

[Yan03] Bin Ju Yan. Research for power quality disturbance recognition and disturbance data compression. M.Eng., Sichuan University, Chengdu, People’s Republic of China, 2003. URL

http://search.proquest.com/docview/1024711625.

Yang:2010:PII

[Yan10] Yixian Yang, editor. Proceedings 2010 IEEE International Conference on Information Theory and Information Security: December 17–19, 2010, Beijing, China. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2010. ISBN 1-4244-6942-2. LCCN

QA76.9.A25. URL http://ieeexplore.ieee.org/servlet/
opac?punumber=5680738.

Yao:2003:PEG

[Yao03] Zhen Yao. Phrase elimination in greedy parsing dictionary coders with deferred innovation. In Storer and Cohn [SC03], page ?? ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2003. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1194075. IEEE Computer Society Order number PR01896.

Yazdi:1981:ELD

[Yaz81] Mohammad Ali Jezu Tabatabai Yazdi. Edge Location and Data Compression for Digital Imagery. Ph.D. thesis, Purdue University, West Lafayette, IN, USA, 1981. 178 pp. URL http:/

/search.proquest.com/docview/303174746.

Yeates:2000:UCI

[YBW00] S. Yeates, D. Bainbridge, and I. H. Witten. Using compression to identify acronyms in text. In Storer and Cohn [SC00], page ?? ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2000. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=838229. This

millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Yan:2018:SMR

[YCMR18] Yizhou Yan, Lei Cao, Samuel Madden, and Elke A. Rundensteiner. SWIFT: mining representative patterns from large event streams. Proceedings of the VLDB Endowment, 12(3): 265–277, November 2018. CODEN ???? ISSN 2150-8097.

Yu:2015:ESS

[YDDB15] William Yu, Noah M. Daniels, David Christian Danko, and Bonnie Berger. Entropy-scaling search of massive biological data. Cell Systems, 1(2):130–140, August 26,

2015. ISSN 2405-4712. URL http://gems.csail.mit.edu/;
 http://www.sciencedirect.com/science/article/pii/
S2405471215000587.

Yang:2008:RF

[YDLC08] L. Yang, R. P. Dick, H. Lekatsas, and S. Chakradhar. RAM for free. IEEE Spectrum, 45(8):38–43, August 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Yu:2011:EVC

[YDZL11] Like Yu, Feng Dai, Yongdong Zhang, and Shouxun Lin. Efficient video coding optimization using a novel perceptual distortion model. In Storer and Marcellin [SM11b], page 487. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5749544. IEEE Computer Society

Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Ye:2001:HTD

[Ye01] Zhong A. Ye. Hypothesis testing and data compression. M.S., Nankai University, Nankai, Peoples Republic of China, 2001.

URL http://search.proquest.com/docview/1027128794.

Yong:2002:NTB

[YG02] Xuerong Yong and M. J. Golin. New techniques for bounding the channel capacity of read/write isolated memory. In Storer and Cohn [SC02], page ?? ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2002. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1000025. IEEE Computer Society Order Number PR01477.

Yung-Gi:2003:DWS

[YG03] Wu Yung-Gi. Dynamic window search and probability dominated algorithm for fast vector quantization encoding. In Storer and Cohn [SC03], page ?? ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 2003. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1194072. IEEE Computer

Society Order number PR01896.

Yun:1997:LCC

[YGM97] Hee Cheol Yun, Brian K. Guenter, and Russell M. Mersereau. Lossless compression of computer generated animation frames. ACM Transactions on Graphics, 16(4):359–396, October 1997.

CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL http://www.acm.org:80/pubs/citations/
journals/tog/1997-16-4/p359-yun/.

Ye:2007:MEI

[YGNM07] Linning Ye, Jiangling Guo, B. Nutter, and S. Mitra. Memoryefficient image codec using line-based backward coding of wavelet trees. In Storer and Cohn [SC07], pages 213–222. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=4148760. IEEE Computer

Society Order Number P2791.

Yang:2001:CPD

[YGSR01] Dow-Yung Yang, Ananth Grama, Vivek Sarin, and Naren Ramakrishnan. Compression of particle data from hierarchical approximate methods. ACM Transactions on Mathematical Software, 27(3):317–339, September 2001. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Yang:2000:SSC

[YH00] Yan Yang and S. S. Hemami. Separate source and channel rate selection for video over ATM. In Storer and Cohn [SC00], page ?? ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2000. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=838228. This

millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Yamagiwa:2020:SBL

[YHM20] Shinichi Yamagiwa, Eisaku Hayakawa, and Koichi Marumo. Stream-based lossless data compression applying adaptive entropy coding for hardware-based implementation. Algorithms (Basel), 13(7), July 2020. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/13/7/
159.

Yousefizadeh:2004:DOT

[YJE04] H. Yousefi’zadeh, H. Jafarkhani, and F. Etemadi. Distortionoptimal transmission of progressive images over channels with random bit errors and packet erasures. In Storer and Cohn [SC04], pages 132–141. ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ????

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=1281458. IEEE Computer Society Order Number: P2082.

Yang:2000:SSF

[YJGS00] Dow-Yung Yang, A. Johar, A. Grama, and W. Szpankowski. Summary structures for frequency queries on large transaction sets. In Storer and Cohn [SC00], pages 420–429. ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2000. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=838182. This millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Yoshida:2010:EAA

[YK10] S. Yoshida and T. Kida. An efficient algorithm for almost instantaneous VF code using multiplexed parse tree. In Storer and Marcellin [SM10b], pages 219–228. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=5453464.

Yoshida:2011:PCP

[YK11] S. Yoshida and T. Kida. On performance of compressed pattern matching on VF codes. In Storer and Marcellin [SM11b], page 486. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=5749543. IEEE

Computer Society Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Yu:2020:OUR

[YK20] Heejung Yu and Taejoon Kim. Optimization of uplink rate and fronthaul compression in cloud radio access networks. Future Generation Computer Systems, 102(??):465–471, January 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-

7115 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167739X18318867.

Yan:1999:RVL

[YKLK99] Z. Yan, S. Kumar, J. Li, and C. C. J. Kuo. Reversible variable length codes (RVLC) for robust coding of 3D topological mesh data. In Storer and Cohn [SC99], page ?? ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1999. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=785717. IEEE Computer Society Or-

der Number PR00096. IEEE Order Plan Catalog Number PR00096.

Yoo:1998:EID

[YKO98] Youngjun Yoo, Younggap Kwon, and Antonio Ortega. Embedded image-domain adaptive compression of simple images. In Matthews et al. [M^{+}98], page ?? ISBN 0-7803-5148-7, 0-7803-5149-5, 0-7803-5150-9. LCCN TK5101.A1 A85 1998; TK454.2 .A8 1998.

Young-Lai:1999:ASM

[YL99] Matthew Young-Lai. Adding state merging to the DMC data compression algorithm. Information Processing Letters, 70(5): 223–228, June 21, 1999. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://www.elsevier.

nl/gej-ng/10/23/20/48/21/18/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/23/20/48/21/18/article.
pdf.

Yang:2005:CBT

[YL05] Qiaofeng Yang and S. Lonardi. A compression-boosting transform for 2D data. In Storer and Cohn [SC05], page ?? ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1402249.

Yang:2011:ENA

[YLZ11] Meng Yang, Xuguang Lan, and Nanning Zheng. Explicit network-adaptive robust multiple description coding. In Storer and Marcellin [SM11b], page 485. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=5749542. IEEE Computer Society Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Yu:1997:FRQ

[YM97] Dongchang Yu and M. W. Marcellin. A fixed-rate quantizer using block-based entropy-constrained quantization and runlength coding. In Storer and Cohn [SC97], pages 310–316. ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33

D232 1997. URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=582054. IEEE Computer Society

Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Yoon:2008:COR

[YM08] Byung-Jun Yoon and H. S. Malvar. Coding overcomplete representations of audio using the MCLT. In Storer and Marcellin [SM08], pages 152–161. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008;

QA76.9.D33 INTERNET. URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=4483293.

Ye:2010:EAI

[YMK10] Yan Ye, G. Motta, and M. Karczewicz. Enhanced adaptive interpolation filters for video coding. In Storer and Marcellin [SM10b], pages 435–444. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ????

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=
&arnumber=5453489.

Yu:1997:PBE

[YML97] G.-S. Yu, M. W. Marcellin, and M. M.-K. Liu. POCS based error concealment for packet video. In Storer and Cohn [SC97], page ?? ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1997. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=582151. IEEE

Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Yoshida:1999:AWB

[YMYS99] S. Yoshida, T. Morihara, H. Yahagi, and N. Satoh. Application of a word-based text compression method to Japanese and Chinese texts. In Storer and Cohn [SC99], page ?? ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1999. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=785718. IEEE Computer Society Or-

der Number PR00096. IEEE Order Plan Catalog Number PR00096.

Young:2017:DCD

[YNQ17] Vinson Young, Prashant J. Nair, and Moinuddin K. Qureshi. DICE: Compressing DRAM caches for bandwidth and capacity. ACM SIGARCH Computer Architecture News, 45(2):627– 638, May 2017. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Yang:2007:FLC

[YNXC07] Chun Yang, Yan Niu, Yubin Xia, and Xu Cheng. A fast lossless codec of continuous-tone images for thin client computing. In Storer and Cohn [SC07], page 409. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4148810. IEEE Computer Society

Order Number P2791.

Yamamoto:1991:NAO

[YO91] Hirosuke Yamamoto and Hiroshi Ochi. A new asymptotically optimal code for the positive integers. IEEE Transactions on Information Theory, 37(5):1420–1429, September 1991. CODEN IETTAW. ISSN 0018-9448 (print), 1557-9654 (electronic).

Yokoo:1991:IDH

[Yok91] Hidetoshi Yokoo. An improvement of dynamic Huffman coding with a simple repetition finder. IEEE Transactions on Communications, 39(1):8–10, January 1991. CODEN IECMBT. ISSN 0090-6778 (print), 1558-0857 (electronic).

Yokoo:1993:AAT

[Yok93] H. Yokoo. Application of AVL trees to adaptive compression of numerical data. In Storer and Cohn [SC93], pages 310– 319. ISBN 0-8186-3391-3 (microfiche), 0-8186-3392-1 (casebound). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 1993. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=253118. IEEE Computer

Society Press order number 3392-02. IEEE catalog number 93TH0536-3.

Yokoo:1996:ADC

[Yok96] Hidetoshi Yokoo. An adaptive data compression method based on context sorting. In Storer and Cohn [SC96], pages 160–169. ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1996. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=488321. IEEE

Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Yokoo:1997:DCU

[Yok97] Hidetoshi Yokoo. Data compression using a sort-based context similarity measure. The Computer Journal, 40(2–3):94–102, ???? 1997. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/40/2_and_3/94.full.pdf+html;   http://
www3.oup.co.uk/computer_journal/Volume_40/Issue_02/
Vol40_02.body.html#AbstractYokoo;   http://www3.oup.
co.uk/computer_journal/Volume_40/Issue_02/Vol40_03.
body.html#AbstractYokoo.

Yokoo:1998:CTT

[Yok98] H. Yokoo. Context tables: a tool for describing text compression algorithms. In Storer and Cohn [SC98], pages 299–308. ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33

D232 1998. URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=672158. IEEE Computer Society

Press order number PR08406. IEEE order plan catalog number 98TB100225.

Yokoo:1999:DDS

[Yok99a] Hidetoshi Yokoo. A dynamic data structure for reverse lexicographically sorted prefixes. In M. Crochemore and M. Paterson, editors, Combinatorial Pattern Matching, volume 1645 of Lecture Notes in Computer Science, pages 150–162. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer-

ny.com/link/service/series/0558/bibs/1645/16450150.
htm;   http://link.springer-ny.com/link/service/series/
0558/papers/1645/16450150.pdf.

Yokoo:1999:PC

[Yok99b] Hidetoshi Yokoo. Private communication, 1999.

Yokoo:2010:FSP

[Yok10] H. Yokoo. File-size preserving LZ encoding for reversible data embedding. In Storer and Marcellin [SM10b], page 559. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5453534.

Yoon:2006:VQM

[Yoo06] Sangho Yoon. Vector quantization with model selection. In Storer and Cohn [SC06], pages 233–241. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=1607258.

York:1995:SCU

[Yor95] Kyle A. York. Sound compression using quantized deltas. Dr. Dobb’s Journal of Software Tools, 20(6):121–122, June 1995. CODEN DDJOEB. ISSN 1044-789X.

Young:1985:MFF

[You85] D. M. Young. MacWrite file format. Wheels for the Mind, 1: 34, Fall 1985. CODEN ???? ISSN 0891-8805.

Youssef:1997:ERI

[You97] A. Youssef. Error resiliency issues in wavelet compression. In Storer and Cohn [SC97], page ?? ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1997. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=582150. IEEE Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Youssef:1998:ACV

[You98a] A. Youssef. Analysis and comparison of various image downsampling and upsampling methods. In Storer and Cohn [SC98], page ?? ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1998. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=672325. IEEE

Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Youssef:1998:PAM

[You98b] A. Youssef. Parallel algorithms for multi-indexed recurrence relations with applications to DPCM image compression. In Storer and Cohn [SC98], page ?? ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1998. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=672326. IEEE Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Yahampath:2003:SDB

[YP03] P. Yahampath and M. Pawlak. Soft-decoding based vector quantization for hidden-Markov channels. In Storer and Cohn [SC03], page ?? ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2003. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1194074. IEEE Computer Society Order number PR01896.

Yang:2005:PLS

[YQ05] Shengtian Yang and Peiliang Qiu. On the performance of linear Slepian–Wolf codes for correlated stationary memoryless sources. In Storer and Cohn [SC05], pages 53–62. ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1402166.

Yu:1987:RDI

[YR87] C. T. Yu and C. J. Rijsbergen, editors. Research and Development in Information Retrieval: Proceedings of the Tenth Annual International ACM SIGIR Conference. ACM Press, New York, NY 10036, USA, 1987. ISBN 0-89791-232-2. LCCN Z 699 A1 I58 1987. US$19.

Yang:1996:MMF

[YR96] Xuguang Yang and K. Ramchandran. Morphological motion field representation for region-based image sequence coding. In Storer and Cohn [SC96], pages 112–121. ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1996. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=488316. IEEE Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Yao:2002:LRC

[YR02] Zhen Yao and N. Rajpoot. Less redundant codes for variable size dictionaries. In Storer and Cohn [SC02], page ?? ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2002. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=1000024. IEEE

Computer Society Order Number PR01477.

Yovanof:1991:LCT

[YS91] G. S. Yovanof and J. R. Sullivan. Lossless coding techniques for color graphical images. In Storer and Reif [SR91b], page ?? ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1990. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=213322. IEEE

Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Yu:1992:DCH

[YS92] Bin Yu and T. P. Speed. Data compression and histograms. Probability Theory and Related Fields, 92(2):195–229, ???? 1992. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064

(electronic). URL http://link.springer.com/article/10.
1007/BF01194921.

Yang:2007:DBE

[YS07] Jeehong Yang and S. A. Savari. Dictionary-based English text compression using word endings. In Storer and Cohn [SC07], page 410. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=4148811. IEEE

Computer Society Order Number P2791.

Yang:2010:LCL

[YS10] Jeehong Yang and S. A. Savari. A lossless circuit layout image compression algorithm for maskless lithography systems. In Storer and Marcellin [SM10b], pages 109–118. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5453455.

Ye:2000:SDD

[YSCK00] Yan Ye, D. Schilling, P. Cosman, and Hyung Hwa Ko. Symbol dictionary design for the JBIG2 standard. In Storer and Cohn [SC00], pages 33–42. ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2000. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=838143. This

millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Yang:2008:AGN

[YSM08] Jeehong Yang, S. A. Savari, and O. Mencer. An approach to graph and netlist compression. In Storer and Marcellin [SM08], pages 33–42. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008;

QA76.9.D33 INTERNET. URL http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=4483281.

Yang:2004:ACD

[YSXZ04] Y. Yang, V. Stankovic, Zixiang Xiong, and W. Zhao. Asymmetric code design for remote multiterminal source coding. In Storer and Cohn [SC04], page ?? ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1281548. IEEE Computer Society Order Number: P2082.

Yang:2005:MSC

[YSXZ05] Yang Yang, V. Stankovic, Zixiang Xiong, and Wei Zhao. On multiterminal source code design. In Storer and Cohn [SC05], pages 43–52. ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=
1402165.

Yang:2011:CPM

[YSZ11] Xiaoyuan Yang, Yan Shi, and Wanlu Zhou. Construction of parameterizations of masks for tight wavelet frames with two symmetric/antisymmetric generators and applications in image compression and denoising. Journal of Computational and Applied Mathematics, 235(8):2112–2136, February 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-

1778 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0377042710005650.

Yoshizawa:2012:AIS

[YTY12] Sayuri Yoshizawa, Takao Terano, and Atsushi Yoshikawa. Assessing the impact of student peer review in writing instruction by using the normalized compression distance. IEEE Transactions on Professional Communication, 55(1):85–96, March 2012. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Yu:1995:SWC

[Yu95] Tong Lai Yu. Sliding-window compression for PC software distribution. In Storer and Cohn [SC95], page 468. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=515578. IEEE catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Yu:1996:DCP

[Yu96a] Tong Lai Yu. Data compression for PC software distribution. Software — Practice and Experience, 26(11):1181–1195, November 1996. CODEN SPEXBL. ISSN 0038-0644 (print),

1097-024X (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract?ID=16769.

Yu:1996:DMC

[Yu96b] Tong Lai Yu. Dynamic Markov compression. Dr. Dobb’s Journal of Software Tools, 21(1):30, 32, 96, 98–100, January 1996. CODEN DDJOEB. ISSN 1044-789X.

Yu:1996:SDC

[Yu96c] Tong Lai Yu. Speech data compression for embedded systems. In Storer and Cohn [SC96], page ?? ISBN 0-8186-7358-3 (case), 0-8186-7359-1 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1996. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=488392. IEEE Order Plan catalog number 96TB100013. IEEE Computer Society Press order number PR07358.

Yu:2004:IRC

[Yu04] W. Yu. Integrated rate control and entropy coding for JPEG 2000. In Storer and Cohn [SC04], pages 152–161. ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1281460. IEEE Computer Society

Order Number: P2082.

Yuan:2002:SDP

[Yua02] Ding Ya Yuan. Study on Data Processing and Compression of Cobalt-60 Digital Radiographic System. D.Eng., Tsinghua University, Tsinghua, Peoples Republic of China, 2002. URL

http://search.proquest.com/docview/1024721621.

Yao:2004:HFV

[YW04] Z. Yao and R. Wilson. Hybrid fractal video coding with neighbourhood vector quantisation. In Storer and Cohn [SC04], page ?? ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=1281549. IEEE

Computer Society Order Number: P2082.

Yeates:2001:TIC

[YWB01] S. Yeates, I. H. Witten, and D. Bainbridge. Tag insertion complexity. In Storer and Cohn [SC01c], pages 243–252. ISBN 0-7695-1031-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2001. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=917155. IEEE

CSP number 01PR1031.

Yen:2009:IDC

[YWC09] Kuang-Yi Yen, Chun-Feng Wu, and Wen-Whei Chang. Iterative decoding of convolutionally encoded multiple descriptions. In Storer and Marcellin [SM09], page 470. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4976524. IEEE Computer Society

Order Number P3592. BMS Part Number CFP09DCC-PRT.

Yang:2008:USN

[YWMS08] Allen Y. Yang, John Wright, Yi Ma, and S. Shankar Sastry. Unsupervised segmentation of natural images via lossy data compression. Computer Vision and Image Understanding: CVIU, 110(2):212–225, May 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Yoon:2003:ICU

[YWPG03] Sangho Yoon, Chee Sun Won, Kyungsuk Pyun, and R. M. Gray. Image classification using GMM with context information and with a solution of singular covariance problem. In Storer and Cohn [SC03], page ?? ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 2003. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1194076. IEEE Computer

Society Order number PR01896.

Ye:2012:OST

[YX12] Xinwei Ye and Hongkai Xiong. Optimal spatio-temporal projections with holo-Kronecker compressive sensing of video acquisition. In Storer and Marcellin [SM12], page 412. ISBN 1-4673-0715-7 (paperback). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=6189293. IEEE Com-

puter Society order number P4656.

Yang:2011:AQD

[YXPM11] Chun-Ling Yang, Dong-Qin Xiao, Lai-Man Po, and Wang-Hua Mo. Adaptive quantization in DCT domain for distributed video coding. In Storer and Marcellin [SM11b], page 484. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5749541. IEEE Computer Society

Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Yoshida:2004:CCI

[YYO04] S. Yoshida, H. Yahagi, and J. Odagiri. CSV compaction to improve data-processing performance for large XML documents. In Storer and Cohn [SC04], page ?? ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1281550. IEEE Computer Society

Order Number: P2082.

Yang:2012:TTP

[YYZ12] Zhuoluo Yang, Jinguo You, and Min Zhou. TDSC: a twophase duplicate string compression algorithm. Lecture Notes in Computer Science, 7234:27–34, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-29426-6_5/.

Yiannis:2007:CTF

[YZ07] John Yiannis and Justin Zobel. Compression techniques for fast external sorting. VLDB Journal: Very Large Data Bases, 16(2):269–291, April 2007. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Yan:2009:IIL

[YZ09] Yusong Yan and Hongmei Zhu. Invertible integer Lie group transforms. In Storer and Marcellin [SM09], page 469. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4976523. IEEE Computer Society

Order Number P3592. BMS Part Number CFP09DCC-PRT.

Ye:2016:PUL

[YZLH16] Fang Ye, Xun Zhang, Yibing Li, and Hui Huang. Primary user localization algorithm based on compressive sensing in cognitive radio networks. Algorithms (Basel), 9(2), June 2016. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https:

//www.mdpi.com/1999-4893/9/2/25.

Yang:2014:SCB

[YZZ^{+}14] Chi Yang, Xuyun Zhang, Changmin Zhong, Chang Liu, Jian Pei, Kotagiri Ramamohanarao, and Jinjun Chen. A spatiotemporal compression based approach for efficient big data processing on Cloud. Journal of Computer and System Sciences, 80(8):1563–1583, December 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002200001400066X.

Zhu:2013:TSD

[Z^{+}13] Ce Zhu et al., editors. 3D-TV system with depth-image-based rendering. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 1-4419-9963-9, 1-4419-9964-7 (e-book). LCCN ????

Zhang:2005:PPA

[ZA05] Yong Zhang and D. Adjeroh. Prediction by partial approximate matching for lossless image compression. In Storer and Cohn [SC05], page ?? ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=1402251.

Zhao:2000:NMO

[ZAA00] Minyi Zhao, A. A. Alatan, and A. N. Akansu. A new method for optimal rate allocation for progressive image transmission over noisy channels. In Storer and Cohn [SC00], pages 213–222. ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2000. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=838161. This millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Zalta:1988:AP

[Zal88] Edward N. Zalta. Are algorithms patentable? Notices of the American Mathematical Society, 35(6):796–799, 1988. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Zandi:1992:ANB

[Zan92] Ahmad Zandi. Adaption to nonstationary binary sources for data compression. Ph.D. thesis, University of California, Santa Cruz, Santa Cruz, CA, USA, 1992. 65 pp. URL http://

search.proquest.com/docview/303987140.

Zandi:1995:CCR

[ZASB95] A. Zandi, J. D. Allen, E. L. Schwartz, and M. Boliek. CREW: Compression with Reversible Embedded Wavelets. In Storer and Cohn [SC95], pages 212–221. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=515511. IEEE catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Zhang:2012:TSB

[ZB12] Mengmeng Zhang and Huihui Bai. Temporal sampling based multiple description video coding for scenes switching. In Storer and Marcellin [SM12], page 413. ISBN 1-4673-0715-7 (paperback). ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=6189294. IEEE Computer Society

order number P4656.

Zingirian:1997:CMV

[ZBMM97] N. Zingirian, P. Baglietto, M. Maresca, and M. Migliardi. Customizing MPEG video compression algorithms to specific application domains: The case of highway monitoring. Lecture Notes in Computer Science, 1311:46–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zue:1991:IIV

[ZC91] K. Zue and J. M. Crissey. An iteratively interpolative vector quantization algorithm for image data compression. In Storer and Reif [SR91b], pages 139–148. ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1990. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=213381. IEEE Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Zeng:1999:MSE

[ZC99] Z. Zeng and I. Cumming. Modified SPIHT encoding for SAR image data. In Storer and Cohn [SC99], page ?? ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1999. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=785719. IEEE Computer Society Or-

der Number PR00096. IEEE Order Plan Catalog Number PR00096.

Zhao:1998:FFC

[ZCG98] D. Zhao, Y. K. Chan, and Wen Gao. FACOLA — FAce COder based on Location and Attention. In Storer and Cohn [SC98], page ?? ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1998. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=672328. IEEE

Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Zhao:1999:EDN

[ZCG99] D. Zhao, Y. K. Chan, and W. Gao. Extending DACLIC for near-lossless compression with postprocessing of greyscale images. In Storer and Cohn [SC99], page ?? ISBN 0-7695-0096-X, 0-7695-0098-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1999. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=785720. IEEE Computer Society Or-

der Number PR00096. IEEE Order Plan Catalog Number PR00096.

Zheng:2020:SDR

[ZCJ^{+}20] Qing Zheng, Charles D. Cranor, Ankush Jain, Gregory R. Ganger, Garth A. Gibson, George Amvrosiadis, Bradley W. Settlemyer, and Gary Grider. Streaming data reorganization at scale with DeltaFS indexed massive directories. ACM Transactions on Storage, 16(4):23:1–23:31, November 2020. CODEN ???? ISSN 1553-3077 (print), 1553-3093 (electronic). URL

https://dl.acm.org/doi/10.1145/3415581.

Zhang:2011:CSR

[ZCW^{+}11] Xinyu Zhang, Zhuoyuan Chen, Jiangtao Wen, Jianwei Ma, Yuxing Han, and J. Villasenor. A compressive sensing reconstruction algorithm for trinary and binary sparse signals using pre-mapping. In Storer and Marcellin [SM11b], pages 203–212. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=5749478. IEEE

Computer Society Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Zaniolo:1981:VLD

[ZD81] Carlo Zaniolo and C. Delobel, editors. Very large data bases: proceedings: seventh International Conference on Very Large Data Bases, Cannes, France, September 9–11, 1981. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1981. LCCN QA76.9.D3 I559 1981. IEEE catalog no. 81CH1701-2.

Zhang:1995:PBLa

[ZD95a] Q. Zhang and J. M. Danskin. A pattern-based lossy compression scheme for document. In Brown et al. [BBKF95], pages 221–234. ISSN 0894-3982.

Zhang:1995:PBLb

[ZD95b] Qin Zhang and John M. Danskin. A pattern-based lossy compression scheme for document images. Electronic Publishing —Origination, Dissemination, and Design, 8(2–3):221–233, June/September 1995. CODEN EPODEU. ISSN 0894-3982.

Ziviani:2000:CKN

[ZdMNBY00] Nivio Ziviani, Edleno Silva de Moura, Gonzalo Navarro, and Ricardo Baeza-Yates. Compression: a key for nextgeneration text retrieval systems. Computer, 33(11):37– 44, November 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

computer.org/co/books/co2000/pdf/ry037.pdf;   http:
//www.computer.org/computer/co/ry037abs.htm.

Zheng:2006:AMA

[ZDR06] J. Zheng, E. Duni, and B. D. Rao. Analysis of multiple antenna systems with finite-rate feedback using high resolution quantization theory. In Storer and Cohn [SC06], pages 73– 82. ISBN 0-7695-2545-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=1607242.

Zhang:1997:CGA

[ZDY97] Qin Zhang, J. M. Danskin, and N. E. Young. A codebook generation algorithm for document image compression. In Storer and Cohn [SC97], pages 300–309. ISBN 0-8186-7761-9 (case), 0-8186-7763-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1997. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=582053. IEEE Computer Society Press order number PR07761. IEEE Order Plan catalog number 96TB100108.

Zhao:2000:LLB

[ZE00] Qian Zhao and M. Effros. Lossless and lossy broadcast system source codes: theoretical limits, optimal design, and empirical performance. In Storer and Cohn [SC00], pages 63–72. ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2000. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=838146. This millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Zhao:2001:OCD

[ZE01] Qian Zhao and M. Effros. Optimal code design for lossless and near lossless source coding in multiple access networks. In Storer and Cohn [SC01c], pages 263–272. ISBN 0-7695-1031-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 2001. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=917157. IEEE CSP num-

ber 01PR1031.

Zhao:2003:LCC

[ZE03] Qian Zhao and M. Effros. Low complexity code design for lossless and near-lossless side information source codes. In Storer and Cohn [SC03], pages 3–12. ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 2003. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1193991. IEEE Computer

Society Order number PR01896.

Zeckendorf:1972:RNN

[Zec72] E. Zeckendorf. Représentation des nombres naturels par une somme de nombres de Fibonacci ou de nombres de Lucas. (French) [Representation of the natural numbers by a sum of Fibonacci or Lucas numbers]. Bulletin de la Société royale des sciences de Liège, 41:179–182, 1972. CODEN BSRSA6. ISSN 0037-9565.

Zeevi:1998:PVQ

[Zee98] A. J. Zeevi. On the performance of vector quantizers empirically designed from dependent sources. In Storer and Cohn [SC98], pages 73–82. ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232 1998. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=672133. IEEE

Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Zeilberger:1993:TPT

[Zei93] D. Zeilberger. Theorems for a price: Tomorrow’s semi-rigorous mathematical culture. Notices of the American Mathematical Society, 40(8):978–981, Fall 1993. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Zeilberger:1994:TPT

[Zei94] D. Zeilberger. Theorems for a price: Tomorrow’s semi-rigorous mathematical culture. The Mathematical Intelligencer, 16(4): 11–14, ???? 1994. CODEN MAINDC. ISSN 0343-6993 (print), 1866-7414 (electronic). Reprint of [Zei93].

Zamir:1992:UCB

[ZF92] R. Zamir and M. Feder. Universal coding of band-limited sources by sampling and dithered quantization. In Storer and Cohn [SC92], pages 329–338. ISBN 0-8186-2717-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ????

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=227448. IEEE catalog number 91TH0436-6.

Zamir:1994:LQN

[ZF94] R. Zamir and M. Feder. On lattice quantization noise. In Storer and Cohn [SC94], pages 380–389. ISBN 0-8186-5636-0, 0-8186-5637-9 (case). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 1994. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=305946. IEEE

Catalog Number 93TH0626-2. IEEE Computer Society Press Order Number 5637-02.

Zhao:2004:MSC

[ZFE04] Q. Zhao, H. Feng, and M. Effros. Multiresolution source coding using entropy constrained dithered scalar quantization. In Storer and Cohn [SC04], pages 22–31. ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1281447. IEEE Computer Society Order Number: P2082.

Zhang:2002:PMC

[ZG02a] Youtao Zhang and R. Gupta. Path matching in compressed control flow traces. In Storer and Cohn [SC02], pages 132– 141. ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2002. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=999951. IEEE

Computer Society Order Number PR01477.

Zhang:2002:DCT

[ZG02b] Youtao Zhang and Rajiv Gupta. Data compression transformations for dynamically allocated data structures. Lecture Notes in Computer Science, 2304:14–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer-ny.com/link/service/series/
0558/bibs/2304/23040014.htm;   http://link.springer-
ny.com/link/service/series/0558/papers/2304/23040014.
pdf.

Zhao:2004:FNL

[ZGB04] R. Zhao, M. Gabriel, and G. G. Belford. Fast near-lossless or lossless compression of large 3D neuro-anatomical images. In Storer and Cohn [SC04], page ?? ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1281551. IEEE Computer Society Order Number: P2082.

Zhao:2002:DCC

[ZGF02] Ying Zhao and J. Garcia-Frias. Data compression of correlated non-binary sources using punctured turbo codes. In Storer and Cohn [SC02], pages 242–251. ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2002. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=999962. IEEE Computer Society Order Number PR01477.

Zyto:2002:SDM

[ZGS02] S. Zyto, A. Grama, and W. Szpankowski. Semi-discrete matrix transforms (SDD) for image and video compression. In Storer and Cohn [SC02], page ?? ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2002. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1000027. IEEE Computer Society Order Number PR01477.

Zhang:1990:JID

[Zha90] Manyun Zhang. The JPEG and Image Data Compression Algorithms. PhD thesis, ????, ????, 1990. ?? pp.

Zhang:1998:CMC

[Zha98] Tong Zhang. Compression by model combination. In Storer and Cohn [SC98], pages 319–328. ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1998. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=672160. IEEE Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Zhao:1999:RDC

[Zha99] Ping Hong Zhao. Research on data compression of industrial TV images based on wavelet transformation. M.Eng., China University of Mining and Technology, Beijing, People’s Republic of China, 1999.

Zhang:2011:RRV

[Zha11] Guoqiang Zhang. On the rate region of the vector Gaussian one-helper distributed source-coding problem. In Storer and Marcellin [SM11b], pages 263–272. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=5749484. IEEE Computer Society Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Zhen:2003:DIF

[Zhe03] Cheng Zhen. Design and implementation of filesystem with automatic data compression and decompression. M.Eng., Tsinghua University, Tsinghua, Peoples Republic of China, 2003.

URL http://search.proquest.com/docview/1024722244.

Zhu:2002:SDN

[Zhu02] Ping Zhu. Special data network protocol analysis, data compression and interface circuit. M.Eng., University of Electronic Science and Technology of China, Chengdu, Peoples Re-

public of China, 2002. URL http://search.proquest.com/
docview/1024708479.

Zhou:2019:RBC

[ZHX^{+}19] S. Zhou, Y. He, S. Xiang, K. Li, and Y. Liu. Region-based compressive networked storage with lazy encoding. IEEE Transactions on Parallel and Distributed Systems, 30(6):1390–1402, June 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhao:2005:NAD

[ZIIK05] Bo Zhao, K. Iwata, S. Itoh, and T. Kato. A new approach of DCA by using BWT. In Storer and Cohn [SC05], page ?? ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1402252.

Zandi:1993:SOP

[ZIL93] A. Zandi, B. Iyer, and G. Langdon. Sort order preserving data compression for extended alphabets. In Storer and Cohn [SC93], pages 330–339. ISBN 0-8186-3391-3 (microfiche), 0-8186-3392-1 (casebound). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1993. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=253116. IEEE Computer Society Press order number 3392-02. IEEE catalog number 93TH0536-3.

Zipstein:1991:DCF

[Zip91] M. Zipstein. Data compression with factor automata. In Storer and Reif [SR91b], page ?? ISBN 0-8186-9202-2 (case), 0-8186-6202-6 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

1990. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=213333. IEEE Computer Society Press order number 2202. IEEE catalog number 91TH0373-1.

Zipstein:1992:DCF

[Zip92] M. Zipstein. Data compression with factor automata. Theoretical Computer Science, 92(1):213–221, January 06, 1992. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zirkind:2007:ADC

[Zir07] Givon Zirkind. AFIS data compression: an example of how domain specific compression algorithms can produce very high compression ratios. Computer Graphics, 41(4):1–36, November 2007. CODEN CGRADI, CPGPBZ. ISSN 0097-8930 (print), 1558-4569 (electronic).

Zhang:2012:CAS

[ZJW^{+}12] Yongbing Zhang, Xiangyang Ji, Haoqian Wang, Lei Zhang, and Qionghai Dai. Content adaptive subsampling for stereo interleaving video coding. In Storer and Marcellin [SM12], pages 179–188. ISBN 1-4673-0715-7 (paperback). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http:

//ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=

6189249. IEEE Computer Society order number P4656.

Zhang:2011:BRR

[ZK11] Guoqiang Zhang and W. B. Kleijn. Bounding the rate region of the two-terminal vector Gaussian CEO problem. In Storer and Marcellin [SM11b], page 488. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=5749545. IEEE Computer Society Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Zuk:2005:AER

[ZKD05] O. Zuk, I. Kanter, and E. Domany. Asymptotics of the entropy rate for a hidden Markov process. In Storer and Cohn [SC05], pages 173–182. ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ????

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=
&arnumber=1402178.

Zhang:2009:AKC

[ZKK09] Guoqiang Zhang, J. Klejsa, and W. B. Kleijn. Analysis of K-channel multiple description quantization. In Storer and Marcellin [SM09], pages 53–62. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=4976449. IEEE Computer Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

Zhang:2010:BRR

[ZKØ10] Guoqiang Zhang, W. B. Kleijn, and J. Østergaard. Bounding the rate region of vector Gaussian multiple descriptions with individual and central receivers. In Storer and Marcellin [SM10b], pages 13–19. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=5453438.

Ziv:1977:UAS

[ZL77] Jacob Ziv and Abraham Lempel. A universal algorithm for sequential data compression. IEEE Transactions on Information Theory, IT-23(3):337–343, May 1977. CODEN IETTAW. ISSN 0018-9448 (print), 1557-9654 (electronic).

Ziv:1978:CIS

[ZL78] Jacob Ziv and A. Lempel. Compression of individual sequences via variable-rate coding. IEEE Transactions on Information Theory, IT-24(5):530–536, September 1978. CODEN IET-TAW. ISSN 0018-9448 (print), 1557-9654 (electronic).

Zampunieris:1995:EHL

[ZL95] D. Zampunieris and B. Le Charlier. Efficient handling of large sets of tuples with sharing trees. In Storer and Cohn [SC95], page 428. ISBN 0-8186-7012-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=515538. IEEE

catalog number 95TH8037. IEEE Computer Society Press order number PR07010.

Zhang:2000:CLM

[ZL00] Cha Zhang and Jin Li. Compression of Lumigraph with multiple reference frame (MRF) prediction and just-in-time rendering. In Storer and Cohn [SC00], pages 253–262. ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2000. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=838165. This millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Zhao:2011:CIP

[ZL11] Hua Zhao and Songfeng Lu. Compressed index for property matching. In Storer and Marcellin [SM11b], pages 133– 142. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=5749471. IEEE

Computer Society Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Zhang:2019:CCF

[ZL19] X. Zhang and Y. Lao. On the construction of composite finite fields for hardware obfuscation. IEEE Transactions on Computers, 68(9):1353–1364, September 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhao:2015:BSB

[ZLC^{+}15] Jishen Zhao, Sheng Li, Jichuan Chang, John L. Byrne, Laura L. Ramirez, Kevin Lim, Yuan Xie, and Paolo Faraboschi. Buri: Scaling big-memory computing with hardwarebased memory expansion. ACM Transactions on Architecture and Code Optimization, 12(3):31:1–31:??, October 2015. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Zheng:2006:IFI

[ZLN06] Yang Zheng, Guanrong Liu, and Xiaoxiao Niu. An improved fractal image compression approach by using iterated function system and genetic algorithm. Computers and Mathematics with Applications, 51(11):1727–1740, June 2006. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-

7668 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0898122106001398.

Zhao:2009:FPS

[ZLRZ09] Fan Zhao, Guizhong Liu, Feifei Ren, and Na Zhang. Flexible predictions selection for multi-view video coding. In Storer and Marcellin [SM09], page 471. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=4976525. IEEE Computer Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

Zhou:2018:AMR

[ZLT^{+}18] Ke Zhou, Wenjie Liu, Kun Tang, Ping Huang, and Xubin He. Alleviating memory refresh overhead via data compression for high performance and energy efficiency. IEEE Transactions on Parallel and Distributed Systems, 29(7):1469–1483, July 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL https://www.computer.org/csdl/trans/td/
2018/07/08068250-abs.html.

Zou:2020:POR

[ZLX^{+}20] X. Zou, T. Lu, W. Xia, X. Wang, W. Zhang, H. Zhang, S. Di, D. Tao, and F. Cappello. Performance optimization for relative-error-bounded lossy compression on scientific data. IEEE Transactions on Parallel and Distributed Systems, 31 (7):1665–1680, July 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2003:APM

[ZMAB03] Nan Zhang, A. Mukherjee, D. Adjeroh, and T. Bell. Approximate pattern matching using the Burrows–Wheeler transform. In Storer and Cohn [SC03], page ?? ISBN 0-7695-1896-6. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

QA76.9.D33 D37 2003. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1194077. IEEE Computer

Society Order number PR01896.

Zordan:2014:PLC

[ZMVR14] Davide Zordan, Borja Martinez, Ignasi Vilajosana, and Michele Rossi. On the performance of lossy compression schemes for energy constrained sensor networking. ACM Transactions on Sensor Networks, 11(1):15:1–15:??, August 2014. CODEN ???? ISSN 1550-4859 (print), 1550-4867 (electronic).

Zhang:2008:FSE

[ZN08] Sen Zhang and Ge Nong. Fast and space efficient linear suffix array construction. In Storer and Marcellin [SM08], page 553. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33

INTERNET. URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4483380.

Zhang:2004:OBA

[ZO04] J. Zhang and C. B. Owen. Octree-based animated geometry compression. In Storer and Cohn [SC04], pages 508–517. ISBN 0-7695-2082-0. ISSN 1068-0314 (print), 2375-0359 (elec-

tronic). LCCN ???? URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1281496. IEEE Computer

Society Order Number: P2082.

Zorpette:1988:FJA

[Zor88] Glenn Zorpette. Fractals: Not just another pretty picture. IEEE Spectrum, 25(10):29–31, October 1988. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zhao:2002:CCM

[ZR02] Yafan Zhao and I. E. G. Richardson. Computational complexity management of motion estimation in video encoders. In Storer and Cohn [SC02], page ?? ISBN 0-7695-1477-4, 0-7695-1478-2 (case), 0-7695-1479-0 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37

2002. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=1000026. IEEE Computer Society Order Number PR01477.

Zhang:2000:RVC

[ZRR00] Rui Zhang, S. L. Regunathan, and K. Rose. Robust video coding for packet networks with feedback. In Storer and Cohn [SC00], pages 450–459. ISBN 0-7695-0592-9, 0-7695-0594-5 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D37 2000. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=838185. This

millennial edition of the DCC proceedings is dedicated to the memory of David A. Huffman, 1925–1999. IEEE Computer Society order number PR00592.

Zhang:2019:CCL

[ZRZ^{+}19] Jeff (Jun) Zhang, Parul Raj, Shuayb Zarar, Amol Ambardekar, and Siddharth Garg. CompAct: On-chip compression of activations for low power systolic array based CNN acceleration. ACM Transactions on Embedded Computing Systems, 18(5s): 47:1–47:??, October 2019. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https://dl.acm.org/

ft_gateway.cfm?id=3358178.

Zanon:2019:FKC

[ZSP^{+}19] G. H. M. Zanon, M. A. Simplicio, G. C. C. F. Pereira, J. Doliskani, and P. S. L. M. Barreto. Faster key compression for isogeny-based cryptosystems. IEEE Transactions on Computers, 68(5):688–701, May 2019. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2005:FCT

[ZTSM05] Nan Zhang, Tao Tao, R. V. Satya, and A. Mukherjee. A flexible compressed text retrieval system using a modified LZW algorithm. In Storer and Cohn [SC05], page ?? ISBN 0-7695-2309-9. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1402250.

Zurek:1989:TCC

[Zur89] Wojciech Zurek. Thermodynamic cost of computation, algorithmic complexity, and the information metric. Nature, 341 (6238):119–124, September 14, 1989. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Zhang:2007:ICQ

[ZW07] Xiangjun Zhang and Xiaolin Wu. Image coding on quincunx lattice with adaptive lifting and interpolation. In Storer and Cohn [SC07], pages 193–202. ISBN 0-7695-2791-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=4148758. IEEE Computer Society Order Number P2791.

Zhang:2008:CLR

[ZW08] Xiangjun Zhang and Xiaolin Wu. Can lower resolution be better? In Storer and Marcellin [SM08], pages 302–311. ISBN 0-7695-3121-0. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D372 2008; QA76.9.D33 INTER-

NET. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=4483308.

Zhou:2012:CMC

[ZW12] Jiantao Zhou and Xiaolin Wu. Context modeling and correction of quantization errors in prediction loop. In Storer and Marcellin [SM12], pages 82–88. ISBN 1-4673-0715-7 (paperback). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=6189239. IEEE Computer Society order number P4656.

Zhu:1992:IRH

[ZWS92] Q.-F. Zhu, Y. Wang, and L. Shaw. Image reconstruction for hybrid video coding systems. In Storer and Cohn [SC92], pages 229–238. ISBN 0-8186-2717-4. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=227458. IEEE

catalog number 91TH0436-6.

Zhang:2011:SDF

[ZWZ11] Yongbing Zhang, Haoqian Wang, and Debin Zhao. Upsampling dependent frame rate reduction for low bit-rate video coding. In Storer and Marcellin [SM11b], page 489. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5749546. IEEE Computer Society

Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Zhang:2012:CSV

[ZWZ^{+}12] Zhebin Zhang, Ronggang Wang, Chen Zhou, Yizhou Wang, and Wen Gao. A compact stereoscopic video representation for 3D video generation and coding. In Storer and Marcellin [SM12], pages 189–198. ISBN 1-4673-0715-7 (paperback). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=6189250. IEEE Computer Society order number P4656.

Zhang:2018:DSD

[ZWZ^{+}18] Binqi Zhang, Chen Wang, Bing Bing Zhou, Dong Yuan, and Albert Y. Zomaya. DCDedupe: Selective deduplication and delta compression with effective routing for distributed storage. Journal of Grid Computing, 16(2):195–209, June 2018. CODEN ???? ISSN 1570-7873 (print), 1572-9184

(electronic). URL https://link.springer.com/article/
10.1007/s10723-018-9429-3.

Zou:2012:PIR

[ZWZW12] Bin Zou, Dewu Wang, Ye Zhang, and Zhilu Wu. Phase information reserved polarimetric SAR raw data compression. In Storer and Marcellin [SM12], page 414. ISBN 1-4673-0715-7 (paperback). ISSN 1068-0314 (print), 2375-0359 (electronic).

LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=6189295. IEEE Computer Society

order number P4656.

Zhu:2011:ITB

[ZX11] Lingchen Zhu and Hongkai Xiong. Iterative thresholding-based sparse directional representation for efficient low bit-rate embedded video coding. In Storer and Marcellin [SM11b], pages 93–102. ISBN 1-61284-279-8. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=5749467. IEEE

Computer Society Order Number P4352; BMS Part Number: CFP11DCC-PRT.

Zhang:2010:ARM

[ZXM^{+}10] Yongbing Zhang, Xinguang Xiang, Siwei Ma, Debin Zhao, and Wen Gao. Auto regressive model and weighted least squares based packet video error concealment. In Storer and Marcellin [SM10b], pages 455–464. ISBN 0-7695-3994-7. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN

???? URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=5453491.

Zhang:2020:IRPb

[ZYF^{+}20] Y. Zhang, Y. Yuan, D. Feng, C. Wang, X. Wu, L. Yan, D. Pan, and S. Wang. Improving restore performance for inline backup system combining deduplication and delta compression. IEEE Transactions on Parallel and Distributed Systems, 31(10):2302–2314, 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zheng:2015:DGC

[ZYT^{+}15] Haifeng Zheng, Feng Yang, Xiaohua Tian, Xiaoying Gan, Xinbing Wang, and Shilin Xiao. Data gathering with compressive sensing in wireless sensor networks: A random walk based approach. IEEE Transactions on Parallel and Distributed Systems, 26(1):35–44, January 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL http://www.computer.org/csdl/trans/td/
2015/01/06748092-abs.html.

Zhang:1998:PTT

[ZZ98] Bo Zhang and Y. F. Zheng. Packed-TS transform [for image compression]. In Storer and Cohn [SC98], page ?? ISBN 0-8186-8406-2 (case), 0-8186-8408-9 (microfiche). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN QA76.9.D33 D232

1998. URL http://ieeexplore.ieee.org/stamp/stamp.

jsp?tp=&arnumber=672327. IEEE Computer Society Press order number PR08406. IEEE order plan catalog number 98TB100225.

Zhang:2021:TTA

[ZZD21] Feng Zhang, Jidong Zhai, and Xiaoyong Du. TADOC: Text analytics directly on compression. VLDB Journal: Very Large Data Bases, 30(2):163–188, March 2021. CO-DEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (elec-

tronic). URL https://link.springer.com/article/10.
1007/s00778-020-00636-3.

Zhang:2013:FBA

[ZZJB13] Yan Zhang, Fan Zhang, Zheming Jin, and Jason D. Bakos. An FPGA-Based accelerator for frequent itemset mining. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 6(1):2:1–2:??, May 2013. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Zhao:2015:GSB

[ZZL^{+}15] Wayne Xin Zhao, Xudong Zhang, Daniel Lemire, Dongdong Shan, Jian-Yun Nie, Hongfei Yan, and Ji-Rong Wen. A general SIMD-based approach to accelerating compression algorithms. ACM Transactions on Information Systems, 33(3):15:1–15:??, March 2015. CODEN ATISET. ISSN 1046-8188.

Zribi:2009:LCJ

[ZZPB09] A. Zribi, S. Zaibi, R. Pyndiah, and A. Bouallegue. Lowcomplexity joint source/channel turbo decoding of arithmetic codes with image transmission application. In Storer and Marcellin [SM09], page 472. ISBN 0-7695-3592-5. ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ????

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=

&arnumber=4976526. IEEE Computer Society Order Number P3592. BMS Part Number CFP09DCC-PRT.

Zhang:2022:PHP

[ZZS^{+}22] Feng Zhang, Jidong Zhai, Xipeng Shen, Onur Mutlu, and Xiaoyong Du. POCLib: a high-performance framework for enabling near orthogonal processing on compression. IEEE Transactions on Parallel and Distributed Systems, 33(2):459– 475, February 2022. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2012:CSR

[ZZZ^{+}12] Jian Zhang, Debin Zhao, Chen Zhao, Ruiqin Xiong, Siwei Ma, and Wen Gao. Compressed sensing recovery via collaborative sparsity. In Storer and Marcellin [SM12], pages 287–296. ISBN 1-4673-0715-7 (paperback). ISSN 1068-0314 (print), 2375-0359 (electronic). LCCN ???? URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=6189260. IEEE Com-

puter Society order number P4656.